A Complete Bibliography of Publications in IEEE Transactions on Parallel and Distributed Systems:

2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 2.01

Title word cross-reference

(e, d) [LC12a]. (K) [WWLX13, GLM13]. (t, k) [Cha11]. (UCON_{ABC}) [MSSB14]. 2 [AVA^{+}17, HWZE10, KGI17, LSWR16]. 3 [AAB16, BKF^{+}16, CLHW13, CCLW15, Che18b, DWH^{+}18, GRUMG17, GAB18, LWSM19, SWT^{+}19, WJTZ14, WCYL19, YTL^{+}19, ZM13, ZYX^{+}10]. 4 [Has16, IGEN11]. ^{cc} [MRH^{+}16]. g [YLM^{+}15]. K [KPA13, WHC^{+}14, YPL^{+}17, Amm12, AH10, FMY^{+}18, JRAS17, JCW^{+}12, LL12, MLT^{+}13, MDM13, PSMD18, SLL16, SRB14, THE^{+}15, XS11, XHHC13, XQL^{+}14, YLM^{+}15, ZZQ18].

LU [HAZ^{+}18, KLFD13]. n [LL12, SLM^{+}10, WCZ^{+}19a, XS11, YLM^{+}15]. O(1) [ACS13, XL10]. QR [MVC^{+}18]. S^{2} [YXWW14]. ε [LLG15a].

-Anycast [WWLX13]. -Approximate [LC12a]. -Approximation [LLG15a, LSWR16]. -Ary [XS11, YLM^{+}15, LL12]. -CAS [AH10]. -Clique [GLM13]. -Clustering [PSMD18]. -Core [MDM13]. -Coverage [MLT^{+}13, YPL^{+}17]. -Covered [Amm12]. -Cube [LL12]. -Cubes [XS11, YLM^{+}15]. -D [GRUMG17, AVA^{+}17, KGI17]. -Diagnosis [Cha11]. -Good-Neighbor [YLM^{+}15]. -Means [XQL^{+}14, KPA13].

-NN [THE^{+}15]. -NN-Based [XHHC13]. -Pancyclicity [LL12]. -RAID [YXWW14]. -Set [JRAS17, SRB14]. -Shortest [FMY^{+}18]. -Step [WHC^{+}14]. -Tier [WCZ^{+}19a]. -Way [SLL16, SLM^{+}10].

1 [ATZZ14].

2019 [Par19b]. 2PASS [HX10].

3.42-Approximation [CC13b]. 360 [RSSC15].

4 [ZWL^{+}15]. 4.0 [dOSMM^{+}16]. 4K [BB15].

5 [MWZX14].

6 [SSF16a, ZWL^{+}16a].

802.11 [FLH13, GYX^{+}10]. 802.11e [MRM12]. 802.15.4 [PDFJ13, TMMN15]. 802.15.6 [RMM16].

A-WiNoC [DKM^{+}15]. Abstraction [LN17, MBH^{+}10, RHT13]. Abstractions [MR16, UDH^{+}17]. AC [APPG16]. AC-WAR [APPG16]. ACC [APJ^{+}16]. Accelerate [RLVTMG^{+}16]. Accelerated [AHJ^{+}11, AJK^{+}17, CRWY15, DB18, EADT19, HKE^{+}16, LLL^{+}14a, LWCL18, SLX19, SCHT16]. Accelerating [AC19, CDPM18, FHLG11, JNL^{+}15, KBS11, LZWZ19, LAFA15, OOA^{+}14, RZB^{+}18, SZA11, SVL^{+}16, TLP12, TZT^{+}16, vdLJR11]. Acceleration [FJV^{+}18, MVC^{+}18, Mur12, OZMC^{+}16, PRS^{+}11, RRM^{+}15, RMB^{+}16, TP18, US16, WTD17, WH16]. Accelerator [APJ^{+}16, APCH^{+}11, CGS^{+}15, LNMMA15, SHY14]. Accelerator-Aware [APJ^{+}16]. Accelerator-Based [APCH^{+}11, LNMMA15]. Accelerators [AKGR13, ALI^{+}17, BKK11, KLFD13, TCM18, WGHP11, WM18]. Access [ALLR14, AJM12, Ano12i, ALI^{+}17, CGKP11, DGI^{+}19, FCM14, HDL^{+}15, HN11, IdM12, JSMK11, Jun17, KZW^{+}12, LXXH16, LTGI16, LWCG10, LLK13, LTM11, LS19, MLC^{+}19, RAS17, RSN14, SMS^{+}13, SGC14, SLS^{+}16, WWW^{+}18, WS14, WWL^{+}17, WWH^{+}17, XLM^{+}11a, XZT^{+}13, XHZ^{+}13, YJ14, YJR15, YRL11, ZZR12]. Accesses [HTA10, WVT13]. Accountable [RYLZ10]. Accrual [KM10]. Accrued [LSWR16]. Accumulative [ZGGW14]. Accuracy [HHWZ17, ITW^{+}14, WYX^{+}15, XSYY13, ZLY^{+}14]. Accurate [DO13, Liu14, MJM16, VTSM12, YGS^{+}19, ZS17, ZLGN13]. Achieve [SL16]. Achieving [GCN^{+}14, HAZ^{+}18, KN16, LC12b, LY11, XSL^{+}16, YYL^{+}13, ZH11]. Acid [LPSS19]. Acoustic [LLZ14]. ACPN [LLG15b]. Acquiring [ZSH^{+}11]. Acquisition [WNLL15, WLL15b]. Across [DWH^{+}18, LGL^{+}18b, LSW17b, Man18, XBZL17, HLL18, LMZG15]. ACStor [WWL^{+}17]. Activation [RCC^{+}14]. Active [CB16, HD15, KTK12, hKYY11, SVK^{+}19, YOK^{+}17]. Activity [LWY^{+}15, LZC^{+}12, SAH15, ZZG^{+}11]. Actor [WMT^{+}11]. Actors [HCC^{+}12]. Acyclic [YWJJ11]. Ad [AE12, BZA10, CCFS11, CLM^{+}15, CPM^{+}10, CYL^{+}14, CLJ11, DMR16, GLJ^{+}15, HCJ^{+}10, JJ11, JGG^{+}11, LLGP13, LYW^{+}12, LMSRSR13, LNA^{+}13, LHYW15, MM10, MY11, She14, SCC11, SZZF10, SJ14, WJTL13, WL14, WCF13, XAY^{+}14, ZZF10, ZHCW12]. Ad-Hoc [SJ14, XAY^{+}14]. ADAPT [ZJTZ14]. ADAPT-POLICY [ZJTZ14]. Adaptable [GFMR13, MLK15, SPH^{+}18]. Adaptation [CRRR15, DK17, LV15, MPS15, RPYO11, yWeH11, YZS13, ZSY14, ZHZL17]. Adapting [ScFRdS15]. Adaption [LSL^{+}14a]. Adaptive [APMG12, AIAD^{+}18, BKH18, CGH13, CLHW13, CSY15, CWZ^{+}15, Che18b, CRG^{+}17, CZWJ18, CLJ11, DC16, DKM^{+}15, DG15, Dua95, EHNS13b, FHW11, FC18, GAB18, HLZ^{+}19, HW13, HZJ^{+}11, JFP^{+}17, JJ11, KPB19, LXHS12, LLY^{+}14, LLK13, LS17c, LCL^{+}15, LX12, MWJ^{+}14, NCM^{+}17, RVCT15, SHG13, She10a, SLGW14, SCH11, TKC^{+}15, TGFFP^{+}19, VS11a, WTD17, WMW11, WMHX12, WPMX18, WCZ^{+}19b, WMC^{+}19, WHYZ10, YGL^{+}15, YL15, YXG12, ZTG^{+}18, ZZF10, ZYQ^{+}14, ZCC^{+}17, ZHZL17, Dua93]. Adaptive-Tree [APMG12]. Adding [ZDF^{+}15]. Additional [AJMW14]. Additions [GLGLBM13]. Address [LZW^{+}17]. Addressable [NTDZ19]. Adjustable [ZZF10]. Adjustment [CCL13, CYL^{+}14]. Administration [HFY^{+}14]. Admission [MSB11, PH11]. Admitting [MLXG19]. Advanced [ZHQ12]. Advances [BBD^{+}19, RBH^{+}14]. Advertising [QZZ^{+}16]. Aerodynamics [AK18]. AES [HMKG19]. Affine [KAC^{+}15]. Affinitizing [HT16]. Affinity [DCA^{+}16]. Against [AGG17, FWH^{+}18, ZYL^{+}17, MS12, QLC13, WMGA15, WXYX14, YYY^{+}14]. Agent [CWZ^{+}15, CBK^{+}10, HPG14, TCZL11, XVC17, YZS13, ZSY14]. Agent-Based [HPG14, XVC17]. Agents [MKOK14]. Aggregate [sCCyW14, CCT^{+}14]. Aggregated [NLY15, SML13]. Aggregated-Proof [NLY15]. Aggregating [Guo17, LZY12]. Aggregation [CC10, CLLS12, DN19, FC10, HJPL14, LC12a, LWY^{+}13, LLL^{+}12, MLL14, RZW^{+}13, SP15, TKS11, TWL^{+}15, WJTL12, WHN^{+}19, WLLL10, XLM^{+}11b, XGZW14, YRLY16, YXG12]. Agile [ZJLG14]. Aging [GAB18, LSL^{+}17, PAB13]. Aging-Aware [GAB18, PAB13]. Agnostic [DED^{+}19, FSM^{+}12]. Agreement [HCL^{+}14, JKT11, JRAS17, MR16, SRB14]. aHDFS [CZT^{+}17]. Ahead [MV18]. Aho [TVCM12]. Aho-Corasick [TVCM12]. Aided [SLL13a, TLJ^{+}14, WCF13]. Air [PT15, ZLZ^{+}14]. Airport [AOW^{+}12]. Algebra [LLCH12]. Algorithm [ADLM19, BKY15, BMB^{+}10, BB16, CC13b, CCH^{+}17, CLT^{+}17, CPL^{+}18, DLZH16, DSM19, EKNS17, GAB18, GLC^{+}15, GHW^{+}16, HWC15, HWG^{+}19, Has16, HMP^{+}19, HH11, HLY10, Hu14, IPQ19, IGEN11, JFP^{+}17, JSK18, JGHD10, KKW13, Kum14, LLCH12, LLW^{+}15, LSWR16, LYL16, LKT11, LY14, LLCL12, LX12, MVC^{+}18, RH16, SDV18, SJVR19, SAM14b, SLG10, She10a, SSM^{+}18, SKA15, TLP15, TCZL11, VMP17, WPKL13, WJTZ14, WQZ^{+}16, WYLH18, WYL19, WSS15, WKW19, XL10, XLM^{+}11b, XZT^{+}13, YXSS13, YN17, ZG11, ZLZ^{+}17, ZYQ^{+}14, ZBS15, ZJZ^{+}16, ZD16b, Zou14]. Algorithm-Architecture [MVC^{+}18]. Algorithm-Based [HWC15]. Algorithm/Architecture [LLCH12]. Algorithmic [Man16, TMJ14, WZGR10]. Algorithmics [PCFP16]. Algorithms [AS16, ABF12, BBCB15, CP17a, CCM^{+}17, CL17, CTX^{+}11, DWW^{+}11, DPRT11, EJRB13, FSM^{+}12, GGS10, GS17, HWZE10, HZJ16, Ksh10, LM17, LPSS19, LVA^{+}11, LC12a, LCGC14, LLLC17, LAD16, Lou14, LSW^{+}15, LHCM^{+}17, LXBZ13, MG18, MV12, MMSAZ11, Ozd19, PSL^{+}11, SJVR19, SLSG18, SLSG19, SSW^{+}17, SBC^{+}19, SZ12, SM16, SDL^{+}15, TKC^{+}15, Tsa13, WVT13, WG13, WZLC15, XZX^{+}17, XLX^{+}16, YTL^{+}10, ZWD^{+}10, ZD12, ZT14, ZCX15]. AliCloud [RSW^{+}17]. Alignment [dOSMM^{+}16, WH16]. Alignments [dOSdM13]. All-Around [SSF16a]. All-Flash [KZK^{+}19, KZK^{+}20]. All-Pairs [MBH^{+}10]. All-Path [LZB14]. All-Port [ZD12]. All-Prefix-Sum [KPA13]. All-To-All [Tou15a, LZH18, Tou15b]. Alleviate [RHDL11]. Alleviating [LA12, ZLT^{+}18]. Alleviation [BSL^{+}17]. Allocate [CW15]. Allocation [ASBL15, AIAD^{+}18, BEDCR13, BSM^{+}11, CB13, Che14, CFLL18, CP17c, DW13a, DW13b, DG15, FDFZB13, GLBJ18, GLC^{+}15, HCyW^{+}17, HKH^{+}10, HYX11, HKkY^{+}16, JWK^{+}16, JZW13, Jia16, JAJ^{+}19, JWNS19, JJG^{+}12, KALK^{+}18, LCG^{+}16, LTC16, LRJX13, LMAS17, LCW11, LLZ18b, LGG^{+}14, MNG15a, MRM12, NMG15, OPM^{+}15, PAB13, PCP14, ST10, SZR17, TFLL18, WLL15a, WKW16, WHGS17, WK11, WW12, WZL^{+}19, XAY^{+}14, XSC13, YQZC12, YLC^{+}16, ZWFX17, ZYL^{+}16]. Allocations [AT12]. Allocator [LGD14]. ALOHA [WZFG13]. Alternating [FXL17, LZXW15]. Amazon [MHL^{+}16, TYWL14]. Ameliorate [CL13]. AMI [DN19]. Amorphous [HH12]. Analysis [AHTD18, ATZZ14, BTL^{+}19, CLLX18, CHW^{+}17, CLL^{+}17, CCW^{+}12, Di 17, DLA^{+}18, DLB^{+}19, DY16, EJRB13, ECV16, FQWL12, GFS^{+}10, GZZ^{+}13, GD16, GWC14, HCH^{+}12, IOY^{+}11, KMM12, KMMR13, KAC^{+}15, Li13, LQK^{+}13, LYL15, LSJ^{+}19, LL11, LLC10, LLY^{+}15, LLG^{+}13, LLH^{+}15a, LWZ^{+}16b, MS15, MC10, MRM12, MSB11, PHGR17, PJAGW14, RMM16, RS12, RLVTMG^{+}16, SJVR19, SVK^{+}19, SWT^{+}19, SILJ11, SYXL16, SOTN12, SZ11, TXWL11, TJH^{+}14, VM12, WYW13, WGZ16, WKK17, WRL15, WMLJ12, WYCZ14, XXWY10, XLY^{+}17, XDLZ19, Yan14, YL11a, YNKD18, YZFZ10, YLR12, ZJLS12, ZD12, ZT14, ZFT^{+}15, ZTH17, ZHL19, ZCXF16, ZCX15, ZLLD18, ZFG^{+}10]. Analytical [Bar10, PFAF16, SV19]. Analytics [AHSK17, DED^{+}19, JZW^{+}17, KKE19, LGM^{+}17, LLY^{+}17, NCM^{+}17, SMS^{+}13, XXL^{+}19, XGL^{+}16, XZJ^{+}19, ZZSZ18]. Analyze [PWRL18]. Analyzing [BM12, CAC^{+}19, NL11, QPB^{+}17, SJR17]. Anchor [KSP10, XL13]. Anchor-Free [KSP10]. Annual [Ano11a, Ano12a, Ano13a, Ano14a]. Anomaly [DNW^{+}16, DLC^{+}16, LZL10, TP18, XHHC13, XHG15, YL16]. Anonymity [ZFG^{+}10]. Anonymization [ZYLC14]. Anonymizing [LHW11].

Anonymous [HX10, IPQ19, LZCK14, MKOK14, RSN14, Tan12]. Answer [XZH14]. Answering [LCL^{+}16a]. Antennas [CWJS11, JWA10, KCYM10, YW10]. Anti [XTFC17, ZJ16, WZS^{+}18]. Anti-Collusion [ZJ16, WZS^{+}18]. Anti-Colocation [XTFC17]. Anticollision [GFMR13, WZFG13]. Anycast [WWLX13]. AP [HST^{+}11]. Aperiodic [ZGL10]. APIs [ECW^{+}18]. AppBooster [LCY^{+}17]. Appearing [AJMW14]. APPLES [SDG17]. Appliance [KTK12]. Appliances [BRX13, CJZ12]. Application [AA14, CDPM18, CCCB14, DGC17, GFLL15, HJS^{+}11, KEGM12, LWC^{+}17, LWSM19, MHL^{+}16, MKVL12, NSLV16, PWRL18, RMB^{+}16, RS12, SV19, STK^{+}19, TWL^{+}15, TSN10, Ven14, VLP16, WMZ^{+}15, WCZ^{+}19a, WRL15, WKW19, WYZ^{+}19, XLT^{+}14, XSTZ10, Zha12]. Application-Aware [WMZ^{+}15, XLT^{+}14]. Application-Layer [TSN10]. Application-Level [STK^{+}19]. Application-Specific [CDPM18]. Applications [APJ^{+}16, ASBL15, BWB^{+}19, BMR15, BBGD^{+}17, CGS^{+}15, CB16, CSV^{+}17, CCT10, CCNMF18, CPH^{+}18, DLZH16, DB18, DLM^{+}17, DC16, DZLC15, EGQ11, FPRG16, GTM^{+}17, GFS^{+}10, GIX^{+}12, Goh14, GKT^{+}17, HAD12, HL12b, HC14, HKkY^{+}16, JHYK11, KOPS10, KL16, LAdS^{+}15, Lai12, LGJZ16, LM17, LZB14, LSWR16, LHJ12, LTBN^{+}12, LJB^{+}13, LH15, LCY^{+}17, LdSB19, LSW^{+}15, LHCM^{+}17, MZLT19, MHL^{+}16, MPM17, MNG^{+}15b, MDZC14, MLVD12, MVML11, NCGP19, NTWL11, RBSS11, RCV^{+}13, RGRM14, RGLDM17, SKGC14, SMS^{+}13, SVL^{+}16, SCH^{+}15, SPH^{+}18, SJAdCL19, SLM^{+}10, TCDMRP17, TP18, VMN^{+}16, VNA^{+}16, WJTZ14, WSC^{+}14, Wan19, WCYL19, WGHP11, XZX^{+}17, YQLS14, YC12, ZSH^{+}11, ZLJ^{+}15a, ZJS12, ZT14, ZYW^{+}14a, ZJZ^{+}16, ZLK^{+}16, ZT16, dBK11]. Applied [GS11b]. Applying [CWLS19]. Approach [BN12, Bar10, BYZ^{+}16, BZA10, BRX13, BZBP10, BB17, CAD^{+}18, CJW^{+}15, CCJ19, CHCC14, CYC^{+}15, CCW^{+}12, DDW^{+}19, DN19, DLM^{+}17, DSJ16, DIAR16, EN12, FYH^{+}15, FXL17, GG10, GTS^{+}15, GRB^{+}19, GV15, HMKG19, HKH^{+}10, HLYJ19, ITL17, IdM12, Iye14, JWY^{+}18, KN12, KKC17, KEGM12, KP12, KPG^{+}12, LTW^{+}14, LV15, LLC^{+}15, LLAL18, LLZ14, LCYW16, NN10, PGP^{+}17, QP16a, SG16b, SSPG17, SCL^{+}15, SKP12, TWW^{+}18, TXL^{+}14, TWL16, TKP12, VLP16, WHF^{+}19, WDL^{+}17, XYT^{+}15, XWL^{+}19, XSTZ10, YY10, YLZ^{+}15a, YLC^{+}16, YHS^{+}14, YZSC14, YPL13, YC14, YZT^{+}17, YYL^{+}13, ZLN^{+}13, ZYLC14, ZLS^{+}18, ZYW^{+}16, ZCLS14, ZYT^{+}15, dSLMM11, KLL^{+}17]. Approach-Based [BZA10]. Approaches [BKL11, KM19, MVL15, MV16a, WIZ^{+}17]. Appropriate [SP15]. Approximate [DFGG13, HHWZ17, HK18, HXLF15, HJF16, LC12a, LCGC14, LWJ^{+}15, MIH17, WMHX12, XXL^{+}19]. Approximated [XHG15]. Approximation [AF19, CC13b, Che18a, DPRT11, GS17, LLG15a, LSWR16, LY14, SP12, XQL^{+}14, ZWJ^{+}19]. Arbitrage [TWT16]. Arbitrary [CHTW12, DWF12, HV11, JVW10, LQW^{+}18, LWJ^{+}15, SOK^{+}19, ZD16b]. Arbitrary-Shaped [LWJ^{+}15]. Arbitrating [Jia14a]. Arbitration [QLNN13]. Architecting [APPG16, MV16c, Mit17]. Architectural [EHM^{+}17, MVL15, MV16a, SKGC14, WM18]. Architecture [ATACA18, AAB16, AF18, ACV17, ASD^{+}18, BICK^{+}15, BBM16, CGS^{+}15, CHM^{+}13, CLO^{+}18, CLB^{+}19, CP17c, DLXS19, DKM^{+}15, DBG^{+}14, EMW16, FC11, JHR^{+}14, JPG14, KBS11, KGR16, KJvR^{+}15, LLCH12, LJ15, LSLD17, LWT^{+}18, LWZ^{+}16a, MVC^{+}18, MB12, MJM16, MKSN18, NTA^{+}16, NHN17, NHN18, Nov15, PL16, RGRM14, SEA18, SHA19, STMM17, USP^{+}12, VMP17, VGMA10, WCLK12, WFZ^{+}17, WLC^{+}17, Wan19, WCYL19, XHC16, YXWW14, YJC^{+}16, YYL^{+}17, YTL^{+}19, ZTG^{+}18, ZDM^{+}19, ZL10]. Architecture-Based [Wan19]. Architectures [AAA19, AA17, AK18, BS15, BB15, BB16, BB17, CSV^{+}17, CGH13, CVM^{+}15, CDPM18, EJGYAM14, FSS11, FFC17, GDRTS16, Has16, Ian14, IGEN11, JSMK11, KGI17, Kao15, KPA13, KAG17, LWLZ17, LAD16, LKD10, PCL15, RH16, THB^{+}14, TVCM12, WYY^{+}12, WWLJ14, WYZ^{+}19, YKW^{+}18, YCMX17, YLLW16, ZTD19, ZZH^{+}17, ZHQ12]. Archival [CZT^{+}17, HWQ^{+}15]. Area [CH13, IvS10, LNL^{+}19, LZCK14, SLGW14, YYK11a, ZWWF15, CDR15]. AREA-Oriented [CDR15]. Argobots [SAB^{+}18]. AROMa [GAB18]. Array [HWZE10, IGEN11, KGI17, KZK^{+}19, KZK^{+}20, PH18, WQZ^{+}15]. Array-Based [PH18]. Arrays [CLLX18, JWJS14, LZC^{+}12, LLD^{+}18, SVK^{+}19, WHH^{+}13, WLX13, XS10, YLL^{+}17, ZZG^{+}11]. Arrivals [ABBCT16, KMM13b]. Articles [Sto10f]. Artificial [LLK^{+}14]. Ary [XS11, YLM^{+}15, LL12]. ASAP [QLNN13]. ASM [LXHS12]. ASN [CJW^{+}15]. Assembly [LPMB13, MTY^{+}12]. Assessing [APCH^{+}11, CP17a]. Asset [BN12]. Assignment [CTA14, CAJ^{+}16, CYC^{+}15, CZL^{+}18, CLHK11, GZY^{+}15, GHW^{+}16, HxjGG19, JSC^{+}17, JRP^{+}10, LC15, NN13, NLGQ14, PSMD18, RCV^{+}13, RGPH15, WYL19, WZQ10, YWC11, ZT14, ZJZ^{+}16, ZJTZ14]. Assignments [AD19]. Assimilation [ELX^{+}11]. Assisted [CCS^{+}12, CMG^{+}14, HWC^{+}14, LAMJ12, LFLW10, LSL^{+}10, SAM14b, SLLL14, SLLZ16, WMT^{+}11, YWC11]. Association [XLM^{+}11a]. Associative [QZW14, YMG15]. Associativity [DK17]. Assumption [XS11]. Assurance [RQZ^{+}16]. Astro [CC17]. Astronomy [FJV^{+}18]. Asymmetric [CLJ11, CRC^{+}17, GCN^{+}14, SHM^{+}12, TLSL15]. Asymmetrically [HZW^{+}19]. Asymmetry [QGPZ13]. Asymptotic [FWJ18]. Asymptotically [AD19]. Asynchronous [AR10, CJH^{+}14, CLSZ12, DGFRR18, HH11, HYC^{+}12, JMA^{+}18, JZZ^{+}15, LRYJ17, LJL^{+}11, Lu14, MGB18, SJVR17, SLG10, SPH^{+}18, WGG^{+}18, YHC^{+}13, ZGGW14]. At-a-Glance [LLY^{+}17]. Athanasia [JHYK11]. Atmospheric [AC19]. ATOM [DL17]. Atomic [GLGLBM13, IPQ19, LAFA15, ZCZ^{+}12]. Attached [WWH13]. Attached-RTS [WWH13]. Attack [MS12, TJH^{+}14, WMGA15, WXYX14]. Attackers [YCTC13]. Attacking [HLY10]. Attacks [ALLR14, AGG17, CDS15, CQZ^{+}12, DMT12, HPG14, LJG12, QLC13, SILJ11, WXTL13, Wu14, XTXH13, XSTZ10, YYY^{+}14, YZDJ11, YZJ^{+}12, YLR12, ZYL^{+}17, ZFG^{+}10]. Attention [SytL19]. Attention-Based [SytL19]. Attribute [CLH^{+}14, GZZ^{+}13, HSMY12, HN11, Hur13, LYZ^{+}13, LHL^{+}14, RZW^{+}13, SYL^{+}16, XWLJ16, XWS17]. Attribute-Aware [RZW^{+}13]. Attribute-Based [CLH^{+}14, GZZ^{+}13, HSMY12, HN11, Hur13, LYZ^{+}13, LHL^{+}14, SYL^{+}16, XWLJ16, XWS17]. Auction [CZWZ14, Guo14, HLeS^{+}15, JWNS19, LZY^{+}18, LYZL18, SWL17, TLL^{+}16, WKW16]. Auditability [WWR^{+}11]. Auditing [Rao14, SYZ18, Xia14, YJ13]. Augmented [GFJT19]. Aurora [LdSB19]. Authenticated [HCL^{+}14, LY16b, TW14, YLW13]. Authentication [DBAT11, FLH13, HXC^{+}11, LLG15b, LNZ^{+}13, LZCK14, LNXY15, LLZ^{+}12b, NLY15, RWLL14, RSN14, SGC14, ZLDC15]. Authority [LXXH16, LNXY15, YJ14].

Authorization [KB13, MSSB14, SYL^{+}16]. Authorized [LLC^{+}15, Rao14]. Auto [BHS^{+}19, BYZ^{+}16, CC17]. Auto-Generation [CC17]. Auto-Scaling [BHS^{+}19]. Auto-Tuning [BYZ^{+}16]. AutoBoT [VS19]. Autogeneration [ZM13]. Automata [DBG^{+}14]. Automated [CCW^{+}12, LZL10, LNL^{+}19, RAS17, TPRH16, ZJLG14]. Automatic [GETFL14, HWF18, JEW^{+}18, LMVS11, RKZC14, STK^{+}19, TDL^{+}19, VGMA10, ZLJ^{+}15a]. Automation [HH15]. Autonomic [CSW^{+}12, LGJZ16, PKS14, PVQ15, VLRP15]. Autonomous [BKH18, PJC^{+}13, YSDQ11, YQ11]. Autonomy [GLZ11]. Autonomy-Oriented [GLZ11]. Autopipelining [TG13]. Autoselection [KKE19]. Autotuning [GIX^{+}12, KKE19, KTD12, ZM13]. AUVs [YQ11]. Availability [AKT^{+}15, CL13, CZX^{+}19b, FHW11, JKVA11, KKC17, LSL^{+}17, MWJ16, TP14, YJC^{+}16, ZYZC12]. Available [AEM17, SBC^{+}10]. Average [CIH13]. Avionics [HL12b]. Avoidance [SCC11, TLP16]. Avoiding [KZW17, SOA15]. Aware [AAB16, APJ^{+}16, ADZZM15, Amm12, ABN19, ARM16, BBCB15, CAD^{+}18, CJ16, CAJ^{+}16, CCT10, CTX^{+}12, CGH13, CLHW13, sCCyW14, CLYR16, CCH^{+}17, Che18b, CLY^{+}19, CNC^{+}14, CL15, CZL^{+}18, CZX^{+}19b, CZD^{+}19, CVM^{+}15, CLKR15, CTP^{+}17, CHLY18, DN19, DGF12, DGI^{+}19, DLZ^{+}14, DZLC15, EHNS13b, ERG^{+}17, GTS^{+}15, GAB18, GHZ15, GHZZ16, GYLW18, GGF^{+}14, Guo14, HLZY15, HWG^{+}19, HAZ17, Has16, HWS16a, HWS16b, HWL^{+}17a, HV11, HJZ^{+}12, HL12b, HjZ^{+}14, HXLF15, HC14, HT16, HPP15, IZA18, JWK^{+}16, JMS^{+}18, JKP12, KAA16, KZW17, LMM18, LSL^{+}14a, LC15, LMZG15, LCG^{+}16, LRYJ17, LGM^{+}17, LZL^{+}18, LWSM19, LIWJ15, LSN19, MZL^{+}19, MNG^{+}15b, MTMR18, MMSS15, MKVL12, MDZC14, Pan14, PAB13, QF14, RBM15, RH16, RG17, RSSC15, RHDL11, RZW^{+}13, RLY^{+}15, SAEH19, SHG13, SWT^{+}17, SL13, SLW15, SLSG18, SLSG19, SZR17, SVK^{+}19, SBMA15, SJ14, TYLG13, TLP15, THT^{+}15, TOA13, VLRP15, WHH^{+}13, WWCZ11]. Aware [WWL11, WTL^{+}14, WSC^{+}14, WL14, WMZ^{+}15, WWZ^{+}16, WKW16, WGCG18, WCZ^{+}19b, WMC^{+}19, WDOX15, yWeH11, WYC^{+}15, WCD^{+}15, WMLJ17, XXLZ16, XXL^{+}19, XBZL17, XLT^{+}14, XFL15, XHZ^{+}13, YTL^{+}10, YLC^{+}16, YLL^{+}17, YGL^{+}15, YN17, ZTA^{+}15, ZWFX17, ZQL^{+}16, ZCG^{+}17, ZCC^{+}17, ZSSR19, ZHZC15, ZWL^{+}18, ZLL^{+}17b, ZYX^{+}10, ZWZ^{+}15, ZLZ^{+}16, ZHW^{+}19, ZXG^{+}19, Zou14, LSL14b, MCMR12, TLRW15]. Awareness [CSY16, LGJ^{+}17, PFMR13, RKGS16, ZHS^{+}19]. Axis [OMMZ14].

Back [WX15, YY14]. Back-Propagation [YY14]. Backbone [DWX14, DWY^{+}13, WTL^{+}14, ZWLL12, AO12]. Backed [CSC16]. Backend [XGL^{+}16]. Backfilling [GLRT18]. Background [LLRP18]. Backup [TWW^{+}18, XLL^{+}18, XLT^{+}14]. Bag [OPM^{+}15, VS19, TLH^{+}14]. Bag-of-Tasks [OPM^{+}15]. Balance [HLCH11, LX10, PCFP16, RKGS16, SSPG17, ZWL^{+}15]. Balanced [CHLC15, HWJ18, HJPL14, HW13, LHCM^{+}17, RZH^{+}11, WPT10, WWJ^{+}18]. Balancing [APG12, CMG17, CL16b, CLHK11, DY17, FSSZ16, GKL^{+}17, Gua14, HT16, HPP15, ITW^{+}14, Jia16, KKK^{+}15, KTK11, LGOB17, LSW17c, NOR16, Ren14, RRS12, SPS18, SLS^{+}16, Tse13, WTXG19, YGL^{+}15, ZYZC12, ZLJ^{+}15b, ZYW^{+}16]. Ballooning [LJL^{+}15]. Band [AA14, LKD10]. Bandit [ZLYL19]. Bands [YTW^{+}19]. Bandwidth [CIP^{+}17, CFLL18, CJW^{+}19, DG15, HX10, HKH^{+}10, LGL^{+}18b, LHM12, SHG13, SHY14, SAA17, SL16, TWL^{+}15, XLSR13, YSS^{+}17, ZJZ^{+}16, ZHS^{+}19, LLZ^{+}12b]. Bandwidth-Aware [SHG13]. Bandwidth-Efficient [LLZ^{+}12b]. Bandwidth-Intensive [ZJZ^{+}16]. Bandwidth-Optimized [HX10]. Bandwidths [LMM18]. Bank [YYL^{+}17]. Bargaining [WS14]. Barnes [ZBS15]. Barrier [AFA12, CJW^{+}15, LLK^{+}14, XLLZ11]. Barrier-Based [CJW^{+}15]. Based [AHSH^{+}16, AAB^{+}17, AWZ15, ABLS16, AGG17, APCH^{+}11, ACV17, BCQ^{+}10, BJ13, BWB^{+}19, BCF13, BTG^{+}18, BZA10, BDLS13, CJW^{+}15, CHCC14, CATC11, CSZ^{+}12, CTX^{+}11, CCKF15, CL14, CLH^{+}14, CYC^{+}15, CHD^{+}15, CCLW15, CSSL15, CP15, CCT16, CCCY16, CYW^{+}18, CHHK19, CH13, CFJ15, CLZ^{+}18, DMR16, DA16, DT14, DTLC19, DCA^{+}16, DWY^{+}13, ECW^{+}18, ET10, EHWX10, EHI11, EN12, ESGQ^{+}13, ERG^{+}17, ERRG18, FC10, FCD^{+}13, FFMR10, FC18, GG13, GTM^{+}17, GGY^{+}19, GRUMG17, GZZ^{+}13, GBFS16, GHZZ16, GZW^{+}18, GFJT19, GHL14, HWC15, HST^{+}11, HSMY12, HLZY15, HZJ16, HMP^{+}19, HWF18, HX10, HCZ12, HLWV14, HPG14, HYX11, HCL^{+}14, HLY^{+}14, HN11, Hur13, IZA18, IvS10, JWE15, JGG^{+}11, JLW^{+}10, JTS^{+}11, JJW11, JZH^{+}14, JWY^{+}18, KHN16, KZW^{+}12, KKW15, KKY^{+}14, KPG^{+}12, LSW17a, LM17, LW11, LJ16, LLLG13, LPP13, LC10, Li13, LYZ^{+}13, LHL^{+}14, LWY^{+}15, LW15, LY16a]. Based [LSLD17, LZH18, LWY^{+}19, LGZ^{+}19, kL11a, LWCG10, LT12, LW14, LLLC17, LZN10, LNA^{+}13, LJB^{+}13, LNZ^{+}13, LWZ^{+}13, LNXY15, LZW^{+}17, LNMMA15, LAFA15, LLG14, MWZ^{+}14, MMYES^{+}18, MGB18, MS12, MWZX14, MA14, Mis14, MPS15, MY11, MMSAZ11, MGR12, NSLV16, NOR16, NML^{+}14, NLY15, NLC12, NFFK14, NTK^{+}15, NSY^{+}16, OOA^{+}14, PFAF16, PH18, PGP^{+}17, PSMD18, QZW14, QCZ^{+}15, QFZZ15, RMG14, RVCT15, RSSC15, RZW^{+}13, RGLDM17, SDV18, SG16a, SY17, SKGC14, ST10, Seh15, SL13, SLGW14, SLC15, SSM^{+}18, SytL19, SCC11, SP15, SS17, SPB^{+}10, SKA15, SYXL16, SZZF10, SWC^{+}14, SYL^{+}16, SJ14, TXWL11, TJH^{+}14, TWW^{+}15, TWW^{+}18, TBÁ^{+}19, TXL^{+}14, TWSW17, TZC19, TNLM17, TKR14, TLSL15, TBC12, TCDMRP17, TCZL11, TFLL18, TRD13, Van14, VM12, WH16, WTTH17, WHH^{+}13, WKK11, WYW13, WPKL13, WJTZ14, WJWX14, WSC^{+}14, WSWY15, WM15, WHB16, WZHZ16, WLC^{+}17]. Based [Wan19, WVM19, WTXG19, WMC^{+}19, WXY^{+}13, WJB14, WMLJ17, WWH^{+}17, XX16, XWH15a, XWH15b, XBZ^{+}16, XTXH13, XHHC13, XHG15, XTHD10, XLLZ11, XLM^{+}12b, XSYY13, XWLJ16, XVC17, XWL^{+}19, XDLZ19, XSTZ10, YLSQ13, YL10, YGL13, YLW^{+}14, YRLY16, YPL^{+}17, YLJ^{+}17, YJC^{+}16, YCMX17, YTW^{+}19, YZS13, YWW^{+}15, YQH16, YPL13, YK14, YCW12, ZJL^{+}12, ZY13, ZLN^{+}13, ZGGW14, ZYW^{+}14a, ZWWF15, ZGL^{+}15, ZQCZ16, ZD16a, ZYL^{+}17, ZJL^{+}17a, ZLS^{+}18, ZX13, ZL14, ZJZ^{+}16, ZYW^{+}16, ZYT^{+}15, ZHX^{+}19, ZFG^{+}10, ZCX^{+}14, ZLYL19, ZD16b, ZASA10, ZFF16, ZBK^{+}15, dSLMM11, dCAB19, BAAT16, EALM15, HDL^{+}15, KLL^{+}17, LSL14b, WDL^{+}17, WMJ^{+}19, XWS17, UBC13, Che18b]. Basis [MKSN18]. Batch [CSW^{+}12, JAJ^{+}19, KMM13b, LNL^{+}19, LNK17, MZK19, SVC12, ZYL^{+}16]. Batched [HAZ^{+}18, JCW^{+}19, KAGD16, ZTZ^{+}18a]. Batching [CZWJ18, WW13]. Battery [LSL^{+}17, TWT16, YJCQ15]. Bayes [ZYW^{+}16]. Bayesian [AF19, WQZ^{+}16, YGL13]. Bayesian-Inference-Based [YGL13]. BCube [Guo17]. BCube-Like [Guo17]. Be [Hen14, SA11, VGMA10]. Beacon [LMSRSR12, TMMN15]. Beacon-Enabled [TMMN15]. Beaconless [ZS10]. Beamforming [SG16b]. Beat [Wu14]. Beats [TGNA^{+}13]. BECAN [LLZ^{+}12b]. Beehive [LL17]. BEES [AO12]. Before [XLL^{+}18]. Behavior [XHX^{+}13, XTXH13]. Behavioral [PLZW14, ZLJ^{+}15b]. Behaviors [DIAR16, ZZH^{+}17]. Belief [GG13]. Bellman [BB16]. Benchmark [HBS^{+}16]. Benchmarking [HCA16, RSW^{+}17, TFPK13]. Benefit [SME10, WZSL12, XZH14]. Benefits [HN10]. Best [GHW^{+}16, LS17a, MLT^{+}13, MPHR17, QGZP17]. Best-Effort [MPHR17, QGZP17]. Best-Harmonically-Fit [GHW^{+}16]. Better [CP15, LZWY14, LGJ^{+}17]. Between [AAB^{+}17, ZLJL17, TDL^{+}19, ZYZC12]. Betweenness [JSK18]. Beyond [YHL^{+}18, ZH11]. BFS [BB15]. BFS-4K [BB15]. BGP [BKL11]. Bi [CLB^{+}19, MZL^{+}19]. Bi-Index [MZL^{+}19]. Bi-layered [CLB^{+}19]. Bias [CP17a]. Biclustering [Yan14]. Bidding [DM11, LLLZ16, TYWL14]. Bidiagonal [LKD10]. Big [CHW^{+}17, CLT^{+}17, CLO^{+}18, CHHK19, CDPM18, CSR^{+}17, DLZH16, DED^{+}19, DZLC15, JZW^{+}17, KAV^{+}17, LGM^{+}17, MPM17, MNG^{+}15b, MDZC14, NCM^{+}17, Rao14, SMB^{+}18, SYZ18, TFLL18, VPS17, WMWW19, XXLZ16, XXL^{+}19, XBZL17, XL17, XZJ^{+}19, XWL^{+}19, YJR15, YLZ^{+}15a, YGS^{+}19, YWZ17]. BiGNoC [CDPM18]. Bijective [CFJ15]. Billion [ZML^{+}17]. Billion-Node [ZML^{+}17]. BiloKey [MZL^{+}19]. Bimatrix [RMG14]. Bin [LTC16]. Binary [KAC^{+}15, SS17, WZFG13, YRLY16]. Binary-Tree [SS17]. Bioinformatics [EGQ11, SJVR17, SJVR19]. BioinspirEd [AO12]. Biological [MC10, dOSdM13]. Biology [TYS^{+}12]. Biomedical [LAT^{+}15]. Biophysical [OOA^{+}14]. Bipancyclicity [CH15, XS11]. Bipartite [ABP17].

Bipartitioning [SAA17]. Bisection [AA14]. Bit [BKL11, KKK11, TTG^{+}15b]. Bit-Representation-Optimized [TTG^{+}15b]. Bit-Split [KKK11]. Bitier [CGH13]. Bitplane [EALM17]. Bitsliced [HMKG19]. BitTorrent [CL13, CNMA11, IRPvdS12, LXBZ13, SYL^{+}14, ZDWR11]. Black [SZL^{+}12]. Blending [FGEL14]. Blind [CZZ^{+}16]. BlindDate [WML15]. BLISS [SLS^{+}16]. Block [AAW^{+}17, BKH18, KN16, NVBH18, QFZZ15, ZL14]. Block-Space [NVBH18]. Blockchain [JWNS19, LWY^{+}19, XWL^{+}19]. Blockchain-Based [LWY^{+}19, XWL^{+}19]. Blocking [DLA^{+}18, HTZY17, NFD10]. Blocks [CL13, FWH^{+}18, LTGI16, SY17]. Bloom [RCM16, AKC^{+}15, GHL14, LWL^{+}19, MLVD12, QZW14, QLC14, XH10, ZS17]. BloomCast [CJL^{+}12]. BLOT [ZLYL19]. Blue [CSR^{+}17, IBC^{+}11, ZYL^{+}16]. Body [CH13, LZCK14, RQZ^{+}16, ZWWF15, ZQH13]. Bodyguard [FDFZB13]. Boltzmann [TS18]. Boost [HWQ^{+}15]. Boosting [HPPR17, HWS16a, LCY^{+}17, WSH^{+}19]. Bootstrapping [SAH15]. BOT [LMPR12]. Both [NZWL14]. Bottleneck [RTZ^{+}18, ZSSR19]. Bottleneck-Aware [ZSSR19]. Bound [Che11, CBF^{+}17, DTLC19, HZW^{+}14, HTZY17, HWG^{+}19, HCyW^{+}17, LZ10, WYX13, XCZ^{+}15, ZLN^{+}13]. Boundaries [DRK11]. Boundary [WJTZ14]. Bounded [Agr14, BV10, CZL^{+}16, DC18, GS17, LAV^{+}10, LMSRSR13, LLY^{+}17, ZGY15]. Bounded-Degree [LMSRSR13]. Bounded-Reorder [ZGY15]. Bounding [DMT12]. Bounds [AH10, CYX^{+}14, FWJ18, Fre13, HGS^{+}19, YNKD18]. Boxes [SZL^{+}12]. Branch [CBF^{+}17]. Branch-and-Bound [CBF^{+}17]. Branching [YLSQ13]. Branching-Router-Based [YLSQ13]. Breadth [BBM16]. Breaking [LKM10]. Bridge [LLY^{+}15, LYZL18]. Bridging [AAB^{+}17]. Brief [YZS13]. Broadband [IG11, KBS11, LLK13]. Broadcast [AMN^{+}16, ATACA18, BV10, CCFS11, CLA^{+}19, GSH^{+}19, HWI12, JLM^{+}12, LRT19, MR16, SLM^{+}10, Tou15b, Tou15a, WTL^{+}14, XL16, ZD12, ZLZ^{+}14]. Broadcast-Oriented [ATACA18]. Broadcasting [Agr14, FCD^{+}13, LC10, YW10]. Broker [LLLH19, TKR14]. Broker-Less [TKR14]. Brokerage [WNLL15]. Brokers [MLT^{+}19]. Brooks [Kum14]. Browsing [SLLZ16, ZHZC15]. Bruijn [FMY^{+}18]. BSN [LQK^{+}13]. BT [DR16]. Budget [ABN19]. Buffer [DSJ16, LDLL18, LN17, NFD10, SML13, TLH^{+}14, WYX13, ZFF16]. Buffered [CCLW11, LY11, XHC16]. Buffering [CJZ12]. Bufferless [SKL^{+}15]. Buffers [LHM12, LW14]. Bug [CWLS19, DDW^{+}19]. Bugs [LPZ12]. Building [CZRB18, FKMC15, NZM^{+}16, ZMTL15, ZLL^{+}17b]. Bulk [VBC19, ZGH14]. Bulk-Data [ZGH14]. Bump [TLJ^{+}14]. Bump-Aided [TLJ^{+}14]. Bumping [TLJ^{+}14]. Bundles [CC10]. Burrows [WH16]. BURSE [YLZ^{+}15b]. Burstiness [ZQL^{+}16]. Burstiness-Aware [ZQL^{+}16]. Bursting [CCNMF18, Zom14]. Bursts [LVD11]. Bursty [MTDD17, WMLJ12, YLZ^{+}15b]. Bus [HWZE10, WSC^{+}14]. Bus-Based [WSC^{+}14]. Bypass [ZPD11, ZD12, ZDF^{+}15]. Byte [NTDZ19]. Byte-Addressable [NTDZ19]. Byzantine [ALLR14, MT15, MR16].

C [AFT^{+}16, PB19, TFPK13]. C-MART [TFPK13]. CA [RMM16]. Cable [TFKN17]. CACAO [YWC11]. Cache [APPG16, AJM12, CAD^{+}18, CGH13, CP17c, Dan11, DGI^{+}19, FPGAD10, HLS^{+}12, HWS16b, HWL^{+}17b, HCJ^{+}10, KZW17, KPB19, KAC^{+}15, LSL^{+}14a, LZH18, LGJ^{+}17, MWJ^{+}14, MV16a, NVS16, PD14, PCP14, RH16, RLY^{+}15, RJ16, SEA18, SPS18, TLH^{+}14, WHH^{+}13, WHC^{+}14, WMLJ17, XX16, YLL^{+}17, ZJS12, ZSW^{+}19, ZH18]. Cache-Aware [DGI^{+}19]. Cache-Coherent [MWJ^{+}14, RH16]. Cache-Oblivious [LZH18]. Cache-Optimized [ZH18]. Cache-to-Cache [Dan11]. Caches [AHS^{+}15, AFMM17, DKS^{+}15, MVL15, MV16c, NVS16, SSPG17, ZML13]. Caching [AKC^{+}15, ARM16, BJ13, CE17, DD11, DSASSLP12, ET10, GKKW16, HN10, HGC12, HLWV14, HGL^{+}16, LGZ^{+}19, SHA19, WHF^{+}19, WCF13, WML14, XX16, YLC^{+}19, ZZCD10, ZTZQ19]. Calculating [AI15]. Calculation [MYPL18]. Calculations [AAW^{+}17]. Calibrate [XYT^{+}15]. Calibrating [BCTB13, XYT^{+}15]. Calibration [LGXL19]. Call [Ano11d, Ano11c, Ano12c, MSB11]. Calls [TTG^{+}15a]. CAM [EHI11]. CAM-Based [EHI11]. CAMF [WDOX15]. Campus [MBH^{+}10]. Can [LGOB17, WZSL12, Wu14, XSZ^{+}10, XZH14]. Canceling [QPB^{+}17]. Cancellation [LWY^{+}13]. Capability [LNA^{+}13]. Capacitated [XLX^{+}16]. Capacity [CHTW12, HLS^{+}15, JCLJ12, LG13, MLVD12, QTC^{+}14, RX11, SKL^{+}15, WBPF11, Wan14, WSL^{+}15, XHC16]. Capping [ZMW17]. Capture [HCY^{+}12, RCC^{+}14]. Carbon [ZLZ^{+}16]. Carbon-Aware [ZLZ^{+}16]. Card [HCL^{+}14]. Cardinality [ABP17, QNLN11]. Carlo [NSLV16, OZMC^{+}16]. Carried [LCL^{+}15]. Carrier [KZW^{+}12, SCC11]. Carrier-Sense-Based [SCC11]. Carry [ZLL17c]. Carry-in [ZLL17c]. Cartesian [CLH13, CH15]. CAS [AH10]. Cascadia [ZL10]. Cascading [HWNS15, YQZC12, ZYSH14]. Case [HAZ^{+}18, Ian14, JKS13, PKG14, PS19, SJVR17, VMN^{+}16, WGHP11, dCAB19].

CASER [TLRW15]. Cashing [HLS^{+}12]. Cast [Wan14]. Catching [WFK^{+}12]. Cauchy [LWCL18]. Causal [HK18]. Caused [LLXC12]. CCD [JHMV12]. CDC [LZK^{+}15]. CDN [LSCL16]. CDO [KBHS14]. CDS [DWY^{+}13]. CDS-Based [DWY^{+}13]. Cell [IG11, KBS11, SZA11, SA11, VGMA10]. Cellular [PKG14]. Censorship [SLLZ16]. Censorship-Resistant [SLLZ16]. Center [Bru14, CWC^{+}13, DY16, DY17, GXZ^{+}15, LYH^{+}15, LWLZ17, LYZ^{+}16, MBV11, PR19, QFZZ15, SJR17, SSW^{+}17, Sto11a, TP13, WWZ^{+}16, WTXG19, XWJX15, YJCQ15, YQ16, ZJLS12, ZRTL15, ZQWL17, ZDM^{+}17, ZDM^{+}19, ZMW17]. Centers [AA14, ABBCT16, BB13, CTP^{+}17, DGC17, FYH^{+}15, GKL^{+}17, GF13, GGF^{+}14, Guo17, HLCB^{+}17, HLL18, KMM12, KMMR13, KMM13a, KMM13b, LMM18, LGD14, LY16a, LLL18, LCA13, LGW^{+}17, PYHY16, Ren14, TZC19, WCLK12, WWX^{+}13, WW13, WZS^{+}19, XDMZ17, XFL15, YLC^{+}16, YHS^{+}14, YGL^{+}15, YWW^{+}15, YJC15, ZCJY14, ZRS18, ZSSR19, ZHL19, ZWY^{+}17, ZXG^{+}19, ZGKB16]. Centrality [JSK18, KI14]. Centralized [WWT^{+}19]. Centric [AHSK17, ASG^{+}14, ACNP11, CFLL18, GHL14, HL12a, LY16a, Ozd19, PG16, PCP14, QFZZ15, SMS^{+}13, WX15, YHL^{+}18, YXLJ16, ZBK^{+}15]. Certificate [LNA^{+}13]. Certificateless [LZCK14]. CFD [RMB^{+}16]. CFS [Tak14]. CGRAs [GYLW18]. Chaining [CTG^{+}19, JY15]. Chains [PHXL19]. Challenges [LHL^{+}13a, WWL^{+}15]. Chameleon [GZX14, BHS^{+}19]. Chance [TUS13]. Changes [LLXC12]. Changing [ACE^{+}19]. Channel [Bis18, Che14, CYC^{+}15, CGKP11, DWX14, GSH^{+}19, GCL14, Mis14, NZWL14, SDL^{+}15, TTH^{+}19, TLP15, TCS13, WZQ10, YTL^{+}10, YWC11]. Channel-Aware [YTL^{+}10]. Channel-Hopping [Mis14]. Channel-Oblivious [SDL^{+}15].

Channel-Related [TLP15]. Channelization [KL11b]. Channelless [SHG11]. Channels [XL16, ZSW^{+}15]. Chaos [LGOB17]. Characteristic [YDH17]. Characteristics [LLZ^{+}12a, MM15, MWJ16, MNE14, TP14]. Characterization [CSM^{+}13, CPH^{+}18, SEA18, WV17, WL12b]. Characterized [MP16]. Charging [TRT19, WPT17, YLH^{+}16]. Chasing [CRRR15]. Cheat [ZY14]. Cheat-Proof [ZY14]. Checkability [LHL^{+}14]. Checked [Hen14]. Checking [CGZQ13, LTW^{+}14]. Checkpoint [DRVC17, STK^{+}19, ST18, TZY^{+}18]. Checkpoint/Restart [STK^{+}19]. Checkpointing [MMBdS14, ST18, XDLZ19, ZXL^{+}17, ZLW^{+}19]. Chemical [AC19, KEGM12, LMVS11, XLL11, XLH^{+}15]. China [TDLR13]. Chip [AMN^{+}16, ATACA18, AJM12, AAB16, ADMX^{+}12, AF18, Bis18, CHM^{+}13, CLT13, CCH^{+}17, CIP^{+}17, Che18b, CDPM18, CP17c, DKM^{+}15, EHM^{+}17, HD15, HYZ15, HGC12, HRGE17, JWK^{+}16, JTS^{+}11, JKP12, LKBK11, LAMJ12, LDLL18, LWW^{+}13, LCL^{+}15, MB12, MVL15, Oru17, PL16, RKGS16, RAG10, SV19, SHG11, SHG13, SKL^{+}15, Sib12, TLP16, TWSW17, Tou15b, Tou15a, VNA^{+}16, WMW11, WWJ^{+}18, WYZ^{+}19, YLJ^{+}17, ZMF10]. Chip-Multiprocessors [CIP^{+}17, CP17c, EHM^{+}17]. Chips [JIP14, TWSW17, WSC^{+}14]. Cholesky [HAZ^{+}18, HWC15, KAGD16]. Chord [YL11b]. Chromosome [dOSMM^{+}16]. Chromosome-Wide [dOSMM^{+}16]. Chunk [SLL13a, dSLMM11]. Chunk-Driven [SLL13a]. Churn [BBR12, LXHL11, YCWL14]. Churn-Resilient [LXHL11]. CIACP [YLL^{+}17]. Ciphertext [XWLJ16, XWS17]. Ciphertext-Policy [XWLJ16, XWS17]. Circuit [CRWY15, WYX^{+}19, XWS17, XWLJ16, YWWR18]. Circuit-Switched [Bis18]. Circuits [HA13]. Circulant [TWL12]. Cities [Iye14]. Claims [HWSX17]. CLAP [HHWZ17]. Clarify [WJX^{+}14]. Classification [AM19, Di 17, ERG^{+}17, ERRG18, JPG14, PT11, QP16c, RJ16, WXZ^{+}14, ZXW^{+}13]. Classifier [MKSN18, YDC^{+}17]. Classifiers [LG10]. Client [CSW^{+}17, LC15, LS17b, NVS16, NN13, WX11, YWC11, ZT14]. Client-Assisted [YWC11]. Client-Driven [CSW^{+}17]. Client-Perceived [WX11]. Client-Server [NN13]. Climate [AC19]. Clip [AZW^{+}19]. Clique [GLM13]. Cloaking [HX10]. Clock [BCQ^{+}10, CLSZ12]. Clocks [JZZ^{+}15, YNKD18]. Cloning [XLY^{+}17, ZSY14]. Clos [CMB18, WYL19, XHC16]. Clos-Network [XHC16]. Close [RGBC11]. Closer [YNK^{+}17, MCMR12]. Closure [ADMX^{+}12]. Cloud [AHSH^{+}16, ASBL15, AIAD^{+}18, ACC^{+}17, AGG15, AGG17, Ano11d, ASSB18, ABBCT16, BM12, BH13, BB13, BMR15, BHEP14, Bru14, CHLZ13, CWL^{+}14a, CL16a, CSC16, CL14, Che15, CWL16, CMG17, CLT^{+}17, CZX^{+}19a, CW19, CCT^{+}14, CCNMF18, CLZ^{+}18, CTP^{+}17, DDW^{+}19, DGC17, DHTZ15, DW13b, DZLC15, DL17, ECW^{+}18, EGQ11, FXL17, FCM14, GHZ15, GYQW15, GRJZ17, HLS^{+}12, HHWZ17, HLZ^{+}19, HWSX17, HH15, HLCB^{+}17, HCH^{+}19, HLWV14, HBS^{+}16, IOY^{+}11, ITW^{+}14, JRZ^{+}18, JWNS19, JRO^{+}17, KKE19, KMM12, KMMR13, KMM13a, KMM13b, KDCR19, LLJ^{+}13, LYZ^{+}13, LLC^{+}15, LCG^{+}16, LTC16, LLLZ16, LXXH16, LSB^{+}18, LGJ^{+}18, LHXP18, LSJ^{+}19, LT12, LS17b, LZWY13, LWZ^{+}13, LS14, LDYZ15, LNXY15, LH15, LSW16, LSW17b, LLL18, LSN19, LYZL18, MZLT19, MS12, Man18, MWJ16, MSSB14, MPM17, MNG15a, MLT^{+}19, MWZ^{+}13, MBMC13, NZM^{+}16, NSY^{+}16, PSL15, PR19, PHXL19, Rao14, RSW^{+}17, RM17, RTZ^{+}18, SDV18, Sam14a]. Cloud [SLG^{+}18, SL16, SZR17, SYZ18, SWC^{+}14, SYL^{+}16, TL14, TRT19, TFPK13, VLRP15, WWR^{+}11, WZSL12, WCRL12, WRWW13, WNLL15, WWL^{+}15, WLL15a, WKW16, WHGS17, WWW^{+}18, WGG^{+}18, Wan19, WCZ^{+}19a, WK11, WSS15, WCD^{+}15, Xia14, XWSW16, XSC13, XBZ^{+}16, XWJX15, XL13, XLT^{+}14, XWLJ16, XALS17, XZJ^{+}19, YJ13, YJ14, YJR15, YLZ^{+}15a, YPL^{+}17, YDH17, YLZ^{+}15b, YZH^{+}19, YSS^{+}17, YYL^{+}13, YY14, YWWR18, ZLJ^{+}15a, ZLZ^{+}17, ZLN^{+}13, ZYLC14, ZYQ^{+}14, ZSW^{+}15, ZQCZ16, ZWLW16, ZCG^{+}17, ZJL^{+}17a, ZZSZ18, ZJZ^{+}16, ZYW^{+}16, ZXL^{+}17, ZTZ^{+}18a, ZWL^{+}18, ZWZ^{+}13, ZWY^{+}17, ZLDC15, ZLZ^{+}16, ZXG^{+}19, ZZLL16, ZJ16, ZLW^{+}18, Zom14, WZS^{+}18]. Cloud-Backed [CSC16]. Cloud-Based [HLWV14, MS12, XBZ^{+}16]. Cloud-Edge [ZZSZ18]. Cloud-Friendly [WSS15]. Cloud-Service [WHGS17]. Cloud/Fog [JWNS19]. CloudArmor [NSY^{+}16]. Cloudde [ZLZ^{+}17]. CloudFog [LS17b]. Cloudlet [CCCY16, XLX^{+}16]. Cloudlet-Based [CCCY16]. Clouds [ALZ17, ABN19, BWB^{+}19, BLLP15, CB14, CPGT14, CZQ^{+}17, CRZH15, DNW^{+}16, DW13a, DG15, GS17, HCSC13, Jia14a, LPP13, LMZG15, LZY^{+}18, LH16, LLZ18b, MLXG19, MTY^{+}12, NMG15, PGP^{+}17, RG17, RSN14, SWL17, SB19, SCJ^{+}17, TRD13, TVRD17, WVT13, WLL15b, WUH^{+}17, Wu14, WWL^{+}17, WIZ^{+}17, XXLZ16, YWY^{+}17, ZTG^{+}18, ZQL^{+}16, ZHCL17, ZWG^{+}16]. CloudScout [YZT^{+}17]. Cloudy [TUS13]. Cluster [FHW11, HPH^{+}12, HWNS15, LNA^{+}13, LN17, LSW17c, LLG14, MB12, SWL17, TMMN15, WS18, WRB11, XHL^{+}11, ZWWF15, ZCG^{+}17, ZWJ^{+}19, Zou14]. Cluster-Aware [ZCG^{+}17]. Cluster-Based [LNA^{+}13, LLG14, ZWWF15].

Cluster-Head [TMMN15]. Cluster-on-a-Chip [MB12]. Cluster-Scheduling [WS18]. Cluster-Tree [HPH^{+}12]. Clustered [CLJ11, DHBB12, KP12, LHL17, PPS^{+}17, SLW15, WWLJ14]. Clustering [DO13, GBP17, GRB^{+}19, GV15, HMP^{+}19, JY15, JJW11, KHN16, PSMD18, SKA15, THE^{+}15, WXZ^{+}14, WSS15, XJ14, YN17, ZYW^{+}16]. Clustering-Based [JJW11, KHN16, ZYW^{+}16]. Clusters [BBK17, CAJ^{+}16, CZT^{+}17, CLO^{+}18, CRG^{+}17, CZL^{+}18, CHPY17, FA19, HLQ^{+}15a, HLQ^{+}15b, HLK^{+}19, JNL^{+}15, KZK^{+}19, KZK^{+}20, KOKA11, LZ12, LM17, LLY16, LS17c, LNK17, Man16, MVML11, MTY^{+}12, NZM^{+}16, Pan14, PGGS19, RGLDM17, dOSMM^{+}16, SJVR15, TMJ14, WW11, uRILP17, WCD^{+}15, XP12, XLY^{+}17, XL17, XZQZ17, YTMS16, ZM13, ZLW^{+}14, ZBS15, ZTZQ19]. CMP [APMG12, APG12, JHR^{+}14, ZJS12]. CMPs [DK17, ERG^{+}17, FPGAD10, AFA12]. CNN [LZWZ19, WCYL19]. Co [GHZZ16, HZJ16, JTS^{+}11, LGJZ16, MVC^{+}18, RSNV18, TZT^{+}16, ZZH^{+}17]. Co-Design [MVC^{+}18]. Co-Located [LGJZ16]. Co-Processor [TZT^{+}16]. Co-Processors [GHZZ16]. Co-Running [ZZH^{+}17]. Co-Scheduling [HZJ16, JTS^{+}11, RSNV18]. Coalesced [HTA10]. Coalescing [AFT^{+}16, GDM^{+}13, ST18]. Coalition [DMR16, Tak14, YZS13]. Coallocation [SME10]. Coarse [CA13, LLD^{+}18, YLL^{+}17, YLLW16, YYL^{+}17]. Coarse-Grain [CA13]. Coarse-Grained [LLD^{+}18, YLL^{+}17, YLLW16, YYL^{+}17]. CoCloud [ECW^{+}18]. Code [AAH15, BKH18, DLZ^{+}14, FGJ^{+}15, LT10, LT12, MG18, MLK15, SSF16a, TTG^{+}15a, ZLL^{+}17b, ZR18, ZWL^{+}16a]. Code-Based [LT12]. Codec [GIP^{+}13]. Coded [CZT^{+}17, FSSZ16, HWQ^{+}15, HLQ^{+}15a, HLQ^{+}15b, KN16, LLRP18, LNK17, She14, SSF16b, SSLF17, SLSG18, SLSG19, WPMX18, ZLL17a, ZHL19, ZLX^{+}14]. Codes [AGG15, KZK^{+}19, KZK^{+}20, KBHS14, LL17, LC14, RMG18, SGGB14, WXLY16, ZM13, ZL14]. Codesign [AJM12, HGY^{+}14, LTW^{+}14]. Coding [AGG17, CL13, CL14, CHD^{+}15, CWL16, CW19, CFLL19, EALM17, JN16, KLWK12, KKW13, KKW15, KL11b, LLLG13, LG13, LGYV14, LHL17, LLK13, LWCL18, NL11, PPR10, TYLG13, TYG^{+}14, WWL^{+}13, WTL^{+}14, WL14, WXYX14, XSZ13, YW10, YY10, YWJJ11, ZJL^{+}12, ZGXJ14, ZL11]. Coding-Aware [TYLG13]. Coding-Based [AGG17, CHD^{+}15, KKW15, LLLG13]. Coefficient [EALM17, YZJ^{+}12]. Coexploration [LLCH12]. Coflow [LYZ^{+}16, WYX^{+}19, ZSSR19]. Coflows [WZS^{+}19]. Cognitive [AKP14, CJH^{+}14, CLM^{+}15, DWX14, HWC^{+}14, JZY^{+}15, LCL^{+}14, LLCL12, MS13b, Mis14, WJTL13, XJL^{+}14, ZY14]. Cognizant [ZSB^{+}13]. Coherence [CAD^{+}18, CRD11, FPGAD10, KPKH16, LSL^{+}14a, RAG10, RJ16, TGFFP^{+}19, YCMX17, dLMPG19]. Coherent [MWJ^{+}14, RH16]. Collaboration [ECW^{+}18, SLG10, XXLZ16]. Collaboration- [XXLZ16]. Collaborations [LTW^{+}14]. Collaborative [BZA10, CC15, HFY^{+}14, HL12a, LCL^{+}14, LLAL18, LC11, LS14, MZLT19, MM10, SLLZ16, WXL10, WHF^{+}19, WLL^{+}19, WUM10, XSZ13, ZFG^{+}14, ZCG^{+}17, MCMR12]. Collecting [XHL^{+}15]. Collection [CJH^{+}14, CHTW12, GFLL15, JCLJ12, JJW11, KPG^{+}12, LLL^{+}13, LZK^{+}15, RY14, WWL11, WMHX12, WLLL10, XSZ13, YQLS14, ZT13]. Collective [GHZ15, LS17d, NCM^{+}17, dBK11]. Collectives [Trä19]. Collision [QLNN13, SCC11, SHF^{+}17]. Collision-Mitigation [SHF^{+}17]. Collisions [KWZ^{+}12]. Collusion [SLSL16, ZJ16, WZS^{+}18]. Colocating [PR19]. Colocation [XTFC17]. Color [Has16]. Coloring [CH13, LRT19, WYLH18]. Coloring-Based [CH13]. Colorings [LHCM^{+}17]. CoMan [LGL^{+}18b]. Combination [TGFFP^{+}19]. Combinatorial [QFZZ15]. Combined [KHK15, KKW18]. Combining [AHSK17, AFT^{+}16, ME15b]. COMIC [YZL^{+}15]. Comment [CL16a, FYH^{+}15, Man16, RCM16, VS11a]. Comments [WZS^{+}18, XWS17]. Commerce [WMGA15]. Commercial [FPF13]. Commodity [MYPL18, VNA^{+}16]. Common [DWX14, YXSS13]. Communication [APMG12, AVA^{+}17, ABF12, ACS13, ACV17, CL17, CBK^{+}10, CCK12, CCV19, DSM19, GHZ15, JKP12, KOPS10, KGR16, LRT19, Li13, LQK^{+}13, LWSM19, LGG^{+}14, MS13a, MG18, PT15, RCK15, RGLDM17, Seh15, SLGW14, SPS18, VS15, WSC^{+}14, WMLJ12, YN17, YDC^{+}17, YLT15, ZSH^{+}11, ZHQ12]. Communication-Aware [JKP12, YN17]. Communication-Efficient [DSM19, YLT15]. Communication-Optimal [YDC^{+}17]. Communications [CJW16, CCD^{+}15, GZX14, GCL14, LAK11, LZH18, LA12, LLL^{+}12, PDFJ13, XLM12a, Zhu14]. Communities [JRV^{+}13, OMMZ14, RKZC14, WZSL12]. Community [ADZZM15, BJ13, DO13, GLM13, LS17d, LH17, LSW^{+}15, SM16]. Community-Based [BJ13]. Comparative [LJL^{+}15, SJVR19]. Comparing [DD17, WGHP11]. Comparison [Di 17, EN12, WKK17, ZD16b]. Comparison-Based [EN12, ZD16b]. Compensation [ZWL17]. Competition [CRZH15, CE10]. Competitive [XLY^{+}17]. Competitiveness [NVBH18]. Compilation [MGS12]. Compiler [DED^{+}19, LAMJ12, MRH^{+}16].

Compiler-Assisted [LAMJ12]. Complement [Van14]. Complete [FLH13, Has16, LVA^{+}11, LG10, LXZB15]. Completion [LGM^{+}17, LLpC15]. Complex [CWZ^{+}15, LLZ14, MSS17, PGGS19, TXZ^{+}11, WYLH18, KLL^{+}17]. Complexities [LC14]. Complexity [CWC11, JTS^{+}11, KKW13, NL11, SLS^{+}16]. Component [GCG^{+}18, HHWZ17, PB12, YLW^{+}14, ZLS^{+}18]. Component-Based [YLW^{+}14]. Component-Level [HHWZ17]. Components [JFP^{+}17, LCD^{+}17]. Composing [TW14]. Composite [LAV^{+}10]. Composition [CP15, DZLC15, HJS^{+}11, KN12, MZLT19, SCL^{+}15, TCZL11, WMC^{+}19, YWZ17]. Comprehensive [LHD^{+}14, uRILP17]. Compress [DC18]. Compressing [LTM11]. Compression [CMK^{+}16, DC18, DTLC19, EALM17, KGK^{+}13, MV16a, Tan12, TDL^{+}19, VPS17, WHB16, ZLT^{+}18]. Compressive [CIH13, LZK^{+}15, LLH^{+}15a, TVG13, XJ14, ZYT^{+}15, ZHX^{+}19]. Computation [CWL14b, CATC11, Che15, CJW^{+}19, CIH13, DHTZ15, FWZ^{+}16, GCG^{+}18, HCH^{+}19, HZW^{+}19, KGI17, LMLM13, LZWZ19, LMFS11, LCY^{+}17, LCD^{+}17, SG16a, SHY14, TZT^{+}16, WTTH17, XAG17, XVC17, YTMS16, ZGGW14]. Computational [CCJ19, CL17, CB13, TYS^{+}12, WZGR10]. Computations [AK18, ARM15, BBP17, CT12, DW10, HWSX17, SZA11, WGG^{+}18, ZGGW13, ZR18]. Compute [DGI^{+}19, WV17, WVM19]. Computer [BKF^{+}16, CMVB17, CP17b, LNK17, MA13, TKC^{+}15]. Computers [DDP^{+}19, EAMEG11, HZJ16]. Computing [ABE^{+}11, ACC^{+}17, Ano11d, ABBCT16, ABP17, BKK11, BM12, BH13, BBD^{+}19, BMR15, BHEP14, BBL^{+}16, Bru14, CHLZ13, CPGT14, Che15, CLYR16, CLT^{+}17, CLO^{+}18, CDPM18, CDR15, DCC^{+}19, DHTZ15, ELX^{+}11, GDRTS16, GG11, HP14, HMP^{+}19, HCH^{+}19, HZW^{+}19, HYC^{+}12, HKkY^{+}16, ITL17, IOY^{+}11, JFP^{+}17, JWNS19, JRO^{+}17, KMM12, KMMR13, KMM13a, KCW11, KPB19, LGOB17, LZ11, LMD16, LYZ^{+}13, LTL14, LCG^{+}16, LLLZ16, LGL^{+}18b, LS14, LNXY15, LSW17c, LM16, LNMMA15, LLS13, LHCM^{+}17, MSSB14, MLT^{+}19, MC10, MWZ^{+}13, MBMC13, MV16d, MVML11, MBH^{+}10, MRGR12, NLC12, OPM^{+}15, PH11, PDH10, PH12, RFZ11, RMG14, RM17, RLVTMG^{+}16, SMB^{+}18, SWT^{+}17, SCL^{+}15, SYZ18, Sto10f, SP12, TKS11, TNH^{+}18, TFM^{+}16, TP14, VLRP15, WWR^{+}11, WWL^{+}15, WLL15a, WKL^{+}16, WMWW19, WDL^{+}17, XSC13, XLL11, XLH^{+}15, XWLJ16, YJ13, YHC^{+}13, YC18, YYK^{+}11b, YLZ^{+}15b, YL16, YY14, ZQL^{+}16, ZWLW16]. Computing [ZTH17, ZXL^{+}17, ZTZ^{+}18a, ZLDC15]. Concave [ZWLW16]. Concealed [CLLS12]. Concurrency [AA12, CWLS19, GBD^{+}13, GTT^{+}17, HYC^{+}12, LPZ12, MLC^{+}15, WCZ^{+}19a]. Concurrency-Bug [CWLS19]. Concurrent [Ara11, FCM14, HAU19, LS19, PLG19, SV19, ZWJ^{+}18, ZTZ18b]. CONDESA [THB^{+}14]. Condition [Dua95, LZL^{+}18, VS11a]. Conditional [Cha11, CH14, CLH13, LL12, LLG15b, LAT^{+}15, LKT11, LZXW15, LXZH16, XS11, YLM^{+}15]. Conditional-Fault [LKT11]. Conditions [LZ11, NX95, VS11b, WHH^{+}13]. Conduct [NCKL14]. Conferencing [ZLCZ14]. Confidence [WHYZ10, YL10]. Confidence-Based [YL10]. Confident [DWLY15]. Configurable [ZGL10]. Configurated [ZDF^{+}15]. Configuration [AAW^{+}17, BYZ^{+}16, BRX13, CHLZ13, CAKRY16, GKT^{+}17, HLZ^{+}19, LAMJ12, LLL18]. Configurations [LLLZ16]. Configurator [ZLJ^{+}15a]. Confirmation [CJW^{+}15]. Conflict [JEW^{+}18, KZW17, KB17, YYL^{+}17]. Conflict-Avoiding [KZW17]. Conflict-Free [KB17, YYL^{+}17]. Conflicting [ZLJL17]. Conflicts [CLL11, TGAG13]. Congestion [ESGQ^{+}13, ESGG^{+}15, LCL^{+}15, LA12, RKGS16, RHDL11, SX10, TLP16, THL13, TCT16]. CongraPlus [PLG19]. Conjugate [CCV19]. Conjunctive [SK14]. Connected [EHNS13b, HWC^{+}14, JFP^{+}17, LCG^{+}13, LHYW15, LCD^{+}17, MM10, WYX13, YNW13, ZLS^{+}18]. Connection [CFJ15, DLXS19]. Connectivity [BBCB15, HCS12, JLW^{+}10, LWZ^{+}15, LZXW15, LXZH16, WMT^{+}11, WJTZ14, ZH11]. Connectivity-Based [JLW^{+}10, WJTZ14]. Connectivity-Coverage [BBCB15]. Conquer [SZWX15, SYZ18]. Conscious [LZ11, XTHD10]. Consensus [AE12, CHCC14, CZL^{+}16, CGKP11, GBFS16, ZGL^{+}15]. Consensus-Based [CHCC14, GBFS16, ZGL^{+}15]. Consequence [ZBK^{+}15]. Consequence-Centric [ZBK^{+}15]. Conservation [WQZ^{+}15, WW13]. Conservative [CW15]. Conserve [CDBQ12]. Consideration [CJH^{+}14]. Considering [Che16, LHXP18, YTL^{+}19, YJC15]. Considers [WYZ^{+}19]. Consistency [CLC^{+}12, HK18, HBF12, HCJ^{+}10, LC15, LSCL16, RJ16, She10b, SL13, TZ10, WDH^{+}16, XHL^{+}11]. Consistency-Aware [LC15]. Consistent [AEM17, NX95, TGT10, TPRH16, USP^{+}12]. Consolidated [HPP15, KL16]. Consolidating [HMS^{+}18]. Consolidation [BB13, HLCB^{+}17, LWZ^{+}13, WWZ^{+}16, YWW^{+}15, ZQL^{+}16]. Constant [ACCP12, SHY14, VBC19]. Constant-Time [ACCP12, VBC19]. Constrained [CBF^{+}17, JRP^{+}10, LG13, MHL^{+}16, RBSS11, SWRQ18, TNZ^{+}12, WXZ^{+}14, WYY^{+}12, WIZ^{+}17, ZCJY14].

Constraint [BBL^{+}16, DOLG16, GJLZ13, JSC^{+}17, KN12, ZLN^{+}13]. Constraint-Based [ZLN^{+}13]. Constraints [BEDCR13, BB13, CC13b, Che18a, DWW^{+}11, FWJ18, GXW^{+}17, HCyW^{+}17, LRT19, NLGQ14, TYWL14, TCS11, TVRD17, XTFC17, ZMLT13, ZYL^{+}16]. Constructed [ZLL^{+}15]. Constructing [BS14, HJPL14, JWJS14, LY14, WJ12]. Construction [DWX14, DWY^{+}13, Lai12, LC10, PYH19, WKC12, YCPC15, ZASA10]. Constructive [WLH^{+}15]. Consumer [MBF19]. Consumption [CB16, CM10, CCD^{+}15, DSM14, GIRT19, LW15, LLpC15, NTKK15, aaGZ19]. Contact [CSY16, ZMF10]. Contained [ZS13]. Container [LCYW16]. Containerized [ALZ17]. Containing [MT15]. Contemporary [ZJS12]. Content [AKT^{+}15, BTG^{+}18, BFPB10, CL13, CE17, CSM^{+}13, CSY16, CL15, CE10, Dan11, HLWV14, JHMV12, JKS13, JWE15, KLWK12, LLLG13, LHL^{+}13a, LSCL16, NFFK14, QCZ^{+}15, RVCT15, WM15, YZL^{+}15, ZL11, ZY13, ZJL^{+}17a, ZCX10, ZCX15, ZWZ^{+}15]. Content-Based [BTG^{+}18, JWE15, QCZ^{+}15, WM15, ZJL^{+}17a]. Contention [ASG^{+}14, CCK12, CWCS15, EHNS13b, GGGA18, HLZY15, RPYO11, SHG13, SBMA15, ZWJ^{+}18]. Contention-Aware [HLZY15]. Contentions [LZH^{+}16]. Contents [CSZ^{+}12]. Context [PD14, RSSC15, SJ14, WDOX15]. Context-Aware [RSSC15, SJ14, WDOX15]. Contexts [BN12]. Contextual [SytL19]. Contiguous [ACS13, MLL14]. Continuous [BBR12, DWW^{+}15, JCLJ12, JCW^{+}12, LCL^{+}16a, MTDD17, SBC^{+}19, ZT14, ZTH17, ZT16]. Continuously [WWT^{+}19]. Continuum [LZB14]. Contour [GFJT19]. Contract [WMC^{+}19]. Contract-Based [WMC^{+}19]. Contrast [SZC^{+}17]. Contribution [NN10]. Control [Ara11, AA12, BKY15, CTX^{+}12, CSP13, CCL13, CFLL19, DWX14, DWW^{+}11, EHWX10, ESGG^{+}15, GLJ^{+}15, HDL^{+}15, HN11, JJ11, LZ12, LGJZ16, LXXH16, LZN10, LTM11, LLZ14, LWZ^{+}16a, LLD^{+}18, LGG^{+}14, MWJ^{+}14, MT15, MSB11, NTK^{+}15, PH11, RSN14, SS12, SX10, SCC11, SGC14, SB19, TCDMRP17, THL13, TKP12, WCF10, WMW11, WW11, WWCZ11, WCLK12, WCZ^{+}19a, WZLC15, XLM^{+}11a, XJY^{+}10, XLM^{+}12b, YJ14, YJR15, YLH^{+}16, YRL11, YXG12, ZJLS12, ZZF10, ZZR12, ZWWF15, ZWY^{+}17, ZJLG14, ZLZ^{+}16, ZHCW12]. Control-Intensive [LWZ^{+}16a]. Controllable [ZLDC15]. Controlled [LLD^{+}18, PNAK11]. Controller [BCTB13, HZT18]. Controllers [GKL^{+}17]. Controlling [THB^{+}14]. Controls [RAS17]. Convergecast [FQWL12]. Convergence [MGB18, SSM^{+}18]. Convergent [LLL^{+}14b]. Convex [GCZ15, LWJ^{+}15, TKP12]. Convolution [IG11]. Convolutional [CLB^{+}19, ZL14]. Cooling [ZTA^{+}15]. Cooper [LNK17]. Cooperate [Dan11]. Cooperation [JKS13]. Cooperative [AKC^{+}15, Cha14, CW15, CLL11, CSSL15, CMC^{+}15, DMR16, DSASSLP12, ERRG18, GCL14, GLJ^{+}15, HLS^{+}15, HN10, HGC12, HCH^{+}19, JZY^{+}15, KL11b, LHF^{+}15, LLZ^{+}12b, MPS15, MY11, NVS16, NTKK15, NTK^{+}15, WZ14, WL14, WL15, yWeH11, WS14, XZH14, YQ11, YSS^{+}17, ZGL^{+}15, ZZCD10, ZMTL15, ZY14, ZLDC15, ZHAY12, ZHCW12, Zhu14]. Cooperatively [TP14]. Coordinated [BRX13, HCY^{+}12, JKP12, MLC^{+}19, YSDQ11, ZJLG14]. Coordinating [LMFS11, LH15, WW11]. Coordination [ZCZ^{+}12]. Cope [SAH15]. Coprocessors [LLH^{+}15b]. Copy [DMS^{+}12, VMB17, WX15, XWH15a, XWH15b]. Copy-Back [WX15]. Coral [CSC16]. Corasick [TVCM12]. Core [AFA12, AAA19, AZW^{+}19, AA17, ASD^{+}18, AFMM17, CCKF15, CCC^{+}16, CRC^{+}17, CLL^{+}17, DMCN12, GZY^{+}15, HT16, IPQ19, JWY^{+}18, KAA16, KPKH16, LJ16, ME15a, MDM13, PCL15, PRS^{+}11, PJAGW14, PGGS19, QF14, RRM^{+}15, RGRM14, RAG10, SEAH16, ScFRdS15, SWRQ18, SAF16, SL14, SVK^{+}19, SKKK16, TCM18, WFZ^{+}17, YLJ^{+}17, YCMX17, YN17, ZJL^{+}17b, ZJS^{+}17, ZCXF16, ZWL17, AK18, KLL^{+}17, YSS^{+}17]. Cores [BHKS^{+}17, HGS^{+}19, HxjGG19, KCH19, MMNN16, PB19, Sib12]. CoreTSAR [ScFRdS15]. CoreVA [ASD^{+}18]. CoreVA-MPSoC [ASD^{+}18]. Correcting [KBHS14]. Correction [Ano11e, LMR10]. Correlated [HP14, HAY^{+}18, HKA12]. Correlation [CJ16, MFO^{+}13, SLSG19, TJH^{+}14, WWZ^{+}16, YLL^{+}17, YZJ^{+}12, ZXW^{+}13, ZHW^{+}19, ZFG^{+}10]. Correlation- [YLL^{+}17]. Correlation-Aware [CJ16, SLSG19, WWZ^{+}16, ZHW^{+}19]. Correlation-Based [ZFG^{+}10]. Correlations [CLL^{+}19, DGG^{+}19]. Correlations-Directed [CLL^{+}19]. Corroboration [OMMZ14]. Corruption [BBGD^{+}17, DC16]. COSE [HL12a]. Cost [APG12, ANE12, ARM16, BFFG11, CP17a, CJZ12, CZX^{+}19a, DLXS19, DWT^{+}16, DWW^{+}11, DWY^{+}13, DY18, ESGG^{+}15, FYH^{+}15, Fre13, GF13, HWJ18, HWL^{+}17a, HLL18, HGL^{+}16, KTK11, LSB^{+}18, LCLD13, LDYZ15, LCL^{+}16b, LLL18, LSN19, MHL^{+}16, NZM^{+}16, PSL15, Ren14, RGLDM17, SSW^{+}17, SYL^{+}14, SRG19, TUS13, VS19, WHF^{+}19, WIZ^{+}17, WHZS19, XBZL17, XDMZ17, XCZ^{+}15, XZJ^{+}19, YTZ^{+}11, YHS^{+}14, YZL^{+}15, YJC15, YJCQ15, YSS^{+}17, YYL^{+}13, ZS13, ZLN^{+}13, ZDM^{+}17, ZMW17, TLRW15]. Cost- [HLL18]. Cost-Aware [ARM16, HWL^{+}17a, XBZL17, TLRW15]. Cost-Driven [ANE12, WHF^{+}19]. Cost-Effective [ESGG^{+}15, KTK11, LSN19, MHL^{+}16, NZM^{+}16, PSL15, VS19, XZJ^{+}19, YTZ^{+}11, ZLN^{+}13, ZDM^{+}17].

Cost-Efficient [CZX^{+}19a, DLXS19, DY18, LSB^{+}18, WHZS19, XDMZ17, ZMW17]. Cost-Sensitive [XCZ^{+}15]. Costly [ARM16]. Costs [Dan11, SAA17, WUH^{+}17, YC18]. Could [Dan11]. Count [ZMA12]. Counter [WPKL13, WLX13]. Counter-Based [WPKL13]. Countermeasures [LJG12, YYY^{+}14, YZFZ10]. Counters [DSASSLP12, RX11]. Counting [BF17, FC10, IPQ19, SDL^{+}15]. Coupled [ASD^{+}18, HKkY^{+}16, MVML11, ZWL^{+}16b]. Coupling [BCQ^{+}10]. Coupling-Based [BCQ^{+}10]. COUPON [ZMTL15]. Covariance [XHG15]. Cover [Amm12, MM10]. Cover-Sense-Inform [Amm12]. Cover1 [Ano12d]. Cover2 [Ano12e]. Cover3 [Ano12f]. Cover4 [Ano12g]. Coverage [BBCB15, CMC^{+}15, DWLY15, GCN^{+}14, HCS12, HCY^{+}12, HCL^{+}12, HA10, JZH^{+}14, KZLL14, LVA^{+}11, LWZ^{+}15, LM12, LDNT13, LWZ12, MLT^{+}13, YPL^{+}17, ZYW^{+}14b]. Covered [Amm12]. Covering [ERSR13, GJLZ12]. Covert [ZSW^{+}15]. CPS [PKL^{+}12, Ano11c, Ano12h, LTW^{+}14, TWW^{+}15]. CPU [BBK17, CLO^{+}18, JHWY19, KLL^{+}17, LWC^{+}17, PD14, VNA^{+}16, WRB11, XCZ^{+}15, ZZH^{+}17]. CPU-Bound [XCZ^{+}15]. CPU-Intensive [JHWY19]. CPU/GPU [ZZH^{+}17]. CRAFT [STK^{+}19]. Crash [DGFRR18]. Crash-Prone [DGFRR18]. Cray [VTSM12]. Creation [LLGP13, SytL19]. CRED [XALS17]. Credibility [LTBN^{+}12]. CRESP [CPGT14]. Criteria [LT16, Tse13, aaGZ19]. Critical [ANE12, FWH^{+}18, GJZZ12, RLSK17, ZLJL17]. Criticality [BJC^{+}18, HTZY17, LGOB17, SAEH19, Wan19]. Cross [AKP14, BZA10, CLM^{+}15, DZLC15, ECW^{+}18, GIRT19, SF10, THL13, ZSW^{+}15, ZCXF16, ZCLS14]. Cross-Cloud [DZLC15, ECW^{+}18]. Cross-Core [ZCXF16]. Cross-Domain [GIRT19, SF10]. Cross-Layer [AKP14, BZA10, CLM^{+}15, THL13, ZCLS14]. Cross-VM [ZSW^{+}15]. Crossed [LMLM13, Wan12]. Crowd [CTLH14, LLZ^{+}12a, TNLM17, RSSC15]. CROWD-PAN-360 [RSSC15]. CrowdBC [LWY^{+}19]. Crowds [YZJ^{+}12]. Crowdsourcing [DLZH16, JAJ^{+}19, LWY^{+}19, OKT^{+}16, RSSC15, WGCG18, ZYZ^{+}14, ZYW^{+}14a]. Crowdsourcing-Based [ZYW^{+}14a]. CRS [LWCL18, WXLY16]. Cruising [ZHL^{+}15]. Cryptographic [SP15]. Cryptography [ARM15]. Cryptosystem [CCT^{+}14]. CSI [Amm12, WXY^{+}13]. CSI-Based [WXY^{+}13]. CSMA [RMM16]. CSMA/CA [RMM16]. Cube [LL12, LMLM13, LY16a]. Cube-Based [LY16a]. Cubes [Wan12, XS11, YLM^{+}15]. Cuckoo [SHF^{+}17]. CUDA [EADT19, LAD16, NSLV16, WJB14, vdLJR11]. CUDAlign [dOSMM^{+}16]. CUP [ERRG18]. Curious [XAG17]. Curve [ARM15]. Customer [AHSH^{+}16, WWL^{+}15]. Customer-Provided [WWL^{+}15]. Customers [GPF12]. Customizable [KGR16, LSL^{+}18]. Cut [BCSKN12, ZGY15]. Cut-Through [ZGY15]. CUTBUF [ZFF16]. CUTS [NZWL14]. Cutting [HMS^{+}18, QPB^{+}17]. Cyber [Ano11c, CTX^{+}12, HGY^{+}14, HWNS15, LQY^{+}12, LCSC12, MV12, RXD12, YQZC12, ZYL^{+}17, PKL^{+}12]. Cyber-Physical [Ano11c, CTX^{+}12, HGY^{+}14, LQY^{+}12, LCSC12, MV12, RXD12, YQZC12, ZYL^{+}17, PKL^{+}12]. Cycle [CH15, IMH12, XWH15b]. Cycled [GCN^{+}14, HCS12, JLM^{+}12]. Cycles [LKM10, LHJ12]. Cyclic [GS11b]. Cycling [Li14b].

D [CCLW15, GRUMG17, GAB18, AAB16, AVA^{+}17, BKF^{+}16, CLHW13, Che18b, DWH^{+}18, Has16, HWZE10, KGI17, LWSM19, SWT^{+}19, WJTZ14, WCYL19, YTL^{+}19, ZM13, ZYX^{+}10]. D2P [MBO15]. DAG [CJ10, CJ16, HGS^{+}19, HxjGG19, MWZ^{+}14]. DAGs [CDR15, SFL^{+}14]. Dark [LODB17, WFZ^{+}17, YLJ^{+}17]. Data [AHSK17, ASG^{+}14, ACNP11, AFT^{+}16, AB14, AA14, AEM17, ASD^{+}18, BM12, BG13, BH13, BB13, BBGD^{+}17, BSM^{+}11, Bru14, BAAT16, BSL^{+}17, BZBP10, CGS^{+}15, CJH^{+}14, CWL^{+}14a, CDBQ12, CZZ^{+}16, CHTW12, CLLS12, sCCyW14, CL14, CYX15, CZT^{+}17, CHW^{+}17, CLT^{+}17, CZQ^{+}17, CLO^{+}18, CHHK19, CLL^{+}19, CMK^{+}16, CDPM18, CIH13, CCT^{+}14, CSR^{+}17, CAKRY16, CWC^{+}13, CTP^{+}17, DGC17, DLZH16, DN19, DRRCB18, DWW^{+}15, DC16, DC18, DTLC19, DED^{+}19, DCA^{+}16, DZLC15, DY16, DY17, EHWX10, ELX^{+}11, FC10, FCD^{+}13, FGJ^{+}15, FYH^{+}15, FGEL14, FRS^{+}16, GFLL15, GXW^{+}17, GKL^{+}17, GSH^{+}19, GETFL14, GYX^{+}10, GG11, GZY^{+}15, GTT^{+}17, GZW^{+}18, GJPPM^{+}12, GF13, GGF^{+}14, GHL14, GXZ^{+}15, Guo17, HOZ12, HJY16, HJPL14, HCG^{+}15, HWS16a, HWL^{+}17a, HLCB^{+}17, HBF12, HLL18, HC14, HWQ^{+}15, HLK^{+}19, HN11, Hur13, IPQ19, IBC^{+}11, IdM12, JRZ^{+}18, JSMK11, JDB^{+}14, JGG^{+}11, JCLJ12, JZW^{+}17, JJW11, JEW^{+}18, JRO^{+}17, Jun17, KCW11]. Data [KAV^{+}17, KXL^{+}14, KPG^{+}12, LMM18, LM17, LGD14, LKE16, LMD16, LXHL11, LAMJ12, LYGX12, LLL^{+}13, LCS14, LWZ14, LWY^{+}15, LLG15a, LYH^{+}15, LY16a, LRYJ17, LGM^{+}17, LLY^{+}17, LWLZ17, LLZ^{+}18a, LT12, LS17c, LZWY13, LZK^{+}15, LLH^{+}15a, LHYW15, LSC16, LNK17, LN17, LS19, LSN19, LLZ^{+}12b, LCA13, LLG14, LTMD11, LYZ^{+}16, LGW^{+}17, MLL14, MZLT19, MMdE19, MPM17, MNG^{+}15b, MTDD17, MDZC14, MP16, MV12, MV16a, MBV11, MBV13, MBH^{+}10, NCGP19, NNKL13, NCM^{+}17, NTWL11, PYHY16, PR19, PD14, PRR^{+}16, PG16, QFZZ15, QGPZ13, Rao14, RGK15, RZH^{+}11, RZW^{+}13, Ren14, RGLDM17, RSN14, SML13, SMS^{+}13, SMTZ17, SMB^{+}18, SJR17, SSF16b, SLSG18, SSW^{+}17, SP15, SYZ18, SS17, SF10, TKS11, TG13, TZC19, Tic14, THB^{+}14, TFLL18, TP13, TPRH16, UDH^{+}17, VMB17, WWR^{+}11, WWL11, WMHX12, WCRL12, WJTL12, WCLK12, WVT13, WWX^{+}13, WW13, WZ14, WMZ^{+}15, WWZ^{+}16, WHF^{+}19, WWT^{+}19, WCZ^{+}19b, WZS^{+}19]. Data [WHN^{+}19, WTXG19, WMWW19, WK11, uRILP17, WDH^{+}16, WLL^{+}19, WLLL10, WCF13, WSSZ13, XXLZ16, XWSW16, XXL^{+}19, XSZ^{+}10, XBZL17, XS10, XWJX15, XDMZ17, XLM^{+}11b, XSZ13, XLSR13, XHQ^{+}15, XFL15, XALS17, XL17, XZJ^{+}19, XWL^{+}19, XZQZ17, Yan14, YNW13, YJ13, YJ14, YYY^{+}14, YXWW14, YJR15, YLZ^{+}15a, YLC^{+}16, YGS^{+}19, YHS^{+}14, YGL^{+}15, YWW^{+}15, YYK11a, YYK^{+}11b, YRL11, YXG12, YQLS14, YJC15, YJCQ15, YQ16, YYL^{+}13, YWZ17, ZJL^{+}12, ZGH14, ZYZC12, ZLN^{+}13, ZLZ^{+}14, ZCJY14, ZYLC14, ZRTL15, ZWL^{+}15, ZWL^{+}16a, ZYL^{+}17, ZQWL17, ZDM^{+}17, ZZSZ18, ZRS18, ZSSR19, ZDM^{+}19, ZHL19, ZT13, ZLK^{+}16, ZZQ18, ZRQA14, ZYT^{+}15, ZMW17, ZWY^{+}17, ZLT^{+}18, ZXG^{+}19, ZHAY12, ZJ16, ZLW^{+}18, ZWT^{+}19, ZGKB16, WZS^{+}18]. Data-Centric [ASG^{+}14, GHL14, PG16, SMS^{+}13]. Data-Injection [YYY^{+}14]. Data-Intensive [HC14, KCW11, LS17c, MZLT19, MBH^{+}10, WCZ^{+}19b, ZLK^{+}16]. Data-Parallel [FGJ^{+}15, GTT^{+}17, JSMK11, RGLDM17, ZWT^{+}19]. Data-Race-Free [JEW^{+}18]. Data-Resolution [WMWW19]. Database [DRSL15, XCZ^{+}15, ZTZQ19]. Databases [FCM14]. Datacenter [AOW^{+}12, CFLL19, EKNS17, LHG^{+}17, LGL^{+}18a, LHXP18, LSL^{+}18, YMHL16]. Datacenters [CMB18, CZD^{+}19, LGJZ16, LGL^{+}18b, LGJ^{+}18, LSC16, LSL^{+}17, LCZ^{+}19, XBZL17, YPL^{+}17]. Dataflow [EJGYAM14, FA19, PBD^{+}13, WZL^{+}16, WM18]. Dataflow/von [EJGYAM14]. Dataflow/von-Neumann [EJGYAM14]. Datasets [KJN15, VPS17]. Datastores [MA14]. Datatype [KB17]. Datatypes [JDB^{+}14]. Day [MV18]. Day-Ahead [MV18]. DC [XLL^{+}18]. DCloud [LCG^{+}16]. DCN [ZDM^{+}17]. DCNS [GFMR13]. DCS [CLSZ12]. DDC [KWZ^{+}12]. DDFCharts [RSR11]. DDoS [Wu14, YZDJ11, YZJ^{+}12]. Deadline [ABN19, CZX^{+}19b, GXW^{+}17, LCG^{+}16, LSW16, RGPH15, WIZ^{+}17]. Deadline-Aware [CZX^{+}19b, LCG^{+}16]. Deadline-Constrained [WIZ^{+}17]. Deadlines [CB14, LMAS17, PP12, XALS17]. Deadlock [ADMX^{+}12, Dua95, GAB18, LX12, MMYES^{+}18, RGBC11, SHG11, VS11a, VS11b, VS14, XL16, XL10, Dua93]. Deadlock-Free [Dua95, GAB18, LX12, MMYES^{+}18, RGBC11, SHG11, VS11a, VS11b, VS14, XL16, Dua93]. Deadlocks [CJW^{+}15]. Deal [QGPZ13]. Dealing [ACNP11, FPGAD10]. Debugging [DAJ14, LZH^{+}16]. Decentralized [Che15, GZZ^{+}13, HSMY12, LWY^{+}19, LT10, LDYZ15, RSN14, She10a, TLL^{+}16, XZT^{+}13, YLT15]. Decision [KM19, LLN^{+}19, LJ15, VS14, WSH^{+}19]. Decision-Making [KM19, LJ15]. Decisions [CAKRY16]. Declarative [EADT19, ZHCL17]. Declustering [TOA13]. Decode [KWZ^{+}12]. Decoder [TBC12]. Decoders [LJ16, NGJ^{+}19, ZL14]. Decoding [BSD^{+}18, FSS11]. Decomposing [LVD11]. Decomposition [ATA18, HWC15, JP12, LWJ^{+}15, MDM13, SSM^{+}18, SS18, Van14, VMP17, XTFC17, YRLY16]. Decompositions [JHR15]. Decoupled [CSW^{+}17]. Decoupling [GIRT19]. Decrease [Dan11]. Deduplicating [ST18]. Deduplication [HL12b, Li14a, LLC^{+}15, LCYW16, LLL^{+}14b, TWW^{+}18, WHZS19, WMJ^{+}19, XLT^{+}14].

Deduplication-Based

[TWW^{+}18, WMJ^{+}19]. Dedupped [YZHZ17]. Deep [CCHH19, CSR^{+}17, LHHR18, WZL^{+}19, YP13, YZH^{+}19]. Deeper [GGGA18]. Deeply [TLP12]. Defending [CDS15, QLC13]. Defense [SILJ11, WXTL13]. Deferred [KKW18, WKK17]. Deferred-Update [KKW18, WKK17]. Deficit [MMACS10]. Defined [BTG^{+}18, HGL^{+}16, SB19]. Deformable [HKE^{+}16]. Defragmentation [LWSM19, TWW^{+}18]. Degradable [JWJS14]. Degraded [SLSG18]. Degree [BEDCR13, KMM13b, LMSRSR13, LY14, TFKN17]. Degree-Dependent [LY14]. Delaunay [SZ12]. Delay [ANN^{+}13, CL17, CSY16, Che18b, CCCB14, CLSZ12, DOLG16, EHNS13a, FYH^{+}15, FWJ18, FQWL12, GJLZ13, HL12b, JZY^{+}15, LAV^{+}10, LCZZ13, LW12, NTK^{+}15, PKCB11, PLZW14, PNAK11, RBSS11, RS12, RKRK17, TFLL18, WBPF11, WYW13, XLM^{+}11b, XGZW14, YHS^{+}14, YXG12, YJCQ15, ZGH14, ZYZC12, ZMLT13, ZDG^{+}14]. Delay-Aware [HL12b]. Delay-Bounded [LAV^{+}10]. Delay-Capacity [WBPF11]. Delay-Controlled [PNAK11]. Delay-Efficient [XLM^{+}11b]. Delay-Sensitive [TFLL18]. Delay-Tolerant [NTK^{+}15, XGZW14, ZDG^{+}14]. Delayed [LCYW16]. Delays [VS15]. Delegation [FGLP10, KSW18, NLC12, XWLJ16, XAG17, XWS17]. Delegation-Based [NLC12]. Deletion [QZW14]. Deletions [Tse13]. Deliberate [WLH^{+}15]. Delivery [AKT^{+}15, CE10, LWZ14, NFFK14, TCS13, WLH^{+}15]. Delta [ZGGW14]. Delta-Based [ZGGW14]. Demand [CE17, CZWZ14, KCK14, LLY^{+}14, LTC16, LSB^{+}18, LFLW10, NSH15, YQH16, ZLZ^{+}14]. Demand-Side [YQH16]. Demands [LZY^{+}18, LLL18]. Denial [TJH^{+}14, XSTZ10]. Denial-of-Service [TJH^{+}14, XSTZ10]. Dense [AHTD18, FGEL14, PSMD18, Tou15b]. Density [WCF10]. Departures [LW14]. Dependability [dCCF15, ZJLS12]. Dependable [FLLS17, HWG^{+}19, VMN^{+}16]. Dependence [KAC^{+}15, LAdS^{+}15, ZR18]. Dependencies [SML13, ZGKB16]. Dependency [WCZ^{+}19b, YZT^{+}17, ZWT^{+}19]. Dependency-Aware [WCZ^{+}19b]. Dependent [AOW^{+}12, Fre13, LY14]. Deployment [CCS^{+}12, DED^{+}19, DLC^{+}16, MVML11, SAM14b, SHX^{+}10, WWL11, WSWY15, ZYW^{+}16]. DepSpawn [FA19]. Depth [HH13, Hen14]. Depth-Optimal [HH13]. Deregulated [Ren14, ZCJY14]. Derivation [SV19]. Derivative [SLG^{+}18]. Derived [JDB^{+}14]. Design [AHTD18, AVA^{+}17, AKP14, ASYK^{+}19, CAC^{+}19, CLM^{+}15, CLLX18, CCS^{+}12, DLXS19, DA16, EAMEG11, Fen14, GG10, HXLF15, HSX^{+}12, HA13, IBC^{+}11, JZZ^{+}15, KGI17, KM18, KE16, KI14, LRW12, LL11, kL11a, LLC10, LLZ^{+}12a, LLH^{+}15a, Lu14, LWZ^{+}16c, MVC^{+}18, NHN17, NHN18, Pan14, PSL^{+}11, RSR11, RH16, RVCT15, RLY^{+}15, SDV18, SHX^{+}10, SZ11, TWW^{+}15, TLRW15, THL13, WWL^{+}13, WL15, WKL^{+}16, WVM19, WZGR10, WCF13, WML14, XXWY10, Yan14, YDC^{+}17, YWWR18, ZD12, ZYZ^{+}14, ZGL^{+}15, ZBS15, ZLL^{+}15, ZD16a, ZZCD10, ZW14, ZWY^{+}17, ZFF16]. Designing [LSLD17, LWLZ17, LAD16, YKN^{+}19]. Designs [CP17b, HYX11, LHL^{+}13a, QFZZ15, QGPZ13]. Desired [LTMD11]. Desktop [ICN18]. Destination [TCS13]. Destination-Oriented [TCS13]. Detailed [MMBdS14]. Details [Ano12h]. Detected [JMA^{+}18]. Detecting [CQZ^{+}12, LPZ12, MLML15, WWCB14, XSTZ10, YLZ^{+}15a, YL16, ZRQA14]. Detection [ALLR14, ADMX^{+}12, ABLS16, BCSKN12, BBGD^{+}17, CWS12, CC15, CPL^{+}18, DC16, DO13, DLC^{+}16, EK10, GDRTS16, GLM13, HST^{+}11, HYC^{+}12, HH12, JEW^{+}18, KKK11, LLZ^{+}18a, LSW^{+}15, LWG^{+}12, MGB18, NFFK14, PLZW14, SAM14b, SK14, SM16, TXWL11, TJH^{+}14, Tic14, TP18, WFA13, WWX^{+}13, XL10, XHHC13, XHG15, XWY^{+}10, XGZW14, YCTC13, YHC^{+}13, ZYW^{+}14b, ZDG^{+}14]. Detector [SRB14, YTZ^{+}11]. Detectors [JRAS17]. Determination [sFC12, LAFA15]. Deterministic [FSM^{+}12, HA10, LW14, MMYES^{+}18]. Developer [DWT^{+}16]. Developing [CLZ^{+}18, HZJ16]. Development [HAD12, WZGR10]. Device [KN12, LTW^{+}14, ZYW^{+}14b]. Device-Free [ZYW^{+}14b]. Devices [KHK15, LLG^{+}13, ZLL^{+}17b]. Devolved [GKL^{+}17]. DGLB [CMG17]. DHT [RVCT15, SX10, SLL13a]. DHT-Aided [SLL13a]. DHTs [AAAK^{+}14, YL11a, TXZ^{+}11]. Diagnosability [CH14, LKT11, LZXW15, LXZH16, YLM^{+}15]. Diagnosing [DD17, TKC^{+}15]. Diagnosis [Cha11, DDW^{+}19, DLC^{+}16, DWF12, EN12, GLL15, LAdS^{+}15, LKT11, MWZ^{+}13, PWT^{+}17, TAZ^{+}19, YL15, ZD16b]. Diameter [Sib12, TFKN17]. Diameters [TCT14]. Diamond [BBP17, YGS^{+}19]. Dictionary [WHB16]. Difference [LC10, PBD^{+}13]. Different [LZ11]. Differential [BTL^{+}19, WHN^{+}19, ZLZ^{+}17]. Differentiated [LV15]. Differentiation [XZSG12]. Diffusive [MM15]. Digit [LAD16]. Digital [MT12, WMZ^{+}15]. Digraphs [GWL^{+}11]. Dimensional [BSF16, CFJ15, KJN15, NGJ^{+}19, SMB^{+}18, Sib12, ZD16a]. Dimensional-Permutation-Based [CFJ15]. Diophantine [ZTD19]. Direct [RAG10, WJB14]. Directed [ADLM19, CLL^{+}19, ZLS^{+}18]. Direction [FXL17]. Directional [CWJS11, GLL15, JWA10, KCK14, YW10]. Directly [KWZ^{+}12]. Directory [ACV17, TGFFP^{+}19]. Disappearing [AJMW14]. Disaster [LODB17, WMWW19]. Disasters [XLL^{+}18, ZHS^{+}19]. Disciplinary [YZFZ10]. Disciplines [Sto10f]. Disco [WLH^{+}15]. Disconnection [SAH15, YL11b]. Discoverability [RXD12]. Discovering [JKVA11]. Discovery [CC10, HCG^{+}15, LLY^{+}15, OKT^{+}16, RVW^{+}15, SGGB14, WML15, WRB11, YZT^{+}17, ZWT^{+}19]. Discrete [PF12, PJAGW14, QJ16, TRT19, WHN^{+}19, XAK17]. Discriminating [YZJ^{+}12]. Disjoint [Lai12, XBL15]. Disk [AZW^{+}19, AT12, CLKR15, FSSZ16, LLJ^{+}13, LIWJ15, LWZ^{+}16c, WHH^{+}13, WWL^{+}17, XTFC17, XS10, ZLS^{+}18]. Disk-Based [ZLS^{+}18]. Disks [HYZ15]. Dispatch [WPT10]. Disruption [LHF^{+}15, YCW12, ZCLS14]. Disruption-Tolerant [YCW12]. Disruptive [GBFS16]. Dissecting [MC17]. Dissemination [CL15, DLZ^{+}14, FCD^{+}13, GBD^{+}13, HCG^{+}15, LXHL11, LSKZ13, LNK17, RVCT15, TYG^{+}14, TZB^{+}14, Ven14, ZGH14, ZCJ19, ZWZ^{+}15]. Distance [ABLS16, Fre13, GC16, GV15, KGI17, Li13, LJB^{+}13, SV19, HZB^{+}16]. Distance-Based [ABLS16, Li13]. Distances [LAFA15]. Distinct [WWT^{+}19]. Distortion [LCW11, TDL^{+}19]. DistR [CYC^{+}16]. Distributed [ALLR14, AB14, Ano11d, Ano11c, Ano12c, Ano15a, Ano16, Ano17a, Ano18, Ano19a, AGJ^{+}16, ADLM19, AF19, BKY15, BGHG16, BG13, BCQ^{+}10, BBR12, BAMJ12, BCTB13, BVEAGVA10, BVFGSFAF17, BBK17, BFPB10, BBM16, CJH^{+}14, CLL^{+}14, CYZ^{+}13, CADK19, CMT^{+}17, CLL11, CCL13, sCCyW14, Che14, CCT16, Che16, CMG17, CLL^{+}19, CYC^{+}16, CZD^{+}19, CLSZ12, dCCF15, DBAT11, DD11, DSM19, DBA17, DHBB12, DGF12, DRRCB18, DRSL15, DCA19, DWF12, ET10, ESGQ^{+}13, FYH^{+}15, FWH^{+}18, FCM14, FHH^{+}15, GG10, GLZ11, GGY^{+}19, GGS10, GKKW16, GD16, GCG^{+}18, GG11, GHZZ16, GTT^{+}17, GLJ^{+}15, GCZ15, GYW^{+}19, HGY^{+}14, HOZ12, HP14, HCG^{+}15, HGC12, HCSC13, HKH^{+}10, HLL18, HXC^{+}11, HPH^{+}12, HCL^{+}14, IdM12, JHMV12, JKS13, JKVA11, JSLD19, JGJF18, JZW^{+}14, JHW^{+}15, Jia16, JMS^{+}18, JCWB10, JRO^{+}17, KN12, KM19, KKC17, KTK11, KDCR19, Ksh10]. Distributed [Kum14, LHHR18, LZ11, LKE16, LC15, LJ15, LL17, LODB17, LGM^{+}17, LLS^{+}18, LL11, LT10, LHM12, LNZ^{+}13, LCS^{+}15, LSC16, LS17d, LH17, LM16, Lu14, LLD^{+}18, MLXG19, MGB18, MB13, MP16, MPS15, MDM13, MLVD12, NCGP19, NNKL13, NSH15, NCKL14, NN13, PHGR17, PKS14, Par19b, PHXL19, PDH10, PH12, PWT^{+}17, PSMD18, RSR11, RAS17, RGPH15, SZC^{+}17, SS12, SMTZ17, SL13, SLGW14, SWL17, SPS18, SLM^{+}10, STMM17, SF10, TNZ^{+}12, TWT16, TZ10, TWL16, TCZL11, TFKN17, Tsa13, TKP12, TVCM12, TS16, VWDM14, WZQY14, WWT^{+}19, WUM10, WZGR10, WSSZ13, WML14, WYCZ14, WZLC15, WXLY16, XP12, XXLZ16, XBZL17, XZX^{+}17, XLL^{+}18, XJY^{+}10, XFL15, YHL^{+}18, YNW13, YLH^{+}16, YHS^{+}14, YZS13, YC14, YYK^{+}11b, YDC^{+}17, YZH^{+}19, YRL11, YJC15, YLC^{+}19, YWC11, YC12, ZG11, ZJL^{+}12, ZLZ^{+}17, ZGL10, ZZR12, ZGGW13, ZT14, ZSY14, ZGL^{+}15, ZTH17, ZLL17a, ZSSR19, ZWJ^{+}19, ZJZ^{+}16]. Distributed [ZT16, ZHQ12, ZLDC15, ZLZ^{+}16, ZHCL17, ZXG^{+}19, ZWT^{+}19, Zou14, MBO15]. Distributed-Healthcare [ZLDC15]. Distributed-Memory [TVCM12]. Distribution [BKH18, BMB^{+}10, CJ16, CTLH14, Dan11, HLWV14, KLWK12, LLLG13, LAMJ12, LHL^{+}13a, LLC10, LSL^{+}18, LA12, PG16, PNAK11, SCCC11, THB^{+}14, WFA13, WHN^{+}19, WYC^{+}15, XHL^{+}11, XZH14, ZL11, ZY13, ZCX10, ZCX15, ZJTZ14, dSLMM11]. Divergence [AB14, Nov15]. Diverse [CSY15, THT^{+}15]. Diversity [CCH^{+}17, MWJ16, MY11, ZWE19]. Diversity-Based [MY11]. Divide [SZWX15, SYZ18, YPL13]. Divide-and-Conquer [SZWX15]. Divide-and-Merge-Based [YPL13]. Dividing [KKK11]. Divisible [DW10, HV11, JVW10]. DMA [MLC^{+}19]. Does [LHL^{+}13b]. Domain [ADZZM15, GJLZ12, GIRT19, ITL17, kL11a, LLLH19, MRH^{+}16, NZWL14, SS18, SF10, XXWY10, ZX13]. Domain-Decomposition [SS18]. Domain-Specific [MRH^{+}16]. Domains [NVBH18]. Dominating [CHD^{+}15, MM10, YC14]. Double [ARM15, CZWZ14, GYX^{+}10, LYZL18, LWZ12, TTJX12]. Double-Edged [GYX^{+}10, TTJX12]. Down [PT11, SKP12, WQZ^{+}15, ZYLC14]. Down* [RGBC11]. Downgrade [RLSK17]. DP [XZQZ17, ZZQ18]. DPillar [EKNS17]. DRAGON [HH12]. Dragonfly [MMYES^{+}18, XL16]. DRAM [KHK15, MLC^{+}19, MVL15, MV16c]. DREAM [ZJZ^{+}16]. DREAM- [ZJZ^{+}16]. Driven [ANE12, AF18, CCJ19, CSW^{+}17, CWCS15, DWT^{+}16, DC16, EHM^{+}17, GIX^{+}12, HZW^{+}19, LLY16, LH17, SLL13a, SHM^{+}12, TZB^{+}14, WHF^{+}19, ZWZ^{+}15]. Drivers [LQY^{+}12]. Drives [YTW^{+}19]. Droppers [WFK^{+}12]. DSP [SY17]. DSystemJ [MGS12]. DTN [CSY15]. Dual [ATACA18, JCLJ12, RJ16]. Dual-Consistency [RJ16]. Dual-Plane [ATACA18]. Dual-Radio [JCLJ12]. Duplex [Zhu14]. Duplication [CZQ^{+}17, HMP^{+}19, TWSW17]. Duplication-Based [TWSW17].

Durability [LSN19]. Durable [LZW^{+}17]. Duration [XHX^{+}13]. during [SAH15, ZWL^{+}15]. Duty [GCN^{+}14, HCS12, JLM^{+}12, Li14b, XWH15b]. Duty-Cycled [HCS12, JLM^{+}12]. Duty-Cycling [Li14b]. DVFS [BSD^{+}18, CZL^{+}18, GIRT19]. DWT [EALM15]. Dynamic [AKC^{+}15, AFT^{+}16, AGJ^{+}16, BCQ^{+}10, BH13, BB13, BS15, BB17, CJW^{+}15, sCCyW14, CYC^{+}15, CCLW15, CJZ^{+}16, CZWJ18, CZX^{+}19b, CCCB14, CCK12, CWC^{+}13, DM11, DK17, DWW^{+}15, DW13a, DWF12, DRK11, EHWX10, FPF13, GKT^{+}17, GBFS16, GYLW18, GZWN14, HLWV14, HH12, JRAS17, JCWB10, JWY^{+}18, KM10, KKE19, KKK^{+}15, LC12a, LMSRSR12, LTC16, LDNT13, LZWY13, LJL^{+}15, LCA13, MWZ^{+}14, MG14, ME15a, MBO15, MGR12, NIP11, NMG15, NTK^{+}15, NL11, PPR10, PHXL19, QZZ^{+}16, Rao14, RHDL11, RZW^{+}13, RCC^{+}14, RGBC11, RJ16, SJR17, SWL17, SGC14, SPH^{+}18, SVC12, TWT16, TYS^{+}12, TZC19, TJLL12, Van14, WZQY14, WNLL15, WUH^{+}17, WWW^{+}18, WGCG18, WK11, yWeH11, WS14, Xia14, XWSW16, XSC13, XBZ^{+}16, XS10, XML^{+}18, YJ13, YHC^{+}13, YTW^{+}19, YZS13, YOK^{+}17, ZFG^{+}14, ZX13, ZT13, ZH14a, ZJ16, ZL10, WZS^{+}18]. Dynamically [AJMW14, HZG^{+}17, LX10, QP16c]. Dynamics [KAG17, RXD12, WWR^{+}11, WZZ^{+}13].

E-Commerce [WMGA15]. e-Science [ABN19]. E-SmallTalker [CYZ^{+}13]. EAFR [LS17c]. Eager [TGNA^{+}13, TGAG13]. EAP [FLH13]. Early [FWH^{+}18]. Earth [HZB^{+}16, WMZ^{+}15, ZWQ^{+}15]. Earth-Observation [ZWQ^{+}15]. Easier [STK^{+}19]. Easy [FA19, HCA16, GLRT18]. EASY-Backfilling [GLRT18]. EasyPDP [TYS^{+}12]. Eavesdropping [CWL16]. EB [XAYM14]. EB-Scale [XAYM14]. EBRP [RZH^{+}11]. EC2 [MHL^{+}16, TYWL14]. Economic [Sam14a]. Economical [LSW17b, YMHL16]. Economically [LHG^{+}17]. Economies [CB13, WZSL12]. Ecosystem [ZDWR11]. EcoUp [YMHL16]. EDCA [MRM12]. EDF [ATZZ14, CLL^{+}17, RGPH15]. Edge [CE17, CLH13, CH15, DLL^{+}11, FWZ^{+}16, HCH^{+}19, HZW^{+}19, JRO^{+}17, LGOB17, LGZ^{+}19, TCT16, WWW^{+}18, YZL^{+}17, ZZSZ18]. Edge-Bipancyclicity [CH15]. Edge-Fault [CLH13]. Edge-Pancyclicity [CH15]. Edged [GYX^{+}10, TTJX12]. Edges [CH15, XS11, XWL^{+}19]. Editing [WUM10]. Editor [Sto11c, Ano11e, BKK11, Bad15, Bad16, Bad17a, Bad17b, Par19b, Par19c, Par19a, Sto10f, Sto10a, Sto10b, Sto10c, Sto10d, Sto10e, Sto11b, Sto12a, Sto12b, Sto13c, Sto13a, Sto13b]. Editorial [Sto11a]. Editors [CLL^{+}14, MBMC13, PKL^{+}12, RFZ11]. Effective [BFD19, CJL^{+}12, ESGQ^{+}13, ESGG^{+}15, JWE15, JLKG17, KTK11, LW11, LQY^{+}12, LWC^{+}17, LSN19, LCA13, MHL^{+}16, NZM^{+}16, PSL15, PNAK11, SP12, VS19, XZJ^{+}19, YTZ^{+}11, ZLN^{+}13, ZDM^{+}17]. Effects [PB12]. Efficiency [CZL^{+}18, DGC17, EK10, FBCB18, FRS^{+}16, HD15, KPB19, MLC^{+}19, MJK14, Ozd19, PCL15, PPS^{+}17, SKKK16, WKK11, WMWW19, XLM^{+}11a, ZTA^{+}15, ZQSY13, ZLT^{+}18, dLMPG19]. Efficient [APMG12, AHTD18, AFA12, ABF12, ACV17, AFMM17, BN12, BWH^{+}19, BSD^{+}18, BBK17, BGE^{+}16, Bis18, BXXC12, BS12, BB15, BB16, CGS^{+}15, CSV^{+}17, CDBQ12, CJL^{+}12, CSY16, CZS^{+}16, CP17b, CBF^{+}17, CFLL18, CZX^{+}19a, CIH13, CLA^{+}19, CMG^{+}14, CRD11, CHPY17, DWX14, DLXS19, DM11, DSM19, DWH^{+}18, DTLC19, DWW^{+}11, DBG^{+}14, DSASSLP12, DL17, DY18, ECW^{+}18, EHI11, ESGG^{+}15, FC10, FLH13, FVLD16, FHW11, Fen14, GBD^{+}13, GGY^{+}19, GSH^{+}19, GGS10, GC16, GG11, GJLZ13, GZW^{+}18, GDM^{+}13, GYQW15, GXZ^{+}15, GS17, HH13, HAU19, HML^{+}14, HJY16, HCY^{+}12, HA10, HGC12, HLL18, HLeS^{+}15, HLQ^{+}15b, HZB^{+}16, HN11, JHR^{+}14, JJW11, JCW^{+}12, JGZZ14, JHW^{+}15, JHWY19, KLWK12, KALK^{+}18, KKW13, KXC11, KKK11, KPG^{+}12, Ksh10, LZ12, LGOB17, Lee12, LPP13, LAV^{+}10, LC10, LdSS^{+}13, LLY^{+}14, LTL14, LHL17, LSB^{+}18, LTC^{+}19, LHR^{+}15, LHZJ19, LKT11, LS17c, LWW^{+}13, LZP^{+}13, LS14]. Efficient [LLM^{+}14, LHYW15, LXZB15, LWZ^{+}16a, LAD16, LLL^{+}14b, LVD11, LLL^{+}12, LLG14, LX12, LGXL19, MTX^{+}11, MA14, MZK19, MG18, MBF19, MVC^{+}18, MRGR12, NCGP19, NLGQ14, PLG19, PF12, PAB13, PWJ16, QCZ^{+}15, QP16a, Rao14, RLSK17, SDV18, SEA18, SHA19, SY17, SMTZ17, She10a, SLL13a, SLGW14, SSLF17, SLSG18, SLSG19, SBMA15, SS17, SYXL16, SCH11, TKS11, TYS^{+}12, TWL^{+}15, TZY^{+}18, TFM^{+}16, TMMN15, TGAG13, Ven14, VBC19, WHH^{+}13, WLS^{+}11, WCRL12, WQZ^{+}16, WHGS17, WVM19, WZS^{+}19, WK11, WSH^{+}19, WLLL10, WKC12, WSSZ13, WHC^{+}14, WXLY16, WWH^{+}17, WKW19, WHZS19, XAY^{+}14, Xia14, XXL^{+}19, XJ14, XHL^{+}15, XZX^{+}17, XDMZ17, XJY^{+}10, XLM^{+}11b, XLM^{+}12b, XLM12a, XL13, XQL^{+}14, XAYM14, XLX^{+}16, XWL^{+}19, YW10, YJ13, YXSS13, YJ14, YLZ^{+}15a, YPL^{+}17, YCMX17, YTW^{+}19, YLW13, YQLS14, YCW12, YLT15, ZWD^{+}10, ZS10, ZPD11, ZY13, ZJKQ16, ZQWL17, ZDM^{+}17, ZLS^{+}18, ZWJ^{+}19, ZCJ19, ZQH13, ZMW17, ZLLD18]. Efficient [ZTZQ19, ZHCW12, ZDG^{+}14, ZWT^{+}19, ZGKB16, ZR18, ZHS^{+}19, LLZ^{+}12b]. Efficiently [CJW^{+}19, PHXL19, SDG17, ZSH^{+}11].

Effort [MPHR17, QGZP17]. Egalitarian [PR19]. EIC [Sto13c]. Eigensolver [AAW^{+}17]. Eikonal [HJ17, SS18]. Eisenstein [FB10]. Elastic [CLLX18, sCCyW14, GJPPM^{+}12, HBS^{+}16, LZY^{+}18, LABQ18, NZM^{+}16, NCB17, SX10, THB^{+}14, WM15, WLL^{+}19, YJC^{+}16, ZXL^{+}17, ZWG^{+}16, YJC^{+}16]. Elastic-RAID [YJC^{+}16]. Elasticity [MMdE19]. Election [SOK^{+}19]. Elections [dCCF15]. Electric [QLC13, WPT17, YLH^{+}16]. Electrical [JMZD12]. Electricity [CJZ12, GF13, LYY16, MV18, Ren14, ZCJY14]. Electronic [SF10]. Elements [LLH14]. ELIAS [KXC11]. Eliminating [WWH13]. Elimination [SCHT16, YSS^{+}17]. Elliptic [ARM15]. Elman [BS15]. Embarrassingly [SZR17]. Embedded [ADMX^{+}12, ASYK^{+}19, CCT10, CCL13, DLC^{+}16, GG10, GHZZ16, MVL15, MRGR12, NLGQ14, PG16, PGGS19, RSR11, RGRM14, TCM18, VMB17, XZX^{+}17]. Embedding [CH15, EMW16, LHJ12, PHXL19, TWW^{+}15, Wan12]. Emergency [CCT16, LLS13, WZQY14]. Emerging [Jun17, WFZ^{+}17]. Empirical [AF19, JKVA11, KCYM10, LLY^{+}15]. Emulating [ACE^{+}19]. En-Route [GKKW16, LYGX12]. Enable [XAY^{+}14, ZJL^{+}17a]. Enabled [GTM^{+}17, LDLL18, LGW^{+}17, Pan14, TMMN15, WKW16, WMS^{+}19, ZLYL19]. Enabling [BH13, CL14, CTG^{+}19, ECW^{+}18, FRS^{+}16, GYW^{+}19, KPG^{+}12, LHL17, LLS14, LH16, LGXL19, MCRC17, PG16, WWR^{+}11, WCRL12, WWL^{+}15, WZL^{+}19, ZY13, ZLCZ14, ZLW^{+}18]. Enclosure [WCF10]. Encoding [HW13, HWQ^{+}15, IZA18, SLSG18, WXYX14, ZHX^{+}19]. Encoding-Aware [SLSG18]. Encrypted [CWL^{+}14a, CWL16, FCM14, FRS^{+}16, GYW^{+}19, XWSW16]. Encryption [GZZ^{+}13, HSMY12, LYZ^{+}13, LHL^{+}14, She14, TKR14, XWLJ16, XWS17]. End [HKA12, HWX12, KOPS10, KAV^{+}17, LZ12, LCZZ13, YSS^{+}17]. End-to-End [HWX12, KAV^{+}17, LZ12, LCZZ13, YSS^{+}17]. Endurable [XX16]. Endurance [APPG16]. Endurance-Limited [APPG16]. Energy [AHTD18, AAB16, Amm12, ACV17, BCTB13, BSD^{+}18, BLLP15, CCJ19, CJZ12, CDBQ12, CLYR16, CZL^{+}18, CM10, CLKR15, CLHK11, CCD^{+}15, DCW^{+}15, DGF12, FBCB18, GFS^{+}10, GYQW15, GF13, GGF^{+}14, HLZY15, HAZ17, HCY^{+}12, HA10, HJS^{+}11, HGC12, JHR^{+}14, JJW11, JGZZ14, JHWY19, KKC18, KPB19, KPG^{+}12, LMM18, LGOB17, LM17, LZ11, Lee12, LAV^{+}10, LWY^{+}13, LQK^{+}13, LG13, LdSS^{+}13, LTL14, LCLL15, LW15, LYH^{+}15, LGJ^{+}18, LLpC15, LS17c, LSL^{+}17, LH17, LA12, LGG^{+}14, MTX^{+}11, MZK19, MNG^{+}15b, MBF19, MJK14, MRGR12, NSH15, NTKK15, NLGQ14, OPM^{+}15, PCL15, PPS^{+}17, PD14, PAB13, RZH^{+}11, Ren14, SAEH19, SEAH16, SAF16, SZR17, SBMA15, SOTN12, TWT16, TWL^{+}15, TFM^{+}16, TMMN15, WQZ^{+}15, WPT10, WLS^{+}11, WW13, WVM19, WLLL10, XZX^{+}17, XLM^{+}12b, XLM12a, YLC^{+}16, YPL^{+}17, YJC15, YJCQ15, ZTA^{+}15, ZS10, ZYL^{+}17, ZDM^{+}17, ZQH13, ZHZC15, ZMW17, ZLT^{+}18, ZTZQ19, ZHCW12, ZSB^{+}13, ZGKB16]. Energy [ZR18, ZHS^{+}19, dLMPG19]. Energy-Aware [Amm12, CLYR16, CLKR15, HAZ17, LMM18, MNG^{+}15b, SAEH19, SZR17, YLC^{+}16, ZHZC15]. Energy-Balanced [RZH^{+}11, WPT10]. Energy-Based [ZYL^{+}17]. Energy-Cognizant [ZSB^{+}13]. Energy-Constrained [LG13]. Energy-Efficiency [MJK14]. Energy-Efficient [AHTD18, ACV17, GYQW15, HCY^{+}12, HA10, JHR^{+}14, JJW11, JGZZ14, JHWY19, KPG^{+}12, LGOB17, Lee12, LAV^{+}10, LdSS^{+}13, LTL14, LS17c, MTX^{+}11, MZK19, MBF19, MRGR12, PAB13, TWL^{+}15, TMMN15, WLS^{+}11, WLLL10, XZX^{+}17, XLM^{+}12b, XLM12a, YPL^{+}17, ZS10, ZDM^{+}17, ZTZQ19, ZHCW12, ZGKB16, ZR18]. Enforced [SYL^{+}16]. Enforcement [LC11]. Enforcements [HZT18]. Engine [IG11, MMYES^{+}18, QP16c, WTL10, WZL^{+}16, ZHCL17, ZKSY14, KBS11]. Engineering [ABE^{+}11, SBC^{+}19, SM16, Sto10f, TP13, XSL^{+}16]. Engines [DSASSLP12, FHW11, LTC^{+}19]. Enhance [MNZ^{+}15, ZWL17]. Enhanced [AAAK^{+}14, BJ13, CMV^{+}10, LYGX12, RYLZ10, YCPC15]. Enhancement [GDM^{+}13, IB14]. Enhancements [SKP12]. Enhances [WYX^{+}15]. Enhancing [AKT^{+}15, BCF13, LGJ^{+}17, RPYO11, SJR17, SLSL16, WSWY15]. Enough [BKL11, CL13]. Ensembles [LLN^{+}19]. Ensuring [CLHK11]. Enterprise [sCCyW14, XHZ^{+}13]. Entities [GLZ11]. Entity [LAT^{+}15, LGZ^{+}19]. Entropy [GIP^{+}13, LZL^{+}18, YZDJ11]. Enumeration [RMG14]. Environment [CLT^{+}17, GZWN14, HH13, LLJ^{+}13, LWC^{+}17, LZZP13, LIWJ15, WL12a, XSC13, XBZ^{+}16, YSG^{+}14, YLC^{+}19, ZYW^{+}16]. Environments [AIAD^{+}18, BZA10, CJ10, CBK^{+}10, EHI11, FPF13, FGLP10, HMP^{+}19, HCH^{+}19, HYC^{+}12, HC14, JRP^{+}10, KL16, LC15, LSKZ13, LH15, PWJ16, RM17, SWT^{+}17, SCL^{+}15, TNZ^{+}12, TZ10, WTL10, WGG^{+}18, WZGR10, yWeH11, WSS15, WKW19, XTHD10, YHC^{+}13, ZWFX17, ZFG^{+}14]. Ephemeral [CE17]. Epidemic [ZWWF15]. Epistasis [GDRTS16]. EPPA [LLL^{+}12]. EPPDR [LLY^{+}14]. Equality [Hen14]. Equation [SS18]. Equations [HJ17, LYL16, WRWW13, CL16a]. Equilibria [RMG14]. Equivalent [AT12, KLWK12]. Era [DMCN12, YLJ^{+}17]. Erasure [CZT^{+}17, FSSZ16, HWQ^{+}15, HLQ^{+}15a, KZK^{+}19, KZK^{+}20, LL17, LHL17, LLRP18, LT10, LT12, SLSG19, WPMX18, XSZ13, ZLL17a, ZHL19, ZLX^{+}14]. Erasure-Coded [CZT^{+}17, HWQ^{+}15, HLQ^{+}15a, LLRP18, SLSG19, WPMX18, ZLL17a, ZHL19, ZLX^{+}14]. Erlang [CMT^{+}17]. ERPOT [aaGZ19]. Errata [KZK^{+}20, NHN18]. Error [DB18, DW13b, DC18, DTLC19, FPRG16, JHR^{+}14, KBHS14, KSP10, LLXC14, MGB18, ZFG^{+}14, ZWL17]. Error-Bounded [DC18]. Error-Correcting [KBHS14]. Error-Minimizing [LLXC14]. Error-Tolerant [DW13b]. Errors [JMA^{+}18, YLZ^{+}15a]. eScience [Li10]. Establishing [RM11]. Establishment [ZDG^{+}14]. Estimating [MM15]. Estimation [AB14, BAMJ12, DSM14, GCZ15, JIP14, KJL^{+}16, KCW11, QNLN11, RGLDM17, SMTZ17, SS17, WMW11, WHN^{+}19, YYY^{+}14, YZSC14, ZMLT13, ZYW^{+}14a, ZLL17c]. ESWC [GJLZ13]. Ethernet [KOKA11]. EULAG [LSW17a]. Eunomia [ZWJ^{+}18]. Evacuate [XLL^{+}18]. Evacuation [CWZ^{+}15, CCT16]. Evaluating [CJ16, CMT^{+}17, EAMEG11, FPRG16, LSCL16, QP16a, RS10]. Evaluation [ABBCT16, BSP10, BDLS13, BLLP15, CJ10, CB16, DLZH16, DCA19, FSM^{+}12, HBS^{+}16, IBC^{+}11, IG11, KCYM10, LT16, LZY^{+}19, LLS14, kL11a, LLY^{+}15, MSSB14, MMBdS14, NHN17, NHN18, Pan14, PSL^{+}11, PT15, RLY^{+}15, SLSG18, SLSG19, WJWX14, WL12b, WCF13, XXL^{+}19, ZT14, ZDF^{+}15, ZJKQ16, ZZCD10, ZW14, ZL10]. Event [CWCS15, GJZZ12, GCZ15, HCS12, LAV^{+}10, Lu14, NSLV16, PF12, PJAGW14, QCZ^{+}15, RKZC14, RCC^{+}14, SHM^{+}12, WLT^{+}12, YLT15, ZCJ19]. Event-Based [NSLV16]. Event-Driven [CWCS15, SHM^{+}12]. Event-Level [WLT^{+}12]. Events [DGG^{+}19, DWF12, HCY^{+}12, HH12]. Eventual [AR10]. Eventually [AEM17, BBR12]. Eventually-Consistent [AEM17]. Eviction [CHHK19]. Evictions [VBC19]. Evidence [MLML15, XP12]. Evolution [LLY^{+}14, MM15, Wan14, ZLZ^{+}17, KLL^{+}17]. Evolution-Cast [Wan14]. Evolutionary [SJVR19, SAF16, ZZLL16]. Evolutive [DSASSLP12]. Evolving [CMPS11, LHZJ19]. Exact [BF17, JMA^{+}18, LC14, MIH17, dOSMM^{+}16]. Exact-MBR [LC14]. ExaGeoStat [ALS^{+}18]. Example [PYH19]. Examples [SS12]. ExCCC [ZDM^{+}17]. ExCCC-DCN [ZDM^{+}17]. Exchange [CGS^{+}15, LY16b, YLW13, ZSY14]. Exchanged [LMLM13, TCT14, TCT16]. Exclusion [CGKP11, WZLC15]. Execution [BTL^{+}19, dCCF15, DD17, GTT^{+}17, GRJZ17, HLK^{+}19, KBS11, LWC^{+}17, LLD^{+}18, MGS12, MHL^{+}16, SP12, TRD13, WZL^{+}16, XALS17, XL17, ZLLD18, aaGZ19]. Execution-Efficient [ZLLD18]. Executions [ZH14a]. Expand [MWZX14]. Expansion [TL14, ZQWL17]. Expectations [SBC^{+}19]. Expedite [LNK17]. Expenditures [ARM16]. Experience [TWL^{+}15]. Experimental [CAC^{+}19]. Experts [ZLL^{+}15]. Explicit [CLL^{+}19]. Exploit [ZWE19, ZHW^{+}19]. Exploitation [LYW^{+}12]. Exploiting [AA17, AGG15, BS12, CZYL14, CJW16, CRZH15, CLKR15, CLA^{+}19, DT14, FFC17, GBD^{+}13, GHL^{+}13, GXZ^{+}15, HYZ15, HWQ^{+}15, JSMK11, JZH^{+}14, JZWN15, JN16, KJN15, LLL^{+}13, LG13, LLXC12, MWJ16, MHL^{+}16, QZZ^{+}16, WLT^{+}12, WCYL19, WK11, WSH^{+}19, XAY^{+}14, XGL^{+}16, YLLW16, ZLJL17]. Exploration [ABE^{+}11, KGI17, KM18, LSLD17, LZY^{+}19, SAEH19, Yan14]. Explorations [EHM^{+}17]. Exploring [CSV^{+}17, DGG^{+}19, HHK10, Jun17, KZK^{+}19, KZK^{+}20, LSL^{+}18, SLX19, SKKK16, WL12a, WKL^{+}16, WL12b, ZLK^{+}16]. Exponential [BCP^{+}14, ZLF^{+}11]. Exponentiations [Lou14]. Exposed [WWH13]. Express [ST18]. Expression [CJBW16, WPKL13]. Expressive [BTG^{+}18, YJ14]. Extend [LS17b]. Extended [CRS^{+}17, JEW^{+}18, KGK^{+}13]. Extending [HMS^{+}18, MJK14]. Extensibility [FGEL14]. Extensible [BFD19, GETFL14]. Extension [AELGE16, CMC^{+}15, HYX11]. Extensive [LLY^{+}15]. Extent [kL11a]. Extent-Based [kL11a]. External [ZML^{+}17]. Externally [LMR10]. Extra [LZXW15, LXZH16]. Extracting [FWZ^{+}16]. Extraction [JLW^{+}10, LJB^{+}13, WJTZ14]. Extrema [BAMJ12]. Extreme [GTM^{+}17, HAY^{+}18, WKL^{+}16, YC18, ZLK^{+}16]. Extreme-Scale [HAY^{+}18, WKL^{+}16, YC18]. Eyeball [XZH14]. Eyes [LODB17].

F2C [LH16]. FA [PH18]. FA-Stack [PH18]. Fabric [AVA^{+}17]. Face [MMNN16, WWCB14]. Factor [CHW^{+}17, HXC^{+}11]. Factorization [AHJ^{+}11, CRWY15, KLFD13, KAGD16, LLAL18, MVC^{+}18, OPJ^{+}19, ZHZL17]. Factorizations [HAZ^{+}18]. Fading [THL13, ZMA12]. Fail [HWC15]. Fail-Stop [HWC15]. Failed [Wan12]. Failure [CWLS19, FSSZ16, GTM^{+}17, HWC15, JRAS17, LSN19, PWT^{+}17, SSLF17, YTZ^{+}11, ZLL17a, ZYSH14]. Failures [BV10, FWH^{+}18, HP14, HWNS15, LL17, MLML15, MT15, PDH10, TKC^{+}15, YQZC12]. Fair [CFLL19, HSN17, HWL^{+}17b, KALK^{+}18, KCH19, LRJX13, LH16, MYPL18, TTH^{+}19, TYLG13, TCS11, WLL15a, WPT17, WLX^{+}15]. Fair-Progress [WLX^{+}15]. FairGV [HSN17]. Fairly [SSPG17]. Fairness [CJH^{+}14, CFLL18, hKYY11, KCH19, LZWY14, NN10, SLS^{+}16, TNH^{+}18, XXLZ16, XLM^{+}11a]. Fairness-Aware [XXLZ16]. False [LYGX12, LLZ^{+}12b, YYY^{+}14]. Family [BGE^{+}16, TTG^{+}15b]. Farewell [Sto13c]. Farm [HJS^{+}11]. Farms [ZJTZ14]. Farther [XSZ^{+}10]. Fast [AHS^{+}15, BAMJ12, CSS^{+}13, CZL^{+}16, CMK^{+}16, CJW^{+}19, CHPY17, DSM19, EHM^{+}17, GBFS16, HMKG19, HSN17, HJ17, JZW^{+}14, KTK11, Ksh10, LAK11, LWT^{+}18, LWL^{+}19, LCD^{+}17, MJM16, PH18, PYH19, QLC14, QP16b, QJ16, RCM16, SLG10, SS18, SRG19, TTG^{+}15b, TCS13, THL13, VTSM12, YXWW14, YZH^{+}19, ZS17, ZLW^{+}14, ZLL17a, ZLW^{+}19, KLL^{+}17, AAB^{+}17]. Fast-Fading [THL13]. Fast-Sweeping [SS18]. Fat [AP17, DY16, KEGM12, MYPL18]. Fat-Tree [DY16, MYPL18]. Fatal [DGG^{+}19]. Fault [AP17, ASYK^{+}19, BKY15, BG13, BGE^{+}16, Che16, CCH^{+}17, CYW^{+}18, CLH13, CH15, DNW^{+}16, EN12, GLJ^{+}15, GLC^{+}15, HWC15, JHYK11, KTK12, KZK^{+}19, KZK^{+}20, LMR10, LL12, LKT11, MNZ^{+}15, PWT^{+}17, SAEH19, STK^{+}19, TZY^{+}18, TCT14, WGG^{+}18, Wan19, WYL19, XS11, YDH17, ZJL^{+}12, ZCX^{+}14, ZWQ^{+}15, ZWG^{+}16]. Fault-Tolerance [CYW^{+}18, Wan19]. Fault-Tolerant [ASYK^{+}19, BKY15, BGE^{+}16, CCH^{+}17, CH15, GLJ^{+}15, GLC^{+}15, JHYK11, KZK^{+}19, KZK^{+}20, SAEH19, TZY^{+}18, WGG^{+}18, WYL19, YDH17, ZCX^{+}14, ZWQ^{+}15, ZWG^{+}16]. Fault/Intrusion [ZJL^{+}12]. Fault/Intrusion-Tolerant [ZJL^{+}12]. Faults [CIH13, FPGAD10, LAdS^{+}15]. Faulty [CH15, LLH14, WHH^{+}13, XS11]. Favors [JKS13]. FCoE [WWH^{+}17]. FCoE-Based [WWH^{+}17]. FDAC [YRL11]. Feasibility [CL13, GHL14, IIKO13, LLLH19]. Feasible [ESGQ^{+}13]. Feature [EK10, WYW13, WJWX14].

Feature-Based [WJWX14]. Federated [CSP13, WSSZ13]. Federation [Sam14a]. Feedback [LZY12, PH11, SCH11, TCDMRP17]. Feedback-Control [TCDMRP17]. Feeding [LGYV14]. Fellow [DK17]. Femtocells [AJMW14]. Femtocellular [PSMD18]. Fence [HZG^{+}17]. Fence-Free [HZG^{+}17]. Fermi [KTD12]. Fetching [WMS^{+}19]. FFT [WJB14]. FFT-Based [WJB14]. Fibonacci [GFJT19]. Fidelity [CTX^{+}12, SHX^{+}10]. Fidelity-Aware [CTX^{+}12]. FiDoop [XZQZ17]. FiDoop-DP [XZQZ17]. Field [BHS^{+}19, LC14]. Fields [LAT^{+}15, LWJ^{+}15]. FIFO [ME15b]. File [CTLH14, CSC16, CAJ^{+}16, CSSL15, CSY16, CLL^{+}19, ECW^{+}18, FHH^{+}15, GGY^{+}19, HWS16a, HCSC13, HZJ^{+}11, HJZ^{+}12, HjZ^{+}14, IRSNF11, Li14a, LHL17, LS17a, LLS^{+}18, kL11a, LY16b, LLC10, LS17c, Mit17, RSW^{+}17, She10a, She10b, SL13, SLW15, SLC15, SS17, STMM17, TCYF16, WMZ^{+}15, WYCZ14, WMJ^{+}19, XHL^{+}11, XAYM14, YZHZ17]. Files [FHH^{+}15, RY14, ST18, WJ12]. Filter [LGXL19, QZW14, XXWY10]. Filtered [AKC^{+}15]. Filtering [Has16, LYGX12, LLAL18, LLZ^{+}12b, THE^{+}15]. Filters [AKC^{+}15, BGHG16, GHL14, LWL^{+}19, MLVD12, QLC14, RCM16, XH10, ZS17]. Finding [ACS13, KBHS14, MLT^{+}13, ZLL^{+}15]. Findings [HSX^{+}12]. Fine [HAU19, IMH12, KMM13a, LKBK11, LZWZ19, LH16, MWZ^{+}13, NML^{+}14, Rao14, SYL^{+}16, TCM18, TWW^{+}18, WJWX14, YRL11, YBY^{+}18]. Fine-Grained [HAU19, KMM13a, LKBK11, LZWZ19, LH16, MWZ^{+}13, NML^{+}14, Rao14, SYL^{+}16, TCM18, TWW^{+}18, WJWX14, YRL11, YBY^{+}18]. Finessing [GAKR11]. Fingerprinting [LJG12, SL11, SCHT16, ZJL^{+}12]. Finite [PBD^{+}13, XFL^{+}19]. Finite-Difference [PBD^{+}13]. Firewall [LC11, LDYZ15]. Firework [ZZSZ18]. First [BBM16]. Fit [GHW^{+}16, DCL^{+}10]. Fitness [WKW16]. Fitness-Enabled [WKW16]. Five [YL15]. Five-Round [YL15]. FiWi [NTKK15]. Fixed [MG18, QF14, WGZ16]. Fixed-Priority [QF14]. Fixing [LL17]. Flash [CTLH14, HYZ15, KZK^{+}19, KZK^{+}20, LLZ^{+}12a, LZW^{+}17, SVK^{+}19, Ven14, WX15, WMLJ17, XX16, YZJ^{+}12]. Flash-Based [WMLJ17, XX16]. Flattened [LLS^{+}18]. Flexible [DG15, DCL^{+}10, GS17, GRJZ17, HCJ^{+}10, JKT11, Tsa13, WZL^{+}19, YZL^{+}17, YQ16, YBY^{+}18]. FlexRay [Fen14, GHZZ16]. FlexRay-Based [GHZZ16]. Flip [KSP10]. Flip-Error-Resistant [KSP10]. Floating [SY17]. Floating-Point [SY17]. Flood [rCHG10]. Flooding [BCP^{+}14, FFC17, GS11a, KCK14, YK14]. Flow [CGZQ13, EHWX10, HH11, hKYY11, MWJ^{+}14, QZG^{+}16, SILJ11, WL13, XJY^{+}10, YZJ^{+}12, ZQWL17, ZRS18, ZBK^{+}15, YGS^{+}19]. Flow-Based [ZBK^{+}15]. Flows [DWW^{+}15, HL12b, LYH^{+}15, MYPL18, WSSZ13]. Fluid [SY17, dSLMM11]. fMRI [Has16]. Focused [AZW^{+}19]. Fog [JWNS19, LS17b, WMWW19, ZLYL19]. Fog-Computing [WMWW19]. Fog-Enabled [ZLYL19]. Folded [CMB18, Tan12, YLJ^{+}17]. Footprint [VBC19, CQZ^{+}12]. Force [ADLM19]. Force-Directed [ADLM19]. Forced [SL14]. Ford [BB16]. Forecasting [TZC19]. Forest [BYZ^{+}16, CLT^{+}17]. Fork [Che11]. Fork/Join [Che11]. Form [HCH^{+}12, LKD10]. Formal [DIAR16, MGS12, RAS17, SL11, YHC^{+}13]. Format [KGK^{+}13]. Formation [DMR16, KP12, MG14, SLM^{+}10, YZS13, YC14]. Formats [JHMV12, LT16, TTG^{+}15b]. Formed [MSB11]. Formulation [Tak14]. Formulations [VS15]. Forward [FLH13, JMA^{+}18]. Forwarding [Cha14, Fre13, HWX12, JGG^{+}11, LWY^{+}15, LT12, LW12, NTK^{+}15, WDOX15, YXG12]. FoToNoC [YLJ^{+}17]. Fourier [XAK17]. FP [AHS^{+}15]. FP-NUCA [AHS^{+}15]. FPGA [CP17b, OZMC^{+}16, QP16b, QP16c, SHY14, SY17, TZT^{+}16, TP18, WTTH17, WZL^{+}16, WLC^{+}17, WM18, ZTZ^{+}18a, ZKP^{+}19]. FPGA-Based [SY17, WLC^{+}17]. FPGA-Platform [WTTH17]. FPGAs [ECV16, HA13, MS15, RCK15, SLX19, WZHZ16]. FPS [WLX^{+}15]. Fractional [SVC12]. Frame [GYX^{+}10, LW15, dLMPG19]. Frame-Based [LW15]. Framework [AAAK^{+}14, Amm12, AKP14, CJZ12, CC18, CLL11, sCCyW14, CJZ^{+}16, CMG^{+}14, DLB^{+}19, DY17, EAMEG11, EHNS13a, GZW^{+}18, HL12a, HWF18, HXC^{+}11, JHMV12, JJW11, JCW^{+}12, LPP13, LLG15b, LLLZ16, LWY^{+}19, LGZ^{+}19, LZH^{+}16, LLXC14, LDYZ15, LLS13, LLH^{+}15b, MTY^{+}12, RAS17, RYLZ10, RS12, SS12, SAA17, SAB^{+}18, TTG^{+}15a, TYWL14, TLL^{+}16, THB^{+}14, VBC19, VT19, WZHZ16, WGG^{+}18, XL13, XSL^{+}16, YBY^{+}18, ZTG^{+}18, ZWFX17, ZGGW13, ZGGW14, ZWL^{+}16b, ZJS^{+}17, ZWJ^{+}19, ZMTL15, ZKP^{+}19]. Frameworks [LGL^{+}18b, LN17]. Fréchet [GV15]. Free [AS16, BRX13, BS14, Dua95, FVLD16, GAB18, HZG^{+}17, IPQ19, JEW^{+}18, KCK14, KB17, KWG17, KSP10, LX12, MMYES^{+}18, ME15b, NML^{+}14, PH18, PYH19, RGBC11, SHG11, VS11a, VS11b, VS14, XL16, YYL^{+}17, ZZG^{+}11, ZLGN13, ZYZ^{+}14, ZD16a, ZH11, ZYW^{+}14b, Dua93]. FreeRider [LCL^{+}15]. Freeweb [SLLZ16]. Frequencies [ZLY^{+}14]. Frequency [CCL13, LYW^{+}12, LZC^{+}12, XXWY10]. Frequency-Temporal [LYW^{+}12]. Frequent [LZC^{+}12, OUA11, RGK15, SZ11, XZQZ17]. Freshness [ZWZ^{+}15]. Freshness-Aware [ZWZ^{+}15]. Friendly [LLC10, WDC12, WSS15, ZH18].

Friendship [BS12]. Frugal [CSC16]. FS2You [LSL^{+}10]. FTL [ZHW^{+}19]. Full [CJL^{+}12, CPH^{+}18, RMB^{+}16, ZWL^{+}16b, Zhu14]. Full-Duplex [Zhu14]. Full-Scale [RMB^{+}16]. Full-System [CPH^{+}18, ZWL^{+}16b]. Full-Text [CJL^{+}12]. Fully [HA13, MWJ^{+}14]. Function [CWL14b, CTG^{+}19, LHXP18, MLXG19, MKSN18, PHXL19, RKRK17, SG16a, TZC19, XDLZ19]. Functional [JSC^{+}17, SytL19, ZWT^{+}19]. Functional-Unit [JSC^{+}17]. Functions [Fre13]. Fundamental [LLZ^{+}12a]. Further [HCL^{+}14]. Fused [BG13]. Fusion [ALI^{+}17, CTX^{+}11, LTMD11, MLML15, MV12, TXWL11]. Fusion-Based [CTX^{+}11, TXWL11]. Future [GXZ^{+}15, WUH^{+}17]. Fuzzy [HML^{+}14, PGP^{+}17].

G [ATZZ14, KMM12, DWH^{+}18, LWCL18, ZJZ^{+}16]. G-CRS [LWCL18]. G-ML-Octree [DWH^{+}18]. GALS [MGS12]. Game [Che15, HLYJ19, KP12, LLW^{+}15, Tak14, TKP12, XZSG12, YLC^{+}16, YC14, ZKSY14, Che18b, ZCJ19]. Game-Based [Che18b]. Game-Theoretic [KP12, YC14, ZKSY14]. Games [GE12, NIP11, RMG14]. Gaming [CZX^{+}19a, GYQW15, LS17b, ZYQ^{+}14, ZQCZ16, CZX^{+}19a]. Gap [AAB^{+}17]. Gateways [AJMW14]. Gather [Trä19]. Gathering [IIKO13, LKE16, MKOK14, RZH^{+}11, XHQ^{+}15, ZYT^{+}15]. Gating [LWW^{+}13]. Gating-Induced [LWW^{+}13]. Gaussian [BSF16, FB10, Tou15b, WFA13]. GBC3 [LY16a]. GC [WMLJ17]. GC-Aware [WMLJ17]. GCA [RKGS16]. GCache [YLC^{+}19]. Gearing [SCH^{+}15]. GEMM [KTD12]. Gene [ZASA10, CSR^{+}17, IBC^{+}11, ZYL^{+}16]. Gene/Q [CSR^{+}17, ZYL^{+}16]. General [ABBCT16, BBGD^{+}17, DSJ16, DLB^{+}19, JCW^{+}12, LCL^{+}11, OOA^{+}14, STMM17, WJTL13, WM15]. General-Purpose [STMM17]. Generalization [QLC14, RCM16]. Generalized [FMY^{+}18, GS11a, SRB14, TWL12, XSL^{+}16]. Generated [CSZ^{+}12]. Generation [AAB16, CC17, CP17b, FBCB18, HJZ^{+}12, LMVS11, LPMB13, LLFL15, PT15, RSSC15, TTG^{+}15a, VPS17]. Generational [SJVR17]. Generator [YLZ^{+}15b, ZR18]. Generic [HXC^{+}11, ZLW^{+}18]. Genome [LPMB13, MTY^{+}12, ZASA10]. Genome-Wide [ZASA10]. Genuine [PRR^{+}16]. Geo [HLL18, LGM^{+}17, LV17, PHXL19, SWL17, THT^{+}15, XBZL17, XFL15, ZLZ^{+}16, ZHCL17, ZXG^{+}19]. Geo-Distributed [HLL18, LGM^{+}17, PHXL19, SWL17, XBZL17, XFL15, ZLZ^{+}16, ZHCL17, ZXG^{+}19]. Geo-Diverse [THT^{+}15]. Geo-Replicated [LV17]. Geocast [JZH^{+}14]. Geocommunity [FCD^{+}13]. Geocommunity-Based [FCD^{+}13]. Geographic [CNC^{+}14, RRS12, WWLX13, ZS10]. Geographical [CMG17]. Geographically [SL13]. Geolocating [TDLR13]. Geolocation [LCG^{+}13]. Geometric [CCFS11, DDP^{+}19, LMSRSR13, Yan14]. Geometries [TS18]. Geostatistics [ALS^{+}18]. GFlink [CLO^{+}18]. GKAR [WWLX13]. Glance [LLY^{+}17]. gLite [BSP10]. Global [CP15, CLL^{+}17, GGS10, HH11, KCH19, Ksh10, LS17d, MGB18, NX95, NN10, Tsa13, WGZ16, WYX^{+}15, XLT^{+}14, ZLL17c, ZLLD18, KLL^{+}17, RKGS16]. Global-Snapshot [Tsa13]. Globally [FC11, JKP12]. Globally-Coordinated [JKP12]. GMRace [ZRQA14]. GMU [PRR^{+}16]. Gnutella [BZA10]. Go [XSZ^{+}10, BWH^{+}19]. Good [YLM^{+}15]. Goodput [WYC^{+}15]. Goodput-Aware [WYC^{+}15]. GOP [HW13]. Gossip [IvS10, KN16, SRG19]. Gossip-Based [IvS10]. Gossiping [HWDP10]. Gossips [LNK17]. GPGPU [AHJ^{+}11, FPRG16, HH13, HA11, KZW17, KPB19, LLW^{+}15]. GPGPUs [TCYF16, WWJ^{+}18]. GPU [ABLS16, BBK17, BB15, BB16, BB17, CC18, CRWY15, CLO^{+}18, CEK16, DB18, EALM15, EALM17, GRUMG17, Goh14, GLGLBM13, GC16, GRB^{+}19, GYQW15, GV15, GIRT19, HAZ^{+}18, HSN17, JDB^{+}14, JNL^{+}15, KLL^{+}17, KJN15, KTD12, LYL15, LYL16, LHR^{+}15, LLL^{+}14a, LWCL18, LLK^{+}14, LAD16, MC17, MIH17, Mit17, MLK15, Mur12, NVBH18, OOA^{+}14, Pan14, PS19, RRM^{+}15, RMG14, RSNV18, RBH^{+}14, dOSdM13, dOSMM^{+}16, SLX19, SA11, SKA15, SYXL16, SCHT16, SFA^{+}17, TLH^{+}14, TTH^{+}19, TTG^{+}15b, VMP17, VNA^{+}16, VT19, WTD17, WYZ^{+}19, XML^{+}18, ZM13, ZYQ^{+}14, ZZH^{+}17, ZWE19, ZRQA14, ZH14a]. GPU-Accelerated [CRWY15, LLL^{+}14a, SLX19]. GPU-Architecture [VMP17]. GPU-Aware [Pan14]. GPU-Based [GRUMG17, RMG14, SKA15]. GPU-Job [PS19]. GPU-Resident [JDB^{+}14]. GPUDirect [CLA^{+}19]. GPUs [AHTD18, AKGR13, BFD19, BF17, BHKS^{+}17, DKS^{+}15, DWH^{+}18, GS11b, GWC14, HKE^{+}16, IMH12, KEGM12, KAGD16, LLAL18, Nov15, PSL^{+}11, PB19, QJ16, RCK15, TS16, WQZ^{+}16, WSH^{+}19, WJB14, XFL^{+}19, YNK^{+}17, YOK^{+}17, YBY^{+}18, ZL14, ZH14b, ZSC^{+}17, dLMPG19, JMZD12]. GPUSCAN [SKA15]. Gradient [CCV19, GHL14, WSH^{+}19]. Gradient-Based [GHL14]. Grafting [ABP17]. Grain [ATA18, CA13]. Grained [HAU19, IMH12, KMM13a, LKBK11, LZWZ19, LH16, LLD^{+}18, MWZ^{+}13, NML^{+}14, Rao14, SYL^{+}16, TCM18, TWW^{+}18, WJWX14, YLL^{+}17, YLLW16, YYL^{+}17, YRL11, YBY^{+}18]. Grammars [DIAR16]. Graph [AHSK17, AAA19, AZW^{+}19, CYW^{+}18, CJW^{+}19, DCC^{+}19, DO13, EJRB13, FMY^{+}18, HZJ16, Hen14, HLYJ19, LC10, LGX^{+}11, LHCM^{+}17, LGXL19, MSSV18, MTMR18, MSS17, PLG19, YTMS16, YHL^{+}18, YXLJ16, YLC^{+}19, ZGGW14, ZLS^{+}18, ZH14b, ZSC^{+}17, ZKP^{+}19, ZYSH14, MTMR18]. Graph-Based [HZJ16]. Graph-Parallel [YTMS16]. GraphCT [EJRB13]. GraphD [YHL^{+}18]. Graphic [DFGG13, LLLC17, TS18]. Graphics [CCHH19, FHLG11, XML^{+}18, vdLJR11]. Graphine [YTMS16]. Graphs [ABP17, CMB15, CH14, CLH13, CH15, CYC^{+}16, CCK12, CMPS11, FWZ^{+}16, JSK18, JLKG17, JWY^{+}18, LKM10, LMSRSR13, LCD^{+}17, Ozd19, RGBC11, SOK^{+}19, TWL12, WKC12, YTMS16, YCWL14, YKN^{+}19, YN17, ZML^{+}17]. Greater [ZDM^{+}19]. Greedy [CNMA11, HWX12, NMG15]. Green [BLLP15, FBCB18, LSL^{+}17, LGG^{+}14, YXWL16, YC18]. GreenDB [ZTZQ19]. Greening [GTS^{+}15]. GreenOrbs [LHL^{+}13b]. Grid [ANE12, BMR15, BMJ^{+}17, DM11, DN19, FGLP10, HCZ12, Hur13, ICN18, LLY^{+}14, LYY16, LLFL15, LA12, MSW^{+}12, NSH15, PCFP16, SME10, WRB11, WHYZ10, XLL11, YQH16, dBK11, CJZ12, GPF12, LJ15, LLL^{+}12, MBO15, ZJLS12, ZHQ12]. Grids [BMJ^{+}17, BSP10, HPG14, Li10, MG14, MBH^{+}10, MTY^{+}12, QLC13, SGGB14, Tak14, ZYSH14]. Ground [LWW^{+}13, ZS13]. Group [GLL11, HJ17, HCyW^{+}17, JKT11, JN16, LC12b, LZXW15, SPB^{+}10, TXL^{+}14, TW14, XSTZ10]. Group-Based [SPB^{+}10]. Group-Ordered [HJ17]. Group-Strategyproof [LC12b]. Group-Testing-Based [XSTZ10]. Grouping [ANN^{+}13, LWX^{+}11, LYGX12, LNZ^{+}13, ZJZ^{+}16]. Grouping-Based [ZJZ^{+}16]. Grouping-Enhanced [LYGX12]. Grouping-Proofs-Based [LNZ^{+}13]. Groups [JCWB10, LZWY13, ZJ16, WZS^{+}18].

GroupTrust [FLLS17]. GSPNs [BSP10]. GT [Tak14]. GT-CFS [Tak14]. GTDAR [Che18b]. Guarantee [LZ12, LZWY14, LCW11, NTWL11, PYHY16, PH18, SAEH19]. Guaranteed [DWY^{+}13, HLCB^{+}17, LGD14, LSW16, LSW17b, NLGQ14, TWL^{+}15, ZWL^{+}18]. Guarantees [DG15, GYQW15, LCSC12, PFAF16, YJCQ15]. Guest [BKK11, CLL^{+}14, MBMC13, PKL^{+}12, RFZ11]. Guide [HAZ^{+}18]. Guided [YLC^{+}19]. Guidelines [TGT10]. Guiding [CCT16]. GVTS [KCH19].

H [CHW^{+}17, QCZ^{+}15]. H-PARAFAC [CHW^{+}17]. H-Tree [QCZ^{+}15]. Hadoop [BYZ^{+}16, CZT^{+}17, CZL^{+}18, GLBJ18, GRCZ17, GRJZ17, HZB^{+}16, JHWY19, KJL^{+}16, LAT^{+}15, LSLD17, LS17a, SCH^{+}15, XZQZ17]. Hamiltonian [JP12, Wan12, YL15]. Hamiltonicity [CLH13, LLH14]. Handheld [JGZZ14]. Handles [Ano12h]. Handling [SKGC14, SDG17, TS18, WV17, ZZQ18]. Handoff [MM12]. Hard [DC18, SEAH16]. Hard-to-Compress [DC18]. Hardware [AFA12, ASG^{+}14, CHM^{+}13, CSV^{+}17, CWS12, CD13, CLA^{+}19, EADT19, EHI11, GHZZ16, HT16, LZL^{+}18, MC14, MKSN18, OZMC^{+}16, PGGS19, QGPZ13, RX11, SAA18, SSPG17, TCYF16, TBÁ^{+}19, TGNA^{+}13, TGAG13, WH16, WZL^{+}16, WGHP11, XL10, ZS17, vdLJR11]. Hardware-Acceleration [WH16]. Hardware-Oriented [LZL^{+}18]. Hardware-Transactional-Memory [SAA18]. Harmonic [QF14]. Harmonic-Aware [QF14]. Harmonically [GHW^{+}16]. Harnessing [HLK^{+}19, WRWW13, CL16a]. HaRP [PT11]. Harvesting [LRJX13]. Hash [KHK15, RRS12]. Hashing [GZX14, LLLC17, PT11, RRS12, SHF^{+}17, ZH18]. Hazards [MM15]. HDR [YTL^{+}10].

HDR-WPAN [YTL^{+}10]. Head [TMMN15]. HEADS [HZB^{+}16]. HEADS-JOIN [HZB^{+}16]. Healing [SAM14b]. Health [HGY^{+}14, LYZ^{+}13, LCS^{+}15, SF10]. Healthcare [LLS13, ZLDC15]. Heaps [GFJT19]. HeteroCore [ZWE19]. Heterogeneity [CP17a, CZD^{+}19, FBCB18, HWS16a, LCLL15, SKKK16, ZFT^{+}15]. Heterogeneity-Aware [HWS16a]. Heterogeneous [Agr14, BKY15, BEDCR13, BBD^{+}19, BICK^{+}15, BSM^{+}11, BBL^{+}16, CJ10, CWL14b, CLT13, CZWZ14, CLYR16, Che16, CLO^{+}18, CRG^{+}17, CZL^{+}18, CVM^{+}15, CTG^{+}19, ECV16, GDRTS16, HGS^{+}19, HP14, HL12a, HMP^{+}19, HL12b, HKkY^{+}16, ITL17, JWK^{+}16, JZY^{+}15, JSC^{+}17, KHN16, KALK^{+}18, KAG17, LMM18, LMD16, LAV^{+}10, LTL14, LW15, LSB^{+}18, LZY^{+}18, LSJ^{+}19, LSL^{+}18, LLZ18b, MLS15, MNG15a, MC10, MA13, NHN17, NHN18, OPJ^{+}19, OOA^{+}14, OPM^{+}15, PPS^{+}17, PGP^{+}17, PH12, RSR11, RG17, RGLDM17, RDG12, SHA19, SG16b, SVL^{+}16, SP15, SBMA15, TFM^{+}16, TL16, VMB17, WTD17, WLL15a, WV17, XBZ^{+}16, XZX^{+}17, XLH^{+}15, YJCQ15, ZLZ^{+}17, ZM13]. Heuristic [CDR15, HH11, MM10, ZYW^{+}16, ZSW^{+}19, aaGZ19]. Heuristics [BSM^{+}11, CTA14, CJ16, CLYR16, CBF^{+}17, JTS^{+}11]. Heuristics-Based [JTS^{+}11]. HEVC [IZA18]. Hexagonal [ABF12, Tou15a]. hiCUDA [HA11]. Hidden [Hur13, XHX^{+}13]. Hide [YOK^{+}17]. Hiding [CCV19]. Hierarchical [CHM^{+}13, CWC11, CHW^{+}17, sFC12, FC11, HLZ^{+}19, JY15, LJ15, LWT^{+}18, LYPL19, NLY15, SMB^{+}18, SK14, VMP17, WCLK12, XTFC17]. Hierarchically [PHGR17]. Hierarchically-Scheduled [PHGR17]. Hierarchize [WCD^{+}11]. Hierarchy [APPG16, sCCyW14, CPH^{+}18, IvS10, MC17]. High [AHTD18, ALS^{+}18, AAW^{+}17, ARM15, BKK11, BCTB13, BKF^{+}16, BF17, BBL^{+}16, BSL^{+}17, CMB15, CB13, CP17b, CJW^{+}19, dCCF15, DCC^{+}19, DRRCB18, EHWX10, EAMEG11, EALM17, ESGQ^{+}13, FHW11, GFMR13, GFS^{+}10, HAZ^{+}18, HMKG19, HA11, HHWZ17, ITL17, JPG14, KOPS10, KMM13b, KL16, LJ16, LWT^{+}18, LHM12, LS17b, LCS^{+}15, LCL^{+}16b, LSL^{+}17, LSN19, MC14, MC10, MB12, MA13, MRGR12, NGJ^{+}19, NLC12, OPJ^{+}19, PH11, PB19, QZG^{+}16, QP16c, SG16b, SWT^{+}17, SLL13b, SHX^{+}10, WCF10, WL13, WKL^{+}16, WWJ^{+}18, WWT^{+}19, WJ12, WWLJ14, WZQ10, XX16, XSYY13, XLSR13, YKN^{+}19, YQ16, YWZ17, YR14, ZH14a, ZLT^{+}18, ZKP^{+}19]. High-Accuracy [XSYY13]. High-Availability [FHW11]. High-Bandwidth [CJW^{+}19, LHM12, XLSR13]. High-Density [WCF10]. High-Dimensional [NGJ^{+}19]. High-End [KOPS10]. High-Fidelity [SHX^{+}10]. High-Level [EAMEG11, HA11, PB19, YR14]. High-Performance [AHTD18, BKK11, BCTB13, BBL^{+}16, DCC^{+}19, EAMEG11, ESGQ^{+}13, GFS^{+}10, JPG14, LCL^{+}16b, MC14, MC10, MA13, MRGR12, OPJ^{+}19, PH11, QZG^{+}16, QP16c, WKL^{+}16, XX16, YKN^{+}19, YQ16, YWZ17]. High-QoS [SLL13b]. High-Quality [LCS^{+}15]. High-Scale [CMB15]. High-Speed [ARM15, BKF^{+}16, EHWX10]. High-Throughput [BSL^{+}17, HMKG19, LJ16, MB12, WJ12, WZQ10, ZH14a, ZKP^{+}19]. High-Utilization [WWLJ14]. High-Velocity [DRRCB18]. Higher [BSF16]. Highly [AEM17, KGR16, SBC^{+}10, TPRH16, YYL^{+}13, ZDM^{+}17, ZDM^{+}19]. Highly-Available [AEM17]. Hint [TRD13, WHC^{+}14]. Hint- [WHC^{+}14]. Hint-Based [TRD13]. Hints [AAH15, WHC^{+}14]. HIPA [MRH^{+}16]. HireSome [DZLC15]. HireSome-II [DZLC15]. Histograms [XHL^{+}15]. Historical [AHSH^{+}16, GZW^{+}18]. HitGraph [ZKP^{+}19]. HL [AJK^{+}17]. HL-PCM [AJK^{+}17]. Hoc [AE12, BZA10, CCFS11, CLM^{+}15, CPM^{+}10, CYL^{+}14, CLJ11, DMR16, GLJ^{+}15, HCJ^{+}10, JJ11, JGG^{+}11, LLGP13, LYW^{+}12, LMSRSR13, LNA^{+}13, LHYW15, MM10, MY11, She14, SCC11, SZZF10, SJ14, WJTL13, WL14, WCF13, ZZF10, ZHCW12, XAY^{+}14]. Hodgkin [CRS^{+}17]. HoL [NFD10]. Hole [SAM14b]. Holistic [Fen14, LGJ^{+}18, LCL^{+}16a, MZK19]. Home [LJ15, LLFL15, XWH15a, JKVA11]. Home-Based [XWH15a]. Homeomorphism [RBSS11]. Homogeneous [CYX^{+}14, Che11, LM17, MMNN16, ZM13]. Homology [IMH12, WKC12]. Homomorphic [ZJL^{+}12]. Hong [TTJX12]. Hop [HCH^{+}19, RWLL14, YXWL16, ZMA12, ZQSY13]. Hop-by-Hop [RWLL14, YXWL16]. Hopping [Mis14]. Host [YKN^{+}19]. Host-Switch [YKN^{+}19]. Hosting [LSL^{+}10, TVG13]. Hosts [BB13, HKA12]. Hot [WWX^{+}13]. Hotplug [LJL^{+}15]. Hotspot [MS12]. Hotspot-Locating [MS12]. Householder [MVC^{+}18]. HPC [APCH^{+}11, CB16, DLXS19, DC16, DRVC17, DC18, DGG^{+}19, DIAR16, ECV16, ESGG^{+}15, FKMC15, MHL^{+}16, MBV11, MBV13, MCRC17, MV18, NZM^{+}16, PGGS19, SMS^{+}13, TAZ^{+}19, UDH^{+}17, uRILP17, XGL^{+}16, ZTG^{+}18]. HPL [TZY^{+}18]. HSDC [ZDM^{+}19]. HSPA [TTJX12]. HTM [MPHR17, ZWJ^{+}18]. HTTP [XTXH13]. Huge [SJAdCL19]. Hull [GCZ15]. Human [LQY^{+}12, WYX^{+}15, ZW14, ZYW^{+}14b]. Hut [ZBS15]. Huxley [CRS^{+}17]. HV [SSF16a]. Hybrid [AVA^{+}17, ARM15, BHS^{+}19, BBK17, Bis18, CP17c, CCNMF18, DDW^{+}19, ESGG^{+}15, EJGYAM14, FFC17, GRB^{+}19, Hsi14, HXLF15, KKW18, LLY16, LdSS^{+}13, LTW^{+}14, LSL^{+}14a, LLC^{+}15, LYL16, LSLD17, LYPL19, LWZ^{+}16c, LGW^{+}17, MZLT19, PRS^{+}11, QJ16, RGLDM17, RJ16, SHA19, TWW^{+}15, VPS17, WPT10, XS10, XLH^{+}15, XWLJ16, YNKD18, YWWR18, ZZSZ18, ZMW17, ZWY^{+}17, XWS17, Gua14]. Hybrid-Double [ARM15]. HyConv [LZWZ19]. Hydrodynamics [RBH^{+}14]. Hydrology [LMD16]. Hyper [CLYR16, TXL^{+}14]. Hyper-Heuristics [CLYR16]. Hyper-Sphere [TXL^{+}14]. Hyperbolic [CYX^{+}14]. Hypercube [WYW13]. Hypercube-Based [WYW13]. Hypercubes [Lai12, TCT14, TCT16, YCPC15]. Hypergraph [AAA19, SAA17, YY10, YPL^{+}17]. Hypergraphs [QFZZ15]. Hypergrid [XHHC13]. Hypermesh [MS15]. Hypervisor [CL16b]. Hypocomb [LMSRSR13]. Hysteresis [BBCTA18]. Hytrace [DDW^{+}19]. HYVI [Gua14].

I/O [AZW^{+}19, BHEP14, CAC^{+}19, CRZH15, DIAR16, GDM^{+}13, HWS16b, HWL^{+}17a, LLJ^{+}13, LMFS11, NCM^{+}17, NLC12, PYHY16, SHA19, SSLF17, WXLY16, WWH^{+}17, ZWFX17, ZLJ^{+}15a, ZWJ^{+}19]. I/O-Efficient [WXLY16]. I/Of [HLQ^{+}15a]. IaaS [Bru14, LZY^{+}18, LH16, SLG^{+}18, TVRD17, WNLL15, WLL15b]. IaaS-Clouds [TVRD17]. iASK [LS17d]. IBM [FES^{+}17]. IBOM [WWJ^{+}18]. iDaaS [LGL^{+}18a]. Identification [ACCP12, FWH^{+}18, GG13, GIP^{+}13, JGZZ14, LZL10, LLM^{+}14, LXZB15, RX11, YQH^{+}15]. Identify [BTL^{+}19]. Identity [SZZF10, TKR14]. Identity-Based [SZZF10, TKR14]. Idle [IMH12]. IDM [LSKZ13]. IEEE [Ano11d, Ano11c, Ano12i, Ano15a, Ano16, Ano17a, FLH13, GYX^{+}10, MRM12, PDFJ13, RMM16, TMMN15, WYW^{+}14, ZZ15, Ano18, Ano19a, Par19b]. II [DZLC15]. ILBO [LX10]. ILP [VS15]. Image [Bar10, DB18, EALM17, GRUMG17, MRH^{+}16, MLK15, PSL^{+}11, WMZ^{+}15, WYZ^{+}19, ZJL^{+}17a, ZHS^{+}19]. Imageries [MWZ^{+}14]. Images [Li14a, WWL^{+}17]. Imaging [BKF^{+}16, RLVTMG^{+}16, WZQY14]. Imbalance [YDH17]. Imbalancing [LSW17a]. IMGPU [LLL^{+}14a]. Immersive [VMN^{+}16]. Immunization [GLZ11]. Impact [DC16, DMT12, EK10, FBCB18, Kum14, LLpC15, MRM12, PP12, TCYF16, Wan14, ZLF^{+}11]. Impact-Driven [DC16]. Impacts [Li10]. Imperfect [HLCH11, YLLW16]. Implement [SAA18]. Implementation [BB15, BB16, CLLX18, CL14, DLXS19, EALM15, EALM17, Fen14, HMKG19, KAGD16, LLC10, LWZ^{+}16c, Pan14, PDH10, RLY^{+}15, SLL16, SA11, SYXL16, TS18, WWL^{+}15, WZL^{+}16, WQZ^{+}16, XL10, YDC^{+}17, ZTG^{+}18, ZZCD10, ZL14]. Implementations [AH10, CHM^{+}13, DMS^{+}12, HXLF15]. Implementing [AHS^{+}15, BBR12, DGFRR18]. Implication [WFZ^{+}17]. Implications [BMJ^{+}17, CE17, LLZ^{+}12a, SJVR19]. Importance [TNLM17]. Improve [APPG16, HCL^{+}12, HWSX17, JSMK11, LCYW16, MWJ16, WHH^{+}13, XZT^{+}13, YLL^{+}17, ZQSY13]. Improved [Che18a, DCA^{+}16, MBV11, SSM^{+}18, SKKK16, TLP12, YJC^{+}16, ZLL17c]. Improvement [FRS^{+}16, SL14]. Improves [LWZ14, WBPF11]. Improving [ATA18, BHEP14, CTA14, CGZQ13, CRG^{+}17, CD13, DBAT11, FES^{+}17, GTT^{+}17, GRCZ17, HYZ15, HWS16b, HWX12, KPB19, LLK^{+}14, LXBZ13, MLC^{+}19, MV16d, NZWL14, Ozd19, PPR10, TWW^{+}18, TZ10, TSN10, TGNA^{+}13, TP13, WLH^{+}15, WL15, WMLJ17, WHZS19, WMJ^{+}19, ZTA^{+}15, ZYL^{+}16, ZSW^{+}19, dLMPG19]. IMR [LCL^{+}16b]. IMS [BCF13]. IMS-Based [BCF13]. In-Home [LLFL15]. In-Memory [CLO^{+}18, CHHK19, CRRR15, HWSX17, MZL^{+}19, TZY^{+}18]. In-Network [CCCY16, PCP14, ZMLT13]. In-Place [SLL16]. In-Situ [HLK^{+}19, HHK10, VLP16]. Inbound [LX10]. Inc-Part [ZLJ^{+}15b]. Incast [Guo17, ZRTL15]. Incentive [CSY15, TZB^{+}14, WGCG18, WZQ10, WML14, ZYZ^{+}14, ZWZ^{+}15]. Incentive-Driven [TZB^{+}14, ZWZ^{+}15]. Incentives [CLL11, XZSG12]. Incentivized [LFLW10]. Inclusion [SYXL16]. Inclusion-Based [SYXL16]. Inclusive [MIH17]. Incomplete [NCKL14]. Incorporating [LCLL15, LS17d]. Increase [CIP^{+}17]. Incremental [JSK18, PB12, dOSMM^{+}16, ST18, ZLJ^{+}15b, ZDM^{+}19]. Incrementally [XDMZ17]. Indefinite [YKW^{+}18]. Independent [BHKS^{+}17, CTA14, CFJ15, FCM14, Tic14, Tse13, YCPC15]. Index [Ano11a, Ano12a, Ano13a, Ano14a, Ano15a, Ano16, Ano17a, Ano18, Ano19a, DWH^{+}18, DR16, Hsi14, LTC^{+}19, LAD16, MZL^{+}19, QCZ^{+}15, TXZ^{+}11, ZWJ^{+}18]. Index-Digit [LAD16]. Indexed [SLL16]. Indexing [GC16, KJN15, WL13, ZLZ^{+}14]. Indices [Has16]. Indirect [ALI^{+}17, BH13, BGE^{+}16, LSKZ13]. Indistinguishability [LWL^{+}17]. Indoor [GZWN14, TLJ^{+}14, WXY^{+}13, WYLX13]. Induced [GGGA18, LWW^{+}13]. Industrial [HH15, HCH^{+}19, RMB^{+}16, SS12]. Inefficient [ECW^{+}18]. Inertial [TLJ^{+}14]. Inference [AF19, BFFG11, DNW^{+}16, HML^{+}14, LAdS^{+}15, YGL13, ZFG^{+}14]. Inferring [SJVR15]. InfiniBand [MMYES^{+}18]. InfiniBand-Based [MMYES^{+}18]. Infinite [CEK16]. Influence [LLL^{+}14a, SZWX15, WJWX14]. Influxes [ZLF^{+}11]. Inform [Amm12]. Information [AB14, CZYL14, CMPS11, DWLY15, GCZ15, HLCH11, JMS^{+}18, LTBN^{+}12, LCL^{+}15, MPS15, PCP14, SGC14, TL14, TYG^{+}14, TNLM17, US16, YQH16, ZXW^{+}13, ZW14, ZASA10, ZBK^{+}15]. Information-Based [MPS15]. Information-Centric [PCP14]. Information-Theory-Based [ZASA10]. Informed [HZW^{+}19, KI14]. Infrastructure [KAV^{+}17, LPSS19, PJC^{+}13, PT15, QTC^{+}14, SLGW14, ZX13, ZHQ12, DNW^{+}16]. Infrastructure-as-a-Service [DNW^{+}16]. Infrastructures [DDW^{+}19, TVG13, Zou14]. Infusion [HDL^{+}15]. Inhomogeneous [AAB16]. Initiated [dBK11]. Initiative [Par19b]. Injected [LYGX12, LLZ^{+}12b]. Injection [KTK12, PWT^{+}17, YYY^{+}14]. Innocuous [PFMR13]. Innovative [ASBL15]. Input [LY11, MBV13, WYLH18]. Input-Buffered [LY11]. Input-Queued [WYLH18]. InSAR [RZB^{+}18]. Insights [GGGA18]. Inspection [YP13]. Inspired [CLYR16]. Installment [WVM19]. Instance [TZC19, WNLL15, WLL15b]. Instant [HPP15]. Instruction [ZJL^{+}17b]. Instruction-Oriented [ZJL^{+}17b]. Instructions [LWZ^{+}16a, USP^{+}12]. Insulin [HDL^{+}15]. Integer [XTFC17]. Integrated [KPB19, LGD14, She10b, SB19, WWJ^{+}18, YWWR18, ZZH^{+}17]. Integrating [DD11, ME15b, WCZ^{+}19a, WYX^{+}19]. Integration [JMS^{+}18, LLFL15]. Integrative [ZSY14]. Integrators [Mur12]. Integrity [CLLS12, CL14, ZYL^{+}17, ZHAY12]. Intel [LSW17a, LLH^{+}15b]. Intelligence [LS17d]. Intelligent [JJG^{+}12, LZY^{+}19, WWX^{+}13]. Intensive [CAKRY16, GG11, HYZ15, HC14, JHWY19, JRO^{+}17, KCW11, LLZ^{+}18a, LS17c, LWZ^{+}16a, MZLT19, MBH^{+}10, NTWL11, SCH^{+}15, WCZ^{+}19b, ZLJ^{+}15a, ZJZ^{+}16, ZLK^{+}16]. Intentions [LPZ12]. Inter [ADZZM15, CJW16, CH13, DLXS19, KKW13, LGL^{+}18a, LLLH19, LAFA15, SSPG17, XLL^{+}18]. Inter-Atomic [LAFA15]. Inter-Connection [DLXS19]. Inter-Datacenter [LGL^{+}18a]. Inter-DC [XLL^{+}18]. Inter-Domain [ADZZM15, LLLH19]. Inter-Server [CJW16]. Inter-Thread [SSPG17]. Inter-WBAN [CH13]. Interaction [AAW^{+}17, LSKZ13, NSLV16, ZTH17]. Interactive [KLWK12, LJ15, LCY^{+}17, ZT14, ZTH17, ZT16]. Interactivity [TNZ^{+}12]. Interactivity-Constrained [TNZ^{+}12]. Interbatch [LG13]. Intercloud [DCA19]. Interconnect [KOPS10]. Interconnecting [Sib12, YQZC12]. Interconnection [APG12, ABF12, CMV^{+}10, CMB15, DY18, ESGG^{+}15, FPGAD10, FB10, LMLM13, XDMZ17, YKN^{+}19]. Interconnects [FKMC15, JWJS14, LY11]. Intercontact [BCP^{+}14, ZLF^{+}11]. Interdependence [HWNS15, YQZC12]. Interest [AKC^{+}15, ERSR13, MFO^{+}13, SLW15]. Interest-Clustered [SLW15]. Interest-Tagged [AKC^{+}15]. Interference [BTL^{+}19, BSL^{+}17, HC14, LWY^{+}13, Li14c, SSPG17, TCS11, WLH^{+}15, YQH^{+}15, ZCXF16]. Interference-Aware [HC14]. Interferences [HZT18]. Interlaced [ZD12]. Interlacing [ZPD11]. Interleaved [WLX13]. Interlocking [TWW^{+}15]. Intermediate [CZQ^{+}17, uRILP17, ZLN^{+}13]. Intermittent [AR10]. Intermittently [EHNS13b, HWC^{+}14, LHYW15, WYX13, YNW13]. Intermittently-Connected [LHYW15]. Internal [BCQ^{+}10, PYH19]. Internet [TW14, AJMW14, HKA12, IB14, LCG^{+}13, LLG^{+}13, NLY15, NN13, PKS14, Ren14, TDLR13, WXZ^{+}14, WSWY15, WX11, XLLZ11, YXWL16, YGL^{+}15, YZL^{+}15, YJC15, ZCJY14, ZX13]. Internet-Based [ZX13]. Internet-Scale [WSWY15]. Interplay [CM10]. Interpolation [MSW^{+}12]. Interrupt [CL16b, GDM^{+}13, HT16]. Intersection [QP16b, WZLC15]. Interval [AD19, XJL^{+}14]. Intra [HxjGG19, RSNV18, SJVR19]. Intra-Algorithm [SJVR19]. Intra-Node [RSNV18]. Intra-Task [HxjGG19]. Intrabatch [LG13]. Intradomain [BCF13]. Intrasession [KKW13]. Intrinsic [LLCH12]. Introduction [BKK11, CLL^{+}14, MBMC13, PKL^{+}12, RFZ11, Sto13c]. Intrusion [EK10, KKK11, MR16, SBC^{+}10, WFA13, ZKSY14]. Intrusion-Tolerant [MR16, SBC^{+}10, ZJL^{+}12]. Intrusive [TWL16, YZT^{+}17]. Invalidation/Self [RLSK17]. Inverse [DFGG13]. Inversion [RDG12]. Inverted [WJ12]. Inverting [CCT10]. Investigate [Bru14]. IoT [HCH^{+}19]. IP [LCG^{+}13, RHT13, TCS13, ZCLS14]. IP-Geolocation [LCG^{+}13]. IP-VPNs [RHT13]. IPv6 [WCD^{+}11]. IRM [She10b]. Irregular [CSV^{+}17, CLHW13, ME15a, MMSAZ11, Ozd19, TZT^{+}16, Trä19, SA11]. Irregularity [HHK10]. Irrevocability [QGZP17]. Irrevocable [KWG17]. IRRWBF [TBC12]. iSCSI [RLY^{+}15]. ISEE [LZY^{+}19]. iShuffle [GRCZ17]. Ising [OZMC^{+}16]. Island [CCKF15]. Island-Based [CCKF15]. Islands [PCL15]. Isoefficiency [DW10]. Isogeometric [SWT^{+}19]. Isolation [JEW^{+}18]. ISP [LLC10]. ISP-Friendly [LLC10]. ISPs [ARM16, Dan11, XZH14]. Issuance [LLD^{+}18]. Issue [Ano11d, Ano11c, Ano12c, BKK11, CLL^{+}14, MBMC13, PKL^{+}12]. Issues [Man16, TMJ14]. ITA [PFMR13]. Item [OUA11]. Items [ARM16]. Itemset [XZQZ17]. Iterated [LPP13]. Iteration [LWS^{+}12, YLL^{+}17]. Iteration- [YLL^{+}17]. Iteration-Level [LWS^{+}12]. Iterations [MGB18]. Iterative [AI15, Che18a, CCNMF18, HJ17, JMA^{+}18, MA13, RCK15, SOA15, WGG^{+}18, XYT^{+}15, YL10, YPL13, ZGGW13, ZGGW14]. ITM [SA11]. Iyengar [Kum14].

Jacobi [FB10, MA13]. Jammer [LLXC12]. Jammers [LLXC14]. Jamming [HLS^{+}15, LLXC12]. Jamming-Caused [LLXC12]. Java [BVEAGVA10]. Jitter [SKGC14]. Job [CZX^{+}19b, CVM^{+}15, FES^{+}17, JTS^{+}11, KJL^{+}16, LLY16, LZWY14, LGM^{+}17, LLpC15, LM16, MBV13, PS19, XDLZ19]. Job-Driven [LLY16]. Jobs [CZWJ18, JHWY19, LCG^{+}16, LMAS17, MNG^{+}15b, MV18, QP16a, SZR17, WCZ^{+}19b]. Join [Che11, HZB^{+}16]. Joins [HZB^{+}16, YNK^{+}17, ZZQ18]. Joint [BBCB15, BSD^{+}18, CWC11, CTP^{+}17, DOLG16, KKW13, LQK^{+}13, LLRP18, RPYO11, XHQ^{+}15, YQH^{+}15, YJCQ15]. Journal [Bad14, Par18]. JSensor [SJAdCL19]. JSON [KB17]. Julia [BFD19]. Jump [LLCL12]. Jump-Stay [LLCL12]. Junction [XP12]. June [Par19b].

KAD [CSM^{+}13]. KASR [MDZC14]. Kautz [GWL^{+}11]. Kepler [BBM16, BB15, BB16]. Kerberos [TW14]. Kernel [DCA^{+}16, GD16, LSW17a, MLK15, SFA^{+}17, YDC^{+}17, ZH14a, KJvR^{+}15]. Kernel-Based [DCA^{+}16]. Kernelet [ZH14a]. Kernels [ALI^{+}17, KTD12, LMVS11, LWZ^{+}16a]. Key [BKL11, CSW^{+}17, CCT^{+}14, GZZ^{+}13, GYW^{+}19, HSMY12, HCL^{+}14, JKT11, LLY^{+}14, LY16b, LLL^{+}14b, MZL^{+}19, MCJT19, RM11, TXL^{+}14, YLW13, ZQH13]. Key-Aggregate [CCT^{+}14]. Key-Policy [GZZ^{+}13, HSMY12]. Key-Value [CSW^{+}17, MCJT19]. KEYing [TW14]. Keys [OMMZ14, RM11, TW14]. Keyword [CWL^{+}14a, CZS^{+}16, MDZC14, RVCT15, SWC^{+}14, SYL^{+}16, WCRL12, XWSW16]. Keyword-Aware [MDZC14]. Keyword-Based [RVCT15]. Kinetics [AC19]. Knowledge [JLKG17, WZ14, XWH15a]. Known [ZJTZ14]. Kong [TTJX12]. Kutta [Mur12]. KV [YTW^{+}19].

L [ZJZ^{+}16]. Label [MMSAZ11]. Label-Based [MMSAZ11]. Lambda [HZT18]. Language [MGS12, MRH^{+}16]. Languages [KBS11]. LANs [FLH13, XHZ^{+}13]. Large [AHSK17, Agr14, AHS^{+}15, BGHG16, BCQ^{+}10, BXXC12, CJW^{+}15, CMVB17, CL16a, CC10, CYW^{+}18, CLB^{+}19, CYC^{+}16, CMK^{+}16, CPL^{+}18, DGI^{+}19, DGG^{+}19, GGY^{+}19, GGS10, GLM13, GTT^{+}17, GZW^{+}18, Guo14, HWJ18, HjZ^{+}14, HJF16, IvS10, JMZD12, JSK18, JKVA11, JGZZ14, JEW^{+}18, KHN16, KCW11, Ksh10, LZL10, LMD16, LLN^{+}19, Li10, LZY12, LHL^{+}13a, LCS14, LLY^{+}17, LLAL18, LZY^{+}19, LTC^{+}19, LLY^{+}15, LSL^{+}10, LLM^{+}14, LLL^{+}14a, LLH^{+}15a, LXZB15, LSCL16, LCD^{+}17, MWZ^{+}14, MCJT19, MCRC17, OKT^{+}16, QNLN11, QLNN13, RMG18, SK14, SZWX15, SHF^{+}17, SDL^{+}15, TNZ^{+}12, TVG13, TKC^{+}15, TZB^{+}14, Tsa13, TTJX12, Van14, WCLK12, WRWW13, WJTZ14, WV17, WVM19, WXTL13, WKC12, XHC16, XTFC17, XHL^{+}15, XHL^{+}11, YTMS16, YQH^{+}15, YC18, YPL13, YQLS14, YL16, ZSH^{+}11, ZLW^{+}14, ZLJ^{+}15b, ZHL^{+}15, ZJL^{+}17a, ZSW^{+}19, ZLX^{+}14, dSLMM11]. Large-Capacity [XHC16]. Large-Scale [AHSK17, BGHG16, BCQ^{+}10, CJW^{+}15, CL16a, CC10, CYW^{+}18, CLB^{+}19, CPL^{+}18, DGG^{+}19, GGY^{+}19, GLM13, GTT^{+}17, Guo14, HWJ18, HJF16, JMZD12, JGZZ14, KCW11, Ksh10, LZL10, LMD16, Li10, LZY12, LHL^{+}13a, LCS14, LLAL18, LZY^{+}19, LLM^{+}14, LLL^{+}14a, LLH^{+}15a, LSCL16, MWZ^{+}14, MCJT19, MCRC17, OKT^{+}16, QNLN11, RMG18, SK14, SZWX15, SHF^{+}17, SDL^{+}15, TNZ^{+}12, TVG13, TKC^{+}15, TZB^{+}14, Tsa13, TTJX12, Van14, WCLK12, WRWW13, WJTZ14, WV17, WVM19, WKC12, XTFC17, XHL^{+}15, XHL^{+}11, YQH^{+}15, YC18, YPL13, YQLS14, ZSH^{+}11, ZLW^{+}14, ZLJ^{+}15b, ZHL^{+}15, ZJL^{+}17a, ZSW^{+}19, ZLX^{+}14, dSLMM11, LLY^{+}15].

LargeScale [LAdS^{+}15]. LASEC [SCL^{+}15]. LASS [LWY^{+}15]. Last [AFMM17]. Late [XLL^{+}18]. Latency [AJM12, ACV17, ASSB18, BSD^{+}18, BSL^{+}17, CC15, FKMC15, HHWZ17, HWDP10, JLM^{+}12, JCW^{+}19, KGR16, LWY^{+}13, LDLL18, LV17, NTKK15, QPB^{+}17, RS10, SOA15, SAA17, TFKN17]. Latency-Energy [LWY^{+}13]. Lattice [CMB15, FC18, TS18]. Lattice-Based [FC18]. Law [BCP^{+}14, FW13]. Laws [WJTL13, ZMF10]. Layer [AKP14, AHS^{+}15, BZA10, CLM^{+}15, TWL^{+}15, TSN10, THL13, Ven14, WX15, ZCLS14]. Layered [XSZ^{+}10, CLB^{+}19]. Layout [HWS16a, HWS16b, WMZ^{+}15]. Layout-Aware [HWS16b]. Lazy [QGZP17, TGNA^{+}13, ZHX^{+}19, SLL16]. Lazy-Merge [SLL16]. LazyCtrl [ZWY^{+}17]. LBMP [XLLZ11]. LCMT [LKBK11]. LDPC [FSS11, LJ16, TBC12, ZL14]. Lead [LGOB17]. Leader [AR10, LV17, SOK^{+}19]. Leading [MSW^{+}12]. Leakage [NFFK14, ZTA^{+}15, ZLN^{+}13]. Leakage-Aware [ZTA^{+}15]. Leaky [LN17]. Learn [BWH^{+}19]. Learn-as-you-go [BWH^{+}19]. Learning [BS15, BRX13, GGGA18, HCZ12, HZW^{+}19, IZA18, IRPvdS12, JGJF18, KKE19, LHHR18, TFLL18, TAZ^{+}19, WQZ^{+}16, WZL^{+}19, YY14, ZJLG14, ZLYL19]. Learning-Based [HCZ12, IZA18, TFLL18, ZLYL19]. Lease [TWW^{+}15]. Least [YPL13]. LEISURE [CHLC15]. Length [CJ16, hKYY11, TFKN17]. Less [ARM16, TKR14]. Lessons [RSW^{+}17]. Level [AELGE16, AFMM17, BHS^{+}19, BBGD^{+}17, BMJ^{+}17, DN19, DMS^{+}12, DRVC17, DD17, EAMEG11, EN12, FPGAD10, FSSZ16, HA11, HHWZ17, HWL^{+}17a, HZT18, IBC^{+}11, JRV^{+}13, JN16, KKC18, KWG17, LWS^{+}12, MMdE19, PB19, SAA18, SAB^{+}18, STK^{+}19, SS18, ST18, WLT^{+}12, WZL^{+}16, YYK11a, YR14, ZQCZ16, ZLDC15, ZHW^{+}19]. Level-Playing [BHS^{+}19]. Levels [BBCTA18]. Leveraging [CCD^{+}15, HCL^{+}12, KI14, LS17b, NCM^{+}17, ZWL17]. LIBRA [CYX15]. Libraries [CGZQ13]. Library [PB19, STK^{+}19, TTG^{+}15a, Tic14]. Library-Independent [Tic14]. Lifetime [APPG16, DOLG16, EMTX15, GCL14, HYX11, LCL^{+}11, LCLD13, WWL11, WL15, ZWLL12]. Lifetimes [YL11a]. Lifting [vdLJR11]. Light [JRZ^{+}18, JGG^{+}11, ZLLZ13]. Light-Traffic [JGG^{+}11]. Lightly [Lee12]. Lightweight [CYX15, DCL^{+}10, KL16, SAB^{+}18, She14, TCM18, TXZ^{+}11, VMB17, WG13, ZWL^{+}16a, ZLW^{+}19, LKBK11]. Like [Guo17, PYH19, RTZ^{+}18, YLJ^{+}17]. Limit [YHL^{+}18]. Limitation [MPHR17, YLH^{+}16]. Limitations [AEM17]. Limited [APPG16, BS14, LYH^{+}15]. Line [RH16]. Linear [AHTD18, AF19, CL16a, FC10, LLCH12, LYL16, NVBH18, TFM^{+}16, WRWW13, WWL^{+}13, WXYX14, YKW^{+}18, YY10, ZTD19]. Link [DGF12, DLZ^{+}14, GHL^{+}13, Li14c, MLL14, MFO^{+}13, SDV18, XBL15, YL11a]. Link-Stability [DGF12]. Linked [ZD16a]. Links [SRG19, Wan12, YQZC12, ZDF^{+}15, ZHW^{+}19]. LINPACK [JNL^{+}15]. Liquid [Li14a]. List [Ano10, Ano11b, Ano12b, Ano15b, Ano19b, Ano19c, WS18, Ano13b, Ano14b, Ano17c]. List* [Ano17b]. List-Scheduling [WS18]. Lists [LTM11, ZD16a]. Little [BKL11]. Live [BWH^{+}19, LJL^{+}11, LLZ^{+}12a, LH15, LSCL16, SLL13a, TVRD17, ZML13]. Live-Time [ZML13]. Lived [TWZW11]. Load [BMJ^{+}17, CHLC15, CMG17, CL16b, DBA17, DY17, FGLP10, FSSZ16, GKL^{+}17, HJPL14, HLCH11, HCSC13, Jia16, KKK^{+}15, KTK11, LGOB17, LSW17a, LSW17c, MRM12, NOR16, PNAK11, RKGS16, Ren14, RRS12, SPS18, TWL16, WTXG19, WYC^{+}15, YLR12, ZLJ^{+}15b, ZWL^{+}15, ZYW^{+}16]. Load-Balanced [CHLC15, HJPL14]. Load-Balancing [KTK11]. Loadable [SFA^{+}17]. Loaded [Lee12]. Loads [HV11, JVW10]. LOBOT [ZS13]. Local [ASD^{+}18, GTM^{+}17, LPP13, LWY^{+}15, LS17a, LKT11, LCL^{+}15, MLML15, TLP16, WHN^{+}19, XLT^{+}14]. Local-Activity [LWY^{+}15]. Local-Global [XLT^{+}14]. Locality [AA17, HAU19, HXLF15, KAA16, LIWJ15, MZL^{+}19, MCMR12, SYL^{+}14, UDH^{+}17, WL12a, XTXH13, XALS17]. Locality-Aware [HXLF15, KAA16, MZL^{+}19, MCMR12]. Localization [CYL^{+}14, DNW^{+}16, KCYM10, KSP10, LMSRSR12, LZZP13, LLXC12, Liu14, LWJ^{+}15, NML^{+}14, SHM^{+}12, WXY^{+}13, WYX^{+}15, XSYY13, YL10, YCTC13, YLW^{+}14, ZS13, ZLY^{+}14, ZH11, ZCX^{+}14, WYLX13]. Localization-Oriented [CYL^{+}14]. Localized [LMSRSR13, Li14c, SAM14b, SCL^{+}15, TKS11, WLS^{+}11]. Localizing [GZWN14, LLXC14]. Locally [BV10, ZZF10, ZLL^{+}15]. Locally-Adjustable [ZZF10]. Located [LGJZ16]. Locating [MS12]. Location [CCT10, CZYL14, DT14, GCZ15, HX10, KCK14, LRW12, Li13, MS12, WG13, XTHD10, ZFT^{+}15, ZX13, LSL14b]. Location-Aware [CCT10]. Location-Based [DT14, HX10, XTHD10, LSL14b]. Location-Free [KCK14]. Locations [WLL^{+}13]. LocaWard [LSL14b]. Lock [AS16, CC13a, CWCS15, LZH^{+}16, LLZ^{+}18a, ME15b, ZD16a, ZCC^{+}17, SDG17]. Lock-Free [AS16, ME15b, ZD16a]. Lock-Intensive [LLZ^{+}18a]. Locking [KSW18, kL11a]. Locks [DLA^{+}18]. LockSim [CWCS15]. Locomotion [YSDQ11]. Log [TOA13]. Logarithm [XLLZ11]. Logarithm-Barrier-Based [XLLZ11]. Logging [CLLX18]. Logic [MT12]. Logoot [WUM10]. Logoot-Undo [WUM10]. Loneliness [SRB14]. Long [HSX^{+}12, LWZ^{+}16a, LSW17c, TNH^{+}18, TWZW11, WGCG18]. Long-Lived [TWZW11]. Long-Term [HSX^{+}12, TNH^{+}18, WGCG18]. Long-View [LSW17c]. Longest [CJ16, YXSS13]. Look [YNK^{+}17]. Lookup [BJ13, Hsi14]. Loop [LWS^{+}12, MG18, Nov15, RMG18, YYL^{+}17]. Loops [GYLW18, YLLW16]. Loose [UBC13]. Loosely [HKkY^{+}16, MVML11, SOK^{+}19, ZWL^{+}16b]. Loosely-Coupled [ZWL^{+}16b]. Loosely-Stabilizing [SOK^{+}19]. Loss [KXL^{+}14, SA11, Tak14, TL16]. Lossy [DC18, DTLC19, LG13, TDL^{+}19]. Lot [AOW^{+}12]. Low [BSD^{+}18, BSL^{+}17, CZZ^{+}16, FKMC15, HHWZ17, JCW^{+}19, KKW13, KCK14, KGR16, LXHS12, LDLL18, LCL^{+}16b, LV17, MS13a, SEAH16, SAB^{+}18, Sib12, TFKN17, XWH15b, ZS13, ZRQA14]. Low-Cost [LCL^{+}16b, ZS13]. Low-Degree [TFKN17]. Low-Diameter [Sib12]. Low-Duty-Cycle [XWH15b]. Low-Energy [SEAH16]. Low-Latency [BSL^{+}17, FKMC15, JCW^{+}19, KGR16, LV17, TFKN17]. Low-Level [SAB^{+}18]. Low-Overhead [ZRQA14]. Low-Power [LXHS12]. Low-Rate [KCK14]. Lower [AH10, Fre13, HCyW^{+}17, LC14, WYX13]. LPM [LS19]. LSM [MCJT19, YTW^{+}19]. LSM-Tree [MCJT19, YTW^{+}19]. LSTM [SytL19]. LU [CRWY15]. Luopan [WTXG19]. Lustre [uRILP17]. LVRM [SDV18]. LvtPPP [ZML13].

m [KMM12, ATZZ14, KMM12, SWRQ18]. M-Oscillating [SWRQ18]. M/G/1 [ATZZ14]. M/G/m/m [KMM12]. M2M [SJ14]. M2M-Based [SJ14]. MAC [MLC^{+}15, MY11, SCC11, WL14, WL15]. Machine [BM12, CRZH15, CHPY17, DSM19, GGGA18, GLBJ18, HCZ12, IZA18, JGJF18, KKE19, KKW18, LMM18, LW11, Li14a, LGJ^{+}18, LJL^{+}11, LV17, NMG15, NCB17, RG17, TAZ^{+}19, VMP17, WKK17, XWJX15, YWY^{+}17, ZLW^{+}14, ZCG^{+}17, ZWL^{+}18]. Machine-Based [LW11]. Machines [ASSB18, BWH^{+}19, BB13, BBL^{+}16, BRX13, CWS12, CSS^{+}13, CL16b, CHLY18, DSM14, sFC12, GCG^{+}18, HPP15, Ian14, IPQ19, LJL^{+}15, LLZ18b, PLG19, PBD^{+}13, TTH^{+}19, XSC13, YDC^{+}17]. MACS [KGR16]. Made [YY14]. Main [APPG16, AJK^{+}17, MV16a, MV16b]. Maintain [NN10]. Maintaining [HCC^{+}12, HBF12]. Maintenance [BM12, HCJ^{+}10, JWY^{+}18, She10b, SL13]. Maiter [ZGGW14]. Make [ZTZ^{+}18a]. Make-span [ZTZ^{+}18a]. Makespan [OPM^{+}15, TFM^{+}16]. Making [FWH^{+}18, KM19, LJ15, XWL^{+}19]. Malicious [GG13]. Malleable [CC13b, Che18a, MSSV18]. Malloc [LGJ^{+}17]. Malware [PLZW14]. Mammoth [SCH^{+}15]. Manage [DN19]. Manageability [Gua14]. Managed [LMR10, MCJT19]. Management [ASG^{+}14, ASYK^{+}19, BCTB13, CC10, CSM^{+}13, CDS15, CCLW15, CZD^{+}19, CCCB14, CLJ11, CTP^{+}17, DK17, DRSL15, DSJ16, ESGQ^{+}13, FLLS17, FXL17, FGEL14, GGY^{+}19, GPF12, GGF^{+}14, GRJZ17, HLZY15, HAZ17, HZJ^{+}11, IZA18, IvS10, KK10, KZW17, KMMR13, hKYY11, KKC18, KPB19, KL16, LMD16, LZY12, Li13, LdSS^{+}13, LODB17, LCSC12, LWW^{+}13, LJL^{+}15, LSL^{+}18, LYPL19, LLL^{+}14b, LVD11, MA14, MBO15, NFD10, NSH15, NSY^{+}16, PR19, PD14, PVQ15, PCP14, Ren14, SDV18, SML13, SSW^{+}17, SFA^{+}17, TXL^{+}14, TGNA^{+}13, TGAG13, TCDMRP17, WW11, WL13, WCZ^{+}19a, WMLJ17, XXLZ16, XX16, XCZ^{+}15, XLLZ11, XL13, XAYM14, XFL15, YGL^{+}15, YQH16, YBY^{+}18, ZTA^{+}15, ZX13, ZQH13, ZFF16]. Managing [BB13, DCC^{+}19, FHH^{+}15, HZT18, LGL^{+}18b, LSL^{+}17, LCZ^{+}19, MVL15, Mit17, MPHR17, SLG^{+}18, TLH^{+}14]. MANET [QTC^{+}14]. MANETs [TYG^{+}14, WL15, YW10, ZYZC12]. Manual [NSLV16]. Many [AFA12, ABE^{+}11, AAA19, AA17, AK18, ASD^{+}18, CCC^{+}16, DMCN12, ELX^{+}11, HAY^{+}18, IOY^{+}11, KAA16, KCH19, ME15a, PGGS19, RRM^{+}15, RFZ11, RAG10, TCM18, YLJ^{+}17, YCMX17, YYK^{+}11b, ZJL^{+}17b, KLL^{+}17]. Many-Core [AFA12, AAA19, AA17, ASD^{+}18, CCC^{+}16, DMCN12, KAA16, ME15a, PGGS19, RRM^{+}15, RAG10, TCM18, YLJ^{+}17, YCMX17, ZJL^{+}17b, AK18, KLL^{+}17]. Many-Particle [HAY^{+}18]. Many-Task [ABE^{+}11, RFZ11, YYK^{+}11b]. Many-Tasks [IOY^{+}11]. Manycore [ALS^{+}18, CSV^{+}17]. Manycores [HPP15]. Map [GYLW18, KSP10, NVBH18, RSSC15]. Mapping [DDP^{+}19, DCA^{+}16, Goh14, GETFL14, GHZZ16, GYLW18, HZW^{+}14, LPP13, LCG^{+}13, LC15, LWSM19, LGX^{+}11, MG18, MA13, TZT^{+}16, TDLR13, VNA^{+}16, WDL^{+}17, YYL^{+}17, ZXG^{+}19, Zou14]. MapReduce [CPGT14, CYX15, CRG^{+}17, CZX^{+}19b, CCNMF18, DLZH16, EADT19, FHLG11, FWZ^{+}16, LLY16, LMAS17, LLpC15, LLH^{+}15b, MNG^{+}15b, MDZC14, PSL15, SMS^{+}13, SCH^{+}15, WZHZ16, uRILP17, XQL^{+}14, XGL^{+}16, XLY^{+}17, ZYLC14, ZJKQ16, ZZQ18]. Maps [DW10, ZMTL15]. Market [CLL11, ZL11, ZYZ^{+}14]. Markets [CLZ^{+}18, DM11, LYY16, LYZL18, MV18, Ren14, ZCJY14]. Markov [XHX^{+}13]. Markovian [BZBP10, CMPS11, PH12, Seh15]. Mars [FHLG11]. MART [TFPK13]. Mashup [DWT^{+}16, SytL19]. Masking [GTM^{+}17, IB14]. Massive [BM12, EJRB13, FHH^{+}15, KJN15, LXHL11, LQW^{+}18, MWZ^{+}14, SM16, TZT^{+}16, WMZ^{+}15, ZCX10]. Massively [BKH18, CCM^{+}17, CPL^{+}18, FSS11, GE12, KAG17, LMFS11, NIP11, NGJ^{+}19, RRM^{+}15, XLSR13]. Master [PF12]. Master/Worker [PF12]. MATCH [PCFP16]. Matching [CYC^{+}15, CJBW16, JWY^{+}18, KKK11, LLLC17, LS19, MC14, MIH17, NCKL14, QCZ^{+}15, TLSL15, TVCM12, WPKL13, YP13, ZS17]. Matchings [ABP17]. Matchmaking [LMZG15]. Matrices [CP17a, HAZ^{+}18, YZSC14]. Matrix [AAA19, AA17, BBD^{+}19, CLY^{+}19, GTT^{+}17, GWC14, KGK^{+}13, KAA16, KBS11, LT16, LLAL18, LKD10, MLC^{+}19, RCK15, RDG12, SOA15, TLP12, TTG^{+}15b, XHG15, YMG15, YR14, Zha12, ZML^{+}17, ZHZL17]. Matrix-Transpose [KAA16]. Matrix-Vector [GWC14, KGK^{+}13, RCK15, YR14, Zha12]. Max [GCL14, MYPL18, TCS11, WPKL13]. Max-Min [GCL14, MYPL18, TCS11]. Maximal [ACS13, LCL^{+}11]. Maximization [CHLZ13, LZL^{+}18, LRJX13, LLL^{+}14a, MLXG19, MLT^{+}19, SWRQ18, SZWX15, VWDM14]. Maximize [BBP17, LSWR16]. Maximized [CLJ11]. Maximizing [CCFS11, Che16, EMTX15, KHK15, LKBK11, LWS^{+}12, PDH10, WWL11, ZWLL12]. Maximum [ABP17, CHCC14, HH11, WMWW19]. MaxMin [CTA14]. MBR [LC14]. MCL [DY18]. MDP [MGR12]. MDP-Based [MGR12]. MDS [SSF16a]. Means [KPA13, XQL^{+}14]. Measure [Wan19]. Measurement [CB16, CHLC15, LRW12, LHD^{+}14, LHL^{+}13b, LLG^{+}13, YGS^{+}19]. Measurements [LSLD17]. Measuring [LS17a, LSCL16, WX11]. Mechanism [BHS^{+}19, CRD11, FPF13, HML^{+}14, JRZ^{+}18, KALK^{+}18, LSKZ13, LLZ18b, LYZL18, MG14, MNG15a, NLC12, RMM16, SWL17, WGCG18, WXTL13, YXWL16, YLL^{+}17, YZS13, ZSY14, ZYZ^{+}14, ZLL^{+}15]. Mechanisms [BFFG11, CHHK19, DD11, HLeS^{+}15, JWNS19, NMG15]. Media [ASBL15, CDBQ12, Sto11a, yWeH11, ZCG^{+}17]. Medical [BKF^{+}16, LTW^{+}14, WYZ^{+}19]. Medium [ATA18]. Medium-Grain [ATA18]. Medusa [ZH14b]. Meet [ASYK^{+}19]. Meeting [CB14, LLL18, PP12]. Mega [GKL^{+}17]. Megabase [dOSdM13]. Megh [BWH^{+}19]. Melia [WZHZ16]. MeLoDy [WGCG18]. Memories [ASD^{+}18, Di 17, MV16b, WLX13]. Memory [APPG16, ASG^{+}14, AJK^{+}17, ALI^{+}17, AA12, BBK17, CB16, CSV^{+}17, CLC^{+}12, CP17b, CLO^{+}18, CWLS19, CHHK19, CCHH19, CCC^{+}16, CD13, CPH^{+}18, CRRR15, DD11, DGI^{+}19, DCA^{+}16, EADT19, FFMR10, FJV^{+}18, GLGLBM13, GCG^{+}18, GBP17, GGGA18, HTA10, HWSX17, HGC12, HPP15, JSMK11, JSLD19, Jun17, KHK15, KKK11, KL16, KWG17, LW11, LAK11, LJL^{+}15, LN17, LYPL19, LLK^{+}14, MZL^{+}19, MC17, MV16a, MV16b, MJK14, NTDZ19, QGZP17, RSNV18, SAA18, SG16a, SHY14, SKGC14, SCH^{+}15, SLS^{+}16, TZY^{+}18, TGNA^{+}13, TGAG13, TFLL18, TVCM12, VBC19, VMB17, WSC^{+}14, WWJ^{+}18, WCYL19, WLX^{+}15, XML^{+}18, YHL^{+}18, YYL^{+}17, YTL^{+}19, YR14, ZML^{+}17, ZLT^{+}18, ZH18, dCAB19]. Memory-Aware [WSC^{+}14]. Memory-Efficient [KKK11]. Memory-Intensive [SCH^{+}15]. Memoryless [SZ12]. Merge [MG14, YPL13, SLL16]. Merge-and-Split [MG14]. Merging [SLL16, WZQY14]. Mesh [AJMW14, ABF12, CLHW13, CHD^{+}15, DDP^{+}19, KCK14, LGG^{+}14, NTDZ19, RYLZ10, WXL10, ZWD^{+}10, ZX13, dSLMM11]. Mesh-Based [dSLMM11]. Mesoscopic [VT19]. Message [CGZQ13, DMR16, DGFRR18, EHNS13a, Ksh10, MB13, RWLL14, TZB^{+}14, WDOX15]. Message-Efficient [Ksh10]. Message-Passing [DGFRR18]. Messages [HD15]. Messaging [JWE15]. Meta [CZRB18]. Meta-Platform [CZRB18]. Metacomputing [PF12]. Metadata [CLL^{+}19, DCC^{+}19, GGY^{+}19, HZJ^{+}11, HJZ^{+}12, LLS^{+}18, STMM17, XHL^{+}11, XAYM14]. Metaheuristics [SJVR15, SJVR17]. Metascheduling [MV18]. Metering [LA12, ZHQ12]. Method [AI15, CZS^{+}16, Che18a, CLL^{+}19, CYC^{+}16, CCV19, EHWX10, FLH13, FXL17, FKMC15, HJ17, KE16, LHHR18, MWZX14, MDZC14, NTKK15, SOA15, SL13, SS18, SZWX15, SP12, TLJ^{+}14, TTH^{+}19, TZY^{+}18, TS18, Van14, XJ14, XFL^{+}19, YL16, AAB^{+}17]. Methodologies [EAMEG11]. Methodology [FPRG16, HJF16, KM18, KOKA11, LS19, MGR12, PWRL18, SL11, WTTH17]. Methods [Jun17, LM17, kL11a]. Metric [TLP15, ZH11, ZBK^{+}15]. Metrics [LCZZ13, PGP^{+}17, WTL^{+}14]. Metropolitan [RYLZ10]. Micro [Tak14, WUH^{+}17, YSG^{+}14]. Micro-Clouds [WUH^{+}17]. Micro-Environment [YSG^{+}14]. Microarchitecture [CA13, HZT18]. Microarchitecture-Level [HZT18]. Microarray [Yan14]. Microbenchmarking [MC17]. Microblog [WSWY15]. Microblog-Based [WSWY15]. Microgrids [YJC15]. Microservice [BWB^{+}19]. Microservice-Based [BWB^{+}19]. Microtask [TNLM17]. Microtask-Based [TNLM17]. Middleware [SJ14, ZGL10]. Migratable [MNZ^{+}15]. Migration [APCH^{+}11, BWH^{+}19, CDBQ12, DBA17, LJL^{+}11, LH15, MWZX14, PS19, TVRD17, XWJX15, YWW^{+}15, ZCG^{+}17, ZLL^{+}17b]. MIKEY [TW14]. Mile [ZHL^{+}15]. MIMO [FQWL12, GHL^{+}13, WCF10, XHQ^{+}15]. Min [GCL14, MYPL18, TCS11, WPKL13]. MIN-MAX [WPKL13]. Minimal [LKM10, LLL18, MMYES^{+}18, NTA^{+}16, YC14]. Minimal-Path [MMYES^{+}18].

Minimization

[DW13b, HJS^{+}11, HGL^{+}16, SSW^{+}17, WSC^{+}14, WGZ16, YJC15, YJCQ15]. Minimize [ACV17, LLD^{+}18, ZTZ^{+}18a]. Minimizing [CJW16, CCD^{+}15, LCZZ13, LW15, LGJ^{+}18, LWZ^{+}15, LLXC14, RKRK17, ZCLS14]. Minimum [HWJ18, HWDP10, JLM^{+}12, LCLD13, MB13, MM10, PKCB11, YYL^{+}13, ZTH17, ZGKB16]. Minimum-Cost [HWJ18, LCLD13]. Minimum-Delay [PKCB11]. Mining [ACC^{+}17, DLC^{+}16, HLY^{+}14, LTGI16, LZC^{+}12, OUA11, RGK15, SZC^{+}17, SCJ^{+}17, SZ11, XZQZ17, Yan14]. MinMax [HWSX17]. MinMax-Memory [HWSX17]. MinMin [CTA14]. MINs [ESGQ^{+}13]. Mirroring [YJC^{+}16]. Misbehavior [ZDG^{+}14]. Mismatch [HLY10]. Misplaced [BXXC12]. Misplaced-Tag [BXXC12]. Miss [PD14]. Mission [JRP^{+}10]. Mitigating [ASSB18, PB12, TCYF16, XLY^{+}17, ZSW^{+}15]. Mitigation [CYX15, SHF^{+}17]. Mixed [BJC^{+}18, CSW^{+}12, GS11b, HTZY17, JZZ^{+}15, MZK19, SAEH19, XTFC17]. Mixed-Criticality [BJC^{+}18, HTZY17, SAEH19]. Mixed-Precision [GS11b]. Mixing [ZFF16]. ML [DWH^{+}18]. MLC [AJK^{+}17]. MM* [YLM^{+}15]. MMOG [LS17b]. MOANA [CAC^{+}19]. Mobi [LZP^{+}13]. Mobi-Sync [LZP^{+}13]. MobiFuzzyTrust [HML^{+}14]. Mobile [ALLR14, AE12, AKT^{+}15, BN12, BZA10, BS12, CYZ^{+}13, CW15, Che15, CH13, CBK^{+}10, DHTZ15, EMTX15, EHNS13b, ERSR13, FCD^{+}13, GXW^{+}17, GJLZ13, HML^{+}14, HWC^{+}14, Iye14, IIKO13, JJ11, KK10, KXC11, KKC18, KPG^{+}12, LJG12, LLL^{+}13, LCS14, LWY^{+}15, LLS14, LWZ^{+}15, LNA^{+}13, LDNT13, LLG^{+}13, LZP^{+}13, LHYW15, LCY^{+}17, LLS13, LWZ12, MKOK14, MS13b, MPS15, MSB11, PJC^{+}13, PAB13, QZZ^{+}16, RBM15, RM11, RM12, RKZC14, SLY^{+}14, SLG10, She14, SZWX15, SJ14, TZB^{+}14, TTJX12, VLRP15, VLP16, WPT10, WUH^{+}17, WWW^{+}18, WDOX15, yWeH11, WYX^{+}15, WKW19, WZL^{+}19, XWH15a, XWY^{+}10, XTHD10, YSDQ11, YQLS14, ZYZ^{+}14, ZYW^{+}14a, ZMTL15, ZLL^{+}17b, ZLW^{+}19, ZWZ^{+}15, dLMPG19]. Mobile-Application [VLP16]. Mobile-Healthcare [LLS13]. Mobility [HWC^{+}14, LMSRSR12, LCS14, LWZ12, TTJX12, WCD^{+}11, WYX^{+}15, YLSQ13, ZFT^{+}15]. Mobility-Assisted [HWC^{+}14]. Mobility-Resilient [LCS14]. MoD [Hu14]. Modal [DWLY15]. Model [Agr14, BCTB13, Bru14, BRX13, Cha11, CH14, CRS^{+}17, CLY^{+}19, CCNMF18, DCC^{+}19, DKS^{+}15, DBA17, DGI^{+}19, DRVC17, FC18, HKA12, HZT18, JGJF18, JHW^{+}15, KMM13a, LSW17a, LM17, LL12, LLJ^{+}13, LTW^{+}14, Li14c, LKT11, MSSV18, NSLV16, OZMC^{+}16, RS10, RGLDM17, Sam14a, SJVR17, SPH^{+}18, SA11, SOK^{+}19, TCZL11, WMW11, WHF^{+}19, WDL^{+}17, WYZ^{+}19, XZSG12, XHX^{+}13, YZSC14, YLM^{+}15]. Model-Based [LSW17a, LM17, RGLDM17]. Model-Free [BRX13]. Model-Predictive [BCTB13]. Modeland [YLM^{+}15]. Modeling [AJMW14, BWB^{+}19, BLLP15, CAC^{+}19, CTLH14, CZZ^{+}16, CRWY15, CMG^{+}14, CWCS15, DGI^{+}19, GTM^{+}17, GLGLBM13, GIRT19, GWC14, HBS^{+}16, KJL^{+}16, KKC17, LKM10, Li10, LQK^{+}13, LYL15, LSJ^{+}19, LNMMA15, MNE14, MV16d, MMBdS14, PDFJ13, PBD^{+}13, PGGS19, VMN^{+}16, WWL^{+}13, WZZ^{+}13, WMLJ12, WSSZ13, WYCZ14, XHX^{+}13, YYY^{+}14, YZFZ10, ZRTL15, ZMF10]. Models [AAA19, AF19, CWZ^{+}15, CLZ^{+}18, DSM14, DMCN12, HAY^{+}18, HKE^{+}16, JKVA11, KKE19, LHHR18, LdSS^{+}13, OOA^{+}14, SRB14, Seh15, TBÁ^{+}19, WJTL13, YCWL14, ZFT^{+}15]. Moderately [LCG^{+}13]. Modern [CMB18, JZW^{+}17, PB19]. Modifiers [WFK^{+}12]. MODLoc [GZWN14]. moDNN [CCHH19]. Modular [HA13, IGEN11, JPG14, Lou14, SEAH16]. Module [ZS17]. Modules [SFA^{+}17]. Modulo [LGX^{+}11, VGMA10]. Moldable [BHKS^{+}17]. Molecular [KAG17, LAFA15]. Mona [LZWY13]. Monitor [CHLC15]. Monitoring [CADK19, DLL^{+}11, DL17, GAB18, GJZZ12, HGY^{+}14, HCS12, HCZ12, HSX^{+}12, KJvR^{+}15, LAV^{+}10, LRJX13, LZC^{+}12, LCS^{+}15, MKVL12, PM13, SHX^{+}10, TVG13, TWL16, YRLY16, YSDQ11, YQLS14, YLT15, YC12]. Monotonic [CYX^{+}14]. Monte [NSLV16, OZMC^{+}16]. Montgomery [IGEN11]. Morton [LZH18]. Mosaicking [MWZ^{+}14]. Mostly [CZL^{+}16]. Motion [CEK16]. MotionCast [WBPF11]. Movement [HLK^{+}19, LKE16, LWZ^{+}15, SAM14b, WMT^{+}11, YWZ17]. Movement-Assisted [SAM14b, WMT^{+}11]. Movements [WWCB14]. Mover [HZB^{+}16]. Moving [DWH^{+}18, GRJZ17]. mPath [XLSR13]. MPCA [LHHR18]. MPI [APJ^{+}16, CGZQ13, CC17, DLM^{+}17, GHZ15, HCA16, JDB^{+}14, JNL^{+}15, LAdS^{+}15, LZH18, kL11a, Pan14, SPH^{+}18, TGT10, VPS17]. MPI-ACC [APJ^{+}16]. MPI-OpenCL [JNL^{+}15]. MPSoC [ASD^{+}18, HYX11, WLC^{+}17]. MPSoCs [JIP14]. MRCP [LMAS17]. MRCP-RM [LMAS17]. MrPhi [LLH^{+}15b]. MSGD [LLAL18]. mSNP [CPL^{+}18]. MTAF [RVCT15]. MTC [MVML11]. mTreebone [WXL10]. Much [XZSG12]. Multi [ATZZ14, ALZ17, Agr14, AIAD^{+}18, AFMM17, BHKS^{+}17, CGS^{+}15, CWL^{+}14a, Cha14, CCKF15, CZWZ14, CRC^{+}17, CLL^{+}17, DN19, DWLY15, DMCN12, DD17, DY17, FWJ18, GFLL15, GSH^{+}19, GZY^{+}15, GYLW18, GLBJ18, GCL14, HGS^{+}19, HJY16, HYZ15, HWL^{+}17a, HxjGG19, HCH^{+}19, Hsi14, HT16, IPQ19, JY15, JNL^{+}15, KALK^{+}18, KJN15, KPKH16, LJ16, LKBK11, Li14b, LC15, LXXH16, LZWZ19, LZWY13, LH15, LSW16, LH16, LCL^{+}16b, LCZ^{+}19, LSN19, MMdE19, MYPL18, PCL15, PB19, PJAGW14, QF14, RGRM14, RM17, RBH^{+}14, SV19, SEAH16, SHY14, SWRQ18, SAF16, SL14, SS18, ST18, SVK^{+}19, SWC^{+}14, TTH^{+}19, TWSW17, TNH^{+}18, VNA^{+}16, VLP16, WLL15a, WLL15b, WFZ^{+}17, WPT17, WVM19, WDL^{+}17, WM18, XWSW16, XWH15a, YJ14, YC14, YYL^{+}17, YN17, ZD16a, ZJS^{+}17, ZLDC15, ZZLL16, ZWL17, KLL^{+}17]. Multi-Accelerator [CGS^{+}15]. Multi-Application [GFLL15]. Multi-Authority [LXXH16, YJ14]. Multi-Bank [YYL^{+}17]. Multi-Channel [GSH^{+}19, GCL14, TTH^{+}19]. Multi-Chip [HYZ15]. Multi-Commodity [MYPL18]. Multi-Copy [XWH15a]. Multi-Core [AFMM17, CCKF15, CRC^{+}17, CLL^{+}17, GZY^{+}15, HT16, IPQ19, KPKH16, LJ16, PCL15, PJAGW14, QF14, RGRM14, SEAH16, SWRQ18, SAF16, SL14, SVK^{+}19, WFZ^{+}17, YN17, ZJS^{+}17, ZWL17, KLL^{+}17]. Multi-Cores [BHKS^{+}17, HGS^{+}19, HxjGG19, PB19]. Multi-CPU [VNA^{+}16]. Multi-Demand [CZWZ14]. Multi-Dimensional [KJN15, ZD16a]. Multi-Dominating [YC14]. Multi-Failure [LSN19]. Multi-FPGA [SHY14]. Multi-GPU [JNL^{+}15, RBH^{+}14]. Multi-Hop [HCH^{+}19]. Multi-Index [Hsi14]. Multi-Installment [WVM19]. Multi-Instance [WLL15b]. Multi-Keyword [CWL^{+}14a, SWC^{+}14, XWSW16]. Multi-Level [DN19, DD17, MMdE19, SS18, ST18, ZLDC15]. Multi-Map [GYLW18]. Multi-Modal [DWLY15]. Multi-Objective [GLBJ18, VLP16, WDL^{+}17, ZZLL16]. Multi-Owner [LZWY13]. Multi-Path [Cha14]. Multi-Phase [LZWZ19]. Multi-Port [Agr14, GZY^{+}15].

Multi-Priority [ATZZ14].

Multi-Processor [SV19, TWSW17]. Multi-Queue [HT16]. Multi-Resource [KALK^{+}18, TNH^{+}18, WLL15a]. Multi-Ring [LCL^{+}16b]. Multi-Sensor [HJY16]. Multi-Server [FWJ18, LC15, WPT17]. Multi-Service [AIAD^{+}18]. Multi-SIMD [WM18]. Multi-Task [Li14b]. Multi-Tenancy [DY17]. Multi-Tenant [LSW16, LH16, LCZ^{+}19, RM17]. Multi-Threaded [JY15, SV19]. Multi-Threading [LKBK11]. Multi-Tier [ALZ17, LH15]. Multi-Tiered [HWL^{+}17a]. Multi-Word [IPQ19]. Multiagent [JZW13, Jia14b]. Multiattribute [DW13a, XH10]. Multicast [APMG12, ADZZM15, CLA^{+}19, DY16, DY18, FW13, GLL11, JZWN15, LXHS12, LC12b, LG13, LGYV14, LY14, QTC^{+}14, SHG11, TSN10, TCS13, Ven14, WXL10, XJY^{+}10, YLSQ13, YY10, ZWD^{+}10, ZCX15, dBK11]. Multicasting [Fre13]. Multicent [CSY15]. Multichannel [FW13, JCLJ12, LYW^{+}12, LCZZ13, ZWD^{+}10]. Multiclock [GG10]. Multicloud [FPF13, MVML11, WZ14, ZHAY12]. Multicluster [SME10, WMLJ12]. Multicopy [LW12]. Multicore [AM19, ACV17, CGH13, CLT13, CVM^{+}15, FSS11, HLZY15, HTZY17, HWG^{+}19, HZJ16, Ian14, IZA18, JHR^{+}14, KM18, KLFD13, LM17, Lee12, LRYJ17, LMVS11, LKD10, MSW^{+}12, Man16, MRGR12, NHN17, NHN18, PD14, PVS18, RCV^{+}13, RDG12, SAEH19, SJVR15, THE^{+}15, TMJ14, WTD17, WLT^{+}12, WYY^{+}12, WW12, WDC12, YKW^{+}18, YTMS16, YP13, Zha12, ZBS15, ZWL^{+}16b, ZCXF16, ZML13, ZYX^{+}10]. Multicore/Multiprocessor [WDC12]. Multicore/Multithreaded [RCV^{+}13]. Multicores [BCTB13, LWZ^{+}16b, MJK14, PPS^{+}17, aaGZ19]. Multidestination [APMG12]. Multidimensional [CHW^{+}17, JCW^{+}12, TXZ^{+}11]. Multifunctional [CSY15]. Multigrid [GS11b]. Multihomed [LX10]. Multihoming [YZL^{+}15]. Multihop [CWJS11, GHL^{+}13, JLM^{+}12, JJG^{+}12, Li14c, MS13a, MLS15, MLT^{+}13, SKP12, TCS11, WLS^{+}11, XLM^{+}11b, ZMA12]. Multilevel [ADLM19, ERG^{+}17, GETFL14, MMBdS14, WHC^{+}14]. Multimedia [CSZ^{+}12, GSH^{+}19, LWCG10, LWZ^{+}16b, PAB13, TW14]. Multinode [CSV^{+}17]. Multiobjective [SJVR15, SJVR19]. Multiorganization [DPRT11]. Multipacket [CWJS11, RVW^{+}15]. Multiparty [ZLCZ14]. Multipath [BZBP10, CFLL19, PNAK11, TCS11, WYW13, WYC^{+}15, XBL15, XLLZ11, XLM^{+}12b, XLM12a, XLSR13]. Multiphase [SPH^{+}18]. Multiplayer [GE12, NIP11]. Multiple [BBCTA18, CGKP11, DED^{+}19, GTM^{+}17, GZWN14, GHW^{+}16, HV11, IBC^{+}11, JZZ^{+}15, KZW^{+}12, KCYM10, LMZG15, LL17, LLLC17, LSW17b, MBF19, NML^{+}14, PCL15, RQZ^{+}16, TTG^{+}15a, WL12a, WWL^{+}13, YCTC13, YXSS13, YLH^{+}16, YLL^{+}17, ZLY^{+}14, ZCX15, ZWQ^{+}15]. Multiple-Level [IBC^{+}11]. Multiplexed [LGL^{+}18b]. Multiplication [AAA19, AA17, CLY^{+}19, GTT^{+}17, GWC14, IGEN11, KGK^{+}13, KAA16, MLC^{+}19, TTG^{+}15b, YMG15, YR14, Zha12, ZML^{+}17]. Multipliers [ARM15]. Multiplies [SOA15]. Multiply [RCK15]. Multipole [AAB^{+}17]. Multiprocessor [BJC^{+}18, HZW^{+}14, LAK11, Lee17, LW15, LKT11, LHJ12, LGX^{+}11, LWW^{+}13, USP^{+}12, WSC^{+}14, WDC12, ZLL17c, ZLLD18, JIP14]. Multiprocessors [AJM12, CYX^{+}14, CIP^{+}17, CP17c, CCK12, DKM^{+}15, EHM^{+}17, HGC12, JTS^{+}11, LAMJ12, PL16, RAG10, SBMA15, SCH11, TL16, WMW11, WLX^{+}15]. Multiradio [FW13, LCZZ13]. Multirate [XJY^{+}10].

Multiregion [CBK^{+}10]. Multirobot [PM13]. Multiserver [CHLZ13]. Multiservice [TKP12]. Multisocket [CGH13]. Multisource [HWI12, JVW10]. Multispanning [MMSAZ11]. Multistep [LYY16]. Multistream [IZA18]. Multitasking [LHR^{+}15]. Multithreaded [CC13a, CJW^{+}15, EJRB13, HH11, RCV^{+}13, VTSM12, ZJS12, ZBS15]. Multithreading [ZL10]. Multitier [LZ12, RX11, SZL^{+}12]. Multivariate [TJH^{+}14]. Multiversion [PRR^{+}16]. Multiview [JN16]. Must [Hen14]. Mutual [BH13, CGKP11, WZLC15].

N [SEAH16]. N-Modular [SEAH16]. Named [LAT^{+}15, XWJX15]. Namespace [HjZ^{+}14]. Nanophotonic [MJK14]. Narrowband [SG16b]. Nash [RMG14, WS14]. Natural [YTMS16]. Navigation [CCS^{+}12, TLJ^{+}14, WLL^{+}13]. NDFT [XAK17]. Near [FJV^{+}18, HLY10, LYZ^{+}16, TP13]. Near-Memory [FJV^{+}18]. Near-Optimal [HLY10, LYZ^{+}16, TP13]. Nearest [JY15]. Nearest-Neighbor [JY15]. Nearly [ZD16b]. Nebula [JRO^{+}17]. Necessary [Dua95, NX95, VS11a, VS11b]. Nefeli [TRD13]. Negotiation [WMC^{+}19]. Negotiations [SPB^{+}10]. Neighbor [JY15, KKY^{+}14, LLXC12, RVW^{+}15, WML15, WMGA15, YL11a, YLM^{+}15]. Neighborhood [YLC^{+}19]. Neighborhood-Guided [YLC^{+}19]. Nessie [CSW^{+}17]. Nested [XHX^{+}13, YLLW16]. Nets [MSB11, ZJLS12]. Network [AMN^{+}16, ATACA18, AJMW14, ADMX^{+}12, AF18, BAMJ12, Bis18, BFFG11, BHEP14, CL13, CHM^{+}13, CHLC15, CYL^{+}14, CHD^{+}15, CSSL15, CP15, CWL16, CCCY16, CCH^{+}17, Che18b, CCHH19, CW19, CFLL19, CDPM18, CE10, CSR^{+}17, CTP^{+}17, CTG^{+}19, DKM^{+}15, DY18, DRK11, EMTX15, EN12, EKNS17, EMW16, FPGAD10, GLZ11, GHZ15, GGGA18, GDM^{+}13, GGF^{+}14, HCY^{+}12, HH11, HSX^{+}12, HWNS15, JGHD10, KHK15, KLWK12, KKE19, KN16, KKW13, KKW15, KCW11, KAV^{+}17, KL11b, KSP10, LMR10, LLLG13, LAMJ12, LMLM13, LG13, LGYV14, LCLL15, LYH^{+}15, LY16a, LWLZ17, LGL^{+}18a, LDLL18, LHXP18, LWZ^{+}15, LY16b, LLK13, LTM11, LWW^{+}13, LHL^{+}13b, LLZ14, LWJ^{+}15, LCL^{+}15, LCZ^{+}19, LGW^{+}17, MLML15, MLXG19, MRM12, MKSN18, MCRC17, NL11, Oru17, PPR10, PHXL19, PL16, PCP14, QZG^{+}16, QFZZ15, QP16b, RCV^{+}13, RAS17, RGK15, RKZC14, RCC^{+}14, RKRK17, She14, SLC15, SSM^{+}18, SL11, Sib12, SLM^{+}10]. Network [SHX^{+}10, SZWX15, SOTN12, SCHT16, TYG^{+}14, TLP16, TWSW17, TZC19, TP18, Tou15b, TP13, WPT10, WXL10, WCD^{+}11, WLT^{+}12, WWL^{+}13, WJTL13, WLL^{+}13, WL14, WL15, WWW^{+}18, WCZ^{+}19b, WVM19, WZZ^{+}13, WXYX14, WMS^{+}19, XHC16, XYT^{+}15, XH10, XHX^{+}13, XSZ13, XAK17, YW10, YY10, YLJ^{+}17, YZS13, YZH^{+}19, YQ16, YWJJ11, YY14, ZTG^{+}18, ZJL^{+}12, ZGXJ14, ZWFX17, ZL11, ZMLT13, ZXW^{+}13, ZSY14, ZDM^{+}19, ZCJ19, ZWY^{+}17, ZYL^{+}16]. Network-Aware [CTP^{+}17]. Network-Based [WVM19]. Network-Coded [She14]. Network-Induced [GGGA18]. Network-Limited [LYH^{+}15]. Network-on-Chip [AMN^{+}16, ATACA18, Bis18, CHM^{+}13, CCH^{+}17, Che18b, CDPM18, DKM^{+}15, LDLL18, LCL^{+}15, PL16, TLP16, TWSW17, YLJ^{+}17]. Network-Wide [CHLC15]. Networked [DLC^{+}16, HOZ12, LPP13, LSKZ13, LT10, RY14, WV17, ZHX^{+}19]. Networking [CYZ^{+}13, HGL^{+}16, Iye14, TL14, XWJX15, XGZW14]. Networks [APG12, AO12, ALLR14, ANN^{+}13, AAB16, ADZZM15, ADMX^{+}12, ABF12, ACNP11, AE12, AKT^{+}15, Amm12, AKP14, AA14, BBCB15, BKY15, BCSKN12, BCP^{+}14, BV10, BTG^{+}18, BS15, BSF16, BZA10, BZBP10, BS12, BS14, CJH^{+}14, CCFS11, CMV^{+}10, CMVB17, CMB18, CLM^{+}15, CWL14b, CHCC14, CPM^{+}10, Cha14, CWC11, CTX^{+}11, CQZ^{+}12, CW15, CJL^{+}12, CHTW12, CLLS12, Che14, CYL^{+}14, CYC^{+}15, CHD^{+}15, CCT16, CSY16, CJW16, CMG17, CLB^{+}19, CFLL19, CH13, CNC^{+}14, CFJ15, CJW^{+}19, CC15, CCCB14, rCHG10, CLSZ12, CLJ11, CIH13, CLHK11, CWJS11, CWC^{+}13, CMC^{+}15, DWX14, DLXS19, DMR16, DGF12, DWLY15, DDP^{+}19, DRSL15, DWW^{+}11, DLL^{+}11, DLZ^{+}14, DOLG16, DWY^{+}13, DY16, DWF12, Dua95, EHNS13a, EHNS13b, ESGG^{+}15, FCD^{+}13, sFC12, FB10, FQWL12, FW13, GS11a, GBD^{+}13, GFLL15, GTS^{+}15, GLL15, GLL11, GLM13, GJLZ12, GJLZ13, GCN^{+}14, GWL^{+}11, GJZZ12, GHL^{+}13, GCL14]. Networks [Guo14, GLJ^{+}15, GCZ15, GXZ^{+}15, GLC^{+}15, HGY^{+}14, HWJ18, HML^{+}14, HCS12, HL12a, HCL^{+}12, HCC^{+}12, HJPL14, HCG^{+}15, HA10, HRGE17, HLY10, HS12, HLWV14, HCJ^{+}10, HWDP10, HPH^{+}12, HWX12, HWI12, HWC^{+}14, HH12, HHK10, JWA10, JRAS17, JJ11, JGG^{+}11, JCLJ12, JVW10, JZY^{+}15, JLW^{+}10, JJW11, JCW^{+}12, JZW13, JZH^{+}14, JZW^{+}14, Jia14b, JHW^{+}15, JZWN15, JLM^{+}12, JWNS19, JKP12, JJG^{+}12, KK10, KKW13, KCK14, KKY^{+}14, KP12, KXL^{+}14, KZLL14, KL11b, LLGP13, LRT19, LMR10, LLH14, LKE16, LMPR12, LKM10, LAV^{+}10, LXHL11, LVA^{+}11, LC12a, LXHS12, LJG12, LYW^{+}12, LL12, LRW12, Li13, LWY^{+}13, LQK^{+}13, LLL^{+}13, LMSRSR13, LG13, LCZZ13, LCGC14, LHD^{+}14, LCL^{+}14, LCS14, LWZ14, Li14c, Li14b, LHF^{+}15, LWY^{+}15, LLG15a, LL11, LRJX13, LLS14, LWZ^{+}15, LWCG10, LCW11, LHJ12, LLK13, LZXW15, LXZH16, LX10, LZN10, LC11, LZNX11, LM12, LLCL12, LW12, LNA^{+}13, LDNT13, LJB^{+}13, LCLD13, LZP^{+}13, LLZ14, LZCK14, LLXC14, LLL^{+}14a]. Networks [LZK^{+}15, LLH^{+}15a, LHYW15, LCL^{+}16a, LSC16, LWL^{+}17, LLZ^{+}12b, LLG14, LSW^{+}15, LTMD11, LWZ12, LX12, LWG^{+}12, LGG^{+}14, LYZ^{+}16, MM12, MLL14, MLC^{+}15, MMYES^{+}18, MS12, MS13a, MLS15, MM15, MTX^{+}11, MLT^{+}13, MKOK14, MMSS15, MSS17, MS13b, Mis14, MM10, MPS15, MY11, MSB11, MYPL18, MMSAZ11, MGR12, NVS16, NN10, NFFK14, NTKK15, NTK^{+}15, NL11, PB12, PFMR13, PKCB11, PKG14, PLZW14, PNAK11, PSMD18, PCP14, QZZ^{+}16, RBM15, RKGS16, RGRM14, RCFW10, RVCT15, RM11, RM12, RYLZ10, RZH^{+}11, RHDL11, RZW^{+}13, RWLL14, RQZ^{+}16, RS12, RGBC11, RXD12, RVW^{+}15, SHG11, SHG13, Seh15, SX10, SLL13b, She14, SLLL14, SCC11, SP15, SKL^{+}15, SPS18, SJAdCL19, SILJ11, SKP12, SM16, SHM^{+}12, SKA15, SZZF10, SJ14, TKS11, TXWL11, TYLG13, TLRW15, Tan12, TKC^{+}15, TMMN15, TZB^{+}14, TLSL15, TLL^{+}16, TCS11, TJLL12, TWZW11, Tou15a, THL13, TFKN17]. Networks [TKP12, TTJX12, UBC13, VM12, VWDM14, VS11a, VS11b, VS14, WLS^{+}11, WMT^{+}11, WWL11, WMHX12, WFK^{+}12, WJTL12, WYW13, WWH13, WWLX13, WFA13, WYX13, WJTL13, WJTZ14, WTL^{+}14, Wan14, WJWX14, WL14, WSL^{+}15, WWZ^{+}16, WHB16, WQZ^{+}16, WYX^{+}19, WTXG19, WYL19, WRB11, WG13, WXTL13, WDOX15, WUM10, WJX^{+}14, WZQ10, WMLJ12, WCF13, WWCB14, WYC^{+}15, XAY^{+}14, XL16, XSZ^{+}10, XWH15a, XWH15b, XHHC13, XJ14, XBL15, XHG15, XLL^{+}18, XWY^{+}10, XJL^{+}14, XJY^{+}10, XLM^{+}11b, XLM^{+}12b, XLM12a, XHQ^{+}15, YOWA14, YY10, YGL13, YNW13, YCTC13, YLW^{+}14, YKN^{+}19, YL15, YK14, YRL11, YWJJ11, YCW12, YLT15, YWZ17, ZWD^{+}10, ZJLS12, ZGH14, ZGXJ14, ZS10, ZZF10, ZPD11, ZD12, ZZR12, ZMA12, ZMLT13, ZWWF15, ZDF^{+}15, ZRTL15, ZHL^{+}15, ZZCD10, ZWLL12, ZX13, ZQH13, ZW14, ZMTL15, ZCLS14, ZYT^{+}15, ZY14, ZWZ^{+}15, ZFG^{+}10, ZHCW12, ZDG^{+}14, ZLYL19, ZASA10, Dua93]. Networks-in-Package [Seh15]. Networks-on-Chip [AAB16, ADMX^{+}12, HRGE17, RKGS16, SHG11, SHG13, SKL^{+}15]. Neumann [EJGYAM14]. Neural [BS15, CHM^{+}13, CLB^{+}19, CCHH19, CSR^{+}17, EN12, MKSN18, YTL^{+}19, YZH^{+}19, YY14]. Neuron [CRS^{+}17]. Never [ACE^{+}19]. Newsletter [Ano12j]. Next [FBCB18, HJZ^{+}12, LPSS19, LPMB13, PT15, VPS17]. Next-Generation [FBCB18, HJZ^{+}12, VPS17]. NIC [WDC12]. NN [XHHC13, THE^{+}15, ZZQ18]. NN-DP [ZZQ18]. No [TL16]. NOC [AHS^{+}15, AJM12, AVA^{+}17, BICK^{+}15, CLHW13, FFC17, HLZY15, LWSM19, WDL^{+}17]. NoC-Based [HLZY15, WDL^{+}17]. NoCs [CCLW15, GAB18, LCL^{+}16b, MWJ^{+}14, MS15, ZFF16]. Node [CRS^{+}17, EMTX15, FWH^{+}18, Lai12, LY14, NTK^{+}15, PDH10, RSNV18, SHM^{+}12, TWZW11, TP14, WWL11, WYX13, XBL15, ZML^{+}17]. Node-Disjoint [Lai12, XBL15]. Node-Weighted [LY14]. Nodes [DGI^{+}19, GG13, JNL^{+}15, YSDQ11, ZQSY13]. Noise [LWW^{+}13]. Non [APPG16, BJC^{+}18, Cha14, DGI^{+}19, FWJ18, GBFS16, Jun17, KKC17, LLG15b, LCL^{+}15, MVL15, MV16b, NVBH18, NTDZ19, PH12, RMM16, SJVR17, SL14, TFKN17, YZT^{+}17, YL16, ZH18]. Non-Asymptotic [FWJ18]. Non-Cooperative [Cha14]. Non-Disruptive [GBFS16]. Non-Generational [SJVR17]. Non-Intrusive [YZT^{+}17]. Non-Linear [NVBH18]. Non-Local [LCL^{+}15]. Non-Markovian [PH12]. Non-Parametric [YL16]. Non-Preemption [SL14].

Non-Preemptive [BJC^{+}18]. Non-Random [TFKN17]. Non-Repudiation [LLG15b]. Non-Saturation [RMM16]. Non-Stationary [KKC17]. Non-Uniform [DGI^{+}19]. Non-Volatile [APPG16, Jun17, MVL15, MV16b, NTDZ19, ZH18]. Nonblocking [DY18, HH11]. Noncontiguous [JDB^{+}14]. Noncooperative [RS12, WZQ10]. Nondeterministic [LW12]. Nonlinear [CEK16]. Nonnegative [AHJ^{+}11]. Nonstationary [CLHW13]. Nonuniform [XAK17]. Nonuniformity [ACNP11]. Normalization [JWE15]. NoSQL [CPH^{+}18]. Note [Ano11e, Bad15, Bad17a, Bad17b, Par19b, Par19c, Par19a, Sto10f, Sto10a, Sto10b, Sto10c, Sto10d, Sto10e, Sto11b, Sto11c, Sto12a, Sto12b, Sto13c, Sto13a, Sto13b, Bad16]. Nothing [TVRD17]. Novel [EHNS13a, HMP^{+}19, KWZ^{+}12, LMLM13, LLG15b, LLG15a, LLAL18, LC14, LN17, MWJ^{+}14, PYHY16, RYLZ10, SLL16, Sam14a, SOA15, WWLX13, YLSQ13, ZWFX17, Zha12, ZX13]. NTC [WFZ^{+}17]. NUCA [AHS^{+}15]. Nuclear [AAW^{+}17]. Null [GYX^{+}10]. NUMA [CAD^{+}18, PLG19, RLY^{+}15, ZCC^{+}17]. NUMA-Aware [CAD^{+}18, RLY^{+}15, ZCC^{+}17]. Number [CCFS11, KHN16, US16]. Numbers [ACS13, FHH^{+}15]. Numerically [CCV19]. NVIDIA [KAGD16, XFL^{+}19]. NVM [CP17c]. NVRAM [ZLL^{+}17b, ZLW^{+}19]. NVRAM-Aware [ZLL^{+}17b].

O [AZW^{+}19, BHEP14, CAC^{+}19, CRZH15, DIAR16, GDM^{+}13, HWS16b, HWL^{+}17a, LLJ^{+}13, LMFS11, NCM^{+}17, NLC12, PYHY16, SHA19, SSLF17, WWH^{+}17, ZWFX17, ZLJ^{+}15a, ZWJ^{+}19]. O-Efficient [WXLY16]. Oasis [LHG^{+}17]. Obfuscation [RBM15]. Object [ET10, HJY16, Liu14, WSSZ13].

Object-Tracking [HJY16]. Objective [GLBJ18, VLP16, WDL^{+}17, ZZLL16]. Objectives [CSY15]. Objects [DGFRR18, DWH^{+}18, GZWN14, MNZ^{+}15, NML^{+}14, PYH19, ZLGN13]. Oblivious [IIKO13, LZH18, SDL^{+}15]. Observation [ZWQ^{+}15]. Observations [HCL^{+}14]. Occupancy [AOW^{+}12, HLY^{+}14]. Ocean [ELX^{+}11]. Octree [DWH^{+}18]. Octrees [NTDZ19]. ODE [OOA^{+}14]. ODE-Based [OOA^{+}14]. OFDM [NHN18, NHN17]. OFDMA [TYLG13]. Off [CDS15, CIP^{+}17, OMMZ14, TFKN17, WBPF11, SPS18]. Off-Axis [OMMZ14]. Off-Chip [CIP^{+}17]. Offline [HWJ18, LTW^{+}14]. Offloading [CL17, Che15, CL15, CL16b, DHTZ15, GXW^{+}17, HCH^{+}19, HZW^{+}19, LCY^{+}17, MBV11, ZLYL19]. Offs [GAKR11, ZYZC12]. OLAP [DRRCB18]. Omni [KJvR^{+}15]. Omni-Kernel [KJvR^{+}15]. Omnidirectional [ZYW^{+}14b]. Omnisc’IO [DIAR16]. On-Chip [HD15, JKP12, LKBK11, LWW^{+}13, MVL15, Sib12, Tou15b, Tou15a, VNA^{+}16, WWJ^{+}18, WYZ^{+}19, Oru17]. On-Demand [CE17, KCK14, LTC16, LSB^{+}18, LFLW10, ZLZ^{+}14]. On-Off [CDS15]. One [ZLCZ14]. One-View [ZLCZ14]. Online [BSL^{+}17, CL17, CLT13, CJW16, CCK12, DNW^{+}16, DRVC17, GLRT18, GAB18, GKKW16, GE12, HWJ18, HHWZ17, HCZ12, IdM12, IRPvdS12, KTK11, LGD14, LZY^{+}18, LSL^{+}10, LSC16, NIP11, NVS16, QP16b, RG17, RX11, SEA18, SZL^{+}12, SLLL14, SLC15, SWL17, SZ12, TDL^{+}19, TLSL15, TLL^{+}16, THT^{+}15, Tse13, TAZ^{+}19, WMW11, WJWX14, WLL15b, WJX^{+}14, XHHC13, XDLZ19, YGL13, ZHL^{+}15, ZWLW16, ZWL^{+}16a, ZCJ19, ZLZ^{+}16, ZHZL17]. Only [YLW13, ZQSY13]. onto [Goh14, MA13]. ONU [NTKK15]. OP2 [RMB^{+}16]. Open [Ano12i, CCCY16, XWL^{+}19, LHL^{+}13a]. Open-P2SP [LHL^{+}13a]. OpenCL [JNL^{+}15, LAFA15, WTTH17, WZHZ16].

OpenCL-Based [WTTH17, WZHZ16]. OpenMP [AAB^{+}17, AELGE16, LdSB19, TCM18, VPS17, YKW^{+}18]. OpenStack [RTZ^{+}18]. Opera [VMN^{+}16]. Operating [BBCTA18, KJvR^{+}15, LZ11, TLH^{+}14]. Operation [KWG17, SOTN12, TWT16, YOK^{+}17, ZCJY14]. Operation-Level [KWG17]. Operational [ARM16, SLG10]. Operations [CCFS11, GHZ15, KWG17, PKG14, TLP12, WX15]. Operator [LMZG15, NCGP19, TZC19]. Operator-Aware [LMZG15]. Operators [LABQ18]. Opportunistic [BCP^{+}14, CNC^{+}14, GXW^{+}17, KKW15, LGYV14, LW12, LLS13, MLC^{+}15, MTX^{+}11, MPS15, PKCB11, RBM15, XSZ13, XDLZ19, ZMTL15, ZWZ^{+}15]. Opportunities [YC18]. Opportunity [LYW^{+}12, LZN10, WTL^{+}14]. Opportunity-Based [LZN10]. OPS [RMG18]. Optical [LY11, WYX^{+}19, ZGY15]. Optimal [AWZ15, AD19, BMB^{+}10, CLM^{+}15, CHLZ13, CPGT14, CCHH19, CLJ11, DRVC17, EKNS17, GPF12, HH13, HWZE10, HLY10, HWL^{+}17b, JWK^{+}16, JTS^{+}11, JSC^{+}17, KN16, Lai12, LMR10, LKE16, LWX^{+}11, LYW^{+}12, LLFL15, LYZ^{+}16, NN13, QZG^{+}16, RCFW10, Ren14, TWT16, TYG^{+}14, Trä19, TCT16, TP13, VS15, WWL^{+}13, WLL15b, WHGS17, WL12b, XJL^{+}14, XSL^{+}16, YQZC12, YYK11a, YDC^{+}17, ZCX10, Zhu14, ZD16b, Zom14]. Optimally [LWS^{+}12]. Optimising [JHR15]. Optimistic [HPPR17, JZW^{+}14, PVQ15, PGGS19]. Optimization [ALI^{+}17, CJ10, CWC11, CCT16, CWJS11, DW13a, DC18, DOLG16, FC11, FHH^{+}15, GLBJ18, GCL14, GWC14, HLS^{+}15, HPH^{+}12, IB14, IdM12, KOPS10, KM18, KGK^{+}13, KTK12, LSW17a, LM17, LW11, LMPR12, LQK^{+}13, LYL15, LHXP18, LGZ^{+}19, LCW11, LDYZ15, LS19, LdSB19, MZLT19, MSW^{+}12, Man18, MP16, MGR12, Nov15, PDFJ13, PT15, PJAGW14, RCK15, SSLF17, TWSW17, TFLL18, WTD17, WTTH17, WWZ^{+}16, WIZ^{+}17, WWH^{+}17, XXWY10, XLL11, XLH^{+}15, XL17, YZL^{+}15, YYK^{+}11b, YWC11, YWZ17, ZXL^{+}17, ZHCL17, KLL^{+}17]. Optimizations [AK18, FGJ^{+}15, GIX^{+}12, NSLV16, dOSdM13]. Optimize [NCM^{+}17, aaGZ19]. Optimized [GLC^{+}15, HX10, LLH^{+}15b, SAF16, TTG^{+}15a, TTG^{+}15b, TS16, VMP17, WJ12, WJB14, ZH18]. Optimizing [Bar10, CRS^{+}17, CJBW16, FSSZ16, GBP17, GRB^{+}19, GZY^{+}15, HS12, KAV^{+}17, KBHS14, Li14c, LTBN^{+}12, MT12, PR19, SSF16b, TDL^{+}19, WHGS17, WWL^{+}17, XFL^{+}19, ZSC^{+}17]. Optimum [CRRR15]. OptiTuner [HJS^{+}11]. Orchestration [DL17]. Order [CA13, LZH18, MTDD17, PYH19, SLY^{+}14, TYG^{+}14, USP^{+}12, dLMPG19]. Order-Optimal [TYG^{+}14]. Ordered [HJ17, MMSAZ11]. Organisation [ZSY14]. Organization [AJM12, HJZ^{+}12, LCYW16, MG14, SLSG19]. Organized [KN16, LGOB17]. Organizing [DW13a]. Oriented [ATACA18, CYL^{+}14, CDR15, DBA17, DY17, GLZ11, Kao15, LZL^{+}18, LLS14, LZNX11, MM12, TCS13, WLC^{+}17, WDL^{+}17, ZJL^{+}17b]. OrthoNoC [ATACA18]. OS-Level [KKC18]. Oscillating [SWRQ18]. Oscillation [XHX^{+}13]. OTN [WYL19]. OTrack [SLY^{+}14]. Out-of-Core [AZW^{+}19]. Out-of-Order [CA13, MTDD17, USP^{+}12]. Outages [YJC15]. Outlier [ABLS16]. Output [CCLW11]. Outsourced [CT12, CLH^{+}14, FRS^{+}16, WCRL12]. Outsourcing [CL16a, HN11, LHL^{+}14, Lou14, WRWW13, XAG17, YJR15]. Overcommitted [CWS12]. Overflow [CW19]. Overhead [CWC11, HTZY17, MS13a, SOA15, WSC^{+}14, XVC17, ZRQA14, ZLT^{+}18]. Overheads [LLLG13]. Overhearing [WCF13]. Overhearing-Aided [WCF13].

Overlaid [FC11]. Overlapping [JAJ^{+}19]. Overlay [BZBP10, GJC^{+}13, HS12, KP12, LMR10, LMPR12, LC10, LZY12, MM12, MCMR12, SLL13a, WXL10]. Overlays [MFO^{+}13, TSN10]. Overload [YLH^{+}16]. Overloaded [BB13]. Own [CZRB18]. Owner [LZWY13, SYL^{+}16]. Owner-Enforced [SYL^{+}16].

P [XAK17]. P-NDFT [XAK17]. P2P [BJ13, CSZ^{+}12, CJL^{+}12, CSSL15, FC11, HBF12, Hu14, JRV^{+}13, LXHL11, LZY12, LWCG10, LLZ^{+}12a, NN10, NL11, PFMR13, ST10, SGGB14, She10a, She10b, SL13, SLGW14, SLLL14, SLW15, SLC15, SLLZ16, SPB^{+}10, WMGA15, WUM10, WL12b, WML14, XZH14, YCWL14, ZL11, ZZCD10, ZLCZ14, dSLMM11]. P2P-Assisted [SLLL14, SLLZ16]. P2P-Based [CSZ^{+}12, SLGW14]. P2P-VoD [WL12b]. P2SP [LHL^{+}13a]. P3S [PWRL18]. Package [Has16, Seh15]. Packet [Bis18, HT16, JPG14, LQK^{+}13, LHM12, LW14, LG10, LY11, LCL^{+}15, PT11, QP16c, SML13, WFK^{+}12, WL13, WLH^{+}15, WW12, YP13, ZGY15]. Packet-Carried [LCL^{+}15]. Packet/Circuit [Bis18]. Packet/Circuit-Switched [Bis18]. PacketCloud [CCCY16]. Packing [LTC16, RG17]. Page [CHLY18, ERRG18, ZHW^{+}19]. Page-Aware [CHLY18]. Page-Level [ZHW^{+}19]. PageRank [CATC11]. Pageview [WX11]. Pairs [MBH^{+}10]. Pairwise [GDRTS16, RM11, SZA11]. PAN [RSSC15]. Pancyclicity [CH15, LL12]. Panoramic [RSSC15]. Papers [Ano11d, Ano11c, Ano12c]. Paradigm [HJZ^{+}12, LLD^{+}18, WMS^{+}19]. PARAFAC [CHW^{+}17]. Parallel [AZW^{+}19, AFT^{+}16, Ano11d, Ano11c, Ano12c, Ano15a, Ano16, Ano17a, Ano18, Ano19a, ABP17, ARM15, Bar10, BBD^{+}19, BKH18, BBGD^{+}17, BS15, BBM16, BSM^{+}11, CMVB17, CAC^{+}19, CLL^{+}14, CATC11, CCM^{+}17, CWZ^{+}15, CBF^{+}17, CHW^{+}17, CLT^{+}17, CLB^{+}19, CZWJ18, CSR^{+}17, CLL^{+}17, CCD^{+}15, CPL^{+}18, DFGG13, DWW^{+}15, DDP^{+}19, DMCN12, DLA^{+}18, DBG^{+}14, EALM17, FGJ^{+}15, sFC12, FGEL14, GDRTS16, GCG^{+}18, GBP17, GRB^{+}19, GLM13, GTT^{+}17, HH13, HGS^{+}19, Has16, HWS16a, HWS16b, HWL^{+}17a, HAD12, HWF18, HW13, JFP^{+}17, JMZD12, JSK18, JSMK11, JY15, JN16, JWY^{+}18, JHYK11, Jun17, Kao15, KM10, KAA16, KKK11, KKK^{+}15, KPA13, KBHS14, KAG17, LM17, LLN^{+}19, LMLM13, LZWY14, LLW^{+}15, LSWR16, LYL16, LY16b, LMFS11, LLLC17, LWZ^{+}13, LPMB13, LKD10, LHCM^{+}17, MSW^{+}12, MC14, MTDD17, MT12, MSS17, MNE14, MJM16, MCRC17, NTDZ19, NGJ^{+}19, OKT^{+}16, Ozd19]. Parallel [OUA11, Par19b, PF12, PVS18, PJAGW14, QP16a, RRM^{+}15, RMG14, RGLDM17, RLVTMG^{+}16, SFL^{+}14, SLL16, SJVR15, SJVR19, SKGC14, SG16b, SOA15, SAF16, SZR17, SSM^{+}18, SJAdCL19, SWT^{+}19, SA11, SM16, SKA15, SP12, TYS^{+}12, TBC12, TVCM12, Van14, VS15, WLT^{+}12, WMZ^{+}15, WZL^{+}16, WYLH18, WCZ^{+}19b, WYL19, WK11, WRL15, WKC12, XL10, XH10, XZX^{+}17, XAK17, XVC17, YTMS16, YXWW14, YCPC15, YTL^{+}19, YR14, ZTD19, ZSH^{+}11, ZLJ^{+}15a, Zha12, ZJKQ16, ZJL^{+}17b, ZJS^{+}17, ZWL17, ZWT^{+}19, ZASA10, KLL^{+}17]. Parallelism [BSD^{+}18, BBP17, HYZ15, JN16, KJN15, LLCH12, LKBK11, LWS^{+}12, QJ16, SCH11, TCM18, WTD17, WLT^{+}12, WCYL19, YYK11a, YLLW16, ZLJL17]. Parallelization [AAH15, CM10, KAC^{+}15, YXSS13, ZR18]. Parallelize [SJVR17]. Parallelized [PPR10]. Parallelizing [JSLD19, MIH17]. Parameter [ABE^{+}11, KM18, LCY^{+}17, ZJLG14]. Parameters [CJBW16, sFC12].

Parametric [YL16]. Parana [YTL^{+}19]. Pareto [TWT16, Zom14]. Pareto-Optimal [Zom14]. Parity [CLLX18, MWZX14, SSF16b, WHH^{+}13, WMJ^{+}19, YJC^{+}16]. Parity-Based [MWZX14, WHH^{+}13, YJC^{+}16]. Parity-Switched [SSF16b]. Parking [AOW^{+}12]. Parsing [EHI11]. Part [HKE^{+}16, ZLJ^{+}15b]. Partial [ANE12, GJC^{+}13, HLY^{+}14, KLFD13, LVA^{+}11, PRR^{+}16, SSF16b, ZLJL17, Zou14]. Partially [HK18, YZHZ17]. PARTIC [WWCZ11]. Participatory [CZZ^{+}16, XYT^{+}15]. Particle [BGHG16, HAY^{+}18, MSW^{+}12, MLK15, NSLV16, RBH^{+}14, WTD17]. Particle-to-Grid [MSW^{+}12]. Parties [KDCR19]. Partition [GETFL14]. Partitionable [DWF12, WV17]. Partitioned [PHGR17, PG16]. Partitioning [AAA19, ATA18, BBD^{+}19, BB17, CATC11, GKT^{+}17, HWJ18, HWG^{+}19, HLYJ19, Ian14, Kao15, KKK^{+}15, LPP13, LZL^{+}18, kL11a, MSS17, PPR10, SVL^{+}16, TPRH16, WKK11, WKW19, XZQZ17, YLL^{+}17, ZLJ^{+}15b]. Party [CRZH15]. PASQUAL [LPMB13]. Passing [DGFRR18]. Passive [KCW11, LZZP13, WRB11, WZFG13, YNW13, ZYW^{+}14b, ZCX^{+}14]. Password [HCL^{+}14, YLW13]. Password-Authenticated [HCL^{+}14]. Password-Only [YLW13]. Past [HK18]. Path [CJ16, CCM^{+}17, Cha14, CCH^{+}17, EKNS17, FMY^{+}18, FFC17, LHD^{+}14, LZB14, MMYES^{+}18, YXLJ16]. Path-Diversity-Aware [CCH^{+}17]. PathGraph [YXLJ16]. Paths [ANE12, Lai12, LHJ12, MLT^{+}13]. Patient [HDL^{+}15, ZLDC15]. Patron [HCyW^{+}17]. Pattern [ACC^{+}17, CC17, HDL^{+}15, HLY^{+}14, HPP15, KKK11, LS19, NCKL14, NFFK14, SZ11, TWW^{+}15, YP13]. Pattern-Aware [HPP15]. Pattern-Based [NFFK14]. Patterns [ALI^{+}17, BVFGSFAF17, CSV^{+}17, HAD12, JSMK11, LTGI16, LZC^{+}12, NCM^{+}17, RGK15, SZC^{+}17, SMS^{+}13, ZT13]. Pay [TNH^{+}18]. Pay-as-you [TNH^{+}18]. Payment [DW13b, MS13a]. Payments [CT12]. PC [KOKA11]. PCFTL [WX15]. PCID [PSMD18]. PCM [AJK^{+}17, LZW^{+}17]. PCM-Based [LZW^{+}17]. PDFS [YZHZ17]. PEACE [RYLZ10]. Peak [ASYK^{+}19]. Pec [WLL^{+}19]. Peer [BFPB10, BMB^{+}10, BS14, CTLH14, CE10, CMG^{+}14, Dan11, GS11a, GG13, GE12, GIP^{+}13, GWL^{+}11, HN10, HLY10, HLCH11, HS12, JCWB10, KLWK12, KXC11, KI14, LWX^{+}11, LFLW10, LSL^{+}10, LHW11, RS10, RCFW10, SLL13a, SLL13b, TJLL12, WL12a, XSZ^{+}10, XZSG12, YTZ^{+}11, YZSC14, ZXL^{+}17]. Peer-Assisted [CMG^{+}14, LFLW10, LSL^{+}10]. Peer-to-Peer [BFPB10, BMB^{+}10, BS14, CTLH14, CE10, Dan11, GS11a, GG13, GE12, GIP^{+}13, GWL^{+}11, HN10, HLY10, HLCH11, HS12, JCWB10, KLWK12, KXC11, KI14, LWX^{+}11, LHW11, RS10, RCFW10, SLL13a, SLL13b, TJLL12, WL12a, XSZ^{+}10, XZSG12, YTZ^{+}11, YZSC14, ZXL^{+}17]. Peers [CNMA11]. Penalty [WHH^{+}13]. Penalty-Aware [WHH^{+}13]. Pending [LLZ^{+}18a]. Penguin [GZW^{+}18]. PEPS [HAY^{+}18]. Per-File [WMJ^{+}19]. Per-Flow [WL13, YGS^{+}19]. Perceived [CLZ^{+}18, WX11]. PerfCompass [DNW^{+}16]. Perfect [LC10, LLLC17, NTA^{+}16]. Performance [APG12, AMN^{+}16, AHTD18, ALS^{+}18, ATZZ14, AJMW14, AAB16, AAW^{+}17, ABBCT16, BKK11, BJ13, BWB^{+}19, BCTB13, BF17, BBL^{+}16, BSP10, Bru14, CTA14, CTLH14, CB13, CRWY15, CSY15, Che16, CFLL18, CLY^{+}19, CRG^{+}17, CE10, CM10, CCNMF18, CCW^{+}12, dCCF15, DBAT11, DDW^{+}19, DCC^{+}19, DKS^{+}15, DNW^{+}16, DWT^{+}16, DD17, DY16, EHWX10, EAMEG11, EALM17, ESGQ^{+}13, FES^{+}17, FGEL14, FHH^{+}15, GFS^{+}10, GLGLBM13, GHZ15, GDM^{+}13, Gua14, GWC14, GRCZ17, HAZ^{+}18, Has16, HWS16a, HWS16b, HJS^{+}11, HWX12, HBS^{+}16, ITL17, IOY^{+}11, ITW^{+}14, IG11, JHR15, JSMK11, JIP14, JRV^{+}13, Jia14a, JPG14, Kao15, KM19, KJL^{+}16, KMM12, KMM13a, KMM13b, KCW11, KPB19, KL16, LAdS^{+}15, LGJZ16, LM17, LGD14, Li10, LYL15, LSLD17, LSJ^{+}19, LZH^{+}16, LGJ^{+}17, LCL^{+}16b, LCY^{+}17, LLK^{+}14, LNMMA15, LWZ^{+}16b, MC14, MC10, MWZ^{+}13, MSB11, MA13, MJK14, MRGR12, NSLV16, NLC12, NTWL11]. Performance [OPJ^{+}19, PSL^{+}11, PH11, PT15, PR19, PH12, QZG^{+}16, QP16c, RX11, RPYO11, RS12, SG16b, SWRQ18, SWT^{+}17, SAF16, SX10, SBC^{+}19, SLS^{+}16, TXWL11, TWW^{+}18, TGT10, TAZ^{+}19, WHH^{+}13, WW11, WKK11, WKL^{+}16, WKW16, WHGS17, WV17, WWJ^{+}18, WRL15, WHYZ10, WCF13, WYCZ14, WWL^{+}17, WMLJ17, WHZS19, WYZ^{+}19, XX16, XZJ^{+}19, YTL^{+}10, YLL^{+}17, YKN^{+}19, YL16, YQ16, YWJJ11, YWZ17, ZJS^{+}17, ZCXF16, ZWL^{+}18, ZLT^{+}18, ZSW^{+}19, ZL10]. Performance-Aware [CLY^{+}19, Has16, WKW16]. Performance-Based [EHWX10]. Performance-Centric [CFLL18]. Performance-Energy-Temperature [SAF16]. Performance-Guaranteed [ZWL^{+}18]. Performance-Oriented [Kao15]. Performances [LHL^{+}13a]. Period [LLZ^{+}18a]. Periodic [CPM^{+}10, GHW^{+}16, HCY^{+}12, HLY^{+}14, Lee12, ZGL10]. PeriSCOPE [FGJ^{+}15]. Permutation [CFJ15]. Persistence [LLH^{+}15a]. Persistency [GE12]. Persistent [NTDZ19, RZB^{+}18]. Personal [KDCR19, LYZ^{+}13, XLT^{+}14]. Personalized [FRS^{+}16]. Perspective [DWT^{+}16, Jia14b, LS19, LCZ^{+}19, WFZ^{+}17].

Perspectives [LPZ12]. Pervasive [HYC^{+}12, KJvR^{+}15, SCL^{+}15, WTL10, YHC^{+}13]. Pesky [CJBW16]. Petri [MSB11, ZJLS12]. PF [PKG14]. PFP [WMJ^{+}19]. pGraph [WKC12]. Phase [CBF^{+}17, LZWZ19, SEAH16, ZYLC14, dCAB19]. Phase-Based [dCAB19]. PHAST [PB19]. PHEVs [MBO15]. Phi [CRS^{+}17, LSW17a, LLH^{+}15b]. Phoenix [PJC^{+}13]. Phone [WYX^{+}15]. Photo [ZSW^{+}19]. Photonic [CDPM18]. Phylogenies [SJVR15]. Phylogeny [MB12]. Physical [Ano11c, CYZ^{+}13, CTX^{+}12, HGY^{+}14, HWNS15, LQY^{+}12, LCGC14, Li14c, LCSC12, MV12, RXD12, SCC11, YQZC12, ZYL^{+}17, PKL^{+}12]. Physical/Virtual [SCC11]. PIC [ZJL^{+}17a]. Piccolo [CHPY17]. Picking [CJBW16]. Pictures [JN16]. Piece [LXBZ13]. Piece-Related [LXBZ13]. Piecewise [AF19]. Piggyback [ZCJ19]. Pinpointing [BXXC12]. Pins [CIP^{+}17]. Pipeline [LLD^{+}18]. Pipelined [CCV19, HWZE10, HA13, HWQ^{+}15, HLQ^{+}15a, JIP14, LGYV14, TLP12, WDH^{+}16, ZD12]. Pipelined-RAM [WDH^{+}16]. Pipelines [FGJ^{+}15, RKRK17]. Pipelining [KN16, MG18, WYY^{+}12]. Pivoting [KLFD13]. Pixel [RZB^{+}18]. Place [SLL16]. Placement [CSW^{+}12, CTX^{+}11, CHLC15, DGC17, DY16, GLBJ18, HWL^{+}17a, LPSS19, LHXP18, LCLD13, Man16, NCGP19, NVS16, PKS14, PHXL19, RCFW10, SSF16b, SLSG18, TZC19, TMJ14, WWX^{+}13, WUH^{+}17, uRILP17, XTFC17, XXL^{+}19, YWY^{+}17, YZL^{+}17, ZG11, ZWL^{+}18]. Placements [Tse13, XLX^{+}16]. PLAN [CTP^{+}17]. Planar [LMSRSR13, ZZF10]. Plane [ATACA18, WX15, ZWY^{+}17]. Plane-Centric [WX15]. Planning [CEK16]. Platform [CZRB18, CCCY16, EHM^{+}17, HZT18, HYX11, LS17a, LZY^{+}19, LS14, MC10, SB19, SZ11, WTTH17]. Platform-Based [HYX11]. Platforms [Agr14, AKT^{+}15, BBD^{+}19, CCKF15, CLL^{+}17, CDR15, CRRR15, DED^{+}19, DCL^{+}10, ECV16, GTT^{+}17, LMD16, LW15, MSW^{+}12, OPJ^{+}19, PAB13, PVS18, PVQ15, PGGS19, RRM^{+}15, SDV18, SDG17, SVL^{+}16, TTG^{+}15a, TP14, WV17, WVM19, ZLLD18, ZWT^{+}19]. Play [LTW^{+}14]. Playback [Hu14]. Playback-Rate [Hu14]. Playing [BHS^{+}19]. Plug [LTW^{+}14]. Plug-and-Play [LTW^{+}14]. PMC [Cha11, CH14, LKT11, YLM^{+}15]. Pocket [MMSS15]. Point [SY17, XZT^{+}13, XHZ^{+}13]. Pointer [HCH^{+}12, SYXL16, VMB17]. Pointer-Rich [VMB17]. Points [ERSR13]. Pointwise [DTLC19]. Poisson [WJB14]. Policies [BLLP15, DBA17, GLRT18, Hur13, HKkY^{+}16, LLpC15, LC11, RCC^{+}14, SL16, VM12, WMZ^{+}15]. Policy [CTP^{+}17, EMW16, GGY^{+}19, GZZ^{+}13, HSMY12, HFY^{+}14, LZWZ19, LLFL15, SJR17, WLX^{+}15, XWLJ16, YJR15, XWS17, ZJTZ14, MBO15]. Policy- [CTP^{+}17]. Poll [SL13]. Poll-Based [SL13]. Pollution [AGG17, LGJ^{+}17, WXYX14]. Polynomial [IIKO13]. Polynomial-Time [IIKO13]. Pool [DSJ16, KMMR13, PYH19]. Pool-Like [PYH19]. Pooling [ZTZ^{+}18a]. Popular [CSM^{+}13]. Popularity [CE17, LSN19]. Popularity-Aware [LSN19]. Population [SOK^{+}19]. Port [Agr14, GZY^{+}15, ZD12]. Portable [AWWS19, HAU19, PB19]. Position [CCT10]. Positioning [LHF^{+}15, WYX^{+}15]. Positions [LJG12]. Possession [WZ14, ZHAY12]. Post [DLC^{+}16, QZZ^{+}16]. Post-Deployment [DLC^{+}16]. Potential [RZW^{+}13]. Potential-Based [RZW^{+}13]. Potentials [WWL^{+}15]. POVA [ZLLZ13]. Power [ASYK^{+}19, BCP^{+}14, CVM^{+}15, CLJ11, DCW^{+}15, DGC17, DSM14, FYH^{+}15, GIRT19, GPF12, HTA10, IZA18, JMZD12, JWK^{+}16, Jia14a, LGJZ16, LXHS12, LWW^{+}13, LSL^{+}18, LCA13, LGG^{+}14, PCFP16, PD14, QLC13, RPYO11, SY17, SCC11, SKKK16, Tak14, TKS11, THL13, TKP12, Van14, WCF10, WMW11, WW11, WWCZ11, WKK11, WCLK12, WWZ^{+}16, XLM^{+}12b, YYY^{+}14, YC18, YHS^{+}14, YGL^{+}15, YJC15, YLR12, ZL11, ZWL^{+}18, ZMW17, ZYSH14, aaGZ19]. Power-Aware [CVM^{+}15, WWCZ11, ZWL^{+}18]. Power-Efficient [SY17, TKS11]. Power-Performance [Jia14a, WKK11]. Power-Proportional [LCA13]. Power/Energy [PD14]. Power/Ground [LWW^{+}13]. POWER8 [FES^{+}17]. PowerPack [GFS^{+}10]. Powers [ZLY^{+}14]. PPS [HLeS^{+}15]. Practicable [CMB18]. Practical [CJZ^{+}16, HLWV14, LYZ^{+}16, Man16, MSSV18, ME15b, TMJ14, WYCZ14, XHC16, YYL^{+}13, YY14]. Practice [CJBW16, TZY^{+}18]. Practices [RSW^{+}17]. Pre [JCW^{+}19]. Pre-Scheduling [JCW^{+}19]. Precedence [CC13b, Che18a]. Precise [SZL^{+}12]. Precision [GS11b, ITW^{+}14]. Preconditioning [DFGG13]. Predicates [SK14]. Predict [DIAR16, PWRL18]. Predictable [LGM^{+}17, PH11, XZJ^{+}19]. Predicted [WUH^{+}17]. Predicting [ZCXF16]. Prediction [AM19, BMJ^{+}17, CCLW15, DBA17, ERRG18, ELX^{+}11, HCL^{+}12, HCZ12, HLY^{+}14, IdM12, JJW11, KKC17, LZWY14, LWC^{+}17, SA11, WSWY15, WRL15, WHYZ10, YYK11a, YYK^{+}11b, YCW12, ZWZ^{+}13, ZWL17, ZHZL17]. Prediction-Based [YCW12]. Predictive [BCTB13, HZT18]. Predistribution [RM11]. Preemption [SL14, WGZ16]. Preemptive [ATZZ14, BJC^{+}18]. Preference [CL15, MTDD17]. Preference-Aware [CL15]. Prefetch [VGMA10]. Prefetching [CHHK19, CLL^{+}19, DD11, LTGI16, ZTZQ19, ZSW^{+}19]. Prefix [KPA13]. Prefixes [PT11]. Presence [CIH13, HP14, MR16]. Preservation [LLG15b]. Preserving [ACCP12, CWL^{+}14a, CZS^{+}16, GZZ^{+}13, GZX14, HSMY12, HXC^{+}11, HLeS^{+}15, JGJF18, LLY^{+}14, LC11, LNXY15, LWL^{+}17, LLL^{+}12, LLS13, SWC^{+}14, TZB^{+}14, YRLY16, YY14, ZZR12, ZLN^{+}13, ZJL^{+}17a, ZLDC15]. Pressure [LN17, TLP15]. Prevention [CWL16, CRD11]. Price [LLLZ16]. Prices [LYY16]. Pricing [AHSH^{+}16, CLL11, CLZ^{+}18, DG15, LH17, MBO15, SL16, TWT16, TKP12, WS14, ZWLW16, ZYL^{+}17]. Primary [MS13b, WJTL13, YZHZ17]. Primitives [SP15]. Priori [ZJTZ14]. Prioritizing [ZGGW13]. Priority [ATZZ14, HxjGG19, LWZ^{+}13, QF14, WL13, WGZ16, ZD16a]. Priority-Based [LWZ^{+}13]. PrIter [ZGGW13]. Privacy [ACCP12, Ano12c, BMJ^{+}17, CLL^{+}14, CWL^{+}14a, CZS^{+}16, DT14, DZLC15, DCA19, GZZ^{+}13, GZX14, HSMY12, HXC^{+}11, HLeS^{+}15, IB14, JGJF18, LRW12, LLY^{+}14, LLG15b, LC11, LNXY15, LWL^{+}17, LLL^{+}12, LLS13, MS12, RYLZ10, RWLL14, SWT^{+}17, SILJ11, SZZF10, SWC^{+}14, TZB^{+}14, XTHD10, YOWA14, YRLY16, YY14, ZZR12, ZLN^{+}13, ZJL^{+}17a, ZLDC15, ZXG^{+}19, LSL14b]. Privacy-Aware [DZLC15, SWT^{+}17]. Privacy-Conscious [XTHD10]. Privacy-Enhanced [RYLZ10]. Privacy-Preservation [LLG15b]. Privacy-Preserving [ACCP12, CWL^{+}14a, CZS^{+}16, GZZ^{+}13, GZX14, HSMY12, HLeS^{+}15, LLY^{+}14, LNXY15, LLL^{+}12, LLS13, SWC^{+}14, TZB^{+}14, YRLY16, ZZR12, ZLDC15]. Private [JRV^{+}13, LC11, TLSL15, TLL^{+}16, US16, WHN^{+}19]. Proactive [BHS^{+}19, CCLW15, NVS16, SBC^{+}10, WLL^{+}19, WS14]. Proactive-Reactive [SBC^{+}10]. Probabilistic [BBCTA18, GS11a, HJPL14, HA10, HCH^{+}12, LAdS^{+}15, LYGX12, LYL15, LWL^{+}17, Mis14, PFAF16, YTZ^{+}11, ZDG^{+}14]. Probabilities [KKC17]. Probability [NLGQ14]. Probe [ZLLZ13]. Probing [GJC^{+}13]. Problem [Ara11, CW19, DWW^{+}11, DPRT11, FDFZB13, GS17, HMS^{+}18, HH11, HLY10, KN12, LCL^{+}11, LLZ14, LWZ12, RBSS11, WWH13, WRB11, YXSS13, YTL^{+}19, ZG11, ZT14, ZRTL15, ZT16]. Problems [YPL13]. Procedure [VS14]. Process [HWQ^{+}15, Man16, TMJ14, WLX^{+}15, ZXG^{+}19]. Processes [MR16, RLVTMG^{+}16]. Processing [AHSK17, AZW^{+}19, BVFGSFAF17, BSM^{+}11, BSL^{+}17, CC18, sCCyW14, CYW^{+}18, CHHK19, CCHH19, DB18, DFGG13, DWW^{+}15, DBG^{+}14, EALM17, FHW11, GRUMG17, HHWZ17, HT16, JDB^{+}14, JCW^{+}12, JCW^{+}19, KKC18, LLLG13, LHZJ19, LLLC17, LN17, LABQ18, MS13a, MMdE19, MZK19, MTMR18, MRH^{+}16, MP16, NCGP19, PLG19, PSL^{+}11, PRS^{+}11, QP16b, RZB^{+}18, TG13, TFLL18, TS16, VLP16, WMZ^{+}15, WCYL19, WMWW19, WK11, WLL^{+}19, WW12, XBZL17, XL17, YHL^{+}18, YXLJ16, ZGGW14, ZLS^{+}18, ZZSZ18, ZH14b, ZSC^{+}17, ZKP^{+}19]. Processing-In-Memory [WCYL19]. Processor [AM19, CA13, HV11, IGEN11, IG11, KGI17, LJ16, SV19, SME10, TZT^{+}16, TWSW17, TBC12, YMG15]. Processors [AFMM17, DSM14, DDP^{+}19, FHLG11, GHZZ16, HWF18, JWK^{+}16, JZW^{+}17, KHN16, KM18, KAA16, Lee12, PD14, RCV^{+}13, SZA11, SWRQ18, SAF16, SA11, TS18, VNA^{+}16, WKK11, YP13, Zha12, ZCXF16, ZYX^{+}10]. Producer [MBF19]. Product [AA14, CLH13, CH15, LLH14, LHJ12]. Production [CWLS19, DDW^{+}19, MWZ^{+}13]. Profile [SV19]. Profiling [DLC^{+}16, GFS^{+}10, YWW^{+}15]. Profiling-Based [YWW^{+}15]. Profit [CHLZ13, MLXG19, MLT^{+}19, XZH14]. Program [CLC^{+}12, CM10, DLC^{+}16]. Programmability [EMW16]. Programming [AWWS19, Ara11, BBK17, BFD19, BBL^{+}16, CEK16, DED^{+}19, DMCN12, FA19, HA11, LdSS^{+}13, MGS12, PB19, TBÁ^{+}19, TYS^{+}12, TFM^{+}16, XTFC17, YTMS16]. Programs [CC13a, CJW^{+}15, LLZ^{+}18a, ME15a, WYY^{+}12, WWLJ14, ZRQA14]. Progress [LAdS^{+}15, LSL^{+}14a, PH18, SPH^{+}18, WLX^{+}15]. Progress-Dependence [LAdS^{+}15]. Progressive [CW15, HOZ12, XLL^{+}18, YXSS13]. Project [SOTN12]. Projective [CMVB17]. Promenade [CFLL19]. Prone [BBR12, DGFRR18]. Proof [LLZ18b, NLY15, ZY14]. Proofs [DKL^{+}19, LNZ^{+}13]. Propagation [BAMJ12, GG13, LCL^{+}15, PBD^{+}13, WZZ^{+}13, XP12, YY14]. Propagation-Based [GG13]. Proper [TWW^{+}15]. Proper-Temporal-Embedding [TWW^{+}15]. Properties [CH14, DGG^{+}19, GIP^{+}13, TL14, TCT14, YHC^{+}13]. Property [HYC^{+}12]. Prophet [ZJL^{+}17b]. Proportional [HKH^{+}10, LCA13, TYLG13]. Proportional-Fair [TYLG13]. Proportionally [CFLL19]. PROSA [AF18]. Prosumer [PCFP16]. Protected [ZML13]. Protecting [MS12, SYL^{+}16]. Protection [AFMM17, Bis18, CL14, DHBB12, DCA19]. Protector [YTZ^{+}11]. Protein [WKC12]. Protocol [ANN^{+}13, ACCP12, AF18, CBK^{+}10, CRRR15, DGF12, DCA19, EHNS13b, FLH13, GFMR13, HA10, JCWB10, LLGP13, LMPR12, LXHL11, kL11a, LLC10, LNZ^{+}13, LWJ^{+}15, LNXY15, LXBZ13, MLC^{+}15, MY11, PDFJ13, RZH^{+}11, RAG10, SCC11, SL11, SOK^{+}19, TWL^{+}15, TLRW15, WL14, WML15, WL15, Xia14, XLLZ11, YLSQ13, YJ13, YCMX17]. Protocol-Driven [AF18]. Protocols [Che14, rCHG10, CLJ11, LSL^{+}14a, LY16b, LW12, LLM^{+}14, MLS15, SS12, TLSL15, TJLL12, XHL^{+}15]. Provable [SX10, WZ14, ZHAY12]. Provably [KKW13, TXL^{+}14]. Provenance [WHB16]. Provided [WWL^{+}15]. Provider [LPSS19, SL16]. Providers [LSW17b, LYZL18, Sam14a]. Provides [MLK15]. Providing [CSP13, MMACS10, YOWA14]. Provision [CSP13]. Provisioning [ALZ17, AIAD^{+}18, CPGT14, CAKRY16, DCW^{+}15, HLWV14, KJL^{+}16, LZ12, LWC^{+}17, LDYZ15, LLZ18b, LCA13, MNG15a, MBV11, NIP11, NMG15, NZM^{+}16, PSL15, PKCB11, SWL17, TNZ^{+}12, TCS11, VLRP15, WHGS17, XBZ^{+}16, XZJ^{+}19, YZL^{+}17, ZLW^{+}14, ZT16, ZHCL17, ZWG^{+}16]. Proxies [DBAT11]. Proximity [CYZ^{+}13, SLW15, TLSL15]. Proximity-Aware [SLW15]. Proxy [HN10, XTXH13]. Proxy-Based [XTXH13]. Pruning [JLKG17, LCD^{+}17]. Pseudopartitioning [ZML13]. PSMPA [ZLDC15]. PSO [GLC^{+}15]. PSO-Optimized [GLC^{+}15]. PTAS [MNG15a]. Public [CB14, CPGT14, JWNS19, LXXH16, PGP^{+}17, Rao14, WWR^{+}11, ZSW^{+}15]. Publicity [OMMZ14]. Publish [JHMV12, MC14, MFO^{+}13, QCZ^{+}15, TKR14, WM15]. Publish-Subscribe [MC14]. Publish/Subscribe [JHMV12, MFO^{+}13, QCZ^{+}15, TKR14, WM15]. Publishing [Ano12i]. Pump [HDL^{+}15]. Puppet [KE16]. PURE [CZZ^{+}16]. Purpose [PBD^{+}13, STMM17]. Pursuing [XLM^{+}11a]. PUSH [HLQ^{+}15a]. Putting [LPSS19].

Q [CC18, CSR^{+}17, ZYL^{+}16]. Q&A [LS17d]. qcAffin [HT16]. QoE [VMN^{+}16]. QoF [LHD^{+}14]. QoS [ADZZM15, BDLS13, Bru14, sCCyW14, CZYL14, CNC^{+}14, Guo14, HLCB^{+}17, HZT18, KM19, LCSC12, MM12, MMACS10, SLL13b, TCS11, WMC^{+}19, yWeH11, XXL^{+}19, ZWZ^{+}13, ZHZL17]. QoS-Aware [ADZZM15, sCCyW14, Guo14, WMC^{+}19, yWeH11, XXL^{+}19]. Quad [aaGZ19]. Quad-Criteria [aaGZ19]. Quadboost [ZTZ18b]. Quadtree [ZTZ18b]. Quality [BB13, CZZ^{+}16, CP15, CLHK11, DCW^{+}15, DLZH16, DN19, DLZ^{+}14, HCC^{+}12, IZA18, JMS^{+}18, LHD^{+}14, LV15, LRJX13, LS17b, LCS^{+}15, TWL^{+}15, WGCG18, YL10]. Quality-Aware [DN19, IZA18, WGCG18]. Quality-of-Experience [TWL^{+}15]. Quantifying [FBCB18, LLCH12, OMMZ14]. Quantitative [LRW12, OKT^{+}16, YLR12]. Quantum [CLYR16, HAY^{+}18]. Quantum-Inspired [CLYR16]. Quasi [CCLW11, GWL^{+}11, HLYJ19, LYL16]. Quasi-Kautz [GWL^{+}11]. Quasi-Output-Buffered [CCLW11]. Quasi-Streaming [HLYJ19]. Quasi-Tridiagonal [LYL16]. Queriable [KTK11]. Queries [DWW^{+}15, DT14, GYW^{+}19, HXLF15, LCL^{+}16a, MTDD17, PLG19, TXZ^{+}11, XTHD10]. Query [CC18, CYC^{+}16, GZW^{+}18, HL12a, JCW^{+}12, LHYW15, SMTZ17, TJLL12, TOA13, XXL^{+}19, YNW13, ZYZC12]. Query-Based [GZW^{+}18]. Query-Centric [HL12a]. Query-Log [TOA13]. Querying [JLKG17]. Queue [ATZZ14, GLRT18, HT16, hKYY11, KSW18, ME15b, WL13, ZD16a]. Queued [WYLH18]. Queueing [TCDMRP17, WPT17]. Queuing [Che11, KMM12, RS10].

R [BFPB10, KMM12]. R-Trees [BFPB10]. Rabin [SCHT16]. Raccoon [ZWFX17]. Race [JEW^{+}18, LZL^{+}18, LLZ^{+}18a, Tic14]. Race-Condition-Aware [LZL^{+}18]. Races [ZRQA14]. Radar [GRUMG17, LL11, PRS^{+}11]. Radial [MKSN18]. Radio [AKP14, BV10, CJH^{+}14, CLM^{+}15, DWX14, FJV^{+}18, HWDP10, HWC^{+}14, JCLJ12, JZY^{+}15, LCL^{+}14, LCSC12, LLCL12, LZC^{+}12, MS13b, SA11, WWW^{+}18, XJL^{+}14, ZY14]. Radix [IGEN11]. RAID [CLLX18, LWT^{+}18, MWZX14, SSF16a, WQZ^{+}15, YXWW14, YJC^{+}16, ZWL^{+}15, ZWL^{+}16a]. RAID-4 [ZWL^{+}15]. RAID-5 [MWZX14]. RAID-6 [SSF16a, ZWL^{+}16a]. RAIDs [YJC^{+}16]. Rail [ZMF10]. RAM [AFMM17, WDH^{+}16]. RAMPS [NTA^{+}16]. RAMSYS [LRYJ17]. Random [BYZ^{+}16, CCFS11, CMB18, CJ16, CLT^{+}17, CFLL19, FMY^{+}18, LAT^{+}15, PDH10, SGGB14, TFKN17, WLS^{+}11, XAK17, ZFT^{+}15, ZYT^{+}15]. Random-Forest [BYZ^{+}16]. Randomize [FKMC15]. Randomized [IIKO13, MKOK14, PSMD18]. Range [KTK11, MA14, ZH11]. Range-Based [MA14]. Range-Free [ZH11]. Range-Queriable [KTK11]. Ranked [CWL^{+}14a, CZS^{+}16, WCRL12, XWSW16]. Ranking [SWC^{+}14, ZWZ^{+}13]. Rapid [MYPL18, PT11, AWWS19]. RASS [ZLGN13]. Rate [CYX^{+}14, CCL13, CFLL19, Hu14, KCK14, LRJX13, LCW11, LGG^{+}14, TDL^{+}19]. Rate-Distortion [TDL^{+}19]. Rateless [AGG15, SGGB14]. Rates [MYPL18]. Rating [AI15]. Ratio [ZQWL17]. Rational [ST10]. Rationally [CW15]. RC [CCLW15]. RC-Based [CCLW15]. RCDA [CLLS12]. RCSMA [KZW^{+}12]. rCUDA [PS19]. RDF [AHSK17, CC18]. RDMA [CSW^{+}17, CLA^{+}19, Pan14, WMS^{+}19]. RDMA-Enabled [Pan14, WMS^{+}19]. Reachability [CYC^{+}16]. Reaction [XLL11, XLH^{+}15]. Reactions [KEGM12]. Reactive [KAG17, SBC^{+}10]. Read [AJK^{+}17, CZL^{+}16, DMS^{+}12, WH16, WDH^{+}16, XX16]. Read-Copy [DMS^{+}12]. Read-Mostly [CZL^{+}16]. Read/Write [WDH^{+}16]. Reader [GFMR13, JGZZ14, ZCX^{+}14]. Reader-to-Reader [GFMR13]. Readers [IPQ19]. Reads [SLSG18, TZT^{+}16]. Real [BJC^{+}18, BVEAGVA10, BVFGSFAF17, BMB^{+}10, CCKF15, CLT13, CCL13, CCC^{+}16, DRRCB18, DLA^{+}18, DCL^{+}10, DLB^{+}19, ELX^{+}11, GRUMG17, GLC^{+}15, HZW^{+}14, HLZY15, HWG^{+}19, HAZ17, HxjGG19, HRGE17, HKH^{+}10, HJF16, HSX^{+}12, KM10, Kum14, Lee12, Lee17, LTW^{+}14, MZK19, NSLV16, PCFP16, PFAF16, PVS18, PM13, QF14, RGPH15, SFL^{+}14, SEAH16, SS12, SL14, SHX^{+}10, SFA^{+}17, TXWL11, TL16, VLP16, WYC^{+}15, XZX^{+}17, YRLY16, YQH16, YC12, ZGL10, ZLGN13, ZYL^{+}17, ZMF10, ZWQ^{+}15, ZWG^{+}16]. Real-Time [BJC^{+}18, BVEAGVA10, BVFGSFAF17, BMB^{+}10, CCKF15, CLT13, CCL13, CCC^{+}16, DRRCB18, DLA^{+}18, DCL^{+}10, DLB^{+}19, ELX^{+}11, GRUMG17, GLC^{+}15, HZW^{+}14, HLZY15, HWG^{+}19, HAZ17, HxjGG19, HRGE17, HKH^{+}10, HJF16, KM10, Lee12, Lee17, MZK19, PCFP16, PFAF16, PVS18, PM13, QF14, RGPH15, SFL^{+}14, SEAH16, SS12, SL14, SHX^{+}10, SFA^{+}17, TXWL11, TL16, VLP16, WYC^{+}15, XZX^{+}17, YRLY16, YQH16, YC12, ZGL10, ZLGN13, ZYL^{+}17, ZMF10, ZWQ^{+}15, ZWG^{+}16]. Real-World [HSX^{+}12, NSLV16]. Realistic [Li10, MNE14, RSW^{+}17]. Realization [MVC^{+}18]. Reallocation [XS10]. Reasoning [AOW^{+}12]. Rebalancing [HCSC13]. ReCA [SEA18]. Receive [GDM^{+}13]. Receive-Side [GDM^{+}13]. Receiver [KZW^{+}12, NHN17, NHN18, dBK11]. Receiver-Based [KZW^{+}12]. Receiver-Initiated [dBK11]. Reception [CWJS11, RVW^{+}15]. Rechargeable [RCC^{+}14]. Reclamation [TWZW11]. Recognition [CC17, LAT^{+}15, MMNN16]. Recommendation [CZYL14, MDZC14, SytL19, YGL13]. Recommender [LLAL18]. Reconfigurable [EAMEG11, HWZE10, Kao15, LWZ^{+}16a, LLD^{+}18, NTA^{+}16, SEA18, SZ11, YLL^{+}17, YLLW16, YYL^{+}17, YN17]. Reconfiguration [GYLW18, QZG^{+}16, RGBC11].

Reconfigurations [GBFS16].

Reconsidering [FSSZ16]. Reconstruction [HLQ^{+}15a, KXL^{+}14, LCGC14]. Record [AHSH^{+}16, LZH^{+}16, SF10]. Record/Replay [LZH^{+}16]. Records [LYZ^{+}13]. Recoverable [CLLS12]. Recovery [Che16, CWLS19, FSSZ16, GTM^{+}17, JMA^{+}18, LWT^{+}18, SSLF17, SBC^{+}10, YXWW14, ZLX^{+}14, ZKSY14]. Rectangular [JP12]. Recurrence [CCV19]. Recurrent [LCZ^{+}19, PVS18]. Recursive [IvS10, SAA17, TWL12]. REDEFINE [MMNN16]. Redirection [XBZ^{+}16]. Redistribute [ZWL^{+}15]. Redistribution [YLR12]. ReDS [AAAK^{+}14]. Reduce [CP17c, NFD10]. Reduced [VBC19]. Reducing [AJM12, CAD^{+}18, CJZ12, NTKK15, SAA17, Tak14, XVC17, YSS^{+}17]. Reduction [CC13a, EK10, FYH^{+}15, GS11b, HA13, LKD10, Nov15, SYL^{+}14, TLP12, YHS^{+}14, ZHL^{+}15]. Redundancy [LG10, MHL^{+}16, SEAH16, XDLZ19, YSS^{+}17]. Redundant [SCHT16]. Reed [LWCL18]. Reference [HPP15]. Referral [ZLL^{+}15]. Refined [SWT^{+}19]. Refinement [NTDZ19, RAS17]. Refining [SLL13b]. Refresh [ZLT^{+}18, MMNN16]. Regain [ZWL^{+}15]. Regenerating [CL14]. Regenerating-Coding-Based [CL14]. Regime [RMM16]. Region [GCZ15, HWL^{+}17a, VWDM14, ZHX^{+}19]. Region-Based [ZHX^{+}19]. Region-Level [HWL^{+}17a]. Regions [JEW^{+}18, LCG^{+}13]. Register [ACE^{+}19, BBR12, EALM15, IPQ19, Mit17, TCYF16]. Register-based [EALM15]. Registration [Bar10, WYZ^{+}19]. Registration/Retrieval [Bar10]. Regression [CZZ^{+}16, ZCXF16]. Regret [CYC^{+}15]. Regular [BBR12, CJBW16, FMY^{+}18, WPKL13]. Regularization [CLC^{+}12]. Regulation [ZXG^{+}19]. Regulatory [ZASA10]. Relabeling [HH11]. Related [LXBZ13, TLP15]. Relation [ZSY14].

Relational [YNK^{+}17]. Relations [BS12, CCV19]. Relationship [XAY^{+}14]. Relative [DTLC19, DAJ14]. Relaxation [SSM^{+}18]. Relaxation-Based [SSM^{+}18]. Relaxed [AA12, RLSK17]. Relaxing [ZYL^{+}16]. Relay [CMC^{+}15, GTS^{+}15, TYLG13, WWL11, ZGXJ14, ZY14, Zhu14]. Relay-Union [CMC^{+}15]. Relaying [CLL11, HLS^{+}15]. Relays [PM13]. Release [HV11, YCMX17]. Reliability [CMT^{+}17, CGZQ13, Che16, DOLG16, GAKR11, HAZ17, HP14, JHR^{+}14, LWT^{+}18, LLpC15, LZNX11, LTMD11, MV16d, PDH10, PH12, TSN10, Wan19, WMJ^{+}19, ZHL19, ZQSY13, ZXL^{+}17, aaGZ19]. Reliability-Oriented [LZNX11]. Reliability-Sensitive [Wan19]. Reliable [BV10, CBK^{+}10, FWH^{+}18, LGYV14, LHL17, LLL^{+}14b, MLS15, PDFJ13, PL16, Ven14, XLM12a, YWY^{+}17, ZGH14]. Relieving [LN17]. Remote [LS17a, LZCK14, MWZ^{+}14, PM13, WMZ^{+}15, WMS^{+}19]. Removal [LG10]. Rendering [dLMPG19]. Rendezvous [KPG^{+}12, LLCL12, Mis14]. Rendezvous-Based [KPG^{+}12]. Reneging [HLCB^{+}17]. Renewable [CCJ19, LQW^{+}18, LLFL15, LH17, LGG^{+}14]. Reorder [LDLL18, ZGY15]. Reordering [GLRT18]. Reorganization [ZWL^{+}16a]. Repair [LC14, ZLL17a]. Repair-by-Transfer [LC14]. Repartitioning [CATC11]. Repeated [XZSG12]. Replacement [TWZW11]. Replay [LZH^{+}16]. Replaying [GZW^{+}18]. Replenishment [NNKL13]. Replica [GLBJ18, XAY^{+}14, ZG11]. Replicas [WDH^{+}16]. Replicated [CRRR15, FWH^{+}18, GAKR11, HK18, KB17, LV17, STMM17, TOA13]. Replication [BAAT16, CB14, CYW^{+}18, CLKR15, FHW11, HAZ17, JKS13, KKW18, LHL17, LTC^{+}19, LS17c, LJL^{+}11, LSC16, LSN19, NOR16, NTK^{+}15, NCB17, NTWL11, OUA11, PRR^{+}16, QP16a, QPB^{+}17, She10a, She10b, SS17, THT^{+}15, WKK17, WL12b, XXL^{+}19, XVC17]. Replication-Based [CYW^{+}18, NOR16]. Representation [LZ10, LLZ^{+}18a, TTG^{+}15b, XH10]. Reproducibility [Par19b]. Reproducible [HCA16]. Reprogramming [PB12]. Repudiation [LLG15b]. Reputation [AAAK^{+}14, CSSL15, dCCF15, NSY^{+}16, RBM15, ST10, SLL13b, SLSL16, TNLM17]. Reputation-Based [NSY^{+}16, ST10]. Reputation-Enhanced [AAAK^{+}14]. Request [HLCB^{+}17, LPP13, SZL^{+}12, WW13, XBZ^{+}16]. Requests [LHXP18, MLXG19, SS17, ZTZ^{+}18a]. Required [LCLD13]. Requirement [HV11]. Requirement-Aware [HV11]. Requirements [JAJ^{+}19, KOPS10, LYZL18]. Research [Sto10f]. Reservation [LW14, MPM17, PFAF16, VM12, ZQL^{+}16]. Reservation-Based [LW14, VM12]. Resident [JDB^{+}14]. Residential [GPF12]. Residual [MGB18]. Resilience [FPRG16, HLWV14, NL11, SLSL16, YCWL14]. Resilience-Complexity [NL11]. Resilient [AVA^{+}17, DB18, LMPR12, LXHL11, LYGX12, LCS14, LSN19, MSSB14, TVG13, VS19]. Resistant [KZW17, KSP10, SLLZ16]. Resisting [XTXH13]. Resizing [YOK^{+}17]. Resolution [WMWW19]. Resource [AHSH^{+}16, ALZ17, ASBL15, AIAD^{+}18, BEDCR13, BSM^{+}11, CC10, CB16, CB13, CPGT14, CBF^{+}17, CZX^{+}19b, DW13a, DW13b, HTZY17, HWG^{+}19, HKA12, HCZ12, HLWV14, HKkY^{+}16, JWA10, JWNS19, KALK^{+}18, KJL^{+}16, KKC17, LGD14, LPP13, LdSS^{+}13, LMZG15, LCG^{+}16, LTC16, LLLZ16, LRYJ17, LHXP18, LTC^{+}19, LCSC12, LMAS17, LS14, LH16, LLL18, LVD11, Man18, MKVL12, MPHR17, NIP11, NZM^{+}16, OPM^{+}15, PSL15, PR19, PCP14, RG17, RCFW10, SDV18, ST10, SGGB14, SZR17, SVC12, SFA^{+}17, TNH^{+}18, TCDMRP17, TP14, VLRP15, WKK11, WLL15a, WKW16, WHGS17, WWW^{+}18, WVM19, WK11, WRB11, WYY^{+}12, WLL^{+}19, WS14, WZL^{+}19, XL10, XSC13, XBZ^{+}16, XL13, XWL^{+}19, YLC^{+}16, YBY^{+}18, ZSY14, ZYQ^{+}14, ZQL^{+}16, ZQCZ16, ZWLW16, ZJZ^{+}16, ZWE19, ZHCL17, ZWG^{+}16]. Resource-Aware [HWG^{+}19, LRYJ17, MKVL12]. Resource-Efficient [LTC^{+}19, XWL^{+}19]. Resources [CRZH15, DL17, GHW^{+}16, HZW^{+}14, LDYZ15, LABQ18, MNG15a, MP16, WWL^{+}15, WYZ^{+}19, XCZ^{+}15, LYZL18]. Response [AWZ15, HGS^{+}19, LZ12, LLY^{+}14, PHGR17, WWCZ11, WX11, ZLLD18, ZKSY14]. Response-Time [PHGR17]. Restart [STK^{+}19]. Restoration [WMT^{+}11]. Restore [LCYW16, ST18, TWW^{+}18, WHZS19]. Restore-Express [ST18]. Restraining [WJX^{+}14]. Restricted [LXZH16]. Restructuring [SMS^{+}13]. Resubmission [PP12]. Result [HHWZ17, MBV11]. Result-Data [MBV11]. Retirement [USP^{+}12]. Retrieval [Bar10, CJL^{+}12, HOZ12, LWZ^{+}16b, US16]. Retrieving [dOSdM13]. Retroscope [CADK19]. Retrospective [CADK19]. Reuse [GHL^{+}13, Guo14, SV19]. Revealing [ZLF^{+}11, ZYSH14]. Reverse [APCH^{+}11]. Reversibility [Lee17]. Reversion [ZLJL17]. Reviewer [Ano11b, Ano13b]. Reviewers [Ano10, Ano12b, Ano14b, Ano15b, Ano17b, Ano19b, Ano19c, Ano17c]. Revisiting [TJLL12]. Revocable [YJ14]. Revocation [HN11, LNA^{+}13]. Rewarding [WML14, LSL14b]. Rewriter [KAC^{+}15]. Rewriting [WHZS19]. RF [NML^{+}14, WMS^{+}19]. RF-Based [NML^{+}14]. RF-RPC [WMS^{+}19]. RFHOC [BYZ^{+}16]. RFID [ACCP12, BXXC12, sCCyW14, CCS^{+}12, GFMR13, JGZZ14, KWZ^{+}12, KZW^{+}12, LNZ^{+}13, LLM^{+}14, LXZB15, QNLN11, QLNN13, SLY^{+}14, SDL^{+}15, WZFG13, WSSZ13, WSS15, XHL^{+}15, YNW13, YQH^{+}15, ZZG^{+}11, ZCX^{+}14]. Rich [DCC^{+}19, GYW^{+}19, JHMV12, VMB17]. Riding [LHW11]. Right [SYL^{+}16, XALS17]. Right-Sizing [XALS17]. Ring [GGY^{+}19, LCL^{+}16b, MKOK14]. Ring-Based [GGY^{+}19]. Rings [ZPD11]. Risk [JRV^{+}13, SLG^{+}18, ZCJY14, ZSW^{+}15, ZYSH14]. Risk-Constrained [ZCJY14]. Risk-Graph [ZYSH14]. RMA [SPH^{+}18]. Road [JGHD10, XVC17]. RoB [LDLL18]. RoB-Router [LDLL18]. Robinhood [PWJ16]. Robot [CEK16]. Robotic [ZS13]. Robots [IIKO13]. Robust [AI15, BSM^{+}11, CIH13, DKL^{+}19, FC10, FGLP10, JKT11, LCL^{+}14, LXXH16, LSB^{+}18, MS13b, MY11, OPM^{+}15, WLX13, YOWA14, YP13, YLW^{+}14, ZYW^{+}14a]. Robustness [CJ10, CNMA11, MLVD12, YQZC12]. Rogue [HST^{+}11]. Rollback [CHPY17]. Rolling [GBFS16, LM12]. Rollup [GBFS16]. Roofline [DGI^{+}19]. Rotating [AR10]. Rotation [EMTX15, TMMN15]. Roughly [MP16]. Round [BAAT16, PT11, YL15]. Round-based [BAAT16]. Round-Down [PT11]. Rounds [ACS13]. Route [FC11, GKKW16, LYGX12, WYL19]. Router [BICK^{+}15, DLXS19, LDLL18, PL16, YLSQ13, ZFF16, LDLL18]. Routers [ACV17, LHM12]. Routing [ANN^{+}13, AP17, BGHG16, BTG^{+}18, BFPB10, BS12, Cha14, CWC11, CLHW13, CHD^{+}15, CSY15, CCH^{+}17, Che18b, CNC^{+}14, CCCB14, DGC17, DGF12, DWW^{+}11, DWY^{+}13, Dua95, EHNS13a, EHNS13b, EKNS17, ESGQ^{+}13, FMY^{+}18, FSM^{+}12, FC18, GAB18, GJC^{+}13, GHL14, HWX12, HWC^{+}14, KM10, KKW15, KCK14, KKY^{+}14, KOKA11, LCZZ13, LGYV14, LZB14, LNL^{+}19, LCL^{+}15, LLLH19, LX12, LGG^{+}14, MWJ^{+}14, MMYES^{+}18, MLS15, MTX^{+}11, MMSAZ11, Oru17, RZH^{+}11, RHDL11, RZW^{+}13, RS12, RGBC11, SHG11, SHG13, SX10, SLX19, SZ12, TLRW15, TLP15, TCS11, VS11a, VS11b, VS14, WLS^{+}11, WYW13, WWLX13, XL16, XWH15a, XWH15b, XLSR13, XGZW14, XSL^{+}16, YLSQ13, YXWL16, YWY^{+}17, YCW12, ZS10, ZCLS14, Dua93]. Routingin [MMSS15]. RPC [CSS^{+}13, WMS^{+}19]. RRE [ZKSY14]. RS [HLQ^{+}15b]. RS-Coded [HLQ^{+}15b]. RSD [ZH11]. RTRN [BS15]. RTS [WWH13]. Rule [HGL^{+}16]. Rumor [LHW11]. Rumors [WJX^{+}14]. Run [RMG18]. Run-Time [RMG18]. Runge [Mur12]. Running [LWZ^{+}16b, SV19, ZZH^{+}17]. Runs [CWLS19]. Runtime [AAB^{+}17, ADMX^{+}12, BBK17, CGS^{+}15, CAD^{+}18, DK17, HYC^{+}12, LWSM19, SHG13, ScFRdS15, TBÁ^{+}19, TYS^{+}12, WW12, YHC^{+}13, ZHZL17].

SaaS [Jia14a, SWT^{+}17]. SACAT [KZW17]. Safe [Iye14, RSNV18]. Sample [CLHK11, XHG15]. Sampling [WWT^{+}19, WTXG19, ZWJ^{+}19]. Sampling-Based [WTXG19]. SAN [WWH^{+}17]. SANE [HjZ^{+}14]. Satellite [BSM^{+}11]. Satellites [WZQY14, ZWQ^{+}15]. Satisfaction [KN12]. Satisfiability [LGX^{+}11]. Satisfying [NLGQ14]. Saturation [RMM16]. Saving [GF13, LYH^{+}15]. Savings [TUS13]. Scalability [AMN^{+}16, BCF13, CMT^{+}17, HD15, MHL^{+}16, ME15a, PWRL18, SBC^{+}19, US16, ZDM^{+}19, ZWL17]. Scalable [ATA18, ADZZM15, ACCP12, CHM^{+}13, CCM^{+}17, CMT^{+}17, CZL^{+}16, CCT^{+}14, CRD11, DRRCB18, DR16, DSJ16, DAJ14, DO13, DBG^{+}14, GJPPM^{+}12, GXZ^{+}15, HH13, HWJ18, HZJ^{+}11, IPQ19, IGEN11, JPG14, KZK^{+}19, KZK^{+}20, LZY12, LYZ^{+}13, Li14a, LCS14, LV15, LLZ^{+}18a, MZL^{+}19, MMYES^{+}18, MA14, MG18, MWZ^{+}13, ME15b, MMBdS14, MTY^{+}12, QLNN13, RSW^{+}17, SZL^{+}12, SHY14, SY17, SLL13a, Sib12, THE^{+}15, TWL16, TGAG13, TGFFP^{+}19, TPRH16, WJTL12, WCLK12, WM15, WWH^{+}17, XHHC13, XDMZ17, XAYM14, XML^{+}18, YOWA14, YP13, YL16, YQ16, YC12, ZLGN13, ZYLC14, ZDM^{+}17, ZCC^{+}17, ZDM^{+}19, ZWY^{+}17, ZHQ12, ZTZ18b, ZWT^{+}19]. Scale [AHSK17, Agr14, BGHG16, BCQ^{+}10, BS14, CJW^{+}15, CMB15, CL16a, CC10, CYW^{+}18, CLB^{+}19, CZX^{+}19a, CPL^{+}18, DGG^{+}19, FYH^{+}15, GGY^{+}19, GLM13, GTT^{+}17, GZW^{+}18, Guo14, HWJ18, HAY^{+}18, HZJ^{+}11, HjZ^{+}14, HJF16, JMZD12, JGZZ14, JLKG17, KCW11, Ksh10, LZL10, LMD16, Li10, LZY12, LHL^{+}13a, LCS14, LSLD17, LS17a, LLAL18, LZY^{+}19, LTC^{+}19, LHZJ19, LSL^{+}10, LHL^{+}13b, LLM^{+}14, LLL^{+}14a, LLH^{+}15a, LSCL16, MWZ^{+}14, MCJT19, MS13b, MCRC17, OKT^{+}16, QNLN11, RMB^{+}16, RMG18, SK14, SZWX15, SHF^{+}17, SDL^{+}15, TNZ^{+}12, TVG13, TKC^{+}15, TZB^{+}14, Tsa13, TTJX12, Van14, WZSL12, WCLK12, WRWW13, WJTZ14, WSWY15, WKL^{+}16, WFZ^{+}17, WV17, WVM19, WKC12, XTFC17, XHL^{+}15, XHL^{+}11, XAYM14, YQH^{+}15, YC18, YHS^{+}14, YPL13, YQLS14, YL16, ZSH^{+}11, ZLW^{+}14, ZLJ^{+}15b, ZHL^{+}15, ZJL^{+}17a, ZSW^{+}19, ZLX^{+}14, dSLMM11, LLY^{+}15, HLQ^{+}15b]. Scale-Free [BS14]. Scale-Out [LS17a, WFZ^{+}17]. Scale-RS [HLQ^{+}15b]. Scale-Up [LSLD17, LS17a]. Scale-Up/Out [LSLD17]. Scales [GTM^{+}17, ZLK^{+}16]. Scaling [BHS^{+}19, CC17, FW13, GDM^{+}13, GJC^{+}13, HLQ^{+}15b, HBS^{+}16, LHG^{+}17, LABQ18, MFO^{+}13, PGP^{+}17, SOA15, TZC19, WJTL13, WSL^{+}15, WCZ^{+}19a, WXLY16, ZWL^{+}15, ZWL^{+}16a, ZWJ^{+}18]. Scan [HH13, MIH17, Zha12]. Scanning [JGHD10]. Scatter [Trä19]. Scatterer [RZB^{+}18]. SCBXP [EHI11]. Scenarios [CWZ^{+}15, SJVR19, WMWW19]. Scene [LODB17, LZY^{+}19]. Schedulability [CLL^{+}17, SL14, WGZ16]. Schedulabilityin [Li14b]. Scheduled [PHGR17]. Scheduler [BBL^{+}16, MZK19, MMACS10, PYHY16, PKG14, YOK^{+}17]. Schedulers [HLZ^{+}19, RGPH15]. Schedules [CJ10, TWSW17]. Scheduling [ATZZ14, ANE12, AS16, ABN19, BJC^{+}18, BWB^{+}19, BVEAGVA10, BHKS^{+}17, BSL^{+}17, CP17a, CYX^{+}14, CCKF15, CC13b, CLT13, CLYR16, Che16, CBF^{+}17, CZQ^{+}17, Che18a, CHHK19, CH13, CCC^{+}16, CZWJ18, CZX^{+}19b, CVM^{+}15, CLKR15, CCK12, CRC^{+}17, CDR15, CWC^{+}13, DGC17, DWLY15, Di 17, DCL^{+}10, DPRT11, EHNS13a, FPF13, FES^{+}17, Fen14, GJLZ13, GHZZ16, GJZZ12, GHL^{+}13, GS17, HWG^{+}19, HZJ16, HxjGG19, HW13, HV11, Hu14, HWL^{+}17b, HLL18, HL12b, HYX11, HC14, ICN18, JWA10, JVW10, JTS^{+}11, JLM^{+}12, JHWY19, JCW^{+}19, KHN16, Kao15, KCH19, KJvR^{+}15, LMM18, LZ11, Lee12, LLY16, Lee17, LMSRSR12, LQY^{+}12, LTL14, LZWY14, Li14c, LSWR16, LGJ^{+}18, LLRP18, LZL^{+}18, LQW^{+}18, LMAS17, LIWJ15, LGX^{+}11, LH17, LM16, LYZ^{+}16, MLL14, MWZ^{+}14, MSSV18, MB13, MNG^{+}15b, MG18, ME15a, PD14, PVS18, QF14, RKZC14, RSNV18, RM17, SAEH19, SFL^{+}14, SMS^{+}13, SWT^{+}17, SAF16, SZR17, SBMA15, SVC12]. Scheduling [SLS^{+}16, SOTN12, SCH11, TRT19, TTH^{+}19, TZ10, TYLG13, VS19, VM12, VS15, VGMA10, WL13, WZQY14, WSC^{+}14, WGZ16, WPT17, WWW^{+}18, WS18, WYLH18, WCZ^{+}19b, WVM19, WZS^{+}19, WYX^{+}19, WWLJ14, WLL^{+}19, WLLL10, WLX^{+}15, WCD^{+}15, WIZ^{+}17, XSZ^{+}10, XZX^{+}17, XWY^{+}10, XXWY10, XLL11, XLH^{+}15, XDLZ19, YTL^{+}10, YDH17, YN17, YJCQ15, ZWFX17, ZFG^{+}14, ZYQ^{+}14, ZGY15, ZQCZ16, ZWLW16, ZQWL17, ZRS18, ZSSR19, ZWLL12, ZT13, ZH14a, ZYX^{+}10, ZYL^{+}16, ZLL17c, ZWQ^{+}15, ZZLL16, ZWG^{+}16, Zhu14, ZSB^{+}13, ZGKB16, aaGZ19]. Scheme [CSY15, CCLW15, CP17c, GZZ^{+}13, HST^{+}11, HLZY15, HGC12, HLQ^{+}15b, HT16, JJG^{+}12, KWZ^{+}12, KLWK12, KZW17, KMMR13, LC10, LLY^{+}14, LMZG15, LLL^{+}12, MM12, MS12, MS13a, NLY15, RM12, RGBC11, She14, SP15, SHF^{+}17, WJTL12, WZ14, WPMX18, WML14, WXYX14, WHZS19, XWSW16, XJY^{+}10, XLH^{+}15, ZJL^{+}12, ZQH13, ZRQA14, ZSW^{+}19, ZDG^{+}14, ZJ16, ZH18, WZS^{+}18, LLZ^{+}12b]. Schemeof [WWLJ14]. Schemes [ASBL15, FC10, GKL^{+}17, HDL^{+}15, LRW12, LCL^{+}14, LZCK14, MNZ^{+}15, RM11, WXLY16, YRLY16]. Schur [Van14]. Schur-Complement-Based [Van14]. Science [ABE^{+}11, ABN19]. Scientific [APJ^{+}16, CB14, DTLC19, IOY^{+}11, KOPS10, NTWL11, PP12, SCJ^{+}17, WZSL12, WGHP11, ZLK^{+}16, ZHCL17, ZWG^{+}16]. Scores [AI15]. Scratch [MBV11]. Scratchpad [CCC^{+}16, GLGLBM13]. SDCon [SB19]. SEALDB [YTW^{+}19]. Seamless [LdSB19, XWJX15]. Search [BBM16, CWL^{+}14a, CSY16, CZS^{+}16, DT14, DSASSLP12, FRS^{+}16, HAU19, HS12, HJF16, IMH12, JLKG17, KBHS14, LPP13, LCS14, LLW^{+}15, LTC^{+}19, LMFS11, MB12, NGJ^{+}19, PM13, SWC^{+}14, SYL^{+}16, THE^{+}15, WTL10, WCRL12, XWSW16, YQ11, ZJL^{+}17a, ZLW^{+}18]. Search-Based [LPP13]. Searches [GC16]. Searching [RY14]. Seclius [ZBK^{+}15]. Secondary [JZY^{+}15]. Secrecy [HLS^{+}15]. Section [RFZ11]. Sections [RLSK17, ZLJL17]. SEcure [TLRW15, CHCC14, CPM^{+}10, CLH^{+}14, CW19, CCCB14, FLH13, GZX14, HCG^{+}15, Hur13, ITW^{+}14, LLGP13, LAK11, LYZ^{+}13, LLC^{+}15, LT10, LT12, LZWY13, Lou14, LLL^{+}14b, LLL^{+}12, LLS13, LLG14, MS13a, MLS15, SP15, TXL^{+}14, TLL^{+}16, UBC13, WCRL12, WWL^{+}13, WHB16, XWSW16, YJ13, YJR15, YWWR18, ZJ16, ZLW^{+}18, WZS^{+}18]. Securely [CL16a, LHL^{+}14, WRWW13]. Securing [AGG17, BKL11, TKR14]. Security [Ano12c, CLL^{+}14, CZQ^{+}17, GZZ^{+}13, GHZZ16, HXC^{+}11, RM12, RYLZ10, RXD12, SZZF10, WWR^{+}11, Xia14, ZBK^{+}15, LSL14b]. Security-Aware [GHZZ16]. Security-Sensitive [CZQ^{+}17]. Seek [SSLF17]. Seek-Efficient [SSLF17]. Seer [BMJ^{+}17]. Segment [Hu14, XHG15]. Segment-Based [XHG15]. Select [SLL13b]. SelectCast [WJTL12]. Selecting [HAD12, LS17a]. Selection [AWZ15, CL15, LLRP18, LZWZ19, LV17, RZB^{+}18, SHG13, SJ14, TDL^{+}19, TP14, XZT^{+}13, XHZ^{+}13, YL11a, AO12]. Selective [CZQ^{+}17, HWS16a, HWS16b, LSC16, OUA11]. Self [BCTB13, BRX13, CJW16, DHBB12, DA16, DW13a, DLL^{+}11, EHNS13b, IvS10, KE16, LGOB17, Oru17, RLSK17, TVG13, TGT10, YZS13, YC14, YLZ^{+}15b, YZFZ10, ZTG^{+}18, ZS13, ZSY14, ZLDC15]. Self-Adaptation-Based [YZS13]. Self-Adaptive [EHNS13b, ZTG^{+}18]. Self-Calibrating [BCTB13]. Self-Compressive [TVG13]. Self-Configuration [BRX13]. Self-Consistent [TGT10]. Self-Contained [ZS13]. Self-Controllable [ZLDC15]. Self-Disciplinary [YZFZ10]. Self-Downgrade [RLSK17]. Self-Invalidation [RLSK17]. Self-Invalidation/Self-Downgrade [RLSK17]. Self-Management [IvS10]. Self-Monitoring [DLL^{+}11]. Self-Organisation [ZSY14]. Self-Organized [LGOB17]. Self-Organizing [DW13a]. Self-Protection [DHBB12]. Self-Routing [Oru17]. Self-Similar [YLZ^{+}15b].

Self-Similarity [CJW16].

Self-Stabilization [DA16, KE16]. Self-Stabilizing [YC14]. Selfish [Sam14a, ZWZ^{+}15]. Semantic [EADT19, HJZ^{+}12, HjZ^{+}14, HJF16, ZHW^{+}19, CMK^{+}16]. Semantic-Aware [HJZ^{+}12, HjZ^{+}14]. Semantics [ET10, MGS12, RLSK17]. Semantics-Based [ET10]. Semi [CL17, CEK16, KCK14, NZM^{+}16, TWL16, ZML^{+}17]. Semi-Directional-Flooding [KCK14]. Semi-Elastic [NZM^{+}16]. Semi-External [ZML^{+}17]. Semi-Infinite [CEK16]. Semi-Intrusive [TWL16]. Semi-Online [CL17]. Semiconductor [DBG^{+}14]. Semipersistent [LSL^{+}10]. Sense [Amm12, KZW^{+}12, SCC11]. Sensed [MWZ^{+}14]. Sensing [CZZ^{+}16, CIH13, CLHK11, GCN^{+}14, HCC^{+}12, HHK10, JMS^{+}18, Kum14, LCL^{+}14, LCS^{+}15, PM13, WMZ^{+}15, XYT^{+}15, XJ14, XJL^{+}14, YSG^{+}14, ZZG^{+}11, ZYZ^{+}14, ZGL^{+}15, ZMTL15, ZYT^{+}15, ZLLZ13]. Sensitive [CZQ^{+}17, LSWR16, TFLL18, Wan19, XWH15b, XCZ^{+}15]. Sensor [AO12, ALLR14, ACNP11, Amm12, BBCB15, BKY15, BCSKN12, CWL14b, CHCC14, CTX^{+}11, CHTW12, CLLS12, Che14, CYL^{+}14, CYC^{+}15, CCT16, CNC^{+}14, CC15, rCHG10, CIH13, CLHK11, DWLY15, DRSL15, DCL^{+}10, DLL^{+}11, DLZ^{+}14, DOLG16, DWY^{+}13, DRK11, FC10, GBD^{+}13, GFLL15, GLL15, GJLZ12, GJLZ13, GCN^{+}14, GJZZ12, GCZ15, GLC^{+}15, HGY^{+}14, HJY16, HCS12, HL12a, HCL^{+}12, HCC^{+}12, HJPL14, HCG^{+}15, HA10, HWX12, HSX^{+}12, HH12, HHK10, JCLJ12, JLW^{+}10, JJW11, JCW^{+}12, JZW^{+}14, JHW^{+}15, JRP^{+}10, KK10, KXL^{+}14, KZLL14, KSP10, LKE16, LAV^{+}10, LVA^{+}11, LC12a, LMSRSR12, LJG12, LRW12, LWY^{+}13, LLL^{+}13, LCGC14, LHD^{+}14, Li14b, LCLL15, LLG15a, LL11, LRJX13, LWZ^{+}15, LCW11, LZN10, LCL^{+}11, LZNX11, LM12, LWW^{+}13, LDNT13, LJB^{+}13, LHL^{+}13b, LCLD13, LZP^{+}13, LLZ14, LWJ^{+}15, LZK^{+}15, LLH^{+}15a, LCL^{+}16a, LLZ^{+}12b, LLG14, LTMD11, LWZ12, LWG^{+}12, MLL14, MLC^{+}15, MS12, MM15, MTX^{+}11, MLT^{+}13, MV12, MM10, MGR12, PB12, RGRM14]. Sensor [RM11, RM12, RGK15, RZH^{+}11, RHDL11, RZW^{+}13, RCC^{+}14, RWLL14, RQZ^{+}16, SAM14b, SP15, SJAdCL19, SHX^{+}10, SHM^{+}12, TKS11, TXWL11, TLRW15, TWZW11, UBC13, WPT10, WMT^{+}11, WWL11, WMHX12, WFK^{+}12, WJTL12, WWLX13, WFA13, WWX^{+}13, WLL^{+}13, WJTZ14, WHB16, WG13, WWCB14, XWH15b, XHHC13, XJ14, XHG15, XWY^{+}10, XLM^{+}11b, XLM^{+}12b, XLM12a, XHQ^{+}15, XAK17, YLZ^{+}15a, YK14, YSDQ11, YRL11, YLT15, ZJL^{+}12, ZS10, ZZR12, ZMLT13, ZWLL12, ZQH13, ZT13, ZYT^{+}15]. Sensor-Mission [JRP^{+}10]. Sensor-Target [LCL^{+}11, LCLD13]. SensorNets [IvS10]. Sensors [CCT10, ERSR13, WPT10]. Sensory [KPG^{+}12, SGC14]. Sequence [ACS13, IMH12, LMFS11, LPMB13, MC10, Mis14, WKC12]. Sequence-Based [Mis14]. Sequences [dOSdM13]. Sequencing [rCHG10, NTA^{+}16, VPS17]. Sequentially [USP^{+}12]. Serializable [PRR^{+}16]. Serialized [HZG^{+}17]. Series [DBA17]. Series-Oriented [DBA17]. Serve [JCWB10]. Server [CJW16, FWJ18, HJS^{+}11, LZ12, LC15, LY16a, LLL18, NN13, TNZ^{+}12, THB^{+}14, WW11, WWX^{+}13, WW13, WPT17, XXWY10, XZJ^{+}19, YLW13, YZL^{+}17, ZTA^{+}15, ZQL^{+}16, ZT16, ZJLG14, ZJTZ14]. Server-Centric [LY16a]. Servers [DSM14, IZA18, KALK^{+}18, LTGI16, RLY^{+}15, SLL13b, WCF10, WWCZ11, XGL^{+}16]. Service [AWZ15, AIAD^{+}18, ABBCT16, BVEAGVA10, BB13, BDLS13, CSP13, CZYL14, CP15, DMR16, DHTZ15, DWT^{+}16, DT14, DZLC15, FGLP10, HH15, LQY^{+}12, LMZG15, LLS^{+}18, LGL^{+}18a, LLS14, LGZ^{+}19, LS17b, LZNX11, LLG^{+}13, LSW16, LSW17b, MZLT19, MWJ16, MDZC14, PKCB11, PHXL19, PDH10, RHT13, SAEH19, SytL19, SCL^{+}15, SJ14, TJH^{+}14, TCZL11, WSWY15, WM15, WUH^{+}17, WHGS17, WLC^{+}17, WMC^{+}19, XZSG12, XLY^{+}17, XSTZ10, YYK^{+}11b, YZT^{+}17, YJCQ15, ZHZL17, ZJTZ14, MCMR12, DNW^{+}16]. Service-Based [BDLS13, DMR16]. Service-Oriented [LLS14, WLC^{+}17]. Serviceability [MBV11]. Services [ALZ17, BCF13, CCCY16, HMS^{+}18, HHWZ17, HCyW^{+}17, HX10, HKH^{+}10, Hu14, IOY^{+}11, LV15, LSB^{+}18, LSJ^{+}19, LFLW10, MLXG19, NSY^{+}16, PKS14, SZL^{+}12, SBC^{+}10, STMM17, WX11, XH10, XBZ^{+}16, XCZ^{+}15, XLT^{+}14, ZCZ^{+}12, ZWZ^{+}13, ZLZ^{+}16, ZLW^{+}18]. Sessions [GIP^{+}13]. Set [CHD^{+}15, JRAS17, LV17, LLLH19, MM10, OUA11, QP16b, SRB14]. SETI [JKVA11]. Sets [DK17, LKM10, QGPZ13, YTW^{+}19, YC14, YYL^{+}13, ZLN^{+}13]. Setup [FFC17, NSLV16]. SFA [LZY12]. SFC [LHXP18]. SGBR [ANN^{+}13]. SGD [LHHR18]. Shadow [KE16]. Shadow/Puppet [KE16]. Shaped [LWJ^{+}15]. Shard [LTC^{+}19]. Share [RGK15, TVRD17, XZSG12]. Share-Frequent [RGK15]. Share-Nothing [TVRD17]. Shared [ACE^{+}19, ASD^{+}18, BBK17, GBP17, HZW^{+}14, HWL^{+}17b, JSLD19, LAK11, LNXY15, MJK14, RKRK17, SKGC14, SSPG17, TP14, TVCM12, WVT13, WLX^{+}15, YR14, ZML13, Zou14]. Shared-Memory [WLX^{+}15, YR14]. Sharing [CSZ^{+}12, CSSL15, CCT^{+}14, DMR16, GFLL15, HTZY17, Hur13, IRSNF11, IPQ19, IMH12, LYZ^{+}13, LZWY13, LS14, LH16, MFO^{+}13, RG17, Sam14a, She10a, SLLL14, SLW15, SLC15, SL16, SF10, VMB17, WS14, XML^{+}18, ZJS12, ZZSZ18, ZW14, ZJ16, ZHS^{+}19, WZS^{+}18]. Sharing-Aware [RG17]. Shaving [ZMW17]. Shelving [YQH^{+}15]. Sherlock [YSG^{+}14, MLML15]. Shield [PL16]. Shifts [PB12]. Shingled [LWZ^{+}16c]. Ship [LWG^{+}12, WCLK12]. Shipping [XGL^{+}16]. Short [JWJS14, TZT^{+}16, WH16]. Short-Read [WH16]. Shortcut [KKY^{+}14, TFKN17]. Shortest [CCM^{+}17, FMY^{+}18, KBHS14, Lai12, LZB14]. Shortest-Path [LZB14]. Shortest-Span [KBHS14]. Shuffle [GXZ^{+}15, GRCZ17, uRILP17, YQ16]. Shuffle-on-Write [GRCZ17]. shuffled [KLL^{+}17]. Shuffling [NCM^{+}17]. Shut [WJX^{+}14]. Side [GDM^{+}13, NSH15, YQH16]. Sided [LKD10, LYZL18]. Signal [GG10, PRS^{+}11]. Signature [QGPZ13, RY14, WRL15]. Signatures [CLH^{+}14, CD13]. Significance [ZJS12]. Silent [BBGD^{+}17, DC16]. Silicon [WFZ^{+}17, YLJ^{+}17]. SIMD [LWL^{+}19, WM18]. SimEDC [ZHL19]. Similar [YLZ^{+}15b]. Similarity [CJW16, DT14, GC16, HZB^{+}16, JKS13, KGI17, LWY^{+}15, SWC^{+}14, WMGA15]. Similarity-Based [SWC^{+}14]. Simple [Ara11, LCA13]. Simplex [NVBH18]. Simplified [GG11, HWZE10, ZH14b]. SIMT [Nov15]. Simulating [DLM^{+}17, DWH^{+}18, GTM^{+}17, RRM^{+}15]. Simulation [CRWY15, CWZ^{+}15, CPH^{+}18, DBA17, EHM^{+}17, JMZD12, JZW^{+}14, KEGM12, LNMMA15, MT12, MCRC17, OOA^{+}14, PF12, PVQ15, PJAGW14, PGGS19, SY17, SSM^{+}18, Van14, VTSM12, VT19, WLT^{+}12, XVC17, ZWL^{+}16b]. Simulations [AC19, CRS^{+}17, HAY^{+}18, OZMC^{+}16, RBH^{+}14, ZLJ^{+}15b, ZLK^{+}16, ZWL17]. Simulator [CWCS15, SJAdCL19, ZJL^{+}17b, ZHL19]. Simulators [MJM16]. Simultaneously [SAA17]. Single [CCM^{+}17, EKNS17, FSSZ16, JWK^{+}16, SSLF17, XL10, XWH15b, ZLL17a, ZQSY13]. Single-Chip [JWK^{+}16]. Single-Copy [XWH15b]. Single-Hop [ZQSY13]. Single-Path [EKNS17]. Single-Unit [XL10]. Sink [GJLZ13, KK10, RM11]. Sinks [KPG^{+}12, RM12]. SIP [DBAT11, FC11]. SIRE [PRS^{+}11]. Site [CATC11, WYLX13]. Site-Based [CATC11]. Situ [HLK^{+}19, HHK10, VLP16]. Situation [SL16, ZHS^{+}19]. Size [AD19, KTK12, LQK^{+}13, RPYO11, ScFRdS15, WYX13, WGZ16]. Size-Interval [AD19]. Sized [ZS13]. Sizes [BAMJ12, LC14]. Sizing [LNL^{+}19, XALS17]. Skeleton [GIX^{+}12, JLW^{+}10, LJB^{+}13]. Skeleton-Driven [GIX^{+}12]. Skeletonization [AAH15]. Sketch [TP18, YGS^{+}19]. Skew [CYX15]. Skewness [ZZQ18]. Skill [JAJ^{+}19]. Skyline [ZJKQ16]. SLA [GYQW15, PYHY16, TYWL14]. Sleep [GJZZ12, HCY^{+}12, NTKK15]. Slicing [AGJ^{+}16, ZH14a]. Sliding [Lu14, VBC19]. SLO [LSW17b]. SLO-Guaranteed [LSW17b]. Slot [AS16, GRJZ17]. Slotted [WZFG13]. Slow [SRG19, YK14]. Slow-Flooding [YK14]. Small [FHH^{+}15, HAZ^{+}18, HWNS15, IRSNF11, LLLH19, MS13b, ZS13]. Small-Scale [MS13b]. Small-Sized [ZS13]. Smaller [LC14]. SmallTalker [CYZ^{+}13]. Smart [BMR15, BMJ^{+}17, CJZ12, DN19, HPG14, HCL^{+}14, Hur13, LLY^{+}14, LYY16, LLFL15, LH17, LA12, NSH15, PCFP16, WMC^{+}19, YQH16, CJZ12, GPF12, LJ15, LLL^{+}12, MBO15, NTKK15, YJC15, ZJLS12, ZHQ12]. Smart-Card-Based [HCL^{+}14]. Smart-FiWi [NTKK15]. Smart-Home [LJ15]. SmartAssoc [XZT^{+}13]. Smartphone [RSSC15, ZWWF15]. Smartphone-Based [ZWWF15]. Smartphones [TLJ^{+}14, CCD^{+}15, Liu14, YSG^{+}14, ZHZC15]. SmartSLA [XCZ^{+}15]. SMGuard [YBY^{+}18]. Smith [dOSdM13]. Smoothed [RBH^{+}14]. SMP [CL16b]. SMPI [DLM^{+}17]. SMR [YTW^{+}19]. SMT [WKK11]. SnapFiner [CHLY18]. Snapshot [CHLY18, DGFRR18, Ksh10, Tsa13]. Snapshots [GGS10, NX95]. Snoogle [WTL10]. Snooping [KPKH16]. SNP [CPL^{+}18]. SNR [GTS^{+}15]. SNR-Aware [GTS^{+}15]. SOBAS [UBC13]. Social [ANN^{+}13, BS12, CYZ^{+}13, CW15, CSSL15, CP15, CSY16, CJW16, FCD^{+}13, HWJ18, HML^{+}14, HLeS^{+}15, Iye14, JKS13, JZW13, Jia14b, LWY^{+}15, LLS14, LWCG10, LTBN^{+}12, LLL^{+}14a, LHYW15, LSC16, LS17d, LWL^{+}17, MMSS15, NVS16, RKZC14, SLLL14, SLC15, SZWX15, TLSL15, TLL^{+}16, THT^{+}15, WYW13, Wan14, WJWX14, WSL^{+}15, WXTL13, WDOX15, WZZ^{+}13, WJX^{+}14, XAY^{+}14, XWH15a, XGZW14, YGL13, ZLL^{+}15, ZCJ19, SLC15]. Social-Aware [MMSS15, THT^{+}15]. Social-Based [LWCG10]. Social-Efficient [HLeS^{+}15]. Social-P2P [SLC15]. Social-Similarity [LWY^{+}15]. Sociality [QZZ^{+}16, XHZ^{+}13]. Sociality-Aware [XHZ^{+}13]. Socially [KI14]. Socially-Informed [KI14]. SocialTube [SLLL14]. SocioNet [LWCG10]. SoCs [VMB17]. Soft [DLB^{+}19, JHR^{+}14, PFAF16, PP12, TL16]. Soft-Error [JHR^{+}14]. Software [ALS^{+}18, AA12, BSD^{+}18, BBGD^{+}17, BTG^{+}18, CJZ^{+}16, Di 17, FFMR10, HGL^{+}16, MV16b, PB12, SB19, WKL^{+}16, WYY^{+}12]. Software-Defined [BTG^{+}18, SB19]. Solar [LA12]. Solomon [LWCL18]. Solution [Ara11, Che11, DRVC17, Gua14, LCL^{+}11, LXZB15, PFAF16, WRB11, WS14, XBL15, ZX13]. Solutions [HWJ18, JTS^{+}11]. Solve [KAGD16]. Solver [BKH18, MA13, WJB14, YKW^{+}18]. Solvers [AHTD18, GS11b, JMA^{+}18, SOA15, XFL^{+}19]. Solving [JRAS17, LYL16, SS18, YPL13, ZTD19, ZRTL15]. Sort [HWF18]. Sorting [CP17b, HWZE10, KPA13]. Soundness [WZ14]. Source [CCM^{+}17, CL15, LRW12, LLRP18, MS12, RWLL14, RGBC11, XLSR13, XLT^{+}14, UBC13]. SOurce-BAsed [UBC13]. Source-Location [LRW12, MS12]. SPA [TLL^{+}16]. Space [AH10, DB18, GJLZ12, JLKG17, KGI17, KM18, NVBH18, XML^{+}18, YQ16]. Span [KBHS14, ZTZ^{+}18a]. Spanning [CFJ15, YCPC15]. Spark [CLT^{+}17, CZWJ18, GKT^{+}17, LHHR18, LNL^{+}19]. SPARQL [AHSK17]. Sparse [AAA19, ATA18, AA17, AE12, AF19, CRWY15, CLY^{+}19, DFGG13, FGEL14, GWC14, JZWN15, KGK^{+}13, KAA16, LT16, RCK15, SOA15, TTG^{+}15b, TS18, YLW^{+}14, YMG15, YR14, Zha12, ZML^{+}17]. Sparse-Matrix [SOA15]. Spatial [BGHG16, GHL^{+}13, Guo14, LSKZ13, LHR^{+}15, LIWJ15, NZWL14, WMWW19, XTXH13]. Spatial-Temporal [LHR^{+}15]. Spatio [AKP14, WMLJ12]. Spatio-Stochastic [AKP14]. Spatio-Temporal [WMLJ12]. Spatiotemporal [AAA19, HAD12, MM15, XWY^{+}10]. Special [Ano11d, Ano11c, BKK11, CLL^{+}14, MBMC13, PKL^{+}12, PBD^{+}13, RFZ11, Ano12c]. Special-Purpose [PBD^{+}13]. Specialization [MLK15, ZYLC14]. Specific [CDPM18, ITL17, MRH^{+}16]. Specification [DA16, YHC^{+}13]. Specification-Based [DA16]. Spectrum [Guo14, HLY^{+}14, HLeS^{+}15, LCL^{+}14, WS14, XJL^{+}14, ZGL^{+}15]. Spectrums [CZWZ14]. Speculation [AELGE16, SAA18]. Speculative [GRJZ17, KB13, dOSMM^{+}16, XL17, ZL10]. Speed [ARM15, BKF^{+}16, EHWX10, HD15, LCYW16, MSSV18, WBPF11, WL13, WWT^{+}19]. Speed-Up [MSSV18]. Speedup [VPS17, XDLZ19, ZLX^{+}14].

Speedup-Function-Based [XDLZ19]. Sphere [NGJ^{+}19, TXL^{+}14]. Spiking [CHM^{+}13]. Spin [CWS12, CWCS15, DLA^{+}18, SDG17]. Spin-lock [SDG17]. Split [KKK11, LXZH16, MG14]. Split-Star [LXZH16]. SPMs [GZY^{+}15]. SpMV [LYL15]. SPOC [LLS13]. Spontaneous [LLGP13]. Spoofing [YCTC13]. Sporadic [TL16]. Spot [VS19]. Spotting [FGJ^{+}15]. Spread [RXD12, WJX^{+}14]. Spreading [CMPS11]. Sprinting [CCJ19]. Squares [YPL13]. SRAM [KHK15]. SRAM/DRAM [KHK15]. SSA [HCH^{+}12]. SSD [CLLX18, HWS16b, PYHY16]. SSDs [SHA19]. STA [NTKK15]. Stability [DGF12, JMZD12, LWX^{+}11, SSM^{+}18, VM12, VWDM14, ZCX15]. Stability-Optimal [LWX^{+}11]. Stabilization [rCHG10, DA16, DMT12, KE16, YL11b]. Stabilizing [BFPB10, SOK^{+}19, YC14]. Stable [CCV19, hKYY11, SCH11, ZRS18]. Stack [FSSZ16, Man18, PH18, WGZ16, WWH^{+}17, PH18]. Stack-Level [FSSZ16]. Stacked [WCYL19, YTL^{+}19]. Stackelberg [YLC^{+}16]. Stacks [PYH19]. Stage [XHC16]. Staging [IBC^{+}11, MBV13, WVT13]. Staleness [CZL^{+}16]. Stalls [LLD^{+}18, YOK^{+}17]. Standby [FFC17]. Star [AR10, CH14, LXZH16]. STARS [PM13]. Starvation [CRD11, ZQWL17]. State [Bad14, GE12, KKW18, LJL^{+}11, LV17, MKVL12, NCB17, Par18, PVQ15, WKK17, XHX^{+}13, YYY^{+}14]. State-Duration [XHX^{+}13]. State-Machine [KKW18, WKK17]. Stateful [FHW11]. Stateless [MMSS15]. Static [AFT^{+}16, OPM^{+}15, PM13, RJ16]. Static-Dynamic [RJ16]. Stationary [CMPS11, KKC17]. Statistical [CC10, JKVA11, SOTN12]. Statistics [WLX13, ZMA12]. Stay [LLCL12]. Stealing [CGH13, PWJ16]. Steering [WZGR10]. Stencil [BKH18, BBP17, GTM^{+}17, RMG18, SHY14, WTTH17, ZM13]. Stencil-Based [GTM^{+}17]. Step [WHC^{+}14]. Steps [KPA13]. Stepwise [KE16]. Stereotypes [SAH15]. STI [DR16]. STI-BT [DR16]. Still [HCA16]. Stitching [KSP10]. Stochastic [ALZ17, AKP14, BDLS13, Bru14, CMG17, CE10, HMS^{+}18, HCY^{+}12, KEGM12, LZ10, LTL14, MSB11, OPM^{+}15, Seh15, YWJJ11, ZJLS12]. Stock [HMS^{+}18]. Stop [HWC15]. Stops [ACE^{+}19]. Storage [AKGR13, ACNP11, AGG15, AGG17, BH13, CDBQ12, CAJ^{+}16, CL14, CWL16, CHHK19, CW19, CLKR15, CCT^{+}14, DKL^{+}19, FWH^{+}18, Fen14, FSSZ16, GAKR11, GF13, GGF^{+}14, HOZ12, HJY16, HXLF15, HJF16, HLQ^{+}15a, HLQ^{+}15b, KM19, KXC11, KDCR19, LT16, LXXH16, LL17, LLRP18, LQW^{+}18, LTGI16, LT10, LT12, LSW16, LZW^{+}17, LSW17b, LSW17c, LSN19, LVD11, MWJ16, MA14, MCJT19, MV16b, NSH15, PJC^{+}13, PYHY16, Rao14, RLY^{+}15, RTZ^{+}18, SEA18, SSF16a, SSF16b, SSLF17, SLSG18, SLSG19, SPS18, SVK^{+}19, SYZ18, SHF^{+}17, TWT16, WWR^{+}11, WZ14, WPMX18, WXLY16, WMLJ17, WWH^{+}17, WMJ^{+}19, Xia14, XLT^{+}14, XGL^{+}16, YTZ^{+}11, YJ13, YJ14, YPL^{+}17, YYL^{+}13, ZJL^{+}12, ZLL17a, ZMW17, ZHX^{+}19, ZHAY12, ZLX^{+}14]. Store [CSW^{+}17, MZL^{+}19, MCJT19, TGNA^{+}13, YTW^{+}19]. Stored [RSN14]. Stores [AEM17, GYW^{+}19, ZSW^{+}19]. Stranded [YC18]. Strategies [ABLS16, CB13, HV11, HBS^{+}16, LdSS^{+}13, NFD10, RLVTMG^{+}16, SHG13, uRILP17, YR14]. Strategy [BAAT16, EALM15, GF13, LKE16, LWX^{+}11, LLZ18b, MPS15, SS18, Tak14, TYWL14, VPS17, WVM19, WJ12, WL12b, YPL^{+}17]. Strategy-Proof [LLZ18b]. Strategyproof [GLL11, HLeS^{+}15, LC12b]. Stream [BVFGSFAF17, FHW11, JCW^{+}19, LXHS12, LHZJ19, LABQ18, MMdE19, MZK19, ME15a, NCGP19, TG13, TBC12, WYY^{+}12, WWLJ14, WLL^{+}19, ZCJ19]. Stream-Based [TBC12]. StreamCloud [GJPPM^{+}12]. Streaming [ASBL15, BMB^{+}10, CDBQ12, CZWJ18, DWW^{+}15, GG13, Goh14, GJPPM^{+}12, Hu14, HLYJ19, JCWB10, JHWY19, KLWK12, KZW17, LV15, LNL^{+}19, LFLW10, LLZ^{+}12a, LLG^{+}13, OKT^{+}16, SML13, SLL13a, SCCC11, TCDMRP17, VNA^{+}16, WXL10, WSC^{+}14, yWeH11, XSZ^{+}10, XZSG12, XBL15, YGS^{+}19, dSLMM11]. Streaming-Aware [KZW17]. Streamline [BMB^{+}10]. Streams [AB14, BSL^{+}17, GSH^{+}19, LLG15a, Lu14, MTDD17, MP16, SMTZ17, SMB^{+}18, WWL^{+}13, WWT^{+}19, WSSZ13]. Stress [GYLW18]. Stress-Aware [GYLW18]. Strict [KCH19, LZWY14]. Strided [ALI^{+}17]. String [KKK11, LLLC17, MIH17, TVCM12, YP13, ZS17]. Stripe [SSF16b, SLSG19]. Strong [SK14, WZQ10]. Strong-Incentive [WZQ10]. Strongly [HAY^{+}18, TPRH16, ZLS^{+}18]. Structural [CH14, HGY^{+}14, LCS^{+}15, SKA15]. Structure [DWH^{+}18, DO13, HW13, LAFA15, LGW^{+}17, QCZ^{+}15, XDMZ17, ZZF10, ZDM^{+}17]. Structured [HLCH11, HBF12, RCFW10, WPMX18]. Structures [BG13, QFZZ15, VMB17, WL13, ZWJ^{+}18]. STT [AFMM17]. STT-RAM [AFMM17]. Stub [LX10]. Study [AF19, BBCTA18, HAZ^{+}18, JKVA11, LHL^{+}13b, LJL^{+}15, NSLV16, SJVR17, VMN^{+}16, uRILP17, WGHP11, ZLY^{+}14]. Sub [JWJS14]. Sub-Arrays [JWJS14]. Subarrays [QZG^{+}16]. Subject [ZMA12]. Subscribe [JHMV12, MC14, MFO^{+}13, QCZ^{+}15, TKR14, WM15]. Subscription [JWE15]. Subsequences [ACS13, YXSS13]. Subspace [THE^{+}15]. Substrate [APMG12]. Subtrajectory [GV15].

Subtree [RBSS11]. Successive [LWY^{+}13]. Succinct [WL13]. Sufferage [CTA14]. Sufficient [Dua95, NX95, VS11a, VS11b]. Suitability [ECV16]. Sum [KPA13]. Summary [DSASSLP12, SMB^{+}18]. Sums [BAMJ12]. Sunway [CLY^{+}19, HAY^{+}18]. Supercapacitor [ZMW17]. Supercomputer [CLY^{+}19, FBCB18, VTSM12]. Supercomputers [MNZ^{+}15]. Superpeer [LC10]. Superscalar [CA13]. Supply [LQW^{+}18]. Support [APMG12, CGS^{+}15, CSV^{+}17, CAKRY16, DSM19, sFC12, HCH^{+}12, KCH19, KWG17, QTC^{+}14, RMG14, SAA18, SKGC14, VMP17, VMB17, YLSQ13, YDC^{+}17, YWZ17, ZHQ12]. Supported [WMWW19]. Supporting [CWS12, HZJ^{+}11, NSY^{+}16, SMS^{+}13, SWC^{+}14, TL16, XWJX15]. Supports [AELGE16]. SURE [MMNN16]. SURF [KKK^{+}15]. Surface [KZLL14, LWZ12]. Surroundings [NTK^{+}15]. Surveillance [CTX^{+}11, CTX^{+}12, CC15, JGHD10, LZY^{+}19, LCL^{+}11, LCLD13]. Survey [AM19, BMR15, DMCN12, FSM^{+}12, GE12, HRGE17, ICN18, Jia16, KM19, KKC18, LNMMA15, MVL15, MV16a, MV16b, MV16c, MV16d, Mit17, WYLX13, YZS13, YQ11, ZSB^{+}13]. Sustainable [CZD^{+}19, GGF^{+}14]. Sustainably [LHG^{+}17]. Swap [FKMC15]. Swap-and-Randomize [FKMC15]. Swapping [ZLL^{+}17b]. Swarm [WTD17]. Swarming [LTBN^{+}12, ZCX10]. Swarms [CL13, CNMA11]. Sweeping [SS18]. Swift [RTZ^{+}18]. Swift-Like [RTZ^{+}18]. Swiper [CRZH15]. Switch [KOKA11, PD14, QFZZ15, XHC16, YKN^{+}19, ZGY15]. Switch-Centric [QFZZ15]. Switch-Tagged [KOKA11]. Switchable [CIP^{+}17]. Switched [Bis18, MMSS15, SHG11, SSF16b]. Switches [CCLW11, LHM12, SJR17, WYLH18, WYL19]. Switching [YL11a].

Sword [GYX^{+}10, TTJX12]. Sybil [CQZ^{+}12, WMGA15, WXTL13]. SybilDefender [WXTL13]. Symbiosis [HWL^{+}17b]. Symbiotic [FES^{+}17, LABQ18]. Symmetric [BKL11, YKW^{+}18]. Symmetrical [Tsa13]. Symptom [DLC^{+}16]. Sync [LZP^{+}13]. Synchronization [AFA12, BCQ^{+}10, CHCC14, CPM^{+}10, CZL^{+}16, CLSZ12, FWJ18, HTA10, HZG^{+}17, JZW^{+}14, LCLL15, LJL^{+}11, LZP^{+}13, LLK^{+}14, LPZ12, MG18, MJM16, RTZ^{+}18, SDG17, UBC13, WCD^{+}15, XSYY13, XVC17, YK14]. Synchronization-Aware [WCD^{+}15]. Synchronized [WLH^{+}15]. Synchronous [BBR12, BVEAGVA10, CCL13, GG10, JZZ^{+}15, LRT19]. Synchronous/Asynchronous [JZZ^{+}15]. Synthesis [KE16, RAS17, WM18]. Synthetic [CC17]. SyRaFa [CCL13]. System [AZW^{+}19, AKGR13, ACE^{+}19, BBR12, BSM^{+}11, CAC^{+}19, CYZ^{+}13, CSC16, CSS^{+}13, CLT13, CSSL15, CZT^{+}17, CZX^{+}19a, CPH^{+}18, CHPY17, CHLY18, DHBB12, DRRCB18, DW13b, DGG^{+}19, DCL^{+}10, EN12, GETFL14, GJPPM^{+}12, HWZE10, HWS16a, HDL^{+}15, HCZ12, JIP14, JHYK11, KLFD13, KJvR^{+}15, Li14a, LCS14, LYL16, LXXH16, LGJ^{+}17, LWCG10, LT12, LS17c, LWW^{+}13, LS17d, LCZ^{+}19, LWZ^{+}16c, MPM17, MMBdS14, OPM^{+}15, PT15, PHXL19, SRB14, SLW15, SLC15, SVK^{+}19, SZZF10, TJH^{+}14, TYS^{+}12, TWSW17, WSC^{+}14, WMZ^{+}15, WKL^{+}16, WUM10, WZGR10, YYY^{+}14, YQH16, YZHZ17, YXLJ16, ZLGN13, ZQCZ16, ZWL^{+}16b, ZW14, ZMF10, ZLDC15, LSL14b]. System-Level [EN12]. System-on-a-Chip [CLT13]. System-On-Chip [ZMF10]. System-on-Chips [JIP14, TWSW17, WSC^{+}14]. Systematic [CCW^{+}12, FPRG16, LC14, LS19]. Systematical [XSZ^{+}10]. Systemic [JRV^{+}13]. Systems [AM19, ALS^{+}18, AAB^{+}17, ACCP12, AGG15, Ano11d, Ano11c, ASYK^{+}19, AGJ^{+}16, BJC^{+}18, BGHG16, BG13, BCQ^{+}10, BJ13, BH13, BBCTA18, BDLS13, Bru14, BXXC12, CMVB17, CLL^{+}14, CL16a, CDBQ12, CCM^{+}17, CADK19, CCT10, Che11, CTX^{+}12, CSP13, CCL13, CLHW13, CWL16, CLYR16, Che16, CCH^{+}17, Che18b, CHHK19, CLL^{+}19, CCS^{+}12, CLKR15, CRC^{+}17, CMG^{+}14, DCC^{+}19, DBA17, DGFRR18, DLC^{+}16, ET10, EADT19, EK10, FWH^{+}18, FWJ18, FHH^{+}15, FSSZ16, GG10, GGY^{+}19, GGS10, GFS^{+}10, GAKR11, GD16, Goh14, GZY^{+}15, GHZZ16, HZW^{+}14, HLZY15, HTZY17, HWG^{+}19, HAZ17, HP14, HWS16b, HWL^{+}17a, HLCH11, HCSC13, HK18, HBF12, HZJ^{+}11, HJZ^{+}12, HjZ^{+}14, HXLF15, HJF16, HXC^{+}11, HCL^{+}14, HWNS15, HT16, HN11, HKkY^{+}16, IBC^{+}11, IdM12, IRPvdS12, JMZD12, JKVA11, JGJF18, JZW13, JGZZ14, JZZ^{+}15, Jia16, JSC^{+}17, JMS^{+}18, Jun17, KWZ^{+}12, KZW^{+}12, KM10, KM19, KMM12, KKC17]. Systems [KXC11, KKK11, KPKH16, KTK11, KI14, Ksh10, Kum14, LZL10, LW11, LZ11, LAK11, Lee17, LWX^{+}11, LQY^{+}12, LTL14, LTW^{+}14, LL17, LHL17, LS17a, LLS^{+}18, LLAL18, LQW^{+}18, LWSM19, LCSC12, LY16b, LKT11, LHJ12, LHW11, LGX^{+}11, LLZ^{+}12a, LNZ^{+}13, LLM^{+}14, LXZB15, LCYW16, LZW^{+}17, LH17, LSW17c, LYPL19, LABQ18, LM16, MWJ16, MB13, MWZ^{+}13, MV12, MV16a, MV16b, MV16d, MCRC17, NLC12, NN13, NLGQ14, PHGR17, PFAF16, PKL^{+}12, Par19b, PF12, PG16, PDH10, PH12, PWT^{+}17, PJAGW14, PGGS19, QNLN11, QLNN13, QCZ^{+}15, QF14, QGZP17, RSR11, RS10, RSW^{+}17, RDG12, RGPH15, RTZ^{+}18, SAEH19, SEAH16, SEA18, ST10, SS12, SLY^{+}14, She10a, She10b, SL13, SK14, SLGW14, SLSL16, SSF16a, SSF16b, SSLF17, SLSG18, SGC14, SME10, SPB^{+}10, STMM17, SF10, SHF^{+}17, SDL^{+}15, TLH^{+}14, TWT16, TWW^{+}18, TNH^{+}18, TNLM17, TKR14, TFM^{+}16, TL16, THT^{+}15, Tsa13, TAZ^{+}19]. Systems [UDH^{+}17, Van14, VS15, WLT^{+}12, WRWW13, WLL15a, WPMX18, WL12b, WMLJ12, WW12, WDC12, WML14, WYCZ14, WXLY16, WMLJ17, WDL^{+}17, WZL^{+}19, WHZS19, WMJ^{+}19, XL10, XHL^{+}15, XZX^{+}17, XHL^{+}11, XAYM14, XLH^{+}15, YQZC12, YQH^{+}15, YRLY16, YLJ^{+}17, YDH17, YN17, YZH^{+}19, YLR12, ZTD19, ZGL10, ZL11, ZYL^{+}17, ZLL17a, ZCJ19, ZLK^{+}16, ZLX^{+}14, ZD16b, ZH18, dSLMM11, LRYJ17, Ano12c, Ano15a, Ano16, Ano17a, Ano18, Ano19a]. Systems-on-Chip [YLJ^{+}17]. SZ [TDL^{+}19].

T [CZX^{+}19a]. T-Gaming [CZX^{+}19a]. TA-Update [WPMX18]. Table [KKY^{+}14, MMACS10, SX10]. Tables [KHK15, RRS12, SYZ18]. Tackling [ZJS^{+}17]. Tag [BXXC12, ESGQ^{+}13, LZC^{+}12, LLM^{+}14, LXZB15, WZFG13, WXYX14, ZZG^{+}11]. Tag-Based [ESGQ^{+}13]. Tag-Free [ZZG^{+}11]. Tagged [AKC^{+}15, KOKA11]. Tags [SLY^{+}14, ZCX^{+}14]. TaihuLight [CLY^{+}19, HAY^{+}18]. Tail [ASSB18, HHWZ17, QPB^{+}17]. TAMES [CZWZ14]. Target [CC15, LWZ^{+}15, LCL^{+}11, LCLD13, WWCB14]. Targeted [PWT^{+}17]. Targeting [TTG^{+}15a, TFKN17]. Targets [GJLZ12]. TASA [ZZG^{+}11]. Task [ABE^{+}11, AAB^{+}17, AD19, CTA14, CL17, CCKF15, CLT13, Che16, CCC^{+}16, CRG^{+}17, CZL^{+}18, CCK12, CRC^{+}17, ELX^{+}11, GZY^{+}15, GFJT19, GLC^{+}15, GHW^{+}16, HAZ17, HxjGG19, HMP^{+}19, HLL18, HYX11, JZW13, Jia16, JJG^{+}12, JHWY19, KHN16, Kao15, KMM13b, LTL14, Li14b, LZL^{+}18, LGX^{+}11, MWZ^{+}14, MSSV18, NLGQ14, RFZ11, ScFRdS15, TBÁ^{+}19, TL16, VS15, WZQY14, WSC^{+}14, WZL^{+}16, WW12, XLL11, XLH^{+}15, XLY^{+}17, YKW^{+}18, YYK^{+}11b, YN17, ZYW^{+}16, ZYX^{+}10, ZJTZ14, LYZL18]. Task-Based [AAB^{+}17, GFJT19, TBÁ^{+}19]. Task-Duplication [HMP^{+}19]. Task-Graph [MSSV18]. Task-Level [WZL^{+}16]. Task-Size [ScFRdS15]. Task-Tree [MWZ^{+}14]. Tasking [SAB^{+}18, TCM18]. Tasks [BHKS^{+}17, CB14, CC13b, CZQ^{+}17, Che18a, CLL^{+}17, DDP^{+}19, DLA^{+}18, DLB^{+}19, HGS^{+}19, HxjGG19, IOY^{+}11, JAJ^{+}19, Lee12, LW15, OPM^{+}15, PR19, PVS18, SAF16, VS19, WZQY14, ZGL10, ZWQ^{+}15, ZLYL19, ZJTZ14]. Taxicab [ZHL^{+}15]. Taxonomy [HPG14, LM16, SJVR19]. TC [YCMX17]. TC-Release [YCMX17]. TCAMs [LG10]. TCP [ZRTL15]. TDOA [XSYY13, LZZP13]. TDOA-Based [XSYY13]. Technique [AFMM17, EHI11, ESGQ^{+}13, GG13, KCK14, LMAS17, SMTZ17, SAF16]. Techniques [AM19, AHTD18, BBP17, CATC11, CRC^{+}17, Di 17, DRSL15, KKC18, LZH^{+}16, LPMB13, LJL^{+}15, LNMMA15, LWL^{+}19, Man16, MT12, ME15b, MV16b, MV16c, MV16d, Mit17, SMS^{+}13, TFM^{+}16, TMJ14, XHL^{+}11, ZSB^{+}13]. Technologies [EGQ11, NML^{+}14]. Technology [MJK14, PG16, XZH14]. Tele [VMN^{+}16]. Tele-Immersive [VMN^{+}16]. Temperature [BBCB15, CCLW15, Che18b, SWRQ18, SAF16, XFL15, aaGZ19]. Temperature-Aware [BBCB15, Che18b]. Temperature-Constrained [SWRQ18]. Temporal [BGHG16, LYW^{+}12, LHR^{+}15, TWW^{+}15, Wan14, WMLJ12, XTXH13]. Temporality [ERG^{+}17]. Temporality-Aware [ERG^{+}17]. Tenancy [DY17]. Tenant [LSW16, LH16, LCZ^{+}19, RM17]. Tenants [SL16]. Tensor [ATA18, AHJ^{+}11]. Terabits [KAV^{+}17]. Term [HSX^{+}12, TNH^{+}18, WGCG18]. Terminal [WWH13]. Terrain [SA11]. Terrains [LM12]. Terrestrial [LZZP13]. TerrierTail [ASSB18]. Test [LPSS19, NHN17, NHN18, TTJX12]. Testbed [CZRB18]. Testing [XSTZ10]. Text [CJL^{+}12, SWC^{+}14]. Their [LHJ12, QLC14, RCM16]. Them [WJX^{+}14]. Theorem [ZYW^{+}16]. Theoretic [KP12, Tak14, TKP12, US16, YC14, ZKSY14]. Theoretical [HLYJ19]. Theory [CL14, LZB14, LGX^{+}11, PDH10, SHG11, TCDMRP17, ZASA10, Dua93]. Theory-Based [TCDMRP17]. Thermal [ASYK^{+}19, BCTB13, CAJ^{+}16, CCLW15, Che18b, GGF^{+}14, IZA18, LWSM19, YGL^{+}15, YTL^{+}19, ZYX^{+}10]. Thermal-Aware [CAJ^{+}16, LWSM19, ZYX^{+}10]. Thermal-Delay-Aware [Che18b]. Thin [KEGM12, LS17b]. Thin-Client [LS17b]. Things [NLY15]. Think [HCA16]. Thinning [WQZ^{+}15]. ThinRAID [WQZ^{+}15]. Third [CRZH15, KDCR19]. Third-Party [CRZH15]. Thousands [Sib12]. Thrashing [KZW17]. Thrashing-Resistant [KZW17]. Thread [AELGE16, BTL^{+}19, DCA^{+}16, LSL^{+}14a, NVBH18, RCV^{+}13, SAA18, SSPG17]. Thread-Level [AELGE16, SAA18]. Threaded [JY15, SV19]. Threading [KEGM12, LKBK11, SAB^{+}18]. Three [HXC^{+}11, RM12, XHC16]. Three-Factor [HXC^{+}11]. Three-Stage [XHC16]. Three-Tier [RM12]. Threshold [GC16, LXXH16, LLFL15, SJR17, WGZ16]. Thresholds [BBCTA18]. ThriftStore [GAKR11]. Throttle [CCLW15]. Throttle-Based [CCLW15]. Throttled [CLHW13]. Throughput [BSL^{+}17, CLM^{+}15, CP17b, CWJS11, FQWL12, GFMR13, GBP17, GRB^{+}19, HMKG19, HPH^{+}12, JZY^{+}15, KHK15, LJ16, Li14c, LY11, MB12, RQZ^{+}16, VWDM14, WJ12, WZQ10, XZT^{+}13, YYK^{+}11b, ZGXJ14, ZH14a, ZKP^{+}19].

Throughput-Optimal [CLM^{+}15]. THz [GRUMG17]. Tie [XGZW14]. Tier [ALZ17, LH15, RM12, WCZ^{+}19a]. Tiered [HWL^{+}17a, LYPL19]. TIGER [CAJ^{+}16]. Tightly [ASD^{+}18]. Tiled [DK17]. Tiling [BBP17, RMG18]. Time [AWZ15, ACCP12, APCH^{+}11, AOW^{+}12, AH10, BJC^{+}18, BVEAGVA10, BVFGSFAF17, BCP^{+}14, BMB^{+}10, CHCC14, CCKF15, CLT13, CCL13, CCT16, CCC^{+}16, DBA17, DRRCB18, DLA^{+}18, DCL^{+}10, DLB^{+}19, ELX^{+}11, FYH^{+}15, FFMR10, GRUMG17, GJLZ12, GLC^{+}15, HZW^{+}14, HLZY15, HWG^{+}19, HGS^{+}19, HAZ17, HxjGG19, HRGE17, HKH^{+}10, HLL18, HJF16, IIKO13, KM10, KCH19, Kum14, LZ12, Lee12, Lee17, LTW^{+}14, LCLL15, LSWR16, LWC^{+}17, LGM^{+}17, LLY^{+}17, LHZJ19, LZP^{+}13, LLL18, MHL^{+}16, MZK19, MB13, NZWL14, PCFP16, PHGR17, PFAF16, PVS18, PM13, QF14, RMG18, RGPH15, SFL^{+}14, SEAH16, SS12, SL14, SHX^{+}10, SP12, SFA^{+}17, TXWL11, TL16, TVRD17, VBC19, VLP16, WWCZ11, WVM19, WX11, WYC^{+}15, XWH15b, XZX^{+}17, XSYY13, YLZ^{+}15a, YRLY16, YHS^{+}14, YQH16, YK14, YC12, ZGL10, ZLGN13, ZTH17, ZYL^{+}17, ZML13, ZMF10, ZLLD18, ZLF^{+}11, ZWQ^{+}15, ZWG^{+}16, ZWL17, ZJTZ14, aaGZ19]. Time- [HLL18]. Time-Based [FFMR10]. Time-Bounded [LLY^{+}17]. Time-Constrained [MHL^{+}16]. Time-Constraints [TVRD17]. Time-Dependent [AOW^{+}12]. Time-Evolving [LHZJ19]. Time-Partitioned [PHGR17]. Time-Reversibility [Lee17]. Time-Sensitive [LSWR16, XWH15b]. Time-Varying [LLL18]. Timely [MBV11, MBV13, PDFJ13]. Times [BCP^{+}14, HV11]. Timestamp [YCMX17]. Timestamp-Based [YCMX17]. Timing [Bis18, HST^{+}11, JSC^{+}17, NLGQ14]. Timing-Based [HST^{+}11]. TLB [ERG^{+}17].

TLB-Based [ERG^{+}17]. TLBs [ERG^{+}17]. TLIA [LWZ^{+}16a]. TLP [ZWE19]. TLP-Resource [ZWE19]. TMACS [LXXH16]. TMC [JZWN15]. Token [CRD11, ERRG18, SG16a]. Token-Based [ERRG18, SG16a]. TokenTLB [ERRG18]. Tolerance [AP17, BG13, CYW^{+}18, HWC15, MNZ^{+}15, STK^{+}19, Wan19]. Tolerant [ANN^{+}13, ASYK^{+}19, BKY15, BGE^{+}16, CSY16, CCH^{+}17, CH15, CCCB14, CLSZ12, DW13b, EHNS13a, FYH^{+}15, GLJ^{+}15, GLC^{+}15, JHYK11, KZK^{+}19, KZK^{+}20, LHF^{+}15, LW12, MR16, NTK^{+}15, PLZW14, RS12, SAEH19, SBC^{+}10, TZY^{+}18, WYW13, WGG^{+}18, WYL19, XGZW14, YHS^{+}14, YDH17, YCW12, ZJL^{+}12, ZGH14, ZCX^{+}14, ZDG^{+}14, ZWQ^{+}15, ZWG^{+}16]. Tomography [BKF^{+}16]. Too [XLL^{+}18]. Tool [GWC14]. Tools [DMCN12]. Top [DGFRR18, JCW^{+}12, SKP12, ZYLC14]. Top- [JCW^{+}12]. Top-Down [SKP12, ZYLC14]. Topological [GCZ15, TCT14]. Topologies [BGE^{+}16, BS14, CMV^{+}10, CMB15, CMVB17, HS12, TFKN17]. Topology [BKY15, CLHW13, DWW^{+}11, DWF12, EMTX15, FB10, FSM^{+}12, GLJ^{+}15, HLY10, HWNS15, HT16, JJ11, LZN10, LLZ14, LGXL19, PFMR13, RHT13, SKP12, TL14, TDLR13, ZZF10, ZHCW12, ZD16b, Zou14]. Topology-Agnostic [FSM^{+}12]. Topology-Aware [CLHW13, Zou14]. Tor [LLY^{+}15]. Torrent [WL12a]. Torus [CMV^{+}10, DDP^{+}19, JP12, LX12, Tou15a, YLJ^{+}17, ZPD11, ZD12, ZDF^{+}15]. Torus-Like [YLJ^{+}17]. Total [LSWR16, LGJ^{+}18]. TPDS [Ano11d, Ano11c, Par19b]. TPUs [YZH^{+}19]. Trace [CC13a, EHM^{+}17]. Trace-Driven [EHM^{+}17]. Traceback [dOSMM^{+}16, YZDJ11]. Traces [CC17, DD17, WDH^{+}16, ZSH^{+}11]. Tracing [GD16, SZL^{+}12, WSSZ13]. Tracking [BN12, DL17, DRK11, HJY16, HH12, LHF^{+}15, MS13b, SLY^{+}14, WSSZ13, WWCB14, ZLGN13]. TRACON [HC14]. Trade [GAKR11, SPS18, TFKN17, WBPF11, ZYZC12]. Trade-Off [TFKN17, WBPF11, SPS18]. Trade-Offs [GAKR11, ZYZC12]. Tradeoff [CFLL18, Jia14a, LWY^{+}13, NL11]. Tradeoffs [IB14, LWLZ17, MLVD12, TFM^{+}16, WKL^{+}16]. Traffic [CAD^{+}18, CHLC15, CL15, DN19, FXL17, GKL^{+}17, HN10, IB14, JGG^{+}11, KK10, LZ10, LGM^{+}17, LLY^{+}17, LLRP18, LX10, MTMR18, NFFK14, RHDL11, SYL^{+}14, SCHT16, TLP15, TZC19, TP13, VT19, WWL11, WXZ^{+}14, WWZ^{+}16, WMLJ12, WZLC15, WYC^{+}15, XHX^{+}13, XLLZ11, XSL^{+}16, XVC17, YZSC14, YSS^{+}17, ZXW^{+}13, ZT13, ZFG^{+}10, ZLF^{+}11, ZLLZ13, ZFF16]. Traffic-Aware [LGM^{+}17, MTMR18, RHDL11, TLP15, WWL11]. Training [CLB^{+}19, CCHH19, CSR^{+}17, DSM19, LHHR18, VMP17, WSH^{+}19, YZH^{+}19]. Trajectories [JZWN15]. Trajectory [ACC^{+}17, GC16, JGG^{+}11, JZH^{+}14, LWZ14, LZC^{+}12, WSS15, ZYW^{+}14a]. Trajectory-Based [JGG^{+}11, JZH^{+}14]. Transactional [ASG^{+}14, AA12, CSW^{+}12, CWLS19, CD13, CRRR15, DD11, Di 17, DR16, FFMR10, GIX^{+}12, HPPR17, KKW18, KWG17, QGPZ13, QGZP17, SAA18, TGNA^{+}13, TGAG13, dCAB19]. Transactions [Ano11d, Ano11c, Ano15a, Ano16, Ano17a, Ano18, Ano19a, ITW^{+}14, Par19b, TPRH16, ZCZ^{+}12, Ano12j]. Transceiver [NML^{+}14, ZLGN13]. Transceiver-Free [NML^{+}14, ZLGN13]. Transcoding [LSB^{+}18, LSJ^{+}19]. Transfer [BZBP10, DCW^{+}15, EHWX10, LRYJ17, LC14, RS10]. Transfers [GXZ^{+}15, Guo17, KAV^{+}17, XLSR13, YYK11a]. Transform [LJB^{+}13, MVC^{+}18, QJ16, WH16, XAK17]. Transform-Based [LJB^{+}13]. Transformation [SLG10].

Transformations [VGMA10]. Transforming [LVA^{+}11]. Transient [FPGAD10, JMZD12, SSM^{+}18]. Transit [SYL^{+}14]. Transition [KKC17, LHL17]. Transitive [ADMX^{+}12]. Translation [LZW^{+}17, WX15]. Transmission [LZNX11, LLG14, RPYO11, SA11, WPMX18, XJ14, Zhu14]. Transmission-Efficient [XJ14]. Transmit [ZQSY13]. Transmit-Only [ZQSY13]. Transparent [JHYK11, TS16]. Transport [DOLG16, TW14, WDC12, YWZ17]. Transport-Friendly [WDC12]. Transport-Support [YWZ17]. Transportation [PT15]. Transpose [KAA16]. Traversal [LLN^{+}19]. Tree [APMG12, AP17, ABP17, DY16, HPH^{+}12, KKY^{+}14, KBHS14, LLW^{+}15, MWZ^{+}14, MCJT19, MYPL18, MMSAZ11, NGJ^{+}19, QCZ^{+}15, SS17, WXL10, WPMX18, WSH^{+}19, WZFG13, XLM^{+}12b, ZLL17a, YTW^{+}19]. Tree-Based [XLM^{+}12b]. Tree-Grafting [ABP17]. Tree-Mesh [WXL10]. Tree-Search [KBHS14]. Tree-Structured [WPMX18]. Trees [BFPB10, CFJ15, GCG^{+}18, GFJT19, HAU19, HJPL14, LLN^{+}19, LY14, TKS11, Trä19, XP12, YCPC15, ZCX15]. Trends [UDH^{+}17]. Triangle [BF17]. Triangular [RDG12]. Triangulations [SZ12]. Tridiagonal [GS11b, LYL16]. Trie [Hsi14]. Triggered [LWZ^{+}16a, LLD^{+}18]. Triggered-Execution [LLD^{+}18]. Triggered-Issuance [LLD^{+}18]. Triggered-Long-Instructions [LWZ^{+}16a]. Trilateration [YL10]. TripleID [CC18]. TripleID-Q [CC18]. True [XL10]. Truly [SLL13b]. Trust [Ano12c, BH13, BKL11, CLL^{+}14, CDS15, CCCB14, DCA19, FLLS17, HML^{+}14, JHW^{+}15, LZY12, LMZG15, NSY^{+}16, OMMZ14, SAH15, WMGA15, ZDG^{+}14]. Trusted [NFFK14]. Trustworthy [LLS14, LS14, PKG14, SLGW14, ZCZ^{+}12].

Truth [OKT^{+}16]. Truthful

[CZWZ14, FPF13, Guo14, NMG15]. Tsumiki [CZRB18]. Tucker [OPJ^{+}19]. Tuning [BYZ^{+}16, CRG^{+}17, CCW^{+}12, GLRT18, HLZ^{+}19, KAGD16, LMD16, LCY^{+}17, ZJLG14]. Tuple [MJM16]. Turn [FC18]. Turns [LKM10]. Twisted [CMV^{+}10, JP12]. Two [BMJ^{+}17, CL13, CBF^{+}17, DRVC17, FYH^{+}15, LKD10, LYZL18, SEAH16, SMB^{+}18, Sib12, Tse13, YHS^{+}14, YLW13, ZGXJ14, ZYLC14]. Two-Dimensional [SMB^{+}18, Sib12]. Two-Level [BMJ^{+}17, DRVC17]. Two-Phase [CBF^{+}17, SEAH16, ZYLC14]. Two-Server [YLW13]. Two-Sided [LKD10, LYZL18]. Two-Time-Scale [YHS^{+}14]. Two-Way [ZGXJ14]. TXOP [MRM12]. Typed [HGS^{+}19]. Types [PR19].

Ubiquitous [LLL^{+}13]. UHF [KWZ^{+}12, KZW^{+}12]. Ultra [FBCB18, HjZ^{+}14, LWL^{+}19, PSMD18]. Ultra-Dense [PSMD18]. Ultra-Fast [LWL^{+}19]. Ultra-Green [FBCB18]. Ultra-Large-Scale [HjZ^{+}14]. Ultralarge [HZJ^{+}11]. Ultralarge-Scale [HZJ^{+}11]. Ultrasound [BKF^{+}16, RLVTMG^{+}16]. UltraWideBand [HKH^{+}10]. Unbalanced [JHR15]. Unbounded [DMT12]. Uncertain [CYC^{+}16, Guo17, WSS15]. Uncertainty [ELX^{+}11, VLRP15]. Uncertainty-Aware [VLRP15]. Uncoordinated [YWC11]. Undependable [JZW13]. Underlay [KXC11]. Understanding [JZW^{+}17, Jia14b, LLLG13, Li14b, LODB17, LXBZ13, YL11b, ZZH^{+}17]. Underwater [LZZP13, LZP^{+}13, LLZ14, XLM^{+}12b, XLM12a, YQ11]. Undo [WUM10]. Unicast [KKW15, WWL^{+}13]. Unicorn [BBK17]. Unidentifiable [QLC13]. Unidirectional [MKOK14]. Unified [ALS^{+}18, AFT^{+}16]. Uniform [DGI^{+}19, TL16, WFA13, ZR18]. Unifying [MG18, YCWL14]. Union [CMC^{+}15]. Unit [JSC^{+}17, XL10]. Units [CCHH19, DFGG13, KKC18, LLLC17, TRT19]. Universal [KKW15]. Unknown [JRAS17, LLM^{+}14, LXZB15]. Unleashing [BFD19, TCM18]. Unnecessary [LZH^{+}16]. Unraveling [GGGA18, ZDWR11]. Unstable [SK14]. Unstructured [AK18, CJL^{+}12, CE10, GS11a, HLY10, HS12, LMPR12, LWCG10, LHW11, PFMR13, TJLL12, YCWL14]. Unsupervised [MWZ^{+}13]. UnSync [JHR^{+}14]. UnSync-CMP [JHR^{+}14]. Up* [RGBC11]. Up*/Down* [RGBC11]. Up/Out [LSLD17]. UPC [FA19]. Updatable [QP16c]. Update [DWH^{+}18, DMS^{+}12, HYZ15, KKW18, PRR^{+}16, TZ10, WPMX18, WKK17, YJR15, WPMX18]. Update-Efficient [DWH^{+}18]. Update-Intensive [HYZ15]. Update-Serializable [PRR^{+}16]. Updates [CPM^{+}10, Hsi14, LCZ^{+}19, Rao14]. Updating [CJZ^{+}16, KPA13]. Upgrade [GBFS16]. Upgrading [YMHL16]. Uplink [TKP12]. upon [TXL^{+}14, Tse13]. Upper [Che11, Fre13, ZLN^{+}13]. Urban [ACC^{+}17, CQZ^{+}12, LWZ14, ZLF^{+}11]. Usage [ERRG18, LLLZ16]. Use [TNH^{+}18]. User [CSZ^{+}12, CLZ^{+}18, DMS^{+}12, FLH13, JRV^{+}13, JHYK11, LJG12, LZY^{+}18, MS13b, SZZF10, ZQCZ16]. User-Level [DMS^{+}12, JRV^{+}13, ZQCZ16]. User-Transparent [JHYK11]. Users [JZY^{+}15, LLL^{+}13, LYZL18, RSSC15, ST10]. Using [ANN^{+}13, ABE^{+}11, ANE12, AKC^{+}15, AHJ^{+}11, AH10, ARM15, BN12, BG13, BTL^{+}19, BRX13, CL13, CC10, CSW^{+}17, CH14, CC18, CZL^{+}16, CC17, CBF^{+}17, CIP^{+}17, CMK^{+}16, CJW^{+}19, CEK16, DCC^{+}19, DSASSLP12, DIAR16, DRK11, EMTX15, GD16, GIP^{+}13, GV15, GF13, GHL14, HAU19, HWSX17, HLCB^{+}17, HJF16, IMH12, JWA10, JRAS17, JZW^{+}14, KMM12, KKK11, KCYM10, KPA13, KAC^{+}15, LLCH12, LYZ^{+}13, LGYV14, LAT^{+}15, LLW^{+}15, LYL15, LSB^{+}18, LZL^{+}18, LSJ^{+}19, LZC^{+}12, LCS^{+}15, LAFA15, MMNN16, MM15, MC14, MFO^{+}13, MNZ^{+}15, MM10, MV16b, MSB11, OOA^{+}14, PJC^{+}13, PH11, PD14, PWT^{+}17, PP12, QJ16, RX11, RZW^{+}13, RGBC11, SAA18, SHA19, dOSdM13, SMS^{+}13, SA11, SYZ18, TLJ^{+}14, TKR14, TP13, TAZ^{+}19, Van14, VWDM14, WXZ^{+}14, WSWY15, WWCB14, WHC^{+}14, XTFC17, XH10, XSC13, XJ14, XSL^{+}16, YKW^{+}18, YW10, YDH17, YSDQ11, YQ11]. Using [YZDJ11, YZJ^{+}12, ZJLS12, ZGXJ14, ZZG^{+}11, ZXW^{+}13, ZFG^{+}14, ZYLC14, ZLL^{+}15, ZJKQ16, ZWL^{+}16a, ZQWL17, ZWJ^{+}18, ZWJ^{+}19, ZWLL12, ZYW^{+}16, ZZQ18, ZLY^{+}14, ZLW^{+}19, ZYSH14, vdLJR11, GLRT18, GRB^{+}19, LWL^{+}19, NML^{+}14, SY17, TFM^{+}16]. Utility [BMJ^{+}17, KM10, LSWR16, LGZ^{+}19, XWH15b]. Utility-Based [LGZ^{+}19, XWH15b]. Utilization [CYX^{+}14, CTX^{+}12, CCL13, CD13, HZW^{+}14, HTZY17, HWG^{+}19, NZWL14, TL16, TP13, WKK11, WWLJ14]. Utilization-Based [WKK11]. Utilize [LZWY14]. Utilizing [WX15]. UVM [NSLV16]. UWB [HKH^{+}10, PRS^{+}11].

Validating [QPB^{+}17]. Validation

[SBC^{+}19]. Value

[CSW^{+}17, CLZ^{+}18, GYW^{+}19, HK18, LSWR16, MZL^{+}19, MCJT19]. VANET [RPYO11, YXG12]. VANETs [LLLG13, LLG15b, SCCC11, ZLF^{+}11, CCS^{+}12]. VarCatcher [ZJS^{+}17]. Variability [CAC^{+}19, FBCB18, TCYF16, XLY^{+}17, ZJS^{+}17]. Variable [MRM12, XHX^{+}13, YPL^{+}17]. Variables [HZG^{+}17]. Variation [TAZ^{+}19]. Variations [DD17, YZDJ11]. Various [ZDF^{+}15]. Varying [LLL18]. Vector [AAA19, DSM19, FVLD16, sFC12, GWC14, KGK^{+}13, KAA16, RCK15, SOA15, TLP12, TTG^{+}15b, VMP17, YNKD18, YDC^{+}17, YR14, Zha12]. Vectorization [HWF18]. Vehicle [WPT17]. Vehicles [TLJ^{+}14, YLH^{+}16, YQ11, ZS13, ZLLZ13]. Vehicular [CQZ^{+}12, DMR16, GZX14, JGG^{+}11, JZH^{+}14, JZWN15, LQY^{+}12, LWZ14, MV12, QZZ^{+}16, SZZF10, XLM^{+}11a, XBL15, YOWA14, ZY13]. Velocimetry [MLK15]. Velocity [DRRCB18]. Verifiable [LXXH16, Rao14, SWC^{+}14, SYL^{+}16, XWLJ16, YJR15, ZLW^{+}18, XWS17]. Verification [CCT10, CLC^{+}12, WG13, XAG17, ZHAY12]. Verifiers [XAG17]. Verifying [OMMZ14, WDH^{+}16]. Versatile [LY16a, XL13]. Version [ZLZ^{+}17]. versus [KEGM12, LZZP13, NSLV16, SVC12, WFA13, WFZ^{+}17, WS18]. Vertex [AHSK17, LRT19, LCD^{+}17, Ozd19, YHL^{+}18]. Vertex-Centric [AHSK17, Ozd19, YHL^{+}18]. Vertical [KKK^{+}15, MM12]. Very [EHM^{+}17, HAZ^{+}18]. vGASA [ZYQ^{+}14]. via [AAH15, ABP17, CJZ12, CB16, CGZQ13, CZYL14, Che18a, CRRR15, HLS^{+}15, HWS16a, LAdS^{+}15, LPP13, LJL^{+}11, LLLH19, LA12, MIH17, PT11, TWW^{+}18, TYG^{+}14, THE^{+}15, TKP12, WNLL15, WLH^{+}15, WKW16, WHGS17, WPT17, WS14, WML14, WHZS19, XWJX15, XLY^{+}17, YWWR18, ZRQA14, ZLT^{+}18, ZSW^{+}19, ZHZL17]. Video [BSD^{+}18, HW13, JN16, LSB^{+}18, LSJ^{+}19, SLLL14, SCCC11, TCS13, WXL10, WSWY15, XBZ^{+}16, XBL15, ZLCZ14]. View [LSW17c, Tan12, ZLCZ14]. Viewpoint [LPSS19]. Views [Hen14]. Vindication [LNA^{+}13]. VINEA [EMW16]. Virtual [ASSB18, BWH^{+}19, BB13, BZA10, BRX13, CWS12, CSS^{+}13, CL16b, CRZH15, CHPY17, CHLY18, DWX14, DSM14, DWY^{+}13, DY16, EMW16, GDM^{+}13, GLBJ18, HLK^{+}19, HPP15, Ian14, JGHD10, KN12, KTK12, KPKH16, KCH19, LMM18, LW11, LLY16, Li14a, LC15, LGJ^{+}18, LHXP18, LSKZ13, LJL^{+}11, LC11, LJL^{+}15, LLZ18b, LCZ^{+}19, MG14, NMG15, NZM^{+}16, PHXL19, RG17, SDV18, SHG11, SWL17, SCC11, TNZ^{+}12, TTH^{+}19, TZ10, VMB17, WW13, WCD^{+}15, WWL^{+}17, XL16, XSC13, XWJX15, XHQ^{+}15, YWY^{+}17, ZWFX17, ZLW^{+}14, ZCG^{+}17, ZWLL12, ZWL^{+}18, Zou14]. Virtual-Channelless [SHG11]. Virtualization [BHEP14, DY17, GDM^{+}13, HSN17, KMM13b, LWC^{+}17, RKRK17, XML^{+}18, ZQCZ16, ZLW^{+}19, Gua14]. Virtualized [GYQW15, GLBJ18, HC14, KPKH16, LGJZ16, LLJ^{+}13, LIWJ15, PYHY16, PWJ16, SDG17, WW11, WWCZ11, WW13, XCZ^{+}15, XGL^{+}16, YWW^{+}15, ZYQ^{+}14, ZWG^{+}16]. Visibility [dLMPG19]. Visual [LLY^{+}17, LZY^{+}19]. VLAN [KOKA11]. VLC [LGW^{+}17]. VLCcube [LGW^{+}17]. VLIW [WWLJ14]. VLSI [JWJS14, QZG^{+}16]. VM [CTP^{+}17, LPSS19, TVRD17, XTFC17, ZWFX17, ZSW^{+}15]. VMbuddies [LH15]. VMs [LZY^{+}18, VS19]. VMThunder [ZLW^{+}14]. VNET [ZFF16]. VNET-Based [ZFF16]. VNF [TZC19]. VoD [CMG^{+}14, KLWK12, WL12b, WML14]. Voice [LXHS12, LSKZ13]. VoIP [GIP^{+}13, SILJ11]. Vol [Ano15a, Ano16, Ano17a, Ano18, Ano19a]. VOLAP [DRRCB18]. Volatile [APPG16, CDR15, Jun17, MVL15, MV16b, NTDZ19, ZH18]. Volcano [HSX^{+}12, SHX^{+}10]. Voltage [PCL15]. Volume [XFL^{+}19]. von-Neumann [EJGYAM14]. VPNs [RHT13]. vs [Mis14]. Vulnerability [CRZH15, ZYSH14].

Wait [AS16, FVLD16, GD16, IPQ19,

KWG17, PH18, PYH19]. Wait-Free [AS16, FVLD16, IPQ19, KWG17, PH18, PYH19]. Waiting [MB13]. Wake [WLLL10]. Wake-Up [WLLL10]. Walk [ZFT^{+}15, ZYT^{+}15]. Walks [SGGB14].

WANETs [HLS^{+}15]. WAR [APPG16]. Warnings [CJW^{+}15]. Warps [YOK^{+}17]. Water [LWZ12]. Waterman [dOSdM13]. Wave [NSLV16, PBD^{+}13]. Wave-Particle [NSLV16]. Wavefront [ZR18]. Waveguide [AVA^{+}17]. Wavelet [QJ16, vdLJR11]. Way [CP17c, SLL16, SLM^{+}10, TGFFP^{+}19, ZGXJ14]. WBAN [CH13]. WDM [LY11]. Weak [SRB14]. Weakened [PYH19]. Weather [BSM^{+}11]. Web [ALZ17, AWZ15, AKC^{+}15, CZYL14, CMK^{+}16, ECW^{+}18, JLKG17, LGJZ16, SLLZ16, TCZL11, WWCZ11, XTXH13, ZCZ^{+}12, ZLL^{+}15, ZHZC15]. Web-Scale [JLKG17]. Websites [RX11]. Weight [FWZ^{+}16, JRZ^{+}18, ZGL^{+}15]. Weighted [FWZ^{+}16, LZY^{+}18, LY14, LWL^{+}17, LSW^{+}15, MJM16, WZS^{+}19]. Weighted-Tuple [MJM16]. Weights [CJ16]. Well [MSB11]. Well-Formed [MSB11]. Wheel [ZMF10]. Wheel-Rail [ZMF10]. Wheeler [WH16]. Whether [WCD^{+}11]. Which [Hen14]. Wide [CHLC15, LNL^{+}19, dOSMM^{+}16, SLGW14, TCT14, YYK11a, ZASA10]. Wide-Area [LNL^{+}19, SLGW14]. Widely [YYK^{+}11b]. Width [AA14]. WiFi [LQK^{+}13, XLM^{+}11a, ZY13]. WiFi-Based [ZY13]. WILL [WYLX13]. Willow [LYH^{+}15]. WiMAX [MM12]. WiMAX/WLAN [MM12]. Win [SL16]. Win-Win [SL16]. Window [Lu14, RPYO11, VBC19]. Windows [WHYZ10]. WiNoC [DKM^{+}15]. Wired [AVA^{+}17]. Wired-Wireless [AVA^{+}17]. Wireless [AMN^{+}16, ATACA18, AO12, ALLR14, AVA^{+}17, ADZZM15, ACNP11, Amm12, ACV17, BBCB15, BKY15, BCSKN12, CCFS11, CWL14b, CHCC14, Cha14, CLL11, CHTW12, CLLS12, Che14, CYL^{+}14, CYC^{+}15, CHD^{+}15, CCT16, CH13, CNC^{+}14, CLJ11, CIH13, CLHK11, CWJS11, CWC^{+}13, DCW^{+}15, DGF12, DKM^{+}15, DRSL15, DWW^{+}11, DCL^{+}10, DLL^{+}11, DLZ^{+}14, DOLG16, DWY^{+}13, EK10, FLH13, sFC12, FQWL12, FW13, GBD^{+}13, GFLL15, GTS^{+}15, GLL15, GLL11, GJLZ13, GCN^{+}14, GJZZ12, GCL14, GLJ^{+}15, GCZ15, GLC^{+}15, HGY^{+}14, HCS12, HCL^{+}12, HCC^{+}12, HJPL14, HCG^{+}15, HDL^{+}15, HCJ^{+}10, HWI12, HLY^{+}14, HH12, HHK10, IvS10, JWA10, JCLJ12, JLW^{+}10, JJW11, JHW^{+}15, JLM^{+}12, JJG^{+}12, KKW13, KCK14, KKY^{+}14, KCYM10, KXL^{+}14, KL11b, KSP10, LLGP13, LKE16, LAV^{+}10, LVA^{+}11, LXHS12, LRW12, Li13, LWY^{+}13, LLL^{+}13, LMSRSR13, LG13, LCZZ13, LHD^{+}14, LCS14, Li14c, LLK13, LZN10, LZNX11, LM12, LHL^{+}13b, LCLD13, LZCK14, LLXC14, LWJ^{+}15, LZK^{+}15]. Wireless [LLH^{+}15a, LLZ^{+}12b, LLG14, LTMD11, LWG^{+}12, LGG^{+}14, MLL14, MLC^{+}15, MS12, MS13a, MLS15, MM15, MTX^{+}11, MLT^{+}13, MY11, MGR12, PB12, RGRM14, RM12, RGK15, RYLZ10, RZH^{+}11, RHDL11, RZW^{+}13, RWLL14, RVW^{+}15, SCC11, SP15, SJAdCL19, SKP12, SHM^{+}12, TWW^{+}15, TLRW15, TCS11, THL13, TKP12, UBC13, VM12, VWDM14, WPT10, WLS^{+}11, WMT^{+}11, WWL11, WMHX12, WFK^{+}12, WJTL12, WWH13, WWLX13, WFA13, WYX13, WTL^{+}14, Wan14, WL14, WSL^{+}15, WHB16, WG13, WZQ10, WCF13, WWCB14, XWH15b, XHHC13, XJ14, XHG15, XWY^{+}10, XLM^{+}11b, XHQ^{+}15, XAK17, XHZ^{+}13, YCTC13, YK14, YRL11, YLT15, ZWD^{+}10, ZS10, ZZF10, ZMA12, ZMLT13, ZZCD10, ZWLL12, ZX13, ZYT^{+}15, WYLX13]. Wiring [CMB18]. within [SKKK16]. Without [ZQWL17, DWX14, GCZ15, KDCR19, QPB^{+}17, WLL^{+}13, WYLX13, XYT^{+}15, XL16, XSYY13]. WLAN [MM12]. WLANs [GYX^{+}10, NZWL14, YWC11]. Word [IPQ19]. Word2Vec [JSLD19]. Work [CW15, CGH13, HH13, PWJ16, TNLM17]. Work-Efficient [HH13]. Work-Stealing [CGH13, PWJ16]. Worker [DLZH16, PF12, TNLM17]. Workflow [ABN19, BWB^{+}19, DHTZ15, FPF13, HWSX17, RM17, SVK^{+}19, SCJ^{+}17, WIZ^{+}17, YDH17, YWZ17, ZZLL16]. Workflow-Aware [SVK^{+}19]. Workflows [ANE12, CB14, CZQ^{+}17, CAKRY16, PP12, VLP16, ZHCL17, ZWG^{+}16]. Worklist [GIX^{+}12]. Workload [BB17, CZD^{+}19, dCCF15, GGF^{+}14, HLZ^{+}19, HLCB^{+}17, JWK^{+}16, Li10, LQW^{+}18, LVD11, MWJ16, MNE14, PAB13, SEA18, SVL^{+}16, WHGS17, WVM19, WHYZ10, XFL15, YGL^{+}15, YWW^{+}15, YLZ^{+}15b, YJCQ15, ZWFX17, ZLL17c]. Workload-Adaptive [HLZ^{+}19]. Workload-Aware [JWK^{+}16, ZWFX17]. Workloads [CSW^{+}12, CC17, FYH^{+}15, HMS^{+}18, HYZ15, JZW^{+}17, LWZ^{+}13, LWZ^{+}16b, MZK19, TRD13, WFZ^{+}17, WV17, YHS^{+}14, YZHZ17, ZJS^{+}17, ZHW^{+}19]. World [HSX^{+}12, IRSNF11, LCGC14, NSLV16, VMN^{+}16]. Wormhole [Dua95, SHG11, VS11a, VS11b, VS14, Dua93]. Wormhole-Switched [SHG11]. Worms [WZZ^{+}13, YZFZ10]. Worst [MLT^{+}13]. WPAN [YTL^{+}10]. WPANs [HKH^{+}10]. Write [GRCZ17, JRZ^{+}18, LWZ^{+}16c, Sto10f, WDH^{+}16, ZH18]. Write-Friendly [ZH18]. Writes [CP17c, SSF16b, SLSG19]. Writing [DKL^{+}19]. WSNs [LYGX12, LCS^{+}15, ZQSY13].

X [LMPR12, ZWL^{+}16a]. X-BOT [LMPR12]. X-Code [ZWL^{+}16a]. X10 [CMK^{+}16]. x86 [HWF18, LJ16]. x86-Based [HWF18]. Xeon [LSW17a, LLH^{+}15b, CRS^{+}17]. Xeon/Xeon [CRS^{+}17]. XML [EHI11, ZLZ^{+}14]. XMT [VTSM12]. XOR [SSF16b, SSLF17, SLSG18]. XOR-Coded [SSF16b, SSLF17, SLSG18]. XPLORE [WYW^{+}14, ZZ15]. Xscale [ZWL^{+}16a].

Z [AP17].  Z-Fat [AP17].  Zapping

[TCS13]. ZEBRA [ASG^{+}14]. Zero [VMB17, XWH15a]. Zero-Copy [VMB17]. Zero-Knowledge [XWH15a]. ZFP [TDL^{+}19]. ZigBee [HPH^{+}12, KKY^{+}14]. Zone [LC15, MMSAZ11]. Zone-Ordered [MMSAZ11]. Zones [MT15].

References

Aydonat:2012:RCC

[AA12] Utku Aydonat and Tarek S. Abdelrahman. Relaxed concurrency control in software transactional memory. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1312–1325, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aroca:2014:BBW

[AA14] Jordi Arjona Aroca and Antonio Fernandez Anta. Bisection (band)width of product networks with application to data centers. IEEE Transactions on Parallel and Distributed Systems, 25(3): 570–580, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Akbudak:2017:ELS

[AA17] Kadir Akbudak and Cevdet Aykanat. Exploiting locality in sparse matrix–matrix multiplication on many-core architectures. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2258–2271, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07829317-
abs.html.

Abubaker:2019:SGH

[AAA19] Nabil Abubaker, Kadir Akbudak, and Cevdet Aykanat. Spatiotemporal graph and hypergraph partitioning models for sparse matrix–vector multiplication on many-core architectures. IEEE Transactions on Parallel and Distributed Systems, 30(2):445– 458, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08432126-
abs.html.

Akavipat:2014:RFR

[AAAK^{+}14] Ruj Akavipat, Mahdi N. Al-Ameen, Apu Kapadia, Zahid Rahman, Roman Schlegel, and Matthew Wright. ReDS: A framework for reputationenhanced DHTs. IEEE Transactions on Parallel and Distributed Systems, 25(2): 321–331, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agyeman:2016:PEA

[AAB16] Michael Opoku Agyeman, Ali Ahmadinia, and Nader Bagherzadeh. Performance and energy aware inhomogeneous 3D networks-on-chip architecture generation. IEEE Transactions on Parallel and Distributed Systems, 27(6):1756–1769, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07161378-
abs.html.

Agullo:2017:BGB

[AAB^{+}17] Emmanuel Agullo, Olivier Aumage, Berenger Bramas, Olivier Coulaud, and Samuel Pitoiset. Bridging the gap between OpenMP and task-based runtime systems for the Fast Multipole Method. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2794–2807, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07912335-
abs.html.

aabdi:2019:EQC

[aaGZ19] a. abdi, A. Girault, and H. R. Zarandi. ERPOT: A quad-criteria scheduling heuristic to optimize execution time, reliability, power consumption and temperature in multicores. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2193–2210, October 2019.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aguston:2015:PHC

[AAH15] Cfir Aguston, Yosi Ben Asher, and Gadi Haber. Parallelization hints via code skeletonization. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3099–3107, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06987337-
abs.html.

Aktulga:2017:HPB

[AAW^{+}17] Hasan Metin Aktulga, Md.

Afibuzzaman, Samuel Williams, Aydin Buluc, Meiyue Shao, Chao Yang, Esmond G. Ng, Pieter Maris, and James P. Vary. A high performance block eigensolver for nuclear configuration interaction calculations. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1550–1563, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07748453-
abs.html.

Anceaume:2014:DID

[AB14] Emmanuelle Anceaume and Yann Busnel. A distributed information divergence estimation over data streams. IEEE Transactions on Parallel and Distributed Systems, 25(2):478– 487, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Atmaca:2016:PEC

[ABBCT16] Tulin Atmaca, Thomas Begin, Alexandre Brandwajn, and Hind Castel-Taleb. Performance evaluation of cloud computing centers with general arrivals and service. IEEE Transactions on Parallel and Distributed Systems, 27(8):2341–2348, August 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07327216-
                                        abs.html.

Abramson:2011:PES

[ABE^{+}11] David Abramson, Blair Bethwaite, Colin Enticott, Slavisa Garic, and Tom Peachey. Parameter exploration in science and engineering using many-task computing. IEEE Transactions on Parallel and Distributed Systems, 22(6):960– 973, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Albader:2012:ECA

[ABF12] Bader Albader, Bella Bose, and Mary Flahive. Efficient communication algorithms in hexagonal mesh interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 23(1):69– 77, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Angiulli:2016:GSD

[ABLS16] Fabrizio Angiulli, Stefano Basta, Stefano Lodi, and Claudio Sartori. GPU strategies for distance-based outlier detection. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3256–3268, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07405341-
abs.html.

Arabnejad:2019:BDA

[ABN19] Vahid Arabnejad, Kris Bubendorfer, and Bryan Ng. Budget and deadline aware escience workflow scheduling in clouds. IEEE Transactions on Parallel and Distributed Systems, 30(1):29– 44, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08400402-
abs.html.

Azad:2017:CMC

[ABP17] Ariful Azad, Aydn Buluc, and Alex Pothen. Computing maximum cardinality matchings in parallel on bipartite graphs via treegrafting. IEEE Transactions on Parallel and Distributed Systems, 28(1):44– 59, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07440857-
abs.html.

Alvanos:2019:AAC

[AC19] M. Alvanos and T. Christoudias. Accelerating atmospheric chemical kinetics for climate simulations. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2396–2407, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Altomare:2017:TPM

[ACC^{+}17] Albino Altomare, Eugenio Cesario, Carmela Comito, Fabrizio Marozzo, and Domenico Talia. Trajectory pattern mining for urban computing in the cloud. IEEE Transactions on Parallel and Distributed Systems, 28(2):586– 599, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07467530-
abs.html.

Alomair:2012:SRS

[ACCP12] Basel Alomair, Andrew Clark, Jorge Cuellar, and Radha Poovendran. Scalable RFID systems: a privacy-preserving protocol with constant-time identification. IEEE Transactions on Parallel and Distributed Systems, 23(8):1536–1550, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Attiya:2019:ESR

[ACE^{+}19] Hagit Attiya, Hyun Chul Chung, Faith Ellen, Saptaparni Kumar, and Jennifer L. Welch. Emulating a shared register in a system that never stops changing. IEEE Transactions on Parallel and Distributed Systems, 30(3):544–559, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08449124-
abs.html.

Albano:2011:DND

[ACNP11] Michele Albano, Stefano Chessa, Francesco Nidito, and Susanna Pelagatti. Dealing with nonuniformity in data centric storage for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1398–1406, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Alves:2013:FAM

[ACS13] C. E. R. Alves, E. N. Caceres, and Siang Wun Song. Finding all maximal contiguous subsequences of a sequence of numbers in O(1) communication rounds. IEEE Transactions on Parallel and Distributed Systems, 24(4):724–733, April 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Asaduzzaman:2017:EED

[ACV17] Abu Asaduzzaman, Kishore K. Chidella, and Divya Vardha. An energy-efficient directory based multicore architecture with wireless routers to minimize the communication latency. IEEE Transactions on Parallel and Distributed Systems, 28(2):374– 385, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/02/07475919-
                                        abs.html.

Anselmi:2019:AOS

[AD19] J. Anselmi and J. Doncel. Asymptotically optimal size-interval task assignments. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2422–2433, November 2019.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Arleo:2019:DMF

[ADLM19] Alessio Arleo, Walter Didimo, Giuseppe Liotta, and Fabrizio Montecchiani. A distributed multilevel forcedirected algorithm. IEEE Transactions on Parallel and Distributed Systems, 30(4):754–765, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8462766/

.

Al-Dujaily:2012:ETC

[ADMX^{+}12] Ra’ed Al-Dujaily, Terrence Mak, Fei Xia, Alexandre (Alex) Yakovlev, and Maurizio Palesi. Embedded transitive closure network for runtime deadlock detection in networks-on-chip. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1205–1215, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Al-Dubai:2015:QAI

[ADZZM15] Ahmed Y. Al-Dubai, Liang Zhao, Albert Y. Zomaya, and Geyong Min. QoSaware inter-domain multicast for scalable wireless community networks. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3136–3148, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/11/06936933-
                                        abs.html.

Alekeish:2012:CSM

[AE12] Khaled Alekeish and Paul Ezhilchelvan. Consensus in sparse, mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 23(3):467– 474, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aldea:2016:OES

[AELGE16] Sergio Aldea, Alvaro Estebanez, Diego R. Llanos, and Arturo Gonzalez-Escribano. An OpenMP extension that supports thread-level speculation. IEEE Transactions on Parallel and Distributed Systems, 27(1):78– 91, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07014262-
                                        abs.html.

Attiya:2017:LHA

[AEM17] Hagit Attiya, Faith Ellen, and Adam Morrison. Limitations of highly-available eventually-consistent data stores. IEEE Transactions on Parallel and Distributed Systems, 28(1):141– 155, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07457259-
abs.html.

Alonso:2018:PPD

[AF18] Miguel Gorgues Alonso and Jose Flich. PROSA: Protocol-driven network on chip architecture. IEEE Transactions on Parallel and Distributed Systems, 29(7): 1560–1574, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08219740-
abs.html.

Asahara:2019:ESD

[AF19] M. Asahara and R. Fujimaki.

An empirical study on distributed Bayesian approximation inference of piecewise sparse linear models. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1481–1493, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abellan:2012:EHB

[AFA12] Jose L. Abellan, Juan Fernandez, and Manuel E. Acacio. Efficient hardware barrier synchronization in many-core CMPs. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1453–1466, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Azad:2017:EPT

[AFMM17] Zahra Azad, Hamed Farbeh, Amir Mahdi Hosseini Monazzah, and Seyed Ghassem Miremadi. An efficient protection technique for last level STT-RAM caches in multi-core processors. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1564–1577, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07744670-
abs.html.

Alvanos:2016:CSD

[AFT^{+}16] Michail Alvanos, Montse Farreras, Ettore Tiotto, Jose Nelson Amaral, and Xavier Martorell. Combining static and dynamic data coalescing in Unified Parallel C. IEEE Transactions on Parallel and Distributed Systems, 27(2):381– 393, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07045602-
abs.html.

Anglano:2015:ERC

[AGG15] Cosimo Anglano, Rossano Gaeta, and Marco Grangetto. Exploiting rateless codes in cloud storage systems. IEEE Transactions on Parallel and Distributed Systems, 26(5):1313–1322, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06810144-
abs.html.

Anglano:2017:SCB

[AGG17] Cosimo Anglano, Rossano Gaeta, and Marco Grangetto. Securing coding-based cloud storage against pollution attacks. IEEE Transactions on Parallel and Distributed Systems, 28(5):1457–1469, May 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07604064-
abs.html.

Anta:2016:DSD

[AGJ^{+}16] Antonio Fernandez Anta, Vincent Gramoli, Ernesto Jimenez, Anne-Marie Kermarrec, and Michel Raynal. Distributed slicing in dynamic systems. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1030–1043, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07103341-
                                        abs.html.

Agrawal:2014:BLS

[Agr14] Shailesh Kumar Agrawal. Broadcasting on large scale heterogeneous platforms under the bounded multi-port model. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2520–2528, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06613483-
                                        abs.html.

Attiya:2010:TSL

[AH10] Hagit Attiya and Danny Hendler. Time and space lower bounds for implementations using k-CAS. IEEE Transactions on Parallel and Distributed Systems, 21(2): 162–173, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Antikainen:2011:NTF

[AHJ^{+}11] Jukka Antikainen, Jiří Havel, Radovan Josth, Adam Herout, Pavel Zemčik, and Markku Hauta-Kasari. Nonnegative tensor factorization accelerated using GPGPU. IEEE Transactions on Parallel and Distributed Systems, 22(7):1135–1141, July 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Arora:2015:FNF

[AHS^{+}15] Anuj Arora, Mayur Harne, Hameedah Sultan, Akriti Bagaria, and Smruti R. Sarangi. FP-NUCA: A fast NOC layer for implementing large NUCA caches. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2465–2478, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06898874.
pdf.

Aazam:2016:CCH

[AHSH^{+}16] Mohammad Aazam, Eui-Nam Huh, Marc St-Hilaire, Chung-Horng Lung, and Ioannis Lambadaris. Cloud customer’s historical record based resource pricing. IEEE Transactions on Parallel and Distributed Systems, 27(7): 1929–1940, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07226829-
abs.html.

Abdelaziz:2017:CVC

[AHSK17] Ibrahim Abdelaziz, Razen Harbi, Semih Salihoglu, and Panos Kalnis. Combining vertex-centric graph processing with SPARQL for large-scale RDF data analytics. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3374–3388, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07959641-
abs.html.

Abdelfattah:2018:ADT

[AHTD18] Ahmad Abdelfattah, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Analysis and design techniques towards high-performance and energy-efficient dense linear solvers on GPUs. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2700–2712, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08370686-
abs.html.

Allahbakhsh:2015:IMC

[AI15] Mohammad Allahbakhsh and Aleksandar Ignjatovic. An iterative method for calculating robust rating scores. IEEE Transactions on Parallel and Distributed Systems, 26(2):340– 350, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2015/02/06762988-
abs.html.

Alsarhan:2018:ARA

[AIAD^{+}18] Ayoub Alsarhan, Awni Itradat, Ahmed Y. Al-Dubai, Albert Y. Zomaya, and Geyong Min. Adaptive resource allocation and provisioning in multi-service cloud environments. IEEE Transactions on Parallel and Distributed Systems, 29(1):31– 42, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08027098-
abs.html.

Arjomand:2017:HPM

[AJK^{+}17] Mohammad Arjomand, Amin Jadidi, Mahmut T. Kandemir, Anand Sivasubramaniam, and Chita R. Das. HL-PCM: MLC PCM main memory with accelerated read. IEEE Transactions on Parallel and Distributed Systems, 28(11):3188–3200, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07930492-
abs.html.

Abousamra:2012:CNC

[AJM12] Ahmed Abousamra, Alex K.

Jones, and Rami Melhem. Codesign of NoC and cache organization for reducing access latency in chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1038–1046, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agrawal:2014:MPM

[AJMW14] Dharma P. Agrawal, Junghyun Jun, Joanna Mitro, and Nishan Weragama. Modeling and performance of a mesh network with dynamically appearing and disappearing femtocells as additional Internet gateways. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1278–1288, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

AlFarhan:2018:OUA

[AK18] Mohammed A. Al Farhan and David E. Keyes. Optimizations of unstructured aerodynamics computations for many-core architectures. IEEE Transactions on Parallel and Distributed Systems, 29(10):2317–2332, October 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/10/08337750-
                                        abs.html.

Alexander:2015:CWC

[AKC^{+}15] Holly Alexander, Ibrahim Khalil, Conor Cameron, Zahir Tari, and Albert Zomaya. Cooperative Web caching using dynamic interest-tagged filtered Bloom filters. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2956–2969, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06926845-
abs.html.

Al-Kiswany:2013:GSS

[AKGR13] Samer Al-Kiswany, Abdullah Gharaibeh, and Matei Ripeanu. GPUs as storage system accelerators. IEEE Transactions on Parallel and Distributed Systems, 24(8):1556–1566, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anifantis:2014:SSF

[AKP14] Evangelos Anifantis, Vasileios Karyotis, and Symeon Papavassiliou. A spatiostochastic framework for cross-layer design in cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2762–2771, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06714573-
abs.html.

Almashor:2015:EAC

[AKT^{+}15] Mahathir Almashor, Ibrahim Khalil, Zahir Tari, Albert Y. Zomaya, and Sartaj Sahni. Enhancing availability in content delivery networks for mobile platforms. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2247–2257, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/08/06863686-
                                        abs.html.

Asahi:2017:OFK

[ALI^{+}17] Yuuichi Asahi, Guillaume Latu, Takuya Ina, Yasuhiro Idomura, Virginie Grandgirard, and Xavier Garbet. Optimization of fusion kernels on accelerators with indirect or strided memory access patterns. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1974–1988, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/07/07762180-
                                        abs.html.

Abdelhakim:2014:DDM

[ALLR14] Mai Abdelhakim, Tongtong Li, Leonard E. Lightfoot, and Jian Ren. Distributed detection in mobile access wireless sensor networks under Byzantine attacks. IEEE Transactions on Parallel and Distributed Systems, 25(4): 950–959, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abdulah:2018:EHP

[ALS^{+}18] Sameh Abdulah, Hatem Ltaief, Ying Sun, Marc G. Genton, and David E. Keyes. ExaGeoStat: A high performance unified software for geostatistics on manycore systems. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2771–2784, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08396303-
abs.html.

Adam:2017:SRP

[ALZ17] Omer Adam, Young Choon Lee, and Albert Y. Zomaya. Stochastic resource provisioning for containerized multi-tier Web services in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(7):2060–2073, July 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07782339-
abs.html.

Ababei:2019:SPC

[AM19] C. Ababei and M. G.

Moghaddam. A survey of prediction and classification techniques in multicore processor systems. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1184–1200, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ammari:2012:CEA

[Amm12] Habib M. Ammari. CSI: An energy-aware cover-sense-inform framework for kcovered wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):651– 658, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abadal:2016:SBP

[AMN^{+}16] Sergi Abadal, Albert Mestres, Mario Nemirovsky, Heekwan Lee, Antonio Gonzalez, Eduard Alarcon, and Albert Cabellos-Aparicio. Scalability of broadcast performance in wireless networkon-chip. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3631–3645, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07423812-
abs.html.

Abrishami:2012:CDS

[ANE12] Saeid Abrishami, Mahmoud Naghibzadeh, and Dick H. J. Epema. Cost-driven scheduling of grid workflows using partial critical paths. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1400–1414, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abdelkader:2013:SRP

[ANN^{+}13] Tamer Abdelkader, Kshirasagar Naik, Amiya Nayak, Nishith Goel, and Vineet Srivastava. SGBR: A routing protocol for delay tolerant networks using social grouping. IEEE Transactions on Parallel and Distributed Systems, 24(12):2472–2481, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2010:RL

[Ano10] Anonymous. 2009 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 21(1):139– 144, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2011:AI

[Ano11a] Anonymous. 2010 annual index. IEEE Transactions on Parallel and Distributed Systems, 22(1):[online only], January 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2011:RL

[Ano11b]  Anonymous.  2010 re-

viewer’s list. IEEE Transactions on Parallel and Distributed Systems, 22(1):185– 191, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2011:CPS

[Ano11c] Anonymous. Call for papers for a special issue of IEEE Transactions on Parallel and Distributed Systems (TPDS) on cyber-physical systems (CPS). IEEE Transactions on Parallel and Distributed Systems, 22(7): 1247, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2011:CPI

[Ano11d] Anonymous. Call for papers for IEEE Transactions on Parallel and Distributed Systems (TPDS) special issue on cloud computing. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2126–2127, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2011:CEN

[Ano11e]  Anonymous.  Correction

to “Editor’s note”. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1246, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:AI

[Ano12a] Anonymous. 2011 annual index. IEEE Transactions on Parallel and Distributed Systems, 23(1):INDEX:1– INDEX:23, January 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:RL

[Ano12b] Anonymous. 2011 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 23(1):185– 191, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:CPS

[Ano12c] Anonymous. Call for papers:

Special Issue on Trust, Security, and Privacy in Parallel and Distributed Systems. IEEE Transactions on Parallel and Distributed Systems, 23(6):1168, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:Ca

[Ano12d] Anonymous. Cover1. IEEE Transactions on Parallel and Distributed Systems, 23(9): c1, September 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:Cb

[Ano12e] Anonymous. Cover2. IEEE Transactions on Parallel and Distributed Systems, 23(9): c2, September 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:Cc

[Ano12f] Anonymous. Cover3. IEEE Transactions on Parallel and Distributed Systems, 23(9): c3, September 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:Cd

[Ano12g] Anonymous. Cover4. IEEE Transactions on Parallel and Distributed Systems, 23(9): c4, September 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:CHD

[Ano12h]  Anonymous.  CPS han-

dles the details. IEEE Transactions on Parallel and Distributed Systems, 23(9):

1804, September 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2012:IOA

[Ano12i] Anonymous. IEEE open access publishing. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2380, December 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.
ieee.org/open-access.

Anonymous:2012:NTN

[Ano12j] Anonymous. New Transactions newsletter. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1803, September 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2013:AI

[Ano13a] Anonymous. 2012 annual index. IEEE Transactions on Parallel and Distributed Systems, 24(1):web, January 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2013:RL

[Ano13b]  Anonymous.  2012 re-

viewer’s list. IEEE Transactions on Parallel and Distributed Systems, 24(1):198– 207, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2014:AI

[Ano14a] Anonymous. 2013 annual index. IEEE Transactions on Parallel and Distributed Systems, 25(1):not in print, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2014:RL

[Ano14b] Anonymous. 2013 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 25(1):268– 278, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2015:IIT

[Ano15a] Anonymous. 2014 index IEEE Transactions on Parallel and Distributed Systems vol. 25. IEEE Transactions on Parallel and Distributed Systems, 26(1):291– 328, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06982263.
pdf.

Anonymous:2015:RL

[Ano15b] Anonymous. 2014 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 26(1):282– 290, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06979387.
pdf.

Anonymous:2016:IIT

[Ano16] Anonymous. 2015 index IEEE Transactions on Parallel and Distributed Systems vol. 26. IEEE Transactions on Parallel and Distributed Systems, 27(1):1– 36, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07350314.
pdf.

Anonymous:2017:IIT

[Ano17a] Anonymous. 2016 index IEEE Transactions on Parallel and Distributed Systems vol. 27. IEEE Transactions on Parallel and Distributed Systems, 28(1):305– 338, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07779225.
pdf.

Anonymous:2017:RLa

[Ano17b] Anonymous. 2016 reviewers list*. IEEE Transactions on Parallel and Distributed Systems, 28(2):600– 611, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/02/07818093.
                                        pdf.

Anonymous:2017:RLb

[Ano17c] Anonymous. 2017 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3684–3689, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/08103837.
                                        pdf.

Anonymous:2018:IIT

[Ano18] Anonymous. 2017 index IEEE Transactions on Parallel and Distributed Systems vol. 28. IEEE Transactions on Parallel and Distributed Systems, 29(1):240– 273, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/01/08172510.
                                        pdf.

Anonymous:2019:IIT

[Ano19a] Anonymous. 2018 index IEEE Transactions on Parallel and Distributed Systems vol. 29. IEEE Transactions on Parallel and Distributed Systems, 30(1):1– 28, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08572814.
pdf.

Anonymous:2019:RLa

[Ano19b] Anonymous. 2018 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 30(5):1201–1207, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2019:RLb

[Ano19c] Anonymous. 2019 reviewers list. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2896–2900, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

AbdelSalam:2012:BBB

[AO12] Hady S. AbdelSalam and Stephan Olariu. BEES: BioinspirEd backbonE Selection in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(1):44– 51, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Arif:2012:DAR

[AOW^{+}12] Samiur Arif, Stephan Olariu, Jin Wang, Gongjun Yan, Weiming Yang, and Ismail Khalil. Datacenter at the airport: Reasoning about time-dependent parking lot occupancy. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2067–2080, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Adda:2017:RFT

[AP17] Mo Adda and Adamantini Peratikou. Routing and fault tolerance in Z-fat tree. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2373–2386, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07850986-
abs.html.

Araya-Polo:2011:AAB

[APCH^{+}11] Mauricio Araya-Polo, Javier Cabezas, Mauricio Hanzich, Miquel Pericas, Felix Rubio, Isaac Gelado, Muhammad Shafiq, Enric Morancho, Nacho Navarro, Eduard Ayguade, Jose Maria Cela, and Mateo Valero. Assessing accelerator-based HPC reverse time migration. IEEE Transactions on Parallel and Distributed Systems, 22(1): 147–162, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abad:2012:BPC

[APG12] Pablo Abad, Valentin Puente, and Jose Angel Gregorio. Balancing performance and cost in CMP interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 23(3):452– 459, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aji:2016:MAA

[APJ^{+}16] Ashwin M. Aji, Lokendra S.

Panwar, Feng Ji, Karthik Murthy, Milind Chabbi, Pavan Balaji, Keith R. Bisset, James Dinan, Wu chun Feng, John Mellor-Crummey, Xiaosong Ma, and Rajeev Thakur. MPI-ACC: Acceleratoraware MPI for scientific applications. IEEE Transactions on Parallel and Distributed Systems, 27(5): 1401–1414, May 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07127020-
abs.html.

Abad:2012:ATM

[APMG12] Pablo Abad, Valentin Puente, Lucia G. Menezo, and Jose Angel Gregorio. Adaptivetree multicast: Efficient multidestination support for CMP communication substrate. IEEE Transactions on Parallel and Distributed Systems, 23(11):2010–2023, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abad:2016:AWA

[APPG16] Pablo Abad, Pablo Prieto, Valentin Puente, and Jose-Angel Gregorio. AC-WAR: Architecting the cache hierarchy to improve the lifetime of a non-volatile endurancelimited main memory. IEEE Transactions on Parallel and Distributed Systems, 27 (1):66–77, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07006735-
                                        abs.html.

Anta:2010:AIR

[AR10] Antonio Fernandez Anta and Michel Raynal. From an asynchronous intermittent rotating star to an eventual leader. IEEE Transactions on Parallel and Distributed Systems, 21(9):1290–1303, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Aravind:2011:YAS

[Ara11] Alex A. Aravind. Yet another simple solution for the concurrent programming control problem. IEEE Transactions on Parallel and Distributed Systems, 22(6): 1056–1063, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Azarderakhsh:2015:PHS

[ARM15] Reza Azarderakhsh and Arash Reyhani-Masoleh. Parallel and high-speed computations of elliptic curve cryptography using hybriddouble multipliers. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1668–1677, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06814322-
abs.html.

Araldo:2016:CAC

[ARM16] Andrea Araldo, Dario Rossi, and Fabio Martignon. Costaware caching: Caching more (costly items) for less (ISPs operational expenditures). IEEE Transactions on Parallel and Distributed Systems, 27(5):1316–1330, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07108052-
abs.html.

Aggarwal:2016:LFW

[AS16] Pooja Aggarwal and Smruti R. Sarangi. Lock-free and wait-free slot scheduling algorithms. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1387–1400, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07110616-
abs.html.

Alasaad:2015:ISR

[ASBL15] Amr Alasaad, Kaveh Shafiee, Hatim M. Behairy, and Victor C. M. Leung. Innovative schemes for resource allocation in the cloud for media streaming applications. IEEE Transactions on Parallel and Distributed Systems, 26(4):1021–1033, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06787111-
abs.html.

Ax:2018:CMM

[ASD^{+}18] Johannes Ax, Gregor Sievers, Julian Daberkow, Martin Flasskamp, Marten Vohrmann, Thorsten Jungeblut, Wayne Kelly, Mario Porrmann, and Ulrich Ruckert. CoreVA-MPSoC: A many-core architecture with tightly coupled shared and local data memories. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1030–1043, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08240949-
abs.html.

Acacio:2014:ZDC

[ASG^{+}14] Manuel E. Acacio, Per Stenstrom, Jose M. Garcia, Ruben Titos-Gil, and Anurag Negi. ZEBRA: Datacentric contention management in hardware transactional memory. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1359–1369, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Asyabi:2018:TMT

[ASSB18] Esmail Asyabi, SeyedAlireza SanaeeKohroudi, Mohsen Sharifi, and Azer Bestavros. TerrierTail: Mitigating tail latency of cloud virtual machines. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2346–2359, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/10/08338088-
abs.html.

Ansari:2019:PPM

[ASYK^{+}19] Mohsen Ansari, Sepideh Safari, Amir Yeganeh-Khaksar, Mohammad Salehi, and Alireza Ejlali. Peak power management to meet thermal design power in faulttolerant embedded systems.

IEEE Transactions on Parallel and Distributed Systems, 30(1):161–173, January 2019. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/01/08419269-
                                        abs.html.

Altiparmak:2012:EDA

[AT12] Nihat Altiparmak and Ali Şaman Tosun. Equivalent disk allocations. IEEE Transactions on Parallel and Distributed Systems, 23(3):538– 546, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Acer:2018:IMG

[ATA18] Seher Acer, Tugba Torun, and Cevdet Aykanat. Improving medium-grain partitioning for scalable sparse tensor decomposition. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2814–2825, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/12/08368258-
                                        abs.html.

Abadal:2018:OBO

[ATACA18] Sergi Abadal, Josep Torrellas, Eduard Alarcón, and Albert Cabellos-Aparicio. OrthoNoC: A broadcastoriented dual-plane wireless network-on-chip architecture. IEEE Transactions on Parallel and Distributed Systems, 29(3):628–641, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8078211/.

Abhaya:2014:PAE

[ATZZ14] Vidura Gamini Abhaya, Zahir Tari, Panlop Zeephongsekul, and Albert Y. Zomaya. Performance analysis of EDF scheduling in a multi-priority preemptive M/G/1 queue. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2149–2158, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Agyeman:2017:RDW

[AVA^{+}17] Michael Opoku Agyeman, Quoc-Tuan Vien, Ali Ahmadinia, Alexandre Yakovlev, Kin-Fai Tong, and Terrence Mak. A resilient 2-D waveguide communication fabric for hybrid wired-wireless NoC design. IEEE Transactions on Parallel and Distributed Systems, 28(2):359– 373, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07484291.
pdf.

Angstadt:2019:PPR

[AWWS19] Kevin Angstadt, Jack Wadden, Westley Weimer, and Kevin Skadron. Portable programming with RAPID. IEEE Transactions on Parallel and Distributed Systems, 30(4):939–952, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

                                        URL https://ieeexplore.
                                        ieee.org/document/8462761/

.

Ahmed:2015:RTB

[AWZ15] Waseem Ahmed, Yongwei Wu, and Weimin Zheng. Response time based optimal Web service selection. IEEE Transactions on Parallel and Distributed Systems, 26(2): 551–561, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06684156-
                                        abs.html.

Ai:2019:CDF

[AZW^{+}19] Zhiyuan Ai, Mingxing Zhang, Yongwei Wu, Xuehai Qian, Kang Chen, and Weimin Zheng. Clip: A disk I/O focused parallel out-of-core graph processing system. IEEE Transactions on Parallel and Distributed Systems, 30(1):45–62, January 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08416754-
abs.html.

Bsoul:2016:RBD

[BAAT16] Mohammad Bsoul, Alaa E.

Abdallah, Khaled Almakadmeh, and Nedal Tahat. A round-based data replication strategy. IEEE Transactions on Parallel and Distributed Systems, 27(1):31– 39, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07004058-
abs.html.

Bader:2014:SJ

[Bad14] David A. Bader. State of the journal. IEEE Transactions on Parallel and Distributed Systems, 25(1):1, January 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bader:2015:EN

[Bad15] David A. Bader. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 26(1):1, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06980173.
pdf.

Bader:2016:EN

[Bad16] David A. Bader. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 27(1):1, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07350369.
                                        pdf.

Bader:2017:ENa

[Bad17a] David A. Bader. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 28(1):1, January 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/01/07779313.
                                        pdf.

Bader:2017:ENb

[Bad17b]  David A. Bader.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3328–3329, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/08103890-
                                        abs.html.

Baquero:2012:EPF

[BAMJ12] Carlos Baquero, Paulo Sergio Almeida, Raquel Menezes, and Paulo Jesus. Extrema propagation: Fast distributed estimation of sums and network sizes. IEEE Transactions on Parallel and Distributed Systems, 23(4): 668–675, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Barlas:2010:AAO

[Bar10] Gerassimos Barlas. An analytical approach to optimizing parallel image registration/retrieval. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1074–1088, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beloglazov:2013:MOH

[BB13] Anton Beloglazov and Rajkumar Buyya. Managing overloaded hosts for dynamic consolidation of virtual machines in cloud data centers under quality of service constraints. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1366–1379, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Busato:2015:BEI

[BB15] Federico Busato and Nicola Bombieri. BFS-4K: An efficient implementation of BFS for Kepler GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 26(7):

1826–1838, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06832649-
abs.html.

Busato:2016:EIB

[BB16] Federico Busato and Nicola Bombieri. An efficient implementation of the Bellman– Ford algorithm for Kepler GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2222–2233, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07287776-
abs.html.

Busato:2017:DAW

[BB17] Federico Busato and Nicola Bombieri. A dynamic approach for workload partitioning on GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 28(6):1535–1549, June 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07750565-
abs.html.

Bachir:2015:JCC

[BBCB15] Abdelmalik Bachir, Walid Bechkit, Yacine Challal, and Abdelmadjid Bouabdallah. Joint connectivitycoverage temperature-aware algorithms for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(7): 1923–1936, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06835181-
abs.html.

Brandwajn:2018:SSM

[BBCTA18] Alexandre Brandwajn, Thomas Begin, Hind Castel-Taleb, and Tulin Atmaca. A study of systems with multiple operating levels, probabilistic thresholds and hysteresis. IEEE Transactions on Parallel and Distributed Systems, 29(4):748–757, April 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08107533-
abs.html.

Beaumont:2019:RAM

[BBD^{+}19] Olivier Beaumont, Brett A.

Becker, Ashley DeFlumere, Lionel Eyraud-Dubois, Thomas Lambert, and Alexey Lastovetsky. Recent advances in matrix partitioning for parallel computing on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 30(1): 218–229, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/01/08404128-
                                        abs.html.

Berrocal:2017:TGS

[BBGD^{+}17] Eduardo Berrocal, Leonardo Bautista-Gomez, Sheng Di, Zhiling Lan, and Franck Cappello. Toward general software level silent data corruption detection for parallel applications. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3642–3655, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/08002625-
                                        abs.html.

Beri:2017:URE

[BBK17] Tarun Beri, Sorav Bansal, and Subodh Kumar. The Unicorn runtime: Efficient distributed shared memory programming for hybrid CPU–GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 28(5):1518–1534, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/05/07588161-
                                        abs.html.

Bridi:2016:CPS

[BBL^{+}16] Thomas Bridi, Andrea Bartolini, Michele Lombardi, Michela Milano, and Luca Benini. A constraint programming scheduler for heterogeneous high-performance computing machines. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2781–2794, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07378987-
abs.html.

Bisson:2016:PDB

[BBM16] Mauro Bisson, Massimo Bernaschi, and Enrico Mastrostefano. Parallel distributed breadth first search on the Kepler architecture. IEEE Transactions on Parallel and Distributed Systems, 27(7):2091–2102, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07234934-
abs.html.

Bondhugula:2017:DTT

[BBP17] Uday Bondhugula, Vinayaka Bandishti, and Irshad Pananilath. Diamond tiling: Tiling techniques to maximize parallelism for stencil computations. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1285–1298, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/05/07582549-
                                        abs.html.

Baldoni:2012:IRR

[BBR12] Roberto Baldoni, Silvia Bonomi, and Michel Raynal. Implementing a regular register in an eventually synchronous distributed system prone to continuous churn. IEEE Transactions on Parallel and Distributed Systems, 23(1):102– 109, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bellavista:2013:EIS

[BCF13] Paolo Bellavista, Antonio Corradi, and Luca Foschini. Enhancing intradomain scalability of IMS-based services. IEEE Transactions on Parallel and Distributed Systems, 24(12):2386–2395, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Becchetti:2014:FTO

[BCP^{+}14] Luca Becchetti, Andrea Clementi, Francesco Pasquale, Giovanni Resta, Paolo Santi, and Riccardo Silvestri. Flooding time in opportunistic networks under power law and exponential intercontact times. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2297–2306, September 2014.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06552196-
abs.html.

Baldoni:2010:CBI

[BCQ^{+}10] Roberto Baldoni, Angelo Corsaro, Leonardo Querzoni, Sirio Scipioni, and Sara Tucci Piergiovanni. Couplingbased internal clock synchronization for large-scale dynamic distributed systems. IEEE Transactions on Parallel and Distributed Systems, 21(5):607–619, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Barooah:2012:CDW

[BCSKN12] Prabir Barooah, Harshavardhan Chenji, Radu Stoleru, and Tamas Kalmar-Nagy. Cut detection in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(3):483– 490, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bartolini:2013:TEM

[BCTB13] Andrea Bartolini, Matteo Cacciari, Andrea Tilli, and Luca Benini. Thermal and energy management of high-performance multicores: Distributed and selfcalibrating model-predictive controller. IEEE Transactions on Parallel and Distributed Systems, 24(1):170– 183, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bruneo:2013:SEQ

[BDLS13] Dario Bruneo, Salvatore Distefano, Francesco Longo, and Marco Scarpa. Stochastic evaluation of QoS in service-based systems. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2090–2099, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Beaumont:2013:HRA

[BEDCR13] Olivier Beaumont, Lionel Eyraud-Dubois, Christopher Thraves Caro, and Hejer Rejeb. Heterogeneous resource allocation under degree constraints. IEEE Transactions on Parallel and Distributed Systems, 24(5): 926–937, May 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bisson:2017:HPE

[BF17] Mauro Bisson and Massimiliano Fatica. High performance exact triangle counting on GPUs. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3501–3510, December 2017.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/08000612-
abs.html.

Besard:2019:EEP

[BFD19] Tim Besard, Christophe Foket, and Bjorn De Sutter. Effective extensible programming: Unleashing Julia on GPUs. IEEE Transactions on Parallel and Distributed Systems, 30(4):827–841, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8471188/

.

Blanton:2011:CNI

[BFFG11] Ethan Blanton, Sonia Fahmy, Greg N. Frederickson, and Sriharsha Gangam. On the cost of network inference mechanisms. IEEE Transactions on Parallel and Distributed Systems, 22(4): 662–672, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bianchi:2010:SDR

[BFPB10] Silvia Bianchi, Pascal Felber, and Maria Gradinariu Potop-Butucaru. Stabilizing distributed R-trees for peer-to-peer content routing. IEEE Transactions on Parallel and Distributed Systems, 21(8):1175–1187, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Balasubramanian:2013:FTD

[BG13] B. Balasubramanian and V. K. Garg. Fault tolerance in distributed systems using fused data structures. IEEE Transactions on Parallel and Distributed Systems, 24(4): 701–715, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

BermudezGarzon:2016:FFT

[BGE^{+}16] Diego F. Bermudez Garzon, Crispin Gomez Requena, Maria Engracia Gomez, Pedro Lo, and Jose Duato. A family of fault-tolerant efficient indirect topologies. IEEE Transactions on Parallel and Distributed Systems, 27(4):927–940, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07103363-
                                        abs.html.

Bai:2016:PRD

[BGHG16] Fan Bai, Feng Gu, Xiaolin Hu, and Song Guo. Particle routing in distributed particle filters for large-scale spatial temporal systems. IEEE Transactions on Parallel and Distributed Systems, 27(2):

481–493, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07046371-
abs.html.

Barsoum:2013:EDD

[BH13] Ayad Barsoum and Anwar Hasan. Enabling dynamic data and indirect mutual trust for cloud computing storage systems. IEEE Transactions on Parallel and Distributed Systems, 24(12): 2375–2385, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bourguiba:2014:INV

[BHEP14] Manel Bourguiba, Kamel Haddadou, Ines El Korbi, and Guy Pujolle. Improving network I/O virtualization for cloud computing. IEEE Transactions on Parallel and Distributed Systems, 25(3): 673–681, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bleuse:2017:SIM

[BHKS^{+}17] Raphael Bleuse, Sascha Hunold, Safia Kedad-Sidhoum, Florence Monna, Gregory Mounie, and Denis Trystram. Scheduling independent moldable tasks on multi-cores with GPUs. IEEE Transactions on Parallel and Distributed Systems, 28(9):2689–2702, September 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07867044-
abs.html.

Bauer:2019:CHP

[BHS^{+}19] André Bauer, Nikolas Herbst, Simon Spinner, Ahmed Ali-Eldin, and Samuel Kounev. Chameleon: A hybrid, proactive auto-scaling mechanism on a level-playing field. IEEE Transactions on Parallel and Distributed Systems, 30(4):800–813, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8465991/

.

Ben-Itzhak:2015:HNR

[BICK^{+}15] Yaniv Ben-Itzhak, Israel Cidon, Avinoam Kolodny, Michael Shabun, and Nir Shmuel. Heterogeneous NoC router architecture. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2479–2492, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06883223.
pdf.

Biswas:2018:ETC

[Bis18] Arnab Kumar Biswas. Efficient timing channel protection for hybrid (packet/ circuit-switched) networkon-chip. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1044–1057, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08207636-
abs.html.

Bandara:2013:CBC

[BJ13] H. M. N. Dilum Bandara and Anura P. Jayasumana. Community-based caching for enhanced lookup performance in P2P systems. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1752–1762, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Baek:2018:NPS

[BJC^{+}18] Hyeongboo Baek, Namyong Jung, Hoon Sung Chwa, Insik Shin, and Jinkyu Lee. Non-preemptive scheduling for mixed-criticality realtime multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 29(8):1766–1779, August 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08292931-
abs.html.

Birk:2016:HSM

[BKF^{+}16] Matthias Birk, Ernst Kretzek, Peter Figuli, Marc Weber, Jurgen Becker, and Nicole V. Ruiter. Highspeed medical imaging in 3D ultrasound computer tomography. IEEE Transactions on Parallel and Distributed Systems, 27(2):455– 467, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07045577-
                                        abs.html.

Berghoff:2018:MPS

[BKH18] Marco Berghoff, Ivan Kondov, and Johannes Hotzer. Massively parallel stencil code solver with autonomous adaptive block distribution. IEEE Transactions on Parallel and Distributed Systems, 29(10):2282–2296, October 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/10/08325488-
                                        abs.html.

Bader:2011:GEI

[BKK11] David A. Bader, David Kaeli, and Volodymyr Kindratenko. Guest Editor’s introduction: Special issue on high-performance computing with accelerators. IEEE Transactions on Parallel and Distributed Systems, 22(1): 3–6, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bruhadeshwar:2011:SKA

[BKL11] Bezawada Bruhadeshwar, Sandeep S. Kulkarni, and Alex X. Liu. Symmetric key approaches to securing BGP — a little bit [of] trust is enough. IEEE Transactions on Parallel and Distributed Systems, 22(9): 1536–1549, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bagci:2015:DFT

[BKY15] Hakki Bagci, Ibrahim Korpeoglu, and Adnan Yazici. A distributed fault-tolerant topology control algorithm for heterogeneous wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(4): 914–923, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06786025-
abs.html.

Bruneo:2015:MEE

[BLLP15] Dario Bruneo, Audric Lhoas, Francesco Longo, and Antonio Puliafito. Modeling and evaluation of energy policies in green clouds. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3052–3065, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/11/06932450-
                                        abs.html.

Bahga:2012:AMM

[BM12] Arshdeep Bahga and Vijay K. Madisetti. Analyzing massive machine maintenance data in a computing cloud. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1831–1843, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bianchi:2010:SOD

[BMB^{+}10] Giuseppe Bianchi, Nicola Blefari Melazzi, Lorenzo Bracciale, Francesca Lo Piccolo, and Stefano Salsano. Streamline: An optimal distribution algorithm for peer-to-peer real-time streaming. IEEE Transactions on Parallel and Distributed Systems, 21(6): 857–871, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Boustani:2017:SGP

[BMJ^{+}17] Arash Boustani, Anindya Maiti, Sina Yousefian Jazi, Murtuza Jadliwala, and Vinod Namboodiri. Seer grid: Privacy and utility implications of two-level load prediction in smart grids. IEEE Transactions on Parallel and Distributed Systems, 28(2):546–557, February 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07466137-
abs.html.

Bera:2015:CCA

[BMR15] Samaresh Bera, Sudip Misra, and Joel J. P. C. Rodrigues. Cloud computing applications for smart grid: A survey. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1477–1494, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06809180-
abs.html.

Balakrishnan:2012:EAM

[BN12] Dineshbalu Balakrishnan and Amiya Nayak. An efficient approach for mobile asset tracking using contexts. IEEE Transactions on Parallel and Distributed Systems, 23(2):211– 218, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bruneo:2014:SMI

[Bru14] Dario Bruneo. A stochastic model to investigate data center performance and QoS in IaaS cloud computing systems. IEEE Transactions on Parallel and Distributed Systems, 25(3):560– 569, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bu:2013:CSC

[BRX13] Xiangping Bu, Jia Rao, and Cheng-Zhong Xu. Coordinated self-configuration of virtual machines and appliances using a model-free learning approach. IEEE Transactions on Parallel and Distributed Systems, 24(4): 681–690, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bulut:2012:EFR

[BS12] Eyuphan Bulut and Boleslaw K. Szymanski. Exploiting friendship relations for efficient routing in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2254–2265, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bulut:2014:CLS

[BS14] Eyuphan Bulut and Boleslaw K. Szymanski. Constructing limited scale-free topologies over peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 25(4):919–928, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bilski:2015:PAL

[BS15] Jaroslaw Bilski and Jacek Smolag. Parallel architectures for learning the RTRN and Elman dynamic neural networks. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2561–2570, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06898879.
pdf.

Benmoussa:2018:JDP

[BSD^{+}18] Yahia Benmoussa, Eric Senn, Nicolas Derouineau, Nicolas Tizon, and Jalil Boukhobza. Joint DVFS and parallelism for energy efficient and low latency software video decoding. IEEE Transactions on Parallel and Distributed Systems, 29(4):858– 872, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08148962-
abs.html.

Bose:2016:HDG

[BSF16] Bella Bose, Arash Shamaei, and Mary Flahive. Higher dimensional Gaussian networks. IEEE Transactions on Parallel and Distributed Systems, 27(9):2628–2638, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/09/07345602-
                                        abs.html.

Buddhika:2017:OSI

[BSL^{+}17] Thilina Buddhika, Ryan Stern, Kira Lindburg, Kathleen Ericson, and Shrideep Pallickara. Online scheduling and interference alleviation for low-latency, highthroughput processing of data streams. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3553–3569, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/07968325-
                                        abs.html.

Briceno:2011:HRR

[BSM^{+}11] Luis Diego Briceno, Howard Jay Siegel, Anthony A. Maciejewski, Mohana Oltikar, Jeff Brateman, Joe White, Jonathan R. Martin, and Keith Knapp. Heuristics for robust resource allocation of satellite weather data processing on a heterogeneous parallel system. IEEE Transactions on Parallel and Distributed Systems, 22(11):1780–1787, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bruneo:2010:PEG

[BSP10] Dario Bruneo, Marco Scarpa, and Antonio Puliafito. Performance evaluation of gLite grids through GSPNs. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1611–1625, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhowmik:2018:ECB

[BTG^{+}18] Sukanya Bhowmik, Muhammad Adnan Tariq, Jonas Grunert, Deepak Srinivasan, and Kurt Rothermel. Expressive content-based routing in software-defined networks. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2460–2477, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08365752-
abs.html.

Bouksiaa:2019:UDE

[BTL^{+}19] M. S. M. Bouksiaa, F. Trahay, A. Lescouet, G. Voron, R. Dulong, A. Guermouche, É. Brunet, and G. Thomas. Using differential execution analysis to identify thread interference. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2866–2878, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bhandari:2010:RBR

[BV10] Vartika Bhandari and Nitin H. Vaidya. Reliable broadcast in radio networks with locally bounded failures. IEEE Transactions on Parallel and Distributed Systems, 21(6): 801–811, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Basanta-Val:2010:SSS

[BVEAGVA10] Pablo Basanta-Val, Iria Estevez-Ayres, Marisol Garcia-Valls, and Luis Almeida. A synchronous scheduling service for distributed realtime Java. IEEE Transactions on Parallel and Distributed Systems, 21(4):506– 519, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Basanta-Val:2017:PDR

[BVFGSFAF17] Pablo Basanta-Val, Norberto Fernandez-Garcia, Luis Sanchez-Fernandez, and Jesus Arias-Fisteus. Patterns for distributed real-time stream processing. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3243–3257, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07953583-
abs.html.

Bao:2019:PMW

[BWB^{+}19] L. Bao, C. Wu, X. Bu, N. Ren, and M. Shen. Performance modeling and workflow scheduling of microservicebased applications in clouds. IEEE Transactions on Parallel and Distributed Systems, 30(9):2114–2129, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Basu:2019:LYG

[BWH^{+}19] D. Basu, X. Wang, Y. Hong, H. Chen, and S. Bressan. Learn-as-you-go with Megh: Efficient live migration of virtual machines. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1786–1801, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bu:2012:EMT

[BXXC12] Kai Bu, Bin Xiao, Qingjun Xiao, and Shigang Chen. Efficient misplaced-tag pinpointing in large RFID systems. IEEE Transactions on Parallel and Distributed Systems, 23(11):2094–2106, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bei:2016:RRF

[BYZ^{+}16] Zhendong Bei, Zhibin Yu, Huiling Zhang, Wen Xiong, Chengzhong Xu, Lieven Eeckhout, and Shengzhong Feng. RFHOC: A randomforest approach to autotuning Hadoop’s configuration. IEEE Transactions on Parallel and Distributed Systems, 27(5):1470–1483, May 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07132754-
abs.html.

Boukerche:2010:CLA

[BZA10] Azzedine Boukerche, Anis Zarrad, and Regina B. Araujo. A cross-layer approach-based Gnutella for collaborative virtual environments over mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 21(7):911– 924, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bui:2010:MAM

[BZBP10] Vinh Bui, Weiping Zhu, Alessio Botta, and Antonio Pescape. A Markovian approach to multipath data transfer in overlay networks. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1398–1411, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Capalija:2013:MCG

[CA13] Davor Capalija and Tarek S.

Abdelrahman. Microarchitecture of a coarse-grain out-of-order superscalar processor. IEEE Transactions on Parallel and Distributed Systems, 24(2):392– 405, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cameron:2019:MMA

[CAC^{+}19] K. W. Cameron, A. Anwar, Y. Cheng, L. Xu, B. Li, U. Ananth, J. Bernard, C. Jearls, T. Lux, Y. Hong, L. T. Watson, and A. R. Butt. MOANA: Modeling and analyzing I/O variability in parallel system experimental design. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1843–1856, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Caheny:2018:RCC

[CAD^{+}18] Paul Caheny, Lluc Alvarez, Said Derradji, Mateo Valero, Miquel Moreto, and Marc Casas. Reducing cache coherence traffic with a NUMA-aware runtime approach. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1174–1187, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08239832-
abs.html.

Charapko:2019:RRM

[CADK19] A. Charapko, A. Ailijiang, M. Demirbas, and S. Kulkarni. Retroscope: Retrospective monitoring of distributed systems. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2582–2594, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chavan:2016:TTA

[CAJ^{+}16] Ajit Chavan, Mohammed I. Alghamdi, Xunfei Jiang, Xiao Qin, Meikang Qiu, Minghua Jiang, and Jifu Zhang. TIGER: Thermalaware file assignment in storage clusters. IEEE Transactions on Parallel and Distributed Systems, 27(2): 558–573, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2016/02/07054504-
abs.html.

Costa:2016:SPC

[CAKRY16] Lauro Beltrao Costa, Samer Al-Kiswany, Matei Ripeanu, and Hao Yang. Support for provisioning and configuration decisions for data intensive workflows. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2725–2739, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07317779-
abs.html.

Cevahir:2011:SBP

[CATC11] Ali Cevahir, Cevdet Aykanat, Ata Turk, and B. Barla Cambazoglu. Site-based partitioning and repartitioning techniques for parallel PageRank computation. IEEE Transactions on Parallel and Distributed Systems, 22(5):786–802, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chard:2013:HPR

[CB13] Kyle Chard and Kris Bubendorfer. High performance resource allocation strategies for computational economies. IEEE Transactions on Parallel and Distributed Systems, 24(1):72–84, January 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Calheiros:2014:MDS

[CB14] Rodrigo N. Calheiros and Rajkumar Buyya. Meeting deadlines of scientific workflows in public clouds with tasks replication. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1787–1796, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Casas:2016:EHA

[CB16] Marc Casas and Greg Bronevetsky. Evaluation of HPC applications memory resource consumption via active measurement. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2560–2573, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/09/07349236-
                                        abs.html.

Chen:2017:ERC

[CBF^{+}17] Mingsong Chen, Yongxiang Bao, Xin Fu, Geguang Pu, and Tongquan Wei. Efficient resource constrained scheduling using parallel two-phase branch-and-bound heuristics. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1299–1314, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07723936-
abs.html.

Choi:2010:RCP

[CBK^{+}10] SungJin Choi, MaengSoon Baik, HongSoo Kim, Eun-Joung Byun, and Hyunseung Choo. A reliable communication protocol for multiregion mobile agent environments. IEEE Transactions on Parallel and Distributed Systems, 21(1):72–85, January 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cardosa:2010:RBU

[CC10] Michael Cardosa and Abhishek Chandra. Resource bundles: Using aggregation for statistical large-scale resource discovery and management. IEEE Transactions on Parallel and Distributed Systems, 21(8):1089–1102, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cai:2013:LTR

[CC13a] Yan Cai and W. K. Chan.

Lock trace reduction for multithreaded programs. IEEE Transactions on Parallel and Distributed Systems, 24(12): 2407–2417, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2013:AAS

[CC13b] Chi-Yeh Chen and Chih-Ping Chu. A 3.42-approximation algorithm for scheduling malleable tasks under precedence constraints. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1479–1488, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chin:2015:LCT

[CC15] Tai-Lin Chin and Wan-Chen Chuang. Latency of collaborative target detection for surveillance sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(2):467– 477, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06674930-
                                        abs.html.

Chen:2017:AAG

[CC17] Jian Chen and Russell M. Clapp. Astro: Autogeneration of synthetic traces using scaling pattern recognition for MPI workloads. IEEE Transactions on Parallel and Distributed Systems, 28(8):2159–2171, August 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/

trans/td/2017/08/07809142-
abs.html.

Chantrapornchai:2018:TQR

[CC18] Chantana Chantrapornchai and Chidchanok Choksuchat. TripleID-Q: RDF query processing framework using GPU. IEEE Transactions on Parallel and Distributed Systems, 29(9):2121–2135, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08314130.
pdf.

Cheng:2016:MCR

[CCC^{+}16] Sheng-Wei Cheng, Che-Wei Chang, Jian-Jia Chen, Tei-Wei Kuo, and Pi-Cheng Hsiu. Many-core real-time task scheduling with scratchpad memory. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2953–2966, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07378514-
abs.html.

Cho:2014:DTM

[CCCB14] Jin-Hee Cho, MoonJeong Chang, Ing-Ray Chen, and Fenye Bao. Dynamic trust management for delay tolerant networks and its application to secure routing. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1200–1210, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2016:PCB

[CCCY16] Yang Chen, Yu Chen, Qiang Cao, and Xiaowei Yang. PacketCloud: A cloudletbased open platform for innetwork services. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1146–1159, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07088634-
                                        abs.html.

Conti:2015:LPC

[CCD^{+}15] Mauro Conti, Bruno Crispo, Daniele Diodati, Jukka K. Nurminen, Cristina M. Pinotti, and Taavi Teemaa. Leveraging parallel communications for minimizing energy consumption on Smartphones. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2778–2790, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06912987.
                                        pdf.

Calamoneri:2011:MNB

[CCFS11] Tiziana Calamoneri, Andrea E. F. Clementi, Emanuele G.

Fusco, and Riccardo Silvestri. Maximizing the number of broadcast operations in random geometric ad hoc wireless networks. IEEE Transactions on Parallel and Distributed Systems, 22(2): 208–216, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2017:PDA

[CCH^{+}17] Yu-Yin Chen, En-Jui Chang, Hsien-Kai Hsin, Kun-Chih Jimmy Chen, and An-Yeu Andy Wu. Path-diversity-aware faulttolerant routing algorithm for network-on-chip systems. IEEE Transactions on Parallel and Distributed Systems, 28(3):838–849, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07506215-
abs.html.

Chen:2019:MMO

[CCHH19] Xiaoming Chen, Danny Ziyi Chen, Yinhe Han, and Xiaobo Sharon Hu. moDNN: Memory optimal deep neural network training on graphics processing units. IEEE Transactions on Parallel and Distributed Systems, 30(3): 646–661, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2019/03/08444714-
                                        abs.html.

Cai:2019:RED

[CCJ19] H. Cai, Q. Cao, and H. Jiang. A renewable energy driven approach for computational sprinting. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1449–1463, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Choudhury:2012:OSD

[CCK12] Pravanjan Choudhury, P. P. Chakrabarti, and Rajeev Kumar. Online scheduling of dynamic task graphs with communication and contention for multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 23(1): 126–133, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2015:RTT

[CCKF15] Che-Wei Chang, Jian-Jia Chen, Tei-Wei Kuo, and Heiko Falk. Real-time task scheduling on island-based multi-core platforms. IEEE Transactions on Parallel and Distributed Systems, 26(2): 538–550, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06714441-
                                        abs.html.

Chen:2013:SSR

[CCL13] Xi Chen, Xiao-Wen Chang, and Xue Liu. SyRaFa: Synchronous rate and frequency adjustment for utilization control in distributed real-time embedded systems. IEEE Transactions on Parallel and Distributed Systems, 24(5):1052–1061, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2011:QOB

[CCLW11] Cheng-Shang Chang, Jay Cheng, Duan-Shin Lee, and Chi-Feung Wu. Quasioutput-buffered switches. IEEE Transactions on Parallel and Distributed Systems, 22(5):833–846, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:RBT

[CCLW15] Kun-Chih Chen, En-Jui Chang, Huai-Ting Li, and An-Yeu Andy Wu. RCbased temperature prediction scheme for proactive dynamic thermal management in throttle-based 3D NoCs. IEEE Transactions on Parallel and Distributed Systems, 26(1):206– 218, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2015/01/06748042-
abs.html.

Chakaravarthy:2017:SSS

[CCM^{+}17] Venkatesan T. Chakaravarthy, Fabio Checconi, Prakash Murali, Fabrizio Petrini, and Yogish Sabharwal. Scalable single source shortest path algorithms for massively parallel systems. IEEE Transactions on Parallel and Distributed Systems, 28(7): 2031–2045, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07763835-
abs.html.

Clemente-Castello:2018:PMM

[CCNMF18] Francisco J. Clemente-Castello, Bogdan Nicolae, Rafael Mayo, and Juan Carlos Fernandez. Performance model of MapReduce iterative applications for hybrid cloud bursting. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1794–1807, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08283575-
abs.html.

Cheng:2012:DDR

[CCS^{+}12] Wei Cheng, Xiuzhen Cheng, Min Song, Biao Chen, and Wendy W. Zhao. On the design and deployment of RFID assisted navigation systems for VANETs. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1267–1274, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2010:ISE

[CCT10] Shu Chen, Yingying Chen, and Wade Trappe. Inverting systems of embedded sensors for position verification in location-aware applications. IEEE Transactions on Parallel and Distributed Systems, 21(5):722–736, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chu:2014:KAC

[CCT^{+}14] Cheng-Kang Chu, Sherman S. M. Chow, Wen-Guey Tzeng, Jianying Zhou, and Robert H. Deng. Keyaggregate cryptosystem for scalable data sharing in cloud storage. IEEE Transactions on Parallel and Distributed Systems, 25(2):468– 477, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2016:DEG

[CCT16] Lien-Wu Chen, Jen-Hsiang Cheng, and Yu-Chee Tseng. Distributed emergency guiding with evacuation time optimization based on wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27(2): 419–427, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07328748-
                                        abs.html.

Cools:2019:NSR

[CCV19] S. Cools, J. Cornelis, and W. Vanroose. Numerically stable recurrence relations for the communication hiding pipelined conjugate gradient method. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2507–2522, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cong:2012:SAT

[CCW^{+}12] Guojing Cong, I-Hsin Chung, Hui-Fang Wen, David Klepacki, Hiroki Murata, Yasushi Negishi, and Takao Moriyama. A systematic approach toward automated performance analysis and tuning. IEEE Transactions on Parallel and Distributed Systems, 23(3):426–435, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Choi:2013:IUH

[CD13] Woojin Choi and Jeffrey Draper. Improving utilization of hardware signatures in transactional memory. IEEE Transactions on Parallel and Distributed Systems, 24(11):2230–2239, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chai:2012:EDM

[CDBQ12] Yunpeng Chai, Zhihui Du, David A. Bader, and Xiao Qin. Efficient data migration to conserve energy in streaming media storage systems. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2081–2093, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chittamuru:2018:BAB

[CDPM18] Sai Vineel Reddy Chittamuru, Dharanidhar Dang, Sudeep Pasricha, and Rabi Mahapatra. BiGNoC: Accelerating big data computing with application-specific photonic network-on-chip architectures. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2402–2415, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08356088-
abs.html.

Cordasco:2015:AOH

[CDR15] Gennaro Cordasco, Rosario De Chiara, and Arnold L. Rosenberg. An AREAoriented heuristic for scheduling DAGs on volatile computing platforms. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2164–2177, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/08/06873341-
                                        abs.html.

Chae:2015:TMD

[CDS15] Younghun Chae, Lisa Cingiser DiPippo, and Yan Lindsay Sun. Trust management for defending on-off attacks. IEEE Transactions on Parallel and Distributed Systems, 26(4):1178–1191, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06800056-
                                        abs.html.

Chiu:2010:PCD

[CE10] Yuh-Ming Chiu and Do Young Eun. On the performance of content delivery under competition in a stochastic unstructured peer-to-peer network. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1487–1500, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Carlsson:2017:ECP

[CE17] Niklas Carlsson and Derek Eager. Ephemeral content popularity at the edge and implications for ondemand caching. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1621–1634, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07581058-
abs.html.

Chretien:2016:GRM

[CEK16] Benjamin Chretien, Adrien Escande, and Abderrahmane Kheddar. GPU robot motion planning using semiinfinite nonlinear programming. IEEE Transactions on Parallel and Distributed Systems, 27(10):2926–2939, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07390287-
abs.html.

Cheng:2015:DPB

[CFJ15] Baolei Cheng, Jianxi Fan, and Xiaohua Jia. Dimensionalpermutation-based independent spanning trees in bijective connection networks. IEEE Transactions on Parallel and Distributed Systems, 26(1):45–53, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06747363-
                                        abs.html.

Chen:2018:EPC

[CFLL18] Li Chen, Yuan Feng, Baochun Li, and Bo Li. Efficient performance-centric bandwidth allocation with fairness tradeoff. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1693–1706, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/08/08295233-
                                        abs.html.

Chen:2019:PPF

[CFLL19] L. Chen, Y. Feng, B. Li, and B. Li. Promenade: Proportionally fair multipath rate control in datacenter networks with random network coding. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2536–2546, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2013:ACA

[CGH13] Quan Chen, Minyi Guo, and Zhiyi Huang. Adaptive cache aware bitier work-stealing in multisocket multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 24(12):2334–2343, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Czyzowicz:2011:CME

[CGKP11] Jurek Czyzowicz, Leszek Gasieniec, Dariusz R. Kowalski, and Andrzej Pelc. Consensus and mutual exclusion in a multiple access channel. IEEE Transactions on Parallel and Distributed Systems, 22(7):1092–1104, July 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cabezas:2015:RAS

[CGS^{+}15] Javier Cabezas, Isaac Gelado, John E. Stone, Nacho Navarro, David B. Kirk, and Wen mei Hwu. Runtime and architecture support for efficient data exchange in multi-accelerator applications. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1405–1418, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803940-
abs.html.

Chen:2013:IRM

[CGZQ13] Zhezhe Chen, Qi Gao, Wenbin Zhang, and Feng Qin. Improving the reliability of MPI libraries via message flow checking. IEEE Transactions on Parallel and Distributed Systems, 24(3):535– 549, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2013:CBI

[CH13] Shih Heng Cheng and Ching Yao Huang. Coloringbased inter-WBAN scheduling for mobile wireless body area networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):250– 259, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2014:SPC

[CH14] Nai-Wen Chang and Sun-Yuan Hsieh. Structural properties and conditional diagnosability of star graphs by using the PMC model. IEEE Transactions on Parallel and Distributed Systems, 25(11): 3002–3011, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06671606-
                                        abs.html.

Cheng:2015:FTC

[CH15] Chia-Wen Cheng and Sun-Yuan Hsieh. Fault-tolerant cycle embedding in Cartesian product graphs: Edgepancyclicity and edge-bipancyclicity with faulty edges. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2997–3011, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06935086-
abs.html.

Chang:2011:CDU

[Cha11] Guey-Yun Chang. Conditional (t, k)-diagnosis under the PMC model. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1797–1803, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chan:2014:MPR

[Cha14] Sammy Chan. Multi-path routing and forwarding in non-cooperative wireless networks. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2638–2647, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06577364-
abs.html.

Cao:2014:STS

[CHCC14] Xianghui Cao, Jianping He, Jiming Chen, and Peng Cheng. Secure time synchronization in wireless sensor networks: A maximum consensus-based approach.

IEEE Transactions on Parallel and Distributed Systems, 25(4):1055–1065, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:DSN

[CHD^{+}15] Jing Chen, Kun He, Ruiying Du, Minghui Zheng, Yang Xiang, and Quan Yuan. Dominating set and network coding-based routing in wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 26(2): 423–433, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06678512-
                                        abs.html.

Chen:2011:UBS

[Che11] Ray Jinzhu Chen. An upper bound solution for homogeneous fork/join queuing systems. IEEE Transactions on Parallel and Distributed Systems, 22(5):874– 878, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2014:DCA

[Che14] Yixin Chen. Distributed channel allocation protocols for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(9):

2264–2274, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06570720-
abs.html.

Chen:2015:DCO

[Che15] Xu Chen. Decentralized computation offloading game for mobile cloud computing. IEEE Transactions on Parallel and Distributed Systems, 26(4):974–983, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06787113-
abs.html.

Chen:2016:TSM

[Che16] Chi-Yeh Chen. Task scheduling for maximizing performance and reliability considering fault recovery in heterogeneous distributed systems. IEEE Transactions on Parallel and Distributed Systems, 27(2):521– 532, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07042339-
abs.html.

Chen:2018:IAS

[Che18a]  Chi-Yeh Chen.  An im-

proved approximation for scheduling malleable tasks with precedence constraints via iterative method. IEEE Transactions on Parallel and Distributed Systems, 29(9): 1937–1946, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08315149.
                                        pdf.

Chen:2018:GBT

[Che18b] Kun-Chih Chen. Game-Based Thermal-Delay-Aware Adaptive Routing (GTDAR) for temperature-aware 3D network-on-chip systems. IEEE Transactions on Parallel and Distributed Systems, 29(9):2018–2032, September 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08306659-
                                        abs.html.

Chen:2019:DPE

[CHHK19] C. Chen, T. Hsia, Y. Huang, and S. Kuo. Data prefetching and eviction mechanisms of in-memory storage systems based on scheduling for big data processing. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1738–1752, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2015:LLB

[CHLC15] Chia-Wei Chang, Guanyao Huang, Bill Lin, and Chen-Nee Chuah. LEISURE: Load-balanced network-wide traffic measurement and monitor placement. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1059–1070, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06573959-
abs.html.

Cui:2018:SPA

[CHLY18] Lei Cui, Zhiyu Hao, Lun Li, and Xiaochun Yun. SnapFiner: A page-aware snapshot system for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2613–2626, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08352722-
abs.html.

Cao:2013:OMC

[CHLZ13] Junwei Cao, Kai Hwang, Keqin Li, and Albert Y. Zomaya. Optimal multiserver configuration for profit maximization in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 24(6):1087–1096, June 2013. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Carrillo:2013:SHN

[CHM^{+}13] Snaider Carrillo, Jim Harkin, Liam J. McDaid, Fearghal Morgan, Sandeep Pande, Seamus Cawley, and Brian McGinley. Scalable hierarchical network-on-chip architecture for spiking neural network hardware implementations. IEEE Transactions on Parallel and Distributed Systems, 24(12):2451–2461, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2017:PFE

[CHPY17] Lei Cui, Zhiyu Hao, Yaqiong Peng, and Xiaochun Yun. Piccolo: A fast and efficient rollback system for virtual machine clusters. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2328–2341, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07852513-
                                        abs.html.

Chen:2012:CDC

[CHTW12] Siyuan Chen, Minsu Huang, Shaojie Tang, and Yu Wang. Capacity of data collection in arbitrary wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(1):52– 60, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2017:HPH

[CHW^{+}17] Dan Chen, Yangyang Hu, Lizhe Wang, Albert Y. Zomaya, and Xiaoli Li. H-PARAFAC: Hierarchical parallel factor analysis of multidimensional big data. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1091–1104, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07575712-
abs.html.

Chou:2013:ECR

[CIH13] Chun Tung Chou, Aleksandar Ignjatovic, and Wen Hu. Efficient computation of robust average of compressive sensing data in wireless sensor networks in the presence of sensor faults. IEEE Transactions on Parallel and Distributed Systems, 24(8):1525–1534, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2017:USP

[CIP^{+}17] Shaoming Chen, Samuel Irving, Lu Peng, Yue Hu, Ying Zhang, and Ashok Srivastava. Using switchable pins to increase offchip bandwidth in chipmultiprocessors. IEEE Transactions on Parallel and Distributed Systems, 28(1): 274–289, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07440872-
abs.html.

Canon:2010:EOR

[CJ10] Louis-Claude Canon and Emmanuel Jeannot. Evaluation and optimization of the robustness of DAG schedules in heterogeneous environments. IEEE Transactions on Parallel and Distributed Systems, 21(4):532– 546, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Canon:2016:CAH

[CJ16] Louis-Claude Canon and Emmanuel Jeannot. Correlationaware heuristics for evaluating the distribution of the longest path length of a DAG with random weights. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3158–3171, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07405345-
abs.html.

Chen:2016:PPP

[CJBW16] Xinming Chen, Brandon Jones, Michela Becchi, and Tilman Wolf. Picking pesky parameters: Optimizing regular expression matching in practice. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1430–1442, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07152954-
abs.html.

Cai:2014:DAD

[CJH^{+}14] Zhipeng Cai, Shouling Ji, Jing He, Lin Wei, and Anu G. Bourgeois. Distributed and asynchronous data collection in cognitive radio networks with fairness consideration. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2020–2029, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2012:BEE

[CJL^{+}12] Hanhua Chen, Hai Jin, Xucheng Luo, Yunhao Liu, Tao Gu, Kaiji Chen, and Lionel M. Ni. BloomCast: Efficient and effective full-text retrieval in unstructured P2P networks. IEEE Transactions on Parallel and Distributed Systems, 23(2):232– 241, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cai:2015:ADB

[CJW^{+}15] Yan Cai, Changjiang Jia, Shangru Wu, Ke Zhai, and Wing Kwong Chan. ASN: A dynamic barrier-based approach to confirmation of deadlocks from warnings for large-scale multithreaded programs. IEEE Transactions on Parallel and Distributed Systems, 26 (1):13–23, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06747310-
                                        abs.html.

Chen:2016:MIS

[CJW16] Hanhua Chen, Hai Jin, and Shaoliang Wu. Minimizing inter-server communications by exploiting self-similarity in online social networks. IEEE Transactions on Parallel and Distributed Systems, 27(4):1116–1130, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07097076-
                                        abs.html.

Cheng:2019:UHB

[CJW^{+}19] Y. Cheng, H. Jiang, F. Wang, Y. Hua, D. Feng, W. Guo, and Y. Wu. Using highbandwidth networks efficiently for fast graph computation. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1170–1183, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2012:REC

[CJZ12] Yang Cao, Tao Jiang, and Qian Zhang. Reducing electricity cost of smart appliances via energy buffering framework in Smart Grid. IEEE Transactions on Parallel and Distributed Systems, 23(9):1572–1582, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2016:FPD

[CJZ^{+}16] Gang Chen, Hai Jin, Deqing Zou, Zhenkai Liang, Bing Bing Zhou, and Hao Wang. A framework for practical dynamic software updating. IEEE Transactions on Parallel and Distributed Systems, 27(4):941– 950, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07103360-
abs.html.

Cai:2013:TBE

[CL13] Qing-Chao Cai and Kwok-Tung Lo. Two blocks are enough: On the feasibility of using network coding to ameliorate the content availability of BitTorrent swarms.

IEEE Transactions on Parallel and Distributed Systems, 24(8):1682–1694, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2014:EDI

[CL14] Henry C. H. Chen and Patrick P. C. Lee. Enabling data integrity protection in regenerating-coding-based cloud storage: Theory and implementation. IEEE Transactions on Parallel and Distributed Systems, 25(2): 407–416, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2015:SSC

[CL15] Hsueh-Hung Cheng and Kate Ching-Ju Lin. Source selection and content dissemination for preference-aware traffic offloading. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3160–3174, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/11/06928487-
                                        abs.html.

Cao:2016:CHC

[CL16a] Zhengjun Cao and Lihua Liu. Comment on ‘Harnessing the Cloud for Securely Outsourcing Large-Scale Systems of Linear Equations’.

IEEE Transactions on Parallel and Distributed Systems, 27(5):1551–1552, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07449034-

abs.html. See [WRWW13].

Cheng:2016:OIL

[CL16b] Luwei Cheng and Francis C. M. Lau. Offloading interrupt load balancing from SMP virtual machines to the hypervisor. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3298–3310, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07425234-
abs.html.

Champati:2017:SOA

[CL17] Jaya Prakash Champati and Ben Liang. Semi-online algorithms for computational task offloading with communication delay. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1189–1201, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07559817-
abs.html.

Chu:2019:EHM

[CLA^{+}19] Ching-Hsiang Chu, Xiaoyi Lu, Ammar A. Awan, Hari Subramoni, Bracy Elton, and Dhabaleswar K. Panda. Exploiting hardware multicast and GPUDirect RDMA for efficient broadcast. IEEE Transactions on Parallel and Distributed Systems, 30(3): 575–588, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08447246-
abs.html.

Chen:2019:BLP

[CLB^{+}19] J. Chen, K. Li, K. Bilal, x. zhou, K. Li, and P. S. Yu. A bi-layered parallel training architecture for large-scale convolutional neural networks. IEEE Transactions on Parallel and Distributed Systems, 30(5): 965–976, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2012:PRM

[CLC^{+}12] Yunji Chen, Lei Li, Tianshi Chen, Ling Li, Lei Wang, Xiaoxue Feng, and Weiwu Hu. Program regularization in memory consistency verification. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2163–2174, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2013:CEF

[CLH13] Chia-Wen Cheng, Chia-Wei Lee, and Sun-Yuan Hsieh. Conditional edgefault Hamiltonicity of Cartesian product graphs. IEEE Transactions on Parallel and Distributed Systems, 24(10): 1951–1960, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2014:SOA

[CLH^{+}14] Xiaofeng Chen, Jin Li, Xinyi Huang, Jingwei Li, Yang Xiang, and Duncan S. Wong. Secure outsourced attributebased signatures. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3285–3294, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714536-
abs.html.

Chu:2011:SAE

[CLHK11] Edward T.-H. Chu, Hsin-Ju Lee, Tai-Yi Huang, and Chung-Ta King. Sample assignment for ensuring sensing quality and balancing energy in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(9):1578–1584, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2013:TAA

[CLHW13] Kun-Chih Chen, Shu-Yen Lin, Hui-Shun Hung, and An-Yeu Andy Wu. Topologyaware adaptive routing for nonstationary irregular mesh in throttled 3D NoC systems. IEEE Transactions on Parallel and Distributed Systems, 24(10):2109–2120, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chou:2011:OAM

[CLJ11] Zi-Tsan Chou, Yu-Hsiang Lin, and Rong-Hong Jan. Optimal asymmetric and maximized adaptive power management protocols for clustered ad hoc wireless networks. IEEE Transactions on Parallel and Distributed Systems, 22(12): 1961–1968, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chou:2015:ERE

[CLKR15] Jerry Chi-Yuan Chou, Ting-Hsuan Lai, Jinoh Kim, and Doron Rotem. Exploiting replication for energy-aware scheduling in disk storage systems. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2734–2749, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06902814.
pdf.

Chen:2011:CIW

[CLL11] Lin Chen, Lavy Libman, and Jean Leneutre. Conflicts and incentives in wireless cooperative relaying: a distributed market pricing framework. IEEE Transactions on Parallel and Distributed Systems, 22(5):758–772, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2014:GEI

[CLL^{+}14] Zhenfu Cao, Keqiu Li, Xu Li, Patrick McDaniel, Radha Poovendran, Guojun Wang, and Yang Xiang. Guest Editors’ introduction: Special issue on trust, security, and privacy in parallel and distributed systems. IEEE Transactions on Parallel and Distributed Systems, 25(2): 279–282, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chwa:2017:GES

[CLL^{+}17] Hoon Sung Chwa, Jinkyu Lee, Jiyeon Lee, Kiew-My Phan, Arvind Easwaran, and Insik Shin. Global EDF schedulability analysis for parallel tasks on multi-core platforms. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1331–1345, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/05/07580608-
                                        abs.html.

Chen:2019:EDC

[CLL^{+}19] Y. Chen, C. Li, M. Lv, X. Shao, Y. Li, and Y. Xu. Explicit data correlationsdirected metadata prefetching method in distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2692–2705, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2012:RRC

[CLLS12] Chien-Ming Chen, Yue-Hsun Lin, Ya-Ching Lin, and Hung-Min Sun. RCDA: Recoverable concealed data aggregation for data integrity in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):727–734, April 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chan:2018:EPL

[CLLX18] Helen H. W. Chan, Yongkun Li, Patrick P. C. Lee, and Yinlong Xu. Elastic parity logging for SSD RAID arrays: Design, analysis, and implementation. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2241–2253, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/10/08322269-
abs.html.

Cammarano:2015:TOC

[CLM^{+}15] Alessandro Cammarano, Francesco Lo Presti, Gaia Maselli, Loreto Pescosolido, and Chiara Petrioli. Throughputoptimal cross-layer design for cognitive radio ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2599–2609, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06881740.
pdf.

Chen:2018:GMC

[CLO^{+}18] Cen Chen, Kenli Li, Aijia Ouyang, Zeng Zeng, and Keqin Li. GFlink: An inmemory computing architecture on heterogeneous CPU– GPU clusters for big data. IEEE Transactions on Parallel and Distributed Systems, 29(6):1275–1288, June 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2018/06/08259455-
                                        abs.html.

Choi:2012:DDA

[CLSZ12] Bong Jun Choi, Hao Liang, Xuemin (Sherman) Shen, and Weihua Zhuang. DCS: Distributed asynchronous clock synchronization in delay tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 23(3): 491–504, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2013:ORT

[CLT13] Ya-Shu Chen, Han Chiang Liao, and Ting-Hao Tsai. Online real-time task scheduling in heterogeneous multicore systemon-a-chip. IEEE Transactions on Parallel and Distributed Systems, 24(1):118– 130, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2017:PRF

[CLT^{+}17] Jianguo Chen, Kenli Li, Zhuo Tang, Kashif Bilal, Shui Yu, Chuliang Weng, and Keqin Li. A parallel random forest algorithm for big data in a spark cloud computing environment. IEEE Transactions on Parallel and Distributed Systems, 28(4): 919–933, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07557062-
abs.html.

Chen:2019:PAM

[CLY^{+}19] Yuedan Chen, Kenli Li, Wangdong Yang, Guoqing Xiao, Xianghui Xie, and Tao Li. Performance-aware model for sparse matrix– matrix multiplication on the Sunway TaihuLight supercomputer. IEEE Transactions on Parallel and Distributed Systems, 30(4):923– 938, April 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8468040/

.

Chen:2016:QIH

[CLYR16] Shaomiao Chen, Zhiyong Li, Bo Yang, and Gunter Rudolph. Quantum-inspired hyper-heuristics for energyaware scheduling on heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 27(6):1796–1810, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07173041-
abs.html.

Cong:2018:DUP

[CLZ^{+}18] Peijin Cong, Liying Li, Junlong Zhou, Kun Cao, Tongquan Wei, Mingsong Chen, and Shiyan Hu. Developing user perceived value based pricing models for cloud markets. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2742–2756, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08370902-
abs.html.

Cho:2010:IPP

[CM10] Sangyeun Cho and Rami G. Melhem. On the interplay of parallelization, program performance, and energy consumption. IEEE Transactions on Parallel and Distributed Systems, 21(3):342– 353, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Camarero:2015:LGH

[CMB15] Cristobal Camarero, Carmen Martinez, and Ramon Beivide. Lattice graphs for high-scale interconnection topologies. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2506–2519, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06894180.
pdf.

Camarero:2018:RWP

[CMB18] Cristobal Camarero, Carmen Martinez, and Ramon Beivide. On random wiring in practicable folded Clos networks for modern datacenters. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1780–1793, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08290692-
abs.html.

Cui:2015:CCE

[CMC^{+}15] Yong Cui, Xiao Ma, Xiuzhen Cheng, Minming Li, Jiangchuan Liu, Tianze Ma, Yihua Guo, and Biao Chen. Cooperative coverage extension for relay-union networks. IEEE Transactions on Parallel and Distributed Systems, 26(2):371– 381, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06748021-
abs.html.

Ciullo:2014:PAV

[CMG^{+}14] Delia Ciullo, Valentina Martina, Michele Garetto, Emilio Leonardi, and Giovanni Luca Torrisi. Peer-assisted VoD systems: An efficient modeling framework. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1852–1863, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2017:DDS

[CMG17] Tianyi Chen, Antonio G. Marques, and Georgios B. Giannakis. DGLB: Distributed stochastic geographical load balancing over cloud networks. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1866–1880, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/07/07775032-
                                        abs.html.

Cheng:2016:FCL

[CMK^{+}16] Long Cheng, Avinash Malik, Spyros Kotoulas, Tomas E. Ward, and Georgios Theodoropoulos. Fast compression of large Semantic Web data using X10. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2603–2617, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/09/07313002-
                                        abs.html.

Clementi:2011:ISS

[CMPS11] Andrea Clementi, Angelo Monti, Francesco Pasquale, and Riccardo Silvestri. Information spreading in stationary Markovian evolving graphs. IEEE Transactions on Parallel and Distributed Systems, 22(9):1425–1432, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chechina:2017:ESD

[CMT^{+}17] Natalia Chechina, Kenneth MacKenzie, Simon Thompson, Phil Trinder, Olivier Boudeville, Viktoria Fordos, Csaba Hoch, Amir Ghaffari, and Mario Moro Hernandez. Evaluating scalable distributed Erlang for scalability and reliability. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2244–2257, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07820204.
pdf.

Camara:2010:TTT

[CMV^{+}10] Jose M. Camara, Miquel Moreto, Enrique Vallejo, Ramon Beivide, Jose Miguel-Alonso, Carmen Martinez, and Javier Navaridas. Twisted torus topologies for enhanced interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 21(12):1765–1778, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Camarero:2017:PNT

[CMVB17] Cristobal Camarero, Carmen Martinez, Enrique Vallejo, and Ramon Beivide. Projective networks: Topologies for large parallel computer systems. IEEE Transactions on Parallel and Distributed Systems, 28(7): 2003–2016, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/07/07769228-
                                        abs.html.

Cheng:2014:QAG

[CNC^{+}14] Long Cheng, Jianwei Niu, Jiannong Cao, Sajal K. Das, and Yu Gu. QoS aware geographic opportunistic routing in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1864–1875, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Carra:2011:RBS

[CNMA11] Damiano Carra, Giovanni Neglia, Pietro Michiardi, and Francesco Albanese. On the robustness of BitTorrent swarms to greedy peers. IEEE Transactions on Parallel and Distributed Systems, 22(12):2071–2078, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:TBQ

[CP15] Wuhui Chen and Incheon Paik. Toward better quality of service composition based on a global social service network. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1466–1476, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06807777-
abs.html.

Canon:2017:HBC

[CP17a] Louis-Claude Canon and Laurent Philippe. On the heterogeneity bias of cost matrices for assessing scheduling algorithms. IEEE Transactions on Parallel and Distributed Systems, 28(6):1675–1688, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07745973-
abs.html.

Chen:2017:CGH

[CP17b] Ren Chen and Viktor K.

Prasanna. Computer generation of high throughput and memory efficient sorting designs on FPGA. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3100–3113, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/11/07930437-
                                        abs.html.

Choi:2017:NWA

[CP17c] Ju-Hee Choi and Gi-Ho Park. NVM way allocation scheme to reduce NVM writes for hybrid cache architecture in chipmultiprocessors. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2896–2910, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07889038-
                                        abs.html.

Chen:2014:CTO

[CPGT14] Keke Chen, James Powers, Shumin Guo, and Fengguang Tian. CRESP: Towards optimal resource provisioning for MapReduce computing in public clouds. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1403–1412, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Colaso:2018:MHC

[CPH^{+}18] Adrian Colaso, Pablo Prieto, Jose Angel Herrero, Pablo Abad, Lucia G. Menezo, Valentin Puente, and Jose Angel Gregorio. Memory hierarchy characterization of NoSQL applications through full-system simulation. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1161–1173, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08239962-
abs.html.

Cui:2018:MMP

[CPL^{+}18] Yingbo Cui, Shaoliang Peng, Yutong Lu, Xiaoqian Zhu, Bingqiang Wang, Chengkun Wu, and Xiangke Liao. mSNP: A massively parallel algorithm for largescale SNP detection. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2557–2567, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08362678-
abs.html.

Carbunar:2010:SSP

[CPM^{+}10] Bogdan Carbunar, Michael Pearce, Shivajit Mohapatra, Loren J. Rittle, Venu Vasudevan, and Octavian Carbunar. Secure synchronization of periodic updates in ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1060–1073, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2012:FDS

[CQZ^{+}12] Shan Chang, Yong Qi, Hongzi Zhu, Jizhong Zhao, and Xuemin (Sherman) Shen. Footprint: Detecting Sybil attacks in urban vehicular networks. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1103–1114, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chronaki:2017:TST

[CRC^{+}17] Kallia Chronaki, Alejandro Rico, Marc Casas, Miquel Moreto, Rosa M. Badia, Eduard Ayguade, Jesus Labarta, and Mateo Valero. Task scheduling techniques for asymmetric multi-core systems. IEEE Transactions on Parallel and Distributed Systems, 28(7):2074–2087, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/07/07762236-
                                        abs.html.

Cuesta:2011:ESS

[CRD11] Blas Cuesta, Antonio Robles, and Jose Duato. Efficient and scalable starvation prevention mechanism for token coherence. IEEE Transactions on Parallel and Distributed Systems, 22(10):1610–1623, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2017:IPH

[CRG^{+}17] Dazhao Cheng, Jia Rao, Yanfei Guo, Changjun Jiang, and Xiaobo Zhou. Improving performance of heterogeneous MapReduce clusters with adaptive task tuning. IEEE Transactions on Parallel and Distributed Systems, 28(3):774–786, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07523426-
abs.html.

Couceiro:2015:COR

[CRRR15] Maria Couceiro, Pedro Ruivo, Paolo Romano, and Luis Rodrigues. Chasing the optimum in replicated inmemory transactional platforms via protocol adaptation. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2942–2955, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06926806-
abs.html.

Chatzikonstantis:2017:OEH

[CRS^{+}17] George Chatzikonstantis, Dimitrios Rodopoulos, Christos Strydis, Chris I. De Zeeuw, and Dimitrios Soudris. Optimizing extended Hodgkin– Huxley neuron model simulations for a Xeon/Xeon Phi node. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2581–2594, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07884986-
                                        abs.html.

Chen:2015:GAS

[CRWY15] Xiaoming Chen, Ling Ren, Yu Wang, and Huazhong Yang. GPU-accelerated sparse LU factorization for circuit simulation with performance modeling. IEEE Transactions on Parallel and Distributed Systems, 26 (3):786–795, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06774937-
                                        abs.html.

Chiang:2015:SEV

[CRZH15] Ron C. Chiang, Sundaresan Rajasekaran, Nan Zhang, and H. Howie Huang. Swiper: Exploiting virtual machine vulnerability in third-party clouds with competition for I/O resources. IEEE Transactions on Parallel and Distributed Systems, 26(6):1732–1742, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06824231-
abs.html.

Chang:2016:CCB

[CSC16] Cheng Chang, Jianhua Sun, and Hao Chen. Coral: A cloud-backed frugal file system. IEEE Transactions on Parallel and Distributed Systems, 27(4):978– 991, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07089277-
abs.html.

Carra:2013:CMP

[CSM^{+}13] D. Carra, M. Steiner, P. Michiardi, E. W. Biersack, W. Effelsberg, and T. En-Najjary. Characterization and management of popular content in KAD. IEEE Transactions on Parallel and Distributed Systems, 24(4): 662–671, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2013:SPC

[CSP13] Gang Chen, Abdolhossein Sarrafzadeh, and Shaoning Pang. Service provision control in federated service providing systems. IEEE Transactions on Parallel and Distributed Systems, 24(3):587– 600, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chung:2017:PDN

[CSR^{+}17] I-Hsin Chung, Tara N. Sainath, Bhuvana Ramabhadran, Michael Picheny, John Gunnels, Vernon Austel, Upendra Chauhari, and Brian Kingsbury. Parallel deep neural network training for big data on Blue Gene/Q. IEEE Transactions on Parallel and Distributed Systems, 28(6):1703–1714, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/06/07738586-
                                        abs.html.

Chen:2013:FRS

[CSS^{+}13] Hao Chen, Lin Shi, Jianhua Sun, Kenli Li, and Ligang He. A fast RPC system for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1267–1276, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:SNB

[CSSL15] Kang Chen, Haiying Shen, Karan Sapra, and Guoxin Liu. A social network based reputation system for cooperative P2P file sharing. IEEE Transactions on Parallel and Distributed Systems, 26(8):2140–2153, August 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06873304-
abs.html.

Ceriani:2017:EEH

[CSV^{+}17] Marco Ceriani, Simone Secchi, Oreste Villa, Antonino Tumeo, and Gianluca Palermo. Exploring efficient hardware support for applications with irregular memory patterns on multinode manycore architectures. IEEE Transactions on Parallel and Distributed Systems, 28(6):1635–1648, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/06871392-
abs.html.

Carrera:2012:APM

[CSW^{+}12] David Carrera, Małgorzata Steinder, Ian Whalley, Jordi Torres, and Eduard Ayguade. Autonomic placement of mixed batch and transactional workloads. IEEE Transactions on Parallel and Distributed Systems, 23(2): 219–231, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cassell:2017:NDC

[CSW^{+}17] Benjamin Cassell, Tyler Szepesi, Bernard Wong, Tim Brecht, Jonathan Ma, and Xiaoyi Liu. Nessie: A decoupled, client-driven key-value store using RDMA. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3537–3552, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/07987083-
                                        abs.html.

Chen:2015:MMI

[CSY15] Kang Chen, Haiying Shen, and Li Yan. Multicent: A multifunctional incentive scheme adaptive to diverse performance objectives for DTN routing. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1643–1653, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/06/06814278-
                                        abs.html.

Chen:2016:EFS

[CSY16] Kang Chen, Haiying Shen, and Li Yan. Efficient file search in delay tolerant networks with social content and contact awareness. IEEE Transactions on Parallel and Distributed Systems, 27(7): 1982–1995, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07219439-
abs.html.

Chandler:2012:TPB

[CSZ^{+}12] Harrison Chandler, Haiying Shen, Lianyu Zhao, Jared Stokes, and Jin Li. Toward P2P-based multimedia sharing in user generated contents. IEEE Transactions on Parallel and Distributed Systems, 23(5): 966–975, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Carbunar:2012:POC

[CT12] Bogdan Carbunar and Mahesh V. Tripunitara. Payments for outsourced computations. IEEE Transactions on Parallel and Distributed Systems, 23(2):313– 320, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cambazoglu:2014:IPI

[CTA14] B. Barla Cambazoglu, E. Kartal Tabak, and Cevdet Aykanat. Improving the performance of independent task assignment heuristics MinMin, MaxMin and sufferage. IEEE Transactions on Parallel and Distributed Systems, 25(5):1244–1256, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2019:EHN

[CTG^{+}19] Lin Cui, Fung Po Tso, Song Guo, Weijia Jia, Kaimin Wei, and Wei Zhao. Enabling heterogeneous network function chaining. IEEE Transactions on Parallel and Distributed Systems, 30(4):842– 854, April 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

                                        URL https://ieeexplore.
                                        ieee.org/document/8470961/

.

Carbunaru:2014:MFC

[CTLH14] Cristina Carbunaru, Yong Meng Teo, Ben Leong, and Tracey Ho. Modeling flash crowd performance in peer-to-peer file distribution. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2617–2626, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06587458-
                                        abs.html.

Cui:2017:PJP

[CTP^{+}17] Lin Cui, Fung Po Tso, Dimitrios P. Pezaros, Weijia Jia, and Wei Zhao. PLAN: Joint policy- and networkaware VM management for cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1163–1175, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07556964-
abs.html.

Chang:2011:SPA

[CTX^{+}11] Xiangmao Chang, Rui Tan, Guoliang Xing, Zhaohui Yuan, Chenyang Lu, Yixin Chen, and Yixian Yang. Sensor placement algorithms for fusion-based surveillance networks. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1407–1414, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2012:FAU

[CTX^{+}12] Jinzhu Chen, Rui Tan, Guoliang Xing, Xiaorui Wang, and Xing Fu. Fidelityaware utilization control for cyber-physical surveillance systems. IEEE Transactions on Parallel and Distributed Systems, 23(9):1739–1751, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chiesi:2015:PAJ

[CVM^{+}15] Matteo Chiesi, Luca Vanzolini, Claudio Mucci, Eleonora Franchi Scarselli, and Roberto Guerrieri. Power-aware job scheduling on heterogeneous multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 26(3):868– 877, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06782408-
abs.html.

Chang:2015:PCR

[CW15] Wei Chang and Jie Wu. Progressive or conservative: Rationally allocate cooperative work in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 26(7): 2020–2035, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06846357-
abs.html.

Chen:2019:OPN

[CW19] Yu-Jia Chen and Li-Chun Wang. An overflow problem in network coding for secure cloud storage. IEEE Transactions on Parallel and Distributed Systems, 30(4):789–799, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.

ieee.org/document/8466924/

.

Chang:2011:JOC

[CWC11] Shih Yu Chang, Hsiao-Chun Wu, and John M. Cioffi. Joint optimization of complexity and overhead for the routing in hierarchical networks. IEEE Transactions on Parallel and Distributed Systems, 22(6): 1034–1041, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2013:DSW

[CWC^{+}13] Yong Cui, Hongyi Wang, Xiuzhen Cheng, Dan Li, and Antti Yla-Jaaski. Dynamic scheduling for wireless data center networks. IEEE Transactions on Parallel and Distributed Systems, 24(12):2365–2374, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cui:2015:LED

[CWCS15] Yan Cui, Yingxin Wang, Yu Chen, and Yuanchun Shi. LockSim: An event-driven simulator for modeling spin lock contention. IEEE Transactions on Parallel and Distributed Systems, 26 (1):185–195, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2015/01/06748069-
abs.html.

Crichigno:2011:TOM

[CWJS11] Jorge Crichigno, Min-You Wu, Sudharman K. Jayaweera, and Wei Shu. Throughput optimization in multihop wireless networks with multipacket reception and directional antennas. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1206–1213, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2014:PPM

[CWL^{+}14a] Ning Cao, Cong Wang, Ming Li, Kui Ren, and Wenjing Lou. Privacy-preserving multi-keyword ranked search over encrypted cloud data. IEEE Transactions on Parallel and Distributed Systems, 25(1):222–233, January 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cao:2014:FCH

[CWL14b] Xuanyu Cao, Xinbing Wang, and Songwu Lu. Function computation over heterogeneous wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1756–1766, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2016:EPN

[CWL16] Yu-Jia Chen, Li-Chun Wang, and Chen-Hung Liao. Eavesdropping prevention for network coding encrypted cloud storage systems. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2261–2273, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07289458-
abs.html.

Chen:2019:ATM

[CWLS19] Y. Chen, S. Wang, S. Lu, and K. Sankaralingam. Applying transactional memory for concurrency-bug failure recovery in production runs. IEEE Transactions on Parallel and Distributed Systems, 30(5):990–1006, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chakraborty:2012:SOV

[CWS12] Koushik Chakraborty, Philip M.

Wells, and Gurindar S. Sohi. Supporting overcommitted virtual machines through hardware spin detection. IEEE Transactions on Parallel and Distributed Systems, 23(2):353– 366, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:PSC

[CWZ^{+}15] Dan Chen, Lizhe Wang, Albert Y. Zomaya, Ming-Gang Dou, Jingying Chen, Ze Deng, and Salim Hariri. Parallel simulation of complex evacuation scenarios with adaptive agent models. IEEE Transactions on Parallel and Distributed Systems, 26(3):847–857, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06774442-
abs.html.

Chen:2015:DCA

[CYC^{+}15] Jiming Chen, Qing Yu, Bo Chai, Youxian Sun, Yanfei Fan, and Xuemin Sherman Shen. Dynamic channel assignment for wireless sensor networks: A regret matching based approach. IEEE Transactions on Parallel and Distributed Systems, 26(1):95–106, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06747351-
abs.html.

Cheng:2016:DDM

[CYC^{+}16] Yurong Cheng, Ye Yuan, Lei Chen, Guoren Wang, Christophe Giraud-Carrier, and Yongjiao Sun. DistR: A distributed method for the reachability query over large uncertain graphs. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3172–3185, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07420710-
abs.html.

Chen:2014:LON

[CYL^{+}14] Tao Chen, Zheng Yang, Yunhao Liu, Deke Guo, and Xueshan Luo. Localizationoriented network adjustment in wireless ad hoc and sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(1):146– 155, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2018:RBF

[CYW^{+}18] Rong Chen, Youyang Yao, Peng Wang, Kaiyuan Zhang, Zhaoguo Wang, Haibing Guan, Binyu Zang, and Haibo Chen. Replicationbased fault-tolerance for large-scale graph processing. IEEE Transactions on Parallel and Distributed Systems, 29(7):1621–1635, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/07927721-
abs.html.

Cao:2014:HUB

[CYX^{+}14] Jiao Cao, Wei Yin, Yingyuan Xiao, LihChyun Shu, and Hongya Wang. Hyperbolic utilization bounds for rate monotonic scheduling on homogeneous multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 25(6):1510–1521, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2015:LLD

[CYX15] Qi Chen, Jinyu Yao, and Zhen Xiao. LIBRA: Lightweight data skew mitigation in MapReduce. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2520–2533, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/09/06882249.
                                        pdf.

Champion:2013:SDM

[CYZ^{+}13] Adam C. Champion, Zhimin Yang, Boying Zhang, Jiangpeng Dai, Dong Xuan, and Du Li. E-SmallTalker: A distributed mobile system for social networking in physical proximity. IEEE Transactions on Parallel and Distributed Systems, 24(8):1535–1545, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2019:HAW

[CZD^{+}19] Dazhao Cheng, Xiaobo Zhou, Zhijun Ding, Yu Wang, and Mike Ji. Heterogeneity aware workload management in distributed sustainable datacenters. IEEE Transactions on Parallel and Distributed Systems, 30(2): 375–387, February 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08439018-
abs.html.

Chen:2016:FCU

[CZL^{+}16] Haibo Chen, Heng Zhang, Ran Liu, Binyu Zang, and Haibing Guan. Fast consensus using bounded staleness for scalable read-mostly synchronization. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3485–3500, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07429771-
abs.html.

Cheng:2018:EEA

[CZL^{+}18] Dazhao Cheng, Xiaobo Zhou, Palden Lama, Mike Ji, and Changjun Jiang. Energy efficiency aware task assignment with DVFS in heterogeneous Hadoop clusters. IEEE Transactions on Parallel and Distributed Systems, 29(1):70–82, January 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08017447-
abs.html.

Chen:2017:SWS

[CZQ^{+}17] Huangke Chen, Xiaomin Zhu, Dishan Qiu, Ling Liu, and Zhihui Du. Scheduling for workflows with securitysensitive intermediate data by selective tasks duplication in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2674–2688, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07872483-
abs.html.

Cappos:2018:TMP

[CZRB18] Justin Cappos, Yanyan Zhuang, Albert Rafetseder, and Ivan Beschastnikh. Tsumiki: A meta-platform for building your own testbed. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2863–2881, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08382260-
abs.html.

Chen:2016:EPP

[CZS^{+}16] Chi Chen, Xiaojie Zhu, Peisong Shen, Jiankun Hu, Song Guo, Zahir Tari, and Albert Y. Zomaya. An efficient privacy-preserving ranked keyword search method. IEEE Transactions on Parallel and Distributed Systems, 27(4):951–963, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07091954.
pdf.

Chen:2017:AEC

[CZT^{+}17] Yuanqi Chen, Yi Zhou, Shubbhi Taneja, Xiao Qin, and Jianzhong Huang. aHDFS: An erasure-coded data archival system for Hadoop clusters. IEEE Transactions on Parallel and Distributed Systems, 28(11):3060–3073, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07932189-
abs.html.

Cheng:2018:ASP

[CZWJ18] Dazhao Cheng, Xiaobo Zhou, Yu Wang, and Changjun Jiang. Adaptive scheduling parallel jobs with dynamic batching in spark streaming. IEEE Transactions on Parallel and Distributed Systems, 29(12):

2672–2685, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/12/08382327-
                                        abs.html.

Chen:2014:TTD

[CZWZ14] Yanjiao Chen, Jin Zhang, Kaishun Wu, and Qian Zhang. TAMES: A truthful double auction for multidemand heterogeneous spectrums. IEEE Transactions on Parallel and Distributed Systems, 25(11): 3012–3024, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06674928-
                                        abs.html.

Chen:2019:GCE

[CZX^{+}19a] H. Chen, X. Zhang, Y. Xu, J. Ren, J. Fan, Z. Ma, and W. Zhang. T-Gaming: A cost-efficient cloud gaming system at scale. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2849–2865, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Cheng:2019:DAM

[CZX^{+}19b] Dazhao Cheng, Xiaobo Zhou, Yinggen Xu, Liu Liu, and Changjun Jiang. Deadline-aware MapReduce job scheduling with dynamic resource availability. IEEE Transactions on Parallel and Distributed Systems, 30(4):814–826, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8478331/

.

Chen:2014:WSR

[CZYL14] Xi Chen, Zibin Zheng, Qi Yu, and Michael R. Lyu. Web service recommendation via exploiting location and QoS information. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1913–1924, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chang:2016:PBR

[CZZ^{+}16] Shan Chang, Hongzi Zhu, Wei Zhang, Li Lu, and Yanmin Zhu. PURE: Blind regression modeling for low quality data with participatory sensing. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1199–1211, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07097702-
abs.html.

Demirbas:2016:SBD

[DA16] Murat Demirbas and Anish Arora. Specification-based design of self-stabilization. IEEE Transactions on Parallel and Distributed Systems, 27(1):263–270, January 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07004085-
                                        abs.html.

Dinh:2014:SRD

[DAJ14] Minh Ngoc Dinh, David Abramson, and Chao Jin. Scalable relative debugging. IEEE Transactions on Parallel and Distributed Systems, 25(3):740–749, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dan:2011:CCC

[Dan11]  Gyorgy Dan.  Cache-to-

cache: Could ISPs cooperate to decrease peer-to-peer content distribution costs? IEEE Transactions on Parallel and Distributed Systems, 22(9):1469–1482, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Davidson:2018:ERG

[DB18] R. L. Davidson and C. P. Bridges. Error resilient GPU accelerated image processing for space applications. IEEE Transactions on Parallel and Distributed Systems, 29(9): 1990–2003, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08307249-
abs.html.

DeGrande:2017:TSO

[DBA17] Robson Eduardo De Grande, Azzedine Boukerche, and Raed Alkharboush. Time series-oriented load prediction model and migration policies for distributed simulation systems. IEEE Transactions on Parallel and Distributed Systems, 28(1):215– 229, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07450171-
abs.html.

Dacosta:2011:IAP

[DBAT11] Italo Dacosta, Vijay Balasubramaniyan, Mustaque Ahamad, and Patrick Traynor. Improving authentication performance of distributed SIP proxies. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1804–1812, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dlugosch:2014:ESS

[DBG^{+}14] Paul Dlugosch, Dave Brown, Paul Glendenning, Michael Leventhal, and Harold Noyes.

An efficient and scalable semiconductor architecture for parallel automata processing. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3088–3098, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/12/06719386-
                                        abs.html.

denBurger:2011:CRI

[dBK11] Mathijs den Burger and Thilo Kielmann. Collective receiver-initiated multicast for grid applications. IEEE Transactions on Parallel and Distributed Systems, 22(2): 231–244, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Di:2016:AID

[DC16] Sheng Di and Franck Cappello. Adaptive impactdriven detection of silent data corruption for HPC applications. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2809–2823, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/10/07393580-
                                        abs.html.

Di:2018:OEB

[DC18] Sheng Di and Franck Cappello. Optimization of errorbounded lossy compression for hard-to-compress HPC data. IEEE Transactions on Parallel and Distributed Systems, 29(1):129–143, January 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08031063-
abs.html.

Diener:2016:KBT

[DCA^{+}16] Matthias Diener, Eduardo H. M. Cruz, Marco A. Z. Alves, Philippe O. A. Navaux, Anselm Busse, and Hans-Ulrich Heiss. Kernelbased thread and data mapping for improved memory affinity. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2653–2666, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07345593-
abs.html.

Dou:2019:DTE

[DCA19] Y. Dou, H. C. B. Chan, and M. H. Au. A distributed trust evaluation protocol with privacy protection for intercloud. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1208–1221, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

deCarvalho:2019:CPB

[dCAB19] Joao P. L. de Carvalho, Guido Araujo, and Alexandro Baldassin. The case for phase-based transactional memory. IEEE Transactions on Parallel and Distributed Systems, 30(2):459– 472, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/02/08423653.
                                        pdf.

Dai:2019:MRM

[DCC^{+}19] D. Dai, Y. Chen, P. Carns, J. Jenkins, W. Zhang, and R. Ross. Managing rich metadata in highperformance computing systems using a graph model. IEEE Transactions on Parallel and Distributed Systems, 30(7):1613–1627, July 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Costa:2015:ERH

[dCCF15] Rogerio Luis de Carvalho Costa and Pedro Furtado. Elections and reputation for high dependability and performance in distributed workload execution. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2233–2246, August 2015.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06868299-
abs.html.

Dong:2010:FFL

[DCL^{+}10] Wei Dong, Chun Chen, Xue Liu, Kougen Zheng, Rui Chu, and Jiajun Bu. FIT: a flexible, lightweight, and real-time scheduling system for wireless sensor platforms. IEEE Transactions on Parallel and Distributed Systems, 21(1):126– 138, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dai:2015:QEP

[DCW^{+}15] Haipeng Dai, Guihai Chen, Chonggang Wang, Shaowei Wang, Xiaobing Wu, and Fan Wu. Quality of energy provisioning for wireless power transfer. IEEE Transactions on Parallel and Distributed Systems, 26(2): 527–537, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06762897-
abs.html.

Dash:2011:ICP

[DD11] Alokika Dash and Brian Demsky. Integrating caching and prefetching mechanisms in a distributed transactional memory. IEEE Transactions on Parallel and Distributed Systems, 22(8):1284–1298, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Doray:2017:DPV

[DD17] Francois Doray and Michel Dagenais. Diagnosing performance variations by comparing multi-level execution traces. IEEE Transactions on Parallel and Distributed Systems, 28(2):462– 474, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/02/07469402-
                                        abs.html.

Deveci:2019:GMT

[DDP^{+}19] M. Deveci, K. D. Devine, K. Pedretti, M. A. Taylor, S. Rajamanickam, and Ü. V. Çatalyürek. Geometric mapping of tasks to processors on parallel computers with mesh or torus networks. IEEE Transactions on Parallel and Distributed Systems, 30(9): 2018–2032, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dai:2019:HHA

[DDW^{+}19] Ting Dai, Daniel Dean, Peipei Wang, Xiaohui Gu, and Shan Lu. Hytrace:

A hybrid approach to performance bug diagnosis in production cloud infrastructures. IEEE Transactions on Parallel and Distributed Systems, 30(1):107– 118, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08417446-
abs.html.

DiMartino:2019:CAP

[DED^{+}19] B. Di Martino, A. Esposito, S. D’Angelo, S. A. Maisto, and S. Nacchia. A compiler for agnostic programming and deployment of big data analytics on multiple platforms. IEEE Transactions on Parallel and Distributed Systems, 30(9): 1920–1931, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dehnavi:2013:PSA

[DFGG13] Maryam Mehri Dehnavi, David M. Fernandez, Jean-Luc Gaudiot, and Dennis D. Giannacopoulos. Parallel sparse approximate inverse preconditioning on graphic processing units. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1852–1862, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Divakaran:2015:TFG

[DG15] Dinil Mon Divakaran and Mohan Gurusamy. Towards flexible guarantees in clouds: Adaptive bandwidth allocation and pricing. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1754–1764, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/06/06817564-
                                        abs.html.

Dalvandi:2017:ASP

[DGC17] Aissan Dalvandi, Mohan Gurusamy, and Kee Chaing Chua. Application scheduling, placement, and routing for power efficiency in cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 28(4): 947–960, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/04/07563845-
                                        abs.html.

DeRango:2012:LSE

[DGF12] Floriano De Rango, Francesca Guerriero, and Peppino Fazio. Link-stability and energy aware routing protocol in distributed wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):713– 726, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Delporte-Gallet:2018:ISO

[DGFRR18] Carole Delporte-Gallet, Hugues Fauconnier, Sergio Rajsbaum, and Michel Raynal. Implementing snapshot objects on top of crash-prone asynchronous message-passing systems. IEEE Transactions on Parallel and Distributed Systems, 29(9):2033–2045, September 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08302604-
abs.html.

Di:2019:EPC

[DGG^{+}19] Sheng Di, Hanqi Guo, Rinku Gupta, Eric R. Pershey, Marc Snir, and Franck Cappello. Exploring properties and correlations of fatal events in a large-scale HPC system. IEEE Transactions on Parallel and Distributed Systems, 30(2):361– 374, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08436427-
abs.html.

Denoyelle:2019:MNU

[DGI^{+}19] N. Denoyelle, B. Goglin, A. Ilic, E. Jeannot, and L. Sousa. Modeling nonuniform memory access on large compute nodes with the cache-aware roofline model. IEEE Transactions on Parallel and Distributed Systems, 30(6):1374–1389, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DePalma:2012:SPC

[DHBB12] Noel De Palma, Daniel Hagimont, Fabienne Boyer, and Laurent Broto. Selfprotection in a clustered distributed system. IEEE Transactions on Parallel and Distributed Systems, 23(2): 330–336, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Deng:2015:COS

[DHTZ15] Shuiguang Deng, Longtao Huang, Javid Taheri, and Albert Y. Zomaya. Computation offloading for service workflow in mobile cloud computing. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3317–3329, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/12/06987330-
                                        abs.html.

DiSanzo:2017:ACC

[Di 17] Pierangelo Di Sanzo. Analysis, classification and comparison of scheduling techniques for software transactional memories. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3356–3373, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/08010909-
abs.html.

Dorier:2016:UFG

[DIAR16] Matthieu Dorier, Shadi Ibrahim, Gabriel Antoniu, and Rob Ross. Using formal grammars to predict I/O behaviors in HPC: The Omnisc’IO approach. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2435–2449, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07289462-
abs.html.

Das:2017:DAM

[DK17] Shirshendu Das and Hemangee K. Kapoor. Dynamic associativity management in tiled CMPs by runtime adaptation of fellow sets. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2229–2243, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2017/08/07831453-
                                        abs.html.

Dobre:2019:PWR

[DKL^{+}19] D. Dobre, G. O. Karame, W. Li, M. Majuntke, N. Suri, and M. Vukoli . Proofs of writing for robust storage. IEEE Transactions on Parallel and Distributed Systems, 30(11):2547–2566, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DiTomaso:2015:WAW

[DKM^{+}15] Dominic DiTomaso, Avinash Kodi, David Matolak, Savas Kaya, Soumyasanta Laha, and William Rayess. A-WiNoC: Adaptive wireless network-on-chip architecture for chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3289–3302, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/12/06991560-
                                        abs.html.

Dao:2015:PMG

[DKS^{+}15] Thanh Tuan Dao, Jungwon Kim, Sangmin Seo, Bernhard Egger, and Jaejin Lee. A performance model for GPUs with caches. IEEE Transactions on Parallel and Distributed Systems, 26(7):1800–1813, July 2015.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06844867-
abs.html.

Du:2017:AET

[DL17] Min Du and Feifei Li.

ATOM: Efficient tracking, monitoring, and orchestration of cloud resources. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2172–2189, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07815424-
abs.html.

Dinh:2018:BAS

[DLA^{+}18] Son Dinh, Jing Li, Kunal Agrawal, Chris Gill, and Chenyang Lu. Blocking analysis for spin locks in real-time parallel tasks. IEEE Transactions on Parallel and Distributed Systems, 29(4):789–802, April 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08122044-
abs.html.

Dong:2019:GAF

[DLB^{+}19] Z. Dong, C. Liu, S. Bateni, Z. Kong, L. He, L. Zhang, R. Prakash, and Y. Zhang. A general analysis framework for soft real-time tasks. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1222–1237, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dong:2016:PDA

[DLC^{+}16] Wei Dong, Luyao Luo, Chun Chen, Jiajun Bu, Xue Liu, and Yunhao Liu. Postdeployment anomaly detection and diagnosis in networked embedded systems by program profiling and symptom mining. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3588–3601, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/12/07434644-
                                        abs.html.

Dong:2011:ESM

[DLL^{+}11] Dezun Dong, Xiangke Liao, Yunhao Liu, Changxiang Shen, and Xinbing Wang. Edge self-monitoring for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(3):514–527, March 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Degomme:2017:SMA

[DLM^{+}17] Augustin Degomme, Arnaud Legrand, George S. Markomanolis, Martin Quinson, Mark Stillwell, and Frederic Suter. Simulating MPI applications: The SMPI approach. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2387–2400, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07855780-
abs.html.

deLucas:2019:VRO

[dLMPG19] Enrique de Lucas, Pedro Marcuello, Joan-Manuel Parcerisa, and Antonio Gonzalez. Visibility rendering order: Improving energy efficiency on mobile GPUs through frame coherence. IEEE Transactions on Parallel and Distributed Systems, 30(2):473– 485, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08440753-
abs.html.

Dai:2019:CER

[DLXS19] Yi Dai, Kai Lu, Liquan Xiao, and Jinshu Su. A costefficient router architecture for HPC inter-connection networks: Design and implementation. IEEE Transactions on Parallel and Distributed Systems, 30(4):738– 753, April 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

                                        URL https://ieeexplore.
                                        ieee.org/document/8478359/

.

Dong:2014:LQA

[DLZ^{+}14] Wei Dong, Yunhao Liu, Zhiwei Zhao, Xue Liu, Chun Chen, and Jiajun Bu. Link quality aware code dissemination in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1776–1786, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dang:2016:CWQ

[DLZH16] Depeng Dang, Ying Liu, Xiaoran Zhang, and Shihang Huang. A crowdsourcing worker quality evaluation algorithm on MapReduce for big data applications. IEEE Transactions on Parallel and Distributed Systems, 27(7):1879–1888, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/07/07273904-
                                        abs.html.

Danak:2011:EBD

[DM11] Amir Danak and Shie Mannor. Efficient bidding in dynamic grid markets. IEEE Transactions on Parallel and Distributed Systems, 22(9): 1483–1496, September 2011.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Diaz:2012:SPP

[DMCN12] Javier Diaz, Camelia Munoz-Caro, and Alfonso Nino. A survey of parallel programming models and tools in the multi and many-core era. IEEE Transactions on Parallel and Distributed Systems, 23(8):1369–1386, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:2016:CFC

[DMR16] Bhaskar Das, Sudip Misra, and Utpal Roy. Coalition formation for cooperative service-based message sharing in vehicular ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 27 (1):144–156, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07001099-
abs.html.

Desnoyers:2012:ULI

[DMS^{+}12] Mathieu Desnoyers, Paul E.

McKenney, Alan S. Stern, Michel R. Dagenais, and Jonathan Walpole. Userlevel implementations of read-copy update. IEEE Transactions on Parallel and Distributed Systems, 23(2): 375–382, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dubois:2012:BIU

[DMT12] Swan Dubois, Toshimitsu Masuzawa, and Sebastien Tixeuil. Bounding the impact of unbounded attacks in stabilization. IEEE Transactions on Parallel and Distributed Systems, 23(3):460– 466, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Das:2019:QAM

[DN19] Uddipan Das and Vinod Namboodiri. A quality-aware multi-level data aggregation approach to manage smart grid AMI traffic. IEEE Transactions on Parallel and Distributed Systems, 30(2): 245–256, February 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08439012-
abs.html.

Dean:2016:POP

[DNW^{+}16] Daniel J. Dean, Hiep Nguyen, Peipei Wang, Xiaohui Gu, Anca Sailer, and Andrzej Kochut. PerfCompass: Online performance anomaly fault localization and inference in Infrastructure-as-a-Service clouds. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1742–1755, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07127024-
abs.html.

Djidjev:2013:SAG

[DO13] Hristo N. Djidjev and Melih Onus. Scalable and accurate graph clustering and community structure detection. IEEE Transactions on Parallel and Distributed Systems, 24(5):1022–1029, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dong:2016:JOL

[DOLG16] Mianxiong Dong, Kaoru Ota, Anfeng Liu, and Minyi Guo. Joint optimization of lifetime and transport delay under reliability constraint wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27 (1):225–236, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07018943-
abs.html.

Sandes:2013:RSW

[dOSdM13] Edans Flavius de O. Sandes and Alba Cristina M. A. de Melo. Retrieving Smith– Waterman alignments with optimizations for megabase biological sequences using GPU. IEEE Transactions on Parallel and Distributed Systems, 24(5):1009–1021, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sandes:2016:CIS

[dOSMM^{+}16] Edans Flavius de Oliveira Sandes, Guillermo Miranda, Xavier Martorell, Eduard Ayguade, George Teodoro, and Alba Cristina Magalhaes Melo. CUDAlign 4.0: Incremental speculative traceback for exact chromosome-wide alignment in GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2838–2850, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/10/07374729-
                                        abs.html.

Dutot:2011:AAM

[DPRT11] Pierre-Francois Dutot, Fanny Pascual, Krzysztof Rzadca, and Denis Trystram. Approximation algorithms for the multiorganization scheduling problem. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1888–1895, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Diegues:2016:SBS

[DR16] Nuno Diegues and Paolo Romano. STI-BT: A scalable transactional index. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2408–2421, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07287754-
abs.html.

Duttagupta:2011:TDB

[DRK11] Subhasri Duttagupta, Krithi Ramamritham, and Purushottam Kulkarni. Tracking dynamic boundaries using sensor network. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1766–1774, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dehne:2018:VSD

[DRRCB18] Frank Dehne, David Edward Robillard, Andrew Rau-Chaplin, and Neil Burke. VOLAP: A scalable distributed real-time OLAP system for high-velocity data. IEEE Transactions on Parallel and Distributed Systems, 29(1):226–239, January 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08014458-
abs.html.

Diallo:2015:DDM

[DRSL15] Ousmane Diallo, Joel Jose P. C. Rodrigues, Mbaye Sene, and Jaime Lloret. Distributed database management techniques for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(2): 604–620, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06582416-
abs.html.

Di:2017:TOO

[DRVC17] Sheng Di, Yves Robert, Frederic Vivien, and Franck Cappello. Toward an optimal online checkpoint solution under a two-level HPC checkpoint model. IEEE Transactions on Parallel and Distributed Systems, 28(1):244– 259, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07442859-
abs.html.

Dominguez-Sal:2012:UES

[DSASSLP12] David Dominguez-Sal, Josep Aguilar-Saborit, Mihai Surdeanu, and Josep Lluis Larriba-Pey. Using evolutive summary counters for efficient cooperative caching in search engines. IEEE Transactions on Parallel and Distributed Systems, 23(4): 776–784, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ding:2016:GAS

[DSJ16] Xiaoning Ding, Jianchen Shan, and Song Jiang. A general approach to scalable buffer pool management. IEEE Transactions on Parallel and Distributed Systems, 27(8):2182–2195, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07286847-
abs.html.

daSilva:2011:CDM

[dSLMM11] Ana Paula Couto da Silva, Emilio Leonardi, Marco Mellia, and Michela Meo. Chunk distribution in mesh-based large-scale P2P streaming systems: a fluid approach. IEEE Transactions on Parallel and Distributed Systems, 22(3):451–463, March 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dargie:2014:PCE

[DSM14] Waltenegus Dargie, Alexander Schill, and Christoph Mobius. Power consumption estimation models for processors, virtual machines, and servers. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1600–1614, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dass:2019:FCE

[DSM19] J. Dass, V. Sarin, and R. N. Mahapatra. Fast and communication-efficient algorithm for distributed support vector machine training. IEEE Transactions on Parallel and Distributed Systems, 30(5):1065–1076, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Dewri:2014:ESS

[DT14] Rinku Dewri and Ramakrisha Thurimella. Exploiting service similarity for privacy in location-based search queries. IEEE Transactions on Parallel and Distributed Systems, 25(2):374– 383, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Di:2019:ELC

[DTLC19] Sheng Di, Dingwen Tao, Xin Liang, and Franck Cappello. Efficient lossy compression for scientific data based on pointwise relative error bound. IEEE Transactions on Parallel and Distributed Systems, 30(2):331– 345, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08421751-
abs.html.

Duato:1993:NTD

[Dua93]  Jose Duato.  New the-

ory of deadlock-free adaptive routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 4(12):1320–1331, December 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [Dua95] and comment [VS11a].

Duato:1995:NSC

[Dua95]  Jose Duato.  A neces-

sary and sufficient condition for deadlock-free adaptive routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 6(10): 1055–1067, October 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/tpds/
td1995/l1055abs.htm. See

[Dua93] and comment [VS11a].

Drozdowski:2010:IMD

[DW10] Maciej Drozdowski and Lukasz Wielebski. Isoefficiency maps for divisible computations. IEEE Transactions on Parallel and Distributed Systems, 21(6): 872–880, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Di:2013:DOM

[DW13a] Sheng Di and Cho-Li Wang. Dynamic optimization of multiattribute resource allocation in self-organizing clouds. IEEE Transactions on Parallel and Distributed Systems, 24(3):464– 478, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Di:2013:ETR

[DW13b] Sheng Di and Cho-Li Wang. Error-tolerant resource allocation and payment minimization for cloud system. IEEE Transactions on Parallel and Distributed Systems, 24(6):1097–1106, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duarte:2012:DDD

[DWF12] Elias P. Duarte, Andrea Weber, and Keiko V. Ono Fonseca. Distributed diagnosis of dynamic events in partitionable arbitrary topology networks. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1415–1426, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Deng:2018:GMO

[DWH^{+}18] Ze Deng, Lizhe Wang, Wei Han, Rajiv Ranjan, and Albert Zomaya. G-ML-Octree:

An update-efficient index structure for simulating 3D moving objects across GPUs. IEEE Transactions on Parallel and Distributed Systems, 29(5):1075–1088, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08241384-
abs.html.

Deng:2015:SSM

[DWLY15] Xianjun Deng, Bang Wang, Wenyu Liu, and Laurence T. Yang. Sensor scheduling for multi-modal confident information coverage in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(3):902– 913, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06782385-
abs.html.

Deng:2016:CPD

[DWT^{+}16] Shuiguang Deng, Hongyue Wu, Javid Taheri, Albert Y. Zomaya, and Zhaohui Wu. Cost performance driven service mashup: A developer perspective. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2234–2247, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/

                                        trans/td/2016/08/07279186-
                                        abs.html.

Ding:2011:EAT

[DWW^{+}11] Ling Ding, Weili Wu, James Willson, Hongjie Du, Wonjun Lee, and Ding-Zhu Du. Efficient algorithms for topology control problem with routing cost constraints in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 22(10):1601–1609, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Deng:2015:PPD

[DWW^{+}15] Ze Deng, Xiaomin Wu, Lizhe Wang, Xiaodao Chen, Rajiv Ranjan, Albert Zomaya, and Dan Chen. Parallel processing of dynamic continuous queries over streaming data flows. IEEE Transactions on Parallel and Distributed Systems, 26(3):834– 846, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06767134-
                                        abs.html.

Dai:2014:EVB

[DWX14] Ying Dai, Jie Wu, and Chunsheng Xin. Efficient virtual backbone construction without a common control channel in cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3156–3166, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714570-
abs.html.

Du:2013:CBV

[DWY^{+}13] Hongwei Du, Weili Wu, Qiang Ye, Deying Li, Wonjun Lee, and Xuepeng Xu. CDS-based virtual backbone construction with guaranteed routing cost in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(4): 652–661, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Duan:2016:PPA

[DY16] Jun Duan and Yuanyuan Yang. Placement and performance analysis of virtual multicast networks in fat-tree data center networks. IEEE Transactions on Parallel and Distributed Systems, 27(10): 3013–3028, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07370928-
abs.html.

Duan:2017:LBM

[DY17] Jun Duan and Yuanyuan Yang. A load balancing and multi-tenancy oriented data center virtualization framework. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2131–2144, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07831441-
abs.html.

Duan:2018:MCE

[DY18] Jun Duan and Yuanyuan Yang. MCL: A cost-efficient nonblocking multicast interconnection network. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2046–2058, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08320837-
abs.html.

Dou:2015:HIT

[DZLC15] Wanchun Dou, Xuyun Zhang, Jianxun Liu, and Jinjun Chen. HireSome-II: Towards privacy-aware crosscloud service composition for big data applications. IEEE Transactions on Parallel and Distributed Systems, 26(2): 455–466, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06613488-
abs.html.

Edmonds:2019:HAS

[EADT19] Mark Edmonds, Tanvir Atahary, Scott Douglass, and Tarek Taha. Hardware accelerated semantic declarative memory systems through CUDA and MapReduce. IEEE Transactions on Parallel and Distributed Systems, 30(3):601–614, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08444694-
abs.html.

Enfedaque:2015:IDG

[EALM15] Pablo Enfedaque, Francesc Auli-Llinas, and Juan C. Moure. Implementation of the DWT in a GPU through a register-based strategy. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3394–3406, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06991550-
abs.html.

Enfedaque:2017:GIB

[EALM17] Pablo Enfedaque, Francesc Auli-Llinas, and Juan Carlos Moure. GPU implementation of bitplane coding with parallel coefficient processing for high performance image compression. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2272–2284, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07833172-
                                        abs.html.

El-Araby:2011:FEH

[EAMEG11] Esam El-Araby, Saumil G. Merchant, and Tarek El-Ghazawi. A framework for evaluating high-level design methodologies for highperformance reconfigurable computers. IEEE Transactions on Parallel and Distributed Systems, 22(1):33– 45, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Escobar:2016:SAF

[ECV16] Fernando A. Escobar, Xin Chang, and Carlos Valderrama. Suitability analysis of FPGAs for heterogeneous platforms in HPC. IEEE Transactions on Parallel and Distributed Systems, 27(2): 600–612, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07051218-
                                        abs.html.

E:2018:CEE

[ECW^{+}18] Jinlong E, Yong Cui, Peng Wang, Zhenhua Li, and Chaokun Zhang. CoCloud:

Enabling efficient cross-cloud file collaboration based on inefficient Web APIs. IEEE Transactions on Parallel and Distributed Systems, 29(1): 56–69, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08030092-
abs.html.

Ekanayake:2011:CTB

[EGQ11] Jaliya Ekanayake, Thilina Gunarathne, and Judy Qiu. Cloud technologies for bioinformatics applications. IEEE Transactions on Parallel and Distributed Systems, 22(6): 998–1011, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

El-Hassan:2011:SEC

[EHI11] Fadi El-Hassan and Dan Ionescu. SCBXP: An efficient CAM-based XML parsing technique in hardware environments. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1879–1887, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Elrabaa:2017:VFT

[EHM^{+}17] Muhammad E. S. Elrabaa, Ayman Hroub, Muhamed F. Mudawar, Amran Al-Aghbari, Mohammed Al-Asli, and Ahmad Khayyat. A very fast trace-driven simulation platform for chip-multiprocessors architectural explorations. IEEE Transactions on Parallel and Distributed Systems, 28(11):3033–3045, November 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07944586-
abs.html.

Elwhishi:2013:NMS

[EHNS13a] Ahmed Elwhishi, Pin-Han Ho, K. Naik, and Basem Shihada. A novel message scheduling framework for delay tolerant networks routing. IEEE Transactions on Parallel and Distributed Systems, 24(5):871– 880, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Elwhishi:2013:SAC

[EHNS13b] Ahmed Elwhishi, Pin-Han Ho, K. Naik, and Basem Shihada. Self-adaptive contention aware routing protocol for intermittently connected mobile networks. IEEE Transactions on Parallel and Distributed Systems, 24(7):1422–1435, July 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Eckart:2010:DPB

[EHWX10] Ben Eckart, Xubin He, Qishi Wu, and Changsheng Xie. A dynamic performance-based flow control method for highspeed data transfer. IEEE Transactions on Parallel and Distributed Systems, 21(1): 114–125, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Etsion:2014:HDN

[EJGYAM14] Yoav Etsion, Daniel Jimenez-Gonzalez, Fahimeh Yazdanpanah, and Carlos Alvarez-Martinez. Hybrid dataflow/ von-Neumann architectures. IEEE Transactions on Parallel and Distributed Systems, 25(6):1489–1509, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ediger:2013:GMA

[EJRB13] David Ediger, Karl Jiang, E. Jason Riedy, and David A. Bader. GraphCT: Multithreaded algorithms for massive graph analysis. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2220–2229, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

El-Khatib:2010:IFR

[EK10] Khalil El-Khatib. Impact of feature reduction on the efficiency of wireless intrusion detection systems. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1143–1149, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Erickson:2017:OSP

[EKNS17] Alejandro Erickson, Abbas E. Kiasari, Javier Navaridas, and Iain A. Stewart. An optimal singlepath routing algorithm in the datacenter network DPillar. IEEE Transactions on Parallel and Distributed Systems, 28(3):689–703, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/03/07511760-
                                        abs.html.

Evangelinos:2011:MTC

[ELX^{+}11] Constantinos Evangelinos, Pierre F. J. Lermusiaux, Jinshan Xu, Patrick J. Haley, and Chris N. Hill. Many task computing for real-time uncertainty prediction and data assimilation in the ocean. IEEE Transactions on Parallel and Distributed Systems, 22(6):1012–1024, June 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

El-Moukaddem:2015:MNT

[EMTX15] Fatme El-Moukaddem, Eric Torng, and Guoliang Xing. Maximizing network topology lifetime using mobile node rotation. IEEE Transactions on Parallel and Distributed Systems, 26(7): 1958–1970, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06828738-
abs.html.

Esposito:2016:VAV

[EMW16] Flavio Esposito, Ibrahim Matta, and Yuefeng Wang. VINEA: An architecture for virtual network embedding policy programmability. IEEE Transactions on Parallel and Distributed Systems, 27(11):3381–3396, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07401050-
abs.html.

Elhadef:2012:CBS

[EN12] Mourad Elhadef and Amiya Nayak. Comparison-based system-level fault diagnosis: a neural network approach. IEEE Transactions on Parallel and Distributed Systems, 23(6):1047–1059, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Esteve:2017:TBT

[ERG^{+}17] Albert Esteve, Alberto Ros, Maria E. Gomez, Antonio Robles, and Jose Duato. TLB-based temporalityaware classification in CMPs with multilevel TLBs. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2401–2413, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07833158-
                                        abs.html.

Esteve:2018:TTB

[ERRG18] Albert Esteve, Alberto Ros, Antonio Robles, and Maria E. Gomez. TokenTLB+CUP: A tokenbased page classification with cooperative usage prediction. IEEE Transactions on Parallel and Distributed Systems, 29(5):1188–1201, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/05/08194915-
                                        abs.html.

Erdelj:2013:CPI

[ERSR13] Milan Erdelj, Tahiry Razafindralambo, and David Simplot-Ryl. Covering points of interest with mobile sensors. IEEE Transactions on Parallel and Distributed Systems, 24(1):32–43, January 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Escudero-Sahuquillo:2015:ECE

[ESGG^{+}15] Jesus Escudero-Sahuquillo, Ernst Gunnar Gran, Pedro J. Garcia, Jose Flich, Tor Skeie, Olav Lysne, Francisco J. Quiles, and Jose Duato. Efficient and costeffective hybrid congestion control for HPC interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 26 (1):107–119, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06747321-
abs.html.

Escudero-Sahuquillo:2013:EFC

[ESGQ^{+}13] Jesus Escudero-Sahuquillo, Pedro J. Garcia, Francisco J. Quiles, Jose Flich, and Jose Duato. An effective and feasible congestion management technique for high-performance MINs with tag-based distributed routing. IEEE Transactions on Parallel and Distributed Systems, 24(10): 1918–1929, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Eberhard:2010:SBO

[ET10] John Eberhard and Anand Tripathi. Semantics-based object caching in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 21(12): 1750–1764, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fraguela:2019:EDP

[FA19] B. B. Fraguela and D. Andrade. Easy dataflow programming in clusters with UPC++ DepSpawn. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1267–1282, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Flahive:2010:TGE

[FB10] Mary Flahive and Bella Bose. The topology of Gaussian and Eisenstein– Jacobi interconnection networks. IEEE Transactions on Parallel and Distributed Systems, 21(8):1132–1142, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fraternali:2018:QIV

[FBCB18] Francesco Fraternali, Andrea Bartolini, Carlo Cavazzoni, and Luca Benini. Quantifying the impact of variability and heterogeneity on the energy efficiency for a next-generation ultragreen supercomputer. IEEE Transactions on Parallel and Distributed Systems, 29(7): 1575–1588, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08081827-
abs.html.

Fan:2010:ERS

[FC10] Yao-Chung Fan and Arbee L. P. Chen. Efficient and robust schemes for sensor data aggregation based on linear counting. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1675–1691, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ferng:2011:GOH

[FC11] Huei-Wen Ferng and Iwan Christanto. A globally overlaid hierarchical P2P–SIP architecture with route optimization. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1826–1833, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fusella:2018:LBT

[FC18] Edoardo Fusella and Alessandro Cilardo. Lattice-based turn model for adaptive routing. IEEE Transactions on Parallel and Distributed Systems, 29(5):1117–1130, May 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08240635-
abs.html.

Fan:2013:GBB

[FCD^{+}13] Jialu Fan, Jiming Chen, Yuan Du, Wei Gao, Jie Wu, and Youxian Sun. Geocommunity-based broadcasting for data dissemination in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 24(4):734– 743, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ferretti:2014:DCI

[FCM14] Luca Ferretti, Michele Colajanni, and Mirco Marchetti. Distributed, concurrent, and independent access to encrypted cloud databases. IEEE Transactions on Parallel and Distributed Systems, 25(2):437–446, February 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fajardo-Delgado:2013:BAP

[FDFZB13] Daniel Fajardo-Delgado, Jose Alberto Fernandez-Zepeda, and Anu G. Bourgeois. The bodyguard allocation problem. IEEE Transactions on Parallel and Distributed Systems, 24(7):1465–1478, July 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Feng:2014:DIH

[Fen14]  Dan Feng.  Design and

implementation of holistic scheduling and efficient storage for FlexRay. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2529–2539, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06579597-
abs.html.

Feliu:2017:IIP

[FES^{+}17] Josue Feliu, Stijn Eyerman, Julio Sahuquillo, Salvador Petit, and Lieven Eeckhout. Improving IBM POWER8 performance through symbiotic job scheduling. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2838–2851, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07893747-
abs.html.

Fusella:2017:PSH

[FFC17] Edoardo Fusella, Jose Flich, and Alessandro Cilardo. Path setup for hybrid NoC architectures exploiting flooding and standby. IEEE Transactions on Parallel and Distributed Systems, 28(5):1403–1416, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07723923-
abs.html.

Felber:2010:TBS

[FFMR10] Pascal Felber, Christof Fetzer, Patrick Marlier, and Torvald Riegel. Time-based software transactional memory. IEEE Transactions on Parallel and Distributed Systems, 21(12):1793–1807, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fresno:2014:BEP

[FGEL14] Javier Fresno, Arturo Gonzalez-Escribano, and Diego R. Llanos. Blending extensibility and performance in dense and sparse parallel data management. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2509–2519, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06616547-
                                        abs.html.

Fan:2015:SCO

[FGJ^{+}15] Xuepeng Fan, Zhenyu Guo, Hai Jin, Xiaofei Liao, Jiaxing Zhang, Hucheng Zhou, Sean McDirmid, Wei Lin, Jingren Zhou, and Lidong Zhou. Spotting code optimizations in data-parallel pipelines through PeriSCOPE. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1718–1731, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06819843-
abs.html.

Folling:2010:RLD

[FGLP10] Alexander Folling, Christian Grimme, Joachim Lepping, and Alexander Papaspyrou. Robust load delegation in service grid environments. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1304–1316, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2015:POM

[FHH^{+}15] Songling Fu, Ligang He, Chenlin Huang, Xiangke Liao, and Kenli Li. Performance optimization for managing massive numbers of small files in distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3433–3448, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06977977-
abs.html.

Fang:2011:MAM

[FHLG11] Wenbin Fang, Bingsheng He, Qiong Luo, and Naga K. Govindaraju. Mars: Accelerating MapReduce with graphics processors. IEEE Transactions on Parallel and Distributed Systems, 22(4): 608–620, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Feng:2011:EAS

[FHW11] Yi-Hsuan Feng, Nen-Fu Huang, and Yen-Min Wu. Efficient and adaptive stateful replication for stream processing engines in highavailability cluster. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1788–1796, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fiorin:2018:NMA

[FJV^{+}18] Leandro Fiorin, Rik Jongerius, Erik Vermij, Jan van Lunteren, and Christoph Hagleitner. Near-memory acceleration for radio astronomy. IEEE Transactions on Parallel and Distributed Systems, 29(1):115– 128, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/01/08027117-
                                        abs.html.

Fujiwara:2015:SRM

[FKMC15] Ikki Fujiwara, Michihiro Koibuchi, Hiroki Matsutani, and Henri Casanova. Swap-and-Randomize: A method for building lowlatency HPC interconnects.

IEEE Transactions on Parallel and Distributed Systems, 26(7):2051–2060, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06860278-
abs.html.

Fan:2013:CEM

[FLH13] Chun-I Fan, Yi-Hui Lin, and Ruei-Hau Hsu. Complete EAP method: User efficient and forward secure authentication protocol for IEEE 802.11 wireless LANs. IEEE Transactions on Parallel and Distributed Systems, 24(4): 672–680, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fan:2017:GDT

[FLLS17] Xinxin Fan, Ling Liu, Mingchu Li, and Zhiyuan Su. GroupTrust: Dependable trust management. IEEE Transactions on Parallel and Distributed Systems, 28(4):1076–1090, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07572209-
abs.html.

Faizian:2018:RRG

[FMY^{+}18] Peyman Faizian, Md Atiqul Mollah, Xin Yuan, Zaid Alzaid, Scott Pakin, and Michael Lang. Random regular graph and generalized de Bruijn graph with k-shortest path routing. IEEE Transactions on Parallel and Distributed Systems, 29(1):144– 155, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08013054-
abs.html.

Fard:2013:TDW

[FPF13] Hamid Mohammadi Fard, Radu Prodan, and Thomas Fahringer. A truthful dynamic workflow scheduling mechanism for commercial multicloud environments. IEEE Transactions on Parallel and Distributed Systems, 24(6):1203–1212, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fernandez-Pascual:2010:DTF

[FPGAD10] Ricardo Fernandez-Pascual, Jose M. Garcia, Manuel E. Acacio, and Jose Duato. Dealing with transient faults in the interconnection network of CMPs at the cache coherence level. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1117–1131, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2016:SME

[FPRG16] Bo Fang, Karthik Pattabiraman, Matei Ripeanu, and Sudhanva Gurumurthi. A systematic methodology for evaluating the error resilience of GPGPU applications. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3397–3411, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07426849-
abs.html.

Fu:2012:TDA

[FQWL12] Luoyi Fu, Yi Qin, Xinbing Wang, and Xue Liu. Throughput and delay analysis for convergecast with MIMO in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):768–775, April 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Frey:2013:LUB

[Fre13] Hannes Frey. Lower and upper bounds for multicasting under distance dependent forwarding cost functions. IEEE Transactions on Parallel and Distributed Systems, 24(5):963–976, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2016:EPS

[FRS^{+}16] Zhangjie Fu, Kui Ren, Jiangang Shu, Xingming Sun, and Fengxiao Huang. Enabling personalized search over encrypted outsourced data with efficiency improvement. IEEE Transactions on Parallel and Distributed Systems, 27(9):2546–2559, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07349214-
abs.html.

Flich:2012:SET

[FSM^{+}12] Jose Flich, Tor Skeie, Andres Mejia, Olav Lysne, Pedro Lopez, Antonio Robles, Jose Duato, Michihiro Koibuchi, Tomas Rokicki, and Jose Carlos Sancho. A survey and evaluation of topology-agnostic deterministic routing algorithms. IEEE Transactions on Parallel and Distributed Systems, 23(3):405–425, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Falcao:2011:MLD

[FSS11] Gabriel Falcao, Leonel Sousa, and Vitor Silva. Massively LDPC decoding on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 22(2):309– 322, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fu:2016:RSD

[FSSZ16] Yingxun Fu, Jiwu Shu, Zhirong Shen, and Guangyan Zhang. Reconsidering single disk failure recovery for erasure coded storage systems: Optimizing load balancing in stack-level. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1457–1469, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07120164-
abs.html.

Feldman:2016:EWF

[FVLD16] Steven Feldman, Carlos Valera-Leon, and Damian Dechev. An efficient waitfree vector. IEEE Transactions on Parallel and Distributed Systems, 27(3):654– 667, March 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/03/07073592-
abs.html.

Fu:2013:MSL

[FW13] Luoyi Fu and Xinbing Wang.

Multicast scaling law in multichannel multiradio wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(12):

2418–2428, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Fang:2018:EIC

[FWH^{+}18] Juntao Fang, Shenggang Wan, Ping Huang, Changsheng Xie, and Xubin He. Early identification of critical blocks: Making replicated distributed storage systems reliable against node failures. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2446–2459, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/11/08355707-
                                        abs.html.

Fidler:2018:NAD

[FWJ18] Markus Fidler, Brenton Walker, and Yuming Jiang. Non-asymptotic delay bounds for multi-server systems with synchronization constraints. IEEE Transactions on Parallel and Distributed Systems, 29(7):1545–1559, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/07/08242677.
                                        pdf.

Feng:2016:EWC

[FWZ^{+}16] Yuhong Feng, Junpeng Wang, Zhiqiang Zhang, Haoming Zhong, Zhong Ming, Xuan Yang, and Rui Mao. The edge weight computation with MapReduce for extracting weighted graphs. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3659–3672, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07422158-
abs.html.

Feng:2017:ADM

[FXL17] Chen Feng, Hong Xu, and Baochun Li. An alternating direction method approach to cloud traffic management. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2145–2158, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07833029-
abs.html.

Fang:2015:CPC

[FYH^{+}15] Weiwei Fang, Yuan Yao, Longbo Huang, Abhishek B. Sharma, Leana Golubchik, and Michael J. Neely. A comment on “Power Cost Reduction in Distributed Data Centers: A Two Time Scale Approach for Delay Tolerant Workloads”. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1495–1496, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/05/06810886-

abs.html. See [YHS^{+}14].

Ghaderi:2018:AAA

[GAB18] Zana Ghaderi, Ayed Alqahtani, and Nader Bagherzadeh. AROMa: Aging-aware deadlockfree adaptive routing algorithm and online monitoring in 3D NoCs. IEEE Transactions on Parallel and Distributed Systems, 29(4): 772–788, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/04/08166793-
                                        abs.html.

Gharaibeh:2011:TFR

[GAKR11] Abdullah Gharaibeh, Samer Al-Kiswany, and Matei Ripeanu. ThriftStore: Finessing reliability trade-offs in replicated storage systems. IEEE Transactions on Parallel and Distributed Systems, 22(6):910–923, June 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2013:ECE

[GBD^{+}13] Yi Gao, Jiajun Bu, Wei Dong, Chun Chen, Lei Rao, and Xue Liu. Exploiting concurrency for efficient dissemination in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(4):691– 700, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gramoli:2016:RND

[GBFS16] Vincent Gramoli, Len Bass, Alan Fekete, and Daniel W. Sun. Rollup: Non-disruptive rolling upgrade with fast consensus-based dynamic reconfigurations. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2711–2724, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07327226-
abs.html.

Gowanlock:2017:OPC

[GBP17] Michael Gowanlock, David M.

Blair, and Victor Pankratius. Optimizing parallel clustering throughput in shared memory. IEEE Transactions on Parallel and Distributed Systems, 28(9):2595–2607, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07865993-
abs.html.

Gowanlock:2016:DTS

[GC16] Michael Gowanlock and Henri Casanova. Distance threshold similarity searches: Efficient trajectory indexing on the GPU. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2533–2545, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07330012-
abs.html.

Gotz:2018:PCC

[GCG^{+}18] Markus Gotz, Gabriele Cavallaro, Thierry Geraud, Matthias Book, and Morris Riedel. Parallel computation of component trees on distributed memory machines. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2582–2598, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08360392.
pdf.

Guo:2014:MML

[GCL14] Song Guo, Zixue Cheng, and Peng Li. Max-min lifetime optimization for cooperative communications in multichannel wireless networks. IEEE Transactions on Parallel and Distributed Systems, 25(6):1533–1542, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2014:AAS

[GCN^{+}14] Yu Gu, Long Cheng, Jianwei Niu, Tian He, and David Hung-Chang Du. Achieving asymmetric sensing coverage for duty cycled wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3076–3087, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06747354-
abs.html.

Guo:2015:DTC

[GCZ15] Peng Guo, Jiannong Cao, and Kui Zhang. Distributed topological convex Hull estimation of event region in wireless sensor networks without location information. IEEE Transactions on Parallel and Distributed Systems, 26(1):85– 94, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748077-
abs.html.

Giraldeau:2016:WAD

[GD16] Francis Giraldeau and Michel Dagenais. Wait analysis of distributed systems using kernel tracing. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2450–2461, August 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07294678.
                                        pdf.

Guan:2013:PEN

[GDM^{+}13] HaiBing Guan, YaoZu Dong, RuHui Ma, Dongxiao Xu, Yang Zhang, and Jian Li. Performance enhancement for network I/O virtualization with efficient interrupt coalescing and virtual receive-side scaling. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1118–1128, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gonzalez-Dominguez:2016:PPE

[GDRTS16] Jorge Gonzalez-Dominguez, Sabela Ramos, Juan Tourino, and Bertil Schmidt. Parallel pairwise epistasis detection on heterogeneous computing architectures. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2329–2340, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07165657-
                                        abs.html.

Gilmore:2012:SSP

[GE12] John S. Gilmore and Herman A. Engelbrecht. A survey of state persistency in peer-to-peer massively multiplayer online games. IEEE Transactions on Parallel and Distributed Systems, 23(5): 818–834, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gonzalez-Escribano:2014:ESM

[GETFL14] Arturo Gonzalez-Escribano, Yuri Torres, Javier Fresno, and Diego R. Llanos. An extensible system for multilevel automatic data partition and mapping. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1145–1154, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2013:ECS

[GF13] Yuanxiong Guo and Yuguang Fang. Electricity cost saving strategy in data centers by using energy storage. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1149–1160, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gueunet:2019:TBA

[GFJT19] C. Gueunet, P. Fortin, J. Jomier, and J. Tierny. Task-based augmented contour trees with Fibonacci heaps. IEEE Transactions on Parallel and Distributed Systems, 30(8):1889–1905, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2015:DCM

[GFLL15] Hong Gao, Xiaolin Fang, Jianzhong Li, and Yingshu Li. Data collection in multiapplication sharing wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(2): 403–412, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06671566-
                                        abs.html.

Gandino:2013:DAH

[GFMR13] Filippo Gandino, Renato Ferrero, Bartolomeo Montrucchio, and Maurizio Rebaudengo. DCNS: An adaptable high throughput RFID reader-to-reader anticollision protocol. IEEE Transactions on Parallel and Distributed Systems, 24(5):893– 905, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ge:2010:PEP

[GFS^{+}10] Rong Ge, Xizhou Feng, Shuaiwen Song, Hung-Ching Chang, Dong Li, and Kirk W. Cameron. Power-Pack: Energy profiling and analysis of high-performance systems and applications. IEEE Transactions on Parallel and Distributed Systems, 21(5):658–671, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gamatie:2010:SSM

[GG10] Abdoulaye Gamatie and Thierry Gautier. The signal synchronous multiclock approach to the design of distributed embedded systems. IEEE Transactions on Parallel and Distributed Systems, 21(5):641–657, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2011:TES

[GG11] Yunhong Gu and Robert Grossman. Toward efficient and simplified distributed data intensive computing. IEEE Transactions on Parallel and Distributed Systems, 22(6):974–984, June 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gaeta:2013:IMN

[GG13] Rossano Gaeta and Marco Grangetto. Identification of malicious nodes in peerto-peer streaming: A belief propagation-based technique. IEEE Transactions on Parallel and Distributed Systems, 24(10):1994–2003, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2014:ENA

[GGF^{+}14] Yuanxiong Guo, Yanmin Gong, Yuguang Fang, Pramod P. Khargonekar, and Xiaojun Geng. Energy and network aware workload management for sustainable data centers with thermal storage. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2030–2042, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Groves:2018:UNI

[GGGA18] Taylor Liles Groves, Ryan E. Grant, Aaron Gonzales, and Dorian Arnold. Unraveling network-induced memory contention: Deeper insights with machine learning. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1907–1922, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/08/08116691-
                                        abs.html.

Garg:2010:EAG

[GGS10] Rahul Garg, Vijay K. Garg, and Yogish Sabharwal. Efficient algorithms for global snapshots in large distributed systems. IEEE Transactions on Parallel and Distributed Systems, 21(5): 620–630, May 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2019:ERB

[GGY^{+}19] Y. Gao, X. Gao, X. Yang, J. Liu, and G. Chen. An efficient ring-based metadata management policy for largescale distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 30(9):1962–1974, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2013:LSE

[GHL^{+}13] Deke Guo, Yuan He, Yunhao Liu, Panlong Yang, Xiang-Yang Li, and Xin Wang. Link scheduling for exploiting spatial reuse in multihop MIMO networks. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1355–1365, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2014:FGB

[GHL14] Deke Guo, Yuan He, and Yunhao Liu. On the feasibility of gradient-based data-centric routing using Bloom filters. IEEE Transactions on Parallel and Distributed Systems, 25(1):180– 190, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2016:BHF

[GHW^{+}16] Chunhui Guo, Xiayu Hua, Hao Wu, Douglas Lautner, and Shangping Ren. Best-harmonically-fit periodic task assignment algorithm on multiple periodic resources. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1303–1315, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07112525-
abs.html.

Gong:2015:NPA

[GHZ15] Yifan Gong, Bingsheng He, and Jianlong Zhong. Network performance aware MPI collective communication operations in the cloud. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3079–3089, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06490322-
abs.html.

Gu:2016:SAM

[GHZZ16] Zonghua Gu, Gang Han, Haibo Zeng, and Qingling Zhao. Security-aware mapping and scheduling with hardware co-processors for FlexRay-based distributed embedded systems. IEEE Transactions on Parallel and Distributed Systems, 27(10): 3044–3057, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07390091-
abs.html.

Gomes:2013:IPP

[GIP^{+}13] Joao V. Gomes, Pedro R. M. Inacio, Manuela Pereira, Mario M. Freire, and Paulo P. Monteiro. Identification of peer-to-peer VoIP sessions using entropy and codec properties. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2004–2014, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guerreiro:2019:MDG

[GIRT19] J. Guerreiro, A. Ilic, N. Roma, and P. Tomás. Modeling and decoupling the GPU power consumption for crossdomain DVFS. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2494–2506, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goes:2012:ASD

[GIX^{+}12] Luis Fabricio Wanderley Goes, Nikolas Ioannou, Polychronis Xekalakis, Murray Cole, and Marcelo Cintra. Autotuning skeleton-driven optimizations for transactional worklist applications. IEEE Transactions on Parallel and Distributed Systems, 23(12):2205–2218, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2013:PPS

[GJC^{+}13] Deke Guo, Hai Jin, Tao Chen, Jie Wu, Li Lu, Dongsheng Li, and Xiaolei Zhou. Partial probing for scaling overlay routing. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2261–2272, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2012:CTS

[GJLZ12] Yu Gu, Yusheng Ji, Jie Li, and Baohua Zhao. Covering targets in sensor networks: From time domain to space domain. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1643–1656, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2013:EES

[GJLZ13] Yu Gu, Yusheng Ji, Jie Li, and Baohua Zhao. ESWC:

Efficient scheduling for the mobile sink in wireless sensor networks with delay constraint. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1310–1320, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gulisano:2012:SES

[GJPPM^{+}12] Vincenzo Gulisano, Ricardo Jimenez-Peris, Marta Patino-Martinez, Claudio Soriente, and Patrick Valduriez. StreamCloud: An elastic and scalable data streaming system. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2351–2365, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2012:SSC

[GJZZ12] Peng Guo, Tao Jiang, Qian Zhang, and Kui Zhang. Sleep scheduling for critical event monitoring in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(2):345– 352, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gharaibeh:2016:DOR

[GKKW16] Ammar Gharaibeh, Abdallah Khreishah, Issa Khalil, and Jie Wu. Distributed online en-route caching. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3455–3468, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07442568-
abs.html.

Gao:2017:TLB

[GKL^{+}17] Xiaofeng Gao, Linghe Kong, Weichen Li, Wanchao Liang, Yuxiang Chen, and Guihai Chen. Traffic load balancing schemes for devolved controllers in mega data centers. IEEE Transactions on Parallel and Distributed Systems, 28(2):572– 585, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07488213-
abs.html.

Gounaris:2017:DCP

[GKT^{+}17] Anastasios Gounaris, Georgia Kougka, Ruben Tous, Carlos Tripiana Montes, and Jordi Torres. Dynamic configuration of partitioning in spark applications. IEEE Transactions on Parallel and Distributed Systems, 28(7):1891–1904, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

trans/td/2017/07/07807262-
abs.html.

Guerrero:2018:MOO

[GLBJ18] Carlos Guerrero, Isaac Lera, Belen Bermejo, and Carlos Juiz. Multi-objective optimization for virtual machine allocation and replica placement in virtualized Hadoop. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2568–2581, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08360490-
abs.html.

Guo:2015:POR

[GLC^{+}15] Wenzhong Guo, Jie Li, Guolong Chen, Yuzhen Niu, and Chengyu Chen. A PSOoptimized real-time faulttolerant task allocation algorithm in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3236–3249, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06998950-
abs.html.

Gomez-Luna:2013:PMA

[GLGLBM13] Juan Gomez-Luna, Jose Maria Gonzalez-Linares, Jose Ignacio Benavides Benitez, and Nicolas Guil Mata. Performance modeling of atomic additions on GPU scratchpad memory. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2273–2282, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2015:DFT

[GLJ^{+}15] Junyao Guo, Xuefeng Liu, Chunxiao Jiang, Jiannong Cao, and Yong Ren. Distributed fault-tolerant topology control in cooperative wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2699–2710, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06922557.
                                        pdf.

Gopinathan:2011:GSM

[GLL11] Ajay Gopinathan, Zongpeng Li, and Baochun Li. Group strategyproof multicast in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 22(5): 708–715, May 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gong:2015:DDW

[GLL15] Wei Gong, Kebin Liu, and Yunhao Liu. Directional diagnosis for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1290–1300, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803037-
abs.html.

Gregori:2013:PCC

[GLM13] Enrico Gregori, Luciano Lenzini, and Simone Mainardi. Parallel (k)-clique community detection on large-scale networks. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1651–1660, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gaussier:2018:OTE

[GLRT18] Eric Gaussier, Jerome Lelong, Valentin Reis, and Denis Trystram. Online tuning of EASY-backfilling using queue reordering policies. IEEE Transactions on Parallel and Distributed Systems, 29(10):2304–2316, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/10/08327923-
abs.html.

Gao:2011:NID

[GLZ11] Chao Gao, Jiming Liu, and Ning Zhong. Network immunization with distributed autonomy-oriented entities. IEEE Transactions on Parallel and Distributed Systems, 22(7):1222–1229, July 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goh:2014:MSA

[Goh14] Rick Siow Mong Goh. Mapping streaming applications onto GPU systems. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2374–2385, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/09/06574848-
                                        abs.html.

Guo:2012:OPM

[GPF12] Yuanxiong Guo, Miao Pan, and Yuguang Fang. Optimal power management of residential customers in the Smart Grid. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1593–1606, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gowanlock:2019:HAO

[GRB^{+}19] Michael Gowanlock, Cody M. Rude, David M. Blair, Justin D. Li, and Victor Pankratius. A hybrid approach for optimizing parallel clustering throughput using the GPU. IEEE Transactions on Parallel and Distributed Systems, 30(4):766–777, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8462789/

.

Guo:2017:IIH

[GRCZ17] Yanfei Guo, Jia Rao, Dazhao Cheng, and Xiaobo Zhou. iShuffle: Improving Hadoop performance with shuffleon-write. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1649–1662, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07506133-
abs.html.

Guo:2017:MHC

[GRJZ17] Yanfei Guo, Jia Rao, Changjun Jiang, and Xiaobo Zhou. Moving Hadoop into the cloud with flexible slot management and speculative execution. IEEE Transactions on Parallel and Distributed Systems, 28(3):798– 812, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07506079-
abs.html.

Garcia-Rial:2017:RTG

[GRUMG17] Federico Garcia-Rial, Luis Ubeda-Medina, and Jesus Grajal. Real-time GPUbased image processing for a 3-D THz radar. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2953–2964, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07887682-
                                        abs.html.

Gaeta:2011:GPF

[GS11a] Rossano Gaeta and Matteo Sereno. Generalized probabilistic flooding in unstructured peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 22(12):2055–2062, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Goddeke:2011:CRT

[GS11b] Dominik Goddeke and Robert Strzodka. Cyclic reduction tridiagonal solvers on GPUs applied to mixed-precision multigrid. IEEE Transactions on Parallel and Distributed Systems, 22(1):22– 32, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2017:EAA

[GS17] Longkun Guo and Hong Shen. Efficient approximation algorithms for the bounded flexible scheduling problem in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3511–3520, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07990567-
abs.html.

Gao:2019:TEM

[GSH^{+}19] X. Gao, A. Song, L. Hao, J. Zou, G. Chen, and S. Tang. Towards efficient multi-channel data broadcast for multimedia streams. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2370–2383, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gamell:2017:MSM

[GTM^{+}17] Marc Gamell, Keita Teranishi, Jackson Mayo, Hemanth Kolla, Michael A. Heroux, Jacqueline Chen, and Manish Parashar. Modeling and simulating multiple failure masking enabled by local recovery for stencilbased applications at extreme scales. IEEE Transactions on Parallel and Distributed Systems, 28(10):

2881–2895, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07908967-
abs.html.

Gao:2015:GWR

[GTS^{+}15] Chenfei Gao, Jian Tang, Xiang Sheng, Weiyi Zhang, and Chonggang Wang. Greening wireless relay networks: An SNR-aware approach. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3027–3039, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06924781-
abs.html.

Gu:2017:IEC

[GTT^{+}17] Rong Gu, Yun Tang, Chen Tian, Hucheng Zhou, Guanru Li, Xudong Zheng, and Yihua Huang. Improving execution concurrency of large-scale matrix multiplication on distributed dataparallel platforms. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2539–2552, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07884988-
abs.html.

Guan:2014:HHV

[Gua14] Haibing Guan. HYVI: A HYbrid VIrtualization solution balancing performance and manageability. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2332–2341, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06605688-
abs.html.

Guo:2014:TQA

[Guo14] Song Guo. A truthful QoSaware spectrum auction with spatial reuse for large-scale networks. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2499–2508, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06594744-
abs.html.

Guo:2017:AUI

[Guo17]  Deke Guo.  Aggregating

uncertain incast transfers in BCube-like data centers. IEEE Transactions on Parallel and Distributed Systems, 28(4):934–946, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07574280-
abs.html.

Gudmundsson:2015:GAS

[GV15] Joachim Gudmundsson and Nacho Valladares. A GPU approach to subtrajectory clustering using the Fréchet distance. IEEE Transactions on Parallel and Distributed Systems, 26(4):924– 937, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06799188-
                                        abs.html.

Guo:2014:PMO

[GWC14] Ping Guo, Liqiang Wang, and Po Chen. A performance modeling and optimization analysis tool for sparse matrix-vector multiplication on GPUs. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1112–1123, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Guo:2011:QKD

[GWL^{+}11] Deke Guo, Jie Wu, Yunhao Liu, Hai Jin, Hanhua Chen, and Tao Chen. Quasi-Kautz digraphs for peerto-peer networks. IEEE Transactions on Parallel and Distributed Systems, 22(6): 1042–1055, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gao:2017:OMD

[GXW^{+}17] Guoju Gao, Mingjun Xiao, Jie Wu, Kai Han, Liusheng Huang, and Zhenhua Zhao. Opportunistic mobile data offloading with deadline constraints. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3584–3599, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07961189-
abs.html.

Guo:2015:EES

[GXZ^{+}15] Deke Guo, Junjie Xie, Xiaolei Zhou, Xiaomin Zhu, Wei Wei, and Xueshan Luo. Exploiting efficient and scalable shuffle transfers in future data center networks. IEEE Transactions on Parallel and Distributed Systems, 26(4):997–1009, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06787046-
abs.html.

Gu:2018:SAL

[GYLW18] Jiangyuan Gu, Shouyi Yin, Leibo Liu, and Shaojun Wei. Stress-aware loops mapping on CGRAs with dynamic multi-map reconfiguration. IEEE Transactions on Parallel and Distributed Systems, 29(9):

2105–2120, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08319523-
                                        abs.html.

Guan:2015:EES

[GYQW15] Haibing Guan, Jianguo Yao, Zhengwei Qi, and Runze Wang. Energy-efficient SLA guarantees for virtualized GPU in cloud gaming. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2434–2443, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/09/06881719.
                                        pdf.

Guo:2019:EER

[GYW^{+}19] Y. Guo, X. Yuan, X. Wang, C. Wang, B. Li, and X. Jia. Enabling encrypted rich queries in distributed key–value stores. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1283–1297, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2010:NDF

[GYX^{+}10] Wenjun Gu, Zhimin Yang, Dong Xuan, Weijia Jia, and Can Que. Null data frame: a double-edged sword in IEEE 802.11 WLANs. IEEE Transactions on Parallel and Distributed Systems, 21(7): 897–910, July 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Gu:2018:PEQ

[GZW^{+}18] Rong Gu, Yufa Zhou, Zhaokang Wang, Chunfeng Yuan, and Yihua Huang. Penguin: Efficient querybased framework for replaying large scale historical data. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2333–2345, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/10/08345689-
abs.html.

Guo:2014:MLM

[GZWN14] Xiaonan Guo, Dian Zhang, Kaishun Wu, and Lionel M. Ni. MODLoc: Localizing multiple objects in dynamic indoor environment. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2969–2980, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06662344-
abs.html.

Guo:2014:CHS

[GZX14] Song Guo, Deze Zeng, and Yang Xiang. Chameleon hashing for secure and privacy-preserving vehicular communications. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2794–2803, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06654169-
abs.html.

Gu:2015:OTD

[GZY^{+}15] Shouzhen Gu, Qingfeng Zhuge, Juan Yi, Jingtong Hu, and Edwin Hsing-Mean Sha. Optimizing task and data assignment on multicore systems with multiport SPMs. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2549–2560, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06894167.
pdf.

Ge:2013:SAP

[GZZ^{+}13] Aijun Ge, Jiang Zhang, Rui Zhang, Chuangui Ma, and Zhenfeng Zhang. Security analysis of a privacypreserving decentralized keypolicy attribute-based encryption scheme. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2319–2321, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hefeeda:2010:EEP

[HA10] Mohamed Hefeeda and Hossein Ahmadi. Energyefficient protocol for deterministic and probabilistic coverage in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(5):579–593, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2011:HHL

[HA11] Tianyi David Han and Tarek S. Abdelrahman. hiCUDA: High-level GPGPU programming. IEEE Transactions on Parallel and Distributed Systems, 22(1):78– 90, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2013:MDF

[HA13] Miaoqing Huang and David Andrews. Modular design of fully pipelined reduction circuits on FPGAs. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1818–1826, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hoffmann:2012:SSP

[HAD12] Henry Hoffmann, Anant Agarwal, and Srinivas Devadas. Selecting spatiotemporal patterns for development of parallel applications. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1970–1982, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hasan:2016:PAA

[Has16] Sami Hasan. Performanceaware architectures for parallel 4D color fMRI filtering algorithm: A complete performance indices package. IEEE Transactions on Parallel and Distributed Systems, 27(7):2116–2129, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07234938-
abs.html.

Ha:2019:ECS

[HAU19] P. H. Ha, O. J. Anshus, and I. Umar. Efficient concurrent search trees using portable fine-grained locality. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1580–1595, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2018:PTE

[HAY^{+}18] Lixin He, Hong An, Chao Yang, Fei Wang, Junshi Chen, Chao Wang, Weihao Liang, Shaojun Dong, Qiao Sun, Wenting Han, Wenyuan Liu, Yongjian Han, and Wenjun Yao. PEPS++: Towards extreme-scale simulations of strongly correlated quantum many-particle models on Sunway TaihuLight. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2838–2848, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08388281-
abs.html.

Haque:2017:RME

[HAZ17] Mohammad A. Haque, Hakan Aydin, and Dakai Zhu. On reliability management of energy-aware real-time systems through task replication. IEEE Transactions on Parallel and Distributed Systems, 28(3):813– 825, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07544521-
abs.html.

Haidar:2018:GAH

[HAZ^{+}18] Azzam Haidar, Ahmad Abdelfattah, Mawussi Zounon, Stanimire Tomov, and Jack Dongarra. A guide for achieving high performance with very small matrices on GPU: A case study of batched LU and Cholesky factorizations. IEEE Transactions on Parallel and Distributed Systems, 29(5):973– 984, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/05/08214236-
                                        abs.html.

Hu:2012:MDC

[HBF12] Yi Hu, Laxmi N. Bhuyan, and Min Feng. Maintaining data consistency in structured P2P systems. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2125–2137, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hwang:2016:CPM

[HBS^{+}16] Kai Hwang, Xiaoying Bai, Yue Shi, Muyang Li, Wen-Guang Chen, and Yongwei Wu. Cloud performance modeling with benchmark evaluation of elastic scaling strategies. IEEE Transactions on Parallel and Distributed Systems, 27 (1):130–143, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07027861-
                                        abs.html.

Huang:2014:TIA

[HC14] H. Howie Huang and Ron C. Chiang. TRACON: Interferenceaware scheduling for dataintensive applications in virtualized environments. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1349–1358, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hunold:2016:RMB

[HCA16] Sascha Hunold and Alexandra Carpen-Amarie. Reproducible MPI benchmarking is still not as easy as you think. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3617–3630, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07426807-
abs.html.

He:2012:MQS

[HCC^{+}12] Shibo He, Jiming Chen, Peng Cheng, Yu (Jason) Gu, Tian He, and Youxian Sun. Maintaining quality of sensing with actors in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1657–1667, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2015:SDD

[HCG^{+}15] Daojing He, Sammy Chan, Mohsen Guizani, Haomiao Yang, and Boyang Zhou. Secure and distributed data discovery and dissemination in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(4):1129–1139, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06787076-
abs.html.

Hung:2012:SPP

[HCH^{+}12] Ming-Yu Hung, Peng-Sheng Chen, Yuan-Shin Hwang, Roy Dz-Ching Ju, and Jenq-Kuen Lee. Support of probabilistic pointer analysis in the SSA form. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2366–2379, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hong:2019:MHC

[HCH^{+}19] Z. Hong, W. Chen, H. Huang, S. Guo, and Z. Zheng. Multihop cooperative computation offloading for industrial IoT edge cloud computing environments. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2759–2774, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2010:FCC

[HCJ^{+}10] Yu Huang, Jiannong Cao, Beihong Jin, Xianping Tao, Jian Lu, and Yulin Feng. Flexible cache consistency maintenance over wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1150–1161, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2012:LPI

[HCL^{+}12] Shibo He, Jiming Chen, Xu Li, Xuemin (Sherman) Shen, and Youxian Sun. Leveraging prediction to improve the coverage of wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(4): 701–712, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2014:FOS

[HCL^{+}14] Xinyi Huang, Xiaofeng Chen, Jin Li, Yang Xiang, and Li Xu. Further observations on smart-card-based password-authenticated key agreement in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 25(7):1767–1775, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2012:CCD

[HCS12] Shibo He, Jiming Chen, and Youxian Sun. Coverage and connectivity in duty-cycled wireless sensor networks for event monitoring. IEEE Transactions on Parallel and Distributed Systems, 23(3): 475–482, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsiao:2013:LRD

[HCSC13] Hung-Chang Hsiao, Hsueh-Yi Chung, Haiying Shen, and Yu-Chang Chao. Load rebalancing for distributed file systems in clouds. IEEE Transactions on Parallel and Distributed Systems, 24(5): 951–962, May 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2012:EEC

[HCY^{+}12] Shibo He, Jiming Chen, David K. Y. Yau, Huanyu Shao, and Youxian Sun. Energy-efficient capture of stochastic events under periodic network coverage and coordinated sleep. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1090–1102, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hong:2017:PAG

[HCyW^{+}17] Hsiang-Jen Hong, Ge-Ming Chiu, Shiow yang Wu, Tien-Ruey Hsiang, and Tai-Lin Chin. Patron allocation for group services under lower bound constraints. IEEE Transactions on Parallel and Distributed Systems, 28(3): 850–862, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07530888-
abs.html.

Hu:2012:OSG

[HCZ12] Liang Hu, Xi-Long Che, and Si-Qing Zheng. Online system for grid resource monitoring and machine learningbased prediction. IEEE Transactions on Parallel and Distributed Systems, 23(1): 134–145, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Harting:2015:CAM

[HD15] R. Curtis Harting and William J. Dally. On-chip active messages for speed, scalability, and efficiency. IEEE Transactions on Parallel and Distributed Systems, 26(2): 507–515, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2015/02/06747410-
                                        abs.html.

Hei:2015:PIP

[HDL^{+}15] Xiali Hei, Xiaojiang Du, Shan Lin, Insup Lee, and Oleg Sokolsky. Patient infusion pattern based access control schemes for wireless insulin pump system. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3108–3121, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/11/06954561-
                                        abs.html.

Hendrickx:2014:VGW

[Hen14] Julien M. Hendrickx. Views in a graph: To which depth must equality be checked? IEEE Transactions on Parallel and Distributed Systems, 25(7):1907–1912, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2014:CPA

[HFY^{+}14] Weili Han, Zheran Fang, Laurence Tianruo Yang, Gang Pan, and Zhaohui Wu. Collaborative policy administration. IEEE Transactions on Parallel and Distributed Systems, 25(2):498– 507, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Herrero:2012:DCC

[HGC12] Enric Herrero, Jose Gonzalez, and Ramon Canal. Distributed cooperative caching: An energy efficient memory scheme for chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 23(5):853– 861, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2016:CMR

[HGL^{+}16] Huawei Huang, Song Guo, Peng Li, Weifa Liang, and Albert Y. Zomaya. Cost minimization for rule caching in software defined networking. IEEE Transactions on Parallel and Distributed Systems, 27(4):1007–1016, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07105417-
abs.html.

Han:2019:RTB

[HGS^{+}19] M. Han, N. Guan, J. Sun, Q. He, Q. Deng, and W. Liu. Response time bounds for typed DAG parallel tasks on heterogeneous multi-cores. IEEE Transactions on Parallel and Distributed Systems, 30(11):2567–2581, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hackmann:2014:CPC

[HGY^{+}14] Gregory Hackmann, Weijun Guo, Guirong Yan, Zhuoxiong Sun, Chenyang Lu, and Shirley Dyke. Cyber-physical codesign of distributed structural health monitoring with wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(1):63–72, January 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hong:2011:AMA

[HH11] Bo Hong and Zhengyu He. An asynchronous multithreaded algorithm for the maximum network flow problem with nonblocking global relabeling heuristic. IEEE Transactions on Parallel and Distributed Systems, 22(6): 1025–1033, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hubbell:2012:DDT

[HH12] Nicholas Hubbell and Qi Han. DRAGON: Detection and tracking of dynamic amorphous events in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1193–1204, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ha:2013:SWE

[HH13] Sang-Won Ha and Tack-Don Han. A scalable work-efficient and depthoptimal parallel scan for the GPGPU environment. IEEE Transactions on Parallel and Distributed Systems, 24(12):2324–2333, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hegazy:2015:IAC

[HH15] Tamir Hegazy and Mohamed Hefeeda. Industrial automation as a cloud service. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2750–2763, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06908023.
pdf.

Hwang:2010:ESS

[HHK10] Joengmin Hwang, Tian He, and Yongdae Kim. Exploring in-situ sensing irregularity in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(4):547–561, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2017:CCL

[HHWZ17] Rui Han, Siguang Huang, Zhentao Wang, and Jianfeng Zhan. CLAP: Componentlevel approximate processing for low tail latency and high result accuracy in cloud online services. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2190–2203, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07812758-
abs.html.

Hong:2017:GOF

[HJ17] Sumin Hong and Won-Ki Jeong. A group-ordered fast iterative method for eikonal equations. IEEE Transactions on Parallel and Distributed Systems, 28(2):318– 331, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07469401-
abs.html.

Hua:2016:RTS

[HJF16] Yu Hua, Hong Jiang, and Dan Feng. Real-time semantic search using approximate methodology for large-scale storage systems. IEEE Transactions on Parallel and Distributed Systems, 27(4):1212–1225, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2016/04/07091938-
abs.html.

He:2014:CLB

[HJPL14] Jing He, Shouling Ji, Yi Pan, and Yingshu Li. Constructing load-balanced data aggregation trees in probabilistic wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(7):1681–1690, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Heo:2011:OPC

[HJS^{+}11] Jin Heo, Praveen Jayachandran, Insik Shin, Dong Wang, Tarek Abdelzaher, and Xue Liu. OptiTuner: On performance composition and server farm energy minimization application. IEEE Transactions on Parallel and Distributed Systems, 22(11):1871–1878, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hao:2016:ESM

[HJY16] Xingjun Hao, Peiquan Jin, and Lihua Yue. Efficient storage of multi-sensor objecttracking data. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2881–2894, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2016/10/07364275-
                                        abs.html.

Hua:2012:SAM

[HJZ^{+}12] Yu Hua, Hong Jiang, Yifeng Zhu, Dan Feng, and Lei Tian. Semantic-aware metadata organization paradigm in next-generation file systems. IEEE Transactions on Parallel and Distributed Systems, 23(2):337– 344, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2014:SSA

[HjZ^{+}14] Yu Hua, Hong jiang, Yifeng Zhu, Dan Feng, and Lei Xu. SANE: Semantic-aware namespace in ultra-large-scale file systems. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1328–1338, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:2018:VRP

[HK18] Ta-Yuan Hsu and Ajay D. Kshemkalyani. Value the recent past: Approximate causal consistency for partially replicated systems. IEEE Transactions on Parallel and Distributed Systems, 29(1):212–225, January 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08010836-
abs.html.

Heien:2012:CRM

[HKA12] Eric M. Heien, Derrick Kondo, and David P. Anderson. A correlated resource model of Internet end hosts. IEEE Transactions on Parallel and Distributed Systems, 23(6):977–984, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hirabayashi:2016:ADP

[HKE^{+}16] Manato Hirabayashi, Shinpei Kato, Masato Edahiro, Kazuya Takeda, and Seiichi Mita. Accelerated deformable part models on GPUs. IEEE Transactions on Parallel and Distributed Systems, 27(6):1589–1602, June 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07152943-
abs.html.

Hu:2010:DAP

[HKH^{+}10] Chunyu Hu, Hwangnam Kim, Jennifer C. Hou, Dennis Chi, and Sai Shankar N. A distributed approach of proportional bandwidth allocation for real-time services in UltraWideBand (UWB) WPANs. IEEE Transactions on Parallel and Distributed Systems, 21(11):

1626–1643, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hwang:2016:RAP

[HKkY^{+}16] Eunji Hwang, Suntae Kim, Tae kyung Yoo, Jik-Soo Kim, Soonwook Hwang, and Young ri Choi. Resource allocation policies for loosely coupled applications in heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 27(8):2349–2362, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07167692-
                                        abs.html.

Kim:2011:AQM

[hKYY11] Jong hwan Kim, Hyunsoo Yoon, and Ikjun Yeom. Active queue management for flow fairness and stable queue length. IEEE Transactions on Parallel and Distributed Systems, 22(4): 571–579, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2012:CQC

[HL12a] Yuan He and Mo Li. COSE: a query-centric framework of collaborative heterogeneous sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1681–1693, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2012:SHF

[HL12b] Yu Hua and Xue Liu.

Scheduling heterogeneous flows with delay-aware deduplication for avionics applications. IEEE Transactions on Parallel and Distributed Systems, 23(9):1790–1802, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Homsi:2017:WCC

[HLCB^{+}17] Soamar Homsi, Shuo Liu, Gustavo A. Chaparro-Baquero, Ou Bai, Shaolei Ren, and Gang Quan. Workload consolidation for cloud data centers with guaranteed QoS using request reneging. IEEE Transactions on Parallel and Distributed Systems, 28(7): 2103–2116, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07792712-
abs.html.

Hsiao:2011:LBI

[HLCH11] Hung-Chang Hsiao, Hao Liao, Ssu-Ta Chen, and Kuo-Chan Huang. Load balance with imperfect information in structured peerto-peer systems. IEEE Transactions on Parallel and Distributed Systems, 22(4): 634–649, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2015:PPP

[HLeS^{+}15] He Huang, Xiang-Yang Li, Yu e Sun, Hongli Xu, and Liusheng Huang. PPS: Privacy-preserving strategyproof social-efficient spectrum auction mechanisms. IEEE Transactions on Parallel and Distributed Systems, 26(5):1393–1404, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/05/06803033-
                                        abs.html.

Huang:2019:HDM

[HLK^{+}19] Dan Huang, Qing Liu, Scott Klasky, Jun Wang, Jong Youl Choi, Jeremy Logan, and Norbert Podhorszki. Harnessing data movement in virtual clusters for in-situ execution. IEEE Transactions on Parallel and Distributed Systems, 30(3):615–629, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/03/08451897-
                                        abs.html.

Hu:2018:TCE

[HLL18] Zhiming Hu, Baochun Li, and Jun Luo. Time- and cost- efficient task scheduling across geo-distributed data centers. IEEE Transactions on Parallel and Distributed Systems, 29(3):705– 718, March 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8107572/.

Huang:2015:PPR

[HLQ^{+}15a] Jianzhong Huang, Xianhai Liang, Xiao Qin, Qiang Cao, and Changsheng Xie. PUSH: A pipelined reconstruction I/Of or erasurecoded storage clusters. IEEE Transactions on Parallel and Distributed Systems, 26(2): 516–526, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06766677-
abs.html.

Huang:2015:SRE

[HLQ^{+}15b] Jianzhong Huang, Xianhai Liang, Xiao Qin, Ping Xie, and Changsheng Xie. Scale-RS: An efficient scaling scheme for RS-coded storage clusters. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1704–1717, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/

                                        trans/td/2015/06/06819450-
                                        abs.html.

Han:2012:CCC

[HLS^{+}12] Hyuck Han, Young Choon Lee, Woong Shin, Hyungsoo Jung, Heon Y. Yeom, and Albert Y. Zomaya. Cashing in on the cache in the cloud. IEEE Transactions on Parallel and Distributed Systems, 23(8):1387–1399, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2015:SCO

[HLS^{+}15] Biao Han, Jie Li, Jinshu Su, Minyi Guo, and Baokang Zhao. Secrecy capacity optimization via cooperative relaying and jamming for WANETs. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1117–1128, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06784367-
                                        abs.html.

Hu:2014:PRP

[HLWV14] Menglan Hu, Jun Luo, Yang Wang, and Bharadwaj Veeravalli. Practical resource provisioning and caching with dynamic resilience for cloud-based content distribution networks. IEEE Transactions on Parallel and Distributed Systems, 25(8):

2169–2179, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsiao:2010:NOA

[HLY10] Hung-Chang Hsiao, Hao Liao, and Po-Shen Yeh. A near-optimal algorithm attacking the topology mismatch problem in unstructured peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 21(7):983–997, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2014:WSO

[HLY^{+}14] Pei Huang, Chin-Jung Liu, Xi Yang, Li Xiao, and Jin Chen. Wireless spectrum occupancy prediction based on partial periodic pattern mining. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1925–1934, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2019:QSG

[HLYJ19] Q. Hua, Y. Li, D. Yu, and H. Jin. Quasi-streaming graph partitioning: a game theoretical approach. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1643–1656, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2019:WAC

[HLZ^{+}19] R. Han, C. H. Liu, Z. Zong, L. Y. Chen, W. Liu, S. Wang, and J. Zhan. Workloadadaptive configuration tuning for hierarchical cloud schedulers. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2879–2895, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2015:CAE

[HLZY15] Jian-Jun Han, Man Lin, Dakai Zhu, and Laurence T. Yang. Contention-aware energy management scheme for NoC-based multicore realtime systems. IEEE Transactions on Parallel and Distributed Systems, 26(3):691– 701, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06774493-
abs.html.

Hajihassani:2019:FAI

[HMKG19] O. Hajihassani, S. K. Monfared, S. H. Khasteh, and S. Gorgin. Fast AES implementation: A highthroughput bitsliced approach. IEEE Transactions on Parallel and Distributed Systems, 30(10):

2211–2222, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hao:2014:MEF

[HML^{+}14] Fei Hao, Geyong Min, Man Lin, Changqing Luo, and Laurence T. Yang. MobiFuzzyTrust: An efficient fuzzy trust inference mechanism in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2944–2955, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06684155-
abs.html.

He:2019:NTD

[HMP^{+}19] Kun He, Xiaozhu Meng, Zhizhou Pan, Ling Yuan, and Pan Zhou. A novel task-duplication based clustering algorithm for heterogeneous computing environments. IEEE Transactions on Parallel and Distributed Systems, 30(1):2– 14, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08399533-
abs.html.

Hahnel:2018:ECS

[HMS^{+}18] Markus Hahnel, John Martinovic, Guntram Scheithauer, Andreas Fischer, Alexander Schill, and Waltenegus Dargie. Extending the cutting stock problem for consolidating services with stochastic workloads. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2478–2488, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/11/08352052.
                                        pdf.

Hefeeda:2010:BCP

[HN10] Mohamed Hefeeda and Behrooz Noorizadeh. On the benefits of cooperative proxy caching for peer-to-peer traffic. IEEE Transactions on Parallel and Distributed Systems, 21(7): 998–1010, July 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hur:2011:ABA

[HN11] Junbeom Hur and Dong Kun Noh. Attribute-based access control with efficient revocation in data outsourcing systems. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1214–1221, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2012:PDR

[HOZ12] Yunghsiang S. Han, Soji Omiwade, and Rong Zheng. Progressive data retrieval for distributed networked storage. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2303–2314, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hayat:2014:RHD

[HP14] Majeed M. Hayat and Jorge E. Pezoa. Reliability of heterogeneous distributed computing systems in the presence of correlated failures. IEEE Transactions on Parallel and Distributed Systems, 25(4): 1034–1043, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2014:TAa

[HPG14] Jiankun Hu, Hemanshu R.

Pota, and Song Guo. Taxonomy of attacks for agentbased smart grids. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1886–1895, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2012:DTO

[HPH^{+}12] Yu-Kai Huang, Ai-Chun Pang, Pi-Cheng Hsiu, Weihua Zhuang, and Pangfeng Liu. Distributed throughput optimization for ZigBee cluster-tree networks. IEEE Transactions on Parallel and Distributed Systems, 23(3): 513–520, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hwang:2015:RPA

[HPP15] Woomin Hwang, Ki-Woong Park, and Kyu Ho Park. Reference pattern-aware instant memory balancing for consolidated virtual machines on manycores. IEEE Transactions on Parallel and Distributed Systems, 26 (7):2036–2050, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06860294-
abs.html.

Hassan:2017:OTB

[HPPR17] Ahmed Hassan, Roberto Palmieri, Sebastiano Peluso, and Binoy Ravindran. Optimistic transactional boosting. IEEE Transactions on Parallel and Distributed Systems, 28(12):3600–3614, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07974764-
abs.html.

Hesham:2017:SRT

[HRGE17] Salma Hesham, Jens Rettkowski, Diana Goehringer, and Mohamed A. Abd El Ghany. Survey on realtime networks-on-chip. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1500–1517, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07728147-
abs.html.

Hsiao:2012:OOT

[HS12] Hung-Chang Hsiao and Hong-Wei Su. On optimizing overlay topologies for search in unstructured peerto-peer networks. IEEE Transactions on Parallel and Distributed Systems, 23(5): 924–935, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsieh:2014:MIH

[Hsi14] Sun-Yuan Hsieh. A multiindex hybrid trie for lookup and updates. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2486–2498, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06585242-
abs.html.

Han:2012:PPD

[HSMY12] Jinguang Han, Willy Susilo, Yi Mu, and Jun Yan.

Privacy-preserving decentralized key-policy attributebased encryption. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2150–2162, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hong:2017:FFF

[HSN17] Cheol-Ho Hong, Ivor Spence, and Dimitrios S. Nikolopoulos. FairGV: Fair and fast GPU virtualization. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3472–3485, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/07954729-
                                        abs.html.

Han:2011:TBS

[HST^{+}11] Hao Han, Bo Sheng, Chiu C. Tan, Qun Li, and Sanglu Lu. A timing-based scheme for rogue AP detection. IEEE Transactions on Parallel and Distributed Systems, 22(11):1912–1925, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2012:RWS

[HSX^{+}12] Renjie Huang, Wen-Zhan Song, Mingsen Xu, Nina Peterson, Behrooz A. Shirazi, and Richard LaHusen. Real-world sensor network for long-term volcano monitoring: Design and findings. IEEE Transactions on Parallel and Distributed Systems, 23(2):321–329, February 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2016:QHT

[HT16] Nen-Fu Huang and Wen-Yen Tsai. qcAffin: A hardware topology aware interrupt affinitizing and balancing scheme for multicore and multi-queue packet processing systems. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1783–1795, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07152926-
abs.html.

Ha:2010:SPC

[HTA10] Phuong Hoai Ha, Philippas Tsigas, and Otto J. Anshus. The synchronization power of coalesced memory accesses. IEEE Transactions on Parallel and Distributed Systems, 21(7):939– 953, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Han:2017:RSM

[HTZY17] Jian-Jun Han, Xin Tao, Dakai Zhu, and Laurence T.

Yang. Resource sharing in multicore mixed-criticality systems: Utilization bound and blocking overhead. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3626–3641, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/07869415-
                                        abs.html.

Hu:2014:PRS

[Hu14] Chia-Cheng Hu. Playbackrate segment scheduling algorithm in MoD P2P streaming services. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1585–1599, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hur:2013:ABS

[Hur13] Junbeom Hur. Attributebased secure data sharing with hidden policies in smart grid. IEEE Transactions on Parallel and Distributed Systems, 24(11):2171–2180, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2011:RAS

[HV11] Menglan Hu and Bharadwaj Veeravalli. Requirementaware strategies with arbitrary processor release times for scheduling multiple divisible loads. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1697–1704, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsiao:2013:BPS

[HW13] Hsu-Feng Hsiao and Chen-Tsang Wu. Balanced parallel scheduling for video encoding with adaptive GOP structure. IEEE Transactions on Parallel and Distributed Systems, 24(12):2355–2364, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2014:MAR

[HWC^{+}14] Jianhui Huang, Shengling Wang, Xiuzhen Cheng, Min Liu, Zhongcheng Li, and Biao Chen. Mobility-assisted routing in intermittently connected mobile cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2956–2968, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06674927-
abs.html.

Hakkarinen:2015:FSF

[HWC15] Doug Hakkarinen, Panruo Wu, and Zizhong Chen. Failstop failure algorithm-based fault tolerance for Cholesky decomposition. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1323–1335, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06805637-
abs.html.

Huang:2010:MLG

[HWDP10] Scott C.-H. Huang, Peng-Jun Wan, Hongwei Du, and E.-K. Park. Minimum latency gossiping in radio networks. IEEE Transactions on Parallel and Distributed Systems, 21(6):790– 800, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hou:2018:FAV

[HWF18] Kaixi Hou, Hao Wang, and Wu-Chun Feng. A framework for the automatic vectorization of parallel sort on x86-based processors. IEEE Transactions on Parallel and Distributed Systems, 29(5): 958–972, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08247204-
abs.html.

Han:2019:RAS

[HWG^{+}19] J. Han, Z. Wang, S. Gong, T. Miao, and L. T. Yang. Resource-aware scheduling for dependable multicore real-time systems: Utilization bound and partitioning algorithm. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2806–2819, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2012:MBW

[HWI12] Scott C.-H. Huang, Hsiao-Chun Wu, and Sundaraja Sitharama Iyengar. Multisource broadcast in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(10):1908–1914, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hada:2018:SMC

[HWJ18] Romas James Hada, Hongyi Wu, and Miao Jin. Scalable minimum-cost balanced partitioning of large-scale social networks: Online and offline solutions. IEEE Transactions on Parallel and Distributed Systems, 29(7): 1636–1649, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/07902221-
abs.html.

He:2017:CAR

[HWL^{+}17a] Shuibing He, Yang Wang, Zheng Li, Xian-He Sun, and Chenzhong Xu. Cost-aware region-level data placement in multi-tiered parallel I/O systems. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1853–1865, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07776880-
abs.html.

Hu:2017:OSF

[HWL^{+}17b] Xiameng Hu, Xiaolin Wang, Yechen Li, Yingwei Luo, Chen Ding, and Zhenlin Wang. Optimal symbiosis and fair scheduling in shared cache. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1134–1148, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07572145-
abs.html.

Huang:2015:SCC

[HWNS15] Zhen Huang, Cheng Wang, Amiya Nayak, and Ivan Stojmenovic. Small cluster in cyber physical systems: Network topology, interdependence and cascading failures. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2340–2351, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06863675-
abs.html.

Huang:2015:EPE

[HWQ^{+}15] Jianzhong Huang, Yanqun Wang, Xiao Qin, Xianhai Liang, Shu Yin, and Changsheng Xie. Exploiting pipelined encoding process to boost erasure-coded data archival. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2984–2996, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06942231-
abs.html.

He:2016:BPF

[HWS16a] Shuibing He, Yang Wang, and Xian-He Sun. Boosting parallel file system performance via heterogeneityaware selective data layout. IEEE Transactions on Parallel and Distributed Systems, 27(9):2492–2505, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07345596-
abs.html.

He:2016:IPP

[HWS16b] Shuibing He, Yang Wang, and Xian-He Sun. Improving performance of parallel I/O systems through selective and layout-aware SSD cache. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2940–2952, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07390309-
abs.html.

He:2017:UMM

[HWSX17] Shuibing He, Yang Wang, Xian-He Sun, and Chengzhong Xu. Using MinMax-memory claims to improve in-memory workflow computations in the cloud. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1202–1214, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07579169-
abs.html.

Huang:2012:IEE

[HWX12] Pei Huang, Chen Wang, and Li Xiao. Improving end-to-end routing performance of greedy forwarding in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(3):556– 563, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2010:OSA

[HWZE10] Min He, Xiaolong Wu, Si Qing Zheng, and Burkhard Englert. Optimal sorting algorithms for a simplified 2D array with reconfigurable pipelined bus system. IEEE Transactions on Parallel and Distributed Systems, 21(3): 303–312, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hu:2010:BOL

[HX10] Haibo Hu and Jianliang Xu. 2PASS: Bandwidthoptimized location cloaking for anonymous locationbased services. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1458–1472, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2011:GFT

[HXC^{+}11] Xinyi Huang, Yang Xiang, Ashley Chonka, Jianying Zhou, and Robert H. Deng. A generic framework for three-factor authentication: Preserving security and privacy in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1390–1397, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2019:ITP

[HxjGG19] Q. He, x. jiang, N. Guan, and Z. Guo. Intra-task priority assignment in real-time scheduling of DAG tasks on multi-cores. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2283–2295, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hua:2015:DIL

[HXLF15] Yu Hua, Bin Xiao, Xue Liu, and Dan Feng. The design and implementations of locality-aware approximate queries in hybrid storage systems. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3194–3207, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/11/06948249-
                                        abs.html.

Huang:2012:RDC

[HYC^{+}12] Yu Huang, Yiling Yang, Jiannong Cao, Xiaoxing Ma, Xianping Tao, and Jian Lu. Runtime detection of the concurrency property in asynchronous pervasive computing environments. IEEE Transactions on Parallel and Distributed Systems, 23(4): 744–750, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Huang:2011:TAS

[HYX11] Lin Huang, Feng Yuan, and Qiang Xu. On task allocation and scheduling for lifetime extension of platformbased MPSoC designs. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2088–2099, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2015:IUI

[HYZ15] Bingsheng He, Jeffrey Xu Yu, and Amelie Chi Zhou. Improving update-intensive workloads on flash disks through exploiting multichip parallelism. IEEE Transactions on Parallel and Distributed Systems, 26 (1):152–162, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748022-
abs.html.

Huang:2016:HJE

[HZB^{+}16] Jin Huang, Rui Zhang, Rajkumar Buyya, Jian Chen, and Yongwei Wu. HEADS-JOIN: Efficient Earth Mover’s Distance similarity joins on Hadoop. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1660–1673, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07172540-
abs.html.

Hong:2017:FFS

[HZG^{+}17] Yang Hong, Yang Zheng, Haibing Guan, Binyu Zang, and Haibo Chen. Fencefree synchronization with dynamically serialized synchronization variables. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3486–3500, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07762194-
abs.html.

Hua:2011:SSA

[HZJ^{+}11] Yu Hua, Yifeng Zhu, Hong Jiang, Dan Feng, and Lei Tian. Supporting scalable and adaptive metadata management in ultralarge-scale file systems. IEEE Transactions on Parallel and Distributed Systems, 22(4):580– 593, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

He:2016:DGB

[HZJ16] Ligang He, Huanzhou Zhu, and Stephen A. Jarvis. Developing graph-based coscheduling algorithms on multicore computers. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1617–1632, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07194818-
abs.html.

HoseinyFarahabady:2018:MPC

[HZT18] M. Reza HoseinyFarahabady, Albert Y. Zomaya, and Zahir Tari. A model predictive controller for managing QoS enforcements and microarchitecture-level interferences in a Lambda platform. IEEE Transactions on Parallel and Distributed Systems, 29(7):1442–1455, July 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08126823-
abs.html.

Han:2014:MRT

[HZW^{+}14] Jian-Jun Han, Dakai Zhu, Xiaodong Wu, Laurence T. Yang, and Hai Jin. Multiprocessor real-time systems with shared resources: Utilization bound and mapping. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2981–2991, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2014/11/06678507-
                                        abs.html.

Hu:2019:LDC

[HZW^{+}19] M. Hu, L. Zhuang, D. Wu, Y. Zhou, X. Chen, and L. Xiao. Learning driven computation offloading for asymmetrically informed edge computing. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1802–1815, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iancu:2014:CPV

[Ian14] Costin Iancu. The case for partitioning virtual machines on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2683–2696, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06613989-
                                        abs.html.

Iacovazzi:2014:ITP

[IB14] Alfonso Iacovazzi and Andrea Baiocchi. Internet traffic privacy enhancement with masking: Optimization and tradeoffs. IEEE Transactions on Parallel and Distributed Systems, 25(2):353– 362, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Isaila:2011:DEM

[IBC^{+}11] Florin Isaila, Javier Garcia Blas, Jesus Carretero, Robert Latham, and Robert Ross. Design and evaluation of multiple-level data staging for Blue Gene systems. IEEE Transactions on Parallel and Distributed Systems, 22(6): 946–959, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ivashko:2018:SDG

[ICN18] Evgeny Ivashko, Ilya Chernov, and Natalia Nikitina. A survey of desktop grid scheduling. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2882–2895, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08395065-
abs.html.

Ishii:2012:ODA

[IdM12] Renato Porfirio Ishii and Rodrigo Fernandes de Mello. An online data access prediction and optimization approach for distributed systems. IEEE Transactions on Parallel and Distributed Systems, 23(6):1017–1029, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ismail:2011:PEC

[IG11] Leila Ismail and Driss Guerchi. Performance evaluation of convolution on the cell broadband engine processor. IEEE Transactions on Parallel and Distributed Systems, 22(2):337– 351, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ibrahim:2011:PAA

[IGEN11] Atef Ibrahim, Fayez Gebali, Hamed Elsimary, and Amin Nassar. Processor array architectures for scalable radix 4 Montgomery modular multiplication algorithm. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1142–1149, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Izumi:2013:FPT

[IIKO13] Taisuke Izumi, Tomoko Izumi, Sayaka Kamei, and Fukuhito Ooshita. Feasibility of polynomial-time randomized gathering for oblivious mobile robots. IEEE Transactions on Parallel and Distributed Systems, 24(4): 716–723, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ino:2012:SHS

[IMH12] Fumihiko Ino, Yuma Munekawa, and Kenichi Hagihara. Sequence homology search using fine grained cycle sharing of idle GPUs. IEEE Transactions on Parallel and Distributed Systems, 23(4): 751–759, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iosup:2011:PAC

[IOY^{+}11] Alexandru Iosup, Simon Ostermann, M. Nezih Yigitbasi, Radu Prodan, Thomas Fahringer, and Dick H. J. Epema. Performance analysis of cloud computing services for many-tasks scientific computing. IEEE Transactions on Parallel and Distributed Systems, 22(6): 931–945, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ianni:2019:ARC

[IPQ19] Mauro Ianni, Alessandro Pellegrini, and Francesco Quaglia. Anonymous readers counting: A wait-free multi-word atomic register algorithm for scalable data sharing on multi-core machines. IEEE Transactions on Parallel and Distributed Systems, 30(2):286– 299, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/02/08438921-
                                        abs.html.

Izhak-Ratzin:2012:OLB

[IRPvdS12] Rafit Izhak-Ratzin, Hyunggon Park, and Mihaela van der Schaar. Online learning in BitTorrent systems. IEEE Transactions on Parallel and Distributed Systems, 23(12):2280–2288, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iamnitchi:2011:SWF

[IRSNF11] Adriana Iamnitchi, Matei Ripeanu, Elizeu Santos-Neto, and Ian Foster. The small world of file sharing. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1120–1134, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Inggs:2017:DSA

[ITL17] Gordon Inggs, David B. Thomas, and Wayne Luk. A domain specific approach to high performance heterogeneous computing. IEEE Transactions on Parallel and Distributed Systems, 28(1): 2–15, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/

trans/td/2017/01/07465804-
abs.html.

Iskander:2014:BPA

[ITW^{+}14] Marian K. Iskander, Tucker Trainor, Dave W. Wilkinson, Adam J. Lee, and Panos K. Chrysanthis. Balancing performance, accuracy, and precision for secure cloud transactions. IEEE Transactions on Parallel and Distributed Systems, 25(2):417– 426, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iwanicki:2010:GBS

[IvS10] Konrad Iwanicki and Maarten van Steen. Gossip-based self-management of a recursive area hierarchy for large wireless SensorNets. IEEE Transactions on Parallel and Distributed Systems, 21(4): 562–576, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Iyengar:2014:TSC

[Iye14] S. S. Iyengar. Towards safe cities: A mobile and social networking approach. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2451–2462, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06574841-
abs.html.

Iranfar:2018:MLB

[IZA18] Arman Iranfar, Marina Zapater, and David Atienza. Machine learning-based qualityaware power and thermal management of multistream HEVC encoding on multicore servers. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2268–2281, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/10/08338403-
                                        abs.html.

Jiang:2019:BAT

[JAJ^{+}19] J. Jiang, B. An, Y. Jiang, P. Shi, Z. Bu, and J. Cao. Batch allocation for tasks with overlapping skill requirements in crowdsourcing. IEEE Transactions on Parallel and Distributed Systems, 30(8):1722–1737, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ji:2012:CDC

[JCLJ12] Shouling Ji, Zhipeng Cai, Yingshu Li, and Xiaohua Jia. Continuous data collection capacity of dual-radio multichannel wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1844–1855, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2012:GFE

[JCW^{+}12] Hongbo Jiang, Jie Cheng, Dan Wang, Chonggang Wang, and Guang Tan. A general framework for efficient continuous multidimensional top-k query processing in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1668–1680, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2019:TLL

[JCW^{+}19] Hai Jin, Fei Chen, Song Wu, Yin Yao, Zhiyi Liu, Lin Gu, and Yongluan Zhou. Towards low-latency batched stream processing by pre-scheduling. IEEE Transactions on Parallel and Distributed Systems, 30(3):710–722, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08444732-
abs.html.

Jin:2010:DPS

[JCWB10] Xing Jin, S.-H. Gary Chan, Wan-Ching Wong, and Ali C. Begen. A distributed protocol to serve dynamic groups for peer-to-peer streaming. IEEE Transactions on Parallel and Distributed Systems, 21(2):216–228, February 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jenkins:2014:PMD

[JDB^{+}14] John Jenkins, James Dinan, Pavan Balaji, Tom Peterka, Nagiza F. Samatova, and Rajeev Thakur. Processing MPI derived datatypes on noncontiguous GPUresident data. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2627–2637, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06600679-
                                        abs.html.

Jimborean:2018:ADL

[JEW^{+}18] Alexandra Jimborean, Per Ekemark, Jonatan Waern, Stefanos Kaxiras, and Alberto Ros. Automatic detection of large extended data-race-free regions with conflict isolation. IEEE Transactions on Parallel and Distributed Systems, 29 (3):527–541, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8100964/.

Jain:2017:APA

[JFP^{+}17] Chirag Jain, Patrick Flick, Tony Pan, Oded Green, and Srinivas Aluru. An adaptive parallel algorithm for computing connected components. IEEE Transactions on Parallel and Distributed Systems, 28(9):2428–2439, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07862286-
abs.html.

Jeong:2011:TBD

[JGG^{+}11] Jaehoon (Paul) Jeong, Shuo Guo, Yu (Jason) Gu, Tian He, and David H. C. Du. Trajectory-based data forwarding for light-traffic vehicular ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 22(5):743–757, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jeong:2010:VSA

[JGHD10] Jaehoon Jeong, Yu Gu, Tian He, and David H. C. Du. Virtual scanning algorithm for road network surveillance. IEEE Transactions on Parallel and Distributed Systems, 21(12):1734–1749, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jia:2018:PMP

[JGJF18] Qi Jia, Linke Guo, Zhanpeng Jin, and Yuguang Fang.

Preserving model privacy for machine learning in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1808–1822, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08302601-
abs.html.

Jiang:2014:EEI

[JGZZ14] Wenchao Jiang, Haibing Guan, Qian Zhang, and Yanmin Zhu. Energy-efficient identification in large-scale RFID systems with handheld reader. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1211–1222, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jafarpour:2012:CDP

[JHMV12] Hojjat Jafarpour, Bijit Hore, Sharad Mehrotra, and Nalini Venkatasubramanian. CCD: a distributed publish/subscribe framework for rich content formats. IEEE Transactions on Parallel and Distributed Systems, 23(5): 844–852, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jeyapaul:2014:UCM

[JHR^{+}14] Reiley Jeyapaul, Fei Hong, Abhishek Rhisheekesan, Aviral Shrivastava, and Kyoungwoo Lee. UnSync-CMP: Multicore CMP architecture for energy-efficient soft-error reliability. IEEE Transactions on Parallel and Distributed Systems, 25(1):254– 263, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jackson:2015:OPT

[JHR15] Adrian Jackson, Joachim Hein, and Colin Roach. Optimising performance through unbalanced decompositions. IEEE Transactions on Parallel and Distributed Systems, 26(10):2863–2873, October 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06883171.
pdf.

Jiang:2015:EDT

[JHW^{+}15] Jinfang Jiang, Guangjie Han, Feng Wang, Lei Shu, and Mohsen Guizani. An efficient distributed trust model for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(5):1228–1237, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06805612-
abs.html.

Jin:2019:EET

[JHWY19] P. Jin, X. Hao, X. Wang, and L. Yue. Energy-efficient task scheduling for CPU-intensive streaming jobs on Hadoop. IEEE Transactions on Parallel and Distributed Systems, 30(6):1298–1311, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jung:2011:AUT

[JHYK11] Hyungsoo Jung, Hyuck Han, Heon Y. Yeom, and Sooyong Kang. Athanasia: a user-transparent and faulttolerant system for parallel applications. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1653–1668, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2014:APP

[Jia14a] Hongbo Jiang. On arbitrating the power-performance tradeoff in SaaS clouds. IEEE Transactions on Parallel and Distributed Systems, 25(10):2648–2658, October 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06582405-
                                        abs.html.

Jiang:2014:USN

[Jia14b] J. C. Jiang. Understanding social networks from a multiagent perspective. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2743–2759, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06620863-
abs.html.

Jiang:2016:STA

[Jia16]  Yichuan Jiang.  A sur-

vey of task allocation and load balancing in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 27(2):585– 599, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07051215-
abs.html.

Javaid:2014:PEP

[JIP14] Haris Javaid, Aleksander Ignjatovic, and Sri Parameswaran. Performance estimation of pipelined MultiProcessor System-on-Chips (MPSoCs). IEEE Transactions on Parallel and Distributed Systems, 25(8):2159–2168, August 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jeng:2011:ATC

[JJ11] Andy An-Kai Jeng and Rong-Hong Jan. Adaptive topology control for mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 22(12): 1953–1960, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2012:ITA

[JJG^{+}12] Yichao Jin, Jiong Jin, Alexander Gluhak, Klaus Moessner, and Marimuthu Palaniswami. An intelligent task allocation scheme for multihop wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(3):444–451, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2011:PEE

[JJW11] Hongbo Jiang, Shudong Jin, and Chonggang Wang. Prediction or not? An energy-efficient framework for clustering-based data collection in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(6): 1064–1071, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2012:CAG

[JKP12] Yuho Jin, Eun Jung Kim, and Timothy Mark Pinkston. Communication-aware globallycoordinated on-chip networks. IEEE Transactions on Parallel and Distributed Systems, 23(2):242– 254, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jaho:2013:SSF

[JKS13] Eva Jaho, Merkourios Karaliopoulos, and Ioannis Stavrakakis. Social similarity favors cooperation: The distributed content replication case. IEEE Transactions on Parallel and Distributed Systems, 24(3):601– 613, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jarecki:2011:FRG

[JKT11] Stanisław Jarecki, Jihye Kim, and Gene Tsudik. Flexible robust group key agreement. IEEE Transactions on Parallel and Distributed Systems, 22(5):879– 886, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Javadi:2011:DSM

[JKVA11] Bahman Javadi, Derrick Kondo, Jean-Marc Vincent, and David P. Anderson. Discovering statistical models of availability in large distributed systems: An empirical study of SETI@home. IEEE Transactions on Parallel and Distributed Systems, 22(11):1896–1903, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2017:QWS

[JLKG17] Jiahui Jin, Junzhou Luo, Samamon Khemmarat, and Lixin Gao. Querying Web-scale knowledge graphs through effective pruning of search space. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2342–2356, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07845718-
                                        abs.html.

Jiao:2012:MLB

[JLM^{+}12] Xianlong Jiao, Wei Lou, Junchao Ma, Jiannong Cao, Xiaodong Wang, and Xingming Zhou. Minimum latency broadcast scheduling in duty-cycled multihop wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(1): 110–117, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2010:CBS

[JLW^{+}10] Hongbo Jiang, Wenping Liu, Dan Wang, Chen Tian, Xiang Bai, Xue Liu, Ying Wu, and Wenyu Liu. Connectivity-based skeleton extraction in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(5):710– 721, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jaulmes:2018:AEF

[JMA^{+}18] Luc Jaulmes, Miquel Moreto, Eduard Ayguade, Jesus Labarta, Mateo Valero, and Marc Casas. Asynchronous and exact forward recovery for detected errors in iterative solvers. IEEE Transactions on Parallel and Distributed Systems, 29(9): 1961–1974, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08320336-
abs.html.

Jiang:2018:TQA

[JMS^{+}18] Wenjun Jiang, Chenglin Miao, Lu Su, Qi Li, Shaohan Hu, Shiguang Wang, Jing Gao, Hengchang Liu, Tarek F. Abdelzaher, Jiawei Han, Xue Liu, Yan Gao, and Lance Kaplan. Towards quality aware information integration in distributed sensing systems. IEEE Transactions on Parallel and Distributed Systems, 29(1):198– 211, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/01/07940106-
                                        abs.html.

Jalili-Marandi:2012:LST

[JMZD12] Vahid Jalili-Marandi, Zhiyin Zhou, and Venkata Dinavahi. Large-scale transient stability simulation of electrical power systems on parallel GPUs. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1255–1266, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2016:PMV

[JN16] Caoyang Jiang and Saeid Nooshabadi. Parallel multiview video coding exploiting group of pictures level parallelism. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2316–2328, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07287762-
                                        abs.html.

Jo:2015:ALM

[JNL^{+}15] Gangwon Jo, Jeongho Nah, Jun Lee, Jungwon Kim, and Jaejin Lee. Accelerating LINPACK with MPI-OpenCL on clusters of multi-GPU nodes. IEEE Transactions on Parallel and Distributed Systems, 26(7):

1814–1825, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06846313-
abs.html.

Jha:2012:HDR

[JP12] Pranava K. Jha and Rachna Prasad. Hamiltonian decomposition of the rectangular twisted torus. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1504–1507, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2014:SMA

[JPG14] Weirong Jiang, Viktor K.

Prasanna, and Thilan Ganegedara. A scalable and modular architecture for high-performance packet classification. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1135–1144, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jeanneau:2017:SSA

[JRAS17] Denis Jeanneau, Thibault Rieutord, Luciana Arantes, and Pierre Sens. Solving k-set agreement using failure detectors in unknown dynamic networks. IEEE Transactions on Parallel and Distributed Systems, 28(5):

1484–1499, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07565628-
abs.html.

Jonathan:2017:NDE

[JRO^{+}17] Albert Jonathan, Mathew Ryden, Kwangsung Oh, Abhishek Chandra, and Jon Weissman. Nebula: Distributed edge cloud for data intensive computing. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3229–3242, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07954728-
abs.html.

Johnson:2010:SMA

[JRP^{+}10] Matthew P. Johnson, Hosam Rowaihy, Diego Pizzocaro, Amotz Bar-Noy, Stuart Chalmers, Thomas F. La Porta, and Alun Preece. Sensor-mission assignment in constrained environments. IEEE Transactions on Parallel and Distributed Systems, 21(11):1692–1705, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jia:2013:SRU

[JRV^{+}13] Adele L. Jia, Rameez Rahman, Tamas Vinko, Johan A. Pouwelse, and Dick H. J. Epema. Systemic risk and user-level performance in private P2P communities. IEEE Transactions on Parallel and Distributed Systems, 24(12):2503–2512, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jahan:2018:LWW

[JRZ^{+}18] Mosarrat Jahan, Mohsen Rezvani, Qianrui Zhao, Partha Sarathi Roy, Kouichi Sakurai, Aruna Seneviratne, and Sanjay Jha. Light weight write mechanism for cloud data. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1131–1146, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08186229-
abs.html.

Jiang:2017:OFU

[JSC^{+}17] Weiwen Jiang, Edwin Hsing-Mean Sha, Xianzhang Chen, Lei Yang, Lei Zhou, and Qingfeng Zhuge. Optimal functional-unit assignment for heterogeneous systems under timing constraint. IEEE Transactions on Parallel and Distributed Systems, 28(9):2567–2580, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2017/09/07867836-
                                        abs.html.

Jamour:2018:PAI

[JSK18] Fuad Jamour, Spiros Skiadopoulos, and Panos Kalnis. Parallel algorithm for incremental betweenness centrality on large graphs. IEEE Transactions on Parallel and Distributed Systems, 29(3):659–672, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8070346/.

Ji:2019:PWS

[JSLD19] S. Ji, N. Satish, S. Li, and P. K. Dubey. Parallelizing Word2Vec in shared and distributed memory. IEEE Transactions on Parallel and Distributed Systems, 30(9): 2090–2100, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jang:2011:EMA

[JSMK11] Byunghyun Jang, Dana Schaa, Perhaad Mistry, and David Kaeli. Exploiting memory access patterns to improve memory performance in data-parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 22(1):105– 118, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2011:COJ

[JTS^{+}11] Yunlian Jiang, Kai Tian, Xipeng Shen, Jinghe Zhang, Jie Chen, and Rahul Tripathi. The complexity of optimal job co-scheduling on chip multiprocessors and heuristics-based solutions. IEEE Transactions on Parallel and Distributed Systems, 22(7):1192–1205, July 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jung:2017:EPD

[Jun17] Myoungsoo Jung. Exploring parallel data access methods in emerging non-volatile memory systems. IEEE Transactions on Parallel and Distributed Systems, 28(3): 746–759, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07514757-
abs.html.

Jia:2010:SMD

[JVW10] Jingxi Jia, Bharadwaj Veeravalli, and Jon Weissman. Scheduling multisource divisible loads on arbitrary networks. IEEE Transactions on Parallel and Distributed Systems, 21(4):520– 531, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jawhar:2010:RSW

[JWA10] Imad Jawhar, Jie Wu, and Dharma P. Agrawal. Resource scheduling in wireless networks using directional antennas. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1240–1253, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jayaram:2015:SNE

[JWE15] K. R. Jayaram, Weihang Wang, and Patrick Eugster. Subscription normalization for effective contentbased messaging. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3184–3193, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/11/06894211-
                                        abs.html.

Jiang:2014:CSA

[JWJS14] Guiyuan Jiang, Kai Wang, Wu Jigang, and Thambipillai Srikanthan. Constructing sub-arrays with short interconnects from degradable VLSI arrays. IEEE Transactions on Parallel and Distributed Systems, 25(4): 929–938, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jang:2016:WAO

[JWK^{+}16] Jae Young Jang, Hao Wang, Euijin Kwon, Jae W. Lee, and Nam Sung Kim. Workload-aware optimal power allocation on single-chip heterogeneous processors. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1838–1851, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07152947-
abs.html.

Jiao:2019:AMC

[JWNS19] Y. Jiao, P. Wang, D. Niyato, and K. Suankaewmanee. Auction mechanisms in cloud/fog computing resource allocation for public blockchain networks. IEEE Transactions on Parallel and Distributed Systems, 30(9): 1975–1989, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jin:2018:CMD

[JWY^{+}18] Hai Jin, Na Wang, Dongxiao Yu, Qiang-Sheng Hua, Xuanhua Shi, and Xia Xie. Core maintenance in dynamic graphs: A parallel approach based on matching. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2416–2428, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/11/08357916-
                                        abs.html.

Jeon:2015:MTH

[JY15] Yongkweon Jeon and Sungroh Yoon. Multi-threaded hierarchical clustering by parallel nearest-neighbor chaining. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2534–2548, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/09/06893001.
                                        pdf.

Jiang:2014:ETB

[JZH^{+}14] Ruobing Jiang, Yanmin Zhu, Tian He, Yunhuai Liu, and Lionel M. Ni. Exploiting trajectory-based coverage for geocast in vehicular networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3177–3189, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/12/06714420-
                                        abs.html.

Jiang:2013:TAU

[JZW13] Yichuan Jiang, Yifeng Zhou, and Wanyuan Wang. Task allocation for undependable multiagent systems in social networks. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1671–1681, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jiang:2014:FDS

[JZW^{+}14] Hao Jiang, Jiannan Zhai, Sally K. Wahba, Biswajit Mazumder, and Jason O. Hallstrom. Fast distributed simulation of sensor networks using optimistic synchronization. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2888–2898, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06636890-
abs.html.

Jia:2017:UBD

[JZW^{+}17] Zhen Jia, Jianfeng Zhan, Lei Wang, Chunjie Luo, Wanling Gao, Yi Jin, Rui Han, and Lixin Zhang. Understanding big data analytics workloads on modern processors. IEEE Transactions on Parallel and Distributed Systems, 28(6):1797–1810, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07736117-
abs.html.

Jiang:2015:TET

[JZWN15] Ruobing Jiang, Yanmin Zhu, Xin Wang, and Lionel M.

Ni. TMC: Exploiting trajectories for multicast in sparse vehicular networks. IEEE Transactions on Parallel and Distributed Systems, 26 (1):262–271, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06747403-
abs.html.

Jia:2015:TDH

[JZY^{+}15] Riheng Jia, Jinbei Zhang, Feng Yang, Xiaoying Gan, Xiaohua Tian, Pengyuan Du, and Xinbing Wang. Throughput and delay in heterogeneous cognitive radio networks with cooperative secondary users. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2586–2598, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06894230.
pdf.

Jiang:2015:DMS

[JZZ^{+}15] Yu Jiang, Hehua Zhang, Huafeng Zhang, Han Liu, Xiaoyu Song, Ming Gu, and Jiaguang Sun. Design of mixed synchronous/ asynchronous systems with multiple clocks. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2220–2232, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06874581-
abs.html.

Karsavuran:2016:LAP

[KAA16] M. Ozan Karsavuran, Kadir Akbudak, and Cevdet Aykanat. Locality-aware parallel sparse matrix–vector and matrixtranspose–vector multiplication on many-core processors. IEEE Transactions on Parallel and Distributed Systems, 27(6):1713–1726, June 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07152923-
abs.html.

Kotha:2015:APU

[KAC^{+}15] Aparna Kotha, Kapil Anand, Timothy Creech, Khaled El-Wazeer, Matthew Smithson, Greeshma Yellareddy, and Rajeev Barua. Affine parallelization using dependence and cache analysis in a binary rewriter. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2154–2163, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06880338-
abs.html.

Kylasa:2017:RMD

[KAG17] Sudhir B. Kylasa, Hasan Metin Aktulga, and Ananth Y. Grama. Reactive molecular dynamics on massively parallel heterogeneous architectures. IEEE Transactions on Parallel and Distributed Systems, 28(1):202– 214, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/01/07444209-
                                        abs.html.

Kurzak:2016:ITB

[KAGD16] Jakub Kurzak, Hartwig Anzt, Mark Gates, and Jack Dongarra. Implementation and tuning of batched Cholesky factorization and solve for NVIDIA GPUs. IEEE Transactions on Parallel and Distributed Systems, 27(7):2036–2048, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/07/07275187-
                                        abs.html.

Khamse-Ashari:2018:EFM

[KALK^{+}18] Jalal Khamse-Ashari, Ioannis Lambadaris, George Kesidis, Bhuvan Urgaonkar, and Yiqiang Zhao. An efficient and fair multi-resource allocation mechanism for heterogeneous servers. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2686–2699, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08368291-
abs.html.

Kao:2015:POP

[Kao15] Chi-Chou Kao. Performanceoriented partitioning for task scheduling of parallel reconfigurable architectures. IEEE Transactions on Parallel and Distributed Systems, 26(3):858–867, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06777320-
abs.html.

Kim:2017:OEE

[KAV^{+}17] Youngjae Kim, Scott Atchley, Geoffroy R. Vallee, Sangkeun Lee, and Galen M. Shipman. Optimizing endto-end big data transfers over terabits network infrastructure. IEEE Transactions on Parallel and Distributed Systems, 28(1):188– 201, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07447811-
abs.html.

Kini:2013:SA

[KB13] P. Kini and K. Beznosov.

Speculative authorization. IEEE Transactions on Parallel and Distributed Systems, 24(4):814–824, April 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kleppmann:2017:CFR

[KB17] Martin Kleppmann and Alastair R. Beresford. A conflict-free replicated JSON datatype. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2733–2746, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07909007-
                                        abs.html.

Kowarzyk:2014:OPT

[KBHS14] Gilbert Kowarzyk, Normand Belanger, David Haccoun, and Yvon Savaria. Optimizing the parallel treesearch for finding shortestspan error-correcting CDO codes. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2992–3001, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06684147-
                                        abs.html.

Khoury:2011:AEM

[KBS11] Raymes Khoury, Bernd Burgstaller, and Bernhard Scholz. Accelerating the execution of matrix languages on the Cell Broadband Engine architecture. IEEE Transactions on Parallel and Distributed Systems, 22(1): 7–21, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2019:GGV

[KCH19] Changdae Kim, Seungbeom Choi, and Jaehyuk Huh. GVTS: Global virtual time fair scheduling to support strict fairness on many cores. IEEE Transactions on Parallel and Distributed Systems, 30(1):79–92, January 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08400397.
pdf.

Kim:2014:LFS

[KCK14] Seong Hoon Kim, Poh Kit Chong, and Daeyoung Kim. A location-free semi-directional-flooding technique for ondemand routing in lowrate wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3066–3075, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2014/12/06740865-
                                        abs.html.

Kim:2011:PNP

[KCW11] Jinoh Kim, Abhishek Chandra, and Jon B. Weissman. Passive network performance estimation for largescale, data-intensive computing. IEEE Transactions on Parallel and Distributed Systems, 22(8):1365–1373, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kleisouris:2010:EEW

[KCYM10] Konstantinos Kleisouris, Yingying Chen, Jie Yang, and Richard P. Martin. Empirical evaluation of wireless localization when using multiple antennas. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1595–1610, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Konstantopoulos:2019:DPC

[KDCR19] M. Konstantopoulos, P. Diamantopoulos, N. Chondros, and M. Roussopoulos. Distributed personal cloud storage without third parties. IEEE Transactions on Parallel and Distributed Systems, 30(11):2434–2448, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Klinkhamer:2016:SPS

[KE16] Alex Klinkhamer and Ali Ebnenasir. Shadow/puppet synthesis: A stepwise method for the design of selfstabilization. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3338–3350, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07422153-
abs.html.

Klingbeil:2012:FVT

[KEGM12] Guido Klingbeil, Radek Erban, Mike Giles, and Philip K. Maini. Fat versus thin threading approach on GPUs: Application to stochastic simulation of chemical reactions. IEEE Transactions on Parallel and Distributed Systems, 23(2): 280–287, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kanan:2017:DSE

[KGI17] Awos Kanan, Fayez Gebali, and Atef Ibrahim. Design space exploration of 2-D processor array architectures for similarity distance computation. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2218–2228, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07829398-
                                        abs.html.

Karakasis:2013:ECF

[KGK^{+}13] Vasileios Karakasis, Theodoros Gkountouvas, Kornilios Kourtis, Georgios Goumas, and Nectarios Koziris. An extended compression format for the optimization of sparse matrix-vector multiplication. IEEE Transactions on Parallel and Distributed Systems, 24(10): 1930–1940, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:2016:MHC

[KGR16] Rohit Kumar and Ann Gordon-Ross. MACS: A highly customizable lowlatency communication architecture. IEEE Transactions on Parallel and Distributed Systems, 27(1):237– 249, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07006979-
                                        abs.html.

Kanizo:2015:MTH

[KHK15] Yossi Kanizo, David Hay, and Isaac Keslassy. Maximizing the throughput of hash tables in network devices with combined SRAM/ DRAM memory. IEEE Transactions on Parallel and Distributed Systems, 26 (3):796–809, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06781627-
abs.html.

Kanemitsu:2016:CBT

[KHN16] Hidehiro Kanemitsu, Masaki Hanada, and Hidenori Nakazato. Clustering-based task scheduling in a large number of heterogeneous processors. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3144–3157, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07401062-
abs.html.

Kourtellis:2014:LPC

[KI14] Nicolas Kourtellis and Adriana Iamnitchi. Leveraging peer centrality in the design of socially-informed peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2364–2374, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06494568-
abs.html.

Khan:2016:HPM

[KJL^{+}16] Mukhtaj Khan, Yong Jin, Maozhen Li, Yang Xiang, and Changjun Jiang. Hadoop performance modeling for job estimation and resource provisioning. IEEE Transactions on Parallel and Distributed Systems, 27(2): 441–454, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07045505-
abs.html.

Kim:2015:EMP

[KJN15] Jinwoong Kim, Won-Ki Jeong, and Beomseok Nam. Exploiting massive parallelism for indexing multidimensional datasets on the GPU. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2258–2271, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06876171-
abs.html.

Kvalnes:2015:OKO

[KJvR^{+}15] Age Kvalnes, Dag Johansen, Robbert van Renesse, Fred B. Schneider, and Steffen Viken Valvag. Omni-Kernel: An operating system architecture for pervasive monitoring and scheduling. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2849–2862, October 2015.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06919315.
pdf.

Karenos:2010:TMS

[KK10] Kyriakos Karenos and Vana Kalogeraki. Traffic management in sensor networks with a mobile sink. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1515–1530, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kianpisheh:2017:RAP

[KKC17] Somayeh Kianpisheh, Mehdi Kargahi, and Nasrolah Moghadam Charkari. Resource availability prediction in distributed systems: An approach for modeling nonstationary transition probabilities. IEEE Transactions on Parallel and Distributed Systems, 28(8):2357–2372, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07835279-
abs.html.

Kim:2018:SRL

[KKC18] Young Geun Kim, Joonho Kong, and Sung Woo Chung. A survey on recent OS-level energy management techniques for mobile processing units. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2388–2401, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/10/08330039-
                                        abs.html.

Karn:2019:DAA

[KKE19] R. R. Karn, P. Kudva, and I. A. M. Elfadel. Dynamic autoselection and autotuning of machine learning models for cloud network analytics. IEEE Transactions on Parallel and Distributed Systems, 30(5):1052–1064, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2011:MEB

[KKK11] Hyun Jin Kim, Hong-Sik Kim, and Sungho Kang. A memory-efficient bit-split parallel string matching using pattern dividing for intrusion detection systems. IEEE Transactions on Parallel and Distributed Systems, 22(11):1904–1911, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2015:DLB

[KKK^{+}15] Deokho Kim, Minwoo Kim, Kyungah Kim, Minyong Sung, and Won Woo Ro. Dynamic load balancing of parallel SURF with vertical partitioning. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3358–3370, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06963467-
abs.html.

Khreishah:2013:LCP

[KKW13] Abdallah Khreishah, Issa Khalil, and Jie Wu. Low complexity and provably efficient algorithm for joint inter and intrasession network coding in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(10):2015–2024, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khreishah:2015:UNC

[KKW15] Abdallah Khreishah, Issa Khalil, and Jie Wu. Universal network coding-based opportunistic routing for unicast. IEEE Transactions on Parallel and Distributed Systems, 26(6):1765–1774, June 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06812230-
abs.html.

Kobus:2018:HTR

[KKW18] Tadeusz Kobus, Maciej Kokocinski, and Pawel T. Wojciechowski. Hybrid transactional replication: State-machine and deferredupdate replication combined. IEEE Transactions on Parallel and Distributed Systems, 29(7):1499–1514, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/07/08265027-
                                        abs.html.

Kim:2014:NTB

[KKY^{+}14] Taehong Kim, Seong Hoon Kim, Jinyoung Yang, Seong eun Yoo, and Daeyoung Kim. Neighbor table based shortcut tree routing in ZigBee wireless networks. IEEE Transactions on Parallel and Distributed Systems, 25(3): 706–716, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2011:DEM

[kL11a] Wei keng Liao. Design and evaluation of MPI file domain partitioning methods under extent-based file locking protocol. IEEE Transactions on Parallel and Distributed Systems, 22(2):260– 272, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:2011:CCW

[KL11b] Raju Kumar and Thomas F.

La Porta. Cooperative channelization in wireless networks with network coding. IEEE Transactions on Parallel and Distributed Systems, 22(7):1073–1084, July 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kocoloski:2016:LMM

[KL16] Brian Kocoloski and John Lange. Lightweight memory management for high performance applications in consolidated environments. IEEE Transactions on Parallel and Distributed Systems, 27(2): 468–480, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07029138-
abs.html.

Kurzak:2013:FPP

[KLFD13] Jakub Kurzak, Piotr Luszczek, Mathieu Faverge, and Jack Dongarra. LU factorization with partial pivoting for a multicore system with accelerators. IEEE Transactions on Parallel and Distributed Systems, 24(8):1613–1621, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kan:2017:MCC

[KLL^{+}17] Guangyuan Kan, Tianjie Lei, Ke Liang, Jiren Li, Liuqian Ding, Xiaoyan He, Haijun Yu, Dawei Zhang, Depeng Zuo, Zhenxin Bao, Mark Amo-Boateng, Youbing Hu, and Mengjie Zhang. A multi-core CPU and manycore GPU based fast parallel shuffled complex evolution global optimization approach. IEEE Transactions on Parallel and Distributed Systems, 28(2):332– 344, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07491261-
abs.html.

Kao:2012:NCE

[KLWK12] Yung-Cheng Kao, Chung-Nan Lee, Peng-Jung Wu, and Hui-Hsiang Kao. A network coding equivalent content distribution scheme for efficient peer-to-peer interactive VoD streaming. IEEE Transactions on Parallel and Distributed Systems, 23(6): 985–994, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kargahi:2010:UAD

[KM10] Mehdi Kargahi and Ali Movaghar. Utility accrual dynamic routing in realtime parallel systems. IEEE Transactions on Parallel and Distributed Systems, 21(12): 1822–1835, December 2010.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kansakar:2018:DSE

[KM18] Prasanna Kansakar and Arslan Munir. A design space exploration methodology for parameter optimization in multicore processors. IEEE Transactions on Parallel and Distributed Systems, 29(1): 2–15, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08017478-
abs.html.

Karniavoura:2019:DMA

[KM19] F. Karniavoura and K. Magoutis.

Decision-making approaches for performance QoS in distributed storage systems: a survey. IEEE Transactions on Parallel and Distributed Systems, 30(8):1906–1919, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khazaei:2012:PAC

[KMM12] Hamzeh Khazaei, Jelena Mišić, and Vojislav B. Mišić. Performance analysis of cloud computing centers using M/G/m/m+r queuing systems. IEEE Transactions on Parallel and Distributed Systems, 23(5): 936–943, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khazaei:2013:FGP

[KMM13a] Hamzeh Khazaei, Jelena Misic, and Vojislav B. Misic. A fine-grained performance model of cloud computing centers. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2138–2147, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khazaei:2013:PCC

[KMM13b] Hamzeh Khazaei, Jelena Misic, and Vojislav B. Misic. Performance of cloud centers with high degree of virtualization under batch task arrivals. IEEE Transactions on Parallel and Distributed Systems, 24(12):2429–2438, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khazaei:2013:APM

[KMMR13] Hamzeh Khazaei, Jelena Misić, Vojislav B. Misić, and Saeed Rashwand. Analysis of a pool management scheme for cloud computing centers. IEEE Transactions on Parallel and Distributed Systems, 24(5):849–861, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Karmouch:2012:DCS

[KN12] Eric Karmouch and Amiya Nayak. A distributed constraint satisfaction problem approach to virtual device composition. IEEE Transactions on Parallel and Distributed Systems, 23(11): 1997–2009, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khabbazian:2016:AOB

[KN16] Majid Khabbazian and Di Niu. Achieving optimal block pipelining in organized network coded gossip. IEEE Transactions on Parallel and Distributed Systems, 27(3):627–639, March 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/03/07069277-
abs.html.

Koibuchi:2011:STR

[KOKA11] Michihiro Koibuchi, Tomohiro Otsuka, Tomohiro Kudoh, and Hideharu Amano. A switch-tagged routing methodology for PC clusters with VLAN Ethernet. IEEE Transactions on Parallel and Distributed Systems, 22(2): 217–230, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kamil:2010:CRI

[KOPS10] Shoaib Kamil, Leonid Oliker, Ali Pinar, and John Shalf. Communication requirements and interconnect optimization for high-end scientific applications. IEEE Transactions on Parallel and Distributed Systems, 21(2):188– 202, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Koloniari:2012:GTA

[KP12] Georgia Koloniari and Evaggelia Pitoura. A game-theoretic approach to the formation of clustered overlay networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):589–597, April 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kohlhoff:2013:MPA

[KPA13] Kai J. Kohlhoff, Vijay S. Pande, and Russ B. Altman. K-means for parallel architectures using all-prefix-sum sorting and updating steps. IEEE Transactions on Parallel and Distributed Systems, 24(8):1602–1612, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2019:IPE

[KPB19] Kyu Yeun Kim, Jinsu Park, and Woongki Baek. Improving the performance and energy efficiency of GPGPU computing through integrated adaptive cache management. IEEE Transactions on Parallel and Distributed Systems, 30(3):630– 645, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08454288-
abs.html.

Konstantopoulos:2012:RBA

[KPG^{+}12] Charalampos Konstantopoulos, Grammati Pantziou, Damianos Gavalas, Aristides Mpitziopoulos, and Basilis Mamalis. A rendezvousbased approach enabling energy-efficient sensory data collection with mobile sinks. IEEE Transactions on Parallel and Distributed Systems, 23(5):809–817, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2016:VSC

[KPKH16] Daehoon Kim, Chang Hyun Park, Hwanju Kim, and Jaehyuk Huh. Virtual snooping coherence for multi-core virtualized systems. IEEE Transactions on Parallel and Distributed Systems, 27(7): 2155–2167, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/

                                        trans/td/2016/07/07225169-
                                        abs.html.

Kshemkalyani:2010:FME

[Ksh10] Ajay D. Kshemkalyani. Fast and message-efficient global snapshot algorithms for large-scale distributed systems. IEEE Transactions on Parallel and Distributed Systems, 21(9):1281–1289, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kwon:2010:AFL

[KSP10] Oh-Heum Kwon, Ha-Joo Song, and Sangjoon Park. Anchor-free localization through flip-error-resistant map stitching in wireless sensor network. IEEE Transactions on Parallel and Distributed Systems, 21(11):1644–1657, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Klaftenegger:2018:QDL

[KSW18] David Klaftenegger, Konstantinos Sagonas, and Kjell Winblad. Queue delegation locking. IEEE Transactions on Parallel and Distributed Systems, 29(3):687– 704, March 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8093701/.

Kurzak:2012:AGK

[KTD12] Jakub Kurzak, Stanimire Tomov, and Jack Dongarra. Autotuning GEMM kernels for the Fermi GPU. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2045–2057, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Konstantinou:2011:FCE

[KTK11] Ioannis Konstantinou, Dimitrios Tsoumakos, and Nectarios Koziris. Fast and cost-effective online loadbalancing in distributed range-queriable systems. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1350–1364, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kecskemeti:2012:VAS

[KTK12] Gabor Kecskemeti, Gabor Terstyanszky, and Peter Kacsuk. Virtual appliance size optimization with active fault injection. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1983–1995, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kumar:2014:IBI

[Kum14] Vijay Kumar. Impact of Brooks–Iyengar distributed sensing algorithm on real time systems. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1370, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Konczak:2017:OLW

[KWG17] Jan Z. Konczak, Pawel T. Wojciechowski, and Rachid Guerraoui. Operation-level wait-free transactional memory with support for irrevocable operations. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3570–3583, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/08000384-
abs.html.

Kang:2012:DNS

[KWZ^{+}12] Lei Kang, Kaishun Wu, Jin Zhang, Haoyu Tan, and Lionel M. Ni. DDC: a novel scheme to directly decode the collisions in UHF RFID systems. IEEE Transactions on Parallel and Distributed Systems, 23(2):263– 270, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kim:2011:EES

[KXC11] Kihwan Kim, Toby Xu, and Ying Cai. ELIAS: An efficient storage underlay for mobile peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 22(11):1851–1861, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kong:2014:DLR

[KXL^{+}14] Linghe Kong, Mingyuan Xia, Xiao-Yang Liu, Guangshuo Chen, Yu Gu, Min-You Wu, and Xue Liu. Data loss and reconstruction in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2818–2828, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06642025-
abs.html.

Koh:2019:EFT

[KZK^{+}19] S. Koh, J. Zhang, M. Kwon, J. Yoon, D. Donofrio, N. S. Kim, and M. Jung. Exploring fault-tolerant erasure codes for scalable allflash array clusters. IEEE Transactions on Parallel and Distributed Systems, 30 (6):1312–1330, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See errata [KZK^{+}20].

Koh:2020:EEF

[KZK^{+}20] S. Koh, J. Zhang, M. Kwon, J. Yoon, D. Donofrio, N. S.

Kim, and M. Jung. Errata to exploring fault-tolerant erasure codes for scalable allflash array clusters. IEEE Transactions on Parallel and Distributed Systems, 31(6): 1460, June 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See [KZK^{+}19].

Kong:2014:SCS

[KZLL14] Linghe Kong, Mingchen Zhao, Xiao-Yang Liu, and Jialiang Lu. Surface coverage in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(1): 234–243, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kang:2012:RRB

[KZW^{+}12] Lei Kang, Jin Zhang, Kaishun Wu, Dian Zhang, and Lionel M. Ni. RC-SMA: Receiver-based carrier sense multiple access in UHF RFID systems. IEEE Transactions on Parallel and Distributed Systems, 23(4): 735–743, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Khairy:2017:SSA

[KZW17] Mahmoud Khairy, Mohamed Zahran, and Amr Wassal. SACAT: Streaming-aware conflict-avoiding thrashingresistant GPGPU cache management scheme. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1740–1753, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07744628-
abs.html.

Lo:2012:ASE

[LA12] Chun-Hao Lo and Nirwan Ansari. Alleviating solar energy congestion in the distribution grid via smart metering communications. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1607–1620, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lombardi:2018:ESS

[LABQ18] Federico Lombardi, Leonardo Aniello, Silvia Bonomi, and Leonardo Querzoni. Elastic symbiotic scaling of operators and resources in stream processing systems. IEEE Transactions on Parallel and Distributed Systems, 29(3):572–585, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8067517/.

Lobeiras:2016:DEI

[LAD16] Jacobo Lobeiras, Margarita Amor, and Ramon Doallo.

Designing efficient indexdigit algorithms for CUDA GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1331–1343, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07138631-
abs.html.

Laguna:2015:DPF

[LAdS^{+}15] Ignacio Laguna, Dong H. Ahn, Bronis R. de Supinski, Saurabh Bagchi, and Todd Gamblin. Diagnosis of performance faults in LargeScale MPI applications via probabilistic progressdependence inference. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1280–1289, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803050-
abs.html.

Lorentz:2015:AMS

[LAFA15] Istvan Lorentz, Razvan Andonie, and Levente Fabry-Asztalos. Accelerating molecular structure determination based on inter-atomic distances using OpenCL. IEEE Transactions on Parallel and Distributed Systems, 26(12):3250–3263, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06995963-
abs.html.

Lai:2012:OCA

[Lai12] Cheng-Nan Lai. Optimal construction of all shortest node-disjoint paths in hypercubes with applications. IEEE Transactions on Parallel and Distributed Systems, 23(6):1129–1134, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2011:FSC

[LAK11] Manhee Lee, Minseon Ahn, and Eun Jung Kim. Fast secure communications in shared memory multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1714–1721, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:CAD

[LAMJ12] Yong Li, Ahmed Abousamra, Rami Melhem, and Alex K. Jones. Compiler-assisted data distribution and network configuration for chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2058–2066, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:HRB

[LAT^{+}15] Kenli Li, Wei Ai, Zhuo Tang, Fan Zhang, Lingang Jiang, Keqin Li, and Kai Hwang. Hadoop recognition of biomedical named entity using conditional random fields. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3040–3051, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06949632-
abs.html.

Li:2010:DBE

[LAV^{+}10] Yingshu Li, Chunyu Ai, Chinh T. Vu, Yi Pan, and Raheem Beyah. Delaybounded and energy-efficient composite event monitoring in heterogeneous wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1373–1385, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2010:ESO

[LC10] Jung-Shian Li and Chih-Hung Chao. An efficient superpeer overlay construction and broadcasting scheme based on perfect difference graph. IEEE Transactions on Parallel and Distributed Systems, 21(5): 594–606, May 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2011:PPC

[LC11] Alex X. Liu and Fei Chen.

Privacy preserving collaborative enforcement of firewall policies in virtual private networks. IEEE Transactions on Parallel and Distributed Systems, 22(5):887– 895, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:AAA

[LC12a] Jianzhong Li and Siyao Cheng. (e, d)-approximate aggregation algorithms in dynamic sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(3):385–396, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:AGS

[LC12b] Zongpeng Li and Xiaowen Chu. On achieving groupstrategyproof multicast. IEEE Transactions on Parallel and Distributed Systems, 23(5): 913–923, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2014:NRT

[LC14] Sian-Jheng Lin and Wei-Ho Chung. Novel repairby-transfer codes and systematic exact-MBR codes with lower complexities and smaller field sizes. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3232–3241, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714425-
abs.html.

Li:2015:CAZ

[LC15] Yusen Li and Wentong Cai. Consistency-aware zone mapping and client assignment in multi-server distributed virtual environments. IEEE Transactions on Parallel and Distributed Systems, 26(6):1570–1579, June 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06814946-
abs.html.

Lu:2013:SED

[LCA13] Tan Lu, Minghua Chen, and Lachlan L. H. Andrew. Simple and effective dynamic provisioning for power-proportional data centers. IEEE Transactions on Parallel and Distributed Systems, 24(6):1161–1171, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lulli:2017:FCC

[LCD^{+}17] Alessandro Lulli, Emanuele Carlini, Patrizio Dazzi, Claudio Lucchese, and Laura Ricci. Fast connected components computation in large graphs by vertex pruning. IEEE Transactions on Parallel and Distributed Systems, 28(3):760–773, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07515231-
abs.html.

Li:2013:IGM

[LCG^{+}13] Dan Li, Jiong Chen, Chuanxiong Guo, Yunxin Liu, Jinyu Zhang, Zhili Zhang, and Yongguang Zhang. IPgeolocation mapping for moderately connected Internet regions. IEEE Transactions on Parallel and Distributed Systems, 24(2):381– 391, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2016:DDA

[LCG^{+}16] Dan Li, Congjie Chen, Junjie Guan, Ying Zhang, Jing Zhu, and Ruozhou Yu. DCloud: Deadline-aware resource allocation for cloud computing jobs. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2248–2260, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07307195-
                                        abs.html.

Li:2014:APW

[LCGC14] Jianzhong Li, Siyao Cheng, Hong Gao, and Zhipeng Cai. Approximate physical world reconstruction algorithms in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3099–3110, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/12/06714527-
                                        abs.html.

Liu:2011:GML

[LCL^{+}11] Hai Liu, Xiaowen Chu, Yiu-Wing Leung, Xiaohua Jia, and Peng-Jun Wan. General maximal lifetime sensortarget surveillance problem and its solution. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1757–1765, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:RCS

[LCL^{+}14] Hongjuan Li, Xiuzhen Cheng, Keqiu Li, Chunqiang Hu, Nan Zhang, and Weilian Xue. Robust collaborative spectrum sensing schemes for cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 25(8):2190–2200, August 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2015:FNL

[LCL^{+}15] Shaoli Liu, Tianshi Chen, Ling Li, Xi Li, Mingzhe Zhang, Chao Wang, Haibo Meng, Xuehai Zhou, and Yunji Chen. FreeRider: Nonlocal adaptive network-on-chip routing with packetcarried propagation of congestion information. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2272–2285, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06871425-
abs.html.

Liu:2016:CAH

[LCL^{+}16a] Kebin Liu, Lei Chen, Yunhao Liu, Wei Gong, and Amiya Nayak. Continuous answering holistic queries over sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27(2): 394–404, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07051241-
abs.html.

Liu:2016:IHP

[LCL^{+}16b] Shaoli Liu, Tianshi Chen, Ling Li, Xiaoxue Feng, Zhiwei Xu, Haibo Chen, Fred Chong, and Yunji Chen. IMR: High-performance lowcost multi-ring NoCs. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1700–1712, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07182353-
abs.html.

Liu:2013:MCS

[LCLD13] Hai Liu, Xiaowen Chu, Yiu-Wing Leung, and Rui Du. Minimum-cost sensor placement for required lifetime in wireless sensor-target surveillance networks. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1783–1796, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:IEH

[LCLL15] Zhenjiangi Li, Wenwei Chen, Mo Li, and Jingsheng Lei. Incorporating energy heterogeneity into sensor network time synchronization. IEEE Transactions on Parallel and Distributed Systems, 26 (1):163–173, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06747315-
abs.html.

Li:2014:SMR

[LCS14] Ze Li, Kang Chen, and Haiying Shen. A scalable and mobility-resilient data search system for large-scale mobile wireless networks. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1124–1134, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2015:DSH

[LCS^{+}15] Xuefeng Liu, Jiannong Cao, Wen-Zhan Song, Peng Guo, and Zongjian He. Distributed sensing for highquality structural health monitoring using WSNs. IEEE Transactions on Parallel and Distributed Systems, 26(3):738–747, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06776483-
abs.html.

Lien:2012:RRM

[LCSC12] Shao-Yu Lien, Shin-Ming Cheng, Sung-Yin Shih, and Kwang-Cheng Chen. Radio resource management for QoS guarantees in cyberphysical systems. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1752–1761, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2011:ORA

[LCW11] Chun-Lung Lin, Chen-Lung Chan, and Jia-Shung Wang. Optimization of rate allocation with distortion guarantee in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1230–1237, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2017:ABP

[LCY^{+}17] Weiqing Liu, Jiannong Cao, Lei Yang, Lin Xu, Xuanjia Qiu, and Jing Li. App-Booster: Boosting the performance of interactive mobile applications with computation offloading and parameter tuning. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1593–1606, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/06/07733131-
                                        abs.html.

Liu:2016:DCO

[LCYW16] Jian Liu, Yunpeng Chai, Chang Yan, and Xin Wang. A delayed container organization approach to improve restore speed for deduplication systems. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2477–2491, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07358159-
abs.html.

Liu:2019:MRV

[LCZ^{+}19] Z. Liu, Y. Cao, X. Zhang, C. Zhu, and F. Zhang. Managing recurrent virtual network updates in multitenant datacenters: a system perspective. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1816–1825, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:RMM

[LCZZ13] Hongkun Li, Yu Cheng, Chi Zhou, and Weihua Zhuang. Routing metrics for minimizing end-to-end delay in multiradio multichannel wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2293–2303, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:RRR

[LDLL18] Cunlu Li, Dezun Dong, Zhonghai Lu, and Xiangke Liao. RoB-Router: A reorder buffer enabled low latency network-on-chip router. IEEE Transactions on Parallel and Distributed Systems, 29(9):2090–2104, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08320326-
                                        abs.html.

Liu:2013:DCM

[LDNT13] Benyuan Liu, Olivier Dousse, Philippe Nain, and Don Towsley. Dynamic coverage of mobile sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):301– 311, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lorenzon:2019:ASO

[LdSB19] A. F. Lorenzon, C. C. de Oliveira, J. D. Souza, and A. C. S. Beck. Aurora: Seamless optimization of OpenMP applications. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1007–1021, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:SEE

[LdSS^{+}13] Dong Li, Bronis R. de Supinski, Martin Schulz, Dimitrios S. Nikolopoulos, and Kirk W. Cameron. Strategies for energy-efficient resource management of hybrid programming models. IEEE Transactions on Parallel and Distributed Systems, 24(1): 144–157, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2015:DCF

[LDYZ15] Meng Liu, Wanchun Dou, Shui Yu, and Zhensheng Zhang. A decentralized cloud firewall framework with resources provisioning cost optimization. IEEE Transactions on Parallel and Distributed Systems, 26(3):621– 631, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06781636-
abs.html.

Lee:2012:EES

[Lee12] Wan Yeon Lee. Energyefficient scheduling of periodic real-time tasks on lightly loaded multicore processors. IEEE Transactions on Parallel and Distributed Systems, 23(3):530– 537, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2017:TRR

[Lee17] Jinkyu Lee. Time-reversibility for real-time scheduling on multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 28(1):230–243, January 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/01/07416238-
                                        abs.html.

Liang:2010:IPA

[LFLW10] Chao Liang, Zhenghua Fu, Yong Liu, and Chai Wah Wu. Incentivized peerassisted streaming for ondemand services. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1354–1367, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2010:CRR

[LG10] Alex X. Liu and Mohamed G. Gouda. Complete redundancy removal for packet classifiers in TCAMs. IEEE Transactions on Parallel and Distributed Systems, 21(4):424–437, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:MCE

[LG13] Peng Li and Song Guo. On the multicast capacity in energy-constrained lossy wireless networks by exploiting intrabatch and interbatch network coding. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2251–2260, November 2013.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Le:2014:OIR

[LGD14] Tho Ngoc Le, Mohan Gurusamy, and Dinil Mon Divakaran. An online integrated resource allocator for guaranteed performance in data centers. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1382–1392, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2014:GCE

[LGG^{+}14] Changqing Luo, Shengyong Guo, Song Guo, Laurence T. Yang, Geyong Min, and Xia Xie. Green communication in energy renewable wireless mesh networks: Routing, rate control, and power allocation. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3211–3220, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714597-
abs.html.

Liao:2017:EMS

[LGJ^{+}17] Xiaofei Liao, Rentong Guo, Hai Jin, Jianhui Yue, and Guang Tan. Enhancing the malloc system with pollution awareness for better cache performance. IEEE Transactions on Parallel and Distributed Systems, 28(3): 731–745, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07506061-
abs.html.

Li:2018:HVM

[LGJ^{+}18] Xiang Li, Peter Garraghan, Xiaohong Jiang, Zhaohui Wu, and Jie Xu. Holistic virtual machine scheduling in cloud datacenters towards minimizing total energy. IEEE Transactions on Parallel and Distributed Systems, 29(6):1317–1331, June 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/06/07888576-
abs.html.

Lama:2016:APP

[LGJZ16] Palden Lama, Yanfei Guo, Changjun Jiang, and Xiaobo Zhou. Autonomic performance and power control for co-located Web applications in virtualized datacenters. IEEE Transactions on Parallel and Distributed Systems, 27(5):1289–1302, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07152934-
abs.html.

Li:2018:IID

[LGL^{+}18a] Wenxin Li, Deke Guo, Keqiu Li, Heng Qi, and Jianhui Zhang. iDaaS: Interdatacenter network as a service. IEEE Transactions on Parallel and Distributed Systems, 29(7):1515–1529, July 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/07347416-
abs.html.

Li:2018:CMB

[LGL^{+}18b] Wenxin Li, Deke Guo, Alex X. Liu, Keqiu Li, Heng Qi, Song Guo, Ali Munir, and Xiaoyi Tao. Co-Man: Managing bandwidth across computing frameworks in multiplexed datacenters. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1013–1029, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08241720-
abs.html.

Li:2017:TAG

[LGM^{+}17] Peng Li, Song Guo, Toshiaki Miyazaki, Xiaofei Liao, Hai Jin, Albert Y. Zomaya, and Kun Wang. Trafficaware geo-distributed big data analytics with predictable job completion time.

IEEE Transactions on Parallel and Distributed Systems, 28(6):1785–1796, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07738559-
abs.html.

Laredo:2017:LBE

[LGOB17] Juan Luis Jimenez Laredo, Frederic Guinand, Damien Olivier, and Pascal Bouvry. Load balancing at the edge of chaos: How self-organized criticality can lead to energyefficient computing. IEEE Transactions on Parallel and Distributed Systems, 28(2): 517–529, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07494682-
abs.html.

Luo:2017:VVE

[LGW^{+}17] Lailong Luo, Deke Guo, Jie Wu, Ting Qu, Tao Chen, and Xueshan Luo. VLCcube: A VLC enabled hybrid network structure for data centers. IEEE Transactions on Parallel and Distributed Systems, 28(7):2088–2102, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07801878-
abs.html.

Liu:2011:SMG

[LGX^{+}11] Weichen Liu, Zonghua Gu, Jiang Xu, Xiaowen Wu, and Yaoyao Ye. Satisfiability modulo graph theory for task mapping and scheduling on multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 22(8):1382–1389, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2019:GFE

[LGXL19] L. Luo, D. Guo, J. Xu, and X. Luo. Graph filter: Enabling efficient topology calibration. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2730–2742, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:RMP

[LGYV14] Peng Li, Song Guo, Shui Yu, and Athanasios V. Vasilakos. Reliable multicast with pipelined network coding using opportunistic feeding and routing. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3264–3273, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714456-
abs.html.

Liang:2019:UBO

[LGZ^{+}19] Y. Liang, J. Ge, S. Zhang, J. Wu, Z. Tang, and B. Luo. A utility-based optimization framework for edge service entity caching. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2384–2395, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2015:VCL

[LH15] Haikun Liu and Bingsheng He. VMbuddies: Coordinating live migration of multi-tier applications in cloud environments. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1192–1205, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06784491-
                                        abs.html.

Liu:2016:FEF

[LH16] Haikun Liu and Bingsheng He. F2C: Enabling fair and fine-grained resource sharing in multi-tenant IaaS clouds. IEEE Transactions on Parallel and Distributed Systems, 27(9):2589–2602, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/

trans/td/2016/09/07327230-
abs.html.

Liu:2017:REP

[LH17] Yang Liu and Shiyan Hu.

Renewable energy pricing driven scheduling in distributed smart community systems. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1445–1456, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07586083-
abs.html.

Lu:2017:ABG

[LHCM^{+}17] Hao Lu, Mahantesh Halappanavar, Daniel Chavarria-Miranda, Assefaw H. Gebremedhin, Ajay Panyala, and Ananth Kalyanaraman. Algorithms for balanced graph colorings with applications in parallel computing. IEEE Transactions on Parallel and Distributed Systems, 28(5):1240–1256, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07605439-
abs.html.

Li:2014:QTC

[LHD^{+}14] Mo Li, Yuan He, Wei Dong, Yunhao Liu, and Jiliang Wang. QoF: Towards comprehensive path quality measurement in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(4): 1003–1013, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:CPT

[LHF^{+}15] Wenzhong Li, Yuefei Hu, Xiaoming Fu, Sanglu Lu, and Daoxu Chen. Cooperative positioning and tracking in disruption tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 26(2):382– 391, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06763003-
                                        abs.html.

Li:2017:OSD

[LHG^{+}17] Chao Li, Yang Hu, Juncheng Gu, Jingling Yuan, and Tao Li. Oasis: Scaling out datacenter sustainably and economically. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1960–1973, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/07/07585018-
                                        abs.html.

Langer:2018:MSM

[LHHR18] Matthias Langer, Ashley Hall, Zhen He, and Wenny Rahayu. MPCA SGD — a method for distributed training of deep learning models on Spark. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2540–2556, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08354695-
abs.html.

Lin:2012:ECP

[LHJ12] Tsong-Jie Lin, Sun-Yuan Hsieh, and Justie Su-Tzu Juan. Embedding cycles and paths in product networks and their applications to multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 23(6):1081–1089, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:CDP

[LHL^{+}13a] Zhenhua Li, Yan Huang, Gang Liu, Fuchen Wang, Yunhao Liu, Zhi-Li Zhang, and Yafei Dai. Challenges, designs, and performances of large-scale Open-P2SP content distribution. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2181–2191, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2013:DWS

[LHL^{+}13b] Yunhao Liu, Yuan He, Mo Li, Jiliang Wang, Kebin Liu, and Xiangyang Li. Does wireless sensor network scale? A measurement study on GreenOrbs. IEEE Transactions on Parallel and Distributed Systems, 24(10): 1983–1993, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:SOA

[LHL^{+}14] Jin Li, Xinyi Huang, Jingwei Li, Xiaofeng Chen, and Yang Xiang. Securely outsourcing attribute-based encryption with checkability. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2201–2210, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2017:EER

[LHL17] Runhui Li, Yuchong Hu, and Patrick P. C. Lee. Enabling efficient and reliable transition from replication to erasure coding for clustered file systems. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2500–2513, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07872497-
                                        abs.html.

Lin:2012:DPB

[LHM12] Dong Lin, Mounir Hamdi, and Jogesh K. Muppala. Distributed packet buffers for high-bandwidth switches and routers. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1178–1192, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2015:EGS

[LHR^{+}15] Yun Liang, Huynh Phung Huynh, Kyle Rupnow, Rick Siow Mong Goh, and Deming Chen. Efficient GPU spatialtemporal multitasking. IEEE Transactions on Parallel and Distributed Systems, 26 (3):748–760, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06777559-
abs.html.

Liu:2011:RRA

[LHW11] Yunhao Liu, Jinsong Han, and Jilong Wang. Rumor riding: Anonymizing unstructured peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 22(3):464–475, March 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:VNF

[LHXP18] Defang Li, Peilin Hong, Kaiping Xue, and Jianing Pei. Virtual network function placement considering resource optimization and SFC requests in cloud datacenter. IEEE Transactions on Parallel and Distributed Systems, 29(7):1664–1677, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08281644-
abs.html.

Liu:2015:EDQ

[LHYW15] Yang Liu, Yanyan Han, Zhipeng Yang, and Hongyi Wu. Efficient data query in intermittently-connected mobile ad hoc social networks. IEEE Transactions on Parallel and Distributed Systems, 26(5):1301–1312, May 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06807752-
abs.html.

Liao:2019:ETE

[LHZJ19] X. Liao, Y. Huang, L. Zheng, and H. Jin. Efficient timeevolving stream processing at scale. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2165–2178, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2010:RWM

[Li10] Hui Li. Realistic workload modeling and its performance impacts in largescale eScience grids. IEEE Transactions on Parallel and Distributed Systems, 21(4): 480–493, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:ADB

[Li13]  Keqin Li.  Analysis of

distance-based location management in wireless communication networks. IEEE Transactions on Parallel and Distributed Systems, 24(2): 225–238, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:LSD

[Li14a] Keqin Li. Liquid: A scalable deduplication file system for virtual machine images. IEEE Transactions on Parallel and Distributed Systems, 25(5):1257–1266, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:UMT

[Li14b]  Xiang-Yang Li.  Under-

standing multi-task schedulabilityin duty-cycling sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(9):

2464–2475, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06471978-
abs.html.

Li:2014:TOL

[Li14c] Zhongcheng Li. Throughput optimizing localized link scheduling for multihop wireless networks under physical interference model. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2708–2720, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06583158-
abs.html.

Liu:2014:TAO

[Liu14] Yunhao Liu. Towards accurate object localization with Smartphones. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2731–2742, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06601603-
abs.html.

Ling:2015:SLA

[LIWJ15] Xiao Ling, Shadi Ibrahim, Song Wu, and Hai Jin. Spatial locality aware disk scheduling in virtualized environment. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2571–2585, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06892999.
pdf.

Li:2015:DSH

[LJ15] Ding Li and Sudharman K.

Jayaweera. Distributed Smart-home decision-making in a hierarchical interactive Smart Grid architecture. IEEE Transactions on Parallel and Distributed Systems, 26(1):75–84, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748053-
abs.html.

LeGal:2016:HTM

[LJ16] Bertrand Le Gal and Christophe Jego. High-throughput multi-core LDPC decoders based on x86 processor. IEEE Transactions on Parallel and Distributed Systems, 27(5):1373–1386, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07110620-
abs.html.

Liu:2013:DTB

[LJB^{+}13] Wenping Liu, Hongbo Jiang, Xiang Bai, Guang Tan, Chonggang Wang, Wenyu Liu, and Kechao Cai. Distance transform-based skeleton extraction and its applications in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(9):1763–1772, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:FMU

[LJG12] Mo Li, Xiaoye (Jonathan) Jiang, and Leonidas J. Guibas. Fingerprinting mobile user positions in sensor networks: Attacks and countermeasures. IEEE Transactions on Parallel and Distributed Systems, 23(4): 676–683, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2011:LVM

[LJL^{+}11] Haikun Liu, Hai Jin, Xiaofei Liao, Chen Yu, and Cheng-Zhong Xu. Live virtual machine migration via asynchronous replication and state synchronization. IEEE Transactions on Parallel and Distributed Systems, 22(12): 1986–1999, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2015:HBC

[LJL^{+}15] Haikun Liu, Hai Jin, Xiaofei Liao, Wei Deng, Bingsheng He, and Cheng zhong Xu. Hotplug or ballooning: A comparative study on dynamic memory management techniques for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1350–1363, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06807799-
abs.html.

Li:2011:LCM

[LKBK11] Sheng Li, Shannon Kuntz, Jay B. Brockman, and Peter M. Kogge. Lightweight Chip Multi-Threading (LCMT): Maximizing fine-grained parallelism on-chip. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1178–1191, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ltaief:2010:PTS

[LKD10] Hatem Ltaief, Jakub Kurzak, and Jack Dongarra. Parallel two-sided matrix reduction to band bidiagonal form on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 21(4):417–423, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2016:TDO

[LKE16] Chul-Ho Lee, Jaewook Kwak, and Do Young Eun. Towards distributed optimal movement strategy for data gathering in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27(2):574– 584, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07051231-
                                        abs.html.

Levitin:2010:MST

[LKM10] Lev Levitin, Mark Karpovsky, and Mehmet Mustafa. Minimal sets of turns for breaking cycles in graphs modeling networks. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1342–1353, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2011:CFD

[LKT11] Cheng-Kuan Lin, Tzu-Liang Kung, and Jimmy J. M. Tan. Conditional-fault diagnosability of multiprocessor systems with an efficient local diagnosis algorithm under the PMC model. IEEE Transactions on Parallel and Distributed Systems, 22(10):1669–1680, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2011:DAD

[LL11] Jing Liang and Qilian Liang.

Design and analysis of distributed radar sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(11):1926–1933, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:PAC

[LL12] Jing Li and Di Liu. kpancyclicity of k-ary n-cube networks under the conditional fault model. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1115–1120, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2017:BEC

[LL17] Jun Li and Baochun Li.

Beehive: Erasure codes for fixing multiple failures in distributed storage systems. IEEE Transactions on Parallel and Distributed Systems, 28(5):1257–1270, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07726018-
abs.html.

Li:2018:MNM

[LLAL18] Hao Li, Kenli Li, Jiyao An, and Keqin Li. MSGD: A novel matrix factorization approach for large-scale collaborative filtering recommender systems on GPUs. IEEE Transactions on Parallel and Distributed Systems, 29(7):1530–1544, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/07955082-
abs.html.

Lin:2010:IFF

[LLC10] Minghong Lin, John C. S. Lui, and Dah-Ming Chiu. An ISP-friendly file distribution protocol: Analysis, design, and implementation. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1317–1329, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:HCA

[LLC^{+}15] Jin Li, Yan Kit Li, Xiaofeng Chen, Patrick P. C. Lee, and Wenjing Lou. A hybrid cloud approach for secure authorized deduplication. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1206–1216, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06802424-
abs.html.

Lee:2012:QIP

[LLCH12] Gwo Giun (Chris) Lee, He-Yuan Lin, Chun-Fu Chen, and Tsung-Yuan Huang. Quantifying intrinsic parallelism using linear algebra for algorithm/architecture coexploration. IEEE Transactions on Parallel and Distributed Systems, 23(5):944– 957, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2012:JSR

[LLCL12] Hai Liu, Zhiyong Lin, Xiaowen Chu, and Yiu-Wing Leung. Jump-stay rendezvous algorithm for cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1867–1881, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2018:TIT

[LLD^{+}18] Yanan Lu, Leibo Liu, Yangdong Deng, Jian Weng, Shouyi Yin, Yiyu Shi, and Shaojun Wei. Triggeredissuance and triggered-execution: A control paradigm to minimize pipeline stalls in distributed controlled coarsegrained reconfigurable arrays. IEEE Transactions on Parallel and Distributed Systems, 29(10):2360–2372, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/10/08330047-
                                        abs.html.

Liu:2015:OTP

[LLFL15] Gi-Ren Liu, Phone Lin, Yuguang Fang, and Yi-Bing Lin. Optimal threshold policy for in-home smart grid with renewable generation integration. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1096–1105, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06797861-
                                        abs.html.

Liu:2013:MAI

[LLG^{+}13] Yao Liu, Fei Li, Lei Guo, Bo Shen, Songqing Chen, and Yingjie Lan. Measurement and analysis of an Internet streaming service to mobile devices. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2240–2250, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2014:SED

[LLG14] Huang Lu, Jie Li, and Mohsen Guizani. Secure and efficient data transmission for cluster-based wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(3): 750–761, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:NAD

[LLG15a] Jianzhong Li, Guohua Li, and Hong Gao. Novel ε-approximation to data streams in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(6):1654–1667, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06814284-
abs.html.

Li:2015:ANA

[LLG15b] Jie Li, Huang Lu, and Mohsen Guizani. ACPN: A novel authentication framework with conditional privacypreservation and non-repudiation for VANETs. IEEE Transactions on Parallel and Distributed Systems, 26(4):938– 948, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06748095-
abs.html.

Lacuesta:2013:SPS

[LLGP13] R. Lacuesta, Jaime Lloret, M. Garcia, and L. Penalver. A secure protocol for spontaneous wireless ad hoc networks creation. IEEE Transactions on Parallel and Distributed Systems, 24(4):629– 641, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2014:HPN

[LLH14] Chia-Wei Lee, Tsong-Jie Lin, and Sun-Yuan Hsieh. Hamiltonicity of product networks with faulty elements. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2318–2331, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/09/06494569-
                                        abs.html.

Liu:2015:DAC

[LLH^{+}15a] Feng Liu, Mu Lin, Yusuo Hu, Chong Luo, and Feng Wu. Design and analysis of compressive data persistence in large-scale wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2685–2698, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06913556.
                                        pdf.

Lu:2015:MOM

[LLH^{+}15b] Mian Lu, Yun Liang, Huynh Phung Huynh, Zhongliang Ong, Bingsheng He, and Rick Siow Mong Goh. MrPhi: An optimized MapReduce framework on Intel Xeon Phi coprocessors. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3066–3078, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06939728-
abs.html.

Li:2013:NDM

[LLJ^{+}13] Dingding Li, Xiaofei Liao, Hai Jin, Bingbing Zhou, and Qi Zhang. A new disk I/O model of virtualized cloud environment. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1129–1138, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2013:ANC

[LLK13] Hui-Tang Lin, Ying-You Lin, and Hung-Jung Kang. Adaptive network coding for broadband wireless access networks. IEEE Transactions on Parallel and Distributed Systems, 24(1):4– 18, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lo:2014:IGM

[LLK^{+}14] Shih-Hsiang Lo, Che-Rung Lee, Quey-Liang Kao, I-Hsin Chung, and Yeh-Ching Chung. Improving GPU memory performance with artificial barrier synchronization. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2342–2352, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06515115-
abs.html.

Lu:2012:EEP

[LLL^{+}12] Rongxing Lu, Xiaohui Liang, Xu Li, Xiaodong Lin, and Xuemin (Sherman) Shen. EPPA: An efficient and privacy-preserving aggregation scheme for secure Smart Grid communications. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1621–1631, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:EUD

[LLL^{+}13] Zhenjiang Li, Yunhao Liu, Mo Li, Jiliang Wang, and Zhichao Cao. Exploiting ubiquitous data collection for mobile users in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):312– 326, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2014:IGA

[LLL^{+}14a] Xiaodong Liu, Mo Li, Shanshan Li, Shaoliang Peng, Xiangke Liao, and Xiaopei Lu. IMGPU: GPUaccelerated influence maximization in large-scale social networks. IEEE Transactions on Parallel and Distributed Systems, 25(1):136– 145, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lou:2014:SDE

[LLL^{+}14b] Wenjing Lou, Patrick P. C.

Lee, Jingwei Li, Mingqiang Li, Xiaofeng Chen, and Jin Li. Secure deduplication with efficient and reliable convergent key management. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1615–1625, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2018:MCS

[LLL18] Chubo Liu, Kenli Li, and Keqin Li. Minimal cost server configuration for meeting time-varying resource demands in cloud centers. IEEE Transactions on Parallel and Distributed Systems, 29(11):2503–2513, November 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2018/11/08359091-
                                        abs.html.

Lin:2017:PHB

[LLLC17] Cheng-Hung Lin, Jin-Cheng Li, Chen-Hsiung Liu, and Shih-Chieh Chang. Perfect hashing based parallel algorithms for multiple string matching on graphic processing units. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2639–2650, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07864442-
                                        abs.html.

Lee:2013:UPO

[LLLG13] Uichin Lee, Seung-Hoon Lee, Kang-Won Lee, and Mario Gerla. Understanding processing overheads of network coding-based content distribution in VANETs. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2304–2318, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2019:FID

[LLLH19] Tingwei Liu, John C. S. Lui, Dong Lin, and David Hui. On the feasibility of interdomain routing via a small broker set. IEEE Transactions on Parallel and Distributed Systems, 30(2):415– 427, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08437158-
abs.html.

Li:2016:FPB

[LLLZ16] Kenli Li, Chubo Liu, Keqin Li, and Albert Y. Zomaya. A framework of price bidding configurations for resource usage in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 27(8):2168–2181, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07307193-
abs.html.

Liu:2014:EUT

[LLM^{+}14] Xiulong Liu, Keqiu Li, Geyong Min, Kai Lin, Bin Xiao, Yanming Shen, and Wenyu Qu. Efficient unknown tag identification protocols in large-scale RFID systems. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3145–3155, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714543-
abs.html.

Lettich:2019:PTL

[LLN^{+}19] F. Lettich, C. Lucchese, F. M. Nardini, S. Orlando, R. Perego, N. Tonellotto, and R. Venturini. Parallel traversal of large ensembles of decision trees. IEEE Transactions on Parallel and Distributed Systems, 30(9): 2075–2089, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2015:IMP

[LLpC15] Jia-Chun Lin, Fang-Yie Leu, and Ying ping Chen. Impact of MapReduce policies on job completion reliability and job energy consumption. IEEE Transactions on Parallel and Distributed Systems, 26(5):1364–1378, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06966761-
abs.html.

Li:2018:JSS

[LLRP18] Shijing Li, Tian Lan, Moo-Ryong Ra, and Rajesh Panta. Joint scheduling and source selection for background traffic in erasurecoded storage. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2826–2837, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08378267-
abs.html.

Lu:2013:SSP

[LLS13] Rongxing Lu, Xiaodong Lin, and Xuemin Shen. SPOC: A secure and privacypreserving opportunistic computing framework for mobilehealthcare emergency. IEEE Transactions on Parallel and Distributed Systems, 24(3): 614–624, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2014:ETS

[LLS14] Xiaohui Liang, Xiaodong Lin, and Xuemin Sherman Shen. Enabling trustworthy service evaluation in service-oriented mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 25(2):310– 320, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:FMS

[LLS^{+}18] Siyang Li, Fenlin Liu, Jiwu Shu, Youyou Lu, Tao Li, and Yang Hu. A flattened metadata service for distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2641–2657, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08370078-
abs.html.

Li:2015:PAG

[LLW^{+}15] Liang Li, Hong Liu, Hao Wang, Taoying Liu, and Wei Li. A parallel algorithm for game tree search using GPGPU. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2114–2127, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/08/06868996-
                                        abs.html.

Liu:2012:EJC

[LLXC12] Zhenhua Liu, Hongbo Liu, Wenyuan Xu, and Yingying Chen. Exploiting jammingcaused neighbor changes for jammer localization. IEEE Transactions on Parallel and Distributed Systems, 23(3): 547–555, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2014:EMF

[LLXC14] Zhenhua Liu, Hongbo Liu, Wenyuan Xu, and Yingying Chen. An errorminimizing framework for localizing jammers in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 25(2):508– 517, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:EEP

[LLY^{+}14] Hongwei Li, Xiaodong Lin, Haomiao Yang, Xiaohui Liang, Rongxing Lu, and Xuemin Shen. EPPDR: An efficient privacy-preserving demand response scheme with adaptive key evolution in smart grid. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2053–2064, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ling:2015:TBD

[LLY^{+}15] Zhen Ling, Junzhou Luo, Wei Yu, Ming Yang, and Xinwen Fu. Tor bridge discovery: Extensive analysis and large-scale empirical evaluation. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1887–1899, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06616552-
abs.html.

Lee:2016:HJD

[LLY16] Ming-Chang Lee, Jia-Chun Lin, and Ramin Yahyapour. Hybrid job-driven scheduling for virtual MapReduce clusters. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1687–1699, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/06/07175043-
                                        abs.html.

Li:2017:TGT

[LLY^{+}17] Gang Li, Xinfeng Li, Fan Yang, Jin Teng, Sihao Ding, Yuan F. Zheng, Dong Xuan, Biao Chen, and Wei Zhao. Traffic at-a-glance: Timebounded analytics on large visual traffic data. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2703–2717, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07880553-
                                        abs.html.

Liu:2012:FCP

[LLZ^{+}12a] Fangming Liu, Bo Li, Lili Zhong, Baochun Li, Hai Jin, and Xiaofei Liao. Flash crowd in P2P live streaming systems: Fundamental characteristics and design implications. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1227–1239, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2012:BBE

[LLZ^{+}12b] Rongxing Lu, Xiaodong Lin, Haojin Zhu, Xiaohui Liang, and Xuemin (Sherman) Shen. BECAN: a Bandwidth-Efficient Cooperative Authentication Scheme for filtering injected false data in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(1):32– 43, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2014:CNA

[LLZ14] Linfeng Liu, Ye Liu, and Ningshen Zhang. A complex network approach to topology control problem in underwater acoustic sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3046–3055, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714535-
abs.html.

Liao:2018:SDR

[LLZ^{+}18a] Xiaofei Liao, Minhao Lin, Long Zheng, Hai Jin, and Zhiyuan Shao. Scalable data race detection for lockintensive programs with pending period representation. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2599–2612, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2018/11/08359198-
                                        abs.html.

Liu:2018:SPM

[LLZ18b] Xi Liu, Weidong Li, and Xuejie Zhang. Strategyproof mechanism for provisioning and allocation virtual machines in heterogeneous clouds. IEEE Transactions on Parallel and Distributed Systems, 29(7): 1650–1663, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/07/08241849-
                                        abs.html.

Liu:2012:CWS

[LM12] Liang Liu and Huadong Ma. On coverage of wireless sensor networks for rolling terrains. IEEE Transactions on Parallel and Distributed Systems, 23(1):118– 125, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lopes:2016:TJS

[LM16] Raquel V. Lopes and Daniel Menasce. A taxonomy of job scheduling on distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3412–3428, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/

trans/td/2016/12/07425222-
abs.html.

Lastovetsky:2017:NMB

[LM17] Alexey Lastovetsky and Ravi Reddy Manumachu. New model-based methods and algorithms for performance and energy optimization of data parallel applications on homogeneous multicore clusters. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1119–1133, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07565748-
abs.html.

Lim:2017:MRT

[LMAS17] Norman Lim, Shikharesh Majumdar, and Peter Ashwood-Smith. MRCP-RM: A technique for resource allocation and scheduling of MapReduce jobs with deadlines. IEEE Transactions on Parallel and Distributed Systems, 28(5):1375–1389, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07590072-
abs.html.

Leonard:2016:THC

[LMD16] Lorne Leonard, Kamesh Madduri, and Christopher J. Duffy. Tuning heterogeneous computing platforms for large-scale hydrology data management. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2753–2765, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/09/07327220-
                                        abs.html.

Lin:2011:CCM

[LMFS11] Heshan Lin, Xiaosong Ma, Wuchun Feng, and Nagiza F. Samatova. Coordinating computation and I/O in massively parallel sequence search. IEEE Transactions on Parallel and Distributed Systems, 22(4):529– 543, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:ECC

[LMLM13] Keqiu Li, Yuanping Mu, Keqin Li, and Geyong Min. Exchanged crossed cube: A novel interconnection network for parallel computation. IEEE Transactions on Parallel and Distributed Systems, 24(11):2211–2219, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lago:2018:EAV

[LMM18] Daniel Guimaraes Lago, Edmundo R. M. Madeira, and Deep Medhi. Energy-aware virtual machine scheduling on data centers with heterogeneous bandwidths. IEEE Transactions on Parallel and Distributed Systems, 29(1): 83–98, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08039210-
abs.html.

Leitao:2012:XBP

[LMPR12] Joao Leitao, Joao Pedro Marques, Jose Pereira, and Luis Rodrigues. X-BOT: a protocol for resilient optimization of unstructured overlay networks. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2175–2188, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2010:TON

[LMR10] Patrick P. C. Lee, Vishal Misra, and Dan Rubenstein. Toward optimal network fault correction in externally managed overlay networks. IEEE Transactions on Parallel and Distributed Systems, 21(3):354–366, March 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:DBM

[LMSRSR12] Xu Li, Nathalie Mitton, Isabelle Simplot-Ryl, and David Simplot-Ryl. Dynamic beacon mobility scheduling for sensor localization. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1439–1452, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:HBD

[LMSRSR13] Xu Li, Nathalie Mitton, Isabelle Simplot-Ryl, and David Simplot-Ryl. Hypocomb: Bounded-degree localized geometric planar graphs for wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 24(7):1341–1354, July 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Linford:2011:AGM

[LMVS11] John C. Linford, John Michalakes, Manish Vachharajani, and Adrian Sandu. Automatic generation of multicore chemical kernels. IEEE Transactions on Parallel and Distributed Systems, 22(1):119–131, January 2011. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:SOA

[LMZG15] Xiaoyong Li, Huadong Ma, Feng Zhou, and Xiaolin Gui. Service operator-aware trust scheme for resource matchmaking across multiple clouds. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1419–1429, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06810192-
abs.html.

Liu:2017:LBN

[LN17] Zhaolei Liu and T. S. Eugene Ng. Leaky buffer: A novel abstraction for relieving memory pressure from cluster data processing frameworks. IEEE Transactions on Parallel and Distributed Systems, 28(1):128– 140, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07442158-
abs.html.

Liu:2013:CBC

[LNA^{+}13] Wei Liu, Hiroki Nishiyama, Nirwan Ansari, Jie Yang, and Nei Kato. Clusterbased certificate revocation with vindication capability for mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):239– 249, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2017:CEB

[LNK17] Yan Liu, Di Niu, and Majid Khabbazian. Cooper: Expedite batch data dissemination in computer clusters with coded gossips. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2204–2217, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07820201-
                                        abs.html.

Li:2019:WAS

[LNL^{+}19] W. Li, D. Niu, Y. Liu, S. Liu, and B. Li. Wide-area Spark streaming: Automated routing and batch sizing. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1434–1448, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lopez-Novoa:2015:SPM

[LNMMA15] Unai Lopez-Novoa, Alexander Mendiburu, and Jose Miguel-Alonso. A survey of performance modeling and simulation techniques for accelerator-based computing. IEEE Transactions on Parallel and Distributed Systems, 26(1):272– 281, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/

trans/td/2015/01/06748067-
abs.html.

Liu:2015:SAB

[LNXY15] Hong Liu, Huansheng Ning, Qingxu Xiong, and Laurence T. Yang. Shared authority based privacypreserving authentication protocol in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 26(1):241– 251, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748054-
abs.html.

Liu:2013:GPB

[LNZ^{+}13] Hong Liu, Huansheng Ning, Yan Zhang, Daojing He, Qingxu Xiong, and Laurence T. Yang. Groupingproofs-based authentication protocol for distributed RFID systems. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1321–1330, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2017:EDD

[LODB17] Liangzhi Li, Kaoru Ota, Mianxiong Dong, and Wuyunzhaola Borjigin. Eyes in the dark: Distributed scene understanding for disaster management. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3458–3471, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/08010892-
                                        abs.html.

Lou:2014:NAS

[Lou14]  Wenjing Lou.  New al-

gorithms for secure outsourcing of modular exponentiations. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2386–2396, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/09/06567860-
                                        abs.html.

Liu:2013:PPT

[LPMB13] Xing Liu, Pushkar R. Pande, Henning Meyerhenke, and David A. Bader. PASQUAL: Parallel techniques for next generation genome sequence assembly. IEEE Transactions on Parallel and Distributed Systems, 24(5):977– 986, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Leivadeas:2013:ERM

[LPP13] Aris Leivadeas, Chrysa Papagianni, and Symeon Papavassiliou. Efficient resource mapping framework over networked clouds via iterated local search-based request partitioning. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1077–1086, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lebre:2019:PNV

[LPSS19] Adrien Lebre, Jonathan Pastor, Anthony Simonet, and Mario Sudholt. Putting the next 500 VM placement algorithms to the acid test: The infrastructure provider viewpoint. IEEE Transactions on Parallel and Distributed Systems, 30(1):204– 217, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08409978-
abs.html.

Lu:2012:DCB

[LPZ12] Shan Lu, Soyeon Park, and Yuanyuan Zhou. Detecting concurrency bugs from the perspectives of synchronization intentions. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1060–1072, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:CEM

[LQK^{+}13] Yantao Li, Xin Qi, Matthew Keally, Zhen Ren, Gang Zhou, Di Xiao, and Shaojiang Deng. Communication energy modeling and optimization through joint packet size analysis of BSN and WiFi networks. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1741–1751, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:WSM

[LQW^{+}18] Daping Li, Xiaoyang Qu, Jiguang Wan, Jun Wang, Yang Xia, Xiaozhao Zhuang, and Changsheng Xie. Workload scheduling for massive storage systems with arbitrary renewable supply. IEEE Transactions on Parallel and Distributed Systems, 29(10):2373–2387, October 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/10/08327505-
abs.html.

Li:2012:TES

[LQY^{+}12] Xu Li, Chunming Qiao, Xuegang Yu, Aditya Wagh, Raghu Sudhaakar, and Sateesh Addepalli. Toward effective service scheduling for human drivers in vehicular cyber-physical systems. IEEE Transactions on Parallel and Distributed Systems, 23(9):1775–1789, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2013:MQM

[LRJX13] Weifa Liang, Xiaojiang Ren, Xiaohua Jia, and Xu Xu. Monitoring quality maximization through fair rate allocation in harvesting sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1827–1840, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lakhlef:2019:VCC

[LRT19] H. Lakhlef, M. Raynal, and F. Taïani. Vertex coloring with communication constraints in synchronous broadcast networks. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1672–1686, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:QMD

[LRW12] Yun Li, Jian Ren, and Jie Wu. Quantitative measurement and design of sourcelocation privacy schemes for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(7):1302–1311, July 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2017:RRA

[LRYJ17] Tan Li, Yufei Ren, Dantong Yu, and Shudong Jin. RAMSYS: Resource-aware asynchronous data transfer with multicore SYStems. IEEE Transactions on Parallel and Distributed Systems, 28(5):1430–1444, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07600392-
abs.html.

Liu:2014:ETR

[LS14] Guoxin Liu and Haiying Shen. An efficient and trustworthy resource sharing platform for collaborative cloud computing. IEEE Transactions on Parallel and Distributed Systems, 25(4): 862–875, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2017:MSS

[LS17a] Zhuozhao Li and Haiying Shen. Measuring scale-up and scale-out Hadoop with remote and local file systems and selecting the best platform. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3201–3214, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

trans/td/2017/11/07940040-
abs.html.

Lin:2017:CLF

[LS17b] Yuhua Lin and Haiying Shen.

CloudFog: Leveraging fog to extend cloud gaming for thin-client MMOG with high quality of service. IEEE Transactions on Parallel and Distributed Systems, 28(2): 431–445, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07465785-
abs.html.

Lin:2017:EEE

[LS17c] Yuhua Lin and Haiying Shen.

EAFR: An energy-efficient adaptive file replication system in data-intensive clusters. IEEE Transactions on Parallel and Distributed Systems, 28(4):1017–1030, April 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07577866-
abs.html.

Liu:2017:IDQ

[LS17d] Guoxin Liu and Haiying Shen. iASK: A distributed Q&A system incorporating social community and global collective intelligence. IEEE Transactions on Parallel and Distributed Systems, 28(5):1346–1360, May 2017.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/05/07600358-
                                        abs.html.

Liu:2019:LSM

[LS19] Y. Liu and X. Sun. LPM: A systematic methodology for concurrent data access pattern optimization from a matching perspective. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2478–2493, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:CER

[LSB^{+}18] Xiangbo Li, Mohsen Amini Salehi, Magdy Bayoumi, Nian-Feng Tzeng, and Rajkumar Buyya. Cost-efficient and robust on-demand video transcoding using heterogeneous cloud services. IEEE Transactions on Parallel and Distributed Systems, 29 (3):556–571, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8081853/.

Liu:2016:SDR

[LSC16] Guoxin Liu, Haiying Shen, and Harrison Chandler. Selective data replication for online social networks with distributed datacenters. IEEE Transactions on Parallel and Distributed Systems, 27(8):2377–2393, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07286841-
abs.html.

Liu:2016:MEL

[LSCL16] Guoxin Liu, Haiying Shen, Harrison Chandler, and Jin Li. Measuring and evaluating live content consistency in a large-scale CDN. IEEE Transactions on Parallel and Distributed Systems, 27(7):2074–2090, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07270312-
abs.html.

Li:2019:PAM

[LSJ^{+}19] Xiangbo Li, Mohsen Amini Salehi, Yamini Joshi, Mahmoud K. Darwich, Brad Landreneau, and Magdy Bayoumi. Performance analysis and modeling of video transcoding using heterogeneous cloud services. IEEE Transactions on Parallel and Distributed Systems, 30(4):910–922, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8466657/

.

Liang:2013:IID

[LSKZ13] Ke Liang, Beomjoo Seo, Andrew Kryczka, and Roger Zimmermann. IDM: An indirect dissemination mechanism for spatial voice interaction in networked virtual environments. IEEE Transactions on Parallel and Distributed Systems, 24(2):356– 367, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2010:FPA

[LSL^{+}10] Fangming Liu, Ye Sun, Bo Li, Baochun Li, and Xinyan Zhang. FS2You: Peer-assisted semipersistent online hosting at a large scale. IEEE Transactions on Parallel and Distributed Systems, 21(10):1442–1457, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:TPA

[LSL^{+}14a] Jianhua Li, Liang Shi, Qing’an Li, Chun Jason Xue, and Yinlong Xu. Thread progress aware coherence adaption for hybrid cache coherence protocols. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2697–2707, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/

trans/td/2014/10/06594732-
abs.html.

Li:2014:LSP

[LSL14b] Ming Li, Sergio Salinas, and Pan Li. LocaWard: A security and privacy aware location-based rewarding system. IEEE Transactions on Parallel and Distributed Systems, 25(2):343– 352, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2017:MBA

[LSL^{+}17] Longjun Liu, Hongbin Sun, Chao Li, Tao Li, Jingmin Xin, and Nanning Zheng. Managing battery aging for high energy availability in green datacenters. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3521–3536, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07942036-
abs.html.

Liu:2018:ECH

[LSL^{+}18] Longjun Liu, Hongbin Sun, Chao Li, Yang Hu, Tao Li, and Nanning Zheng. Exploring customizable heterogeneous power distribution and management for datacenter. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2798–2813, December 2018.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/12/08368308-
                                        abs.html.

Li:2017:EDH

[LSLD17] Zhuozhao Li, Haiying Shen, Walter Ligon, and Jeffrey Denton. An exploration of designing a hybrid scaleup/out Hadoop architecture based on performance measurements. IEEE Transactions on Parallel and Distributed Systems, 28(2):386– 400, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/02/07480403-
                                        abs.html.

Liu:2019:PAM

[LSN19] J. Liu, H. Shen, and H. S. Narman. Popularity-aware multi-failure resilient and cost-effective replication for high data durability in cloud storage. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2355–2369, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2015:AAC

[LSW^{+}15] Zongqing Lu, Xiao Sun, Yonggang Wen, Guohong Cao, and Thomas La Porta. Algorithms and applications for community detection in weighted networks. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2916–2926, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06954543-
abs.html.

Liu:2016:DGS

[LSW16] Guoxin Liu, Haiying Shen, and Haoyu Wang. Deadline guaranteed service for multitenant cloud storage. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2851–2865, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07368211-
abs.html.

Lastovetsky:2017:MBO

[LSW17a] Alexey Lastovetsky, Lukasz Szustak, and Roman Wyrzykowski. Model-based optimization of EULAG kernel on Intel Xeon Phi through load imbalancing. IEEE Transactions on Parallel and Distributed Systems, 28(3):787– 797, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07542201-
abs.html.

Liu:2017:ESG

[LSW17b] Guoxin Liu, Haiying Shen, and Haoyu Wang. An economical and SLO-guaranteed cloud storage service across multiple cloud service providers. IEEE Transactions on Parallel and Distributed Systems, 28(9):2440–2453, September 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07865999-
                                        abs.html.

Liu:2017:TLV

[LSW17c] Guoxin Liu, Haiying Shen, and Haoyu Wang. Towards long-view computing load balancing in cluster storage systems. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1770–1784, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/06/07756366-
                                        abs.html.

Li:2016:AAS

[LSWR16] Shuhui Li, Miao Song, Peng-Jun Wan, and Shangping Ren. A 2-approximation algorithm for scheduling parallel and time-sensitive applications to maximize total accrued utility value. IEEE Transactions on Parallel and Distributed Systems, 27(7):1864–1878, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07229345-
abs.html.

Lin:2010:SDE

[LT10] Hsiao-Ying Lin and Wen-Guey Tzeng. A secure decentralized erasure code for distributed networked storage. IEEE Transactions on Parallel and Distributed Systems, 21(11):1586–1594, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2012:SEC

[LT12] Hsiao-Ying Lin and Wen-Guey Tzeng. A secure erasure code-based cloud storage system with secure data forwarding. IEEE Transactions on Parallel and Distributed Systems, 23(6):995– 1003, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Langr:2016:ECS

[LT16] Daniel Langr and Pavel Tvrdik. Evaluation criteria for sparse matrix storage formats. IEEE Transactions on Parallel and Distributed Systems, 27(2):428– 440, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07036061-
abs.html.

Liu:2012:OIC

[LTBN^{+}12] Bin Liu, Peter Terlecky, Amotz Bar-Noy, Ramesh Govindan, Micheal J. Neely, and Dror Rawitz. Optimizing information credibility in social swarming applications. IEEE Transactions on Parallel and Distributed Systems, 23(6):1147–1158, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2016:DPD

[LTC16] Yusen Li, Xueyan Tang, and Wentong Cai. Dynamic bin packing for on-demand cloud resource allocation. IEEE Transactions on Parallel and Distributed Systems, 27 (1):157–170, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07014234-
abs.html.

Li:2019:REI

[LTC^{+}19] Y. Li, X. Tang, W. Cai, J. Tong, X. Liu, and G. Wang. Resource-efficient index shard replication in large scale search engines. IEEE Transactions on Parallel and Distributed Systems, 30(12):2820–2835, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2016:PSS

[LTGI16] Jianwei Liao, Francois Trahay, Balazs Gerofi, and Yutaka Ishikawa. Prefetching on storage servers through mining access patterns on blocks. IEEE Transactions on Parallel and Distributed Systems, 27(9):2698–2710, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07313010-
abs.html.

Li:2014:EES

[LTL14] Kenli Li, Xiaoyong Tang, and Keqin Li. Energyefficient stochastic task scheduling on heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2867–2876, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06642030-
abs.html.

Liu:2011:CNA

[LTM11] Alex X. Liu, Eric Torng, and Chad R. Meiners. Compressing network access control lists. IEEE Transactions on Parallel and Distributed Systems, 22(12): 1969–1977, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2011:DFD

[LTMD11] Hong Luo, Huixiang Tao, Huadong Ma, and Sajal K. Das. Data fusion with desired reliability in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(3): 501–513, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:OTR

[LTW^{+}14] Tao Li, Feng Tan, Qixin Wang, Lei Bu, Jian-Nong Cao, and Xue Liu. From offline toward real time: A hybrid systems model checking and CPS codesign approach for medical device plug-and-play collaborations. IEEE Transactions on Parallel and Distributed Systems, 25(3): 642–652, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2014:DSW

[Lu14]  Jian Lu.  Design of a

sliding window over distributed and asynchronous event streams. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2551–2560, October 2014.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06600682-
abs.html.

Li:2015:DQA

[LV15] Xiaorong Li and Bharadwaj Veeravalli. A differentiated quality adaptation approach for scalable streaming services. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2089–2099, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06748101-
abs.html.

Liu:2017:LSS

[LV17] Shengyun Liu and Marko Vukolic. Leader set selection for low-latency georeplicated state machine. IEEE Transactions on Parallel and Distributed Systems, 28(7):1933–1946, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07774985-
abs.html.

Li:2011:TCC

[LVA^{+}11] Yingshu Li, Chinh Vu, Chunyu Ai, Guantao Chen, and Yi Zhao. Transforming complete coverage algorithms to partial coverage algorithms for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(4): 695–703, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lu:2011:DWB

[LVD11] Lanyue Lu, Peter J. Varman, and Kshitij Doshi. Decomposing workload bursts for efficient storage resource management. IEEE Transactions on Parallel and Distributed Systems, 22(5):860– 873, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Le:2011:EMO

[LW11] Duy Le and Haining Wang. An effective memory optimization for virtual machinebased systems. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1705–1713, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2012:MOF

[LW12] Cong Liu and Jie Wu. On multicopy opportunistic forwarding protocols in nondeterministic delay tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1121–1128, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2014:RBP

[LW14] Bill Lin and Hao Wang.

Reservation-based packet buffers with deterministic packet departures. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1297–1305, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:MEC

[LW15] Dawei Li Li and Jie Wu.

Minimizing energy consumption for frame-based tasks on heterogeneous multiprocessor platforms. IEEE Transactions on Parallel and Distributed Systems, 26 (3):810–823, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06777565-
abs.html.

Li:2017:AET

[LWC^{+}17] Hong-Wei Li, Yu-Sung Wu, Yi-Yung Chen, Chieh-Min Wang, and Yen-Nun Huang. Application execution time prediction for effective CPU provisioning in virtualization environment. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3074–3088, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07933268-
abs.html.

Lin:2010:SSB

[LWCG10] Kate Ching-Ju Lin, Chun-Po Wang, Cheng-Fu Chou, and Leana Golubchik. SocioNet: a social-based multimedia access system for unstructured P2P networks. IEEE Transactions on Parallel and Distributed Systems, 21(7):1027–1041, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2018:GCG

[LWCL18] Chengjian Liu, Qiang Wang, Xiaowen Chu, and Yiu-Wing Leung. G-CRS: GPU accelerated Cauchy Reed– Solomon coding. IEEE Transactions on Parallel and Distributed Systems, 29(7): 1484–1498, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08252749-
abs.html.

Luo:2012:SDW

[LWG^{+}12] Hanjiang Luo, Kaishun Wu, Zhongwen Guo, Lin Gu, and Lionel M. Ni. Ship detection with wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(7):

1336–1343, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2015:ACD

[LWJ^{+}15] Wenping Liu, Dan Wang, Hongbo Jiang, Wenyu Liu, and Chonggang Wang. An approximate convex decomposition protocol for wireless sensor network localization in arbitrary-shaped fields. IEEE Transactions on Parallel and Distributed Systems, 26(12):3264–3274, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/07000585-
abs.html.

Liu:2017:PPP

[LWL^{+}17] Qin Liu, Guojun Wang, Feng Li, Shuhui Yang, and Jie Wu. Preserving privacy with probabilistic indistinguishability in weighted social networks. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1417–1429, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07582550-
abs.html.

Lu:2019:UFB

[LWL^{+}19] Jianyuan Lu, Ying Wan, Yang Li, Chuwen Zhang, Huichen Dai, Yi Wang, Gong Zhang, and Bin Liu. Ultra-fast Bloom filters using SIMD techniques. IEEE Transactions on Parallel and Distributed Systems, 30(4):953–964, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8462781/

.

Li:2017:TTD

[LWLZ17] Dawei Li, Jie Wu, Zhiyong Liu, and Fa Zhang. Towards the tradeoffs in designing data center network architectures. IEEE Transactions on Parallel and Distributed Systems, 28(1):260– 273, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07779314-
abs.html.

Liu:2012:OMI

[LWS^{+}12] Duo Liu, Yi Wang, Zili Shao, Minyi Guo, and Jingling Xue. Optimally maximizing iteration-level loop parallelism. IEEE Transactions on Parallel and Distributed Systems, 23(3):564– 572, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2019:RCT

[LWSM19] B. Li, X. Wang, A. K. Singh, and T. Mak. On runtime communication and thermalaware application mapping and defragmentation in 3D NoC systems. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2775–2789, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:HRA

[LWT^{+}18] Yongkun Li, Neng Wang, Chengjin Tian, Si Wu, Yueming Zhang, and Yinlong Xu. A hierarchical RAID architecture towards fast recovery and high reliability. IEEE Transactions on Parallel and Distributed Systems, 29(4):734–747, April 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08115198-
abs.html.

Liu:2013:CSN

[LWW^{+}13] Weichen Liu, Yu Wang, Xuan Wang, Jiang Xu, and Huazhong Yang. Onchip sensor network for efficient management of power gating-induced power/ ground noise in multiprocessor system on chip. IEEE Transactions on Parallel and Distributed Systems, 24(4): 767–777, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2011:SOG

[LWX^{+}11] Zhenhua Li, Jie Wu, Junfeng Xie, Tieying Zhang, Guihai Chen, and Yafei Dai. Stability-optimal grouping strategy of peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2079–2087, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:ALE

[LWY^{+}13] Hongxing Li, Chuan Wu, Dongxiao Yu, Qiang-Sheng Hua, and Francis C. M. Lau. Aggregation latencyenergy tradeoff in wireless sensor networks with successive interference cancellation. IEEE Transactions on Parallel and Distributed Systems, 24(11):2160–2170, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:LLA

[LWY^{+}15] Zhong Li, Cheng Wang, Siqian Yang, Changjun Jiang, and Xiangyang Li. LASS: Local-activity and social-similarity based data forwarding in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 26(1):174– 184, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748062-
abs.html.

Li:2019:CBB

[LWY^{+}19] M. Li, J. Weng, A. Yang, W. Lu, Y. Zhang, L. Hou, J. Liu, Y. Xiang, and R. H. Deng. CrowdBC: a blockchain-based decentralized framework for crowdsourcing. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1251–1266, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2012:DMP

[LWZ12] Ji Luo, Dan Wang, and Qian Zhang. On the double mobility problem for water surface coverage with mobile sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(1):146– 159, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2013:PBC

[LWZ^{+}13] Xiaocheng Liu, Chen Wang, Bing Bing Zhou, Junliang Chen, Ting Yang, and Albert Y. Zomaya. Prioritybased consolidation of parallel workloads in the cloud. IEEE Transactions on Parallel and Distributed Systems, 24(9):1874–1883, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:TID

[LWZ14] Bo Li, Yuchen Wu, and Yanmin Zhu. Trajectory improves data delivery in urban vehicular networks. IEEE Transactions on Parallel and Distributed Systems, 25(4): 1089–1100, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2015:MMT

[LWZ^{+}15] Zhuofan Liao, Jianxin Wang, Shigeng Zhang, Jiannong Cao, and Geyong Min. Minimizing movement for target coverage and network connectivity in mobile sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1971–1983, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/07/06846302-
                                        abs.html.

Liu:2016:TER

[LWZ^{+}16a] Leibo Liu, Junbin Wang, Jianfeng Zhu, Chenchen Deng, Shouyi Yin, and Shaojun Wei. TLIA: Efficient reconfigurable architecture for control-intensive kernels with triggered-long-instructions. IEEE Transactions on Parallel and Distributed Systems, 27(7):2143–2154, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07254223.
pdf.

Lu:2016:PAM

[LWZ^{+}16b] Yunping Lu, Xin Wang, Weihua Zhang, Haibo Chen, Lu Peng, and Wenyun Zhao. Performance analysis of multimedia retrieval workloads running on multicores. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3323–3337, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07416213-
abs.html.

Luo:2016:DIH

[LWZ^{+}16c] Dan Luo, Jiguang Wan, Yifeng Zhu, Nannan Zhao, Feng Li, and Changsheng Xie. Design and implementation of a hybrid shingled write disk system. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1017–1029, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2016/04/07091937-
                                        abs.html.

Liu:2010:IBI

[LX10] Xiaomei Liu and Li Xiao. ILBO: Balance inbound traffic dynamically in multihomed stub networks. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1561–1572, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2012:EAD

[LX12] Wei Luo and Dong Xiang. An efficient adaptive deadlock-free routing algorithm for torus networks. IEEE Transactions on Parallel and Distributed Systems, 23(5):800–808, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2013:UIP

[LXBZ13] Jiaqing Luo, Bin Xiao, Kai Bu, and Shijie Zhou. Understanding and improving piece-related algorithms in the BitTorrent protocol. IEEE Transactions on Parallel and Distributed Systems, 24(12):2526–2537, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2011:CRP

[LXHL11] Zhenyu Li, Gaogang Xie, Kai Hwang, and Zhongcheng Li. Churn-resilient protocol for massive data dissemination in P2P networks. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1342–1349, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:AAV

[LXHS12] Liqun Li, Guoliang Xing, Qi Han, and Limin Sun. ASM: Adaptive voice stream multicast over low-power wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(4): 626–633, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2016:TRV

[LXXH16] Wei Li, Kaiping Xue, Yingjie Xue, and Jianan Hong. TMACS: A robust and verifiable threshold multiauthority access control system in public cloud storage. IEEE Transactions on Parallel and Distributed Systems, 27(5):1484–1496, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07130642-
abs.html.

Liu:2015:UTI

[LXZB15] Xuan Liu, Bin Xiao, Shigeng Zhang, and Kai Bu. Unknown tag identification in large RFID systems: An efficient and complete solution. IEEE Transactions on Parallel and Distributed Systems, 26(6):1775–1788, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/06/06820770-
                                        abs.html.

Lin:2016:ERC

[LXZH16] Limei Lin, Li Xu, Shuming Zhou, and Sun-Yuan Hsieh. The extra, restricted connectivity and conditional diagnosability of split-star networks. IEEE Transactions on Parallel and Distributed Systems, 27(2): 533–545, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07031970-
                                        abs.html.

Liu:2011:ATI

[LY11] Lin Liu and Yuanyuan Yang. Achieving 100% throughput in input-buffered WDM optical packet interconnects. IEEE Transactions on Parallel and Distributed Systems, 22(2):273–286, February 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2014:AaC

[LY14] Hwa-Chun Lin and Hsiu-Ming Yang. An approximation algorithm for constructing degree-dependent nodeweighted multicast trees. IEEE Transactions on Parallel and Distributed Systems, 25(8):1976–1985, August 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2016:GVC

[LY16a] Zhenhua Li and Yuanyuan Yang. GBC3: A versatile cube-based servercentric network for data centers. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2895–2910, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07364277-
abs.html.

Lim:2016:AKE

[LY16b] Hoon Wei Lim and Guomin Yang. Authenticated key exchange protocols for parallel network file systems. IEEE Transactions on Parallel and Distributed Systems, 27(1):92–105, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07004049-
abs.html.

Li:2012:GER

[LYGX12] Jianzhong Li, Lei Yu, Hong Gao, and Shuguang Xiong. Grouping-enhanced resilient probabilistic en-route filtering of injected false data in WSNs. IEEE Transactions on Parallel and Distributed Systems, 23(5):881– 889, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2015:WSD

[LYH^{+}15] Dan Li, Yirong Yu, Wu He, Kai Zheng, and Bingsheng He. Willow: Saving data center network energy for network-limited flows. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2610–2620, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06882231.
pdf.

Li:2015:PAO

[LYL15] Kenli Li, Wangdong Yang, and Keqin Li. Performance analysis and optimization for SpMV on GPU using probabilistic modeling. IEEE Transactions on Parallel and Distributed Systems, 26 (1):196–205, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748055-
abs.html.

Li:2016:HPS

[LYL16] Kenli Li, Wangdong Yang, and Keqin Li. A hybrid parallel solving algorithm on GPU for quasitridiagonal system of linear equations. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2795–2808, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07378515-
abs.html.

Liu:2019:HHM

[LYPL19] L. Liu, S. Yang, L. Peng, and X. Li. Hierarchical hybrid memory management in OS for tiered memory systems. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2223–2236, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2012:OFT

[LYW^{+}12] Bowen Li, Panlong Yang, Jinlong Wang, Qihui Wu, Shao-Jie Tang, Xiang-Yang Li, and Yunhao Liu. Optimal frequency-temporal opportunity exploitation for multichannel ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 23(12):2289–2302, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lin:2016:TME

[LYY16] Jie Lin, Wei Yu, and Xinyu Yang. Towards multistep electricity prices in smart grid electricity markets. IEEE Transactions on Parallel and Distributed Systems, 27(1):286– 302, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07004067-
                                        abs.html.

Li:2013:SSS

[LYZ^{+}13] Ming Li, Shucheng Yu, Yao Zheng, Kui Ren, and Wenjing Lou. Scalable and secure sharing of personal health records in cloud computing using attribute-based encryption. IEEE Transactions on Parallel and Distributed Systems, 24(1):131– 143, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2016:TPN

[LYZ^{+}16] Shouxi Luo, Hongfang Yu, Yangming Zhao, Sheng Wang, Shui Yu, and Lemin Li. Towards practical and near-optimal coflow scheduling for data center networks. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3366–3380, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07399419-
abs.html.

Lu:2018:DAM

[LYZL18] Li Lu, Jiadi Yu, Yanmin Zhu, and Minglu Li. A double auction mechanism to bridge users? Task requirements and providers? Resources in two-sided cloud markets. IEEE Transactions on Parallel and Distributed Systems, 29(4):720–733, April 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08170253-
abs.html.

Li:2010:RST

[LZ10] Ming Li and Wei Zhao.

Representation of a stochastic traffic bound. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1368–1372, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2011:ECS

[LZ11] Young Choon Lee and Albert Y. Zomaya. Energy conscious scheduling for distributed computing systems under different operating conditions. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1374–1381, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lama:2012:ESP

[LZ12] Palden Lama and Xiaobo Zhou. Efficient server provisioning with control for endto-end response time guarantee on multitier clusters. IEEE Transactions on Parallel and Distributed Systems, 23(1):78–86, January 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:SPA

[LZB14] Yanhua Li, Zhi-Li Zhang, and Daniel Boley. From shortest-path to all-path: The routing continuum theory and its applications. IEEE Transactions on Parallel and Distributed Systems, 25(7):1745–1755, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2012:MFT

[LZC^{+}12] Yunhao Liu, Yiyang Zhao, Lei Chen, Jian Pei, and Jinsong Han. Mining frequent trajectory patterns for activity monitoring using radio frequency tag arrays. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2138–2149, November 2012.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2014:CRA

[LZCK14] Jingwei Liu, Zonghua Zhang, Xiaofeng Chen, and Kyung Sup Kwak. Certificateless remote anonymous authentication schemes for wireless body area networks. IEEE Transactions on Parallel and Distributed Systems, 25(2): 332–342, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liao:2016:PDF

[LZH^{+}16] Xiaofei Liao, Long Zheng, Bingsheng He, Song Wu, and Hai Jin. A performance debugging framework for unnecessary lock contentions with record/replay techniques. IEEE Transactions on Parallel and Distributed Systems, 27(7): 1889–1901, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07222463-
abs.html.

Li:2018:COM

[LZH18] Shigang Li, Yunquan Zhang, and Torsten Hoefler. Cacheoblivious MPI all-to-all communications based on Morton order. IEEE Transactions on Parallel and Distributed Systems, 29(3):542– 555, March 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8091010/.

Liu:2015:CCD

[LZK^{+}15] Xiao-Yang Liu, Yanmin Zhu, Linghe Kong, Cong Liu, Yu Gu, Athanasios V. Vasilakos, and Min-You Wu. CDC: Compressive data collection for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2188–2197, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/08/06870490-
                                        abs.html.

Lan:2010:TAA

[LZL10] Zhiling Lan, Ziming Zheng, and Yawei Li. Toward automated anomaly identification in large-scale systems. IEEE Transactions on Parallel and Distributed Systems, 21(2): 174–187, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:RCA

[LZL^{+}18] Sizhao Li, Yuanzhi Zhang, Hongyin Luo, Yan Chen, Chao Lu, and Donghui Guo. Race-condition-aware and hardware-oriented task partitioning and scheduling using entropy maximization. IEEE Transactions on Parallel and Distributed Systems, 29(7):1589–1604, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08226841-
abs.html.

Liu:2010:OBT

[LZN10] Yunhuai Liu, Qian Zhang, and Lionel M. Ni. Opportunitybased topology control in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(3):405–416, March 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2011:ROT

[LZNX11] Yunhuai Liu, Yanmin Zhu, Lionel M. Ni, and Guangtao Xue. A reliability-oriented transmission service in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2100–2107, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2013:MSE

[LZP^{+}13] Jun Liu, Zhong Zhou, Zheng Peng, Jun-Hong Cui, Michael Zuba, and Lance Fiondella. Mobi-Sync: Efficient time synchronization for mobile underwater sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):406– 416, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liu:2017:DAT

[LZW^{+}17] Duo Liu, Kan Zhong, Tianzheng Wang, Yi Wang, Zili Shao, Edwin Hsing-Mean Sha, and Jingling Xue. Durable address translation in PCM-based flash storage systems. IEEE Transactions on Parallel and Distributed Systems, 28(2):475– 490, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/02/07501824-
                                        abs.html.

Liu:2013:MSM

[LZWY13] Xuefeng Liu, Yuqing Zhang, Boyang Wang, and Jingbo Yan. Mona: Secure multiowner data sharing for dynamic groups in the cloud. IEEE Transactions on Parallel and Distributed Systems, 24(6):1182–1191, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2014:GSF

[LZWY14] Keqin Li, Weimin Zheng, Yongwei Wu, and Yulai Yuan. Guarantee strict fairness and utilize prediction better in parallel job scheduling. IEEE Transactions on Parallel and Distributed Systems, 25(4):971– 981, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2019:HAM

[LZWZ19] Xiaqing Li, Guangyan Zhang, Zhufan Wang, and Weimin Zheng. HyConv: Accelerating multi-phase CNN computation by fine-grained policy selection. IEEE Transactions on Parallel and Distributed Systems, 30(2):388– 399, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08428450-
abs.html.

Lin:2015:ECC

[LZXW15] Limei Lin, Shuming Zhou, Li Xu, and Dajin Wang. The extra connectivity and conditional diagnosability of alternating group networks. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2352–2362, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06879268-
abs.html.

Li:2012:SFA

[LZY12] Xiaoyong Li, Feng Zhou, and Xudong Yang. Scalable feedback aggregating (SFA) overlay for large-scale P2P trust management. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1944–1957, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2018:OAI

[LZY^{+}18] Juan Li, Yanmin Zhu, Jiadi Yu, Chengnian Long, Guangtao Xue, and Shiyou Qian. Online auction for IaaS clouds: Towards elastic user demands and weighted heterogeneous VMs. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2075–2089, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08314090-
                                        abs.html.

Li:2019:IIS

[LZY^{+}19] D. Li, Z. Zhang, K. Yu, K. Huang, and T. Tan. ISEE: An intelligent scene exploration and evaluation platform for large-scale visual surveillance. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2743–2758, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Liang:2013:TPL

[LZZP13] Qilian Liang, Baoju Zhang, Chenglin Zhao, and Yiming Pi. TDoA for passive localization: Underwater versus terrestrial environment. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2100–2108, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Morris:2013:MJI

[MA13] Gerald R. Morris and Khalid H. Abed. Mapping a Jacobi iterative solver onto a high-performance heterogeneous computer. IEEE Transactions on Parallel and Distributed Systems, 24(1): 85–91, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Margaritis:2014:ERB

[MA14] Giorgos Margaritis and Stergios V. Anastasiadis. Efficient range-based storage management for scalable datastores. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2851–2866, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2014/11/06684149-
                                        abs.html.

Mann:2016:CPP

[Man16] Zoltan Adam Mann. A comment on “Process Placement in Multicore Clusters: Algorithmic Issues and Practical Techniques”. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2475–2476, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07302074-
                                        abs.html.

Mann:2018:ROA

[Man18] Zoltan Adam Mann. Resource optimization across the cloud stack. IEEE Transactions on Parallel and Distributed Systems, 29(1):169– 182, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/01/08016586-
                                        abs.html.

Mintz:2012:CCA

[MB12] Tiffany M. Mintz and Jason D. Bakos. A clusteron-a-chip architecture for high-throughput phylogeny search. IEEE Transactions on Parallel and Distributed Systems, 23(4):579– 588, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Martelli:2013:MMW

[MB13] Francesca Martelli and Maurizio A. Bonuccelli. Minimum message waiting time scheduling in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 24(9):1797–1806, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Medhat:2019:EEM

[MBF19] Ramy Medhat, Borzoo Bonakdarpour, and Sebastian Fischmeister. Energyefficient multiple producer– consumer. IEEE Transactions on Parallel and Distributed Systems, 30(3):560– 574, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08451894-
abs.html.

Moretti:2010:APA

[MBH^{+}10] Christopher Moretti, Hoang Bui, Karen Hollingsworth, Brandon Rich, Patrick Flynn, and Douglas Thain. Allpairs: An abstraction for data-intensive computing on campus grids. IEEE Transactions on Parallel and Distributed Systems, 21(1):33– 46, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Misic:2013:GEI

[MBMC13] Vojislav B. Misic, Rajkumar Buyya, Dejan Milojicic, and Yong Cui. Guest Editors’ introduction: Special issue on cloud computing. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1062–1065, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Misra:2015:DDD

[MBO15] Sudip Misra, Samaresh Bera, and Tamoghna Ojha. D2P: Distributed Dynamic Pricing Policy in Smart Grid for PHEVs management. IEEE Transactions on Parallel and Distributed Systems, 26(3):702–712, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06782392-
                                        abs.html.

Monti:2011:TRD

[MBV11] Henry M. Monti, Ali R. Butt, and Sudharshan S. Vazhkudai. Timely result-data offloading for improved HPC center scratch provisioning and serviceability. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1307–1322, August 2011.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Monti:2013:TSH

[MBV13] Henry M. Monti, Ali R. Butt, and Sudharshan S. Vazhkudai. On timely staging of HPC job input data. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1841–1851, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meng:2010:HPH

[MC10] Xiandong Meng and Vipin Chaudhary. A highperformance heterogeneous computing platform for biological sequence analysis. IEEE Transactions on Parallel and Distributed Systems, 21(9):1267–1280, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Margara:2014:HPP

[MC14] Alessandro Margara and Gianpaolo Cugola. Highperformance publish-subscribe matching using parallel hardware. IEEE Transactions on Parallel and Distributed Systems, 25(1):126– 135, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mei:2017:DGM

[MC17] Xinxin Mei and Xiaowen Chu. Dissecting GPU memory hierarchy through microbenchmarking. IEEE Transactions on Parallel and Distributed Systems, 28(1): 72–86, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07445236-
abs.html.

Mei:2019:LTM

[MCJT19] Fei Mei, Qiang Cao, Hong Jiang, and Lei Tian. LSMtree managed storage for large-scale key-value store. IEEE Transactions on Parallel and Distributed Systems, 30(2):400–414, February 2019. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08428463-
abs.html.

Manzillo:2012:CCL

[MCMR12] Marco Papa Manzillo, Luigi Ciminiera, Guido Marchetto, and Fulvio Risso. CLOSER: a Collaborative Localityaware Overlay SERvice. IEEE Transactions on Parallel and Distributed Systems, 23(6):1030–1037, June 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mubarak:2017:EPS

[MCRC17] Misbah Mubarak, Christopher D. Carothers, Robert B. Ross, and Philip Carns. Enabling parallel simulation of large-scale HPC network systems. IEEE Transactions on Parallel and Distributed Systems, 28(1):87– 100, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07448965-
abs.html.

Montresor:2013:DCD

[MDM13] Alberto Montresor, Francesco De Pellegrini, and Daniele Miorandi. Distributed k-core decomposition. IEEE Transactions on Parallel and Distributed Systems, 24(2):288– 300, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meng:2014:KKA

[MDZC14] Shunmei Meng, Wanchun Dou, Xuyun Zhang, and Jinjun Chen. KASR: A keyword-aware service recommendation method on MapReduce for big data applications. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3221–3231, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/12/06714480-
                                        abs.html.

Min:2015:DSI

[ME15a] Changwoo Min and Young Ik Eom. Dynamic scheduling of irregular stream programs toward many-core scalability. IEEE Transactions on Parallel and Distributed Systems, 26(6):1594–1607, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/06/06824243-
                                        abs.html.

Min:2015:ILF

[ME15b] Changwoo Min and Young Ik Eom. Integrating lock-free and combining techniques for a practical and scalable FIFO queue. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1910–1922, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/07/06844158-
                                        abs.html.

Matos:2013:SPS

[MFO^{+}13] Miguel Matos, Pascal Felber, Rui Oliveira, Jose O. Pereira, and Etienne Riviere. Scaling up publish/ subscribe overlays using interest correlation for link sharing. IEEE Transactions on Parallel and Distributed Systems, 24(12):2462–2471, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mashayekhy:2014:MSM

[MG14] Lena Mashayekhy and Daniel Grosu. A merge-and-split mechanism for dynamic virtual organization formation in grids. IEEE Transactions on Parallel and Distributed Systems, 25(3):540– 549, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mastoras:2018:UFC

[MG18] Aristeidis Mastoras and Thomas R. Gross. Unifying fixed code mapping, communication, synchronization and scheduling algorithms for efficient and scalable loop pipelining. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2136–2149, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08319946-
abs.html.

Magoules:2018:DCD

[MGB18] Frederic Magoules and Guillaume Gbikpi-Benissan. Distributed convergence detection based on global residual error under asynchronous iterations. IEEE Transactions on Parallel and Distributed Systems, 29(4):819– 829, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/04/08169062-
                                        abs.html.

Munir:2012:MBD

[MGR12] Arslan Munir and Ann Gordon-Ross. An MDPbased dynamic optimization methodology for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(4): 616–625, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Malik:2012:FSC

[MGS12] Avinash Malik, Alain Girault, and Zoran Salcic. Formal semantics, compilation and execution of the GALS programming language DSystemJ. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1240–1254, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Marathe:2016:ERA

[MHL^{+}16] Aniruddha Marathe, Rachel Harris, David K. Lowenthal, Bronis R. de Supinski, Barry Rountree, and Martin Schulz. Exploiting redundancy and application scalability for costeffective, time-constrained execution of HPC applications on Amazon EC2. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2574–2588, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07355374-
abs.html.

Mitani:2017:PEA

[MIH17] Yasuaki Mitani, Fumihiko Ino, and Kenichi Hagihara. Parallelizing exact and approximate string matching via inclusive scan on a GPU. IEEE Transactions on Parallel and Distributed Systems, 28(7):1989–2002, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07797444-
abs.html.

Misic:2014:PVS

[Mis14] Vojislav B. Misic. Probabilistic vs. sequence-based rendezvous in channel-hopping cognitive networks. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2418–2427, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2014/09/06574867-
                                        abs.html.

Mittal:2017:STA

[Mit17] Sparsh Mittal. A survey of techniques for architecting and managing GPU register file. IEEE Transactions on Parallel and Distributed Systems, 28(1):16– 28, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/01/07448930-
                                        abs.html.

Morris:2014:EPE

[MJK14] Randy Morris, Evan Jolley, and Avinash Karanth Kodi. Extending the performance and energy-efficiency of shared memory multicores with nanophotonic technology. IEEE Transactions on Parallel and Distributed Systems, 25(1):83– 92, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moeng:2016:WTF

[MJM16] Michael Moeng, Alex K. Jones, and Rami G. Melhem. Weighted-tuple: Fast and accurate synchronization for parallel architecture simulators. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2462–2474, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07308077-
abs.html.

Masuzawa:2014:RGM

[MKOK14] Toshimitsu Masuzawa, Hirotsugu Kakugawa, Fukuhito Ooshita, and Shinji Kawai. Randomized gathering of mobile agents in anonymous unidirectional ring networks. IEEE Transactions on Parallel and Distributed Systems, 25(5):1289–1296, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mohammadi:2018:HAR

[MKSN18] Mahnaz Mohammadi, Akhil Krishna, Nalesh S., and S. K. Nandy. A hardware architecture for radial basis function neural network classifier. IEEE Transactions on Parallel and Distributed Systems, 29(3):481–495, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8094305/.

Meng:2012:RAA

[MKVL12] Shicong Meng, Srinivas Raghav Kashyap, Chitra Venkatramani, and Ling Liu. Resource-aware application state monitoring. IEEE Transactions on Parallel and Distributed Systems, 23(12):

2315–2329, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2015:OCM

[MLC^{+}15] Qiang Ma, Kebin Liu, Zhichao Cao, Tong Zhu, Xin Miao, and Yunhao Liu. Opportunistic concurrency: A MAC protocol for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1999–2008, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/07/06826533-
                                        abs.html.

Ma:2019:CDI

[MLC^{+}19] S. Ma, Z. Liu, S. Chen, L. Huang, Y. Guo, Z. Wang, and M. Zhang. Coordinated DMA: Improving the DRAM access efficiency for matrix multiplication. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2148–2164, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moore:2015:KSP

[MLK15] Nicholas Moore, Miriam Leeser, and Laurie Smith King. Kernel specialization provides adaptable GPU code for particle image velocimetry. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1049–1058, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06798707-
abs.html.

Ma:2014:CLS

[MLL14] Junchao Ma, Wei Lou, and Xiang-Yang Li. Contiguous link scheduling for data aggregation in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1691–1701, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2015:SAD

[MLML15] Qiang Ma, Kebin Liu, Xin Miao, and Yunhao Liu. Sherlock is around: Detecting network failures with local evidence fusion. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1430–1440, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06807767-
abs.html.

Mahmoud:2015:SRR

[MLS15] Mohamed M. E. A. Mahmoud, Xiaodong Lin, and Xuemin Sherman Shen. Secure and reliable routing protocols for heterogeneous multihop wireless networks. IEEE Transactions on Parallel and Distributed Systems, 26(4):1140–1153, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06519238-
                                        abs.html.

Mao:2013:FBW

[MLT^{+}13] Xufei Mao, Yunhao Liu, Shaojie Tang, Huafu Liu, Jiankang Han, and Xiang-Yang Li. Finding best and worst k-coverage paths in multihop wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(12):2396–2406, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mei:2019:PMC

[MLT^{+}19] Jing Mei, Kenli Li, Zhao Tong, Qiang Li, and Keqin Li. Profit maximization for cloud brokers in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 30(1):190– 203, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/01/08399541-
                                        abs.html.

Moreira:2012:CRT

[MLVD12] Marcelo Duffles Donato Moreira, Rafael Pinaud Laufer, Pedro Braconnot Velloso, and Otto Carlos M. B. Duarte. Capacity and robustness tradeoffs in Bloom filters for distributed applications. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2219–2230, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2019:PMA

[MLXG19] Y. Ma, W. Liang, Z. Xu, and S. Guo. Profit maximization for admitting requests with network function services in distributed clouds. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1143–1157, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Misra:2010:MCD

[MM10] Rajiv Misra and Chittaranjan Mandal. Minimum connected dominating set using a collaborative cover heuristic for ad hoc sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(3):292– 302, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2012:QOV

[MM12] Dong Ma and Maode Ma. A QoS oriented vertical handoff scheme for WiMAX/WLAN overlay networks. IEEE Transactions on Parallel and Distributed Systems, 23(4): 598–606, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Manatakis:2015:ESE

[MM15] Dimitris V. Manatakis and Elias S. Manolakos. Estimating the spatiotemporal evolution characteristics of diffusive hazards using wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2444–2458, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/09/06895267.
                                        pdf.

Martinez-Morais:2010:PQD

[MMACS10] Raul Martinez-Morais, Francisco J. Alfaro-Cortes, and Jose L. Sanchez. Providing QoS with the deficit table scheduler. IEEE Transactions on Parallel and Distributed Systems, 21(3):327– 341, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mohror:2014:DME

[MMBdS14] Kathryn Mohror, Adam Moody, Greg Bronevetsky, and Bronis R. de Supinski. Detailed modeling and evaluation of a scalable multilevel checkpointing system. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2255–2263, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06494566-
abs.html.

Marangozova-Martin:2019:MLE

[MMdE19]  V.  Marangozova-Martin,

N. de Palma, and A. El Rheddane. Multi-level elasticity for data stream processing. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2326–2337, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahale:2016:RRF

[MMNN16] Gopinath Mahale, Hamsika Mahale, S. K. Nandy, and Ranjani Narayan. RE-FRESH: REDEFINE for face recognition using SURE homogeneous cores. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3602–3616, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07426839-
abs.html.

Moraveji:2011:MTZ

[MMSAZ11] Reza Moraveji, Parya Moinzadeh, Hamid Sarbazi-Azad, and Albert Y. Zomaya. Multispanning tree zoneordered label-based routing algorithms for irregular networks. IEEE Transactions on Parallel and Distributed Systems, 22(5):817– 832, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mei:2015:SAS

[MMSS15] Alessandro Mei, Giacomo Morabito, Paolo Santi, and Julinda Stefa. Social-aware stateless routingin pocket switched networks. IEEE Transactions on Parallel and Distributed Systems, 26 (1):252–261, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06747341-
                                        abs.html.

Maglione-Mathey:2018:SDF

[MMYES^{+}18] German Maglione-Mathey, Pedro Yebenes, Jesus Escudero-Sahuquillo, Pedro Javier Garcia, Francisco J. Quiles, and Eitan Zahavi. Scalable deadlock-free deterministic minimal-path routing engine for InfiniBand-based Dragonfly networks. IEEE Transactions on Parallel and Distributed Systems, 29(1):

183–197, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08013852-
abs.html.

Minh:2014:PWM

[MNE14] Tran Ngoc Minh, Thoai Nam, and Dick H. J. Epema. Parallel workload modeling with realistic characteristics. IEEE Transactions on Parallel and Distributed Systems, 25(8):2138–2148, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mashayekhy:2015:PMP

[MNG15a] Lena Mashayekhy, Mahyar Movahed Nejad, and Daniel Grosu. A PTAS mechanism for provisioning and allocation of heterogeneous cloud resources. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2386–2399, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06893022.
pdf.

Mashayekhy:2015:EAS

[MNG^{+}15b] Lena Mashayekhy, Mahyar Movahed Nejad, Daniel Grosu, Quan Zhang, and Weisong Shi. Energy-aware scheduling of MapReduce jobs for big data applications. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2720–2733, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06899694.
pdf.

Meneses:2015:UMO

[MNZ^{+}15] Esteban Meneses, Xiang Ni, Gengbin Zheng, Celso L. Mendes, and Laxmikant V. Kale. Using migratable objects to enhance fault tolerance schemes in supercomputers. IEEE Transactions on Parallel and Distributed Systems, 26(7): 2061–2074, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06862914-
abs.html.

Millot:2016:OPD

[MP16] Daniel Millot and Christian Parrot. Optimization of the processing of data streams on roughly characterized distributed resources. IEEE Transactions on Parallel and Distributed Systems, 27(5):1415–1429, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2016/05/07128739-
abs.html.

Mohamedin:2017:MRL

[MPHR17] Mohamed Mohamedin, Roberto Palmieri, Ahmed Hassan, and Binoy Ravindran. Managing resource limitation of best-effort HTM. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2299–2313, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07852519-
abs.html.

Marinescu:2017:CRS

[MPM17] Dan C. Marinescu, Ashkan Paya, and John P. Morrison. A cloud reservation system for big data applications. IEEE Transactions on Parallel and Distributed Systems, 28(3):606–618, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07523396-
abs.html.

Misra:2015:DIB

[MPS15] Sudip Misra, Sujata Pal, and Barun Kumar Saha. Distributed information-based cooperative strategy adaptation in opportunistic mobile networks. IEEE Transactions on Parallel and Distributed Systems, 26(3):724– 737, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06781642-
                                        abs.html.

Mostefaoui:2016:ITB

[MR16] Achour Mostefaoui and Michel Raynal. Intrusiontolerant broadcast and agreement abstractions in the presence of Byzantine processes. IEEE Transactions on Parallel and Distributed Systems, 27(4):1085–1098, April 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07097735-
                                        abs.html.

Munir:2012:HPE

[MRGR12] Arslan Munir, Sanjay Ranka, and Ann Gordon-Ross. Highperformance energy-efficient multicore embedded computing. IEEE Transactions on Parallel and Distributed Systems, 23(4):684– 700, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Membarth:2016:HDS

[MRH^{+}16] Richard Membarth, Oliver Reiche, Frank Hannig, Jurgen Teich, Mario Korner, and Wieland Eckert. HIPA^{cc}: A domain-specific language and compiler for image processing. IEEE Transactions on Parallel and Distributed Systems, 27(1):210– 224, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07017495-
abs.html.

Misic:2012:AIT

[MRM12] Jelena Mišić, Saeed Rashwand, and Vojislav B. Mišić. Analysis of impact of TXOP allocation on IEEE 802.11e EDCA under variable network load. IEEE Transactions on Parallel and Distributed Systems, 23(5):785– 799, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahmoud:2012:CBS

[MS12] Mohamed M. E. A. Mahmoud and Xuemin (Sherman) Shen. A cloudbased scheme for protecting source-location privacy against hotspot-locating attack in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(10):1805–1818, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mahmoud:2013:SPS

[MS13a] Mohamed M. E. A. Mahmoud and Xuemin (Sherman) Shen. A secure payment scheme with low communication and processing overhead for multihop wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):209– 224, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Min:2013:RTS

[MS13b] A. W. Min and K. G. Shin. Robust tracking of smallscale mobile primary user in cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 24(4):778–788, April 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Marvasti:2015:AHN

[MS15] Mohammadreza Binesh Marvasti and Ted H. Szymanski. An analysis of hypermesh NoCs in FPGAs. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2643–2656, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06910284.
                                        pdf.

Mokdad:2011:CAC

[MSB11] Lynda Mokdad, Mbaye Sene, and Azzedine Boukerche.

Call admission control performance analysis in mobile networks using stochastic well-formed Petri nets. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1332–1341, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meyerhenke:2017:PGP

[MSS17] Henning Meyerhenke, Peter Sanders, and Christian Schulz. Parallel graph partitioning for complex networks. IEEE Transactions on Parallel and Distributed Systems, 28(9):2625–2638, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07859409-
abs.html.

Marcon:2014:RAE

[MSSB14] Arlindo Luis Marcon, Altair Olivo Santin, Maicon Stihler, and Juliana Bachtold. A (UCON_{ABC}) resilient authorization evaluation for cloud computing. IEEE Transactions on Parallel and Distributed Systems, 25(2): 457–467, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Marchal:2018:MTG

[MSSV18] Loris Marchal, Bertrand Simon, Oliver Sinnen, and Frederic Vivien. Malleable task-graph scheduling with a practical speed-up model. IEEE Transactions on Parallel and Distributed Systems, 29(6):1357–1370, June 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/06/08259313-
                                        abs.html.

Madduri:2012:OPP

[MSW^{+}12] Kamesh Madduri, Jimmy Su, Samuel Williams, Leonid Oliker, Stephane Ethier, and Katherine Yelick. Optimization of parallel particle-to-grid interpolation on leading multicore platforms. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1915–1922, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Meraji:2012:OTP

[MT12] Sina Meraji and Carl Tropper. Optimizing techniques for parallel digital logic simulation. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1135–1146, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Maurer:2015:CBF

[MT15] Alexandre Maurer and Sebastien Tixeuil. Containing Byzantine failures with control zones. IEEE Transactions on Parallel and Distributed Systems, 26(2):362– 370, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06748070-
abs.html.

Mencagli:2017:PCP

[MTDD17] Gabriele Mencagli, Massimo Torquati, Marco Danelutto, and Tiziano De Matteis. Parallel continuous preference queries over out-of-order and bursty data streams. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2608–2624, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07873332-
abs.html.

Mayer:2018:GTA

[MTMR18] Christian Mayer, Muhammad Adnan Tariq, Ruben Mayer, and Kurt Rothermel. GrapH: Traffic-aware graph processing. IEEE Transactions on Parallel and Distributed Systems, 29(6): 1289–1302, June 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/06/08263157-
abs.html.

Mao:2011:EEO

[MTX^{+}11] Xufei Mao, Shaojie Tang, Xiaohua Xu, Xiang-Yang Li, and Huadong Ma. Energyefficient opportunistic routing in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(11):1934–1942, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moretti:2012:FSG

[MTY^{+}12] Christopher Moretti, Andrew Thrasher, Li Yu, Michael Olson, Scott Emrich, and Douglas Thain. A framework for scalable genome assembly on clusters, clouds, and grids. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2189–2197, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Murray:2012:GAR

[Mur12] Lawrence Murray. GPU acceleration of Runge–Kutta integrators. IEEE Transactions on Parallel and Distributed Systems, 23(1):94– 101, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Miloslavov:2012:SDF

[MV12] Adelin Miloslavov and Malathi Veeraraghavan. Sensor data fusion algorithms for vehicular cyber-physical systems. IEEE Transactions on Parallel and Distributed Systems, 23(9):1762–1774, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mittal:2016:SAA

[MV16a] Sparsh Mittal and Jeffrey S.

Vetter. A survey of architectural approaches for data compression in cache and main memory systems. IEEE Transactions on Parallel and Distributed Systems, 27(5):1524–1536, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07110612-
abs.html.

Mittal:2016:SST

[MV16b] Sparsh Mittal and Jeffrey S.

Vetter. A survey of software techniques for using nonvolatile memories for storage and main memory systems. IEEE Transactions on Parallel and Distributed Systems, 27(5):1537–1550, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07120149-
abs.html.

Mittal:2016:STA

[MV16c] Sparsh Mittal and Jeffrey S. Vetter. A survey of techniques for architecting DRAM caches. IEEE Transactions on Parallel and Distributed Systems, 27(6):1852–1863, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07181712-
abs.html.

Mittal:2016:STM

[MV16d] Sparsh Mittal and Jeffrey S. Vetter. A survey of techniques for modeling and improving reliability of computing systems. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1226–1238, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07094277-
abs.html.

Murali:2018:MHJ

[MV18] Prakash Murali and Sathish Vadhiyar. Metascheduling of HPC jobs in dayahead electricity markets. IEEE Transactions on Parallel and Distributed Systems, 29(3):614–627, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8094010/.

Merchant:2018:ERH

[MVC^{+}18] Farhad Merchant, Tarun Vatwani, Anupam Chattopadhyay, Soumyendu Raha, S. K. Nandy, and Ranjani Narayan. Efficient realization of Householder transform through algorithmarchitecture co-design for acceleration of QR factorization. IEEE Transactions on Parallel and Distributed Systems, 29(8):1707–1720, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08283815-
abs.html.

Mittal:2015:SAA

[MVL15] Sparsh Mittal, Jeffrey S.

Vetter, and Dong Li. A survey of architectural approaches for managing embedded DRAM and nonvolatile on-chip caches. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1524–1537, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06816046-
abs.html.

Moreno-Vozmediano:2011:MDC

[MVML11] Rafael Moreno-Vozmediano, Ruben S. Montero, and Ignacio M. Llorente. Multicloud deployment of computing clusters for loosely coupled MTC applications. IEEE Transactions on Parallel and Distributed Systems, 22(6):924–930, June 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2014:NFC

[MWJ^{+}14] Sheng Ma, Zhiying Wang, Natalie Enright Jerger, Li Shen, and Nong Xiao. Novel flow control for fully adaptive routing in cache-coherent NoCs. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2397–2407, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06549103-
abs.html.

Mao:2016:EWC

[MWJ16] Bo Mao, Suzhen Wu, and Hong Jiang. Exploiting workload characteristics and service diversity to improve the availability of cloud storage systems. IEEE Transactions on Parallel and Distributed Systems, 27(7): 2010–2021, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07234928-
abs.html.

Mi:2013:TFG

[MWZ^{+}13] Haibo Mi, Huaimin Wang, Yangfan Zhou, Michael Rung-Tsong Lyu, and Hua Cai. Toward fine-grained, unsupervised, scalable performance diagnosis for production cloud computing systems. IEEE Transactions on Parallel and Distributed Systems, 24(6):1245–1255, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2014:TTB

[MWZ^{+}14] Yan Ma, Lizhe Wang, Albert Y. Zomaya, Dan Chen, and Rajiv Ranjan. Task-tree based large-scale mosaicking for massive remote sensed imageries with dynamic DAG scheduling. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2126–2137, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mao:2014:NPB

[MWZX14] Yu Mao, Jiguang Wan, Yifeng Zhu, and Changsheng Xie. A new paritybased migration method to expand RAID-5. IEEE Transactions on Parallel and Distributed Systems, 25(8): 1945–1954, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Moh:2011:CDB

[MY11] Sangman Moh and Chansu Yu. A cooperative diversitybased robust MAC protocol in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 22(3):353–363, March 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mollah:2018:RCM

[MYPL18] Md Atiqul Mollah, Xin Yuan, Scott Pakin, and Michael Lang. Rapid calculation of max-min fair rates for multi-commodity flows in fat-tree networks. IEEE Transactions on Parallel and Distributed Systems, 29(1): 156–168, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/01/08017578-
                                        abs.html.

Maroulis:2019:HEE

[MZK19] S. Maroulis, N. Zacheilas, and V. Kalogeraki. A holistic energy-efficient real-time scheduler for mixed stream and batch processing workloads. IEEE Transactions on Parallel and Distributed Systems, 30(12):2624–2635, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2019:BSB

[MZL^{+}19] W. Ma, Y. Zhu, C. Li, M. Guo, and Y. Bao. BiloKey: a scalable bi-index locality-aware in-memory key–value store. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1528–1540, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ma:2019:COS

[MZLT19] H. Ma, H. Zhu, K. Li, and W. Tang. Collaborative optimization of service composition for data-intensive applications in a hybrid cloud. IEEE Transactions on Parallel and Distributed Systems, 30(5):1022–1035, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nogueira:2017:ESM

[NCB17] Andre Nogueira, Antonio Casimiro, and Alysson Bessani. Elastic state machine replication. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2486–2499, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07885120-
abs.html.

Nardelli:2019:EOP

[NCGP19] M. Nardelli, V. Cardellini, V. Grassi, and F. L. Presti. Efficient operator placement for distributed data stream processing applications. IEEE Transactions on Parallel and Distributed Systems, 30(8):1753–1767, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ni:2014:HCD

[NCKL14] Lionel Ni, Lei Chen, Lei Kang, and Siyuan Liu. How to conduct distributed incomplete pattern matching. IEEE Transactions on Parallel and Distributed Systems, 25(4):982–992, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nicolae:2017:LAO

[NCM^{+}17] Bogdan Nicolae, Carlos H. A. Costa, Claudia Misale, Kostas Katrinis, and Yoonho Park. Leveraging adaptive I/O to optimize collective data shuffling patterns for big data analytics. IEEE Transactions on Parallel and Distributed Systems, 28(6):1663–1674, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07740885-
abs.html.

Nachiondo:2010:BMS

[NFD10] Teresa Nachiondo, Jose Flich, and Jose Duato. Buffer management strategies to reduce HoL blocking. IEEE Transactions on Parallel and Distributed Systems, 21(6): 739–753, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nishiyama:2014:TPB

[NFFK14] Hiroki Nishiyama, Desmond Fomo, Zubair Md. Fadlullah, and Nei Kato. Traffic pattern-based content leakage detection for trusted content delivery networks. IEEE Transactions on Parallel and Distributed Systems, 25(2): 301–309, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nikitopoulos:2019:MPT

[NGJ^{+}19] K. Nikitopoulos, G. Georgis, C. Jayawardena, D. Chatzipanagiotis, and R. Tafazolli. Massively parallel tree search for high-dimensional sphere decoders. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2309–2325, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nouri:2017:EHM

[NHN17] Sajjad Nouri, Waqar Hussain, and Jari Nurmi. Evaluation of a heterogeneous multicore architecture by design and test of an OFDM receiver. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3171–3187, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07932115-

abs.html. See errata [NHN18].

Nouri:2018:EEH

[NHN18] Sajjad Nouri, Waqar Hussain, and Jari Nurmi. Errata to “Evaluation of a Heterogeneous Multicore Architecture by Design and Test of an OFDM Receiver”. IEEE Transactions on Parallel and Distributed Systems, 29(3):719, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//ieeexplore.ieee.org/
document/8290633/. See

[NHN17].

Nae:2011:DRP

[NIP11] Vlad Nae, Alexandru Iosup, and Radu Prodan. Dynamic resource provisioning in massively multiplayer online games. IEEE Transactions on Parallel and Distributed Systems, 22(3):380– 395, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Niu:2011:ARC

[NL11] Di Niu and Baochun Li.

Analyzing the resiliencecomplexity tradeoff of network coding in dynamic P2P networks. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1842–1850, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nisar:2012:DBM

[NLC12] Arifa Nisar, Wei-Keng Liao, and Alok Choudhary. Delegationbased I/O mechanism for high performance computing systems. IEEE Transactions on Parallel and Distributed Systems, 23(2):271– 279, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Niu:2014:EET

[NLGQ14] Jianwei Niu, Chuang Liu, Yuhang Gao, and Meikang Qiu. Energy efficient task assignment with guaranteed probability satisfying timing constraints for embedded systems. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2043–2052, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ning:2015:APB

[NLY15] Huansheng Ning, Hong Liu, and Laurence T. Yang. Aggregated-proof based hierarchical authentication scheme for the Internet of Things. IEEE Transactions on Parallel and Distributed Systems, 26(3):657– 667, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06767153-
abs.html.

Nejad:2015:TGM

[NMG15] Mahyar Movahed Nejad, Lena Mashayekhy, and Daniel Grosu. Truthful greedy mechanisms for dynamic virtual machine provisioning and allocation in clouds. IEEE Transactions on Parallel and Distributed Systems, 26(2):594– 603, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06748071-
abs.html.

Ni:2014:FGL

[NML^{+}14] Lionel M. Ni, Zhong Ming, Yunhuai Liu, Yuhong Feng, Rui Mao, Kezhong Lu, and Dian Zhang. Fine-grained localization for multiple transceiver-free objects by using RF-based technologies.

IEEE Transactions on Parallel and Distributed Systems, 25(6):1464–1475, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nishida:2010:GCA

[NN10] Hiroshi Nishida and Thinh Nguyen. A global contribution approach to maintain fairness in P2P networks. IEEE Transactions on Parallel and Distributed Systems, 21(6):812–826, June 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nishida:2013:OCS

[NN13] Hiroshi Nishida and Thinh Nguyen. Optimal clientserver assignment for Internet distributed systems. IEEE Transactions on Parallel and Distributed Systems, 24(3):565–575, March 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nguyen:2013:DDR

[NNKL13] Kien Nguyen, Thinh Nguyen, Yevgeniy Kovchegov, and Viet Le. Distributed data replenishment. IEEE Transactions on Parallel and Distributed Systems, 24(2):275– 287, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Nahir:2016:RBL

[NOR16] Amir Nahir, Ariel Orda, and Danny Raz. Replicationbased load balancing. IEEE Transactions on Parallel and Distributed Systems, 27(2): 494–507, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07031965-
abs.html.

Novak:2015:LOD

[Nov15] Roman Novak. Loop optimization for divergence reduction on GPUs with SIMT architecture. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1633–1642, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06815977-
abs.html.

Nguyen:2015:DDS

[NSH15] Hung Khanh Nguyen, Ju Bin Song, and Zhu Han. Distributed demand side management with energy storage in smart grid. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3346–3357, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/

trans/td/2015/12/06963474-
abs.html.

Nadal-Serrano:2016:PSC

[NSLV16] Jose M. Nadal-Serrano and Marisa Lopez-Vallejo. A performance study of CUDA UVM versus manual optimizations in a real-world setup: Application to a Monte Carlo wave-particle event-based interaction model. IEEE Transactions on Parallel and Distributed Systems, 27(6):1579–1588, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07175058-
abs.html.

Noor:2016:CSR

[NSY^{+}16] Talal H. Noor, Quan Z.

Sheng, Lina Yao, Schahram Dustdar, and Anne H. H. Ngu. CloudArmor: Supporting reputation-based trust management for cloud services. IEEE Transactions on Parallel and Distributed Systems, 27(2):367– 380, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07054490-
abs.html.

Nelson:2016:RRA

[NTA^{+}16] Chad Nelson, Kevin R.

Townsend, Osama G. Attia, Phillip H. Jones, and Joseph Zambreno. RAMPS: A reconfigurable architecture for minimal perfect sequencing. IEEE Transactions on Parallel and Distributed Systems, 27(10): 3029–3043, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07368199-
abs.html.

Nguyen:2019:POP

[NTDZ19] Bao Nguyen, Hua Tan, Kei Davis, and Xuechen Zhang. Persistent octrees for parallel mesh refinement through non-volatile byte-addressable memory. IEEE Transactions on Parallel and Distributed Systems, 30(3):677– 691, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08451966-
abs.html.

Nishiyama:2015:DRF

[NTK^{+}15] Hiroki Nishiyama, Asato Takahashi, Nei Kato, Katsuya Nakahira, and Takatoshi Sugiyama. Dynamic replication and forwarding control based on node surroundings in cooperative delaytolerant networks. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2711–2719, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06908007.
pdf.

Nishiyama:2015:COS

[NTKK15] Hiroki Nishiyama, Ko Togashi, Yuichi Kawamoto, and Nei Kato. A cooperative ONU sleep method for reducing latency and energy consumption of STA in Smart-FiWi networks. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2621–2629, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06915901.
pdf.

Nukarapu:2011:DRD

[NTWL11] Dharma Teja Nukarapu, Bin Tang, Liqiang Wang, and Shiyong Lu. Data replication in data intensive scientific applications with performance guarantee. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1299–1306, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Navarro:2018:CNL

[NVBH18] Cristobal A. Navarro, Matthieu Vernier, Benjamin Bustos, and Nancy Hitschfeld. Competitiveness of a non-linear block-space GPU thread map for simplex domains. IEEE Transactions on Parallel and Distributed Systems, 29(12):2728–2741, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08392762-
abs.html.

Nikolaou:2016:PCP

[NVS16] Stavros Nikolaou, Robbert Van Renesse, and Nicolas Schiper. Proactive cache placement on cooperative client caches for online social networks. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1174–1186, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07091936-
abs.html.

Netzer:1995:NSC

[NX95] Robert H. B. Netzer and Jian Xu. Necessary and sufficient conditions for consistent global snapshots. IEEE Transactions on Parallel and Distributed Systems, 6(2): 165–169, February 1995. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/tpds/
td1995/l0165abs.htm.

Niu:2016:BSE

[NZM^{+}16] Shuangcheng Niu, Jidong Zhai, Xiaosong Ma, Xiongchao Tang, Wenguang Chen, and Weimin Zheng. Building semi-elastic virtual clusters for cost-effective HPC cloud resource provisioning. IEEE Transactions on Parallel and Distributed Systems, 27(7): 1915–1928, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07239625-
abs.html.

Ni:2014:CIC

[NZWL14] Lionel M. Ni, Qian Zhang, Kaishun Wu, and Haochao Li. CUTS: Improving channel utilization in both time and spatial domain in WLANs. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1413–1423, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ouyang:2016:PST

[OKT^{+}16] Robin Wentao Ouyang, Lance M. Kaplan, Alice Toniolo, Mani Srivastava, and Timothy J. Norman. Parallel and streaming truth discovery in large-scale quantitative crowdsourcing. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2984–2997, October 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/10/07373643-
                                        abs.html.

Osterweil:2014:VKT

[OMMZ14] Eric Osterweil, Dan Massey, Danny McPherson, and Lixia Zhang. Verifying keys through publicity and communities of trust: Quantifying off-axis corroboration. IEEE Transactions on Parallel and Distributed Systems, 25(2):283–291, February 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Okuyama:2014:AOB

[OOA^{+}14] Tomohiro Okuyama, Masao Okita, Takeshi Abe, Yoshiyuki Asai, Hiroaki Kitano, Taishin Nomura, and Kenichi Hagihara. Accelerating ODEbased simulation of general and heterogeneous biophysical models using a GPU. IEEE Transactions on Parallel and Distributed Systems, 25(8):1966–1975, August 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Oh:2019:HPT

[OPJ^{+}19] S. Oh, N. Park, J. Jang, L. Sael, and U. Kang. Highperformance Tucker factorization on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2237–2248, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Oxley:2015:MER

[OPM^{+}15] Mark A. Oxley, Sudeep Pasricha, Anthony A. Maciejewski, Howard Jay Siegel, Jonathan Apodaca, Dalton Young, Luis Briceno, Jay Smith, Shirish Bahirat, Bhavesh Khemka, Adrian Ramirez, and Yong Zou. Makespan and energy robust stochastic static resource allocation of a bagof-tasks to a heterogeneous computing system. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2791–2805, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06922558.
pdf.

Oruc:2017:SRC

[Oru17] A. Yavuz Oruc. A selfrouting on-chip network. IEEE Transactions on Parallel and Distributed Systems, 28(5):1229–1239, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2017/05/07723924-
                                        abs.html.

Ozkural:2011:PFI

[OUA11] Eray Ozkural, Bora Ucar, and Cevdet Aykanat. Parallel frequent item set mining with selective item replication. IEEE Transactions on Parallel and Distributed Systems, 22(10):1632–1640, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ozdal:2019:IEP

[Ozd19] M. M. Ozdal. Improving efficiency of parallel vertexcentric algorithms for irregular graphs. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2265–2282, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ortega-Zamorano:2016:FHA

[OZMC^{+}16] Francisco Ortega-Zamorano, Marcelo A. Montemurro, Sergio Alejandro Cannas, Jose M. Jerez, and Leonardo Franco. FPGA hardware acceleration of Monte Carlo simulations for the Ising model. IEEE Transactions on Parallel and Distributed Systems, 27(9):2618–2627, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/

trans/td/2016/09/07347434-
abs.html.

Paterna:2013:AAE

[PAB13] Francesco Paterna, Andrea Acquaviva, and Luca Benini. Aging-aware energy-efficient workload allocation for mobile multimedia platforms. IEEE Transactions on Parallel and Distributed Systems, 24(8):1489–1499, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Panda:2014:GAM

[Pan14] Dhabaleswar K. Panda.

GPU-aware MPI on RDMAenabled clusters: Design, implementation and evaluation. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2595–2605, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06587715-
abs.html.

Parashar:2018:SJ

[Par18] Manish Parashar. State of the journal. IEEE Transactions on Parallel and Distributed Systems, 29(1):1, January 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08173510.
pdf.

Parashar:2019:ENb

[Par19a] M. Parashar. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 30(8):1687–1689, August 2019. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parashar:2019:ENI

[Par19b] M. Parashar. Editor’s note: IEEE Transactions on Parallel and Distributed Systems (TPDS) reproducibility initiative, June 2019. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1690, August 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Parashar:2019:ENa

[Par19c] Manish Parashar. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 30(1):1, January 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/01/08573098.
                                        pdf.

Panta:2012:MES

[PB12] Rajesh Krishna Panta and Saurabh Bagchi. Mitigating the effects of software component shifts for incremental reprogramming of wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1882–1894, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peccerillo:2019:PPH

[PB19] Biagio Peccerillo and Sandro Bartolini. PHAST — a portable high-level modern C++ programming library for GPUs and multicores. IEEE Transactions on Parallel and Distributed Systems, 30(1):174– 189, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08410014-
abs.html.

Pell:2013:FDW

[PBD^{+}13] Oliver Pell, Jacob Bower, Robert Dimond, Oskar Mencer, and Michael J. Flynn. Finite-difference wave propagation modeling on special-purpose dataflow machines. IEEE Transactions on Parallel and Distributed Systems, 24(5):906– 915, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pal:2016:MPS

[PCFP16] Ranjan Pal, Charalampos Chelmis, Marc Frincu, and Viktor Prasanna. MATCH for the prosumer smart grid the algorithmics of realtime power balance. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3532–3546, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07437455-
abs.html.

Pagani:2015:EEM

[PCL15] Santiago Pagani, Jian-Jia Chen, and Minming Li. Energy efficiency on multicore architectures with multiple voltage islands. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1608–1621, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06814918-
abs.html.

Psaras:2014:NCM

[PCP14] Ioannis Psaras, Wei Koong Chai, and George Pavlou. In-network cache management and resource allocation for information-centric networks. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2920–2931, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2014/11/06684140-
abs.html.

Patel:2014:CSP

[PD14] Rajesh Patel and Ajoy K.

Datta. CPU scheduling for power/energy management on multicore processors using cache miss and context switch data. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1190–1199, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2013:MOI

[PDFJ13] Pangun Park, Piergiuseppe Di Marco, Carlo Fischione, and Karl Henrik Johansson. Modeling and optimization of the IEEE 802.15.4 protocol for reliable and timely communications. IEEE Transactions on Parallel and Distributed Systems, 24(3): 550–564, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pezoa:2010:MSR

[PDH10] Jorge E. Pezoa, Sagar Dhakal, and Majeed M. Hayat. Maximizing service reliability in distributed computing systems with random node failures: Theory and implementation. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1531–1544, October 2010.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2012:EMW

[PF12] Alfred J. Park and Richard M. Fujimoto. Efficient master/worker parallel discrete event simulation on metacomputing systems. IEEE Transactions on Parallel and Distributed Systems, 23(5): 873–880, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Palopoli:2016:ASP

[PFAF16] Luigi Palopoli, Daniele Fontanelli, Luca Abeni, and Bernardo Villalba Frias. An analytical solution for probabilistic guarantees of reservation based soft real-time systems. IEEE Transactions on Parallel and Distributed Systems, 27(3):640– 653, March 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/03/07070759-
                                        abs.html.

Papadakis:2013:IIT

[PFMR13] Harris Papadakis, Paraskevi Fragopoulou, Evangelos P. Markatos, and Mema Roussopoulos. ITA: Innocuous topology awareness for unstructured P2P networks. IEEE Transactions on Parallel and Distributed Systems, 24(8):1589–1601, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Perez:2016:EDC

[PG16] Hector Perez and J. Javier Gutierrez. Enabling datacentric distribution technology for partitioned embedded systems. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3186–3198, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07412779-
abs.html.

Poshtkohi:2019:OMS

[PGGS19] Alireza Poshtkohi, M. B.

Ghaznavi-Ghoushchi, and Kamyar Saghafi. Optimistic modeling and simulation of complex hardware platforms and embedded systems on many-core HPC clusters. IEEE Transactions on Parallel and Distributed Systems, 30(2):428– 444, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08419330-
abs.html.

Persico:2017:FAB

[PGP^{+}17] Valerio Persico, Domenico Grimaldi, Antonio Pescape, Alessandro Salvi, and Stefania Santini. A fuzzy approach based on heterogeneous metrics for scaling out public clouds. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2117–2130, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/08/07814313-
                                        abs.html.

Park:2011:PHP

[PH11] Sang-Min Park and Marty A. Humphrey. Predictable highperformance computing using feedback control and admission control. IEEE Transactions on Parallel and Distributed Systems, 22(3):396– 411, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pezoa:2012:PRN

[PH12] Jorge E. Pezoa and Majeed M. Hayat. Performance and reliability of non-Markovian heterogeneous distributed computing systems. IEEE Transactions on Parallel and Distributed Systems, 23(7):1288–1301, July 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peng:2018:FSF

[PH18] Yaqiong Peng and Zhiyu Hao. FA-Stack: A fast arraybased stack with wait-free progress guarantee. IEEE Transactions on Parallel and Distributed Systems, 29(4): 843–857, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08097018-
abs.html.

Palencia:2017:RTA

[PHGR17] J. Carlos Palencia, Michael Gonzalez Harbour, J. Javier Gutierrez, and Juan M. Rivas. Response-time analysis in hierarchically-scheduled time-partitioned distributed systems. IEEE Transactions on Parallel and Distributed Systems, 28(7): 2017–2030, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07792746-
abs.html.

Pei:2019:EES

[PHXL19] J. Pei, P. Hong, K. Xue, and D. Li. Efficiently embedding service function chains with dynamic virtual network function placement in geo-distributed cloud system. IEEE Transactions on Parallel and Distributed Systems, 30(10):2179–2192, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ponomarev:2014:PDE

[PJAGW14] Dmitry Ponomarev, Deepak Jagtap, Nael Abu-Ghazaleh, and Jingjing Wang. Parallel discrete event simulation for multi-core systems: Analysis and optimization. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1574–1584, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Panta:2013:PSU

[PJC^{+}13] Rajesh K. Panta, Rittwik Jana, Fan Cheng, Yih-Farn Robin Chen, and Vinay A. Vaishampayan. Phoenix: Storage using an autonomous mobile infrastructure. IEEE Transactions on Parallel and Distributed Systems, 24(9):1863–1873, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Passarella:2011:MDS

[PKCB11] Andrea Passarella, Mohan Kumar, Marco Conti, and Eleonora Borgia. Minimumdelay service provisioning in opportunistic networks. IEEE Transactions on Parallel and Distributed Systems, 22(8):1267–1275, August 2011. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pelechrinis:2014:TOC

[PKG14] Konstantinos Pelechrinis, Prashant Krishanmurthy, and Christos Gkantsidis. Trustworthy operations in cellular networks: The case of PF scheduler. IEEE Transactions on Parallel and Distributed Systems, 25(2): 292–300, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Papavassiliou:2012:GEI

[PKL^{+}12] Symeon Papavassiliou, Nei Kato, Yunhao Liu, Cheng-Zhong Xu, and Xinbing Wang. Guest Editors’ introduction: Special issue on Cyber-Physical Systems (CPS). IEEE Transactions on Parallel and Distributed Systems, 23(9):1569–1571, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pantazopoulos:2014:DPA

[PKS14] Panagiotis Pantazopoulos, Merkouris Karaliopoulos, and Ioannis Stavrakakis. Distributed placement of autonomic Internet services. IEEE Transactions on Parallel and Distributed Systems, 25(7):1702–1712, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Poluri:2016:SRN

[PL16] Pavan Poluri and Ahmed Louri. Shield: A reliable network-on-chip router architecture for chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 27(10): 3058–3070, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/10/07390298-
                                        abs.html.

Pan:2019:CTE

[PLG19] P. Pan, C. Li, and M. Guo. CongraPlus: Towards efficient processing of concurrent graph queries on NUMA machines. IEEE Transactions on Parallel and Distributed Systems, 30(9): 1990–2002, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peng:2014:BMD

[PLZW14] Wei Peng, Feng Li, Xukai Zou, and Jie Wu. Behavioral malware detection in delay tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 25(1): 53–63, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pei:2013:SSR

[PM13] Yuanteng Pei and Matt W.

Mutka. STARS: Static relays for remote sensing in multirobot real-time search and monitoring. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2079–2089, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Prabhavat:2011:EDC

[PNAK11] Sumet Prabhavat, Hiroki Nishiyama, Nirwan Ansari, and Nei Kato. Effective delay-controlled load distribution over multipath networks. IEEE Transactions on Parallel and Distributed Systems, 22(10):1730–1741, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Plankensteiner:2012:MSD

[PP12] Kassian Plankensteiner and Radu Prodan. Meeting soft deadlines in scientific workflows using resubmission impact. IEEE Transactions on Parallel and Distributed Systems, 23(5):890– 901, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2010:IPN

[PPR10] Karam Park, Joon-Sang Park, and Won W. Ro. On improving parallelized network coding with dynamic partitioning. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1547–1560, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Pagani:2017:EEC

[PPS^{+}17] Santiago Pagani, Anuj Pathania, Muhammad Shafique, Jian-Jia Chen, and Jorg Henkel. Energy efficiency for clustered heterogeneous multicores. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1315–1330, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/05/07728098-
                                        abs.html.

Pascual:2019:OEP

[PR19] F. Pascual and K. Rzadca. Optimizing egalitarian performance when colocating tasks with types for cloud data center resource management. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2523–2535, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Peluso:2016:GGM

[PRR^{+}16] Sebastiano Peluso, Pedro Ruivo, Paolo Romano, Francesco Quaglia, and Luis Rodrigues. GMU: Genuine multiversion update-serializable partial data replication. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2911–2925, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07362215-
abs.html.

Park:2011:HCA

[PRS^{+}11] Song Jun Park, James A.

Ross, Dale R. Shires, David A. Richie, Brian J. Henz, and Lam H. Nguyen. Hybrid core acceleration of UWB SIRE radar signal processing. IEEE Transactions on Parallel and Distributed Systems, 22(1):46– 57, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Prades:2019:GJM

[PS19] J. Prades and F. Silla. GPUjob migration: The rCUDA case. IEEE Transactions on Parallel and Distributed Systems, 30(12):2718–2729, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2011:DPE

[PSL^{+}11] In Kyu Park, Nitin Singhal, Man Hee Lee, Sungdae Cho, and Chris W. Kim. Design and performance evaluation of image processing algorithms on GPUs. IEEE Transactions on Parallel and Distributed Systems, 22(1): 91–104, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Palanisamy:2015:CER

[PSL15] Balaji Palanisamy, Aameek Singh, and Ling Liu. Costeffective resource provisioning for MapReduce in a cloud. IEEE Transactions on Parallel and Distributed Systems, 26(5):1265–1279, May 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06805615-
abs.html.

Pratap:2018:DRC

[PSMD18] Ajay Pratap, Rishabh Singhal, Rajiv Misra, and Sajal K. Das. Distributed randomized k-clustering based PCID assignment for ultradense femtocellular networks. IEEE Transactions on Parallel and Distributed Systems, 29(6):1247–1260, June 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/06/08276600-
abs.html.

Pong:2011:HRP

[PT11] Fong Pong and Nian-Feng Tzeng. HaRP: Rapid packet classification via hashing round-down prefixes. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1105–1119, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2015:PEO

[PT15] Pangun Park and Claire Tomlin. Performance evaluation and optimization of communication infrastructure for the next generation air transportation system. IEEE Transactions on Parallel and Distributed Systems, 26(4):1106–1116, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06799224-
abs.html.

Pellegrini:2015:ASM

[PVQ15] Alessandro Pellegrini, Roberto Vitali, and Francesco Quaglia. Autonomic state management for optimistic simulation platforms. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1560–1569, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/

trans/td/2015/06/06815695-
abs.html.

Pathan:2018:SPR

[PVS18] Risat Pathan, Petros Voudouris, and Per Stenstrom. Scheduling parallel real-time recurrent tasks on multicore platforms. IEEE Transactions on Parallel and Distributed Systems, 29(4):915– 928, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08119822-
abs.html.

Peng:2016:RTE

[PWJ16] Yaqiong Peng, Song Wu, and Hai Jin. Robinhood: Towards efficient work-stealing in virtualized environments. IEEE Transactions on Parallel and Distributed Systems, 27(8):2363–2376, August 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07300465-
abs.html.

Panadero:2018:PMA

[PWRL18] Javier Panadero, Alvaro Wong, Dolores Rexachs, and Emilio Luque. P3S: A methodology to analyze and predict application scalability. IEEE Transactions on Parallel and Distributed Systems, 29(3):642–658, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8068239/.

Pham:2017:FDD

[PWT^{+}17] Cuong Pham, Long Wang, Byung Chul Tak, Salman Baset, Chunqiang Tang, Zbigniew Kalbarczyk, and Ravishankar K. Iyer. Failure diagnosis for distributed systems using targeted fault injection. IEEE Transactions on Parallel and Distributed Systems, 28(2):503– 516, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07484300-
abs.html.

Peng:2019:FWF

[PYH19] Y. Peng, X. Yun, and Z. Hao. Fast wait-free construction for pool-like objects with weakened internal order: Stacks as an example. IEEE Transactions on Parallel and Distributed Systems, 30(7):1596–1612, July 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Park:2016:SSG

[PYHY16] Hyunchan Park, Seehwan Yoo, Cheol-Ho Hong, and Chuck Yoo. Storage SLA guarantee with novel SSD I/O scheduler in virtualized data centers. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2422–2434, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07303970-
abs.html.

Qian:2015:HTE

[QCZ^{+}15] Shiyou Qian, Jian Cao, Yanmin Zhu, Minglu Li, and Jie Wang. H-tree: An efficient index structure for event matching in content-based publish/subscribe systems. IEEE Transactions on Parallel and Distributed Systems, 26(6):1622–1632, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06814863-
abs.html.

Quan:2014:HAM

[QF14] Gang Quan and Ming Fan. Harmonic-aware multi-core scheduling for fixed-priority real-time systems. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1476–1488, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qu:2015:SCD

[QFZZ15] Guannan Qu, Zhiyi Fang, Jianfei Zhang, and Si-Qing Zheng. Switch-centric data center network structures based on hypergraphs and combinatorial block designs. IEEE Transactions on Parallel and Distributed Systems, 26(4):1154–1164, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06802363-
abs.html.

Quislant:2013:HSD

[QGPZ13] Ricardo Quislant, Eladio Gutierrez, Oscar Plata, and Emilio L. Zapata. Hardware signature designs to deal with asymmetry in transactional data sets. IEEE Transactions on Parallel and Distributed Systems, 24(3): 506–519, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Quislant:2017:LIB

[QGZP17] Ricardo Quislant, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Lazy irrevocability for best-effort transactional memory systems. IEEE Transactions on Parallel and Distributed Systems, 28(7):1919–1932, July 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07782383-
abs.html.

Quan:2016:FDW

[QJ16] Tran Minh Quan and Won-Ki Jeong. A fast discrete wavelet transform using hybrid parallelism on GPUs. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3088–3100, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/11/07422119-
                                        abs.html.

Qin:2013:DAU

[QLC13] Zhengrui Qin, Qun Li, and Mooi-Choo Chuah. Defending against unidentifiable attacks in electric power grids. IEEE Transactions on Parallel and Distributed Systems, 24(10):1961–1971, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiao:2014:FBF

[QLC14] Yan Qiao, Tao Li, and Shigang Chen. Fast Bloom filters and their generalization. IEEE Transactions on Parallel and Distributed Systems, 25(1):93–103, January 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See comment [RCM16].

Qian:2013:ASC

[QLNN13] Chen Qian, Yunhuai Liu, Raymond Hoilun Ngan, and Lionel M. Ni. ASAP: Scalable collision arbitration for large RFID systems. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1277–1288, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qian:2011:CEL

[QNLN11] Chen Qian, Hoilun Ngan, Yunhao Liu, and Lionel M. Ni. Cardinality estimation for large-scale RFID systems. IEEE Transactions on Parallel and Distributed Systems, 22(9):1441–1454, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qiu:2016:ERP

[QP16a] Zhan Qiu and Juan F.

Perez. Evaluating replication for parallel jobs: An efficient approach. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2288–2302, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07313012-
abs.html.

Qu:2016:FOS

[QP16b] Yun R. Qu and Viktor K.

Prasanna. Fast online set intersection for network processing on FPGA. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3214–3225, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07425232-
abs.html.

Qu:2016:HPD

[QP16c] Yun R. Qu and Viktor K. Prasanna. High-performance and dynamically updatable packet classification engine on FPGA. IEEE Transactions on Parallel and Distributed Systems, 27(1):197– 209, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07004892-
abs.html.

Qiu:2017:CLT

[QPB^{+}17] Zhan Qiu, Juan F. Perez, Robert Birke, Lydia Chen, and Peter G. Harrison. Cutting latency tail: Analyzing and validating replication without canceling. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3128–3141, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07932099-
abs.html.

Qian:2014:MCM

[QTC^{+}14] Zhenzhi Qian, Xiaohua Tian, Xi Chen, Wentao Huang, and Xinbing Wang. Multicast capacity in MANET with infrastructure support. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1808–1818, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qian:2016:ORH

[QZG^{+}16] Junyan Qian, Zhide Zhou, Tianlong Gu, Lingzhong Zhao, and Liang Chang. Optimal reconfiguration of highperformance VLSI subarrays with network flow. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3575–3587, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07429772-
abs.html.

Qian:2014:BFB

[QZW14] Jiangbo Qian, Qiang Zhu, and Yongli Wang. Bloom filter based associative deletion. IEEE Transactions on Parallel and Distributed Systems, 25(8):1986–1998, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Qin:2016:PED

[QZZ^{+}16] Jun Qin, Hongzi Zhu, Yanmin Zhu, Li Lu, Guangtao Xue, and Minglu Li. POST: Exploiting dynamic sociality for mobile advertising in vehicular networks. IEEE Transactions on Parallel and Distributed Systems, 27(6):1770–1782, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07192632-
abs.html.

Ros:2010:DCP

[RAG10] Alberto Ros, Manuel E. Acacio, and Jose M. Garcia. A direct coherence protocol for many-core chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 21(12): 1779–1792, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rao:2014:APA

[Rao14] Kotagiri Rao. Authorized public auditing of dynamic big data storage on cloud with efficient verifiable finegrained updates. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2234–2244, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2014/09/06574863-
abs.html.

Rahman:2017:ASD

[RAS17] Mohammad Ashiqur Rahman and Ehab Al-Shaer. Automated synthesis of distributed network access controls: A formal framework with refinement. IEEE Transactions on Parallel and Distributed Systems, 28(2): 416–430, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07500078-
abs.html.

Rustico:2014:AMG

[RBH^{+}14] Eugenio Rustico, Giuseppe Bilotta, Alexis Herault, Ciro Del Negro, and Giovanni Gallo. Advances in multi-GPU smoothed particle hydrodynamics simulations. IEEE Transactions on Parallel and Distributed Systems, 25(1):43– 52, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Radenkovic:2015:RAO

[RBM15] Milena Radenkovic, Abderrahim Benslimane, and Derek McAuley. Reputation aware obfuscation for mobile opportunistic networks. IEEE Transactions on Parallel and Distributed Systems, 26(1):230– 240, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06636891-
                                        abs.html.

Radhakrishnan:2011:DCS

[RBSS11] Sridhar Radhakrishnan, Shankar M. Banik, Venkatesh Sarangan, and Chandra N. Sekharan. Delay constrained subtree homeomorphism problem with applications. IEEE Transactions on Parallel and Distributed Systems, 22(12): 1978–1985, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2014:DAP

[RCC^{+}14] Zhu Ren, Peng Cheng, Jiming Chen, David K. Y. Yau, and Youxian Sun. Dynamic activation policies for event capture in rechargeable sensor network. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3124–3134, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/12/06714601-
                                        abs.html.

Rao:2010:ORP

[RCFW10] Weixiong Rao, Lei Chen, Ada Wai-Chee Fu, and Guoren Wang. Optimal resource placement in structured peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 21(7):1011–1026, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Choi:2010:SFS

[rCHG10] Young ri Choi, Chin-Tser Huang, and Mohamed G. Gouda. Stabilization of flood sequencing protocols in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(7): 1042–1055, July 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rafique:2015:COI

[RCK15] Abid Rafique, George A.

Constantinides, and Nachiket Kapre. Communication optimization of iterative sparse matrix-vector multiply on GPUs and FPGAs. IEEE Transactions on Parallel and Distributed Systems, 26 (1):24–34, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06719387-
abs.html.

Reviriego:2016:CFB

[RCM16] P. Reviriego, K. Christensen, and J. A. Maestro. A comment on “Fast Bloom Filters and Their Generalization”. IEEE Transactions on Parallel and Distributed Systems, 27(1):303– 304, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07012061-

abs.html. See [QLC14].

Radojkovic:2013:TAM

[RCV^{+}13] Petar Radojkovic, Vladimir Cakarevic, Javier Verdu, Alex Pajuelo, Francisco J. Cazorla, Mario Nemirovsky, and Mateo Valero. Thread assignment of multithreaded network applications in multicore/multithreaded processors. IEEE Transactions on Parallel and Distributed Systems, 24(12):2513–2525, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ries:2012:TMI

[RDG12] Florian Ries, Tommaso De Marco, and Roberto Guerrieri. Triangular matrix inversion on heterogeneous multicore systems. IEEE Transactions on Parallel and Distributed Systems, 23(1):177– 184, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2014:OLB

[Ren14] Kui Ren. Optimal load balancing and energy cost management for Internet data centers in deregulated electricity markets. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2659–2669, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06594731-
abs.html.

Raicu:2011:GEI

[RFZ11] Ioan Raicu, Ian T. Foster, and Yong Zhao. Guest Editors’ introduction: Special section on many-task computing. IEEE Transactions on Parallel and Distributed Systems, 22(6):897– 898, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rampersaud:2017:SAO

[RG17] Safraz Rampersaud and Daniel Grosu. Sharingaware online virtual machine packing in heterogeneous resource clouds. IEEE Transactions on Parallel and Distributed Systems, 28(7): 2046–2059, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07792170-
abs.html.

Robles-Gomez:2011:DFD

[RGBC11] Antonio Robles-Gomez, Aurelio Bermudez, and Rafael Casado. A deadlockfree dynamic reconfiguration scheme for source routing networks using close up*/ down* graphs. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1641–1652, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rashid:2015:SFS

[RGK15] Md. Mamunur Rashid, Iqbal Gondal, and Joarder Kamruzzaman. Share-frequent sensor patterns mining from wireless sensor network data. IEEE Transactions on Parallel and Distributed Systems, 26(12):3471–3484, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06977946-
abs.html.

Rico-Gallego:2017:MBE

[RGLDM17] Juan-Antonio Rico-Gallego, Alexey L. Lastovetsky, and Juan-Carlos Diaz-Martin. Model-based estimation of the communication cost of hybrid data-parallel applications on heterogeneous clusters. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3215–3228, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07949135-
abs.html.

Rivas:2015:DAE

[RGPH15] Juan M. Rivas, J. Javier Gutierrez, J. Carlos Palencia, and Michael Gonzalez Harbour. Deadline assignment in EDF schedulers for real-time distributed systems. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2671–2684, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06905828.
pdf.

Ranka:2014:MCE

[RGRM14] Sanjay Ranka, Ann Gordon-Ross, and Arslan Munir. Multi-core embedded wireless sensor networks: Architecture and applications. IEEE Transactions on Parallel and Distributed Systems, 25(6):1553–1562, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramos:2016:CLA

[RH16] Sabela Ramos and Torsten Hoefler. Cache line aware algorithm design for cachecoherent architectures. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2824–2837, October 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/10/07378320-
                                        abs.html.

Ren:2011:TAD

[RHDL11] Fengyuan Ren, Tao He, Sajal K. Das, and Chuang Lin. Traffic-aware dynamic routing to alleviate congestion in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(9):1585–1599, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ravindran:2013:TAS

[RHT13] Ravishankar Ravindran, Changcheng Huang, and Krishnaiya Thulasiraman. Topology abstraction service for IP-VPNs. IEEE Transactions on Parallel and Distributed Systems, 24(1):184– 197, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ros:2016:HSD

[RJ16] Alberto Ros and Alexandra Jimborean. A hybrid static-dynamic classification for dual-consistency cache coherence. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3101–3115, November 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07404282-
abs.html.

Ramakrishna:2016:GGC

[RKGS16] Mukund Ramakrishna, Vamsi Krishna Kodati, Paul V. Gratz, and Alexander Sprintson. GCA:Global congestion awareness for load balance in networks-on-chip. IEEE Transactions on Parallel and Distributed Systems, 27(7): 2022–2035, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07254220-
abs.html.

Rottenstreich:2017:MDN

[RKRK17] Ori Rottenstreich, Isaac Keslassy, Yoram Revah, and Aviran Kadosh. Minimizing delay in network function virtualization with shared pipelines. IEEE Transactions on Parallel and Distributed Systems, 28(1):156– 169, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07457328-
abs.html.

Raychoudhury:2014:AES

[RKZC14] Vaskar Raychoudhury, Ajay D.

Kshemkalyani, Daqing Zhang, and Jiannong Cao. Automatic event scheduling in mobile social network communities. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2772–2782, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06714503-
abs.html.

Ros:2017:ESI

[RLSK17] Alberto Ros, Carl Leonardsson, Christos Sakalis, and Stefanos Kaxiras. Efficient self-invalidation/self-downgrade for critical sections with relaxed semantics. IEEE Transactions on Parallel and Distributed Systems, 28(12):3413–3425, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07961235-
abs.html.

Romero-Laorden:2016:APC

[RLVTMG^{+}16] D. Romero-Laorden, J. Villazon-Terrazas, O. Martinez-Graullera, A. Ibanez, M. Parrilla, and M. Santos Penas. Analysis of parallel computing strategies to accelerate ultrasound imaging processes. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3429–3440, December 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07437495-
abs.html.

Ren:2015:DIE

[RLY^{+}15] Yufei Ren, Tan Li, Dantong Yu, Shudong Jin, and Thomas Robertazzi. Design, implementation, and evaluation of a NUMAaware cache for iSCSI storage servers. IEEE Transactions on Parallel and Distributed Systems, 26(2):413– 422, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06767115-
abs.html.

Rasheed:2011:KPS

[RM11] Amar Rasheed and Rabi N.

Mahapatra. Key predistribution schemes for establishing pairwise keys with a mobile sink in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(1):176– 184, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rasheed:2012:TTS

[RM12] Amar Rasheed and Rabi N.

Mahapatra. The three-tier security scheme in wireless sensor networks with mobile sinks. IEEE Transactions on Parallel and Distributed Systems, 23(5):958– 965, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rimal:2017:WSM

[RM17] Bhaskar Prasad Rimal and Martin Maier. Workflow scheduling in multi-tenant cloud computing environments. IEEE Transactions on Parallel and Distributed Systems, 28(1):290– 304, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07457258-
abs.html.

Reguly:2016:AFS

[RMB^{+}16] Istvan Z. Reguly, Gihan R. Mudalige, Carlo Bertolli, Michael B. Giles, Adam Betts, Paul H. J. Kelly, and David Radford. Acceleration of a full-scale industrial CFD application with OP2. IEEE Transactions on Parallel and Distributed Systems, 27(5):1265–1278, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07152942-
abs.html.

Rampersaud:2014:CNE

[RMG14] Safraz Rampersaud, Lena Mashayekhy, and Daniel Grosu. Computing Nash equilibria in bimatrix games: GPU-based parallel support enumeration. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3111–3123, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06747409-
abs.html.

Reguly:2018:LTL

[RMG18] Istvan Z. Reguly, Gihan R.

Mudalige, and Michael B. Giles. Loop tiling in largescale stencil codes at runtime with OPS. IEEE Transactions on Parallel and Distributed Systems, 29(4): 873–886, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08121995-
abs.html.

Rashwand:2016:ACC

[RMM16] Saeed Rashwand, Jelena Misic, and Vojislav B. Misic. Analysis of CSMA/ CA mechanism of IEEE 802.15.6 under non-saturation regime. IEEE Transactions on Parallel and Distributed Systems, 27(5): 1279–1288, May 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/05/07128741-
                                        abs.html.

Rawat:2011:EVP

[RPYO11] Danda B. Rawat, Dimitrie C. Popescu, Gongjun Yan, and Stephan Olariu. Enhancing VANET performance by joint adaptation of transmission power and contention window size. IEEE Transactions on Parallel and Distributed Systems, 22(9): 1528–1535, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2016:TAM

[RQZ^{+}16] Zhen Ren, Xin Qi, Gang Zhou, Haining Wang, and David T. Nguyen. Throughput assurance for multiple body sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27(2): 546–557, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07054555-
                                        abs.html.

Raghav:2015:GAS

[RRM^{+}15] Shivani Raghav, Martino Ruggiero, Andrea Marongiu, Christian Pinto, David Atienza, and Luca Benini.

GPU acceleration for simulating massively parallel many-core platforms. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1336–1349, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803951-
abs.html.

Renda:2012:LBH

[RRS12] M. Elena Renda, Giovanni Resta, and Paolo Santi. Load balancing hashing in geographic hash tables. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1508–1519, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ramachandran:2010:QME

[RS10] Krishna Kumar Ramachandran and Biplab Sikdar. A queuing model for evaluating the transfer latency of peer-to-peer systems. IEEE Transactions on Parallel and Distributed Systems, 21(3): 367–378, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Resta:2012:FRP

[RS12] Giovanni Resta and Paolo Santi. A framework for routing performance analysis in delay tolerant networks with application to noncooperative networks. IEEE Transactions on Parallel and Distributed Systems, 23(1): 2–10, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ruj:2014:DAC

[RSN14] Sushmita Ruj, Milos Stojmenovic, and Amiya Nayak. Decentralized access control with anonymous authentication of data stored in clouds. IEEE Transactions on Parallel and Distributed Systems, 25(2):384– 394, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Reano:2018:INM

[RSNV18] Carlos Reano, Federico Silla, Dimitrios S. Nikolopoulos, and Blesson Varghese. Intranode memory safe GPU coscheduling. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1089–1102, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08219713-
abs.html.

Radojevic:2011:DDH

[RSR11] Ivan Radojevic, Zoran Salcic, and Partha S. Roop. Design of distributed heterogeneous embedded systems in DDFCharts. IEEE Transactions on Parallel and Distributed Systems, 22(2):296– 308, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Raychoudhury:2015:CPC

[RSSC15] Vaskar Raychoudhury, Shikhar Shrivastav, Sandeep Singh Sandha, and Jiannong Cao. CROWD-PAN-360: Crowdsourcing based context-aware panoramic map generation for Smartphone users. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2208–2219, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06871426-
abs.html.

Ren:2017:RSB

[RSW^{+}17] Zujie Ren, Weisong Shi, Jian Wan, Feng Cao, and Jiangbin Lin. Realistic and scalable benchmarking cloud file systems: Practices and lessons from AliCloud. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3272–3285, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07949134-
abs.html.

Ruan:2018:SBO

[RTZ^{+}18] Mingkang Ruan, Thierry Titcheu, Ennan Zhai, Zhenhua Li, Yao Liu, Jinlong E, Yong Cui, and Hong Xu. On the synchronization bottleneck of OpenStack Swift-like cloud storage systems. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2059–2074, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08303732-
                                        abs.html.

Rao:2015:MAD

[RVCT15] Weixiong Rao, Roman Vitenberg, Lei Chen, and Sasu Tarkoma. MTAF: An adaptive design for keywordbased content dissemination on DHT networks. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1071–1084, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06719388-
                                        abs.html.

Russell:2015:NDW

[RVW^{+}15] Alexander Russell, Sudarshan Vasudevan, Bing Wang, Wei Zeng, Xian Chen, and Wei Wei. Neighbor discovery in wireless networks with multipacket reception. IEEE Transactions on Parallel and Distributed Systems, 26(7):1984–1998, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06846356-
abs.html.

Ren:2014:HHM

[RWLL14] Jian Ren, Jie Wu, Yun Li, and Jian Li. Hop-by-hop message authentication and source privacy in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1223–1232, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Rao:2011:OCI

[RX11] Jia Rao and Cheng-Zhong Xu. Online capacity identification of multitier Websites using hardware performance counters. IEEE Transactions on Parallel and Distributed Systems, 22(3): 426–438, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Roy:2012:SDS

[RXD12] Sandip Roy, Mengran Xue, and Sajal K. Das. Security and discoverability of spread dynamics in cyberphysical networks. IEEE Transactions on Parallel and Distributed Systems, 23(9):

1694–1707, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Robertazzi:2014:SSN

[RY14] Thomas G. Robertazzi and Zhongwen Ying. Signature searching in a networked collection of files. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1339–1348, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2010:PNP

[RYLZ10] Kui Ren, Shucheng Yu, Wenjing Lou, and Yanchao Zhang. PEACE: a novel privacy-enhanced yet accountable security framework for metropolitan wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 21(2): 203–215, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Reza:2018:APS

[RZB^{+}18] Tahsin Reza, Aaron Zimmer, Jose Manuel Delgado Blasco, Parwant Ghuman, Tanuj Kr Aasawat, and Matei Ripeanu. Accelerating persistent scatterer pixel selection for InSAR processing. IEEE Transactions on Parallel and Distributed Systems, 29(1):16–30, January 2018.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/07932146-
abs.html.

Ren:2011:EEB

[RZH^{+}11] Fengyuan Ren, Jiao Zhang, Tao He, Chuang Lin, and Sajal K. Das. EBRP: Energybalanced routing protocol for data gathering in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2108–2125, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ren:2013:AAD

[RZW^{+}13] Fengyuan Ren, Jiao Zhang, Yongwei Wu, Tao He, Canfeng Chen, and Chuang Lin. Attribute-aware data aggregation using potential-based dynamic routing in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(5): 881–892, May 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2011:PII

[SA11] Yang Song and Ali Akoglu.

Parallel implementation of the Irregular Terrain Model (ITM) for radio transmission loss prediction using GPU and Cell BE processors. IEEE Transactions on Parallel and Distributed Systems, 22(8):1276–1283, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Selvitopi:2017:RHB

[SAA17] Oguz Selvitopi, Seher Acer, and Cevdet Aykanat. A recursive hypergraph bipartitioning framework for reducing bandwidth and latency costs simultaneously. IEEE Transactions on Parallel and Distributed Systems, 28(2): 345–358, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07485866-
abs.html.

Salamanca:2018:UHT

[SAA18] Juan Salamanca, Jose Nelson Amaral, and Guido Araujo. Using hardwaretransactional-memory support to implement threadlevel speculation. IEEE Transactions on Parallel and Distributed Systems, 29(2): 466–480, February 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/02/08038067-
abs.html.

Seo:2018:ALL

[SAB^{+}18] Sangmin Seo, Abdelhalim Amer, Pavan Balaji, Cyril Bordage, George Bosilca, Alex Brooks, Philip Carns, Adrián Castelló, Damien Genet, Thomas Herault, Shintaro Iwasaki, Prateek Jindal, Laxmikant V. Kalé, Sriram Krishnamoorthy, Jonathan Lifflander, Huiwei Lu, Esteban Meneses, Marc Snir, Yanhua Sun, Kenjiro Taura, and Pete Beckman. Argobots: A lightweight low-level threading and tasking framework. IEEE Transactions on Parallel and Distributed Systems, 29(3):512–526, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8082139/.

Safari:2019:SEA

[SAEH19] S. Safari, M. Ansari, G. Ershadi, and S. Hessabi. On the scheduling of energyaware fault-tolerant mixedcriticality multicore systems with service guarantee exploration. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2338–2354, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sheikh:2016:ETP

[SAF16] Hafiz Fahad Sheikh, Ishfaq Ahmad, and Dongrui Fan. An evolutionary technique for performanceenergy-temperature optimized scheduling of parallel tasks on multi-core processors. IEEE Transactions on Parallel and Distributed Systems, 27(3):668– 681, March 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/03/07083753-
                                        abs.html.

Sanchez-Artigas:2015:ASH

[SAH15] Marc Sanchez-Artigas and Blas Herrera. Activity stereotypes, or how to cope with disconnection during trust bootstrapping. IEEE Transactions on Parallel and Distributed Systems, 26(1):2–12, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06748038-
                                        abs.html.

Samaan:2014:NES

[Sam14a] Nancy Samaan. A novel economic sharing model in a federation of selfish cloud providers. IEEE Transactions on Parallel and Distributed Systems, 25(1):12– 21, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Senouci:2014:LMA

[SAM14b] Mustapha Reda Senouci, Khalid Assnoune, and Abdelhamid Mellouk. Localized movement-assisted sensor deployment algorithm for hole detection and healing. IEEE Transactions on Parallel and Distributed Systems, 25(5):1267–1277, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Son:2019:SIC

[SB19] Jungmin Son and Rajkumar Buyya. SDCon: Integrated control platform for softwaredefined clouds. IEEE Transactions on Parallel and Distributed Systems, 30(1):230– 244, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08409965-
abs.html.

Sousa:2010:HAI

[SBC^{+}10] Paulo Sousa, Alysson Neves Bessani, Miguel Correia, Nuno Ferreira Neves, and Paulo Verissimo. Highly available intrusion-tolerant services with proactivereactive recovery. IEEE Transactions on Parallel and Distributed Systems, 21(4): 452–465, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shudler:2019:EAS

[SBC^{+}19] S. Shudler, Y. Berens, A. Calotoiu, T. Hoefler, A. Strube, and F. Wolf. Engineering algorithms for scalability through continuous validation of performance expectations. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1768–1785, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Singh:2015:CAE

[SBMA15] Jagpreet Singh, Sandeep Betha, Bhargav Mangipudi, and Nitin Auluck. Contention aware energy efficient scheduling on heterogeneous multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 26(5):1251–1264, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/05/06811172-
                                        abs.html.

Shih:2011:PVC

[SCC11] Kuei-Ping Shih, Yen-Da Chen, and Chau-Chieh Chang. A physical/virtual carrier-sense-based power control MAC protocol for collision avoidance in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 22(2): 193–207, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Soldo:2011:VSD

[SCCC11] Fabio Soldo, Claudio Casetti, Carla-Fabiana Chiasserini, and Pedro Alonso Chaparro. Video streaming distribution in VANETs. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1085–1091, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Chen:2014:DAE

[sCCyW14] Wen sheng Chen, Ying-Jun Chen, and Shiow yang Wu. Dynamic aggregate: An elastic framework for QoSaware distributed processing of RFID data on enterprise hierarchy. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1724–1734, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scogland:2015:CCT

[ScFRdS15] Thomas R. W. Scogland, Wu chun Feng, Barry Rountree, and Bronis R. de Supinski. CoreTSAR: Core tasksize adapting runtime. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2970–2983, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06936921-
abs.html.

Sun:2011:EAS

[SCH11] Hongyang Sun, Yangjie Cao, and Wen-Jing Hsu. Efficient adaptive scheduling of multiprocessors with stable parallelism feedback. IEEE Transactions on Parallel and Distributed Systems, 22(4): 594–607, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shi:2015:MGH

[SCH^{+}15] Xuanhua Shi, Ming Chen, Ligang He, Xu Xie, Lu Lu, Hai Jin, Yong Chen, and Song Wu. Mammoth: Gearing Hadoop towards memory-intensive MapReduce applications. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2300–2315, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06869021-
abs.html.

Sun:2016:RNT

[SCHT16] Jianhua Sun, Hao Chen, Ligang He, and Huailiang Tan. Redundant network traffic elimination with GPU accelerated Rabin fingerprinting. IEEE Transactions on Parallel and Distributed Systems, 27(7):2130–2142, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07225160-
abs.html.

Song:2017:SWM

[SCJ^{+}17] Wei Song, Fangfei Chen, Hans-Arno Jacobsen, Xiaoxu Xia, Chunyang Ye, and Xiaoxing Ma. Scientific workflow mining in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2979–2992, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07907335-
abs.html.

Siebert:2015:LLA

[SCL^{+}15] Joanna Siebert, Jiannong Cao, Yi Lai, Peng Guo, and Weiping Zhu. LASEC: A localized approach to service composition in pervasive computing environments. IEEE Transactions on Parallel and Distributed Systems, 26(7):1948–1957, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06846304-
abs.html.

Shan:2017:AEH

[SDG17] Jianchen Shan, Xiaoning Ding, and Narain Gehani.

APPLES: Efficiently handling spin-lock synchronization on virtualized platforms. IEEE Transactions on Parallel and Distributed Systems, 28(7):1811–1824, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07736153-
abs.html.

Sze:2015:COC

[SDL^{+}15] Wai Kit Sze, Yulin Deng, Wing Cheong Lau, Murali Kodialam, Thyaga Nandagopal, and Onching Yue. Channeloblivious counting algorithms for large-scale RFID systems. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3303–3316, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06991608-
abs.html.

Sahoo:2018:LDE

[SDV18] Prasan Kumar Sahoo, Chinmaya Kumar Dehury, and Bharadwaj Veeravalli. LVRM: On the design of efficient link based virtual resource management algorithm for cloud platforms. IEEE Transactions on Parallel and Distributed Systems, 29(4): 887–900, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08169111-
abs.html.

Salkhordeh:2018:RER

[SEA18] Reza Salkhordeh, Shahriar Ebrahimi, and Hossein Asadi. ReCA: An efficient reconfigurable cache architecture for storage systems with online workload characterization. IEEE Transactions on Parallel and Distributed Systems, 29(7): 1605–1620, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08265001-
abs.html.

Salehi:2016:TPL

[SEAH16] Mohammad Salehi, Alireza Ejlali, and Bashir M. Al-Hashimi. Two-phase lowenergy n-modular redundancy for hard real-time multi-core systems. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1497–1510, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07122343-
abs.html.

Sehgal:2015:MMB

[Seh15] Vivek Kumar Sehgal. Markovian models based stochastic communication in networksin-package. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2806–2821, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06898833.
                                        pdf.

Sun:2010:CDD

[SF10] Jinyuan Sun and Yuguang Fang. Cross-domain data sharing in distributed electronic health record systems. IEEE Transactions on Parallel and Distributed Systems, 21(6):754–764, June 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Suzuki:2017:RTG

[SFA^{+}17] Yuhei Suzuki, Yusuke Fujii, Takuya Azumi, Nobuhiko Nishio, and Shinpei Kato. Real-time GPU resource management with loadable kernel modules. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1715–1727, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/06/07748485-
                                        abs.html.

Feng:2012:DWN

[sFC12] Vin sen Feng and Shih Yu Chang. Determination of wireless networks parameters through parallel hierarchical support vector machines. IEEE Transactions on Parallel and Distributed Systems, 23(3):505–512, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Saifullah:2014:PRT

[SFL^{+}14] Abusayeed Saifullah, David Ferry, Jing Li, Kunal Agrawal, Chenyang Lu, and Christopher Gill. Parallel real-time scheduling of DAGs. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3242–3252, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714435-
abs.html.

Salehkaleybar:2016:TBF

[SG16a] Saber Salehkaleybar and S. Jamaloddin Golestani. Token-based function computation with memory. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1811–1823, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07163357-
abs.html.

Sarofeen:2016:HPP

[SG16b] Christian Sarofeen and Philip Gillett. A high performance parallel and heterogeneous approach to narrowband beamforming. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2196–2207, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07305825.
                                        pdf.

Shu:2014:DAS

[SGC14] Yuanchao Shu, Yu Jason Gu, and Jiming Chen. Dynamic authentication with sensory information for the access control systems. IEEE Transactions on Parallel and Distributed Systems, 25(2): 427–436, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sereno:2014:RCR

[SGGB14] Matteo Sereno, Marco Grangetto, Rossano Gaeta, and Valerio Bioglio. Rateless codes and random walks for P2P resource discovery in grids. IEEE Transactions on Parallel and Distributed Systems, 25(4):1014–1023, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Salkhordeh:2019:EHC

[SHA19] R. Salkhordeh, M. Hadizadeh, and H. Asadi. An efficient hybrid I/O caching architecture using heterogeneous SSDs. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1238–1250, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2010:EAD

[She10a]  Haiying Shen.  An effi-

cient and adaptive decentralized file replication algorithm in P2P file sharing systems. IEEE Transactions on Parallel and Distributed Systems, 21(6):827–840, June 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2010:IIF

[She10b] Haiying (Helen) Shen. IRM:

Integrated file replication and consistency maintenance in P2P systems. IEEE Transactions on Parallel and Distributed Systems, 21(1): 100–113, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2014:LES

[She14] Xuemin Shen. A lightweight encryption scheme for networkcoded mobile ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2211–2221, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/09/06559980-
                                        abs.html.

Sun:2017:CMC

[SHF^{+}17] Yuanyuan Sun, Yu Hua, Dan Feng, Ling Yang, Pengfei Zuo, Shunde Cao, and Yuncheng Guo. A collisionmitigation cuckoo hashing scheme for large-scale storage systems. IEEE Transactions on Parallel and Distributed Systems, 28(3):619– 632, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/03/07523403-
                                        abs.html.

Samman:2011:NTD

[SHG11] Faizal Arya Samman, Thomas Hollstein, and Manfred Glesner. New theory for deadlock-free multicast routing in wormhole-switched virtual-channelless networkson-chip. IEEE Transactions on Parallel and Distributed Systems, 22(4):544– 557, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Samman:2013:RCB

[SHG13] Faizal Arya Samman, Thomas Hollstein, and Manfred Glesner. Runtime contention and bandwidth-aware adaptive routing selection strategies for networks-on-chip. IEEE Transactions on Parallel and Distributed Systems, 24(7):1411–1421, July 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stoleru:2012:AED

[SHM^{+}12] Radu Stoleru, Tian He, Siddhartha S. Mathiharan, Stephen M. George, and John A. Stankovic. Asymmetric event-driven node localization in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):634– 642, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Song:2010:DDS

[SHX^{+}10] Wen-Zhan Song, Renjie Huang, Mingsen Xu, Behrooz A. Shirazi, and Richard LaHusen. Design and deployment of sensor network for real-time high-fidelity volcano monitoring. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1658–1674, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sano:2014:MFA

[SHY14] Kentaro Sano, Yoshiaki Hatsuda, and Satoru Yamamoto.

Multi-FPGA accelerator for scalable stencil computation with constant memory bandwidth. IEEE Transactions on Parallel and Distributed Systems, 25(3):695– 705, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sibai:2012:TDL

[Sib12] Fadi N. Sibai. A twodimensional low-diameter scalable on-chip network for interconnecting thousands of cores. IEEE Transactions on Parallel and Distributed Systems, 23(2):193– 201, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Srivatsa:2011:PVN

[SILJ11] Mudhakar Srivatsa, Arun Iyengar, Ling Liu, and Hongbo Jiang. Privacy in VoIP networks: Flow analysis attacks and defense. IEEE Transactions on Parallel and Distributed Systems, 22(4): 621–633, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Surobhi:2014:CAM

[SJ14] Nusrat Ahmed Surobhi and Abbas Jamalipour. A context-aware M2M-based middleware for service selection in mobile ad-hoc networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3056–3065, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06747333-
abs.html.

Silva:2019:JPS

[SJAdCL19] M. L. Silva, L. N. S.

Júnior, A. L. L. Aquino, and J. d. C. Lima. JSensor: A parallel simulator for huge wireless sensor networks applications. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2296–2308, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shan:2017:AED

[SJR17] Danfeng Shan, Wanchun Jiang, and Fengyuan Ren. Analyzing and enhancing dynamic threshold policy of data center switches. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2454–2470, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07859368-
abs.html.

Santander-Jimenez:2015:PMM

[SJVR15] Sergio Santander-Jimenez and Miguel A. Vega-Rodriguez.

Parallel multiobjective metaheuristics for inferring phylogenies on multicore clusters. IEEE Transactions on Parallel and Distributed Systems, 26(6):1678–1692, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06819075-
abs.html.

Santander-Jimenez:2017:ANG

[SJVR17] Sergio Santander-Jimenez and Miguel A. Vega-Rodriguez. Asynchronous non-generational model to parallelize metaheuristics: A bioinformatics case study. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1825–1838, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07801038-
abs.html.

Santander-Jimenez:2019:CAI

[SJVR19] Sergio Santander-Jimenez and Miguel A. Vega-Rodriguez. Comparative analysis of intra-algorithm parallel multiobjective evolutionary algorithms: Taxonomy implications on bioinformatics scenarios. IEEE Transactions on Parallel and Distributed Systems, 30(1):63– 78, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08409336-
abs.html.

Shen:2014:HDS

[SK14] Min Shen and Ajay D.

Kshemkalyani. Hierarchical detection of strong unstable conjunctive predicates in large-scale systems. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2899–2908, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06684148-
abs.html.

Stovall:2015:GGB

[SKA15] Thomas Ryan Stovall, Sinan Kockara, and Recep Avci. GPUSCAN: GPU-based parallel structural clustering algorithm for networks. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3381–3393, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06967853-
abs.html.

Sarangi:2014:ASH

[SKGC14] Smruti R. Sarangi, Prathmesh Kallurkar, Parul Gupta, and Sandeep Chandran. Architectural support for handling jitter in shared memory based parallel applications.

IEEE Transactions on Parallel and Distributed Systems, 25(5):1166–1176, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Srinivasan:2016:EHW

[SKKK16] Sudarshan Srinivasan, Nithesh Kurella, Israel Koren, and Sandip Kundu. Exploring heterogeneity within a core for improved power efficiency. IEEE Transactions on Parallel and Distributed Systems, 27(4): 1057–1069, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07103357-
                                        abs.html.

Shpiner:2015:CBN

[SKL^{+}15] Alexander Shpiner, Erez Kantor, Pu Li, Israel Cidon, and Isaac Keslassy. On the capacity of bufferless networks-on-chip. IEEE Transactions on Parallel and Distributed Systems, 26(2): 492–506, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06762977-
                                        abs.html.

Stai:2012:TEW

[SKP12] Eleni Stai, Vasileios Karyotis, and Symeon Papavassiliou. Topology enhancements in wireless multihop networks: a top-down approach. IEEE Transactions on Parallel and Distributed Systems, 23(7):1344–1357, July 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shu:2011:FMN

[SL11] Guoqiang Shu and David Lee. A formal methodology for network protocol fingerprinting. IEEE Transactions on Parallel and Distributed Systems, 22(11):1813–1825, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2013:GAP

[SL13] Haiying Shen and Guoxin Liu. A geographically aware poll-based distributed file consistency maintenance method for P2P systems. IEEE Transactions on Parallel and Distributed Systems, 24(11):2148–2159, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shin:2014:IRT

[SL14] Kang G. Shin and Jinkyu Lee. Improvement of realtime multi-core schedulability with forced nonpreemption. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1233–1243, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2016:NBS

[SL16] Haiying Shen and Zhuozhao Li. New bandwidth sharing and pricing policies to achieve a win-win situation for cloud provider and tenants. IEEE Transactions on Parallel and Distributed Systems, 27(9):2682–2697, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07317795-
abs.html.

Shen:2015:SPO

[SLC15] Haiying Shen, Ze Li, and Kang Chen. Social-P2P: An online social network based P2P file sharing system. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2874–2889, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06902795.
pdf.

Shao:2010:FOT

[SLG10] Bin Shao, Du Li, and Ning Gu. A fast operational transformation algorithm for mobile and asynchronous collaboration. IEEE Transactions on Parallel and Distributed Systems, 21(12):

1707–1720, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sharma:2018:MRD

[SLG^{+}18] Prateek Sharma, Stephen Lee, Tian Guo, David Irwin, and Prashant Shenoy. Managing risk in a derivative IaaS cloud. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1750–1765, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/07833031-
abs.html.

Shen:2014:PBI

[SLGW14] Haiying Shen, Guoxin Liu, Jill Gemmill, and Lee Ward. A P2P-based infrastructure for adaptive trustworthy and efficient communication in wide-area distributed systems. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2222–2233, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06542624-
abs.html.

Shen:2013:DAC

[SLL13a] Haiying Shen, Ze Li, and Jin Li. A DHT-aided chunkdriven overlay for scalable and efficient peer-to-peer live streaming. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2125–2137, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2013:RRT

[SLL13b] Haiying Shen, Yuhua Lin, and Ze Li. Refining reputation to truly select high-QoS servers in peer-to-peer networks. IEEE Transactions on Parallel and Distributed Systems, 24(12):2439–2450, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Salah:2016:LMN

[SLL16] Ahmad Salah, Kenli Li, and Keqin Li. Lazy-Merge: A novel implementation for indexed parallel k-way in-place merging. IEEE Transactions on Parallel and Distributed Systems, 27(7):2049–2061, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/07/07236913-
                                        abs.html.

Shen:2014:SPA

[SLLL14] Haiying Shen, Ze Li, Yuhua Lin, and Jin Li. SocialTube: P2P-assisted video sharing in online social networks. IEEE Transactions on Parallel and Distributed Systems, 25(9):

2428–2440, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06519241-
abs.html.

Shen:2016:FPA

[SLLZ16] Haiying Shen, Alex X. Liu, Guoxin Liu, and Lianyu Zhao. Freeweb: P2P-assisted collaborative censorshipresistant Web browsing. IEEE Transactions on Parallel and Distributed Systems, 27(11):3226–3241, November 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07194820-
abs.html.

Smaragdakis:2010:DNF

[SLM^{+}10] Georgios Smaragdakis, Nikolaos Laoutaris, Pietro Michiardi, Azer Bestavros, John W. Byers, and Mema Roussopoulos. Distributed network formation for n-way broadcast applications. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1427–1441, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Subramanian:2016:BBP

[SLS^{+}16] Lavanya Subramanian, Donghyuk Lee, Vivek Seshadri, Harsha Rastogi, and Onur Mutlu.

BLISS: Balancing performance, fairness and complexity in memory access scheduling. IEEE Transactions on Parallel and Distributed Systems, 27(10): 3071–3087, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07399423-
abs.html.

Shen:2018:EAD

[SLSG18] Zhirong Shen, Patrick P. C. Lee, Jiwu Shu, and Wenzhong Guo. Encoding-aware data placement for efficient degraded reads in XORcoded storage systems: Algorithms and evaluation. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2757–2770, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08370044-
abs.html.

Shen:2019:CAS

[SLSG19] Z. Shen, P. P. C. Lee, J. Shu, and W. Guo. Correlationaware stripe organization for efficient writes in erasurecoded storage: Algorithms and evaluation. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1552–1564, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2016:ECR

[SLSL16] Haiying Shen, Yuhua Lin, Karan Sapra, and Ze Li. Enhancing collusion resilience in reputation systems. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2274–2287, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07295625-
abs.html.

Shen:2015:PAI

[SLW15] Haiying Shen, Guoxin Liu, and Lee Ward. A proximityaware interest-clustered P2P file sharing system. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1509–1523, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06823748-
abs.html.

Shen:2019:EGA

[SLX19] M. Shen, G. Luo, and N. Xiao. Exploring GPUaccelerated routing for FP-GAs. IEEE Transactions on Parallel and Distributed Systems, 30(6):1331–1345, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shangguan:2014:OTO

[SLY^{+}14] Longfei Shangguan, Zhenjiang Li, Zheng Yang, Mo Li, Yunhao Liu, and Jinsong Han. OTrack: Towards order tracking for tags in mobile RFID systems. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2114–2125, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Staudt:2016:EPA

[SM16] Christian L. Staudt and Henning Meyerhenke. Engineering parallel algorithms for community detection in massive networks. IEEE Transactions on Parallel and Distributed Systems, 27 (1):171–184, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07006796-
abs.html.

Shah:2018:CHS

[SMB^{+}18] Zubair Shah, Abdun Naser Mahmood, Michael Barlow, Zahir Tari, Xun Yi, and Albert Y. Zomaya. Computing hierarchical summary from two-dimensional big data streams. IEEE Transactions on Parallel and Distributed Systems, 29(4):803– 818, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08125170-
abs.html.

Sonmez:2010:BPC

[SME10] Omer Ozan Sonmez, Hashim Mohamed, and Dick H. J. Epema. On the benefit of processor coallocation in multicluster grid systems. IEEE Transactions on Parallel and Distributed Systems, 21(6):778–789, June 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Scalosub:2013:BMA

[SML13] Gabriel Scalosub, Peter Marbach, and Jorg Liebeherr. Buffer management for aggregated streaming data with packet dependencies. IEEE Transactions on Parallel and Distributed Systems, 24(3): 439–449, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sehrish:2013:SHA

[SMS^{+}13] Saba Sehrish, Grant Mackey, Pengju Shang, Jun Wang, and John Bent. Supporting HPC analytics applications with access patterns using data restructuring and data-centric scheduling techniques in MapReduce. IEEE Transactions on Parallel and Distributed Systems, 24(1): 158–169, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shah:2017:TEQ

[SMTZ17] Zubair Shah, Abdun Naser Mahmood, Zahir Tari, and Albert Y. Zomaya. A technique for efficient query estimation over distributed data streams. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2770–2783, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07898394-
abs.html.

Selvitopi:2015:NMS

[SOA15] R. Oguz Selvitopi, Muhammet Mustafa Ozdal, and Cevdet Aykanat. A novel method for scaling iterative solvers: Avoiding latency overhead of parallel sparsematrix vector multiplies. IEEE Transactions on Parallel and Distributed Systems, 26(3):632–645, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06766662-
abs.html.

Sudo:2019:LSL

[SOK^{+}19] Y. Sudo, F. Ooshita, H. Kakugawa, T. Masuzawa, A. K. Datta, and L. L. Larmore. Loosely-stabilizing leader election for arbitrary graphs in population protocol model. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1359–1373, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sugaya:2012:SAO

[SOTN12] Yoshihiro Sugaya, Shinichiro Omachi, Akira Takeuchi, and Yousuke Nozaki. A statistical analysis on operation scheduling for an energy network project. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1583–1592, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2012:EET

[SP12] Junqing Sun and Gregory D.

Peterson. An effective execution time approximation method for parallel computing. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2024–2032, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shim:2015:SDA

[SP15] Kyung-Ah Shim and Cheol-Min Park. A secure data aggregation scheme based on appropriate cryptographic primitives in heterogeneous wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2128–2139, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06875932-
abs.html.

Squicciarini:2010:GBN

[SPB^{+}10] Anna C. Squicciarini, Federica Paci, Elisa Bertino, Alberto Trombetta, and Stefano Braghin. Group-based negotiations in P2P systems. IEEE Transactions on Parallel and Distributed Systems, 21(10):1473–1486, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Si:2018:DAA

[SPH^{+}18] Min Si, Antonio J. Pena, Jeff Hammond, Pavan Balaji, Masamichi Takagi, and Yutaka Ishikawa. Dynamic adaptable asynchronous progress model for MPI RMA multiphase applications. IEEE Transactions on Parallel and Distributed Systems, 29(9): 1975–1989, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08315136-
abs.html.

Siavoshani:2018:SCL

[SPS18] Mahdi Jafari Siavoshani, Ali Pourmiri, and Seyed Pooya Shariatpanahi. Storage, communication, and load balancing trade-off in distributed cache networks. IEEE Transactions on Parallel and Distributed Systems, 29(4):943– 957, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08170231-
abs.html.

Schmid:2014:GLD

[SRB14] Ulrich Schmid, Peter Robinson, and Martin Biely. The generalized loneliness detector and weak system models for k-set agreement. IEEE Transactions on Parallel and Distributed Systems, 25(4): 1078–1088, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sourav:2019:SLF

[SRG19] S. Sourav, P. Robinson, and S. Gilbert. Slow links, fast links, and the cost of gossip. IEEE Transactions on Parallel and Distributed Systems, 30(9):2130–2147, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Schmidt:2012:DRT

[SS12] Klaus Werner Schmidt and Ece Guran Schmidt. Distributed real-time protocols for industrial control systems: Framework and examples. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1856–1866, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Souravlas:2017:BTB

[SS17] Stavros Souravlas and Angelo Sifaleras. Binary-tree based estimation of file requests for efficient data replication. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1839–1852, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07811179-
abs.html.

Shrestha:2018:MLD

[SS18] Anup Shrestha and Inanc Senocak. Multi-level domaindecomposition strategy for solving the eikonal equation with the fast-sweeping method. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2297–2303, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/10/08345737-
abs.html.

Shen:2016:HCA

[SSF16a] Zhirong Shen, Jiwu Shu, and Yingxun Fu. HV code: An all-around MDS code for RAID-6 storage systems. IEEE Transactions on Parallel and Distributed Systems, 27(6):1674–1686, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07180372-
abs.html.

Shen:2016:PSD

[SSF16b] Zhirong Shen, Jiwu Shu, and Yingxun Fu. Parity-switched data placement: Optimizing partial stripe writes in XORcoded storage systems. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3311–3322, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07399425-
abs.html.

Shen:2017:SEO

[SSLF17] Zhirong Shen, Jiwu Shu, Patrick P. C. Lee, and Yingxun Fu. Seek-efficient I/O optimization in single failure recovery for XORcoded storage systems. IEEE Transactions on Parallel and Distributed Systems, 28(3):

877–890, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07511742-
abs.html.

Shi:2018:RBN

[SSM^{+}18] Jian Shi, Brian Sullivan, Mike Mazzola, Babak Saravi, Uttam Adhikari, and Tomasz Haupt. A relaxationbased network decomposition algorithm for parallel transient stability simulation with improved convergence. IEEE Transactions on Parallel and Distributed Systems, 29(3):496–511, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8094963/.

Selfa:2017:HAF

[SSPG17] Vicent Selfa, Julio Sahuquillo, Salvador Petit, and Maria E. Gomez. A hardware approach to fairly balance the inter-thread interference in shared caches. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3021–3032, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07944620-
abs.html.

Shi:2017:CMA

[SSW^{+}17] Lei Shi, Yi Shi, Xing Wei, Xu Ding, and Zhenchun Wei. Cost minimization algorithms for data center management. IEEE Transactions on Parallel and Distributed Systems, 28(1):60– 71, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07445250-
abs.html.

Satsiou:2010:RBR

[ST10] Anna Satsiou and Leandros Tassiulas. Reputation-based resource allocation in P2P systems of rational users. IEEE Transactions on Parallel and Distributed Systems, 21(4):466–479, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sigdel:2018:CDI

[ST18] Purushottam Sigdel and Nian-Feng Tzeng. Coalescing and deduplicating incremental checkpoint files for restore-express multilevel checkpointing. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2713–2727, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

trans/td/2018/12/08372629-
abs.html.

Shahzad:2019:CLE

[STK^{+}19] Faisal Shahzad, Jonas Thies, Moritz Kreutzer, Thomas Zeiser, Georg Hager, and Gerhard Wellein. CRAFT: A library for easier applicationlevel checkpoint/restart and automatic fault tolerance. IEEE Transactions on Parallel and Distributed Systems, 30(3):501–514, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08444763-
abs.html.

Stamatakis:2017:GPA

[STMM17] Dimokritos Stamatakis, Nikos Tsikoudis, Eirini Micheli, and Kostas Magoutis. A general-purpose architecture for replicated metadata services in distributed file systems. IEEE Transactions on Parallel and Distributed Systems, 28(10):2747–2759, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07917318-
abs.html.

Stojmenovic:2010:ENa

[Sto10a]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 21(1):1– 3, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2010:ENb

[Sto10b]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 21(3):289– 291, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2010:ENc

[Sto10c]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 21(5):577– 578, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2010:ENd

[Sto10d]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 21(6):737– 738, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2010:ENe

[Sto10e]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1057–59, August 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2010:ENH

[Sto10f] Ivan Stojmenovic. Editor’s note: How to write research articles in computing and engineering disciplines. IEEE Transactions on Parallel and Distributed Systems, 21(2): 145–147, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2011:EMC

[Sto11a] Ivan Stojmenovic. Editorial: Media center. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1777, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2011:ENa

[Sto11b]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 22(1):1– 2, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2011:ENb

[Sto11c]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 22(5):705– 707, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2012:ENa

[Sto12a] Ivan Stojmenovic. Editor’s note. IEEE Transactions on Parallel and Distributed Systems, 23(1):1, January 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2012:ENb

[Sto12b]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 23(4):577– 578, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2013:ENa

[Sto13a]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 24(1):1– 3, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2013:ENb

[Sto13b]  Ivan Stojmenovic.  Edi-

tor’s note. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2121–2124, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stojmenovic:2013:ENE

[Sto13c]  Ivan Stojmenovic.  Edi-

tor’s note: EIC farewell and new EIC introduction. IEEE Transactions on Parallel and Distributed Systems, 24(12):2322–2323, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sabarimuthu:2019:ADC

[SV19] J. M. Sabarimuthu and T. G. Venkatesh. Analytical derivation of concurrent reuse distance profile for multi-threaded application running on chip multiprocessor. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1704–1721, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Stillwell:2012:DFR

[SVC12] Mark Stillwell, Frederic Vivien, and Henri Casanova. Dynamic fractional resource scheduling versus batch scheduling. IEEE Transactions on Parallel and Distributed Systems, 23(3):521– 529, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sim:2019:AWA

[SVK^{+}19] Hyogi Sim, Geoffroy Vallee, Youngjae Kim, Sudharshan S. Vazhkudai, Devesh Tiwari, and Ali R. Butt. An analysis workflow-aware storage system for multi-core active flash arrays. IEEE Transactions on Parallel and Distributed Systems, 30(2): 271–285, February 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08436425-
abs.html.

Shen:2016:WPA

[SVL^{+}16] Jie Shen, Ana Lucia Varbanescu, Yutong Lu, Peng Zou, and Henk Sips. Workload partitioning for accelerating applications on heterogeneous platforms. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2766–2780, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07360199-
abs.html.

Sun:2014:VPP

[SWC^{+}14] Wenhai Sun, Bing Wang, Ning Cao, Ming Li, Wenjing Lou, Y. Thomas Hou, and Hui Li. Verifiable privacypreserving multi-keyword text search in the cloud supporting similarity-based ranking. IEEE Transactions on Parallel and Distributed Systems, 25(11): 3025–3035, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2014/11/06656804-
                                        abs.html.

Shi:2017:OAM

[SWL17] Weijie Shi, Chuan Wu, and Zongpeng Li. An online auction mechanism for dynamic virtual cluster provisioning in geo-distributed clouds. IEEE Transactions on Parallel and Distributed Systems, 28(3):677–688, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/03/07548335-
                                        abs.html.

Sha:2018:MOP

[SWRQ18] Shi Sha, Wujie Wen, Shaolei Ren, and Gang Quan. M-Oscillating: Performance maximization on temperatureconstrained multi-core processors. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2528–2539, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/11/08359006-
                                        abs.html.

Sharif:2017:PAS

[SWT^{+}17] Shaghahyegh Sharif, Paul Watson, Javid Taheri, Surya Nepal, and Albert Y. Zomaya. Privacy-aware scheduling SaaS in high performance computing environments. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1176–1188, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07552583-
abs.html.

Siwik:2019:PRI

[SWT^{+}19] L. Siwik, M. Wo niak, V. Trujillo, D. Pardo, V. M. Calo, and M. Paszy ski. Parallel refined isogeometric analysis in 3D. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1134–1142, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shen:2010:ERT

[SX10] Haiying Shen and Cheng-Zhong Xu. Elastic routing table with provable performance for congestion control in DHT networks. IEEE Transactions on Parallel and Distributed Systems, 21(2): 242–256, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sano:2017:FBS

[SY17] Kentaro Sano and Satoru Yamamoto. FPGA-based scalable and power-efficient fluid simulation using floatingpoint DSP blocks. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2823–2837, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07893769-
abs.html.

Siganos:2014:BLT

[SYL^{+}14] Georgos Siganos, Xiaoyuan Yang, Nikolaos Laoutaris, Pablo Rodriguez, and Ruben Cuevas. BitTorrent locality and transit traffic reduction: When, why, and at what cost? IEEE Transactions on Parallel and Distributed Systems, 25(5): 1177–1189, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sun:2016:PYR

[SYL^{+}16] Wenhai Sun, Shucheng Yu, Wenjing Lou, Y. Thomas Hou, and Hui Li. Protecting your right: Verifiable attribute-based keyword search with fine-grained owner-enforced search authorization in the cloud. IEEE Transactions on Parallel and Distributed Systems, 27(4):1187–1198, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/06893023-
abs.html.

Shi:2019:FCA

[SytL19] M. Shi, y. tang, and J. Liu.

Functional and contextual attention-based LSTM for service recommendation in mashup creation. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1077–1090, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Su:2016:EGI

[SYXL16] Yu Su, Ding Ye, Jingling Xue, and Xiang-Ke Liao. An efficient GPU implementation of inclusion-based pointer analysis. IEEE Transactions on Parallel and Distributed Systems, 27(2): 353–366, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07029117-
abs.html.

Sookhak:2018:ABD

[SYZ18] Mehdi Sookhak, F. Richard Yu, and Albert Y. Zomaya. Auditing big data storage in cloud computing using divide and conquer tables. IEEE Transactions on Parallel and Distributed Systems, 29(5):999–1012, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2018/05/08219724-
                                        abs.html.

Sun:2011:DAR

[SZ11] Song Sun and Joseph Zambreno. Design and analysis of a reconfigurable platform for frequent pattern mining. IEEE Transactions on Parallel and Distributed Systems, 22(9):1497–1505, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Si:2012:NMO

[SZ12] Weisheng Si and Albert Y. Zomaya. New memoryless online routing algorithms for Delaunay triangulations. IEEE Transactions on Parallel and Distributed Systems, 23(8):1520–1527, August 2012. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Sarje:2011:APC

[SZA11] Abhinav Sarje, Jaroslaw Zola, and Srinivas Aluru. Accelerating pairwise computations on Cell processors. IEEE Transactions on Parallel and Distributed Systems, 22(1):69–77, January 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Savage:2017:DMC

[SZC^{+}17] David Savage, Xiuzhen Zhang, Pauline Chou, Xinghou Yu, and Qingmai Wang. Distributed mining of contrast patterns. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1881–1890, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07782416-
abs.html.

Sang:2012:PSO

[SZL^{+}12] Bo Sang, Jianfeng Zhan, Gang Lu, Haining Wang, Dongyan Xu, Lei Wang, Zhihong Zhang, and Zhen Jia. Precise, scalable, and online request tracing for multitier services of black boxes. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1159–1167, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Shi:2017:EAS

[SZR17] Li Shi, Zhemin Zhang, and Thomas Robertazzi. Energyaware scheduling of embarrassingly parallel jobs and resource allocation in cloud. IEEE Transactions on Parallel and Distributed Systems, 28(6):1607–1620, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07736137-
abs.html.

Song:2015:IML

[SZWX15] Guojie Song, Xiabing Zhou, Yu Wang, and Kunqing Xie. Influence maximization on large-scale mobile social network: A divideand-conquer method. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1379–1392, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/05/06805620-
                                        abs.html.

Sun:2010:IBS

[SZZF10] Jinyuan Sun, Chi Zhang, Yanchao Zhang, and Yuguang Fang. An identity-based security system for user privacy in vehicular ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 21(9):1227–1239, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Takeuchi:2014:GCG

[Tak14] Akira Takeuchi. GT-CFS: A game theoretic coalition formulation strategy for reducing power loss in micro grids. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2307–2317, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06565982-
abs.html.

Tani:2012:CVA

[Tan12] Seiichiro Tani. Compression of view on anonymous networks ‘folded view’. IEEE Transactions on Parallel and Distributed Systems, 23(2): 255–262, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tuncer:2019:ODP

[TAZ^{+}19] Ozan Tuncer, Emre Ates, Yijia Zhang, Ata Turk, Jim Brandt, Vitus J. Leung, Manuel Egele, and Ayse K. Coskun. Online diagnosis of performance variation in HPC systems using machine learning. IEEE Transactions on Parallel and Distributed Systems, 30(4):883–896, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8466019/

.

Tan:2019:HRT

[TBÁ^{+}19] X. Tan, J. Bosch, C. Álvarez, D. Jiménez-González, E. Ayguadé, and M. Valero. A hardware runtime for task-based programming models. IEEE Transactions on Parallel and Distributed Systems, 30(9): 1932–1946, September 2019.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tiwari:2012:PIL

[TBC12] Honey Durga Tiwari, Huynh Ngoc Bao, and Yong Beom Cho. A parallel IRRWBF LDPC decoder based on stream-based processor. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2198–2204, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tolosana-Calasanz:2017:QTB

[TCDMRP17] Rafael Tolosana-Calasanz, Javier Diaz-Montes, Omer F. Rana, and Manish Parashar. Queueing theory-based resource management for streaming applications feedbackcontrol & queueing theorybased resource management for streaming applications. IEEE Transactions on Parallel and Distributed Systems, 28(4):1061–1075, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/04/07553568-
                                        abs.html.

Tagliavini:2018:UFG

[TCM18] Giuseppe Tagliavini, Daniele Cesarini, and Andrea Marongiu. Unleashing fine-grained parallelism on embedded manycore accelerators with lightweight OpenMP tasking. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2150–2163, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/08314096-
abs.html.

Thulasiraman:2011:MRM

[TCS11] Preetha Thulasiraman, Jiming Chen, and Xuemin (Sherman) Shen. Multipath routing and max-min fair QoS provisioning under interference constraints in wireless multihop networks. IEEE Transactions on Parallel and Distributed Systems, 22(5): 716–728, May 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tian:2013:FCZ

[TCS13] Xiaohua Tian, Yu Cheng, and Xuemin (Sherman) Shen. Fast channel zapping with destination-oriented multicast for IP video delivery. IEEE Transactions on Parallel and Distributed Systems, 24(2):327– 341, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsai:2014:TPW

[TCT14] Tsung-Han Tsai, Y-Chuang Chen, and Jimmy J. M.

Tan. Topological properties on the wide and fault diameters of exchanged hypercubes. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3317–3327, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06747316-
abs.html.

Tsai:2016:OEC

[TCT16] Tsung-Han Tsai, Y-Chuang Chen, and Jimmy J. M. Tan. Optimal edge congestion of exchanged hypercubes. IEEE Transactions on Parallel and Distributed Systems, 27(1):250– 262, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07001085-
abs.html.

Tan:2016:MIH

[TCYF16] Jingweijia Tan, Mingsong Chen, Yang Yi, and Xin Fu. Mitigating the impact of hardware variability for GPGPUs register file. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3283–3297, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

trans/td/2016/11/07412772-
abs.html.

Tong:2011:DAW

[TCZL11] Hongxia Tong, Jian Cao, Shensheng Zhang, and Minglu Li. A distributed algorithm for Web service composition based on service agent model. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2008–2021, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tao:2019:OLC

[TDL^{+}19] D. Tao, S. Di, X. Liang, Z. Chen, and F. Cappello. Optimizing lossy compression rate-distortion from automatic online selection between SZ and ZFP. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1857–1871, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tian:2013:TMG

[TDLR13] Ye Tian, Ratan Dey, Yong Liu, and Keith W. Ross. Topology mapping and geolocating for China’s Internet. IEEE Transactions on Parallel and Distributed Systems, 24(9):1908–1917, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Truong:2017:DSN

[TFKN17] Nguyen T. Truong, Ikki Fujiwara, Michihiro Koibuchi, and Khanh-Van Nguyen. Distributed shortcut networks: Low-latency lowdegree non-random topologies targeting the diameter and cable length trade-off. IEEE Transactions on Parallel and Distributed Systems, 28(4):989–1001, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07576694.
pdf.

Tsai:2018:LBM

[TFLL18] Linjiun Tsai, Hubertus Franke, Chung-Sheng Li, and Wanjiun Liao. Learningbased memory allocation optimization for delay-sensitive big data processing. IEEE Transactions on Parallel and Distributed Systems, 29(6): 1332–1341, June 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/06/08276656-
abs.html.

Tarplee:2016:EMT

[TFM^{+}16] Kyle M. Tarplee, Ryan Friese, Anthony A. Maciejewski, Howard Jay Siegel, and Edwin K. P. Chong. Energy and makespan tradeoffs in heterogeneous computing systems using efficient linear programming techniques. IEEE Transactions on Parallel and Distributed Systems, 27(6):1633–1646, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07155583-
abs.html.

Turner:2013:CMB

[TFPK13] Andrew Turner, Andrew Fox, John Payne, and Hyong S. Kim. C-MART: Benchmarking the cloud. IEEE Transactions on Parallel and Distributed Systems, 24(6):1256–1266, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2013:ADS

[TG13] Yuzhe Tang and Bugra Gedik. Autopipelining for data stream processing. IEEE Transactions on Parallel and Distributed Systems, 24(12):2344–2354, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Titos-Gil:2013:EEM

[TGAG13] Ruben Titos-Gil, Manuel E.

Acacio, and Jose M. Garcia. Efficient eager management of conflicts for scalable hardware transactional memory. IEEE Transactions on Parallel and Distributed Systems, 24(1):59–71, January 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Titos-Gil:2019:WCA

[TGFFP^{+}19] R. Titos-Gil, A. Flores, R. Fernández-Pascual, A. Ros, S. Petit, J. Sahuquillo, and M. E. Acacio. Way combination for an adaptive and scalable coherence directory. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2608–2623, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Titos-Gil:2013:EBL

[TGNA^{+}13] Ruben Titos-Gil, Anurag Negi, Manuel E. Acacio, Jose M. Garcia, and Per Stenstrom. Eager beats lazy: Improving store management in eager hardware transactional memory. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2192–2201, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Traff:2010:SCM

[TGT10] Jesper Larsson Traff, William D. Gropp, and Rajeev Thakur. Self-consistent MPI performance guidelines. IEEE Transactions on Parallel and Distributed Systems, 21(5): 698–709, May 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tirado:2014:CFC

[THB^{+}14] Juan M. Tirado, Daniel Higuero, Javier Garcia Blas, Florin Isaila, and Jesus Carretero. CONDESA: A framework for controlling data distribution on elastic server architectures. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2010–2019, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2015:SMN

[THE^{+}15] Xiaoxin Tang, Zhiyi Huang, David Eyers, Steven Mills, and Minyi Guo. Scalable multicore k-NN search via subspace clustering for filtering. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3449–3460, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06963421-
abs.html.

Tran:2013:CLD

[THL13] Nguyen H. Tran, Choong Seon Hong, and Sungwon Lee. Cross-layer design of congestion control and power control in fast-fading wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):260– 274, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Traverso:2015:SAR

[THT^{+}15] Stefano Traverso, Kevin Huguenin, Ionut Trestian, Vijay Erramilli, Nikolaos Laoutaris, and Konstantina Papagiannaki. Social-aware replication in geo-diverse online systems. IEEE Transactions on Parallel and Distributed Systems, 26(2):584– 593, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06774946-
abs.html.

Tichy:2014:LID

[Tic14] Walter F. Tichy. Libraryindependent data race detection. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2606–2616, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06583165-
abs.html.

Tan:2014:SDS

[TJH^{+}14] Zhiyuan Tan, Aruna Jamdagni, Xiangjian He, Priyadarsi Nanda, and Ren Ping Liu. A system for denial-of-service attack detection based on multivariate correlation analysis. IEEE Transactions on Parallel and Distributed Systems, 25(2):447–456, February 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tian:2012:RDQ

[TJLL12] Chen Tian, Hongbo Jiang, Xue Liu, and Wenyu Liu. Revisiting dynamic query protocols in unstructured peerto-peer networks. IEEE Transactions on Parallel and Distributed Systems, 23(1): 160–167, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tati:2015:AAD

[TKC^{+}15] Srikar Tati, Bong Jun Ko, Guohong Cao, Ananthram Swami, and Thomas F. La Porta. Adaptive algorithms for diagnosing largescale failures in computer networks. IEEE Transactions on Parallel and Distributed Systems, 26(3):646– 656, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06767126-
abs.html.

Tsiropoulou:2012:DUP

[TKP12] Eirini Eleni Tsiropoulou, Georgios K. Katsinis, and Symeon Papavassiliou. Distributed uplink power control in multiservice wireless networks via a game theoretic approach with convex pricing. IEEE Transactions on Parallel and Distributed Systems, 23(1):61– 68, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tariq:2014:SBL

[TKR14] Muhammad Adnan Tariq, Boris Koldehofe, and Kurt Rothermel. Securing brokerless publish/subscribe systems using identity-based encryption. IEEE Transactions on Parallel and Distributed Systems, 25(2):518– 528, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2011:CLP

[TKS11] Huseyin Ozgur Tan, Ibrahim Korpeoglu, and Ivan Stojmenovic. Computing localized power-efficient data aggregation trees for sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(3):489– 500, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

TalebiFard:2014:EPT

[TL14] Peyman TalebiFard and Victor C. M. Leung. Expansion properties of topology for networking of information in cloud. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2877–2887, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06627891-
abs.html.

Tong:2016:SSR

[TL16] Guangmo Tong and Cong Liu. Supporting soft realtime sporadic task systems on uniform heterogeneous multiprocessors with no utilization loss. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2740–2752, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07336563-
abs.html.

Tan:2014:BMG

[TLH^{+}14] Huailiang Tan, Kenli Li, Ligang He, Jianhua Sun, and Hao Chen. BAG: Managing GPU as buffer cache in operating systems. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1393–1402, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2014:BBA

[TLJ^{+}14] Guang Tan, Mingming Lu, Fangsheng Jiang, Kongyang Chen, Xiaoxia Huang, and Jie Wu. Bumping: A bump-aided inertial navigation method for indoor vehicles using smartphones. IEEE Transactions on Parallel and Distributed Systems, 25(7):1670–1680, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Thapa:2016:SSP

[TLL^{+}16] Arun Thapa, Weixian Liao, Ming Li, Pan Li, and Jinyuan Sun. SPA: A secure and private auction framework for decentralized online social networks. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2394–2407, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/08/07303967-
                                        abs.html.

Tai:2012:AMO

[TLP12] Yi-Gang Tai, Chia-Tien Dan Lo, and Kleanthis Psarris. Accelerating matrix operations with improved deeply pipelined vector reduction. IEEE Transactions on Parallel and Distributed Systems, 23(2):202– 210, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2015:RPC

[TLP15] Minghua Tang, Xiaola Lin, and Maurizio Palesi. Routing pressure: A channelrelated and traffic-aware metric of routing algorithm. IEEE Transactions on Parallel and Distributed Systems, 26(3):891–901, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06570725-
abs.html.

Tang:2016:LCA

[TLP16] Minghua Tang, Xiaola Lin, and Maurizio Palesi. Local congestion avoidance in network-on-chip. IEEE Transactions on Parallel and Distributed Systems, 27(7): 2062–2073, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07229336-
abs.html.

Tang:2015:CAS

[TLRW15] Di Tang, Tongtong Li, Jian Ren, and Jie Wu. Cost-Aware SEcure Routing (CASER) protocol design for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(4):960–973, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/04/06800079-
                                        abs.html.

Thapa:2015:ASP

[TLSL15] Arun Thapa, Ming Li, Sergio Salinas, and Pan Li. Asymmetric social proximity based private matching protocols for online social networks. IEEE Transactions on Parallel and Distributed Systems, 26(6):1547–1559, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/06/06826575-
                                        abs.html.

Tessier:2014:PPM

[TMJ14] Francois Tessier, Guillaume Mercier, and Emmanuel Jeannot. Process placement in multicore clusters: Algorithmic issues and practical techniques. IEEE Transactions on Parallel and Distributed Systems, 25(4):993– 1002, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tavakoli:2015:EEC

[TMMN15] Hamidreza Tavakoli, Jelena Misic, Vojislav B. Misic, and Majid Naderi. Energyefficient cluster-head rotation in beacon-enabled IEEE 802.15.4 networks. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3371–3380, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06995960-
abs.html.

Tang:2018:LTM

[TNH^{+}18] Shanjiang Tang, Zhaojie Niu, Bingsheng He, Bu-Sung Lee, and Ce Yu. Longterm multi-resource fairness for pay-as-you use computing systems. IEEE Transactions on Parallel and Distributed Systems, 29(5): 1147–1160, May 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08244302-
abs.html.

Tarable:2017:IWR

[TNLM17] Alberto Tarable, Alessandro Nordio, Emilio Leonardi, and Marco Ajmone Marsan. The importance of worker reputation information in microtask-based crowd work systems. IEEE Transactions on Parallel and Distributed Systems, 28(2):558– 571, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07478121-
abs.html.

Ta:2012:ICS

[TNZ^{+}12] Duong Nguyen Binh Ta, Thang Nguyen, Suiping Zhou, Xueyan Tang, Wentong Cai, and Rassul Ayani. Interactivity-constrained server provisioning in large-scale distributed virtual environments. IEEE Transactions on Parallel and Distributed Systems, 23(2):304– 312, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Turk:2013:QLA

[TOA13] Ata Turk, Kerim Yasin Oktay, and Cevdet Aykanat. Query-log aware replicated declustering. IEEE Transactions on Parallel and Distributed Systems, 24(5):987– 995, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Touzene:2015:AABb

[Tou15a] Abderezak Touzene. Allto-all broadcast in hexagonal torus networks onchip. IEEE Transactions on Parallel and Distributed Systems, 26(9):2410–2420, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/09/06881698.
                                        pdf.

Touzene:2015:AABa

[Tou15b] Abderezak Touzene. On all-to-all broadcast in dense Gaussian network on-chip. IEEE Transactions on Parallel and Distributed Systems, 26(4):1085–1095, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06787056-
abs.html.

Tso:2013:IDC

[TP13] Fung Po Tso and Dimitrios P. Pezaros. Improving data center network utilization using near-optimal traffic engineering. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1139–1148, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tripathi:2014:RAC

[TP14] Anand Tripathi and Vinit Padhye. Resource availability characteristics and node selection in cooperatively shared computing platforms. IEEE Transactions on Parallel and Distributed Systems, 25(4):1044–1054, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tong:2018:SAF

[TP18] Da Tong and Viktor K.

Prasanna. Sketch acceleration on FPGA and its applications in network anomaly detection. IEEE Transactions on Parallel and Distributed Systems, 29(4):929– 942, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08085171-
abs.html.

Turcu:2016:ADP

[TPRH16] Alexandru Turcu, Roberto Palmieri, Binoy Ravindran, and Sachin Hirve. Automated data partitioning for highly scalable and strongly consistent transactions. IEEE Transactions on Parallel and Distributed Systems, 27(1):106– 118, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07004087-
abs.html.

Traff:2019:OTI

[Trä19] J. L. Träff. On optimal trees for irregular gather and scatter collectives. IEEE Transactions on Parallel and Distributed Systems, 30(9): 2060–2074, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsakalozos:2013:HBE

[TRD13] Konstantinos Tsakalozos, Mema Roussopoulos, and Alex Delis. Hint-based execution of workloads in clouds with Nefeli. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1331–1340, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2019:CSD

[TRT19] M. M. Tan, R. Ren, and X. Tang. Cloud scheduling with discrete charging units. IEEE Transactions on Parallel and Distributed Systems, 30(7):1541–1551, July 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tupinamba:2016:TOD

[TS16] Andre Luiz Rocha Tupinamba and Alexandre Sztajnberg. Transparent and optimized distributed processing on GPUs. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3673–3686, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07447815-
abs.html.

Tomczak:2018:SGH

[TS18] Tadeusz Tomczak and Roman G. Szafran. Sparse geometries handling in lattice Boltzmann method implementation for graphic processors. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1865–1878, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/08/08303717-
                                        abs.html.

Tsai:2013:FSG

[Tsa13] Jichiang Tsai. Flexible symmetrical global-snapshot algorithms for large-scale distributed systems. IEEE Transactions on Parallel and Distributed Systems, 24(3): 493–505, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tse:2013:OBT

[Tse13] S. S. H. Tse. Online balancing two independent criteria upon placements and deletions. IEEE Transactions on Parallel and Distributed Systems, 24(8):1644–1650, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tian:2010:IRA

[TSN10] Ye Tian, Hong Shen, and Kam-Wing Ng. Improving reliability for applicationlayer multicast overlays. IEEE Transactions on Parallel and Distributed Systems, 21(8):1103–1116, August 2010. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2015:CGF

[TTG^{+}15a] Wen Jun Tan, Wai Teng Tang, Rick Siow Mong Goh, Stephen John Turner, and Weng-Fai Wong. A code generation framework for targeting optimized library calls for multiple platforms. IEEE Transactions on Parallel and Distributed Systems, 26(7):1789–1799, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06827957-
abs.html.

Tang:2015:FBR

[TTG^{+}15b] Wai Teng Tang, Wen Jun Tan, Rick Siow Mong Goh, Stephen John Turner, and Weng-Fai Wong. A family of bit-representation-optimized formats for fast sparse matrix–vector multiplication on the GPU. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2373–2385, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06895301.
pdf.

Tan:2019:VMC

[TTH^{+}19] Huailiang Tan, Yanjie Tan, Xiaofei He, Kenli Li, and Keqin Li. A virtual multichannel GPU fair scheduling method for virtual machines. IEEE Transactions on Parallel and Distributed Systems, 30(2):257– 270, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08434359-
abs.html.

Tso:2012:MDE

[TTJX12] Fung Po Tso, Jin Teng, Weijia Jia, and Dong Xuan. Mobility: a double-edged sword for HSPA networks: a largescale test on Hong Kong mobile HSPA networks. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1895–1907, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tak:2013:CCC

[TUS13] Byung Chul Tak, Bhuvan Urgaonkar, and Anand Sivasubramaniam. Cloudy with a chance of cost savings. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1223–1233, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tumeo:2012:ACS

[TVCM12] Antonino Tumeo, Oreste Villa, and Daniel G. Chavarria-Miranda. Aho-corasick string matching on shared and distributed-memory parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 23(3):436– 443, March 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tan:2013:RSC

[TVG13] Yongmin Tan, Vinay Venkatesh, and Xiaohui Gu. Resilient self-compressive monitoring for large-scale hosting infrastructures. IEEE Transactions on Parallel and Distributed Systems, 24(3):576– 586, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tsakalozos:2017:LVM

[TVRD17] Konstantinos Tsakalozos, Vasilis Verroios, Mema Roussopoulos, and Alex Delis. Live VM migration under time-constraints in sharenothing IaaS-clouds. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2285–2298, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07833184-
abs.html.

Tripunitara:2014:CKM

[TW14] Mahesh V. Tripunitara and Jeffrey Lok Tin Woo. Composing Kerberos and Multimedia Internet KEYing (MIKEY) for authenticated transport of group keys. IEEE Transactions on Parallel and Distributed Systems, 25(4):898–907, April 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2012:GRC

[TWL12] Shyue-Ming Tang, Yue-Li Wang, and Chien-Yi Li. Generalized recursive circulant graphs. IEEE Transactions on Parallel and Distributed Systems, 23(1):87– 93, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2015:ALP

[TWL^{+}15] Zaiyang Tang, Zirui Wang, Peng Li, Song Guo, Xiaofei Liao, and Hai Jin. An application layer protocol for energy-efficient bandwidth aggregation with guaranteed quality-of-experience. IEEE Transactions on Parallel and Distributed Systems, 26(6):1538–1546, June 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06825847-
abs.html.

Tang:2016:DSA

[TWL16] Guoming Tang, Kui Wu, and Jingsheng Lei. A distributed and scalable approach to semi-intrusive load monitoring. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1553–1565, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07210212-
abs.html.

Tang:2017:ODB

[TWSW17] Qi Tang, Shang-Feng Wu, Jun-Wu Shi, and Ji-Bo Wei. Optimization of duplication-based schedules on network-on-chip based multi-processor system-on-chips. IEEE Transactions on Parallel and Distributed Systems, 28(3):826– 837, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07539657-
abs.html.

Tan:2016:POO

[TWT16] Xiaoqi Tan, Yuan Wu, and Danny H. K. Tsang. Pareto optimal operation of distributed battery energy storage systems for energy arbitrage under dynamic pricing. IEEE Transactions on Parallel and Distributed Systems, 27(7):2103–2115, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/07/07265089-
                                        abs.html.

Tan:2015:LBH

[TWW^{+}15] Feng Tan, Yufei Wang, Qixin Wang, Lei Bu, and Neeraj Suri. A lease based hybrid design pattern for propertemporal-embedding of wireless CPS interlocking. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2630–2642, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06898840.
                                        pdf.

Tan:2018:IRP

[TWW^{+}18] Yujuan Tan, Baiping Wang, Jian Wen, Zhichao Yan, Hong Jiang, and Witawas Srisa-an. Improving restore performance in deduplicationbased backup systems via a fine-grained defragmentation approach. IEEE Transactions on Parallel and Distributed Systems, 29(10): 2254–2267, October 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/10/08344493-
                                        abs.html.

Tong:2011:NRR

[TWZW11] Bin Tong, Guiling (Grace) Wang, Wensheng Zhang, and Chuang Wang. Node reclamation and replacement for long-lived sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(9):1550–1563, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2014:PSG

[TXL^{+}14] Shaohua Tang, Lingling Xu, Niu Liu, Xinyi Huang, Jintai Ding, and Zhiming Yang. Provably secure group key management approach based upon hyper-sphere. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3253–3263, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714432-
abs.html.

Tan:2011:PAR

[TXWL11] Rui Tan, Guoliang Xing, Jianping Wang, and Benyuan Liu. Performance analysis of real-time detection in fusionbased sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(9): 1564–1577, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2011:LMI

[TXZ^{+}11] Yuzhe Tang, Jianliang Xu, Shuigeng Zhou, Wang-Chien Lee, Dingxiong Deng, and Yue Wang. A lightweight multidimensional index for complex queries over DHTs. IEEE Transactions on Parallel and Distributed Systems, 22(12):2046–2054, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2014:OOI

[TYG^{+}14] Bin Tang, Baoliu Ye, Song Guo, Sanglu Lu, and Dapeng Oliver Wu. Orderoptimal information dissemination in MANETs via network coding. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1841–1851, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2013:CAP

[TYLG13] Bin Tang, Baoliu Ye, Sanglu Lu, and Song Guo. Coding-aware proportionalfair scheduling in OFDMA relay networks. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1727–1740, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2012:EEP

[TYS^{+}12] Shanjiang Tang, Ce Yu, Jizhou Sun, Bu-Sung Lee, Tao Zhang, Zhen Xu, and Huabei Wu. EasyPDP: An efficient parallel dynamic programming runtime system for computational biology. IEEE Transactions on Parallel and Distributed Systems, 23(5):862– 872, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2014:FAE

[TYWL14] Shaojie Tang, Jing Yuan, Cheng Wang, and Xiang-Yang Li. A framework for Amazon EC2 bidding strategy under SLA constraints. IEEE Transactions on Parallel and Distributed Systems, 25(1):2–11, January 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tang:2010:USI

[TZ10] Xueyan Tang and Suiping Zhou. Update scheduling for improving consistency in distributed virtual environments. IEEE Transactions on Parallel and Distributed Systems, 21(6): 765–777, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Teng:2014:IDP

[TZB^{+}14] Jin Teng, Boying Zhang, Xiaole Bai, Zhimin Yang, and Dong Xuan. Incentive-driven and privacy-preserving message dissemination in largescale mobile networks. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2909–2919, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06532281-
abs.html.

Tang:2019:DNF

[TZC19] Hong Tang, Danny Zhou, and Duan Chen. Dynamic network function instance scaling based on traffic forecasting and VNF placement in operator data centers. IEEE Transactions on Parallel and Distributed Systems, 30(3):530–543, March 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08449985-
abs.html.

Tan:2016:AIC

[TZT^{+}16] Guangming Tan, Chunming Zhang, Wen Tang, Peiheng Zhang, and Ninghui Sun. Accelerating irregular computation in massive short reads mapping on FPGA co-processor. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1253–1264, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2016/05/07122363-
abs.html.

Tang:2018:EMC

[TZY^{+}18] Xiongchao Tang, Jidong Zhai, Bowen Yu, Wenguang Chen, Weimin Zheng, and Keqin Li. An efficient inmemory checkpoint method and its practice on faulttolerant HPL. IEEE Transactions on Parallel and Distributed Systems, 29(4):758– 771, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08170311-
abs.html.

Uluagac:2013:SSB

[UBC13] A. Selcuk Uluagac, R. A.

Beyah, and J. A. Copeland. Secure SOurce-BAsed Loose Synchronization (SOBAS) for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(4):803–813, April 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Unat:2017:TDL

[UDH^{+}17] Didem Unat, Anshu Dubey, Torsten Hoefler, John Shalf, Mark Abraham, Mauro Bianco, Bradford L. Chamberlain, Romain Cledat, H. Carter Edwards, Hal Finkel, Karl Fuerlinger, Frank Hannig, Emmanuel Jeannot, Amir Kamil, Jeff Keasler, Paul H. J. Kelly, Vitus Leung, Hatem Ltaief, Naoya Maruyama, Chris J. Newburn, and Miquel Pericas. Trends in data locality abstractions for HPC systems. IEEE Transactions on Parallel and Distributed Systems, 28(10): 3007–3020, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07924389-
abs.html.

Wasi-ur-Rahman:2017:CSM

[uRILP17] Md. Wasi ur Rahman, Nusrat Sharmin Islam, Xiaoyi Lu, and Dhabaleswar K. DK Panda. A comprehensive study of MapReduce over Lustre for intermediate data placement and shuffle strategies on HPC clusters. IEEE Transactions on Parallel and Distributed Systems, 28(3): 633–646, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07514945-
abs.html.

Unal:2016:ASN

[US16] Ecem Unal and Erkay Savas. On acceleration and scalability of number theoretic private information retrieval. IEEE Transactions on Parallel and Distributed Systems, 27(6):1727–1741, June 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07155582-
abs.html.

Ubal:2012:SCM

[USP^{+}12] Rafael Ubal, Julio Sahuquillo, Salvador Petit, Pedro Lopez, and David R. Kaeli. A sequentially consistent multiprocessor architecture for out-of-order retirement of instructions. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1361–1368, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

VanCutsem:2014:DSL

[Van14] Thierry Van Cutsem. Dynamic simulation of largescale power systems using a parallel Schur-complement-based decomposition method. IEEE Transactions on Parallel and Distributed Systems, 25(10):2561–2570, October 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06619387-
abs.html.

Villalba:2019:CTS

[VBC19] Alvaro Villalba, Josep Lluis Berral, and David Carrera. Constant-time sliding window framework with reduced memory footprint and efficient bulk evictions. IEEE Transactions on Parallel and Distributed Systems, 30(3): 486–500, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/03/08456588.
                                        pdf.

vanderLaan:2011:AWL

[vdLJR11] Wladimir J. van der Laan, Andrei C. Jalba, and Jos B. T. M. Roerdink. Accelerating wavelet lifting on graphics hardware using CUDA. IEEE Transactions on Parallel and Distributed Systems, 22(1):132– 146, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Venkatasubramanian:2014:ERA

[Ven14] Nalini Venkatasubramanian. Efficient and reliable application layer multicast for flash dissemination. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2571–2582, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06656803-
                                        abs.html.

Vujic:2010:APM

[VGMA10] Nikola Vujic, Marc Gonzalez, Xavier Martorell, and Eduard Ayguade. Automatic prefetch and modulo scheduling transformations for the Cell BE architecture. IEEE Transactions on Parallel and Distributed Systems, 21(4): 494–505, April 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Viswanathan:2016:MOA

[VLP16] Hariharasudhan Viswanathan, Eun Kyung Lee, and Dario Pompili. A multi-objective approach to real-time insitu processing of mobileapplication workflows. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3116–3130, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07422143-
abs.html.

Viswanathan:2015:UAA

[VLRP15] Hariharasudhan Viswanathan, Eun Kyung Lee, Ivan Rodero, and Dario Pompili. Uncertainty-aware autonomic resource provisioning for mobile cloud computing. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2363–2372, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2015/08/06869019-
                                        abs.html.

Vejarano:2012:SAR

[VM12] Gustavo Vejarano and Janise McNair. Stability analysis of reservation-based scheduling policies in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 23(4):760–767, April 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vogel:2017:LVM

[VMB17] Pirmin Vogel, Andrea Marongiu, and Luca Benini. Lightweight virtual memory support for zero-copy sharing of pointerrich data structures in heterogeneous embedded SoCs. IEEE Transactions on Parallel and Distributed Systems, 28(7):1947–1959, July 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/07/07797491-
                                        abs.html.

Veeraragavan:2016:MQD

[VMN^{+}16] Narasimha Raghavan Veeraragavan, Leonardo Montecchi, Nicola Nostro, Roman Vitenberg, Hein Meling, and Andrea Bondavalli. Modeling QoE in dependable tele-immersive applications: A case study of world opera. IEEE Transactions on Parallel and Distributed Systems, 27(9):

2667–2681, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07336533-
abs.html.

Vanek:2017:GAO

[VMP17] Jan Vanek, Josef Michalek, and Josef Psutka. A GPUarchitecture optimized hierarchical decomposition algorithm for support vector machine training. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3330–3343, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07990554-
abs.html.

Vilches:2016:MSA

[VNA^{+}16] Antonio Vilches, Angeles Navarro, Rafael Asenjo, Francisco Corbera, Ruben Gran, and Maria J. Garzaran. Mapping streaming applications on commodity multi-CPU and GPU on-chip processors. IEEE Transactions on Parallel and Distributed Systems, 27(4): 1099–1115, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07106505-
abs.html.

Vargas-Perez:2017:HMO

[VPS17] Sandino Vargas-Perez and Fahad Saeed. A hybrid MPI–OpenMP strategy to speedup the compression of big next-generation sequencing datasets. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2760–2769, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07895161-
                                        abs.html.

Verbeek:2011:CNS

[VS11a] Freek Verbeek and Julien Schmaltz. A comment on “A Necessary and Sufficient Condition for Deadlock-Free Adaptive Routing in Wormhole Networks”. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1775–1776, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic). See [Dua93, Dua95].

Verbeek:2011:NSC

[VS11b] Freek Verbeek and Julien Schmaltz. On necessary and sufficient conditions for deadlock-free routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2022–2032, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Verbeek:2014:DPD

[VS14] Freek Verbeek and Julien Schmaltz. A decision procedure for deadlock-free routing in wormhole networks. IEEE Transactions on Parallel and Distributed Systems, 25(8):1935–1944, August 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Venugopalan:2015:IFO

[VS15] Sarad Venugopalan and Oliver Sinnen. ILP formulations for optimal task scheduling with communication delays on parallel systems. IEEE Transactions on Parallel and Distributed Systems, 26(1):142– 151, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748080-
abs.html.

Varshney:2019:ARC

[VS19] P. Varshney and Y. Simmhan.

AutoBoT: Resilient and costeffective scheduling of a bag of tasks on spot VMs. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1512–1527, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vu:2019:FMT

[VT19] V. A. Vu and G. Tan. A framework for mesoscopic traffic simulation in GPU. IEEE Transactions on Parallel and Distributed Systems, 30(8):1691–1703, August 2019. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Villa:2012:FAS

[VTSM12] Oreste Villa, Antonino Tumeo, Simone Secchi, and Joseph B. Manzano. Fast and accurate simulation of the Cray XMT multithreaded supercomputer. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2266–2279, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Vejarano:2014:DTM

[VWDM14] Gustavo Vejarano, Dexiang Wang, Ritwik Dubey, and Janise McNair. Distributed throughput maximization in wireless networks using the stability region. IEEE Transactions on Parallel and Distributed Systems, 25(7):1713–1723, July 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:HEC

[Wan12] Dajin Wang. Hamiltonian embedding in crossed cubes with failed links. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2117–2124, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:ECT

[Wan14] Xinbing Wang. Evolutioncast: Temporal evolution in wireless social networks and its impact on capacity. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2583–2594, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06604392-
abs.html.

Wang:2019:ABR

[Wan19] L. Wang. Architecture-based reliability-sensitive criticality measure for fault-tolerance cloud applications. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2408–2421, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:SID

[WBPF11] Xinbing Wang, Yuanzhe Bei, Qiuyu Peng, and Luoyi Fu. Speed improves delaycapacity trade-off in MotionCast. IEEE Transactions on Parallel and Distributed Systems, 22(5):729– 742, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:MIW

[WCD^{+}11] Shengling Wang, Yong Cui, Sajal K. Das, Wei Li, and Jianping Wu. Mobility in IPv6: Whether and how to hierarchize the network? IEEE Transactions on Parallel and Distributed Systems, 22(10):1722–1729, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2015:SAS

[WCD^{+}15] Song Wu, Haibao Chen, Sheng Di, Bingbing Zhou, Zhenjiang Xie, Hai Jin, and Xuanhua Shi. Synchronizationaware scheduling for virtual clusters in cloud. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2890–2902, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/10/06902819.
                                        pdf.

Wang:2010:MPC

[WCF10] Xiaorui Wang, Ming Chen, and Xing Fu. MIMO power control for highdensity servers in an enclosure. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1412–1426, October 2010.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2013:DPE

[WCF13] Weigang Wu, Jiannong Cao, and Xiaopeng Fan. Design and performance evaluation of overhearing-aided data caching in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 24(3):450– 463, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:SSH

[WCLK12] Xiaorui Wang, Ming Chen, Charles Lefurgy, and Tom W. Keller. SHIP: a scalable hierarchical power control architecture for large-scale data centers. IEEE Transactions on Parallel and Distributed Systems, 23(1):168– 176, January 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:ESE

[WCRL12] Cong Wang, Ning Cao, Kui Ren, and Wenjing Lou. Enabling secure and efficient ranked keyword search over outsourced cloud data. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1467–1479, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2019:EPC

[WCYL19] Yi Wang, Weixuan Chen, Jing Yang, and Tao Li. Exploiting parallelism for CNN applications on 3D stacked processing-in-memory architecture. IEEE Transactions on Parallel and Distributed Systems, 30(3):589– 600, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08452983-
abs.html.

Wang:2019:ICCa

[WCZ^{+}19a] Qingyang Wang, Hui Chen, Shungeng Zhang, Liting Hu, and Balaji Palanisamy. Integrating concurrency control in n-tier application scaling management in the cloud. IEEE Transactions on Parallel and Distributed Systems, 30(4):855–869, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8468089/

.

Wang:2019:DAN

[WCZ^{+}19b] Shaoqi Wang, Wei Chen, Xiaobo Zhou, Liqiang Zhang, and Yin Wang. Dependencyaware network adaptive scheduling of data-intensive parallel jobs. IEEE Transactions on Parallel and Distributed Systems, 30(3):515– 529, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08447511-
abs.html.

Wu:2012:TFN

[WDC12] Wenji Wu, Phil DeMar, and Matt Crawford. A transportfriendly NIC for multicore/multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 23(4):607–615, April 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2016:VPR

[WDH^{+}16] Hengfeng Wei, Marzio De Biasi, Yu Huang, Jiannong Cao, and Jian Lu. Verifying pipelined-RAM consistency over read/write traces of data replicas. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1511–1523, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07152941-
abs.html.

Wu:2017:MOM

[WDL^{+}17] Chen Wu, Chenchen Deng, Leibo Liu, Jie Han, Jiqiang Chen, Shouyi Yin, and Shaojun Wei. A multi-objective model oriented mapping approach for NoC-based computing systems. IEEE Transactions on Parallel and Distributed Systems, 28(3):662– 676, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07508918-
abs.html.

Wei:2015:CCA

[WDOX15] Kaimin Wei, Mianxiong Dong, Kaoru Ota, and Ke Xu. CAMF: Contextaware message forwarding in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 26(8):2178–2187, August 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06873340-
abs.html.

Wang:2013:GVU

[WFA13] Yun Wang, Weihuang Fu, and Dharma P. Agrawal. Gaussian versus uniform distribution for intrusion detection in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):342– 355, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:CPD

[WFK^{+}12] Chuang Wang, Taiming Feng, Jinsook Kim, Guiling Wang, and Wensheng Zhang. Catching packet droppers and modifiers in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(5): 835–843, May 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2017:INV

[WFZ^{+}17] Jing Wang, Xin Fu, Weigong Zhang, Junwei Zhang, Keni Qiu, and Tao Li. On the implication of NTC versus dark silicon on emerging scale-out workloads: The multi-core architecture perspective. IEEE Transactions on Parallel and Distributed Systems, 28(8): 2314–2327, August 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/08/07823012-
abs.html.

Wei:2013:LLV

[WG13] Yawen Wei and Yong Guan.

Lightweight location verification algorithms for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(5):

938–950, May 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2018:MLT

[WGCG18] Hongwei Wang, Song Guo, Jiannong Cao, and Minyi Guo. MeLoDy: A longterm dynamic quality-aware incentive mechanism for crowdsourcing. IEEE Transactions on Parallel and Distributed Systems, 29(4):901– 914, April 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/04/08115213-
abs.html.

Wang:2018:FTF

[WGG^{+}18] Zhigang Wang, Lixin Gao, Yu Gu, Yubin Bao, and Ge Yu. A fault-tolerant framework for asynchronous iterative computations in cloud environments. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1678–1692, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08300647-
abs.html.

Weber:2011:CHA

[WGHP11] Rick Weber, Akila Gothandaraman, Robert J. Hinde, and Gregory D. Peterson. Comparing hardware accelerators in scientific applications: a case study. IEEE Transactions on Parallel and Distributed Systems, 22(1):58– 68, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2016:GFP

[WGZ16] Chao Wang, Zonghua Gu, and Haibo Zeng. Global fixed priority scheduling with preemption threshold: Schedulability analysis and stack size minimization. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3242–3255, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07405349-
abs.html.

Waidyasooriya:2016:HAS

[WH16] Hasitha Muthumala Waidyasooriya and Masanori Hariyama. Hardware-acceleration of short-read alignment based on the Burrows–Wheeler transform. IEEE Transactions on Parallel and Distributed Systems, 27(5): 1358–1372, May 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07122348-
abs.html.

Wang:2016:DBS

[WHB16] Changda Wang, Syed Rafiul Hussain, and Elisa Bertino. Dictionary based secure provenance compression for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 27(2):405–418, February 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07038199-
                                        abs.html.

Wu:2014:HEM

[WHC^{+}14] Chentao Wu, Xubin He, Qiang Cao, Changsheng Xie, and Shenggang Wan. Hint-K: An efficient multilevel cache using K-step hints. IEEE Transactions on Parallel and Distributed Systems, 25(3):653–662, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2019:CDC

[WHF^{+}19] Yang Wang, Shuibing He, Xiaopeng Fan, Chengzhong Xu, and Xian-He Sun. On cost-driven collaborative data caching: A new model approach. IEEE Transactions on Parallel and Distributed Systems, 30(3):662– 676, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08454824-
abs.html.

Wang:2017:OCS

[WHGS17] Zhuoyao Wang, Majeed M.

Hayat, Nasir Ghani, and Khaled B. Shaban. Optimizing cloud-service performance: Efficient resource provisioning via optimal workload allocation. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1689–1702, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07742971-
abs.html.

Wan:2013:EPA

[WHH^{+}13] Shenggang Wan, Xubin He, Jianzhong Huang, Qiang Cao, Shiyi Li, and Changsheng Xie. An efficient penalty-aware cache to improve the performance of parity-based disk arrays under faulty conditions. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1500–1513, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2019:LDP

[WHN^{+}19] S. Wang, L. Huang, Y. Nie, X. Zhang, P. Wang, H. Xu, and W. Yang. Local differential private data aggregation for discrete distribution estimation. IEEE Transactions on Parallel and Distributed Systems, 30(9): 2046–2059, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2010:AWP

[WHYZ10] Yongwei Wu, Kai Hwang, Yulai Yuan, and Weimin Zheng. Adaptive workload prediction of grid performance in confidence windows. IEEE Transactions on Parallel and Distributed Systems, 21(7):925–938, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2019:IRP

[WHZS19] Jie Wu, Yu Hua, Pengfei Zuo, and Yuanyuan Sun. Improving Restore performance in deduplication systems via a cost-efficient rewriting scheme. IEEE Transactions on Parallel and Distributed Systems, 30(1):119– 132, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2019/01/08402122-
                                        abs.html.

Wu:2017:DCC

[WIZ^{+}17] Quanwang Wu, Fuyuki Ishikawa, Qingsheng Zhu, Yunni Xia, and Junhao Wen. Deadline-constrained cost optimization approaches for workflow scheduling in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3401–3412, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/08000634-
abs.html.

Wei:2012:OHT

[WJ12] Zheng Wei and Joseph JaJa. An optimized highthroughput strategy for constructing inverted files. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2033–2044, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2014:OFB

[WJB14] Jing Wu, Joseph JaJa, and Elias Balaras. An optimized FFT-based direct Poisson solver on CUDA GPUs. IEEE Transactions on Parallel and Distributed Systems, 25(3):550–559, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:SSD

[WJTL12] Cheng Wang, Changjun Jiang, Shaojie Tang, and Xiang-Yang Li. SelectCast: Scalable data aggregation scheme in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1958–1969, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:SLC

[WJTL13] Cheng Wang, Changjun Jiang, Shaojie Tang, and Xiang-Yang Li. Scaling laws of cognitive ad hoc networks over general primary network models. IEEE Transactions on Parallel and Distributed Systems, 24(5): 1030–1041, May 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:CBB

[WJTZ14] Chonggang Wang, Hongbo Jiang, Guang Tan, and Shengkai Zhang. Connectivitybased boundary extraction of large-scale 3D sensor networks: Algorithm and applications. IEEE Transactions on Parallel and Distributed Systems, 25(4):908– 918, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:FGF

[WJWX14] Guojun Wang, Wenjun Jiang, Jie Wu, and Zhengli Xiong. Fine-grained featurebased social influence evaluation in online social networks. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2286–2296, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06517181-
abs.html.

Wen:2014:STC

[WJX^{+}14] Sheng Wen, Jiaojiao Jiang, Yang Xiang, Shui Yu, Wanlei Zhou, and Weijia Jia. To shut them up or to clarify: Restraining the spread of rumors in online social networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3306–3316, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714492-
abs.html.

Warneke:2011:EDR

[WK11] Daniel Warneke and Odej Kao. Exploiting dynamic resource allocation for efficient parallel data processing in the cloud. IEEE Transactions on Parallel and Distributed Systems, 22(6): 985–997, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2012:PEP

[WKC12] Changjun Wu, Ananth Kalyanaraman, and William R.

Cannon. pGraph: Efficient parallel construction of large-scale protein sequence homology graphs. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1923–1933, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:UBR

[WKK11] Huaping Wang, Israel Koren, and C. Mani Krishna. Utilization-based resource partitioning for power-performance efficiency in SMT processors. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1150–1163, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wojciechowski:2017:SMD

[WKK17] Pawel T. Wojciechowski, Tadeusz Kobus, and Maciej Kokocinski. State-machine and deferred-update replication: Analysis and comparison. IEEE Transactions on Parallel and Distributed Systems, 28(3):891– 904, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/03/07508994-
                                        abs.html.

Wang:2016:EDT

[WKL^{+}16] Ke Wang, Abhishek Kulkarni, Michael Lang, Dorian Arnold, and Ioan Raicu. Exploring the design tradeoffs for extreme-scale highperformance computing system software. IEEE Transactions on Parallel and Distributed Systems, 27(4): 1070–1084, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07103354-
abs.html.

Wang:2016:PAC

[WKW16] Hongbing Wang, Zuling Kang, and Lei Wang. Performance-aware cloud resource allocation via fitnessenabled auction. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1160–1173, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07094282-
abs.html.

Wu:2019:EAP

[WKW19] H. Wu, W. J. Knottenbelt, and K. Wolter. An efficient application partitioning algorithm in mobile environments. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1464–1480, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:EPP

[WL12a] Haiyang Wang and Jiangchuan Liu. Exploring peer-to-peer locality in multiple torrent environment. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1216–1226, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2012:EOR

[WL12b] Weijie Wu and John C. S. Lui. Exploring the optimal replication strategy in P2P-VoD systems: Characterization and evaluation. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1492–1503, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:PFQ

[WL13] Hao Wang and Bill Lin. Per-flow queue management with succinct priority indexing structures for high speed packet scheduling. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1380–1389, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:NCA

[WL14] Xiaoyan Wang and Jie Li. Network coding aware cooperative MAC protocol for wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 25(1):167– 179, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:INL

[WL15] Xiaoyan Wang and Jie Li.

Improving the network lifetime of MANETs through cooperative MAC protocol design. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1010–1020, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06497046-
abs.html.

Wang:2017:SOA

[WLC^{+}17] Chao Wang, Xi Li, Yunji Chen, Youhui Zhang, Oliver Diessel, and Xuehai Zhou. Service-oriented architecture on FPGA-based MPSoC. IEEE Transactions on Parallel and Distributed Systems, 28(10):2993–3006, October 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07920399-
abs.html.

Wang:2015:DIP

[WLH^{+}15] Yin Wang, Yunhao Liu, Yuan He, Xiang-Yang Li, and Dapeng Cheng. Disco: Improving packet delivery via deliberate synchronized constructive interference. IEEE Transactions on Parallel and Distributed Systems, 26(3):713–723, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06774898-
abs.html.

Wang:2013:SNN

[WLL^{+}13] Jiliang Wang, Zhenjiang Li, Mo Li, Yunhao Liu, and Zheng Yang. Sensor network navigation without locations. IEEE Transactions on Parallel and Distributed Systems, 24(7):1436–1446, July 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:MRF

[WLL15a] Wei Wang, Ben Liang, and Baochun Li. Multi-resource fair allocation in heterogeneous cloud computing systems. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2822–2835, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06919321.
pdf.

Wang:2015:OOM

[WLL15b] Wei Wang, Ben Liang, and Baochun Li. Optimal online multi-instance acquisition in IaaS clouds. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3407–3419, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06995955-
abs.html.

Wei:2019:PPE

[WLL^{+}19] X. Wei, L. Li, X. Li, X. Wang, S. Gao, and H. Li. Pec: Proactive elastic collaborative resource scheduling in data stream processing. IEEE Transactions on Parallel and Distributed Systems, 30(7):1628–1642, July 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2010:EEW

[WLLL10] Yanwei Wu, Xiang-Yang Li, YunHao Liu, and Wei Lou. Energy-efficient wakeup scheduling for data collection and aggregation. IEEE Transactions on Parallel and Distributed Systems, 21(2): 275–287, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:EEL

[WLS^{+}11] Yu Wang, Xiang-Yang Li, Wen-Zhan Song, Minsu Huang, and Teresa A. Dahlberg. Energy-efficient localized routing in random multihop wireless networks. IEEE Transactions on Parallel and Distributed Systems, 22(8):1249–1257, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:EEL

[WLT^{+}12] Shie-Yuan Wang, Chih-Che Lin, Yan-Shiun Tzeng, Wen-Gao Huang, and Tin-Wei Ho. Exploiting event-level parallelism for parallel network simulation on multicore systems. IEEE Transactions on Parallel and Distributed Systems, 23(4):659– 667, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:RSC

[WLX13] Hao Wang, Bill Lin, and Jun Jim Xu. Robust statistics counter arrays with interleaved memories. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1894–1907, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2015:FFP

[WLX^{+}15] Chenggang Wu, Jin Li, Di Xu, Pen-Chung Yew, Jianjun Li, and Zhenjiang Wang. FPS: A fair-progress process scheduling policy on shared-memory multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 26(2):444– 454, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06766707-
abs.html.

Wang:2015:GSE

[WM15] Yijie Wang and Xingkong Ma. A general scalable and elastic content-based publish/subscribe service. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2100–2113, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06876150-
abs.html.

Wu:2018:ASM

[WM18] Yun Wu and John McAllister. Architectural synthesis of multi-SIMD dataflow accelerators for FPGA. IEEE Transactions on Parallel and Distributed Systems, 29(1): 43–55, January 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08017591-
abs.html.

Wang:2019:SCB

[WMC^{+}19] P. Wang, J. Meng, J. Chen, T. Liu, Y. Zhan, W. Tsai, and Z. Jin. Smart contractbased negotiation for adaptive QoS-aware service composition. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1403–1420, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:NST

[WMGA15] Guojun Wang, Felix Musau, Song Guo, and Muhammad Bashir Abdullahi. Neighbor similarity trust against Sybil attack in P2P ecommerce. IEEE Transactions on Parallel and Distributed Systems, 26(3):824– 833, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06776524-
                                        abs.html.

Wang:2012:AAD

[WMHX12] Chao Wang, Huadong Ma, Yuan He, and Shuguagn Xiong. Adaptive approximate data collection for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1004–1016, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2019:PIR

[WMJ^{+}19] S. Wu, B. Mao, H. Jiang, H. Luan, and J. Zhou. PFP: Improving the reliability of deduplication-based storage systems with per-file parity. IEEE Transactions on Parallel and Distributed Systems, 30(9):2117–2129, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2014:DCR

[WML14] Weijie Wu, Richard T. B.

Ma, and John C. S. Lui. Distributed caching via rewarding: An incentive scheme design in P2P–VoD systems. IEEE Transactions on Parallel and Distributed Systems, 25(3):612–621, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:BND

[WML15] Keyu Wang, Xufei Mao, and Yunhao Liu. BlindDate: A neighbor discovery protocol. IEEE Transactions on Parallel and Distributed Systems, 26(4):949–959, April 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06784517-
abs.html.

Wu:2012:MAC

[WMLJ12] Yulei Wu, Geyong Min, Keqiu Li, and Bahman Javadi. Modeling and analysis of communication networks in multicluster systems under spatio-temporal bursty traffic. IEEE Transactions on Parallel and Distributed Systems, 23(5):902–912, May 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2017:IPF

[WMLJ17] Suzhen Wu, Bo Mao, Yanping Lin, and Hong Jiang. Improving performance for flash-based storage systems through GC-aware cache management. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2852–2865, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07895128-
abs.html.

Wu:2019:RRR

[WMS^{+}19] Y. Wu, T. Ma, M. Su, M. Zhang, K. Chen, and Z. Guo. RF-RPC: Remote fetching RPC paradigm for RDMA-enabled network. IEEE Transactions on Parallel and Distributed Systems, 30(7):1657–1671, July 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:MAC

[WMT^{+}11] ShiGuang Wang, XuFei Mao, Shao-Jie Tang, Xiang-Yang Li, JiZhong Zhao, and GuoJun Dai. On “Movement-Assisted Connectivity Restoration in Wireless Sensor and Actor Networks”. IEEE Transactions on Parallel and Distributed Systems, 22(4):687– 694, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:APC

[WMW11] Xiaorui Wang, Kai Ma, and Yefu Wang. Adaptive power control with online model estimation for chip multiprocessors. IEEE Transactions on Parallel and Distributed Systems, 22(10):1681–1696, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2019:MDR

[WMWW19] J. Wang, M. C. Meyer, Y. Wu, and Y. Wang. Maximum data-resolution efficiency for fog-computing supported spatial big data processing in disaster scenarios. IEEE Transactions on Parallel and Distributed Systems, 30(8):1826–1842, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:PFS

[WMZ^{+}15] Lizhe Wang, Yan Ma, Albert Y. Zomaya, Rajiv Ranjan, and Dan Chen. A parallel file system with application-aware data layout policies for massive remote sensing image processing in digital Earth. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1497–1508, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06824837-
abs.html.

Wang:2015:DCI

[WNLL15] Wei Wang, Di Niu, Ben Liang, and Baochun Li. Dynamic cloud instance acquisition via IaaS cloud brokerage. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1580–1593, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06819811-
abs.html.

Wang:2013:MMC

[WPKL13] Hao Wang, Shi Pu, Gabe Knezek, and Jyh-Charn Liu. MIN-MAX: A counter-based algorithm for regular expression matching. IEEE Transactions on Parallel and Distributed Systems, 24(1):92– 103, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2018:UAU

[WPMX18] Yijie Wang, Xiaoqiang Pei, Xingkong Ma, and Fangliang Xu. TA-Update: An adaptive update scheme with tree-structured transmission in erasure-coded storage systems. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1893–1906, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/07954661-
abs.html.

Wang:2010:EBD

[WPT10] You-Chiun Wang, Wen-Chih Peng, and Yu-Chee Tseng. Energy-balanced dispatch of mobile sensors in a hybrid wireless sensor network. IEEE Transactions on Parallel and Distributed Systems, 21(12):1836–1850, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2017:SEV

[WPT17] Xudong Wang, Yibo Pi, and Aimin Tang. Scheduling of electric vehicle charging via multi-server fair queueing. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3298–3312, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07937889-
abs.html.

Wan:2015:TTR

[WQZ^{+}15] Jiguang Wan, Xiaoyang Qu, Nannan Zhao, Jun Wang, and Changsheng Xie. Thin-RAID: Thinning down RAID array for energy conservation. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2903–2915, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06912991.
pdf.

Wang:2016:LAB

[WQZ^{+}16] Yu Wang, Weikang Qian, Shuchang Zhang, Xiaoyao Liang, and Bo Yuan. A learning algorithm for Bayesian networks and its efficient implementation on GPUs. IEEE Transactions on Parallel and Distributed Systems, 27(1):17–30, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07001096-
abs.html.

Watkins:2011:PSC

[WRB11] Lanier Watkins, William H.

Robinson, and Raheem A. Beyah. A passive solution to the CPU resource discovery problem in cluster grid networks. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2000–2007, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wong:2015:PAS

[WRL15] Alvaro Wong, Dolores Rexachs, and Emilio Luque. Parallel application signature for performance analysis and prediction. IEEE Transactions on Parallel and Distributed Systems, 26 (7):2009–2019, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06827943-
abs.html.

Wang:2013:HCS

[WRWW13] Cong Wang, Kui Ren, Jia Wang, and Qian Wang. Harnessing the cloud for securely outsourcing large-scale systems of linear equations. IEEE Transactions on Parallel and Distributed Systems, 24(6):1172–1181, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See comment [CL16a].

Wu:2014:CRS

[WS14] Yuan Wu and Wen-Zhan Song. Cooperative resource sharing and pricing for proactive dynamic spectrum access via Nash bargaining solution. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2804–2817, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06658752-
                                        abs.html.

Wang:2018:LSV

[WS18] Huijun Wang and Oliver Sinnen. List-scheduling versus cluster-scheduling. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1736–1749, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/08/08301529-
                                        abs.html.

Wang:2014:MAT

[WSC^{+}14] Yi Wang, Zili Shao, Henry C. B. Chan, Duo Liu, and Yong Guan. Memory-aware task scheduling with communication overhead minimization for streaming applications on bus-based multiprocessor system-on-chips. IEEE Transactions on Parallel and Distributed Systems, 25(7):1797–1807, July 2014.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wen:2019:EGE

[WSH^{+}19] Z. Wen, J. Shi, B. He, J. Chen, K. Ramamohanarao, and Q. Li. Exploiting GPUs for efficient gradient boosting decision tree training. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2706–2717, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:CSW

[WSL^{+}15] Cheng Wang, Lu Shao, Zhong Li, Lei Yang, Xiang-Yang Li, and Changjun Jiang. Capacity scaling of wireless social networks. IEEE Transactions on Parallel and Distributed Systems, 26(7):1839–1850, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06844853-
abs.html.

Wu:2015:CFR

[WSS15] Yanbo Wu, Hong Shen, and Quan Z. Sheng. A cloudfriendly RFID trajectory clustering algorithm in uncertain environments. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2075–2088, August 2015.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/08/06877686-
                                        abs.html.

Wu:2013:MOF

[WSSZ13] Yanbo Wu, Quan Z. Sheng, Hong Shen, and Sherali Zeadally. Modeling object flows from distributed and federated RFID data streams for efficient tracking and tracing. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2036–2045, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:EIS

[WSWY15] Zhi Wang, Lifeng Sun, Chuan Wu, and Shiqiang Yang. Enhancing Internetscale video service deployment using microblog-based prediction. IEEE Transactions on Parallel and Distributed Systems, 26(3):775– 785, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06779659-
                                        abs.html.

Wachowiak:2017:APS

[WTD17] Mark P. Wachowiak, Mitchell C. Timson, and David J. Du-Val. Adaptive particle swarm optimization with heterogeneous multicore parallelism and GPU acceleration. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2784–2793, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07886331-
abs.html.

Wang:2010:SSE

[WTL10] Haodong Wang, Chiu C.

Tan, and Qun Li. Snoogle: a search engine for pervasive environments. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1188–1202, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:COA

[WTL^{+}14] Shuai Wang, Guang Tan, Yunhuai Liu, Hongbo Jiang, and Tian He. Coding opportunity aware backbone metrics for broadcast in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 25(8): 1999–2009, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Waidyasooriya:2017:OBF

[WTTH17] Hasitha Muthumala Waidyasooriya, Yasuhiro Takei, Shunsuke Tatsumi, and Masanori Hariyama. OpenCLbased FPGA-platform for stencil computation and its optimization methodology. IEEE Transactions on Parallel and Distributed Systems, 28(5):1390–1402, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07582502-
abs.html.

Wang:2019:LSB

[WTXG19] Peng Wang, George Trimponias, Hong Xu, and Yanhui Geng. Luopan: Sampling-based load balancing in data center networks. IEEE Transactions on Parallel and Distributed Systems, 30(1):133– 145, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08417438-
abs.html.

Wu:2014:CWB

[Wu14] Dapeng Oliver Wu. Can we beat DDoS attacks in clouds? IEEE Transactions on Parallel and Distributed Systems, 25(9): 2245–2254, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06567859-
abs.html.

Wang:2017:DSP

[WUH^{+}17] Shiqiang Wang, Rahul Urgaonkar, Ting He, Kevin Chan, Murtaza Zafer, and Kin K. Leung. Dynamic service placement for mobile micro-clouds with predicted future costs. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1002–1016, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07557016-
abs.html.

Weiss:2010:LUD

[WUM10] Stephane Weiss, Pascal Urso, and Pascal Molli. Logoot-Undo: Distributed collaborative editing system on P2P networks. IEEE Transactions on Parallel and Distributed Systems, 21(8): 1162–1174, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2017:PCH

[WV17] Xiaoli Wang and Bharadwaj Veeravalli. Performance characterization on handling large-scale partitionable workloads on heterogeneous networked compute platforms. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2925–2938, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07896640-
                                        abs.html.

Wang:2019:DTR

[WVM19] X. Wang, B. Veeravalli, and H. Ma. On the design of a time, resource and energy efficient multi-installment large-scale workload scheduling strategy for networkbased compute platforms. IEEE Transactions on Parallel and Distributed Systems, 30(5):1120–1133, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:DSA

[WVT13] Yang Wang, B. Veeravalli, and Chen-Khong Tham. On data staging algorithms for shared data accesses in clouds. IEEE Transactions on Parallel and Distributed Systems, 24(4):825– 838, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:CPC

[WW11] Xiaorui Wang and Yefu Wang. Coordinating power control and performance management for virtualized server clusters. IEEE Transactions on Parallel and Distributed Systems, 22(2):245– 259, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2012:RTA

[WW12] Qiang Wu and Tilman Wolf.

Runtime task allocation in multicore packet processing systems. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1934–1943, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:VBR

[WW13] Yefu Wang and Xiaorui Wang. Virtual batching: Request batching for server energy conservation in virtualized data centers. IEEE Transactions on Parallel and Distributed Systems, 24(8):1695–1705, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2014:DMT

[WWCB14] Jie Wu, Guojun Wang, Jiannong Cao, and Md Zakirul Alam Bhuiyan. Detecting movements of a target using face tracking in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(4): 939–949, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:PPA

[WWCZ11] Yefu Wang, Xiaorui Wang, Ming Chen, and Xiaoyun Zhu. PARTIC: Power-aware response time control for virtualized Web servers. IEEE Transactions on Parallel and Distributed Systems, 22(2): 323–336, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:ARE

[WWH13] Lu Wang, Kaishun Wu, and Mounir Hamdi. Attached-RTS: Eliminating an exposed terminal problem in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1289–1299, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2017:SOE

[WWH^{+}17] Yunxiang Wu, Fang Wang, Yu Hua, Dan Feng, Yuchong Hu, Wei Tong, Jingning Liu, and Dan He. I/O stack optimization for efficient and scalable access in FCoEbased SAN storage. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2514–2526, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07882635-
abs.html.

Wang:2018:IIB

[WWJ^{+}18] Jianfei Wang, Qin Wang, Li Jiang, Chao Li, Xiaoyao Liang, and Naifeng Jing. IBOM: An integrated and balanced on-chip memory for high performance GPGPUs. IEEE Transactions on Parallel and Distributed Systems, 29(3):586–599, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8107587/.

Wang:2011:TAR

[WWL11] Feng Wang, Dan Wang, and Jiangchuan Liu. Trafficaware relay node deployment: Maximizing lifetime for data collection wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1415–1423, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:MOD

[WWL^{+}13] Jin Wang, Jianping Wang, Kejie Lu, Bin Xiao, and Naijie Gu. Modeling and optimal design of linear network coding for secure unicast with multiple streams. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2025–2035, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:ECP

[WWL^{+}15] Haiyang Wang, Feng Wang, Jiangchuan Liu, Dan Wang, and Justin Groen. Enabling customer-provided resources for cloud computing: Potentials, challenges, and implementation. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1874–1886, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/07/06857371-
                                        abs.html.

Wu:2017:AOA

[WWL^{+}17] Song Wu, Yihong Wang, Wei Luo, Sheng Di, Haibao Chen, Xiaolin Xu, Ran Zheng, and Hai Jin. ACStor: Optimizing access performance of virtual disk images in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2414–2427, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07866840-
                                        abs.html.

Wei:2014:HUS

[WWLJ14] Shaojun Wei, Fang Wang, Zhaolin Li, and Guoyue Jiang. A high-utilization scheduling schemeof stream programs on clustered VLIW stream architectures. IEEE Transactions on Parallel and Distributed Systems, 25(4): 840–850, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:GNG

[WWLX13] Xiumin Wang, Jianping Wang, Kejie Lu, and Yinlong Xu. GKAR: A novel geographic (K)-anycast routing for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(5):916–925, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2011:EPA

[WWR^{+}11] Qian Wang, Cong Wang, Kui Ren, Wenjing Lou, and Jin Li. Enabling public auditability and data dynamics for storage security in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 22(5): 847–859, May 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2019:CDS

[WWT^{+}19] Pinghui Wang, Xiangyu Wang, Jing Tao, Peng Zhang, and Xiaohong Guan. Continuously distinct sampling over centralized and distributed high speed data streams. IEEE Transactions on Parallel and Distributed Systems, 30(2):300– 314, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08436443-
abs.html.

Wang:2018:DRS

[WWW^{+}18] Xinhou Wang, Kezhi Wang, Song Wu, Sheng Di, Hai Jin, Kun Yang, and Shumao Ou. Dynamic resource scheduling in mobile edge cloud with cloud radio access network. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2429–2445, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08353131-
abs.html.

Wang:2013:ISP

[WWX^{+}13] Xiaodong Wang, Xiaorui Wang, Guoliang Xing, Jinzhu Chen, Cheng-Xian Lin, and Yixin Chen. Intelligent sensor placement for hot server detection in data centers. IEEE Transactions on Parallel and Distributed Systems, 24(8):1577–1588, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2016:CAT

[WWZ^{+}16] Xiaodong Wang, Xiaorui Wang, Kuangyu Zheng, Yanjun Yao, and Qing Cao. Correlation-aware traffic consolidation for power optimization of data center networks. IEEE Transactions on Parallel and Distributed Systems, 27(4):992– 1006, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07084121-
abs.html.

Wei:2011:MCP

[WX11] Jianbin Wei and Cheng-Zhong Xu. Measuring clientperceived pageview response time of Internet services. IEEE Transactions on Parallel and Distributed Systems, 22(5):773–785, May 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2015:PPC

[WX15] Wei Wang and Tao Xie.

PCFTL: A plane-centric flash translation layer utilizing copy-back operations. IEEE Transactions on Parallel and Distributed Systems, 26(12):3420–3432, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06957592-
abs.html.

Wang:2010:MCT

[WXL10] Feng Wang, Yongqiang Xiong, and Jiangchuan Liu. mTreebone: a collaborative tree-mesh overlay network for multicast video streaming. IEEE Transactions on Parallel and Distributed Systems, 21(3):379– 392, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2016:ESS

[WXLY16] Si Wu, Yinlong Xu, Yongkun Li, and Zhijia Yang. I/Oefficient scaling schemes for distributed storage systems with CRS codes. IEEE Transactions on Parallel and Distributed Systems, 27(9): 2639–2652, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/09/07347422-
                                        abs.html.

Wei:2013:SDM

[WXTL13] Wei Wei, Fengyuan Xu, Chiu C. Tan, and Qun Li. SybilDefender: A defense mechanism for Sybil attacks in large social networks. IEEE Transactions on Parallel and Distributed Systems, 24(12):2492–2502, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2013:CBI

[WXY^{+}13] Kaishun Wu, Jiang Xiao, Youwen Yi, Dihu Chen, Xiaonan Luo, and Lionel M. Ni. CSI-based indoor localization. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1300–1309, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2014:TES

[WXYX14] Xiaohu Wu, Yinlong Xu, Chau Yuen, and Liping Xiang. A tag encoding scheme against pollution attack to linear network coding. IEEE Transactions on Parallel and Distributed Systems, 25(1): 33–42, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:ITC

[WXZ^{+}14] Yu Wang, Yang Xiang, Jun Zhang, Wanlei Zhou, Guiyi Wei, and Laurence T. Yang. Internet traffic classification using constrained clustering. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2932–2943, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06684161-
abs.html.

Wu:2015:GAL

[WYC^{+}15] Jiyan Wu, Chau Yuen, Bo Cheng, Yanlei Shang, and Junliang Chen. Goodputaware load distribution for real-time traffic over multipath networks. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2286–2299, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/08/06876149-
                                        abs.html.

Wu:2014:MDF

[WYCZ14] Yongwei Wu, Feng Ye, Kang Chen, and Weimin Zheng. Modeling of distributed file systems for practical performance analysis. IEEE Transactions on Parallel and Distributed Systems, 25(1):156– 166, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2019:PRA

[WYL19] L. Wang, T. Ye, and T. T. Lee. A parallel route assignment algorithm for faulttolerant Clos networks in OTN switches. IEEE Transactions on Parallel and Distributed Systems, 30(5):977– 989, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2018:PCC

[WYLH18] Lingkang Wang, Tong Ye, Tony T. Lee, and Weisheng Hu. A parallel complex coloring algorithm for scheduling of input-queued switches. IEEE Transactions on Parallel and Distributed Systems, 29(7):1456–1468, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/07/08281531-
abs.html.

Wu:2013:WWI

[WYLX13] Chenshu Wu, Zheng Yang, Yunhao Liu, and Wei Xi. WILL: Wireless Indoor Localization without site survey. IEEE Transactions on Parallel and Distributed Systems, 24(4):839–848, April 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2013:AHB

[WYW13] Yunsheng Wang, Wei-Shih Yang, and Jie Wu. Analysis of a hypercube-based social feature multipath routing in delay tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 24(9):1706–1716, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:IX

[WYW^{+}14] Chen Wang, Baoliu Ye, Xiaoliang Wang, Song Guo, and Sanglu Lu. IEEE XPLORE. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2829–2839, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/11/06714437-
abs.html.

Wang:2013:LBN

[WYX13] Xinbing Wang, Tuo Yu, and Yuanzhong Xu. Lower bound for node buffer size in intermittently connected wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(4):754– 766, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2015:HME

[WYX^{+}15] Chenshu Wu, Zheng Yang, Yu Xu, Yiyang Zhao, and Yunhao Liu. Human mobility enhances global positioning accuracy for mobile phone localization. IEEE Transactions on Parallel and Distributed Systems, 26 (1):131–141, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748094-
abs.html.

Wang:2019:ICCb

[WYX^{+}19] H. Wang, X. Yu, H. Xu, J. Fan, C. Qiao, and L. Huang. Integrating coflow and circuit scheduling for optical networks. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1346–1358, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wei:2012:SPS

[WYY^{+}12] Haitao Wei, Junqing Yu, Huafei Yu, Mingkang Qin, and Guang R. Gao. Software pipelining for stream programs on resource constrained multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 23(12):2338–2350, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2019:PMG

[WYZ^{+}19] J. Wu, X. Yang, Z. Zhang, G. Chen, and R. Mao. A performance model for GPU architectures that considers on-chip resources: Application to medical image registration. IEEE Transactions on Parallel and Distributed Systems, 30(9):1947–1961, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:KSC

[WZ14] Huaqun Wang and Yuqing Zhang. On the knowledge soundness of a cooperative provable data possession scheme in multicloud storage. IEEE Transactions on Parallel and Distributed Systems, 25(1):264– 267, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2013:BTS

[WZFG13] Haifeng Wu, Yu Zeng, Jihua Feng, and Yu Gu. Binary tree slotted ALOHA for passive RFID tag anticollision. IEEE Transactions on Parallel and Distributed Systems, 24(1):19–31, January 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2010:SDA

[WZGR10] Qishi Wu, Mengxia Zhu, Yi Gu, and Nageswara S. V. Rao. System design and algorithmic development for computational steering in distributed environments. IEEE Transactions on Parallel and Distributed Systems, 21(4):438–451, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2016:MMF

[WZHZ16] Zeke Wang, Shuhao Zhang, Bingsheng He, and Wei Zhang. Melia: A MapReduce framework on OpenCLbased FPGAs. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3547–3560, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07425227-
abs.html.

Wang:2016:HIF

[WZL^{+}16] Chao Wang, Junneng Zhang, Xi Li, Aili Wang, and Xuehai Zhou. Hardware implementation on FPGA for task-level parallel dataflow execution engine. IEEE Transactions on Parallel and Distributed Systems, 27(8): 2303–2315, August 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/08/07289452-
abs.html.

Wu:2019:EFR

[WZL^{+}19] Chao Wu, Lan Zhang, Qiushi Li, Ziyan Fu, Wenwu Zhu, and Yaoxue Zhang. Enabling flexible resource allocation in mobile deep learning systems. IEEE Transactions on Parallel and Distributed Systems, 30(2):346– 360, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08434315-
abs.html.

Wu:2015:DME

[WZLC15] Weigang Wu, Jiebin Zhang, Aoxue Luo, and Jiannong Cao. Distributed mutual exclusion algorithms for intersection traffic control. IEEE Transactions on Parallel and Distributed Systems, 26(1):65–74, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06747396-
                                        abs.html.

Wu:2010:SIH

[WZQ10] Fan Wu, Sheng Zhong, and Chunming Qiao. Strongincentive, high-throughput channel assignment for noncooperative wireless networks. IEEE Transactions on Parallel and Distributed Systems, 21(12):1808–1821, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2014:DSE

[WZQY14] Jianjiang Wang, Xiaomin Zhu, Dishan Qiu, and Laurence T. Yang. Dynamic scheduling for emergency tasks on distributed imaging satellites with task merging. IEEE Transactions on Parallel and Distributed Systems, 25(9):2275–2285, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06532288-
abs.html.

Wang:2018:CSA

[WZS^{+}18] Qiang Wang, Hao Zhang, Jianfei Sun, Hu Xiong, and Zhiguang Qin. Comments on “A secure anti-collusion data sharing scheme for dynamic groups in the cloud”. Information Processing Letters, 140(??):30–33, December 2018. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019018300851.

See [ZJ16].

Wang:2019:ESW

[WZS^{+}19] Z. Wang, H. Zhang, X. Shi, X. Yin, Y. Li, H. Geng, Q. Wu, and J. Liu. Efficient scheduling of weighted coflows in data centers. IEEE Transactions on Parallel and Distributed Systems, 30(9): 2003–2017, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wang:2012:CCS

[WZSL12] Lei Wang, Jianfeng Zhan, Weisong Shi, and Yi Liang. In cloud, can scientific communities benefit from the economies of scale? IEEE Transactions on Parallel and Distributed Systems, 23(2):

296–303, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wen:2013:MPD

[WZZ^{+}13] Sheng Wen, Wei Zhou, Jun Zhang, Yang Xiang, Wanlei Zhou, and Weijia Jia. Modeling propagation dynamics of social network worms. IEEE Transactions on Parallel and Distributed Systems, 24(8):1633–1643, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2017:DCV

[XAG17] Gang Xu, George T. Amariucai, and Yong Guan. Delegation of computation with verification outsourcing: Curious verifiers. IEEE Transactions on Parallel and Distributed Systems, 28(3):717– 730, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/03/07533495-
                                        abs.html.

Xu:2017:PND

[XAK17] Xi Xu, Rashid Ansari, and Ashfaq Khokhar. Parallel nonuniform discrete Fourier transform (P-NDFT) over a random wireless sensor network. IEEE Transactions on Parallel and Distributed Systems, 28(12):

3615–3625, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07932188-
abs.html.

Xu:2017:CCR

[XALS17] Maotong Xu, Sultan Alamro, Tian Lan, and Suresh Subramaniam. CRED: Cloud right-sizing with execution deadlines and data locality. IEEE Transactions on Parallel and Distributed Systems, 28(12):3389–3400, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07976384-
abs.html.

Xia:2014:ESR

[XAY^{+}14] Feng Xia, Ahmedin Mohammed Ahmed, Laurence Tianruo Yang, Jianhua Ma, and Joel J. P. C. Rodrigues. Exploiting social relationship to enable efficient replica allocation in ad-hoc social networks. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3167–3176, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714465-
abs.html.

Xu:2014:ESM

[XAYM14] Quanqing Xu, Rajesh Vellore Arumugam, Khai Leong Yong, and Sridhar Mahadevan. Efficient and scalable metadata management in EB-scale file systems. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2840–2850, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06674929-
                                        abs.html.

Xie:2015:MVS

[XBL15] Hengheng Xie, Azzedine Boukerche, and Antonio A. F. Loureiro. A multipath video streaming solution for vehicular networks with link disjoint and nodedisjoint. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3223–3235, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/12/06957547-
                                        abs.html.

Xiao:2016:DRR

[XBZ^{+}16] Wenhua Xiao, Weidong Bao, Xiaomin Zhu, Chen Wang, Lidong Chen, and Laurence T. Yang. Dynamic request redirection and resource provisioning for cloudbased video services under heterogeneous environment. IEEE Transactions on Parallel and Distributed Systems, 27(7):1954–1967, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07217812-
abs.html.

Xiao:2017:CAB

[XBZL17] Wenhua Xiao, Weidong Bao, Xiaomin Zhu, and Ling Liu. Cost-aware big data processing across geo-distributed datacenters. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3114–3127, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07934132-
abs.html.

Xiong:2015:SCS

[XCZ^{+}15] Pengcheng Xiong, Yun Chi, Shenghuo Zhu, Hyun Jin Moon, Calton Pu, and Hakan Hacgumus. SmartSLA: Costsensitive management of virtualized resources for CPUbound database services. IEEE Transactions on Parallel and Distributed Systems, 26(5):1441–1451, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803071-
abs.html.

Xu:2019:OJS

[XDLZ19] Huanle Xu, Gustavo De Veciana, Wing Cheong Lau, and Kunxiao Zhou. Online job scheduling with redundancy and opportunistic checkpointing: A speedupfunction-based analysis. IEEE Transactions on Parallel and Distributed Systems, 30(4):897–909, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

                                        URL https://ieeexplore.
                                        ieee.org/document/8468118/

.

Xie:2017:ISC

[XDMZ17] Junjie Xie, Yuhui Deng, Geyong Min, and Yongtao Zhou. An incrementally scalable and cost-efficient interconnection structure for data centers. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1578–1592, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/06/07745911-
                                        abs.html.

Xu:2015:TAW

[XFL15] Hong Xu, Chen Feng, and Baochun Li. Temperature aware workload management in geo-distributed data centers. IEEE Transactions on Parallel and Distributed Systems, 26(6):

1743–1753, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06819031-
abs.html.

Xu:2019:OFV

[XFL^{+}19] J. Xu, H. Fu, W. Luk, L. Gan, W. Shi, W. Xue, C. Yang, Y. Jiang, C. He, and G. Yang. Optimizing finite volume method solvers on Nvidia GPUs. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2790–2805, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2016:EAS

[XGL^{+}16] Cong Xu, Robin Goldstone, Zhuo Liu, Hui Chen, Bryon Neitzel, and Weikuan Yu. Exploiting analytics shipping with virtualized MapReduce on HPC backend storage servers. IEEE Transactions on Parallel and Distributed Systems, 27(1):185– 196, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07004064-
abs.html.

Xu:2014:SDT

[XGZW14] Ke Xu, Song Guo, Deze Zeng, and Kaimin Wei. On social delay-tolerant networking: Aggregation, tie detection, and routing. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1563–1573, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2010:UPB

[XH10] Bin Xiao and Yu Hua. Using parallel Bloom filters for multiattribute representation on network services. IEEE Transactions on Parallel and Distributed Systems, 21(1):20–32, January 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2016:PLC

[XHC16] Yu Xia, Mounir Hamdi, and H. Jonathan Chao. A practical large-capacity threestage buffered Clos-network switch architecture. IEEE Transactions on Parallel and Distributed Systems, 27(2): 317–328, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07054539-
                                        abs.html.

Xie:2015:SBA

[XHG15] Miao Xie, Jiankun Hu, and Song Guo. Segment-based anomaly detection with approximated sample covariance matrix in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(2):574– 583, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06748064-
abs.html.

Xie:2013:SHN

[XHHC13] Miao Xie, Jiankun Hu, Song Han, and Hsiao-Hwa Chen. Scalable hypergrid k-nn-based online anomaly detection in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(8):1661–1670, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiong:2011:MDC

[XHL^{+}11] Jin Xiong, Yiming Hu, Guojie Li, Rongfeng Tang, and Zhihua Fan. Metadata distribution and consistency techniques for large-scale cluster file systems. IEEE Transactions on Parallel and Distributed Systems, 22(5): 803–816, May 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2015:EPC

[XHL^{+}15] Lei Xie, Hao Han, Qun Li, Jie Wu, and Sanglu Lu. Efficient protocols for collecting histograms in largescale RFID systems. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2421–2433, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06895296.
pdf.

Xu:2015:JVM

[XHQ^{+}15] Hongli Xu, Liusheng Huang, Chunming Qiao, Weichao Dai, and Yu e Sun. Joint virtual MIMO and data gathering for wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1034–1048, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06787070-
abs.html.

Xie:2013:MOB

[XHX^{+}13] Yi Xie, Jiankun Hu, Yang Xiang, Shui Yu, Shensheng Tang, and Yu Wang. Modeling oscillation behavior of network traffic by nested hidden Markov model with variable state-duration. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1807–1817, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xue:2013:SAA

[XHZ^{+}13] Guangtao Xue, Qi He, Hongzi Zhu, Tian He, and Yunhuai Liu. Sociality-aware access point selection in enterprise wireless LANs. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2069–2078, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2014:SED

[Xia14]  Qi Xia.  On the secu-

rity of an efficient dynamic auditing protocol in cloud storage. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2760–2761, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06577391-
abs.html.

Xie:2014:TEC

[XJ14] Ruitao Xie and Xiaohua Jia. Transmission-efficient clustering method for wireless sensor networks using compressive sensing. IEEE Transactions on Parallel and Distributed Systems, 25(3): 806–815, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xing:2014:OSS

[XJL^{+}14] Xiaoshuang Xing, Tao Jing, Hongjuan Li, Yan Huo, Xiuzhen Cheng, and Taieb Znati. Optimal spectrum sensing interval in cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2408–2417, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06522414-
abs.html.

Xiong:2010:DEF

[XJY^{+}10] Naixue Xiong, Xiaohua Jia, Laurence T. Yang, Athanasios V. Vasilakos, Yingshu Li, and Yi Pan. A distributed efficient flow control scheme for multirate multicast networks. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1254–1266, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2010:TPD

[XL10] Xiang Xiao and Jaehwan John Lee. A true O(1) parallel deadlock detection algorithm for single-unit resource systems and its hardware implementation. IEEE Transactions on Parallel and Distributed Systems, 21(1): 4–19, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2013:AVE

[XL13] Hong Xu and Baochun Li.

Anchor: A versatile and efficient framework for resource management in the cloud. IEEE Transactions on Parallel and Distributed Systems, 24(6):1066–1076, June 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiang:2016:DFB

[XL16] Dong Xiang and Xiaowei Liu. Deadlock-free broadcast routing in Dragonfly networks without virtual channels. IEEE Transactions on Parallel and Distributed Systems, 27(9):2520–2532, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07337415-
abs.html.

Xu:2017:OSE

[XL17] Huanle Xu and Wing Cheong Lau. Optimization for speculative execution in big data processing clusters. IEEE Transactions on Parallel and Distributed Systems, 28(2): 530–545, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07466828-
abs.html.

Xu:2015:HCR

[XLH^{+}15] Yuming Xu, Kenli Li, Ligang He, Longxin Zhang, and Keqin Li. A hybrid chemical reaction optimization scheme for task scheduling on heterogeneous computing systems. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3208–3222, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06995978-
abs.html.

Xu:2011:CRO

[XLL11] Jin Xu, Albert Y. S. Lam, and Victor O. K. Li. Chemical reaction optimization for task scheduling in grid computing. IEEE Transactions on Parallel and Distributed Systems, 22(10):1624–1631, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2018:EBT

[XLL^{+}18] Xiaokang Xie, Qing Ling, Ping Lu, Wei Xu, and Zuqing Zhu. Evacuate before too late: Distributed backup in inter-DC networks with progressive disasters. IEEE Transactions on Parallel and Distributed Systems, 29(5):1058–1074, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08227003-
abs.html.

Xu:2011:LLB

[XLLZ11] Ke Xu, Hongying Liu, Jiangchuan Liu, and Jixiu Zhang. LBMP: a logarithmbarrier-based multipath protocol for Internet traffic management. IEEE Transactions on Parallel and Distributed Systems, 22(3):476– 488, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2011:ACV

[XLM^{+}11a] Lei Xie, Qun Li, Weizhen Mao, Jie Wu, and Daoxu Chen. Association control for vehicular WiFi access: Pursuing efficiency and fairness. IEEE Transactions on Parallel and Distributed Systems, 22(8):1323–1331, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2011:DEA

[XLM^{+}11b] Xiaohua Xu, Xiang-Yang Li, XuFei Mao, Shaojie Tang, and Shiguang Wang. A delay-efficient algorithm for data aggregation in multihop wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(1):163– 175, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2012:REE

[XLM12a] Junfeng Xu, Keqiu Li, and Geyong Min. Reliable and energy-efficient multipath communications in underwater sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1326–1335, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2012:EET

[XLM^{+}12b] Junfeng Xu, Keqiu Li, Geyong Min, Kai Lin, and Wenyu Qu. Energy-efficient tree-based multipath power control for underwater sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(11): 2107–2116, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2013:MHB

[XLSR13] Yin Xu, Ben Leong, Daryl Seah, and Ali Razeen. mPath: High-bandwidth data transfers with massively multipath source routing. IEEE Transactions on Parallel and Distributed Systems, 24(10):2046–2059, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2014:AaL

[XLT^{+}14] Lei Xu, Fang Liu, Lei Tian, Nong Xiao, Hong Jiang, and Yinjin Fu. Application-aware local-global source deduplication for cloud backup services of personal storage. IEEE Transactions on Parallel and Distributed Systems, 25(5):1155–1165, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2016:EAC

[XLX^{+}16] Zichuan Xu, Weifa Liang, Wenzheng Xu, Mike Jia, and Song Guo. Efficient algorithms for capacitated cloudlet placements. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2866–2880, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07362036-
abs.html.

Xu:2017:MSV

[XLY^{+}17] Huanle Xu, Wing Cheong Lau, Zhibo Yang, Gustavo de Veciana, and Hanxu Hou. Mitigating service variability in MapReduce clusters via task cloning: A competitive analysis. IEEE Transactions on Parallel and Distributed Systems, 28(10):

2866–2880, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07890998-
                                        abs.html.

Xue:2018:SGV

[XML^{+}18] Mochi Xue, Jiacheng Ma, Wentai Li, Kun Tian, Yaozu Dong, Jinyu Wu, Zhengwei Qi, Bingsheng He, and Haibing Guan. Scalable GPU virtualization with dynamic sharing of graphics memory space. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1823–1836, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/08/08247267-
                                        abs.html.

Xia:2012:DEP

[XP12] Yinglong Xia and Viktor K. Prasanna. Distributed evidence propagation in junction trees on clusters. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1169–1177, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2014:EMA

[XQL^{+}14] Yujie Xu, Wenyu Qu, Zhiyang Li, Geyong Min, Keqiu Li, and Zhaobin Liu.

Efficient k-means++ approximation with MapReduce. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3135–3144, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06740074-
abs.html.

Xie:2010:DDR

[XS10] Tao Xie and Yao Sun. Dynamic data reallocation in hybrid disk arrays. IEEE Transactions on Parallel and Distributed Systems, 21(9): 1330–1341, September 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiang:2011:BAC

[XS11] Yonghong Xiang and Iain A.

Stewart. Bipancyclicity in k-ary n-cubes with faulty edges under a conditional fault assumption. IEEE Transactions on Parallel and Distributed Systems, 22(9): 1506–1513, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2013:DRA

[XSC13] Zhen Xiao, Weijia Song, and Qi Chen. Dynamic resource allocation using virtual machines for cloud computing environment. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1107–1117, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2016:AOT

[XSL^{+}16] Ke Xu, Meng Shen, Hongying Liu, Jiangchuan Liu, Fan Li, and Tong Li. Achieving optimal traffic engineering using a generalized routing framework. IEEE Transactions on Parallel and Distributed Systems, 27(1):51– 65, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/01/07010944-
                                        abs.html.

Xuan:2010:DAD

[XSTZ10] Ying Xuan, Incheol Shin, My T. Thai, and Taieb Znati. Detecting application denial-of-service attacks: a group-testing-based approach. IEEE Transactions on Parallel and Distributed Systems, 21(8):1203–1216, August 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2013:HAT

[XSYY13] Bin Xu, Guodong Sun, Ran Yu, and Zheng Yang. Highaccuracy TDOA-based localization without time synchronization. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1567–1576, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiao:2010:TSD

[XSZ^{+}10] Xin Xiao, Yuanchun Shi, Qian Zhang, Jianhua Shen, and Yuan Gao. Toward systematical data scheduling for layered streaming in peer-to-peer networks: Can we go farther? IEEE Transactions on Parallel and Distributed Systems, 21(5): 685–697, May 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2013:CDC

[XSZ13] Mingsen Xu, Wen-Zhan Song, and Yichuan Zhao. Collaborative data collection with opportunistic network erasure coding. IEEE Transactions on Parallel and Distributed Systems, 24(10): 1941–1950, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2017:LSV

[XTFC17] Ye Xia, Mauricio Tsugawa, Jose A. B. Fortes, and Shigang Chen. Large-scale VM placement with disk anti-colocation constraints using hierarchical decomposition and mixed integer programming. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1361–1374, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/05/07586124-
                                        abs.html.

Xu:2010:PCL

[XTHD10] Jianliang Xu, Xueyan Tang, Haibo Hu, and Jing Du. Privacy-conscious locationbased queries in mobile environments. IEEE Transactions on Parallel and Distributed Systems, 21(3):313– 326, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2013:RWP

[XTXH13] Yi Xie, S. Tang, Y. Xiang, and J. Hu. Resisting Web proxy-based HTTP attacks by temporal and spatial locality behavior. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1401–1410, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2017:RSO

[XVC17] Yadong Xu, Vaisagh Viswanathan, and Wentong Cai. Reducing synchronization overhead with computation replication in parallel agent-based road traffic simulation. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3286–3297, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07945281-
abs.html.

Xiao:2015:HBZ

[XWH15a] Mingjun Xiao, Jie Wu, and Liusheng Huang. Homebased zero-knowledge multicopy routing in mobile social networks. IEEE Transactions on Parallel and Distributed Systems, 26(5): 1238–1250, May 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06803962-
abs.html.

Xiao:2015:TSU

[XWH15b] Mingjun Xiao, Jie Wu, and Liusheng Huang. Timesensitive utility-based singlecopy routing in low-duty-cycle wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(5):1452–1465, May 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/05/06808518-
abs.html.

Xie:2015:SSV

[XWJX15] Ruitao Xie, Yonggang Wen, Xiaohua Jia, and Haiyong Xie. Supporting seamless virtual machine migration via named data networking in cloud data center. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3485–3497, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/12/06975186-
                                        abs.html.

Xu:2019:MBD

[XWL^{+}19] Chenhan Xu, Kun Wang, Peng Li, Song Guo, Jiangtao Luo, Baoliu Ye, and Minyi Guo. Making big data open in edges: A resource-efficient blockchain-based approach. IEEE Transactions on Parallel and Distributed Systems, 30(4):870–882, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

                                        URL https://ieeexplore.
                                        ieee.org/document/8469010/

.

Xu:2016:CCP

[XWLJ16] Jie Xu, Qiaoyan Wen, Wenmin Li, and Zhengping Jin. Circuit ciphertextpolicy attribute-based hybrid encryption with verifiable delegation in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 27(1):119– 129, January 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07010954-

abs.html. See comments [XWS17].

Xiong:2017:CCC

[XWS17] Hu Xiong, Qiang Wang, and Jianfei Sun. Comments on “Circuit ciphertext-policy attribute-based hybrid encryption with verifiable delegation”. Information Processing Letters, 127(??):67– 70, November 2017. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0020019017301308.

See [XWLJ16].

Xia:2016:SDM

[XWSW16] Zhihua Xia, Xinhui Wang, Xingming Sun, and Qian Wang. A secure and dynamic multi-keyword ranked search scheme over encrypted cloud data. IEEE Transactions on Parallel and Distributed Systems, 27(2):340– 352, February 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07039216-
abs.html.

Xing:2010:MSS

[XWY^{+}10] Guoliang Xing, Jianping Wang, Zhaohui Yuan, Rui Tan, Limin Sun, Qingfeng Huang, Xiaohua Jia, and Hing Cheung So. Mobile scheduling for spatiotemporal detection in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(12): 1851–1866, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2016:HPE

[XX16] Qianbin Xia and Weijun Xiao. High-performance and endurable cache management for flash-based read caching. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3518–3531, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/12/07425233-
                                        abs.html.

Xia:2019:EDP

[XXL^{+}19] Q. Xia, Z. Xu, W. Liang, S. Yu, S. Guo, and A. Y. Zomaya. Efficient data placement and replication for QoS-aware approximate query evaluation of big data analytics. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2677–2691, December 2019.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xia:2016:CFA

[XXLZ16] Qiufen Xia, Zichuan Xu, Weifa Liang, and Albert Y. Zomaya. Collaborationand fairness-aware big data management in distributed clouds. IEEE Transactions on Parallel and Distributed Systems, 27(7):1941–1953, July 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07225163-
abs.html.

Xu:2010:FDA

[XXWY10] Cheng-Zhong Xu, Minghua Xu, Le Yi Wang, and George Yin. Filter design and analysis in frequency domain for server scheduling and optimization. IEEE Transactions on Parallel and Distributed Systems, 21(11): 1573–1585, November 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xiang:2015:CCI

[XYT^{+}15] Chaocan Xiang, Panlong Yang, Chang Tian, Haibin Cai, and Yunhao Liu. Calibrate without calibrating: An iterative approach in participatory sensing network. IEEE Transactions on Parallel and Distributed Systems, 26(2):351– 361, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06762992-
                                        abs.html.

Xu:2014:CPT

[XZH14] Ke Xu, Yifeng Zhong, and Huan He. Can P2P technology benefit eyeball ISPs? A cooperative profit distribution answer. IEEE Transactions on Parallel and Distributed Systems, 25(11): 2783–2793, November 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06636889-
                                        abs.html.

Xu:2019:CEC

[XZJ^{+}19] F. Xu, H. Zheng, H. Jiang, W. Shao, H. Liu, and Z. Zhou. Cost-effective cloud server provisioning for predictable performance of big data analytics. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1036–1051, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xun:2017:FDD

[XZQZ17] Yaling Xun, Jifu Zhang, Xiao Qin, and Xujun Zhao.

FiDoop-DP: Data partitioning in frequent itemset mining on Hadoop clusters. IEEE Transactions on Parallel and Distributed Systems, 28(1):101–114, January 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07462277-
abs.html.

Xiao:2012:HMS

[XZSG12] Xin Xiao, Qian Zhang, Yuanchun Shi, and Yuan Gao. How much to share: a repeated game model for peer-to-peer streaming under service differentiation incentives. IEEE Transactions on Parallel and Distributed Systems, 23(2):288– 295, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xu:2013:SDA

[XZT^{+}13] Fengyuan Xu, Xiaojun Zhu, Chiu C. Tan, Qun Li, Guanhua Yan, and Jie Wu. SmartAssoc: Decentralized access point selection algorithm to improve throughput. IEEE Transactions on Parallel and Distributed Systems, 24(12):2482–2491, December 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Xie:2017:EES

[XZX^{+}17] Guoqi Xie, Gang Zeng, Xiongren Xiao, Renfa Li, and Keqin Li. Energyefficient scheduling algorithms for real-time parallel applications on heterogeneous distributed embedded systems. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3426–3442, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/12/07990045-
                                        abs.html.

Yan:2014:DEG

[Yan14] Hong Yan. Design exploration of geometric biclustering for microarray data analysis in data mining. IEEE Transactions on Parallel and Distributed Systems, 25(10): 2540–2550, October 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/10/06579602-
                                        abs.html.

Yu:2018:SFF

[YBY^{+}18] Chao Yu, Yuebin Bai, Hailong Yang, Kun Cheng, Yuhao Gu, Zhongzhi Luan, and Depei Qian. SMGuard: A flexible and fine-grained resource management framework for GPUs. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2849–2862, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08388218-
abs.html.

Yuen:2012:SRT

[YC12] C.-H. Philip Yuen and S.- H. Gary Chan. Scalable real-time monitoring for distributed applications. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2330–2337, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yen:2014:GTA

[YC14] Li-Hsing Yen and Zong-Long Chen. Game-theoretic approach to self-stabilizing distributed formation of minimal multi-dominating sets. IEEE Transactions on Parallel and Distributed Systems, 25(12):3201–3210, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714459-
abs.html.

Yang:2018:LSE

[YC18] Fan Yang and Andrew A.

Chien. Large-scale and extreme-scale computing with stranded green power: Opportunities and costs.

IEEE Transactions on Parallel and Distributed Systems, 29(5):1103–1116, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08187696-
abs.html.

Yao:2017:TRE

[YCMX17] Yuan Yao, Wenzhi Chen, Tulika Mitra, and Yang Xiang. TC-Release++: An efficient timestamp-based coherence protocol for manycore architectures. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3313–3327, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07959101-
abs.html.

Yang:2015:PCI

[YCPC15] Jinn-Shyong Yang, Jou-Ming Chang, Kung-Jui Pai, and Hung-Chang Chan. Parallel construction of independent spanning trees on enhanced hypercubes. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3090–3098, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06948321-
abs.html.

Yang:2013:DLM

[YCTC13] Jie Yang, Yingying (Jennifer) Chen, Wade Trappe, and Jerry Cheng. Detection and localization of multiple spoofing attackers in wireless networks. IEEE Transactions on Parallel and Distributed Systems, 24(1):44– 58, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2012:EPB

[YCW12] Quan Yuan, Ionut Cardei, and Jie Wu. An efficient prediction-based routing in disruption-tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 23(1):19–31, January 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2014:UMC

[YCWL14] Zhongmei Yao, Daren B. H.

Cline, Xiaoming Wang, and Dmitri Loguinov. Unifying models of churn and resilience for unstructured P2P graphs. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2475–2485, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06547627-
abs.html.

You:2017:DIC

[YDC^{+}17] Yang You, James Demmel, Kent Czechowski, Le Song, and Rich Vuduc. Design and implementation of a communication-optimal classifier for distributed kernel support vector machines. IEEE Transactions on Parallel and Distributed Systems, 28(4):974–988, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07565530-
abs.html.

Yao:2017:UIC

[YDH17] Guangshun Yao, Yongsheng Ding, and Kuangrong Hao. Using imbalance characteristic for fault-tolerant workflow scheduling in cloud systems. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3671–3683, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07887706-
abs.html.

Yang:2013:BIB

[YGL13] Xiwang Yang, Yang Guo, and Yong Liu. Bayesianinference-based recommendation in online social networks. IEEE Transactions on Parallel and Distributed Systems, 24(4):642– 651, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2015:APM

[YGL^{+}15] Jianguo Yao, Haibing Guan, Jianying Luo, Lei Rao, and Xue Liu. Adaptive power management through thermal aware workload balancing in Internet data centers. IEEE Transactions on Parallel and Distributed Systems, 26(9):2400–2409, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06887332.
pdf.

Yang:2019:DSA

[YGS^{+}19] T. Yang, S. Gao, Z. Sun, Y. Wang, Y. Shen, and X. Li. Diamond sketch: Accurate per-flow measurement for big streaming data. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2650–2662, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2013:FSR

[YHC^{+}13] Yiling Yang, Yu Huang, Jiannong Cao, Xiaoxing Ma, and Jian Lu. Formal specification and runtime detection of dynamic properties in asynchronous pervasive computing environments. IEEE Transactions on Parallel and Distributed Systems, 24(8):1546–1555, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yan:2018:GDV

[YHL^{+}18] Da Yan, Yuzhen Huang, Miao Liu, Hongzhi Chen, James Cheng, Huanhuan Wu, and Chengcui Zhang. GraphD: Distributed vertexcentric graph processing beyond the memory limit. IEEE Transactions on Parallel and Distributed Systems, 29(1):99–114, January 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/01/08016377-
abs.html.

Yao:2014:PCR

[YHS^{+}14] Yuan Yao, Longbo Huang, Abhishek B. Sharma, Leana Golubchik, and Michael J. Neely. Power cost reduction in distributed data centers: A two-time-scale approach for delay tolerant workloads. IEEE Transactions on Parallel and Distributed Systems, 25(1):200– 211, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). See comment [FYH^{+}15].

Yang:2013:ESD

[YJ13] Kan Yang and Xiaohua Jia. An efficient and secure dynamic auditing protocol for data storage in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1717–1726, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2014:EER

[YJ14] Kan Yang and Xiaohua Jia. Expressive, efficient, and revocable data access control for multi-authority cloud storage. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1735–1744, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2015:ECM

[YJC15] Liang Yu, Tao Jiang, and Yang Cao. Energy cost minimization for distributed Internet data centers in Smart microgrids considering power outages. IEEE Transactions on Parallel and Distributed Systems, 26 (1):120–130, January 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748075-
abs.html.

Yao:2016:ERN

[YJC^{+}16] Jie Yao, Hong Jiang, Qiang Cao, Lei Tian, and Changsheng Xie. Elastic-RAID: A new architecture for improved availability of paritybased RAIDs by elastic Mirroring. IEEE Transactions on Parallel and Distributed Systems, 27(4): 1044–1056, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07106529-
abs.html.

Yu:2015:JWB

[YJCQ15] Liang Yu, Tao Jiang, Yang Cao, and Qi Qi. Joint workload and battery scheduling with heterogeneous service delay guarantees for data center energy cost minimization. IEEE Transactions on Parallel and Distributed Systems, 26(7): 1937–1947, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06827218-
abs.html.

Yang:2015:SVP

[YJR15] Kan Yang, Xiaohua Jia, and Kui Ren. Secure and verifiable policy update outsourcing for big data access control in the cloud. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3461–3470, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06987313-
abs.html.

Yildirim:2014:TSB

[YK14] Kasim Sinan Yildirim and Aylin Kantarci. Time synchronization based on slowflooding in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 25(1):244– 253, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yasudo:2019:DHP

[YKN^{+}19] Ryota Yasudo, Michihiro Koibuchi, Koji Nakano, Hiroki Matsutani, and Hideharu Amano. Designing high-performance interconnection networks with hostswitch graphs. IEEE Transactions on Parallel and Distributed Systems, 30(2):315– 330, February 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/02/08428449-
abs.html.

Yamazaki:2018:SIL

[YKW^{+}18] Ichitaro Yamazaki, Jakub Kurzak, Panruo Wu, Mawussi Zounon, and Jack Dongarra. Symmetric indefinite linear solver using OpenMP task on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1879–1892, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/08/08301559-
                                        abs.html.

Yang:2010:QTC

[YL10] Zheng Yang and Yunhao Liu. Quality of trilateration: Confidence-based iterative localization. IEEE Transactions on Parallel and Distributed Systems, 21(5): 631–640, May 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2011:ALL

[YL11a] Zhongmei Yao and Dmitri Loguinov. Analysis of link lifetimes and neighbor selection in switching DHTs. IEEE Transactions on Parallel and Distributed Systems, 22(11):1834–1841, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2011:UDS

[YL11b] Zhongmei Yao and Dmitri Loguinov. Understanding disconnection and stabilization of Chord. IEEE Transactions on Parallel and Distributed Systems, 22(4):650– 661, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ye:2015:FRA

[YL15] Liang-Cheng Ye and Jia-Rong Liang. Five-round adaptive diagnosis in Hamiltonian networks. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2459–2464, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06881704.
pdf.

Yu:2016:SNP

[YL16] Li Yu and Zhiling Lan.

A scalable, non-parametric method for detecting performance anomaly in large scale computing. IEEE Transactions on Parallel and Distributed Systems, 27(7): 1902–1914, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07236888-
abs.html.

Yang:2016:SGA

[YLC^{+}16] Bo Yang, Zhiyong Li, Shaomiao Chen, Tao Wang, and Keqin Li. Stackelberg game approach for energy-aware resource allocation in data centers. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3646–3658, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07425237-
abs.html.

Yuan:2019:GNG

[YLC^{+}19] Y. Yuan, X. Lian, L. Chen, G. Wang, J. X. Yu, Y. Wang, and Y. M. Ma. GCache: Neighborhood-guided graph caching in a distributed environment. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2463–2477, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2016:DCC

[YLH^{+}16] Bo Yang, Jingwei Li, Qiaoni Han, Tian He, Cailian Chen, and Xinping Guan. Distributed control for charging multiple electric vehicles with overload limitation. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3441–3454, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07422114-
abs.html.

Yang:2017:FFT

[YLJ^{+}17] Lei Yang, Weichen Liu, Weiwen Jiang, Mengquan Li, Peng Chen, and Edwin Hsing-Mean Sha. FoToNoC: A folded torus-like networkon-chip based many-core systems-on-chip in the dark silicon era. IEEE Transactions on Parallel and Distributed Systems, 28(7): 1905–1918, July 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/07/07795229-
abs.html.

Yang:2017:CCI

[YLL^{+}17] Chen Yang, Leibo Liu, Kai Luo, Shouyi Yin, and Shaojun Wei. CIACP: A correlation- and iterationaware cache partitioning mechanism to improve performance of multiple coarsegrained reconfigurable arrays. IEEE Transactions on Parallel and Distributed Systems, 28(1):29–43, January 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/01/07452657-
abs.html.

Yin:2016:EPI

[YLLW16] Shouyi Yin, Xinhan Lin, Leibo Liu, and Shaojun Wei. Exploiting parallelism of imperfect nested loops on coarse-grained reconfigurable architectures. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3199–3213, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07412757-
abs.html.

Yuan:2015:GNC

[YLM^{+}15] Jun Yuan, Aixia Liu, Xue Ma, Xiuli Liu, Xiao Qin, and Jifu Zhang. The ggood-neighbor conditional diagnosability of k-ary ncubes under the PMC Modeland MM* model. IEEE Transactions on Parallel and Distributed Systems, 26(4): 1165–1177, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06800081-
abs.html.

Yuan:2012:QAL

[YLR12] Yanling Yuan, Zuyi Li, and Kui Ren. Quantitative analysis of load redistribution attacks in power systems. IEEE Transactions on Parallel and Distributed Systems, 23(9):1731–1738, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yan:2013:NBR

[YLSQ13] Zhiwei Yan, Jong-Hyouk Lee, Sean Shen, and Chunming Qiao. Novel branchingrouter-based multicast routing protocol with mobility support. IEEE Transactions on Parallel and Distributed Systems, 24(10): 2060–2068, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2015:CED

[YLT15] Kun Yuan, Qing Ling, and Zhi Tian. Communicationefficient decentralized event monitoring in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2198–2207, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06881722-
abs.html.

Yi:2013:ETS

[YLW13] Xun Yi, San Ling, and Huaxiong Wang. Efficient twoserver password-only authenticated key exchange. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1773–1782, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2014:RCB

[YLW^{+}14] Zheng Yang, Kai Lu, Xiaoping Wang, Jun Luo, and Yunhao Liu. Robust component-based localization in sparse networks. IEEE Transactions on Parallel and Distributed Systems, 25(5):1317–1327, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2015:TEA

[YLZ^{+}15a] Chi Yang, Chang Liu, Xuyun Zhang, Surya Nepal, and Jinjun Chen. A time efficient approach for detecting errors in big sensor data on cloud. IEEE Transactions on Parallel and Distributed Systems, 26(2): 329–339, February 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06714550-
abs.html.

Yin:2015:BBS

[YLZ^{+}15b] Jianwei Yin, Xingjian Lu, Xinkui Zhao, Hanwei Chen, and Xue Liu. BURSE: A bursty and self-similar workload generator for cloud computing. IEEE Transactions on Parallel and Distributed Systems, 26(3):668– 680, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06782285-
abs.html.

Yavits:2015:SMM

[YMG15] L. Yavits, A. Morad, and R. Ginosar. Sparse matrix multiplication on an associative processor. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3175–3183, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06954570-
abs.html.

Yan:2016:ETE

[YMHL16] Guihai Yan, Jun Ma, Yinhe Han, and Xiaowei Li. EcoUp: Towards economical datacenter upgrading. IEEE Transactions on Parallel and Distributed Systems, 27(7): 1968–1981, July 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07254219-
abs.html.

Yoosefi:2017:CAC

[YN17] Amin Yoosefi and Hamid Reza Naji. A clustering algorithm for communicationaware scheduling of task graphs on multi-core reconfigurable systems. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2718–2732, October 2017.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/10/07924424-
                                        abs.html.

Yabuta:2017:RJG

[YNK^{+}17] Makoto Yabuta, Anh Nguyen, Shinpei Kato, Masato Edahiro, and Hideyuki Kawashima. Relational joins on GPUs: A closer look. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2663–2673, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/09/07869389-
                                        abs.html.

Yingchareonthawornchai:2018:ABH

[YNKD18] Sorrachai Yingchareonthawornchai, Duong N. Nguyen, Sandeep S. Kulkarni, and Murat Demirbas. Analysis of bounds on hybrid vector clocks. IEEE Transactions on Parallel and Distributed Systems, 29(9): 1947–1960, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08323242-
                                        abs.html.

Yang:2013:DDQ

[YNW13] Zhipeng Yang, Ting Ning, and Hongyi Wu. Distributed data query in intermittently connected passive RFID networks. IEEE Transactions on Parallel and Distributed Systems, 24(10): 1972–1982, October 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yoon:2017:DRA

[YOK^{+}17] Myung Kuk Yoon, Yunho Oh, Seung Hun Kim, Sangpil Lee, Deokho Kim, and Won Woo Ro. Dynamic resizing on active warps scheduler to hide operation stalls on GPUs. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3142–3156, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07927466-
abs.html.

Yan:2014:TPS

[YOWA14] Gongjun Yan, Stephan Olariu, Jin Wang, and Samiur Arif. Towards providing scalable and robust privacy in vehicular networks. IEEE Transactions on Parallel and Distributed Systems, 25(7):1896–1906, July 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2013:RSS

[YP13] Yi-Hua E. Yang and Viktor K. Prasanna. Robust and scalable string pattern matching for deep packet inspection on multicore processors. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2283–2292, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yeh:2013:IDM

[YPL13] Chi-Yuan Yeh, Yu-Ting Peng, and Shie-Jue Lee. An iterative divide-and-merge-based approach for solving large-scale least squares problems. IEEE Transactions on Parallel and Distributed Systems, 24(3):428– 438, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2017:EES

[YPL^{+}17] Ting Yang, Haibo Pen, Wei Li, Dong Yuan, and Albert Y. Zomaya. An energyefficient storage strategy for cloud datacenters based on variable K-coverage of a hypergraph. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3344–3355, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/07968334-
abs.html.

Yoon:2011:CSS

[YQ11] Seokhoon Yoon and Chunming Qiao. Cooperative search and survey using autonomous underwater vehicles (AUVs). IEEE Transactions on Parallel and Distributed Systems, 22(3):364– 379, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2016:SSS

[YQ16] Ye Yu and Chen Qian. Space shuffle: A scalable, flexible, and high-performance data center network. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3351–3365, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07416229-
abs.html.

Yang:2015:SIJ

[YQH^{+}15] Lei Yang, Yong Qi, Jinsong Han, Cheng Wang, and Yunhao Liu. Shelving interference and joint identification in large-scale RFID systems. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3149–3159, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

trans/td/2015/11/06654142-
abs.html.

Ye:2016:RTI

[YQH16] Feng Ye, Yi Qian, and Rose Qingyang Hu. A real-time information based demand-side management system in smart grid. IEEE Transactions on Parallel and Distributed Systems, 27(2): 329–339, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/02/07042318-
abs.html.

Yu:2014:EDC

[YQLS14] Lei Yu, Chenxi Qiu, Ze Li, and Haiying Shen. Efficient data collection for large-scale mobile monitoring applications. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1424–1436, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yagan:2012:OAI

[YQZC12] Osman Yagan, Dajun Qian, Junshan Zhang, and D. Cochran. Optimal allocation of interconnecting links in cyberphysical systems: Interdependence, cascading failures, and robustness. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1708–1720, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yzelman:2014:HLS

[YR14] Albert-Jan Nicholas Yzelman and Dirk Roose. Highlevel strategies for parallel shared-memory sparse matrix-vector multiplication. IEEE Transactions on Parallel and Distributed Systems, 25(1):116–125, January 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2011:FTF

[YRL11] Shucheng Yu, Kui Ren, and Wenjing Lou. FDAC: Toward fine-grained distributed data access control in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 22(4): 673–686, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2016:BDB

[YRLY16] Xinyu Yang, Xuebin Ren, Jie Lin, and Wei Yu. On binary decomposition based privacy-preserving aggregation schemes in real-time monitoring systems. IEEE Transactions on Parallel and Distributed Systems, 27(10): 2967–2983, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/10/07379007-
abs.html.

Yoon:2011:CLM

[YSDQ11] Seokhoon Yoon, Onur Soysal, Murat Demirbas, and Chunming Qiao. Coordinated locomotion and monitoring using autonomous mobile sensor nodes. IEEE Transactions on Parallel and Distributed Systems, 22(10): 1742–1756, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2014:SME

[YSG^{+}14] Zheng Yang, Longfei Shangguan, Weixi Gu, Zimu Zhou, Chenshu Wu, and Yunhao Liu. Sherlock: Micro-environment sensing for Smartphones. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3295–3305, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714451-
abs.html.

Yu:2017:CCE

[YSS^{+}17] Lei Yu, Haiying Shen, Karan Sapra, Lin Ye, and Zhipeng Cai. CoRE: Cooperative endto-end traffic redundancy elimination for reducing cloud bandwidth cost. IEEE Transactions on Parallel and Distributed Systems, 28(2): 446–461, February 2017.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07488173-
abs.html.

Yang:2010:PCA

[YTL^{+}10] Guangdi Yang, Dingyuan Tu, Rufeng Lin, Lu Rong, and Yang Du. Performance of channel-aware scheduling algorithms for HDR-WPAN. IEEE Transactions on Parallel and Distributed Systems, 21(2):257– 274, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yin:2019:PPN

[YTL^{+}19] Shouyi Yin, Shibin Tang, Xinhan Lin, Peng Ouyang, Fengbin Tu, Leibo Liu, Jishen Zhao, Cong Xu, Shuangcheng Li, Yuan Xie, and Shaojun Wei. Parana: A parallel neural architecture considering thermal problem of 3D stacked memory. IEEE Transactions on Parallel and Distributed Systems, 30(1): 146–160, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08416708-
abs.html.

Yan:2016:GPG

[YTMS16] Jie Yan, Guangming Tan, Zeyao Mo, and Ninghui Sun. Graphine: Programming graph-parallel computation of large natural graphs for multicore clusters. IEEE Transactions on Parallel and Distributed Systems, 27(6):1647–1659, June 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/06/07152922-
                                        abs.html.

Yao:2019:SEL

[YTW^{+}19] T. Yao, Z. Tan, J. Wan, P. Huang, Y. Zhang, C. Xie, and X. He. SEALDB: An efficient LSM-tree based KV store on SMR drives with sets and dynamic bands. IEEE Transactions on Parallel and Distributed Systems, 30(11):2595–2607, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2011:PPF

[YTZ^{+}11] Zhi Yang, Jing Tian, Ben Y. Zhao, Wei Chen, and Yafei Dai. Protector: a probabilistic failure detector for costeffective peer-to-peer storage. IEEE Transactions on Parallel and Distributed Systems, 22(9):1514–1527, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2010:EBU

[YW10] Shuhui Yang and Jie Wu.

Efficient broadcasting using network coding and directional antennas in MANETs. IEEE Transactions on Parallel and Distributed Systems, 21(2):148–161, February 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yue:2011:CDC

[YWC11] Xiaonan Yue, Chi-Fai Michael Wong, and Shueng-Han Gary Chan. CACAO: Distributed client-assisted channel assignment optimization for uncoordinated WLANs. IEEE Transactions on Parallel and Distributed Systems, 22(9):1433–1440, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Wu:2011:QAD

[yWeH11] Shiow yang Wu and Cheng en He. QoS-aware dynamic adaptation for cooperative media streaming in mobile environments. IEEE Transactions on Parallel and Distributed Systems, 22(3):439– 450, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2011:PAS

[YWJJ11] Yuan Yuan, Kui Wu, Weijia Jia, and Yuming Jiang. Performance of acyclic stochastic networks with network coding. IEEE Transactions on Parallel and Distributed Systems, 22(7):1238–1245, July 2011. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ye:2015:PBW

[YWW^{+}15] Kejiang Ye, Zhaohui Wu, Chen Wang, Bing Bing Zhou, Weisheng Si, Xiaohong Jiang, and Albert Y. Zomaya. Profiling-based workload consolidation and migration in virtualized data centers. IEEE Transactions on Parallel and Distributed Systems, 26(3):878– 890, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06777568-
abs.html.

Yuan:2018:SIC

[YWWR18] Xingliang Yuan, Jian Weng, Cong Wang, and Kui Ren. Secure integrated circuit design via hybrid cloud. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1851–1864, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08295269-
abs.html.

Yang:2017:RVM

[YWY^{+}17] Song Yang, Philipp Wieder, Ramin Yahyapour, Stojan Trajanovski, and Xiaoming Fu. Reliable virtual machine placement and routing in clouds. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2965–2978, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07896612-
abs.html.

Yun:2017:TSW

[YWZ17] Daqing Yun, Chase Q.

Wu, and Michelle M. Zhu. Transport-support workflow composition and optimization for big data movement in high-performance networks. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3656–3670, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/12/08000332-
abs.html.

Yu:2012:AFD

[YXG12] Bo Yu, Cheng-Zhong Xu, and Minyi Guo. Adaptive forwarding delay control for VANET data aggregation. IEEE Transactions on Parallel and Distributed Systems, 23(1):11–18, January 2012.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2016:PPC

[YXLJ16] Pingpeng Yuan, Changfeng Xie, Ling Liu, and Hai Jin. PathGraph: A path centric graph processing system. IEEE Transactions on Parallel and Distributed Systems, 27(10):2998–3012, October 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/10/07384525-
                                        abs.html.

Yang:2013:NPA

[YXSS13] Jiaoyun Yang, Yun Xu, Guangzhong Sun, and Yi Shang. A new progressive algorithm for a multiple longest common subsequences problem and its efficient parallelization. IEEE Transactions on Parallel and Distributed Systems, 24(5):862– 870, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2016:HHR

[YXWL16] Yuan Yang, Mingwei Xu, Dan Wang, and Suogang Li. A hop-by-hop routing mechanism for green Internet. IEEE Transactions on Parallel and Distributed Systems, 27(1):2–16, January 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07017539-
abs.html.

Yang:2014:RPR

[YXWW14] Qing Yang, Changsheng Xie, Jibin Wang, and Jiguang Wan. S^{2}-RAID: Parallel RAID architecture for fast data recovery. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1638–1647, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yang:2010:HAL

[YY10] Min Yang and Yuanyuan Yang. A hypergraph approach to linear network coding in multicast networks. IEEE Transactions on Parallel and Distributed Systems, 21(7):968–982, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2014:PPB

[YY14] Jiawei Yuan and Shucheng Yu. Privacy preserving backpropagation neural network learning made practical with cloud computing. IEEE Transactions on Parallel and Distributed Systems, 25(1): 212–221, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yildirim:2011:POP

[YYK11a] Esma Yildirim, Dengpan Yin, and Tevfik Kosar. Prediction of optimal parallelism level in wide area data transfers. IEEE Transactions on Parallel and Distributed Systems, 22(12): 2033–2045, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yin:2011:DTP

[YYK^{+}11b] Dengpan Yin, Esma Yildirim, Sivakumar Kulasekaran, Brandon Ross, and Tevfik Kosar. A data throughput prediction and optimization service for widely distributed many-task computing. IEEE Transactions on Parallel and Distributed Systems, 22(6): 899–909, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yuan:2013:HPA

[YYL^{+}13] Dong Yuan, Yun Yang, Xiao Liu, Wenhao Li, Lizhen Cui, Meng Xu, and Jinjun Chen. A highly practical approach toward achieving minimum data sets storage cost in the cloud. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1234–1244, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yin:2017:CFL

[YYL^{+}17] Shouyi Yin, Xianqing Yao, Tianyi Lu, Dajiang Liu, Jiangyuan Gu, Leibo Liu, and Shaojun Wei. Conflictfree loop mapping for coarsegrained reconfigurable architecture with multi-bank memory. IEEE Transactions on Parallel and Distributed Systems, 28(9):2471–2485, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07879331-
abs.html.

Yang:2014:FDI

[YYY^{+}14] Qingyu Yang, Jie Yang, Wei Yu, Dou An, Nan Zhang, and Wei Zhao. On false datainjection attacks against power system state estimation: Modeling and countermeasures. IEEE Transactions on Parallel and Distributed Systems, 25(3):717– 729, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2011:TDA

[YZDJ11] Shui Yu, Wanlei Zhou, Robin Doss, and Weijia Jia. Traceback of DDoS attacks using entropy variations. IEEE Transactions on Parallel and Distributed Systems, 22(3): 412–425, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2010:SDW

[YZFZ10] Wei Yu, Nan Zhang, Xinwen Fu, and Wei Zhao. Selfdisciplinary worms and countermeasures: Modeling and analysis. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1501–1514, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

You:2019:FDN

[YZH^{+}19] Y. You, Z. Zhang, C. Hsieh, J. Demmel, and K. Keutzer. Fast deep neural network training on distributed systems and cloud TPUs. IEEE Transactions on Parallel and Distributed Systems, 30(11): 2449–2462, November 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2017:PPD

[YZHZ17] Hongliang Yu, Xu Zhang, Wei Huang, and Weimin Zheng. PDFS: Partially dedupped file system for primary workloads. IEEE Transactions on Parallel and Distributed Systems, 28(3): 863–876, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/

trans/td/2017/03/07522088-
abs.html.

Yu:2012:DDA

[YZJ^{+}12] Shui Yu, Wanlei Zhou, Weijia Jia, Song Guo, Yong Xiang, and Feilong Tang. Discriminating DDoS attacks from flash crowds using flow correlation coefficient. IEEE Transactions on Parallel and Distributed Systems, 23(6): 1073–1080, June 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yao:2015:CCO

[YZL^{+}15] Jianguo Yao, Haihang Zhou, Jianying Luo, Xue Liu, and Haibing Guan. COMIC: Cost optimization for Internet content multihoming. IEEE Transactions on Parallel and Distributed Systems, 26(7):1851–1860, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06832634-
abs.html.

Yin:2017:EPF

[YZL^{+}17] Hao Yin, Xu Zhang, Hongqiang H.

Liu, Yan Luo, Chen Tian, Shuoyao Zhao, and Feng Li. Edge provisioning with flexible server placement. IEEE Transactions on Parallel and Distributed Systems, 28(4):1031–1045, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/04/07556983-
                                        abs.html.

Ye:2013:SAB

[YZS13] Dayong Ye, Minjie Zhang, and Danny Sutanto. Selfadaptation-based dynamic coalition formation in a distributed agent network: A mechanism and a brief survey. IEEE Transactions on Parallel and Distributed Systems, 24(5):1042–1051, May 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ye:2014:MAE

[YZSC14] Mingjiang Ye, Yifeng Zhong, Meng Shen, and Yong Cui. A model approach to the estimation of peer-to-peer traffic matrices. IEEE Transactions on Parallel and Distributed Systems, 25(5): 1101–1111, May 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yin:2017:CNI

[YZT^{+}17] Jianwei Yin, Xinkui Zhao, Yan Tang, Chen Zhi, Zuoning Chen, and Zhaohui Wu. CloudScout: A nonintrusive approach to service dependency discovery. IEEE Transactions on Parallel and Distributed Systems, 28(5):1271–1284, May 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07604145-
abs.html.

Zola:2010:PIT

[ZASA10] Jaroslaw Zola, Maneesha Aluru, Abhinav Sarje, and Srinivas Aluru. Parallel information-theory-based construction of genome-wide gene regulatory networks. IEEE Transactions on Parallel and Distributed Systems, 21(12):1721–1733, December 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zonouz:2015:SIF

[ZBK^{+}15] Saman A. Zonouz, Robin Berthier, Himanshu Khurana, William H. Sanders, and Tim Yardley. Seclius: An information flow-based, consequence-centric security metric. IEEE Transactions on Parallel and Distributed Systems, 26(2):562– 573, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06547151-
abs.html.

Zhang:2015:DMB

[ZBS15] Junchao Zhang, Babak Behzad, and Marc Snir. Design of a multithreaded Barnes–Hut algorithm for multicore clusters. IEEE Transactions on Parallel and Distributed Systems, 26 (7):1861–1873, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06837521-
abs.html.

Zhang:2017:SAN

[ZCC^{+}17] Mingzhe Zhang, Haibo Chen, Luwei Cheng, Francis C. M. Lau, and Cho-Li Wang. Scalable adaptive NUMAaware lock. IEEE Transactions on Parallel and Distributed Systems, 28(6): 1754–1769, June 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07748539-
abs.html.

Zhang:2017:CAV

[ZCG^{+}17] Weizhan Zhang, Yuxuan Chen, Xiang Gao, Zhichao Mo, Qinghua Zheng, and Zongqing Lu. Clusteraware virtual machine collaborative migration in media cloud. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2808–2822, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07909001-
abs.html.

Zhang:2019:PGE

[ZCJ19] Fan Zhang, Hanhua Chen, and Hai Jin. Piggyback Game: Efficient event stream dissemination in online social network systems. IEEE Transactions on Parallel and Distributed Systems, 30(3): 692–709, March 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/03/08440751-
abs.html.

Zhang:2014:RCO

[ZCJY14] Qian Zhang, Yang Cao, Tao Jiang, and Liang Yu. Riskconstrained operation for Internet data centers in deregulated electricity markets. IEEE Transactions on Parallel and Distributed Systems, 25(5):1306–1316, May 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2014:CLA

[ZCLS14] Qiang Zheng, Guohong Cao, Thomas F. La Porta, and Ananthram Swami. Crosslayer approach for minimizing routing disruption in IP networks. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1659–1669, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2010:OSM

[ZCX10] Xiaoying Zheng, Chunglae Cho, and Ye Xia. Optimal swarming for massive content distribution. IEEE Transactions on Parallel and Distributed Systems, 21(6): 841–856, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2014:FTR

[ZCX^{+}14] Weiping Zhu, Jiannong Cao, Yi Xu, Lei Yang, and Junjun Kong. Fault-tolerant RFID reader localization based on passive RFID tags. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2065–2076, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2015:ASA

[ZCX15] Xiaoying Zheng, Chunglae Cho, and Ye Xia. Algorithms and stability analysis for content distribution over multiple multicast trees. IEEE Transactions on Parallel and Distributed Systems, 26(5):1217–1227, May 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2015/05/06804015-
                                        abs.html.

Zhao:2016:PCC

[ZCXF16] Jiacheng Zhao, Huimin Cui, Jingling Xue, and Xiaobing Feng. Predicting crosscore performance interference on multicore processors with regression analysis. IEEE Transactions on Parallel and Distributed Systems, 27(5):1443–1456, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07120141-
abs.html.

Zhang:2012:TCW

[ZCZ^{+}12] Honglei Zhang, Hua Chai, Wenbing Zhao, P. Michael Melliar-Smith, and Louise E. Moser. Trustworthy coordination of Web services atomic transactions. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1551–1565, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2012:DAP

[ZD12] Peng Zhang and Yuefan Deng. Design and analysis of pipelined broadcast algorithms for the all-port interlaced bypass torus networks. IEEE Transactions on Parallel and Distributed Systems, 23(12):2245–2253, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2016:LFP

[ZD16a] Deli Zhang and Damian Dechev. A lock-free priority queue design based on multidimensional linked lists. IEEE Transactions on Parallel and Distributed Systems, 27(3):613–626, March 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/03/07079379-
                                        abs.html.

Ziwich:2016:NOC

[ZD16b] Roverli P. Ziwich and Elias P. Duarte. A nearly optimal comparison-based diagnosis algorithm for systems of arbitrary topology. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3131–3143, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/11/07399770-
                                        abs.html.

Zhang:2015:EVN

[ZDF^{+}15] Peng Zhang, Yuefan Deng, Rui Feng, Xingguo Luo, and Jiangxing Wu. Evaluation of various networks configurated by adding bypass or torus links. IEEE Transactions on Parallel and Distributed Systems, 26(4):

984–996, April 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/04/06782442-
abs.html.

Zhu:2014:PMD

[ZDG^{+}14] Haojin Zhu, Suguo Du, Zhaoyu Gao, Mianxiong Dong, and Zhenfu Cao. A probabilistic misbehavior detection scheme toward efficient trust establishment in delay-tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 25(1):22–32, January 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:EDH

[ZDM^{+}17] Zhen Zhang, Yuhui Deng, Geyong Min, Junjie Xie, and Shuqiang Huang. ExCCC-DCN: A highly scalable, cost-effective and energyefficient data center structure. IEEE Transactions on Parallel and Distributed Systems, 28(4):1046–1060, April 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07567529-
abs.html.

Zhang:2019:HHS

[ZDM^{+}19] Z. Zhang, Y. Deng, G. Min, J. Xie, L. T. Yang, and Y. Zhou. HSDC: A highly scalable data center network architecture for greater incremental scalability. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1105–1119, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2011:UBE

[ZDWR11] Chao Zhang, Prithula Dhungel, Di Wu, and Keith W. Ross. Unraveling the Bit-Torrent ecosystem. IEEE Transactions on Parallel and Distributed Systems, 22(7): 1164–1177, July 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zoni:2016:CBM

[ZFF16] Davide Zoni, Jose Flich, and William Fornaciari. CUT-BUF: Buffer management and router design for traffic mixing in VNET-based NoCs. IEEE Transactions on Parallel and Distributed Systems, 27(6):1603–1616, June 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

                                        csdl.computer.org/csdl/
                                        trans/td/2016/06/07202891-
                                        abs.html.

Zhu:2010:CBT

[ZFG^{+}10] Ye Zhu, Xinwen Fu, Bryan Gramham, Riccardo Bettati, and Wei Zhao. Correlationbased traffic analysis attacks on anonymity networks. IEEE Transactions on Parallel and Distributed Systems, 21(7):954–967, July 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:CSD

[ZFG^{+}14] Qingquan Zhang, Lingkun Fu, Yu Jason Gu, Lin Gu, Qing Cao, Jiming Chen, and Tian He. Collaborative scheduling in dynamic environments using error inference. IEEE Transactions on Parallel and Distributed Systems, 25(3):591– 601, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:ARW

[ZFT^{+}15] Jinbei Zhang, Luoyi Fu, Xiaohua Tian, Ying Cui, and Xinbing Wang. Analysis of random walk mobility models with location heterogeneity. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2657–2670, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06915731.
pdf.

Zaman:2011:DAR

[ZG11] Sharrukh Zaman and Daniel Grosu. A distributed algorithm for the replica placement problem. IEEE Transactions on Parallel and Distributed Systems, 22(9): 1455–1468, September 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2013:PDF

[ZGGW13] Yanfeng Zhang, Qixin Gao, Lixin Gao, and Cuirong Wang. PrIter: A distributed framework for prioritizing iterative computations. IEEE Transactions on Parallel and Distributed Systems, 24(9): 1884–1893, September 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:MAG

[ZGGW14] Yanfeng Zhang, Qixin Gao, Lixin Gao, and Cuirong Wang. Maiter: An asynchronous graph processing framework for delta-based accumulative iterative computation. IEEE Transactions on Parallel and Distributed Systems, 25(8):2091–2100, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2014:RBD

[ZGH14] Deze Zeng, Song Guo, and Jiankun Hu. Reliable bulkdata dissemination in delay tolerant networks. IEEE Transactions on Parallel and Distributed Systems, 25(8): 2180–2189, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zotkiewicz:2016:MDE

[ZGKB16] Mateusz Zotkiewicz, Mateusz Guzek, Dzmitry Kliazovich, and Pascal Bouvry. Minimum dependencies energy-efficient scheduling in data centers. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3561–3574, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07434655-
abs.html.

Zhang:2010:CMD

[ZGL10] Yuanfang Zhang, Christopher D. Gill, and Chenyang Lu. Configurable middleware for distributed real-time systems with aperiodic and periodic tasks. IEEE Transactions on Parallel and Distributed Systems, 21(3):393– 404, March 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:DCB

[ZGL^{+}15] Wenlin Zhang, Yi Guo, Hongbo Liu, Yingying Jennifer Chen, Zheng Wang, and Joseph Mitola. Distributed consensus-based weight design for cooperative spectrum sensing. IEEE Transactions on Parallel and Distributed Systems, 26(1):54– 64, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06747402-
abs.html.

Zeng:2014:TTW

[ZGXJ14] Deze Zeng, Song Guo, Yong Xiang, and Hai Jin. On the throughput of two-way relay networks using network coding. IEEE Transactions on Parallel and Distributed Systems, 25(1):191– 199, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:BRP

[ZGY15] Zhemin Zhang, Zhiyang Guo, and Yuanyuan Yang. Bounded-reorder packet scheduling in optical cut-through switch. IEEE Transactions on Parallel and Distributed Systems, 26(11): 2927–2941, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06928510-
abs.html.

Zhong:2011:RMA

[ZH11] Ziguo Zhong and Tian He.

RSD: a metric for achieving range-free localization beyond connectivity. IEEE Transactions on Parallel and Distributed Systems, 22(11): 1943–1951, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2014:KHT

[ZH14a] Jianlong Zhong and Bingsheng He. Kernelet: Highthroughput GPU kernel executions with dynamic slicing and scheduling. IEEE Transactions on Parallel and Distributed Systems, 25(6): 1522–1532, June 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2014:MSG

[ZH14b] Jianlong Zhong and Bingsheng He. Medusa: Simplified graph processing on GPUs. IEEE Transactions on Parallel and Distributed Systems, 25(6):1543–1552, June 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zuo:2018:WFC

[ZH18] Pengfei Zuo and Yu Hua.

A write-friendly and cacheoptimized hashing scheme for non-volatile memory systems. IEEE Transactions on Parallel and Distributed Systems, 29(5):985– 998, May 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/05/08186236-
abs.html.

Zhang:2012:NPS

[Zha12] Nan Zhang. A novel parallel scan for multicore processors and its application in sparse matrix-vector multiplication. IEEE Transactions on Parallel and Distributed Systems, 23(3):397–404, March 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2012:CPD

[ZHAY12] Yan Zhu, Hongxin Hu, Gail-Joon Ahn, and Mengyang Yu. Cooperative provable data possession for integrity verification in multicloud storage. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2231–2244, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2017:DOE

[ZHCL17] Amelie Chi Zhou, Bingsheng He, Xuntao Cheng, and Chiew Tong Lau. A declarative optimization engine for resource provisioning of scientific workflows in geo-distributed clouds. IEEE Transactions on Parallel and Distributed Systems, 28(3):647–661, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07542141-
abs.html.

Zhu:2012:EET

[ZHCW12] Ying Zhu, Minsu Huang, Siyuan Chen, and Yu Wang. Energy-efficient topology control in cooperative ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1480–1491, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:OCM

[ZHL^{+}15] Desheng Zhang, Tian He, Shan Lin, Sirajum Munir, and John A. Stankovic. Online cruising mile reduction in large-scale taxicab networks. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3122–3135, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06930792-
abs.html.

Zhang:2019:SSR

[ZHL19] M. Zhang, S. Han, and P. P. C. Lee. SimEDC:

A simulator for the reliability analysis of erasurecoded data centers. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2836–2848, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2012:SDC

[ZHQ12] Jiazhen Zhou, Rose Qingyang Hu, and Yi Qian. Scalable distributed communication architectures to support advanced metering infrastructure in Smart Grid. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1632–1642, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zuo:2019:BEE

[ZHS^{+}19] Pengfei Zuo, Yu Hua, Yuanyuan Sun, Xue Liu, Jie Wu, Yuncheng Guo, Wen Xia, Shunde Cao, and Dan Feng. Bandwidth and energy efficient image sharing for situation awareness in disasters. IEEE Transactions on Parallel and Distributed Systems, 30(1):15– 28, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08419321-
abs.html.

Zhuang:2014:OTS

[Zhu14]  Weihua Zhuang.  Opti-

mal transmission scheduling of cooperative communications with a fullduplex relay. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2353–2363, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06579601-
abs.html.

Zhou:2019:CAP

[ZHW^{+}19] Jian Zhou, Dezhi Han, Jun Wang, Xiaobo Zhou, and Changjun Jiang. A correlation-aware page-level FTL to exploit semantic links in workloads. IEEE Transactions on Parallel and Distributed Systems, 30(4):723–737, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO.

URL https://ieeexplore.
ieee.org/document/8470972/

.

Zhou:2019:RBC

[ZHX^{+}19] S. Zhou, Y. He, S. Xiang, K. Li, and Y. Liu. Regionbased compressive networked storage with lazy encoding. IEEE Transactions on Parallel and Distributed Systems, 30(6):1390–1402, June 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2015:EAW

[ZHZC15] Bo Zhao, Wenjie Hu, Qiang Zheng, and Guohong Cao. Energy-aware Web browsing on Smartphones. IEEE Transactions on Parallel and Distributed Systems, 26 (3):761–774, March 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/03/06776557-
abs.html.

Zhu:2017:OQP

[ZHZL17] Jieming Zhu, Pinjia He, Zibin Zheng, and Michael R. Lyu. Online QoS prediction for runtime service adaptation via adaptive matrix factorization. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2911–2924, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07918529-
abs.html.

Zhu:2016:SAC

[ZJ16] Zhongma Zhu and Rui Jiang. A secure anti-collusion data sharing scheme for dynamic groups in the cloud. IEEE Transactions on Parallel and Distributed Systems, 27 (1):40–50, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07004069-

abs.html. See comments [WZS^{+}18].

Zhang:2016:EPS

[ZJKQ16] Ji Zhang, Xunfei Jiang, Wei-Shinn Ku, and Xiao Qin. Efficient parallel skyline evaluation using MapReduce. IEEE Transactions on Parallel and Distributed Systems, 27(7):1996–2009, July 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/07/07219441-
abs.html.

Zeng:2012:DFI

[ZJL^{+}12] Rongfei Zeng, Yixin Jiang, Chuang Lin, Yanfei Fan, and Xuemin (Sherman) Shen. A distributed fault/intrusiontolerant sensor data storage scheme based on network coding and homomorphic fingerprinting. IEEE Transactions on Parallel and Distributed Systems, 23(10): 1819–1830, October 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:PEL

[ZJL^{+}17a] Lan Zhang, Taeho Jung, Kebin Liu, Xiang-Yang Li, Xuan Ding, Jiaxi Gu, and Yunhao Liu. PIC: Enable large-scale privacy preserving content-based image search on cloud. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3258–3271, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07938682-
abs.html.

Zhang:2017:PPI

[ZJL^{+}17b] Weihua Zhang, Xiaofeng Ji, Yunping Lu, Haojun Wang, Haibo Chen, and Pen-Chung Yew. Prophet: A parallel instruction-oriented manycore simulator. IEEE Transactions on Parallel and Distributed Systems, 28(10): 2939–2952, October 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/10/07917344-
abs.html.

Zhou:2014:AaS

[ZJLG14] Xiaobo Zhou, Changjun Jiang, Palden Lama, and Yanfei Guo. Automated and agile server parameter tuning by coordinated learning and control. IEEE Transactions on Parallel and Distributed Systems, 25(4): 876–886, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2012:DAC

[ZJLS12] Rongfei Zeng, Yixin Jiang, Chuang Lin, and Xuemin (Sherman) Shen. Dependability analysis of control center networks in Smart Grid using stochastic Petri nets. IEEE Transactions on Parallel and Distributed Systems, 23(9): 1721–1730, September 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2012:SCC

[ZJS12] Eddy Zheng Zhang, Yunlian Jiang, and Xipeng Shen. The significance of CMP cache sharing on contemporary multithreaded applications. IEEE Transactions on Parallel and Distributed Systems, 23(2):367– 374, February 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:VFT

[ZJS^{+}17] Weihua Zhang, Xiaofeng Ji, Bo Song, Shiqiang Yu, Haibo Chen, Tao Li, Pen-Chung Yew, and Wenyun Zhao. VarCatcher: A framework for tackling performance variability of parallel workloads on multi-core. IEEE Transactions on Parallel and Distributed Systems, 28(4):1215–1228, April 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/04/07576653-
                                        abs.html.

Zomaya:2014:APT

[ZJTZ14] Albert Y. Zomaya, Malith Jayasinghe, Zahir Tari, and Panlop Zeephongsekul. ADAPT-POLICY: Task assignment in server farms when the service time distribution of tasks is not known a priori. IEEE Transactions on Parallel and Distributed Systems, 25(4): 851–861, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2016:DGD

[ZJZ^{+}16] Yuhong Zhao, Hong Jiang, Ke Zhou, Zhijie Huang, and Ping Huang. DREAM-(L)G: A distributed groupingbased algorithm for resource assignment for bandwidthintensive applications in the cloud. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3469–3484, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/12/07423784-
                                        abs.html.

Zhou:2019:HHT

[ZKP^{+}19] S. Zhou, R. Kannan, V. K. Prasanna, G. Seetharaman, and Q. Wu. HitGraph: Highthroughput graph processing framework on FPGA. IEEE Transactions on Parallel and Distributed Systems, 30(10): 2249–2264, October 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zonouz:2014:RGT

[ZKSY14] Saman A. Zonouz, Himanshu Khurana, William H. Sanders, and Timothy M. Yardley. RRE: A gametheoretic intrusion response and recovery engine. IEEE Transactions on Parallel and Distributed Systems, 25(2): 395–406, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zier:2010:PED

[ZL10] David A. Zier and Ben Lee.

Performance evaluation of dynamic speculative multithreading with the Cascadia architecture. IEEE Transactions on Parallel and Distributed Systems, 21(1):47– 59, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2011:MPN

[ZL11] Xinyu Zhang and Baochun Li. On the market power of network coding in P2P content distribution systems. IEEE Transactions on Parallel and Distributed Systems, 22(12):2063–2070, December 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2014:IDL

[ZL14] Yue Zhao and Francis C. M. Lau. Implementation of decoders for LDPC block codes and LDPC convolutional codes based on GPUs. IEEE Transactions on Parallel and Distributed Systems, 25(3):663–672, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2014:EPO

[ZLCZ14] Yongxiang Zhao, Yong Liu, Changjia Chen, and Jianyin Zhang. Enabling P2P oneview multiparty video conferencing. IEEE Transactions on Parallel and Distributed Systems, 25(1):73– 82, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2015:PPS

[ZLDC15] Jun Zhou, Xiaodong Lin, Xiaolei Dong, and Zhenfu Cao. PSMPA: Patient self-controllable and multilevel privacy-preserving cooperative authentication in distributed-healthcare cloud computing system. IEEE Transactions on Parallel and Distributed Systems, 26(6): 1693–1703, June 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/06/06779640-
abs.html.

Zhu:2011:ITI

[ZLF^{+}11] Hongzi Zhu, Minglu Li, Luoyi Fu, Guangtao Xue, Yanmin Zhu, and Lionel M. Ni. Impact of traffic influxes: Revealing exponential intercontact time in urban VANETs. IEEE Transactions on Parallel and Distributed Systems, 22(8): 1258–1266, August 2011. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2013:RRT

[ZLGN13] Dian Zhang, Yunhuai Liu, Xiaonan Guo, and Lionel M. Ni. RASS: A real-time, accurate, and scalable system for tracking transceiverfree objects. IEEE Transactions on Parallel and Distributed Systems, 24(5):996– 1008, May 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhai:2015:ACC

[ZLJ^{+}15a] Jidong Zhai, Mingliang Liu, Ye Jin, Xiaosong Ma, and Wenguang Chen. Automatic cloud I/O configurator for I/O intensive parallel applications. IEEE Transactions on Parallel and Distributed Systems, 26(12): 3275–3288, December 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2015/12/06977978-
abs.html.

Zhang:2015:IPI

[ZLJ^{+}15b] Yu Zhang, Xiaofei Liao, Hai Jin, Guang Tan, and Geyong Min. Inc-Part: Incremental partitioning for load balancing in largescale behavioral simulations. IEEE Transactions on Parallel and Distributed Systems, 26(7):1900–1909, July 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/07/06844883-
abs.html.

Zheng:2017:EPB

[ZLJL17] Long Zheng, Xiaofei Liao, Hai Jin, and Haikun Liu. Exploiting the parallelism between conflicting critical sections with partial reversion. IEEE Transactions on Parallel and Distributed Systems, 28(12): 3443–3457, December 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

trans/td/2017/12/07983450-
abs.html.

Zhao:2016:TED

[ZLK^{+}16] Dongfang Zhao, Ning Liu, Dries Kimpe, Robert Ross, Xian-He Sun, and Ioan Raicu. Towards exploring data-intensive scientific applications at extreme scales through systems and simulations. IEEE Transactions on Parallel and Distributed Systems, 27(6): 1824–1837, June 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07159078-
abs.html.

Zhang:2015:MDF

[ZLL^{+}15] Lan Zhang, Xiang-Yang Li, Jingsheng Lei, Jiaguang Sun, and Yunhao Liu. Mechanism design for finding experts using locally constructed social referral Web. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2316–2326, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06866889-
abs.html.

Zhang:2017:RTF

[ZLL17a] Huayu Zhang, Hui Li, and Shuo-Yen Robert Li. Repair tree: Fast repair for single failure in erasure-coded distributed storage systems. IEEE Transactions on Parallel and Distributed Systems, 28(6):1728–1739, June 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/06/07742352-
abs.html.

Zhong:2017:BNA

[ZLL^{+}17b] Kan Zhong, Duo Liu, Lingbo Long, Jinting Ren, Yang Li, and Edwin Hsing-Mean Sha. Building NVRAMaware swapping through code migration in mobile devices. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3089–3099, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07944530-
abs.html.

Zhou:2017:ICW

[ZLL17c] Quan Zhou, Guohui Li, and Jianjun Li. Improved carry-in workload estimation for global multiprocessor scheduling. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2527–2538, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07873231-
abs.html.

Zhou:2018:EER

[ZLLD18] Quan Zhou, Guohui Li, Jianjun Li, and Chenggang Deng. Execution-efficient response time analysis on global multiprocessor platforms. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2785–2797, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08371627-
abs.html.

Zhu:2013:PTL

[ZLLZ13] Yanmin Zhu, Xuemei Liu, Minglu Li, and Qian Zhang. POVA: Traffic light sensing with probe vehicles. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1390–1400, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2013:PLU

[ZLN^{+}13] Xuyun Zhang, Chang Liu, Surya Nepal, Suraj Pandey, and Jinjun Chen. A privacy leakage upper bound constraint-based approach for cost-effective privacy preserving of intermediate data sets in cloud. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1192–1202, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2018:EDB

[ZLS^{+}18] Yu Zhang, Xiaofei Liao, Xiang Shi, Hai Jin, and Bingsheng He. Efficient diskbased directed graph processing: A strongly connected component approach. IEEE Transactions on Parallel and Distributed Systems, 29(4):830–842, April 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/04/08118091-
                                        abs.html.

Zhou:2018:AMR

[ZLT^{+}18] Ke Zhou, Wenjie Liu, Kun Tang, Ping Huang, and Xubin He. Alleviating memory refresh overhead via data compression for high performance and energy efficiency. IEEE Transactions on Parallel and Distributed Systems, 29(7):1469–1483, July 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/07/08068250-
                                        abs.html.

Zhang:2014:VFP

[ZLW^{+}14] Zhaoning Zhang, Ziyang Li, Kui Wu, Dongsheng Li, Huiba Li, Yuxing Peng, and Xicheng Lu. VMThunder: Fast provisioning of large-scale virtual machine clusters. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3328–3338, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06719385-
abs.html.

Zhu:2018:EGV

[ZLW^{+}18] Jie Zhu, Qi Li, Cong Wang, Xingliang Yuan, Qian Wang, and Kui Ren. Enabling generic, verifiable, and secure data search in cloud services. IEEE Transactions on Parallel and Distributed Systems, 29(8):1721–1735, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/08/08295253-
abs.html.

Zhong:2019:TFL

[ZLW^{+}19] K. Zhong, D. Liu, Y. Wu, L. Long, W. Liu, J. Ren, R. Liu, L. Liang, Z. Shao, and T. Li. Towards fast and lightweight checkpointing for mobile virtualization using NVRAM. IEEE Transactions on Parallel and Distributed Systems, 30(6): 1421–1433, June 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2014:SRL

[ZLX^{+}14] Yunfeng Zhu, Patrick P. C.

Lee, Yinlong Xu, Yuchong Hu, and Liping Xiang. On the speedup of recovery in large-scale erasure-coded storage systems. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1830–1840, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2014:SLA

[ZLY^{+}14] Xiuyuan Zheng, Hongbo Liu, Jie Yang, Yingying Chen, Richard P. Martin, and Xiaoyan Li. A study of localization accuracy using multiple frequencies and powers. IEEE Transactions on Parallel and Distributed Systems, 25(8):1955–1965, August 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhu:2019:BBL

[ZLYL19] Z. Zhu, T. Liu, Y. Yang, and X. Luo. BLOT: Bandit learning-based offloading of tasks in fog-enabled networks. IEEE Transactions on Parallel and Distributed Systems, 30(12): 2636–2649, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:AID

[ZLZ^{+}14] Jian Zhang, Peng Liu, Zhuoyao Zhang, Ping Yu, Baihua Zheng, Jingjing Wu, Yongrui Qin, and Weiwei Sun. Air indexing for ondemand XML data broadcast. IEEE Transactions on Parallel and Distributed Systems, 25(6):1371–1381, June 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2016:CAO

[ZLZ^{+}16] Zhi Zhou, Fangming Liu, Ruolan Zou, Jiangchuan Liu, Hong Xu, and Hai Jin. Carbon-aware online control of geo-distributed cloud services. IEEE Transactions on Parallel and Distributed Systems, 27(9):2506–2519, September 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/09/07345588-
abs.html.

Zhan:2017:CHD

[ZLZ^{+}17] Zhi-Hui Zhan, Xiao-Fang Liu, Huaxiang Zhang, Zhengtao Yu, Jian Weng, Yun Li, Tianlong Gu, and Jun Zhang. Cloudde: A heterogeneous differential evolution algorithm and its distributed cloud version. IEEE Transactions on Parallel and Distributed Systems, 28(3):704–716, March 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2017/03/07530859.
                                        pdf.

Zhang:2013:AAS

[ZM13] Yongpeng Zhang and Frank Mueller. Autogeneration and autotuning of 3D stencil codes on homogeneous and heterogeneous GPU clusters. IEEE Transactions on Parallel and Distributed Systems, 24(3):417–427, March 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2012:HCS

[ZMA12] Zijie Zhang, Guoqiang Mao, and Brian D. O. Anderson. On the hop count statistics in wireless multihop networks subject to fading. IEEE Transactions on Parallel and Distributed Systems, 23(7): 1275–1287, July 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2010:RTM

[ZMF10] Yongji Zhou, T. X. Mei, and Steven Freear. Realtime modeling of wheel-rail contact laws with systemon-chip. IEEE Transactions on Parallel and Distributed Systems, 21(5):672– 684, May 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

ZhiBin:2013:LLT

[ZML13] Huang ZhiBin, Zhu Mingfa, and Xiao Limin. LvtPPP: Live-time protected pseudopartitioning of multicore shared caches. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1622–1632, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2017:SEM

[ZML^{+}17] Da Zheng, Disa Mhembere, Vince Lyzinski, Joshua T. Vogelstein, Carey E. Priebe, and Randal Burns. Semiexternal memory sparse matrix multiplication for billion-node graphs. IEEE Transactions on Parallel and Distributed Systems, 28(5): 1470–1483, May 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/05/07593270-
abs.html.

Zhang:2013:NED

[ZMLT13] Haitao Zhang, Huadong Ma, Xiang-Yang Li, and Shaojie Tang. In-network estimation with delay constraints in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(2):368– 380, February 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2015:CCF

[ZMTL15] Dong Zhao, Huadong Ma, Shaojie Tang, and Xiang-Yang Li. COUPON: A cooperative framework for building sensing maps in mobile opportunistic networks. IEEE Transactions on Parallel and Distributed Systems, 26(2):392– 402, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/02/06748096-
                                        abs.html.

Zheng:2017:HES

[ZMW17] Wenli Zheng, Kai Ma, and Xiaorui Wang. Hybrid energy storage with supercapacitor for cost-efficient data center power shaving and capping. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1105–1118, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/04/07563872-
                                        abs.html.

Zomaya:2014:POC

[Zom14] Albert Y. Zomaya. Paretooptimal cloud bursting. IEEE Transactions on Parallel and Distributed Systems, 25(10):2670–2682, October 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06587242-
abs.html.

Zou:2014:TAP

[Zou14] Lei Zou. Topology-aware partial virtual cluster mapping algorithm on shared distributed infrastructures. IEEE Transactions on Parallel and Distributed Systems, 25(10):2721–2730, October 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/10/06589562-
abs.html.

Zhang:2011:IBR

[ZPD11] Peng Zhang, Reid Powell, and Yuefan Deng. Interlacing bypass rings to torus networks for more efficient networks. IEEE Transactions on Parallel and Distributed Systems, 22(2):287– 295, February 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2016:CGS

[ZQCZ16] Youhui Zhang, Peng Qu, Jiang Cihang, and Weimin Zheng. A cloud gaming system based on user-level virtualization and its resource scheduling. IEEE Transactions on Parallel and Distributed Systems, 27 (5):1239–1252, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/05/07109163.
                                        pdf.

Zhao:2013:EEK

[ZQH13] Huawei Zhao, Jing Qin, and Jiankun Hu. An energy efficient key management scheme for body sensor networks. IEEE Transactions on Parallel and Distributed Systems, 24(11): 2202–2210, November 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2016:BAR

[ZQL^{+}16] Sheng Zhang, Zhuzhong Qian, Zhaoyi Luo, Jie Wu, and Sanglu Lu. Burstinessaware resource reservation for server consolidation in computing clouds. IEEE Transactions on Parallel and Distributed Systems, 27(4):964–977, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/04/07091926-
                                        abs.html.

Zhao:2013:IER

[ZQSY13] Jia Zhao, Chunming Qiao, Raghuram S. Sudhaakar, and Seokhoon Yoon. Improve efficiency and reliability in single-hop WSNs with transmit-only nodes. IEEE Transactions on Parallel and Distributed Systems, 24(3): 520–534, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:EDC

[ZQWL17] Sheng Zhang, Zhuzhong Qian, Hao Wu, and Sanglu Lu. Efficient data center flow scheduling without starvation using expansion ratio. IEEE Transactions on Parallel and Distributed Systems, 28(11): 3157–3170, November 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/11/07932173-
abs.html.

Zou:2018:CGE

[ZR18] Yun Zou and Sanjay Rajopadhye. A code generator for energy-efficient wavefront parallelization of uniform dependence computations. IEEE Transactions on Parallel and Distributed Systems, 29(9):1923–1936, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/09/07935535-
abs.html.

Zheng:2014:GDD

[ZRQA14] Mai Zheng, Vignesh T.

Ravi, Feng Qin, and Gagan Agrawal. GMRace: Detecting data races in GPU programs via a low-overhead scheme. IEEE Transactions on Parallel and Distributed Systems, 25(1):104– 115, January 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2018:TSF

[ZRS18] Tong Zhang, Fengyuan Ren, and Ran Shu. Towards stable flow scheduling in data centers. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2627–2640, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08355701-
abs.html.

Zhang:2015:MST

[ZRTL15] Jiao Zhang, Fengyuan Ren, Li Tang, and Chuang Lin. Modeling and solving TCP incast problem in data center networks. IEEE Transactions on Parallel and Distributed Systems, 26(2):478– 491, February 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/02/06762979-
abs.html.

Zhang:2010:EEB

[ZS10] Haibo Zhang and Hong Shen.

Energy-efficient beaconless geographic routing in wireless sensor networks. IEEE Transactions on Parallel and Distributed Systems, 21(6): 881–896, June 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhan:2013:LLC

[ZS13] Guoxing Zhan and Weisong Shi. LOBOT: Low-cost, self-contained localization of small-sized ground robotic vehicles. IEEE Transactions on Parallel and Distributed Systems, 24(4):744– 753, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zengin:2017:FAH

[ZS17] Salih Zengin and Ece Guran Schmidt. A fast and accurate hardware string matching module with Bloom filters. IEEE Transactions on Parallel and Distributed Systems, 28(2):305– 317, February 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07485864-
abs.html.

Zhuravlev:2013:SEC

[ZSB^{+}13] Sergey Zhuravlev, Juan Carlos Saez, Sergey Blagodurov, Alexandra Fedorova, and Manuel Prieto. Survey of energy-cognizant scheduling techniques. IEEE Transactions on Parallel and Distributed Systems, 24(7): 1447–1464, July 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2017:OGP

[ZSC^{+}17] Wenyong Zhong, Jianhua Sun, Hao Chen, Jun Xiao, Zhiwen Chen, Chang Cheng, and Xuanhua Shi. Optimizing graph processing on GPUs. IEEE Transactions on Parallel and Distributed Systems, 28(4): 1149–1162, April 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/04/07572196-
abs.html.

Zhai:2011:EAC

[ZSH^{+}11] Jidong Zhai, Tianwei Sheng, Jiangzhou He, Wenguang Chen, and Weimin Zheng. Efficiently acquiring communication traces for largescale parallel applications. IEEE Transactions on Parallel and Distributed Systems, 22(11):1862–1870, November 2011. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2019:DBA

[ZSSR19] T. Zhang, R. Shu, Z. Shan, and F. Ren. Distributed bottleneck-aware coflow scheduling in data centers. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1565–1579, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:MRC

[ZSW^{+}15] Rui Zhang, Xiaojun Su, Jianping Wang, Cong Wang, Wenyin Liu, and Rynson W. H. Lau. On mitigating the risk of cross-VM covert channels in a public cloud. IEEE Transactions on Parallel and Distributed Systems, 26(8): 2327–2339, August 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/08/06878465-
abs.html.

Zhou:2019:ICP

[ZSW^{+}19] K. Zhou, S. Sun, H. Wang, P. Huang, X. He, R. Lan, W. Li, W. Liu, and T. Yang. Improving cache performance for large-scale photo stores via heuristic prefetching scheme. IEEE Transactions on Parallel and Distributed Systems, 30(9):

2033–2045, September 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:CRE

[ZSY14] Minjie Zhang, Danny Sutanto, and Dayong Ye. Cloning, resource exchange, and relation adaptation: An integrative self-organisation mechanism in a distributed agent network. IEEE Transactions on Parallel and Distributed Systems, 25(4):887– 897, April 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2013:SSD

[ZT13] Wenbo Zhao and Xueyan Tang. Scheduling sensor data collection with dynamic traffic patterns. IEEE Transactions on Parallel and Distributed Systems, 24(4):789– 802, April 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:CAP

[ZT14] Lu Zhang and Xueyan Tang. The client assignment problem for continuous distributed interactive applications: Analysis, algorithms, and evaluation. IEEE Transactions on Parallel and Distributed Systems, 25(3):785– 795, March 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zheng:2016:SPP

[ZT16] Hanying Zheng and Xueyan Tang. The server provisioning problem for continuous distributed interactive applications. IEEE Transactions on Parallel and Distributed Systems, 27 (1):271–285, January 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/01/07001654-
abs.html.

Zapater:2015:LAC

[ZTA^{+}15] Marina Zapater, Ozan Tuncer, Jose L. Ayala, Jose M. Moya, Kalyan Vaidyanathan, Kenny Gross, and Ayse K. Coskun. Leakage-aware cooling management for improving server energy efficiency. IEEE Transactions on Parallel and Distributed Systems, 26(10): 2764–2777, October 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/10/06915867.
pdf.

Zaitsev:2019:SLD

[ZTD19] D. Zaitsev, S. Tomov, and J. Dongarra. Solving linear Diophantine systems on parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 30(5): 1158–1169, May 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zahid:2018:SAN

[ZTG^{+}18] Feroz Zahid, Amir Taherkordi, Ernst Gunnar Gran, Tor Skeie, and Bjorn Dag Johnsen. A self-adaptive network for HPC clouds: Architecture, framework, and implementation. IEEE Transactions on Parallel and Distributed Systems, 29(12): 2658–2671, December 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/12/08385210-
abs.html.

Zhang:2017:AMI

[ZTH17] Lu Zhang, Xueyan Tang, and Bingsheng He. Analysis of minimum interaction time for continuous distributed interactive computing. IEEE Transactions on Parallel and Distributed Systems, 28(2): 401–415, February 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/02/07500040-
abs.html.

Zhao:2018:MMS

[ZTZ^{+}18a] Yangming Zhao, Chen Tian, Zhuangdi Zhu, Jie Cheng, Chunming Qiao, and Alex X.

Liu. Minimize the makespan of batched requests for FPGA pooling in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 29(11): 2514–2527, November 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/11/08345755-
abs.html.

Zhou:2018:QSC

[ZTZ18b] Keren Zhou, Guangming Tan, and Wei Zhou. Quadboost: A scalable concurrent quadtree. IEEE Transactions on Parallel and Distributed Systems, 29(3):673– 686, March 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

http://ieeexplore.ieee.
org/document/8066340/.

Zhou:2019:GEE

[ZTZQ19] Y. Zhou, S. Taneja, C. Zhang, and X. Qin. GreenDB: Energy-efficient prefetching and caching in database clusters. IEEE Transactions on Parallel and Distributed Systems, 30(5):1091–1104, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2014:DEI

[ZW14] Yaxiong Zhao and Jie Wu.

The design and evaluation of an information sharing system for human networks. IEEE Transactions on Parallel and Distributed Systems, 25(3):796–805, March 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zeng:2010:EMA

[ZWD^{+}10] Guokai Zeng, Bo Wang, Yong Ding, Li Xiao, and Matt W. Mutka. Efficient multicast algorithms for multichannel wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 21(1):86– 99, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2019:HGE

[ZWE19] Xia Zhao, Zhiying Wang, and Lieven Eeckhout. HeteroCore GPU to exploit TLPresource diversity. IEEE Transactions on Parallel and Distributed Systems, 30(1): 93–106, January 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2019/01/08409306-
abs.html.

Zeng:2017:RNN

[ZWFX17] Lingfang Zeng, Yang Wang, Xiaopeng Fan, and Chengzhong Xu. Raccoon: A novel network I/O allocation framework for workload-aware VM scheduling in virtual environments. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2651–2662, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07883963-
abs.html.

Zhu:2016:FTS

[ZWG^{+}16] Xiaomin Zhu, Ji Wang, Hui Guo, Dakai Zhu, Laurence T. Yang, and Ling Liu. Faulttolerant scheduling for realtime scientific workflows with elastic resource provisioning in virtualized clouds. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3501–3517, December 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07435325-
abs.html.

Zhang:2018:ESC

[ZWJ^{+}18] Weihua Zhang, Xin Wang, Shiyu Ji, Ziyun Wei, Zhaoguo Wang, and Haibo Chen. Eunomia: Scaling concurrent index structures under contention using HTM. IEEE Transactions on Parallel and Distributed Systems, 29(8): 1837–1850, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2018/08/07987039-
                                        abs.html.

Zhang:2019:EDA

[ZWJ^{+}19] X. Zhang, J. Wang, S. Ji, J. Yin, R. Wang, X. Zhou, and C. Jiang. An I/O efficient distributed approximation framework using cluster sampling. IEEE Transactions on Parallel and Distributed Systems, 30(7): 1494–1511, July 2019. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:RDR

[ZWL^{+}15] Guangyan Zhang, Jigang Wang, Keqin Li, Jiwu Shu, and Weimin Zheng. Redistribute data to regain load balance during RAID-4 scaling. IEEE Transactions on Parallel and Distributed Systems, 26(1):219– 229, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/01/06748089-
                                        abs.html.

Zhang:2016:XOX

[ZWL^{+}16a] Guangyan Zhang, Guiyong Wu, Yu Lu, Jie Wu, and Weimin Zheng. Xscale: Online X-Code RAID-6 scaling using lightweight data reorganization. IEEE Transactions on Parallel and Distributed Systems, 27(12): 3687–3700, December 2016.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/12/07434645-
abs.html.

Zhang:2016:LCF

[ZWL^{+}16b] Weihua Zhang, Haojun Wang, Yunping Lu, Haibo Chen, and Wenyun Zhao. A loosely-coupled full-system multicore simulation framework. IEEE Transactions on Parallel and Distributed Systems, 27(6):1566–1578, June 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://

csdl.computer.org/csdl/
trans/td/2016/06/07155591-
abs.html.

Zhu:2017:LTP

[ZWL17] Xiaodong Zhu, Junmin Wu, and Tao Li. Leveraging time prediction and error compensation to enhance the scalability of parallel multicore simulations. IEEE Transactions on Parallel and Distributed Systems, 28(9): 2553–2566, September 2017. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/09/07574321-
abs.html.

Zhao:2018:PAP

[ZWL^{+}18] Hui Zhao, Jing Wang, Feng Liu, Quan Wang, Weizhan Zhang, and Qinghua Zheng. Power-aware and performanceguaranteed virtual machine placement in the cloud. IEEE Transactions on Parallel and Distributed Systems, 29(6):1385–1400, June 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2018/06/08259446-
abs.html.

Zhao:2012:MLW

[ZWLL12] Yaxiong Zhao, Jie Wu, Feng Li, and Sanglu Lu. On maximizing the lifetime of wireless sensor networks using virtual backbone scheduling. IEEE Transactions on Parallel and Distributed Systems, 23(8): 1528–1535, August 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2016:ORS

[ZWLW16] Rui Zhang, Kui Wu, Minming Li, and Jianping Wang. Online resource scheduling under concave pricing for cloud computing. IEEE Transactions on Parallel and Distributed Systems, 27 (4):1131–1145, April 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/04/07106504-
abs.html.

Zhu:2015:FTS

[ZWQ^{+}15] Xiaomin Zhu, Jianjiang Wang, Xiao Qin, Ji Wang, Zhong Liu, and Erik Demeulemeester. Fault-tolerant scheduling for real-time tasks on multiple Earthobservation satellites. IEEE Transactions on Parallel and Distributed Systems, 26(11): 3012–3026, November 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/11/06928517-
abs.html.

Zhu:2019:ESF

[ZWT^{+}19] G. Zhu, Q. Wang, Q. Tang, R. Gu, C. Yuan, and Y. Huang. Efficient and scalable functional dependency discovery on distributed data-parallel platforms. IEEE Transactions on Parallel and Distributed Systems, 30(12):2663–2676, December 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:CBE

[ZWWF15] Zhaoyang Zhang, Honggang Wang, Chonggang Wang, and Hua Fang. Cluster-based epidemic control through Smartphone-based body area networks. IEEE Transactions on Parallel and Distributed Systems, 26(3):681– 690, March 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2015/03/06777528-
                                        abs.html.

Zheng:2017:LSH

[ZWY^{+}17] Kai Zheng, Lin Wang, Baohua Yang, Yi Sun, and Steve Uhlig. LazyCtrl: A scalable hybrid network control plane design for cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 28(1):115– 127, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/01/07460089-
                                        abs.html.

Zheng:2013:QRP

[ZWZ^{+}13] Zibin Zheng, Xinmiao Wu, Yilei Zhang, Michael R. Lyu, and Jianmin Wang. QoS ranking prediction for cloud services. IEEE Transactions on Parallel and Distributed Systems, 24(6): 1213–1222, June 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2015:IDF

[ZWZ^{+}15] Huan Zhou, Jie Wu, Hongyang Zhao, Shaojie Tang, Canfeng Chen, and Jiming Chen. Incentive-driven and freshness-aware content dissemination in selfish opportunistic mobile networks. IEEE Transactions on Parallel and Distributed Systems, 26(9): 2493–2505, September 2015. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/09/06942224.
pdf.

Zhao:2013:DND

[ZX13] Weiyi Zhao and Jiang Xie.

DoMaIN: A novel dynamic location management solution for Internet-based infrastructure wireless mesh networks. IEEE Transactions on Parallel and Distributed Systems, 24(8): 1514–1524, August 2013. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2019:PRA

[ZXG^{+}19] A. C. Zhou, Y. Xiao, Y. Gong, B. He, J. Zhai, and R. Mao. Privacy regulation aware process mapping in geo-distributed cloud data centers. IEEE Transactions on Parallel and Distributed Systems, 30(8): 1872–1888, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2017:ERO

[ZXL^{+}17] Juzi Zhao, Yu Xiang, Tian Lan, H. Howie Huang, and Suresh Subramaniam. Elastic reliability optimization through peer-to-peer checkpointing in cloud computing. IEEE Transactions on Parallel and Distributed Systems, 28(2):491–502, February 2017. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2017/02/07475901-
                                        abs.html.

Zhang:2013:NTC

[ZXW^{+}13] Jun Zhang, Yang Xiang, Yu Wang, Wanlei Zhou, Yong Xiang, and Yong Guan. Network traffic classification using correlation information. IEEE Transactions on Parallel and Distributed Systems, 24(1):104– 117, January 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2013:EEW

[ZY13] Da Zhang and Chai Kiat Yeo. Enabling efficient WiFibased vehicular content distribution. IEEE Transactions on Parallel and Distributed Systems, 24(3):479– 492, March 2013. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2014:TCP

[ZY14] Sheng Zhong and Haifan Yao. Towards cheat-proof cooperative relay for cognitive radio networks. IEEE Transactions on Parallel and Distributed Systems, 25(9): 2442–2451, September 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/09/06520841-
abs.html.

Zhou:2016:IBS

[ZYL^{+}16] Zhou Zhou, Xu Yang, Zhiling Lan, Paul Rich, Wei Tang, Vitali Morozov, and Narayan Desai. Improving batch scheduling on Blue Gene/Q by relaxing network allocation constraints. IEEE Transactions on Parallel and Distributed Systems, 27(11): 3269–3282, November 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2016/11/07404249-
abs.html.

Zhang:2017:DIA

[ZYL^{+}17] Xialei Zhang, Xinyu Yang, Jie Lin, Guobin Xu, and Wei Yu. On data integrity attacks against real-time pricing in energy-based cyber-physical systems. IEEE Transactions on Parallel and Distributed Systems, 28(1):170– 187, January 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/

                                        trans/td/2017/01/07440873-
                                        abs.html.

Zhang:2014:STP

[ZYLC14] Xuyun Zhang, Laurence T. Yang, Chang Liu, and Jinjun Chen. A scalable two-phase top-down specialization approach for data anonymization using MapReduce on cloud. IEEE Transactions on Parallel and Distributed Systems, 25(2):363– 373, February 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:VAS

[ZYQ^{+}14] Chao Zhang, Jianguo Yao, Zhengwei Qi, Miao Yu, and Haibing Guan. vGASA: Adaptive scheduling algorithm of virtualized GPU resource in cloud gaming. IEEE Transactions on Parallel and Distributed Systems, 25(11):3036–3045, November 2014. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2014/11/06671583-
                                        abs.html.

Zhu:2014:RCF

[ZYSH14] Yihai Zhu, Jun Yan, Yan Sun, and Haibo He. Revealing cascading failure vulnerability in power grids using risk-graph. IEEE Transactions on Parallel and Distributed Systems, 25(12):

3274–3284, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714491-
abs.html.

Zheng:2015:DGC

[ZYT^{+}15] Haifeng Zheng, Feng Yang, Xiaohua Tian, Xiaoying Gan, Xinbing Wang, and Shilin Xiao. Data gathering with compressive sensing in wireless sensor networks: A random walk based approach. IEEE Transactions on Parallel and Distributed Systems, 26(1):35– 44, January 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2015/01/06748092-
abs.html.

Zhang:2014:RTE

[ZYW^{+}14a] Xinglin Zhang, Zheng Yang, Chenshu Wu, Wei Sun, Yunhao Liu, and Kai Liu. Robust trajectory estimation for crowdsourcing-based mobile applications. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1876–1885, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhou:2014:OCD

[ZYW^{+}14b] Zimu Zhou, Zheng Yang, Chenshu Wu, Longfei Shangguan, and Yunhao Liu. Omnidirectional coverage for device-free passive human detection. IEEE Transactions on Parallel and Distributed Systems, 25(7): 1819–1829, July 2014. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhao:2016:HCB

[ZYW^{+}16] Jia Zhao, Kun Yang, Xiaohui Wei, Yan Ding, Liang Hu, and Gaochao Xu. A heuristic clustering-based task deployment approach for load balancing using Bayes theorem in cloud environment. IEEE Transactions on Parallel and Distributed Systems, 27(2): 305–316, February 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

                                        /www.computer.org/csdl/
                                        trans/td/2016/02/07039230-
                                        abs.html.

Zhou:2010:TAT

[ZYX^{+}10] Xiuyi Zhou, Jun Yang, Yi Xu, Youtao Zhang, and Jianhua Zhao. Thermalaware task scheduling for 3D multicore processors. IEEE Transactions on Parallel and Distributed Systems, 21(1): 60–71, January 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2014:FMC

[ZYZ^{+}14] Xinglin Zhang, Zheng Yang, Zimu Zhou, Haibin Cai, Lei Chen, and Xiangyang Li. Free market of crowdsourcing: Incentive mechanism design for mobile sensing. IEEE Transactions on Parallel and Distributed Systems, 25(12): 3190–3200, December 2014. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2014/12/06714605-
abs.html.

Zhang:2012:BTO

[ZYZC12] Yang Zhang, Liangzhong Yin, Jing Zhao, and Guohong Cao. Balancing the trade-offs between query delay and data availability in MANETs. IEEE Transactions on Parallel and Distributed Systems, 23(4):643– 650, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2015:IX

[ZZ15] Shurong Zhang and Xianwen Zhang. IEEE XPLORE. IEEE Transactions on Parallel and Distributed Systems, 26(2):434–443, February 2015. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/

                                        trans/td/2015/02/06766740-
                                        abs.html.

Zhao:2010:CCW

[ZZCD10] Jing Zhao, Ping Zhang, Guohong Cao, and Chita R. Das. Cooperative caching in wireless P2P networks: Design, implementation, and evaluation. IEEE Transactions on Parallel and Distributed Systems, 21(2):229– 241, February 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2010:LAP

[ZZF10] Guangquan Zhang, Zhaoliang Zhang, and Jianxi Fan. A locally-adjustable planar structure for adaptive topology control in wireless ad hoc networks. IEEE Transactions on Parallel and Distributed Systems, 21(10): 1387–1397, October 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2011:TTF

[ZZG^{+}11] Daqiang Zhang, Jingyu Zhou, Minyi Guo, Jiannong Cao, and Tianbao Li. TASA: Tag-free activity sensing using RFID tag arrays. IEEE Transactions on Parallel and Distributed Systems, 22(4): 558–570, April 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2017:UCR

[ZZH^{+}17] Feng Zhang, Jidong Zhai, Bingsheng He, Shuhao Zhang, and Wenguang Chen. Understanding co-running behaviors on integrated CPU/ GPU architectures. IEEE Transactions on Parallel and Distributed Systems, 28(3): 905–918, March 2017. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

/www.computer.org/csdl/
trans/td/2017/03/07501903-
abs.html.

Zhu:2016:EMO

[ZZLL16] Zhaomeng Zhu, Gongxuan Zhang, Miqing Li, and Xiaohui Liu. Evolutionary multi-objective workflow scheduling in cloud. IEEE Transactions on Parallel and Distributed Systems, 27(5):1344–1357, May 2016. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:/

/www.computer.org/csdl/
trans/td/2016/05/07127017-
abs.html.

Zhao:2018:NDH

[ZZQ18] Xujun Zhao, Jifu Zhang, and Xiao Qin. k NN-DP: Handling data skewness in kNN joins using MapReduce. IEEE Transactions on Parallel and Distributed Systems, 29(3):600–613, March 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/document/8089758/.

Zhang:2012:DPP

[ZZR12] Rui Zhang, Yanchao Zhang, and Kui Ren. Distributed privacy-preserving access control in sensor networks. IEEE Transactions on Parallel and Distributed Systems, 23(8):1427–1438, August 2012. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhang:2018:FDP

[ZZSZ18] Quan Zhang, Qingyang Zhang, Weisong Shi, and Hong Zhong. Firework: Data processing and sharing for hybrid cloud-edge analytics. IEEE Transactions on Parallel and Distributed Systems, 29(9): 2004–2017, September 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https:/

                                        /www.computer.org/csdl/
                                        trans/td/2018/09/08306827-
                                        abs.html.